A Complete Bibliography of Publications in the Journal of Theoretical Biology: 1990–1999

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
             WWW URL: http://www.math.utah.edu/~beebe/

24 February 2023 Version 1.01

Title word cross-reference

2 [SOR91]. 3 [LBND98, Mar92b, SL91a, Van91]. 4 [Lac90]. 6 [RP99]. > [Hir93]. ^{+} [BS92, Cla95, Joh93, NK90]. ^{2+} [AH90, Cha90, CJMBM92, HP97a, KD94, Kum91, LF95, LC97, LH91, LR94c, SW91]. ^{∘} [Fli99]. _{0} [O’N91]. _{1} [O’N91]. _{2} [CLY99]. _{3} [KD94, LR94c, WBC99]. _{cmax} [WBC99]. _{i} [LR94c]. _{ml} [WBC99]. α [LBDDG90, MGL^{+}95, Nak95, SST91, TSC90]. b_{1} [Yar96]. β [Iid90, Rad91, TSC90]. · [SST91, Spe90]. Δ [Fri98]. ϵ [Boe90]. g [MNB97]. K_{m} [RP96]. L = C [Hes90]. μ [RSB92]. N [Dug90, Bla97]. p [DGSK99, Khr98]. Φ [Nöl98, Nöl99]. Φ_{F (0)} [PN98]. → [MR98, O’N91].

-1 [LBND98]. -Adic [DGSK99, Khr98]. -amylase-catalyzed [Nak95]. -ATPase [BS92]. -azaadenine [SOR91]. -chymotrypsin [LBDDG90]. -D [Mar92b, SL91a, Van91]. -dependent [MNB97, CJMBM92]. -globin [Boe90, Iid90]. -helix [TSC90]. -induced [KD94]. -morphogen [Lac90]. -person [Dug90]. -phosphogluconolactone [RP99]. -player [Bla97]. -sarcin [MGL^{+}95]. -sarcin-membrane [MGL^{+}95]. -structure [TSC90].

-structures [Rad91]. -thalassemia [Iid90]. -values [Nöl98, Nöl99].

. [Boe90].

/K [BS92].

1 [Fin92b, Ham93, Mel93a, TCZ95, TL98, WDP98]. 1-dimenthylurea [LBND98]. 100 [CF94b]. 100-Å [CF94b]. 16S [AIK99]. 185 [BBM^{+}13a, BBM^{+}13b]. 19th [Ano99c].

2 [MK93, SST91]. 2- [RSB92]. 2nd [Ano99a].

353-379 [Ano92-29].

4 [LBND98]. 4’-dichlorophenyl [LBND98].

5 [Iid90, Say94]. 5’-3 [Say94]. 5’-splice [Iid90]. 5-trisphosphate-gated [LC97].

A-I [TP94]. Århem [Bec96, Lib97]. ABC [BB97d]. Abies [ISN91, SKI98]. Ability [Hur97a, SNN^{+}97, Beu90]. Able [AKMV96]. Abnormal [VW94a]. Abortive [SS96a]. Above [Wes93]. Above-ground [Wes93]. Abrupt [RV99]. absolute [BH92]. Absorption [LW94, SBM92]. Abundance [EL96]. Abundant [KKN99, Kat90]. Acanthaceae [WS95]. Acceptance [SHB90]. acceptor [GWS90, WAJ91]. Accidents [FL96]. Acclimation [SP94]. according [Sav92]. Account [DvH93, LH94, MCGPP97, PNB92]. Accounting [BF99b, Jon94, RGMG98]. Accounts [KPS97]. Accretive [Kaa94, Kaa95]. Accumulate [Mar97a]. Accumulation [BB97b, FDT^{+}93, YG94, FD92]. Accuracy [CSSF93, CC99, FSPC98, KK93, PGL94]. Accurate [Ker96]. Accurately [DBH94]. Acer [YLT98]. acervorum [THF92]. Acetabularia [CS90]. Acetylcholine [Got94]. achievement [Sav92]. Acid [BMG98, BBL97b, Dav98b, Duf97, FO96, GP93, Kal97, KQ96, NGP97, SS96b, Sta96b, TJ93, VGV93, WW97, YY94, FSD91, YYY91]. Acid-binding [VGV93]. Acidity [WSF99]. Acids [BDK^{+}97, DCM94, RJ96, Rak98, Shc93b, Shc94, JH95, WR93, Zha90]. Acquired [YCM97, YC99, JLL92]. Acquisition [BCV97]. across [Bra90a]. act [WBS97]. Actin [AOL98, Edm96, IET99, MC96, MBSM93, SS94d, Hum92, MB90, Mor90, SHB90]. Actin/Myosin [Edm96]. Action [AOU98, AWA98, BB99a, LBND98, LRMS98, LL93, Mar96, RAGO99, Szá99, Thi93, Ute93, CJ90, HCM^{+}92, Hur95, LG90a, LA90a]. Action-Response [Szá99, Hur95]. Activate [CAPP99]. activating [OHS95]. Activation [FV94, FK98, Kee96, MPSK94, MS94, PA98, ROL99, SL98, SSA94, UP96, WBS97, Zhu96, dLA97, Ale91, Cla95, HV90, VHV^{+}90]. activators [Wan90c].

Active [BLM94, CSSF93, MKGB98, BHVC90, LBDDG90, LRS90, Man95, Mur91]. Actively [Mar97a]. Activities [Gaá93, KV93, MR98]. Activity [ASA98b, BMG98, BD94, BKK96, Dem98, DBH94, DCL96a, GWL94, Gri94a, HK96, KK96, LLS94, LWS97, SSCvdM99, SB97, SHW94, VW94a, YG94, vOvP94, vOvP96, vOvO97, CJ90, FPB99, GE90, Has92, HH91a, MHTS90, Spu91, SV90, Str95, VVC95, Wan90e, ZT92]. Activity-Dependent [vOvP96, vOvP94, vOvO97]. Actomyosin [MHM94, Sud90]. Acts [EEG93, Ing94, SFKD99, Hem91, VHGV92]. Acute [BGE94, EWHB94, EWHB96, EPT97, RBBG98, Vol97, ECB91, MRB91]. Adapative [BP96]. Adaptation [CBS^{+}99, Che93, FSD91, HT96, LJ96, Ned97, MR90]. Adaptational [Gre97]. Adaptations [Mal97]. Adaptedness [Bur92]. Adaptive [AH98a, BN97, Fra97b, Gos92, OK94, SvH99, DvH93, Jam91, JH95, Sak91]. adaptor [Eld91a]. Addition [PSLD96]. Additive [HT97, Hin98]. Adenine [MFG99]. Adenine-Thymine [MFG99]. Adenylate [AKV96, AKMV96]. Adherent [SC99b]. Adhesion [Cum96b, Gra93, GS93, Kuz96, MIT96b, SEB^{+}93, Ste97, CH90a, Zhu91]. adhesions [BC90]. Adhesive [Cum99a]. adiabatic [Nöl95]. Adic [DGSK99, Khr98]. Adjusting [HL96b]. administration [RBR90, SC90]. ADP [CKR95, Fri98, Mat93]. ADP-ribose [Mat93]. ADP-ribosyl [Mat93]. Adult [OSM95, vdHdJ96, McN95]. advanced [RR92]. Advantage [Daw95, GV99, ZJ94, Eas91]. Advantages [NC97, LCD95]. Advection [NYI99]. adventurous [Ng91]. advertisement [Joh95]. aerobic [WW92]. Aerodynamic [Fed93]. affect [Cor95]. Affected [Zon98]. affecting [Has92]. Affects [SL98]. Affine [All93c]. Affinities [JJ98, WRP99]. Affinity [DSP96, DMP99, MCG95, PGBYP97, WMLK96b, CW92, FP95, Mac90a, Wei90a, Yee91]. Affinity-driven [DMP99]. Affordable [FM98]. Africa [McK95]. After [WG97a, AKL97, AM95, DR91, NK90, Por96]. Afterload [Cha98]. Against [KK94b, Møl97, RBB96, HG91, NMS95]. agaric [MR91b]. Age [HLS94, LNF96, McN94a, Rig94, SK99b, WNY98, WG97a, FC92, Kor91, Lor90, Lun90, Ste91b, Tay91, WM95]. Age-structured [HLS94, WM95]. aged [Wes93]. Ageing [All96, KK94e, Mor99]. agent [Bro90, How92, SAS92]. Agents [AWA98, MSHN97]. ages [GW95]. Aggregated [Kok97]. aggregates [ZCS^{+}90]. Aggregation [Bla94, EIU94, VVKZ94, vOPH^{+}96, WWC90]. Aggregations [FGLO99, GLR96]. aggression [Rat91]. Aggressive [HE98]. Aging [Kol97, RV99, vWP98, Aok91, Car90, Zhe91]. Aging-Abrupt [RV99]. Agonist [CS96, WBS97, Jär95, Mac90a]. Agonist-Receptor [CS96]. Agonistic [DS99]. Agonists [Zhu96]. Agro [And98]. Agro-ecosystem [And98]. Aided [Say94, WCPR91]. AIDS [LN95, RBH93, RBM97, Whi97]. Aimed [WDP98]. airway [DF91b]. Alanine [Rot95]. albers [WCW95]. albumin [Gli91]. Alcohol [Cze93, DP94b]. Algae [Hai93, PS93, PP92]. algal [MWAS91]. Algebra [JTC99]. Algebraic [Tyv93]. Algorithm [All93b, BN96, BN97, GMR98, GHW99, Hei94, McN93, Rak97, SC99a, SNH99, VGP95]. Algorithms [KV97, SHMH90]. Alignment [All93b, All93c, Hei94, ZM95, OT90]. All-pay [Kur99]. Allee [Sch99a]. Allele [GV98, Yi98, ZJ94]. Alleles [Get93, Cam95, Has92, Kel95c]. Allelocentric [SC93, Sib94]. Allelopathy [DL97]. Alligator [KCL^{+}96, SAF^{+}93]. alligators [Hai91]. Allocation [BSTV93, EGDG^{+}98, Har99a, McN94b, TN96, VV96, WC94b, Sak95]. Alloenergetic [Mel93b]. allometric [Fel95]. Allometries [Wit95]. Allometry [Dru93, ER94, Jol90, JD92, Spe90, VC95]. Allosteric [HK98, KH99, LC97, WK90]. Along [WK96, AP91a, CB94a, KSF98, LH90, Mir90, MC96, MB90]. Alpha [Miz95]. Alpha-fetoprotein [Miz95]. Alphanumerical [Ver94]. also [Cár91]. Alteration [Kun95, Gli91]. Altered [MSN96, TB91]. Alternans [Adl93, SAGB95]. Alternating [NS94]. Alternative [Edm96, For98, Nab96, Plu94, Wie96, BR90, BBB90, WT90]. Alternatives [Hai93]. altipetens [Ryt90]. altriusm [BK95]. Altruism [CST98, Fra94, Lew98, vBR98, BW91, Sey91]. altruistic [Hem91, Sai95]. Aluminium [EWHB94, EWHB96, EPT97, ECB91, EB92]. Aluminum [FA94, FG92]. alveolar [SC92a]. alveoli [Bla92]. Alzheimer [Gli91]. Amino [BMG98, Dav98b, DCM94, Duf97, FO96, GP93, RJ96, Rak98, Shc93b, Shc94, Sta96b, TJ93, WR93, WW97, YYY91, Zha90]. Amino-acids [WR93]. Amoebae [AKMS97]. Amoeboid [NU96, NYI99]. Among [AH96, BSTV93, CSB98, Mar97a, AH98a, BR90, GC91, JIDH97, Lim90, MM98, MS92b, Sav91a, SM92, Wil95, Yam93]. Amplification [ABK94, DA96, SFKD99, Trz99a, Trz99b, YY99, RSB92]. Amplitude [DPG^{+}94, FTCM90, Mur91]. amylase [Nak95]. anaerobic [FTD90]. analog [BJZ^{+}91]. Analogies [Mat93, Ser91]. analogs [SOR91]. Analogy [YYY91]. analyses [AHA92, CL92]. Analysing [KV97, MPO95]. Analysis [AM98, Adl93, AL96, AB98, AW94, AW93, AOU98, ABK94, BBB91a, BW98, BVvL96, BB97d, BT98, Bra96, BWK96, CPK96, Com94a, CS96, CMB95, DO98, DS99, DYL92, DBH94, DST96, DSD93, DSD94b, DP95b, Eas96, EG98, FV94, FWF97, FS94, Fra97c, FH99, GM94, Gie97, GMR98, GHW99, GET96, HFS97, HL98a, HK96, HH93, HCB96, IWH^{+}99, JTC99, Kaa95, KWB93, KSM97, Kok96, KDBK93, KQ96, LPP^{+}94, LRMS98, LGSM97, LPM95, LPRB96, LJ97, MSHN97, MKM96, Mal94b, Mar93b, Mar93c, MR98, Mog93, MA99, MTLK96, Nak95, NYY96, Nöl99, Oex98, OKAP98, ORG94, PNB92, PD93, PTBL98, Pro94, Rak97, RP99, RH96, SK96a, SvHT97, SM90a, San99, SS96a, SG90, Sch96, Seg94, SGP98, SS96b, Sib94, SM99, SC99b, SVRF94, SL91c, Sug98, TP94, TF94, Tor96]. Analysis [TCLB90, VKH^{+}98, VR94a, Vis93, WD97, WDP98, Whi98, WOQP93, YC99, YYYU97, ZZ94, ZZ99, ZW96, AHL95, BLB95, BCQ92, BPHH90, D.92, Dex95, DR91, DSD92, Enr90, FSDT91, FBH91, GRRB90, GW95, HR91, Iid90, JS92, Kat95, KA91, MPRB91, MMH90, MHTS90, Mir90, MBS92, NT95, Pat95, PVT91, RV90, RRB90, Sak91, Sav92, SP92a, SP92b, Sey91, SC90, Tak95b, Tch91, TCB91, Van91, WCW95, WC95a, Whi91, Wil95, YC91, ZM95]. Analytic [rvdBvdG96]. Analytical [AM95, ND99, Sak97, WB99, DF91b]. Anaplerotic [Kal97]. anastomosing [KH91]. Anastrophes [Bal97]. Anatomical [MCGPP97, SNN^{+}97]. Ancestor [Koc94a]. ancestral [Mil91b]. ancestry [Fin92b]. anchor [Sug98]. Anchorage [Enn93]. Ancient [Eld97, Di 95]. and/or [Mor90]. Androgen [CCK98]. Anemotaxis [Wil94]. Anesthesia [DBH94, Kat94a, Kat94b, Dex95]. aneuploid [For95]. aneuploids [Bar92]. angiogenesis [SL91c]. Angles [ALS93, Mul93a]. Animal [BF99a, GT97, Koo95, SSCvdM99, Smi94, SH95, Ujh96, Kim95, Kor91, McA91, Sav95b, TBD95, WL91]. Animals [Emm97, KD97, McA91]. anisogamy [Hur90]. Anisotropy [BP99a, SG93]. annealing [Sun95a]. Annotated [CB96b]. Announcement [Ano96a, Ano97a, Ano98a, Ano98b, Ano99b, Ano99c]. Announcements [Ano97b]. Annual [Ano97-28, LGM98, SH93a, AC90, Boi95]. Annuus [Han93]. anomalous [KJ95]. Anoxygenic [AIK99]. Ant [ALS93, BTD98, CSB98, SMG93, TT93, BDG92, GGT92, Mil92b, Ryt90, THF92]. Antagonist [Min94, Jär95]. Anterior [SB97]. Anthus [FAB92]. Anti [NBSM97]. Anti-viral [NBSM97]. Antiarrhythmic [Liu93]. Antibiotic [Bas98]. antibiotics [SOR91]. Antibodies [BLGG^{+}93, TD96, OHS91]. Antibody [CS94a, Hav97, Hav99, OP96, Wei90a, Yee91]. Anticodon [KK93]. Antigen [AL96, BK97, CS94a, FV94, FP93, Sas94, SL91b, WR93, Man91]. Antigen-presenting [AL96]. Antigen-Receptor [FV94]. Antigenic [BvHS93, HW98a]. Antigenicity [OHS91]. Antigens [OP96]. Antiparallel [LSB^{+}97]. antiporter [Rad91]. antipredator [Lim90]. Antiretroviral [WDP98]. Antisense [Jai97, RBH98, FO92]. Antiviral [BB99a, BR94, MPRB91, MRB91]. Antiviruses [MSHN97]. Ants [CH94, HC94, ND99, WEK95]. Any [Ujh96]. Aorta [TE96]. Aperture [RGMG98]. Apical [Bar94]. Apis [SDB90]. APL [VGP95]. APL-programmed [VGP95]. Apolipoprotein [CC96, TP94]. Apparent [BW92, Bow99, GH98b, WMLK96b, Mac90a]. Appearance [Blo97, CLF95]. Apples [DR91, Enr90]. applicable [BB91]. Applicaiton [YYYU97]. Application [AB98, BLGG^{+}93, BW98, BT90a, CL98, DK95, FT93, GMR98, JTC99, KS96b, KKH99, KQ96, LH94, LL93, NS96, ONT97, PVT91, Rau94, SK99a, SP92d, TN96, Tor96, WZ96, CS90, Meg90, RCCM92, RBR90, Tak95b, WC95a]. Applications [CFC98a, GH98a, MBM98, MA99, Zha97, Wan90b, WK90, Wan90c, Wan90d]. Applied [Ale97b, CS94b, Kau92, SS96b, Woo99, DM92b, FBH91, KM95, LK91, LG91, Meg90]. Approach [BBPD93, BP98, BBIK97, BP94, CG96, CSSF93, Cho93, Fis99, FO96, GM94, Gho93, GMR98, HBJ^{+}97, HL98b, HU97, HH93, KMJ98, KV93, Lei97, Mot93, Niw96, NH99, Pea96, Pet96b, PDP97, SAB97, SFW^{+}96, Sta96b, Tar94, VK99, WZN97, Whi97, Wil93a, ZLYZ98, Zon98, And95, BP92, Gil92, GR95, HH91a, Her95, LWHZ91, Mar95a, Mar92b, Mil92a, PP92, SBM92, Sun95a, VK91, Wei90b]. Approaches [EEK93, KQ96, WBC99, DL90]. Approximate [BSS96]. Approximations [Hie97, ML94b]. April [Ano90-33, Ano90-39, Ano91-32, Ano91-38, Ano92-33, Ano92-39, Ano93p, Ano93v, Ano94n, Ano94t, Ano95p, Ano95v, Ano96l, Ano96r, Ano97t, Ano97z, Ano98r, Ano98w, Ano99w, Ano99-28]. aquatic [TD92]. Arabidopsis [MAB98]. Araceae [BJ96]. arbitrary [Cor95]. Archaebacteria [HS97]. Architecture [ORM99, Shc93a, SC99b, Sak91]. Architectures [TB95]. Area [DL96, KIF96, NNM99, Wal96, Ben92, Ing90, May90]. area-concentrated [Ben92]. Area-restricted [Wal96]. arginine [Kun95]. arising [DF91a, HW98a]. Arithmetic [Koz96, Gin95]. Arithmetic-Normal [Koz96, Gin95]. Arithmetical [Shc93b, Shc94]. Arm [dLA97]. Arms [HS96, Här99b, dLA97]. array [Spu91]. Arrays [FR93]. Arrested [Mir98]. Arrhythmia [ASA98b, DBH94, DRLB91]. Arrhythmias [RWDM98, BHP92, CL92, DYL92]. Arrival [RWB98]. Arrivals [BCV97]. Arterial [BU98, CK96, KB93b, SNN^{+}96, SNK^{+}99, WGH99, Zam99, GE90, KB91, Pol92b, YH90]. Arteries [Jam99b, Woo99]. arterioles [WPP90]. arteriosclerotic [DB95]. Arthropod [GM94]. Article [Bec96]. Artifactual [Fra97d]. Artificial [GV91, Var93a, Var98]. Asarum [KKA96]. Ascendency [Bur94]. Ascomycetes [NH96]. Asex [KK94b]. Asexual [Daw95, JIDH97, JM98, RBP99, Han92]. Asexuality [Jud97]. asexuals [JM95]. aspect [SH91a]. Aspects [Woo99, ZL96, RKM92, Wes91, Woj92]. Aspiration [Pos99]. Assays [MKM96]. assemblage [JFM90]. assemblies [RB91]. Assembly [CMG^{+}96, Dwy98, KOTL97, ML97a, Dra90, FO92, THS^{+}95]. assembly/disassembly [THS^{+}95]. Assertion [TD96]. Assessing [HH91b, Whi97, Wie90]. Assessment [BBS^{+}98, BP98, Fis99, GGML93, HL98b, MNB97, MGMD96, NS96, Rat90, ES95, MR91a]. Assimilate [Wes93]. Assimilation [GVL93, PM99]. assisted [FC93]. Associated [AB98, FDT^{+}93, PGK94, Pet96a, SW96]. association [GWS90, Hum92, Orp90]. associations [MP91]. assortative [Fia92]. assumption [SP92e]. Assumptions [Kal98, Irv90]. astonishing [Hes90]. Astral [Yos99]. Asymmetric [HLT96, Kim93, Kis99, Li94, Nuy94, RB97b, NS91]. Asymmetrical [LL97]. Asymmetries [LH94]. Asymmetry [MP93, TH98a, ZL97, PW91]. Asymptotic [PD93]. Asynchronous [Sak92, Tho91]. asynchrony [Hai91, Kra95a]. Athletic [FVB97]. Atom [RJ96, Dav98b]. ATP [BM95, Fri98, Kem94, SNH99]. ATP-producing [SNH99]. ATP/ADP [Fri98]. ATPase [BS92, CKR95, Ste90a]. Atrial [ZH97, ZWH98, HZ95, ZCS^{+}90]. atrioventricular [SAGB95]. Attached [MKC^{+}98]. Attachment [KD99a, Sin96]. Attack [PTBL98]. Attacks [PLB96]. Attainable [FS98]. Attentive [Smi94]. Attractor [Hav97, JSR99, Hav91]. Attractors [BG96, BH95]. Attrition [Bla97, Kim93, KK98, MGMD96, Nuy94]. August [Ano90-32, Ano90-38, Ano91-35, Ano91-41, Ano92-31, Ano92-37, Ano93s, Ano93x, Ano94o, Ano94u, Ano95r, Ano95x, Ano96n, Ano96s, Ano97u, Ano97-27, Ano98x, Ano98-28, Ano99t, Ano99z]. Author [Ová91b]. autocatalysis [VHGV92]. Autocatalytic [CK99, FLR93]. Autoimmune [BOS98]. Autoimmunity [RBH93, SvH99]. Automata [Bea93, BGE94, EEK93, KMJ98, KC97, TPAS95, Tho91]. Automatic [CGB98]. Automaton [KM96b, PDP97, QZDA93, Var98]. Automimicry [Gui94]. autonomous [FM91]. Autonomy [SM97e]. autoregression [Sta92]. Autoregulated [OK94]. Autoregulation [KO93]. Autoregulatory [Kel94, Kel95b]. Autorotation [YA97]. Autosterilization [WH94a]. autosynchronization [THF92]. Autosynergism [FPB99]. Autotetraploid [Tyv94]. autotrophy [EGDG^{+}98]. auxiliary [LSM^{+}92]. Availability [Sak96, Sak97, Rob90a]. available [Pet91]. avalanches [KJ91]. avenue [MA95]. Avian [DS99, HP97b, JSB93, WH97, For90]. Avifauna [KM96a]. avoid [Fra90]. Avoidance [BF99a, BB99b, HW98a, CA92, SW95b]. Avshalom [Yoc95]. Axes [Fra97d, Mul93a, Jol90]. axial [FB95, Zah96]. Axiom [Eli96]. Axis [Fra97d, BZGM91, Car90, JD92, Tak95b]. Axon [KSF98, MS97b]. Axonal [KSF98, BH92]. Axoneme [Lin94, Spu91]. azaadenine [SOR91].

B [Bec96, FV94, RDHD98, ALW91, BvHS93, CCFS96, CCS96, CC96, CW92, KP93b, MPRB91, MRB91, PNB92, SST91]. B-Cell [FV94, RDHD98]. B. [Bec96]. Baboons [GG98]. Bacillus [KMM^{+}97, Koc92, Kun95]. back [BN92a]. Backbone [Fer94]. Bacteria [AIK99, Bas98, CFC98a, GPL98, Koc95, KD99a, RBH93]. Bacterial [Bar98, CFC98b, DFG95, HW94, HMUA93, MAGO97, RZF^{+}99, SV97, BL95, Ber95a]. Bacteriophage [Mit96a, Abe90, Kun92]. Bacteriorhodopsin [TP96]. Badger [SAM97]. Bahaviour [SMS98]. Bailey [Doe97]. Balaenoptera [LIWE90]. Balance [EGDG^{+}98, Rog94]. Balanced [CC97a]. Balances [Ver94]. Baldwin [Anc99]. Band [Won94]. bar [Mul93a, MdR98]. Barrier [GM94, SB94, SGG95, SB92]. Basal [GE90, TMII98]. Base [DST96, Fer94, For99, KKS97, LL96, MFG99, ONT97, SW98]. Base-Backbone [Fer94]. Base-Pairs [ONT97]. Based [ASA98b, ASA98a, Alm99, BH94, Cum96b, DY99, DST96, Di 94, ELS97, HL96b, IWH^{+}99, Mat93, MWC^{+}97, ME93, NNM99, Orp94, OK94, Rot95, RDHD98, SK96a, Sea99, WR94, ZLYZ98, ZZ99, BBB91a, BDA99, DH98, Dug90, FPB99, GS95b, Her90, KKS90, Kan92, KV97, LG90b, RLRV97, RS97, Smi90a, Smi91, Smo95, SFW^{+}96, Whi98]. Baseline [Bow99]. Bases [CF94b, ZZ94, CF92, KPK92]. Basic [FCV99a, FF90, Liq91, NLS91, Spa92, Mit96a]. Basis [DRB97, For94a, JW99, KKN99, Plu94, Sac94, SM97d, Won99, BHP92, DRLB91, McC91]. Basket [FL96]. Batch [NKR97]. Bateson [For99]. Bats [SHW94]. Battle [Myl99]. Bayesian [KDBK93, SB94]. Be [Eli94, GS95a, Hur91, KL96, Now90, PA90, SL91b, Smi94, Sre91, Yoc95]. bearing [Ste91b]. beat [CK97b, SH90b]. beaten [Sre91]. beating [FBH91, Mur90]. beats [CL92]. Beck [LA96b]. Bed [KB94]. Beddington [HD97]. Bee [BSTV93, Vis93, Wat97]. Bees [Bou94, CS91, KSG90, Kra95b, LL90, PSE92, TPAS95]. Beetle [PLB96, PTBL98]. before [AM95]. Behave [ON98]. Behavior [BB93, FGLO99, MA97b, NU96, SA97, BW92, Cor95, JFM90, LdQW95, Lim90, Mil92b, Mor90, SG95, SDB90, Ste90b, TPAS95, THS^{+}95]. Behavioral [SG95]. Behaviour [AW93, Bea93, JSR99, KKN98, KK94d, KV97, Ric96, SEB99, TAA^{+}98, vOvP96, vRGTV98, Ben92, BE92, GGT92, Koe90, NZPB90, SF90, Ss90c]. Behavioural [EL97, FS94]. Being [KL96]. Bell [HU97]. Bell-Shaped [HU97]. Below [WHL93]. beluga [Boi95]. bend [Mur91]. Bending [CB94a, ZBB97, SH91a]. benefit [HS95]. Benthic [MNH97]. benzo [Mac90b]. Bet [Hut99]. Bet-hedging [Hut99]. Beta [HDF^{+}99]. Beta-binomial [HDF^{+}99]. Between [AFS95, BOS98, BT99, CG96, CCVH98, DvH93, Daw96, GP93, HA97, JO96, KKN99, LM94, LTZ96, LA96b, Mat93, Mul93b, NT97, Rak98, Rau94, Rux96, SAB97, SLR^{+}98, ST93, SZK98, Ujh96, ZLYZ98, BSGB91, Bea90, BT97, Bow99, CK95a, CL98, DCM94, DR91, Enr90, FGLO99, Fra94, FD92, GL95, Gou94, GW95, Grü95, HW98a, HLG92, HP95, KDG92, LL96, Lib91, Lor90, Man92a, Mar92a, McA91, MN92, MHTS90, Mil91a, Mró95, Nöl95, Now91, PSE92, PGK94, Pet91, RB91, RKM92, RP98, RLF93, San99, Ser91, SGP98, Spi91, Tak95a, Thi90, WKSH92, Wil95, WBC99, WE98, Zha90, vdGvdBB92]. betweenn [TL96]. Beyond [FB97, KC91]. Bi [KK94b]. Bi-gender [KK94b]. Biarticular [dLA97]. Bias [BSGB91, Tak98]. Biased [HH97]. biases [Wol91]. Bibliography [CB96b]. bifurcating [Sta92]. Bifurcation [BBIK97, BT98, RL94, SNN^{+}96, CL92, SH91b, GK90]. Bifurcations [Jos97, GE90, SAGB95]. Bilayers [PGK94, Rau94]. Biliary [GET96]. bimodality [Wys92]. bimodalizable [Wys92]. Binary [Ber95b, Trz99a, RN95]. Binding [BLGG^{+}93, BW98, BB97d, FK98, FR99, Hav97, HMEW96, Mal94b, SS96a, Seg94, SVRF94, XCC94, BW92, BSB95, Boy95, Com90, GWS90, Hum92, Jär95, Mac90b, MBSB95, PST91, TS91, VGV93, WAJ91, WH92, Yee91]. Binding-site [FK98]. binocular [McI95]. Binomial [Ute93, HDF^{+}99, SH90a]. Bioassay [BW98]. Biochemical [BS98a, CSFC91, CPK96, HK98, KH99, Kle98, NS96, Sav91a, VF98, Irv90]. Biochemically [Smi90a]. Biochemically-based [Smi90a]. biochemistry [McC91]. Biocides [JIDH97]. biodegradation [CH90b]. Biodiversity [HL96a]. Bioenergetics [ZL96]. Biogenesis [Eli94, Gán97, Yoc95]. Biol [BBM^{+}13a, BBM^{+}13b]. Biological [Ale97b, Bal97, BC97, CK94, Coc98, DRB97, EEK93, For96, Gra90a, Gra93, HWO90, Has97, HE93, HLT96, JW99, Kre92, Lin97, Mog93, MB93, Nak99, Ned97, PC97, QF94, SS99, Sjö97, VC98, AP91b, BC90, BSM90, GAP95, HE95, JG92, LVL92, LWHZ91, Man95, Mil91a, PO91, Ser91, SV92, WC95a, ZW95].

Biologically [KDBK93]. Biology [AM98, Ano97-28, Ano99a, BRT98, BLLÅ94, Blo94, DBS94, Gat95, HDF^{+}99, Hop94, KV93, MA99, RS98, Spe94, TPN98, vdS90]. Biomass [DK95, HS98a, BDK91, PAS91, TD92]. Biomechanical [MPSK94, Tab98, Zhu91]. biomechanics [De 95]. Biomedical [Ano99c]. Biomembrane [GBD99]. Biomembranes [BLM94, Che94]. Biomolecular [TP93]. Biomolecules [Li94]. Biopanning [MCG95]. Biophysical [PK97, sDA91, sDAB91]. Biophysics [Kol97]. Biopolymer [AH98b, OES93]. Biopolymers [RS98]. bioregulators [SP90]. Biosynthesis [FSPC98, FO96, ML97b, BM95, Mat91]. Biosynthetic [Rak98, VP93a]. Biosystems [DBS97]. Biotechnological [Tor96]. Biotechnologists [WK96]. Biotoxicity [FA94]. biotransformation [CH90b]. Bipedal [McG93, MA97a]. Biphasic [KL99]. Bird [Pen98, Spe94, WS90]. Birds [HA97, JSB93, KL96, LHA94, NM93a, Tho96a, Tho96b, Wag93, AP91a, GH90, Hed92, Hou90, KS91, KB92]. Birhythmicity [LG99]. Birth [Bur99, Emm97, Jam96, Jam98a, PSLB99, Jam90a, Jam92]. Birth-death [Bur99]. bistability [Ste90b]. bistable [Ish90]. bit [Cou91, GN95]. biting [Gal92]. Bivalent [BLGG^{+}93]. Bivariate [Jol90, JD92]. Blissful [Web99]. Block [KL98, LL97, LPBR96]. Blockade [GWL94]. blocking [BLB95, Bro90, PST91]. Blood [HHM98, KB93b, LD94, LPP^{+}94, LBN95, MMB99, MB97, NS96, PA98, WGH99, DB95, KB91]. Blood-Stage [MMB99, MB97]. Blot [Oex98]. Blowfly [FWF97]. blue [CC95, LIWE90]. Blum [Spe90]. Board [Ano90a, Ano90b, Ano90c, Ano90d, Ano90e, Ano90f, Ano91a, Ano91b, Ano91c, Ano91d, Ano91e, Ano91f, Ano92a, Ano92b, Ano92c, Ano92d, Ano92e, Ano92f]. Body [BDK91, Dru93, JE98, LP93, McC94, Mel93b, Wit95, Aok90, Kor91, KS91, Mar93b, Mar95a]. Body-size [BDK91]. Bond [DBB99, AdG90, Mró95]. Bonding [EKK98]. bonds [BC90, BT91]. Bone [LL93, MR90, SH90a, VC95]. bone-marrow [SH90a]. Bonhoeffer [YFG^{+}93]. Boolean [Bea93]. Boost [TCE95]. born [SL91a]. borne [Bas98, NBBH99]. botanic [GR95]. Botanical [KW98c]. both [AFS95, BSB95, RHS^{+}95]. Bottle [BS94]. Bottom [MNH97, Her95]. bottom-up [Her95]. Bound [CAPP99, RKM92]. Boundaries [SB97]. Boundary [MBM98, Ves98, YA97, Ste91a]. boundedness [Van91]. Bourgeois [ES95]. Bout [Mor98]. Boutons [CGB98]. Bouts [TAA^{+}98]. Bovine [MFM97, FBL^{+}95]. Bowel [MAC99a, MAC99b]. boxes [Nus92]. bp [Bai93]. Brahnched [Ost96]. Braid [LBW^{+}95]. Braid-like [LBW^{+}95]. Brain [Bec96, GT97, Hak94, Ing94, LL94, LA96b, Rig94, Sal96, Sve94, Vrb98]. Brain-How [LL94]. branch [FF90, FAF90]. Branches [DM95]. Branching [ALS93, Kis99, TPN98, LQW92, LG91, Spa92, vVvP92]. Breath [WGH98]. breathing [LCD95]. Breeding [HDF^{+}99, Whi94, DH95, SW95b]. bridge [Mur92]. bridges [CKR95]. Broadcast [For96]. Brood [NM93a, GH90, HS95]. Brownian [Ste94, SSGR^{+}98]. bryozoans [Grü95].

Bud [DO99]. Budget [Mor98]. Buffering [Vin98, VF98]. Building [KP93a, TB95, GV91, KV97, Ryt90, SDB90]. Built [MO93]. Built-in [MO93]. bullfrog [SCG91]. Bureaucracy [MHSR^{+}96]. Burn [Pen98]. Bursting [BKK96, Cha90, KK96, SS94b, HI92]. Business [RBP99]. butterflies [WWC90].

C [For96, For99, GRUH99, Yoc95, BW92, Mil91b, LRB94, MR98, WBC99]. Ca [AH90, Cha90, CJMBM92, HP97a, KD94, Kum91, LF95, LC97, LH91, LR94c, Mir90, SW91, VS97]. Cable [MB93, TRVL96]. Cairina [HDF^{+}99]. Calcite [ME93]. Calcium [CGB98, FR99, WS98, CLF95, Jaf95, JLK92, LOLM91, Leo92, PST91, SM90b, SBM92, WAJ91]. calcium-binding [WAJ91]. Calculate [KV93]. Calculated [ML94b]. Calculating [ODA97]. Calculation [BLGG^{+}93, Gaá93, Mar95a, NT97, NT99, OES93, SL91a, Ng91]. Calculations [DY99, HWO90, Nov97, SST91]. Calmer [Kno91]. CAM [BBL97b, MS94]. Cambium [Kra99]. cAMP [LG90b]. Can [FK98, Gra93, GS93, GMN96, JD99, Kok97, LL94, LCD95, Myl99, PDP97, SW98, Ste91a, UJ95, Whi97, Cus95, Hur91, KL96, LH94, MCGPP97, ON98, SL91b, SB90b, TCE95]. Cancer [Mor99, Ros98, vWP98, CSS91]. Cancerous [QZDA93]. cancers [SS90a]. canine [vdGvdBB92]. Cannot [Zon96, Sre91]. Canonical [Shc93b, Shc94]. Canopies [WBC99]. Canopy [Sto94]. Can’t [Bed96]. Cantor [Fu95]. Capabilities [Orp94, VP93a, VP93b]. Capability [FVB97]. Capacitance [GBD99]. Capacity [CST98, JM98, Sak94, Sch91a]. capillaries [WPP90]. capillary [SGP98]. Captured [Sek95]. Carbohydrate [AW94, WA94]. Carbon [EGDG^{+}98, Hir93, Mar97a, SSS93, ZvdBK97, FAF90]. Carboxylation [LAV^{+}97]. Carcinogenesis [Rig94, BCQ92, SS90a, Ste91b]. Carcinoma [TMII98, PNB92, Ste91b]. Cardiac [BH94, DLC98, GWL94, Gri94a, Kee96, LL97, LGSM97, LCR94, MCGPP97, OR97, PK93, PK97, PK99, RWDM98, TGS93, VW94a, VS97, VR94a, VR94b, LRS90, LK91, LRJ92, LG90a]. Cardinal [RLF93]. Cardiomyocytes [VGV93]. cards [Sav92]. Care [BEE97, Bed96, BK90b]. Carlo [AB98, HH91b, KKV95]. Carrier [BE92]. Carrier-like [BE92]. carriers [LVL92]. cascade [HV90, Her95]. Cascades [EEG93, KW91, Pat95, VH90]. Case [ASML97, BJ96, EEG93, For98, KS96a, SP94, Spe94, Wic99, AS92, BS92, FR90, SP92e, TBD95, Van92b, Van91, WBS95]. cases [Fel91b]. Caste [Yam93, Pam91]. Castes [Yam93]. cat [CPLA95]. catabolism [Fie91]. Catalase [YYYU97]. catalysed [Top90]. Catalysing [WBS97]. Catalysis [Bri93, Kal98, AS92, Hav91, PT92b]. Catalytic [BMG98, Dem98, Hav99, NT99, Fer92]. Catalyzed [Nak96, Nak95]. Catastrophe [iT94]. Catastrophes [FSPC98]. catching [CA92, Hut99]. Categorize [CST98]. cations [CH90a]. Causal [Nak98]. Causation [vWP98]. causative [How92]. Cause [Jam98b, Kau98, Fri98, HA90]. Caused [CS94b, IP99a, KKN99, AP96a]. Causes [BDK^{+}97, Thi93]. cautious [Ng91]. CB [Ano99a]. CC [SST91]. CCAAT [Nus92]. CD4 [EP94, RBH93, Wic99]. CD4-like [RBH93]. CD5 [CW92]. Cdk2 [KH98]. Cdk2-kinase [KH98]. cDNA [BHVC90]. Celestial [HP97b]. Cell [ASA98b, ASA98a, AL96, ACC96, Ano99a, BLGG^{+}93, Bje94, Bje96, BGE94, BVW97, BT99, BFK^{+}93, CC97a, CC97b, CCFS96, CCS96, CLMH^{+}96, Cha93, CLW93, DP94a, DP98, DMP99, DCL96a, FV94, FP93, Fli99, GMST99, Gra93, GS93, HFS97, HHM98, HU97, JSR99, KH98, KY99, Kau98, KKV95, KG93, KW98a, KD99a, KV93, Kuz96, LBWP93, LBP^{+}97, MNB97, McL94, MSSG94, MPFHG96, MIT96b, Moc99, Mor99, NYI99, NC97, NAB^{+}94, NT93, Nov97, NTCN^{+}99, NBSM97, OP96, PSLB99, RWN98, SS95, SSM93, SS93b, SSGR^{+}98, TMI98, TMII98, TF93, VR94a, VR94b, Wie96, WLJN99, Yos99, YG94, ZvdBK97, vWP98, Aga95, Bro91, Cha90, CLF95, CM95, DP95a, EP94, FW90, Gra95b, HA90, HLG92, HZ95, JP90a, JP90b, KA91, KCM90, MR92, Man91, MK90, Nor92b, O’N91]. cell [PMA97, RB91, RDHD98, SP92b, SP92d, SCG91, SW95a, Sta92, SL91c, Ss90c, THY91, VR90, Whi91, WG92, ZCS^{+}90, Zhu91]. Cell-Antigen [FP93]. Cell-cell [ACC96, CM95]. Cell-Cycle [HFS97]. Cell-Dependent [OP96]. Cell-differentiation [TMI98]. Cell-rearrangement [Gra93]. Cell-sorting [Gra93]. Cells [BSW97, BD94, Bar94, BvHS93, Ben99, Big93, BG97, DSM99, End97, HLT96, HMEW96, KP93b, KM97, Kuz96, LPB96, LL97, MKZ97, MFHA^{+}95, MM98, NS96, NH99, SH97a, SPS93, SC99b, Tan99, VS97, Wic99, WOQP93, YG94, ALW91, BSOB92, Boe90, Bro91, CJMBM92, DF91b, GW95, HM95, KA91, LRS90, LK91, LRJ92, LG90b, MH95, Mor92, PO91, RGK91, Ste90b, THY91, Whi91, Zhe91]. Cells-Gene [Big93]. Cellular [Aga95, ABW90, AKMS97, BGE94, EEK93, Isr96, KMJ98, KC97, KK93, KM96b, MNB97, OKAP98, PDP97, QZDA93, RBH93, RV99, Ste97, WOI99, CH90a, DRLB91, EB92, LA90a, MU92, MH90, PNB92, Sri91, SV90, WM91]. Center [RDHD98]. Centers [PGBYP97]. Centipede [AF99]. Central [TMZ99, vRGTV98]. centres [Bea90]. Centrifugal [LD94]. Centromeric [Ost96]. Ceramiales [MWAS91]. cereal [Rob90a]. Cerebellar [MKZ97]. Cerebral [Gri94b, Ing94, KS99, LPP^{+}94, Orp94, SLR^{+}98]. cerevisiae [GW95, LPRB96, vRGTV98]. Cervical [JO96]. CG [SST91]. Chain [Col93a, EM95, HH98, MO93, Rot95, SM97b, SM97c, Tyv94, VK99, VGV93, Hav99, JR92, JH95, Mró95, MP91, OHS91]. Chains [Ale97b, HK96, JE98, KK94d, NRD94]. Challenge [BvHS93, Wäc97]. Challenges [Koc93]. Chambers [ODA97]. chance [FO92, HP95]. Change [Cha98, DP94b, HC93, JBSdlG97, JO96, Mel93b, Sea99, Spi93, Kum91, NK90, Wie90]. Changeovers [HMUA93]. Changes [Fri98, HP97b, JO96, MPSK94, ONT97, SH90b, TF96b, TWB^{+}93, YY94]. Changing [Pay97, Tan96, Man90]. Channel [BP94, CAPP99, Che94, Col93b, GBD99, Got94, Gre98, KN93, LC97, MB93, STI93, UP96, Bro90, BP92, CK95b, Cla95, DF91a, FW90, Kno91, LT91, SGG95, Sch91a]. Channeling [Edw96, Sre91, Cle91b, GC91, Her91, Sav91b, Shu91, Spi91, WM91]. Channelling [Bat91, KV93, MHM91, Pet99, Sti91, CB91a, Eas91, Fel91a, HS91, Hof91, Kay92, Kel91, Mat91, Ová91a, Ová91b, Pet91, Ure91, Wes91, WW92]. Channels [BLM94, Gho93, GM93, Ute93, CK95b, DL90, MLS92, SM90a, SW91, Sta90, Ste90b, TCE95]. Chaos [BS98a, BdS97, BC97, CK94, Doe94, End97, Fre94, GM90, KK96, LGSM97, LG99, NSH97, OBGGGRR93, Par97, RL94, áSá96, SMG93, SG98, Tho95, YQsZgZw97, CK95a, GAP95, Ing90, Jør95, KJ91, RM95, Rux95, SV92, SB95]. Chaos-game-derived [OBGGGRR93]. Chaotic [BTM93, HI92, LL97, LG90a, PC97, LT91, McC92]. Character [Bür93a, DM99, GR97b, HL98b, WM98]. Characteristics [HZ95, RJ96, WAJ91, Woj92]. Characterization [BDA99, Ber95b, CPK96, CK96, CL99a, CL99b, CCRM90, Reg93, SSCC93]. Characterize [LT98]. Characterizing [LOB92, Sch96, SS90a]. Characters [Gav97, DS90, JLL92]. Chargaff [BF99c]. Charge [ASML97, CMB95]. charged [RB91, SL91a]. Charging [KK93]. Charnov [LS94]. Chase [Gav97, GH98a]. cheat [PV95]. Cheating [Has94b]. Chemical [Cor98, Dav98b, JD99, Laz97, Rak97, RJ96, SLBT99, SSM93, AP91b, Kan92, SG95]. chemistry [LBDDG90]. Chemostat [Fle97, Nar98]. Chemostat-like [Fle97]. Chemostats [KK94d]. Chemotactic [DO98, EIU94]. Chemotaxis [DFG95, MTM^{+}96, SL91c]. Chemotaxonomic [KS96b]. Cheng [FL95]. Chewing [Dru93]. Chiasma [IWH^{+}99, NS95]. Chiasmata [IWH^{+}99]. Chick [VKH^{+}98]. Chickenpox [KB99]. chickens [PBJ^{+}92]. chills [Aok90]. chirality [LJ92]. Chitons [CSSF93]. Chloride [Sta98]. Chlorophyll [PN98, SGSS98]. Chloroplast [All93a, Mal94a]. Chloroplasts [DPG^{+}94]. Choice [FM98, HW98b, KB98a, Kok97, MHW93, Mar91, Mar97b, Wie90]. Chondral [Ham99]. Choosiness [SR98]. Chosen [SK96a]. Christ [TBD95]. Chromatid [ABK94, FR93]. Chromatin [BB96]. Chromophore [Dem94]. Chromosomal [SK96a, DKL95]. Chromosome [For94b, IP99a, KD99a, Mar96, Nor93, Sem96, HLG92, Mag95, SP99]. Chromosomes [LPRB96, SS93a, SS93c, Dav91]. Chromosomic [MAGO97]. Chronic [BGE94, RBBG98]. Chymotrypsin [Nöl99, Hav91, LBDDG90]. cichlid [Gal92]. Cichlids [TH98a]. Cilia [Lin94]. ciliary [Mur90, Spu91]. Ciliate [RH96]. Ciliated [OPRN93]. Circadian [Dum94, JK98, LG99, Nun98, ORG94, RR96, NLS91]. circle [RSB92]. circles [AP91a]. circuits [Kel95b]. Circular [AM96, AM97, AFM97, FR99]. circulation [GK91]. Circulatory [Mar93b]. Circumferential [ME93]. Circumstantial [Yar96]. Citric [Kal97]. city [DM92b]. Civilization [ZL96]. Clamp [LCR94, Mar92b]. Class [FV93, Mal94b, SV90, VS90, Wil95]. classes [Lor90]. Classic [Ber93]. Classical [Joh93, HV90]. Classical-type [Joh93]. Classification [CV93, NYY96, ZZ99]. Classified [AFS95]. Classifying [BG96]. Clastogen [LL93]. claw [MR91a]. Cleaning [PV95]. clearance [LBN95]. Cleavage [TCZ95]. Clock [Abe98, BD94, Dum94, RR96, NLS91, NSL91]. clocks [Hol91]. Clonal [Jud97, WR93]. Clonality [LBP^{+}97]. Clone [KKA96]. Clones [Bje94, MK90]. Cloning [Jud97, BHVC90]. clonogenic [MR92]. close [GE90]. Closed [Lor94b, SM97a, Tan94a]. closely [NS95]. Closer [Fis99]. CLOV [SK99b]. Clues [vWP98]. Clumped [Sas97]. Cluster [NYY96, WJ97, SST91]. clustered [Gra95a]. Clustering [FV94, Phi99, Sta96b, WJ97, ZGZ94]. Clusters [EIU94, WC95b, LL90]. Clutch [Lin94, SH93b, For90]. Co [CCFS96, CCS96, CLW93, Gaá93, Gie97, HZ93, HCB96, Mot91, Pec93, PT92a, RBH93, Sza93a, TP93, BR91, Dug90, Lah91, LOH92, PA90, PT92b, RV90, Vid91, AVM^{+}99, Com94a, LAV^{+}97]. co-evolution [Lah91]. Co-existence [HZ93]. Co-operation [CCFS96, CCS96, Mot91, Pec93, Sza93a, BR91, Dug90]. Co-operative [PT92a, TP93, LOH92, PA90, PT92b, Vid91]. Co-operativity [CLW93, RV90]. Co-ordinate [Gaá93]. Co-response [Gie97, HCB96]. Co-transmission [RBH93]. Coagulation [PA98, Rau94]. Coat [GRRF93]. coatings [Sav95b]. cobI5 [FLR93]. coccidial [PBJ^{+}92]. Code [AM96, AM97, AFM97, BA94, Dav98b, DCM94, Di 97, Di 98, Doi97, Duf97, KL97, Shc93a, Shc93b, Shc94, Ver94, JO96, KSH95, YYY91]. Codes [WH98b, Nev90]. codifying [Di 95]. Coding [Alm99, AM96, AM97, CCM^{+}97, Chi96, GOM94, HKH93, NSW97, Wil93b, ZZ94, AM90, AMO93, ALM95, BSB90, KSH95, Man92a, Man92b, SW98, TET91, YYY91]. Codon [ASE94, BSN^{+}94b, Doi97, GRRF93, Kat90, KK93, LL96, LRB94, BL95, Kun92, Kun95]. Codons [CZ93, ZGZ94, BSGB91]. Coefficient [Col93b, Mar93c, Sal96, YG94]. Coefficients [EG98, Gie97, GMN96, MRL96, PL93, Pet96a, Sen96, TF94]. Coevolution [ÅL97, CHM98, Dam99, KJ91, Leó90a]. Coevolutionary [Gav97, GH98a, Atk91, KJ91]. Coexistence [AP95, DC96b, GH97, HL96a, HH98, Neu98, NM92a, RP98, VV96, ZJ95, ZH98, CC95, Sza92a]. Cofactors [RBM97, VP93a]. Cognate [KO93, Kat90, SC92b]. Cognition [ÅL97]. Cognitive [Nak99]. Coherence [PDP97]. Coherent [Baj99, LVL92]. cohesion [Tri92, WL91]. cohort [BM90]. Coinfection [MA98]. Coleoptera [TLS94]. coli [AKL97, ASE94, BDA99, GRRF93, HLG92, HB91, KKV95, Kun92, Kuz95a, LL96, Mah93, MBSB95, Nor92b, Nor93, OWM96, RWK93, Ric96, SP92d, Say94, SG90, SLR97, SVRF94, VP93a, VP93b, Yar96, VK98]. Colicin [PH99]. collagen [Fie91]. collagenases [Fie91]. Collective [Her94, SFKD99, TB95, CS91, For95]. Collision [SL98]. colocalization [ALW91]. Colon [Bje96]. Colonial [DEK98]. Colonies [BSTV93, BTD98, TT93, Vis93, CSJM90, GGT92]. Colonization [MSSG94, MFHA^{+}95]. Colony [NG99, Wag93, Grü95, KA91]. Color [Chi96]. colour [KG90]. Column [HVKS99, JO96]. Comb [KP93a]. combinations [Jam91]. Combinatorial [CV93, GCB96, Gie97, KW98c, MS98, WR93]. Combined [HBJ^{+}97, SGG95]. Combining [Hei94, KK94e, MHW93]. combs [CSJM90]. come [Cle91b]. Comma [Nev90, KSH95]. Comma-free [Nev90]. comma-less [KSH95]. Commands [LSG97]. Commensuration [SK99b]. Comment [Koz96]. commentaries [Ová91b]. Comments [Bec96, Joh97, Kir92, LÅ94, MBS92, Yoc95]. commitment [Leo92].

Common [AMO93, ALM95, Cab92, TD92]. Communication [BB97c, HC94, Hur95, Ujh96, AMG95, CM95, Tan91]. Communities [AO96, CHM98, GH98b, ML97a, RB97b, PCN95]. Community [BT97, Has94a, ZL97, Dra90, Van90]. Compact [CM95]. Comparable [ZW96]. Comparative [LSB^{+}97, Mar93c, MAC99b, PGL94, WD97, D.92, RN95]. Comparison [Bar98, BDA99, Ber95b, CL99a, CL99b, DP98, Her90, MAGO97, ODA97, RH92, Tak95a, McK95, NC92, Sta90, TH90]. Comparisons [DR91, Enr90, Wil93a, WAJ91]. Compartment [MS97b, FTL^{+}90, Hag93, LW94, LL93]. Compartmental [FBS93, Mar93b, FSD91]. Compartmentalization [SL98]. compartmentation [Fri91]. Compartments [Nor93, VFS92]. compass [AP91a, Pau95]. Compensated [RR96]. Compensation [For94b]. Compensatory [HT96]. Competing [BTM93, Gat95, HH98, LM94, SH93b, TH98a, JR92]. Competition [BP96, BP97a, BP98, BF99a, BT97, Bow99, FCV99a, FCV99b, GM99, Gan99, Gou94, GH98b, HVKS99, Jon96, Kis99, Lor90, LNF96, MP99, NBBH99, Now91, RB97b, Sas97, SFGG^{+}99, SVRF94, Toq97, VV96, WH94b, ZJ95, Zon96, Arm90, Gro92, NM92a, SBV92, Wie90]. Competitive [Cam93, DP94a, Neu98, SM97b, Whi94, ZL97, HP95, Kat90, PW91, Sak92]. Competitors [ZJ95, ZH98]. complement [HV90]. Complementarity [CS94a, CCG^{+}98, DRB97, RBD97]. Complementary [AM96, AFM97, HBP92, Hon98, KSH95]. complementing [Sza92a]. Complex [CS96, HW98b, Hem98, Ker96, KKN98, KQ96, MHM94, MA97b, Ng96, TB95, Trz99a, Trz99b, WMLK96a, WMLK96b, WMLK96c, WG97b, vOvP96, BS91b, Sol95]. Complexes [Cze93, Dem94, SS96a, Bat91, BS91b, Cár91, Pet91]. Complexities [KYH99]. Complexity [AO96, DC99, Duf97, KB98b, Lon94, PL93, SBV92]. Complicated [Van90, HS91]. Component [CV93, For99, Hol99, KN93, Mar97a, SEB^{+}93]. Components [BW98, Mal94a, BSB95, SV90]. composite [WG92]. Composition [Dav98b, DY99, For99, WW97]. Compounds [DBS97, SV97, tBázsáSH96]. comprehensive [SAS92]. Compressibility [GP93, Nöl95]. Compressive [MM97]. Compromise [RdM91, SAB97]. compulsory [BK95]. Computation [Hop94, Fer92, ZCS^{+}90]. Computational [HC93, KOTL97, CW92, VpM95]. computations [JP90b]. Computed [Joh93, HZ95]. Computer [BLGG^{+}93, BKA97, BB93, FR90, GT97, Hon98, Ing90, Jon97, KH91, LL94, LQL94, LdQW95, MTS97, MK93, PH93, Pen98, RS97, Say94, SS93b, Ste90b, WH92, AGG91, LQW92, QC91a, QC91b, SS95, SRW95, WCPR91, Wei90a]. computer-aided [WCPR91]. Computer-simulations [BB93]. Computing [MWC^{+}97]. Conal [MCGPP97]. concatemers [MS90b]. concentrated [Ben92]. concentrating [Beu90]. Concentration [HU97, Kab94, KB93a, Nov97, NH99, SVRF94, BS91a, Bro91, BP92, Mac90a, MHTS90, SH91b]. concentration-response [Mac90a]. Concentrations [Gie94, GPL98, Ker96, KV93, Mah93, ZW96, ABW90, CB91a, For95, FTD90]. Concentric [NH99]. Concept [Kol97, Kor93, Mel93b, Sem96, Ves98, BW92, Bur92]. Conception [Jam96, Jam99a]. Concepts [Bar98, Cum96a, Gie97, vdS90, Mor91, Ons92]. Conceptual [CST98, Gri94a, Jon96]. concerning [Wan90e]. concrete [PVT91]. Condition [MBM98, Thi93, RG95]. Conditional [GR98]. Conditions [HLS94, JW95, LR94b, LL93, MKM96, ON98, TCE95, Yam93, YZ94, BW92, Com90, HB91, LSM^{+}92, MHTS90]. conduc [SAGB95]. Conductance [MB93, YLT98, OHS95]. Conducted [MKM96]. Conduction [LL97, NH99, RL94, BHP92, LL90, TCE95]. Cone [LQL94, TMI98, TMI99, LdQW95]. Conference [Ano99c]. configuration [GE90]. Conflict [ES95, Här99b, RR92, TS94, Vis93, Zha98, ZJ98, CW91, Eas91, Hai92, MBC95, RdM91]. Conflicts [LS94, MH96, WC94a]. conformal [GKMSS90]. conformation [MGL^{+}95]. Conformational [Che94, Kat96, PG94a, PG94b, TP94, Vol90]. conformations [SST91]. conforms [KSH95]. Confusions [OP94]. connectionist [Kra95b, Mil92b, MSR91]. Connections [vOW99]. connective [Com90]. Connectivity [DSP96, Cou91, dBP91]. Conscious [Lib96, Lib97, LA96b]. Consciousness [ÅL97, Hes94, Lin97, Orp94, Ber92]. Consensus [WR94, DM92a]. Consequence [HMEW96, Jam99b, Mor91]. Consequences [BN96, Daw96, GL95, NM93a, PMA97, PGL94, SBRG99, SA93, Hes90, HP95, Lim90, SW95b]. Conservation [BRT98, KKH99, VS97, RHS^{+}95]. Conservations [CPK96]. conservative [SYK91]. Conserved [Isr96]. Consideration [Coh91, Mró95]. Considerations [ABR99, Hoa90, SM97b, TF99, TS94, Ryt90, Wan91]. Consistent [Bje94]. Conspecific [NM93a]. Constancy [Gou94]. Constant [Gro92, Iwa93, Sun95b, BP92, Van92a, Van92b]. Constant- [Gro92]. Constants [CBFR98, YYYU97, Coh91, JLK92, Mac90a, WT90]. Constitutes [KV93]. Constrained [Ace96a, SNN^{+}96, SNN^{+}97, SNK^{+}99, HH95]. Constraint [AF99, LRB94, ONT97, PSE92, SHW94, Bro91]. Constraints [BJ96, EWHB96, GVL93, KH94, Lor94b, Thi93, Del95, Res95, Sri91]. Construction [KOTL97, WA94, CH91, MPRB91]. Constructive [SNN^{+}96, SNN^{+}97, SNK^{+}99]. Consumer [Mat95]. consumption [PAS91]. Contact [BTD98, PA98]. Contacts [Fer94, Nöl98, Nöl99]. containing [ZLYZ98]. Contamination [Kon95]. Content [AIK99, Sch97a, Sto98, WGH99, ZZ99]. Contest [Kur99]. contests [Kim95, Kor91]. Continuous [AH94, BP97a, BP99a, DT98, EMS97, HMUA93, WN99a, WN99b, Ben92, JG92, SF90]. Continuously [Sas94]. Continuum [Cho93, DO98, RLRV97, TM92b]. Continuums [HW98a]. Contracted [OSM95]. Contraction [Min94, MKC^{+}98, MM95, NU96, NYI99, OSM95, SAB97, San99, BT91, Koc90, MMH90, MBS92, Smi90a, Smi90b, TM92b]. contraction/expansion [Koc90]. Contractions [TGS93]. contrasting [McN95]. Contribution [Dav98b, FHB^{+}97, KV93, SBRG99, Bra90a, LRS90]. Contributions [CS96, OES93, Van91, Fra95a]. Contrition [BNS97]. Control [Ace93, Ace96a, AB98, All93a, BB94, Ben99, BT98, Bra96, BWK96, CPK96, CK94, CC96, DP98, DCL96b, DLC98, Eas96, EG98, Gie97, GET96, GMN96, Ham93, Hav97, HHM98, HK96, HH93, HCB96, KW91, KP93b, KQ96, LRZ^{+}96, MTS97, MRL96, MSN96, McG93, MHTS90, ML94a, NAB^{+}94, NT93, OA98, Pet96a, RB97b, Sal96, SP99, SKW98, SS96a, Sch96, Sen96, SSM93, SvH99, SAH96, TF94, TF96b, Tho96b, Tor96, WH94a, WZN97, Won94, vWP98, BB91, Bro91, BH92, GB95, GAP95, HR91, Irv90, Jam90a, Jam91, Jam92, Lac90, NT95, NLS91, PVT91, Sav92, SP92e, Tch91, Wes91]. Controlled [Jam96, BM98, Hum92, TB91]. Controls [MHSR^{+}96]. Convection [Ves98, WPP90]. Conventional [Boc98, HE98, DS99]. Convergence [EMS97]. Conversion [Bal97, PT92b, UJ95]. Convincing [DvH93, Pet91]. Cooperation [AP99, BKD99, CW98, Fre96, Koe97, NMI97, NNI98]. Cooperative [SR98, SR99]. cooperatively [SW95b]. Coordinates [GE96]. coplanarity [KPK92]. Coprophagy [MA95]. Copulation [NM93b]. Copulations [Wag93]. copulatory [FCV99b]. copy [Fra90, SP92e]. Copying [MDGB94, DH95]. core [Liq91]. Cores [DJZK96]. Corneal [GMST99, HC93, HE94]. coronary [GK91]. correct [DH92]. correcting [SF90]. Correction [MO93]. Correlate [BF99c]. Correlated [Ber92, BS93, Fra97c, Wil92a]. Correlates [Ing94]. Correlation [DLC96, LL96, Nöl98, Nöl99, RLF93, Hem90]. correlational [Sey91]. Correlations [Atk91, CT96, CL98, Man92a, WH98b, Man92b]. correspond [CW92]. Corrigendum [BBM^{+}13a, BBM^{+}13b]. Cortex [Gri94b, KS99, LH94, Orp94, RM93, SLR^{+}98, LH91]. Cortical [RLRV97, FTCM90, LH90, LH91, Rob90a]. corticogeniculate [McI95]. Cortiocotroph [LRMS98]. Cost [Bed96, Dam99, Duf97, Gan99, JM98, AMG95, Hou92]. Costlier [Hur97a]. Costly [NS99]. Costs [All93c, Duk99, Här99b, KL96, MA97a, NM93b, PM96]. could [Fri98]. counter [Nun90]. Counting [BG96, Fel91b]. Coupled [BSW97, Big93, KH99, KJ91, KM97, LL97, Pet99, ROL99, RRG98, SG98, Ste94, TUHT97, AS92, End97, Gor95, Jär95, Llo95, NLS91, NSL91, Sal91, SM92]. coupled-oscillator [SM92]. Coupling [CL98, FT93, Leb96, Mog93, OWM96, RBC98, SL98, YI95, DC90, Kee91, Rad91, VHGV92]. course [BP92]. Covalent [GW90, BC90, Mac90b]. Cow [FHR^{+}99]. Cows [BDK^{+}97]. Crab [Gri94a]. Cracks [DM95]. crassa [MR98, RVMR99]. Crassulacean [BBL97b]. Crawford [BP93]. Crawling [SH97a]. crayfish [SWA91]. creates [Com90]. Creatine [Kem94, CKR95, DC90]. creatures [Koe90]. Creeping [OPRN93]. Crick [Eld91a, KPK92]. Crithidia [SH90b]. Critical [BP97b, GLO98, LGM98, Pro94, RBH98, SFGG^{+}99, WBC99, LIWE90, MS95a, SM95, GLO98]. Criticality [KM96a, KD98, ORM99]. critique [AMW95]. Cross [AP91b, FP94, GR96, GR97c, Sat96, TF99, WW97, CKR95, GV91, Mur92, Ste91a, Yee91]. cross-bridge [Mur92]. cross-bridges [CKR95]. Cross-diffusion [AP91b, Sat96]. cross-entropy [Yee91]. Cross-immunity [TF99]. Cross-Species [WW97, GR96, GR97c]. crossbridge [Mor90]. Crossbridges [MM95]. crossovers [NS95]. Crowded [MGR97]. Crowding [ORM99]. Crown [SB94, GKMSS90, SB92]. Crown-of-thorns [SB94, SB92]. Cruciate [Mul93a, Mul93b, MdR98]. crypsis [BBB90]. Crypt [Bje94, LBWP93]. Crypts [LBP^{+}97, LG91]. CTL [Boc98]. Cubic [WMLK96a, WMLK96b, WMLK96c]. cuckoos [BBB90]. Cue [ALS93]. CUG [KL98]. Cultural [GM98, PB97a, Tak98, Ser91]. cultural/linguistic [Ser91]. Culture [AH94, And98, HMUA93, Fin92a, Zhe91]. cultured [BSOB92, RGK91]. Curculionidae [TLS94]. currency [Cab92]. Current [BS98c, LCR94, Sym94, VS97, Cha90, SM90a]. current-voltage [SM90a]. Currents [DCL96a, FT93, Joh93, RWDM98, Grü95, SGG95]. Curvature [KS98]. Curve [ASA98a, Zon97]. Curves [DL96, LWS97, NNM99, RP96, ZLYZ98, FTCM90, Mac90a, SW95a, Tay91]. cutaneous [Ste91b]. Cuticles [GM94]. Cutting [Duf97]. Cyanobacteria [Mar97a]. Cyanobacterium [WAO^{+}99]. cybernetic [BK90a]. Cycle [CC97a, CC97b, HFS97, JK98, KH98, Kal97, KKV95, MHWM94, NAB^{+}94, NT93, NTCN^{+}99, SS96b, YQsZgZw97, FTD90, GW95, KSG90, Nor92b, Rog95, SAS92, SM92, SS95, Smi90a]. Cycles [BSM97, CCK98, JD99, JSB93, LBP^{+}97, Mor98, Sta96a, Atk91, KS91, Van90]. Cyclic [Yam99, Han91]. Cyclical [HHM98, KKN98, RP98]. Cyclically [BN96]. cycling [DML95, Whi91]. Cyclization [FLR93]. Cycloheximide [RVMR99]. Cylindrical [CFC98b, NH99]. Cysteine [GMU95, NT99]. Cysteine-string [GMU95]. Cytochrome [Mil91b, Yar96, BW92]. Cytokine [BOS98, DJZK96]. Cytokinesis [Yos99]. Cytological [IWH^{+}99]. Cytometric [HFS97]. Cytometry [BFK^{+}93]. Cytonuclear [DA98]. Cytoplasmic [Fra97a, Fra98, Hur97b, VGV93, WS94, Hur91, Mau92]. cytosine [NS91]. Cytoskeletal [SFW^{+}96, SC99b]. Cytoskeleton [SC99b]. cytosolic [Jaf95]. cytotoxic [MGL^{+}95].

D [Mar92b, SL91a, Van91, MK93, SST91]. Daily [Sto94]. Dairy [FHR^{+}99]. Daisyworld [RR98, Sau94, Stö95]. Damage [BB97b, CS94b, BM95, Cus95]. Damped [Rou98]. dams [MA95]. Danaus [WWC90]. Dangers [HL96b]. Dart [AC96, Leo92]. Darwin [Ada98, PSLD96, Rob90b]. Darwinian [RR98, Rob91, RM93]. Data [AB98, ABK94, CCD^{+}98, DS99, DK95, FBS93, GRUH99, GR96, GR97c, JP90b, KS96b, LCR94, Mor96, SSCvdM99, Seg94, SB94, SFHP97, SSCC93, Woo93, YCM97, D.92, Don90, FSD91, JLK92, KA91, Mar92b, McK95, RN95, Sib95, Ste91b, TCB91, Yee91, MPRB91]. David [Koh91]. dawn [Mag95]. DDESS [YZ94]. deactivation [MS90a]. Deacylcortivazol [MK93]. Death [DRB97, KW98a, PSLB99, WW94, Bur99, MR92]. Debate [KV93]. Debris [BFK^{+}93]. Decay [CK97a, Li94, SK99b]. decaying [DSD94a]. December [Ano90-34, Ano90-40, Ano91-34, Ano91-42, Ano92-34, Ano92-29, Ano93o, Ano93y, Ano94s, Ano94x, Ano95n, Ano95u, Ano96p, Ano96u, Ano97p, Ano97v, Ano98t, Ano98z, Ano99r, Ano99s]. deceptive [AMG95]. Deciphering [BKNR97]. Decision [MWC^{+}97, SFKD99, SM97e]. Decision-Making [SM97e, MWC^{+}97, SFKD99]. Decline [WOQP93, Tri92]. Decomposer [NRD94]. Decomposition [Tar94]. Deducing [Sib95]. Defection [Sza93a, Mot91]. Defective [BN94, FM91, Sza92a]. Defects [Kra99]. defence [HG91]. Defences [JD99, TAN94b]. defending [DM92b]. Defibrillation [Kee96, KL99, PK97]. Defibrillator [BH94]. Deficiency [OVH^{+}96]. Deficient [AKL97]. Defined [Gaá93]. Defining [CA92]. Definite [ONT97]. Definition [BP93, Spe94, Cra90b, TCB91]. Deflector [HP97b]. Deformability [SC99b, DS92]. Deformable [Rau94]. Deformation [SZK98, DS92]. Degenerate [SK99a, SH91b]. Degradation [FDT^{+}93, KO93, MS97b, SV97, FTL^{+}90]. degrading [SB92]. Degree [vBW98]. Degrees [Fre96]. Dehydrogenase [Cze93]. dehydrogenases [Spi91]. Delay [MBM98, YZ97, HA92, LGBP92]. Delayed [Baj99, DH95, LCR94, VB93, Whi94]. Delays [MHSR^{+}96, Rod98]. Delbrück [Jon94]. Deleterious [Ano94b, MK97b, Kel95c, Kon95, Sai95]. deletions [AM92]. delivery [PBJ^{+}92]. Delta [CMML96, FTCM90]. Delta-Notch [CMML96]. deltoides [Kra99]. Demersal [DK95]. Demographic [Gil92, KE96, Rod96]. Demography [CCVH98]. Demon [ON98]. denaturants [sDAB91]. denaturation [sDA91, sDAB91]. Dendrites [WOQP93]. Dendritic [NGP97, vP97]. Denitrification [ARC97]. Density [AP95, rvdBvdG96, BBIK97, BN92b, Bur99, CFC98a, CB97b, CCK98, CV97, FK98, GH97, HS98a, JS97, LP93, Nak98, Rod98, Sak94, SW96, TN96, TP94, YZ94, YZ95, Yos99, YG94, ZJ97, Cre90, FTCM90, FC92, Kel92, McN95, PW91, Ryt90, Tak95a, WM95]. Density-and-frequency [Nak98]. Density-Dependent [GH97, AP95, BN92b, CCK98, TN96, YZ94, YZ95, Cre90, FC92, McN95, Ryt90, Tak95a]. density-independent [Tak95a]. Density-sensing [YG94]. Dental [SAF^{+}93]. dentate [AMT91]. Dentition [Lub97]. deoxyribonucleotide [Mat91]. Departure [KL96, Web99]. Dependence [Bur99, HS98a, Rod98, Wal96, ZJ97, LK91, Sme90, WM95]. Dependency [HU97]. Dependent [AH94, Ben99, For99, GH97, HL96a, JS97, LBP^{+}97, LNF96, MBM98, MTM^{+}96, Nak98, OP96, Sak94, SAH96, TRVL96, Yi98, vOvP96, AP95, BN92b, BSB95, CJMBM92, CCK98, Cre90, FC92, LKS98, MNB97, McN95, NNI98, QG92, RBR90, Ryt90, SM97a, SW91, Ste90b, Tak95a, TN96, Tay91, UP96, VVC95, YS92, YZ94, YZ95, vOvP94, vOvO97]. Depending [PSLB99, Nak91]. Dephosphorylation [EEG93]. Depletion [SA93]. Depression [LR94b]. Depth [Ben97]. Derivation [HD97, JLK92, MdR98, TW97, VHG^{+}90]. derivatives [CSFC91]. Derived [LR94c, Ost96, Rig94, FHB^{+}97, KDBK93, OBGGGRR93]. Deriving [SFHP97, TJ93]. Dermal [OSM95]. Descent [WN99c]. describe [BP92, DL90, LPS90]. describes [FTCM90]. Describing [BP94, DH98, Tar94, DP95a, MPO95]. Description [Kaa94, OKAP98, SM97b, SM97c, SC96, WMLK96a, tBázsáSH96, Gra95b, Han95, Sav91b, Tho91, Tch91]. descriptions [Mil91a]. Descriptive [BB97b]. Desensitization [BP94, ROL99, LG90b]. Design [Ace93, Ale97a, BH94, Fra97b, OA98, Pet96b, TF96b, Zhu96, Fra95c, Ste90a, WCPR91, vL91]. Designing [RP96]. Designs [SNH99, Fad92]. Despite [HS96]. Destabilizing [Rod98]. Destruction [BS98b, FK97, Neu98]. Detachment [CLW93, HMEW96]. Detailed [LR94c]. Detecting [Wys92]. Detection [CMPR93, HS97, Wil94]. Determinant [SAF^{+}93, TD96]. determinate [Van92a]. Determination [Dem94, DSD94a, JBSdlG97, Jon97, MNB97, MH96, Pet96b, Rak97, RJ96, Sal96, SP99, Wal99a, For95, Hai91]. Determinations [Pet96a]. Determine [MAGO97, San99, Nor92b, RCCM92, VC95]. Determined [CCI94, SNH99, TF94, WW94, Wit95]. Determines [SH93b]. Determining [Coh91, FLR93, Gie94, GCB96, FTD90, SF90, WT90]. Deterministic [Bar98, EV94, FWF97, KK96, LT91]. detrimental [GWS90]. Detritus [NRD94]. Developing [Con93, KS99, SLR^{+}98, vOvP94, dBP91]. Development [Bur94, DBSV94, Fra97b, Hem98, HC93, KY99, Koc94a, MFHA^{+}95, MPFHG96, Sac94, SvHT97, WCPR91, Won90, vOW99, vOvO97, Aok91, Bak95, GW90, GK90, KG90, MSR91, O’N91, PNB92, PAS91, Ric90, SK92, VVC95, VHSC92, SP92a, SP92c]. Developmental [AF99, DBS94, FL96, Hai91, Møl97, MM98, Res95, VC95, Bur95, Hol91, JG90, Kra95a, Tak95b]. Developmentally [Møl97]. Deviation [Alm99]. Deviations [BF99c]. Device [Joh93]. devices [Kre92]. devil [CLF95]. Devonian [HC93]. DFT [FS94]. diabetes [RBR90]. Diagram [Hag93, JLK92]. diagrams [MU92]. dialects [WS90]. Dialogue [BOS98]. Diapause [McN94b, RRG98, Nun90]. Diastolic [CBS^{+}99]. diazymes [DC90]. dicarboxylic [JH95]. dichlorophenyl [LBND98]. Dichotomous [Lon94]. Dichroism [FR99]. dicotyledonous [VHSC92]. Dictyostelium [AW94, AKMS97, BVW97, BVW99, DO98, LG90b, MPH99, SS96b, VSW97, VVKZ94, WA94, YG94, vOPH^{+}96]. Did [JBSdlG97, Laz97]. died [Kon95]. Diet [BDDF97]. Diet-Specific [BDDF97]. Difference [VR94a, Mar92a]. Differences [DvH93, Jos97, Koc92, KK98, Kur99, HP95, Sav91a]. Different [AMO93, BFG96, BB99b, DS99, EEG93, For96, Hop94, WBC99, Yam93, Coh91, GW95, Pol92b, Sai95, Smi91]. Differential [CB94b, Jon94, SP99, Ste94, LL90]. differentiated [THY91]. Differentiates [Alm99, SW96]. Differentiation [Bje94, Fli99, TF93, vdHdJ96, vWP98, FW90, HA90, MGP95, TMI98, THY91]. Difficult [BMS93]. diffuse [SB90b]. Diffusion [BIV97, BRV99, Con93, DL90, DP95b, GM94, Ham98, KSF98, LLS94, LAV^{+}97, MCGPP97, ML94b, NYI99, SEB^{+}93, Ste94, TMZ99, Yot93, YG94, AMW95, AP91b, CM95, FT93, KJ95, Lac90, LH90, LSM^{+}92, SGG95, Sat96]. diffusion-limited [KJ95]. Diffusivity [LJ95]. DiFrancesco [GLH97].

DiFrancesco-Noble [GLH97]. difuse [Van90]. digesta [FSDT91, SF90]. digestion [Fad92]. Digestive [ABH94]. Digoxin [Hav97]. Dilemma [CCG^{+}98, HS98c, Koe97, Kur99, Leo90b, Lor94a, MI98, Mow97, NS97, NS94, SNT95, TAN94b, WN99a, WN99b, BPHH90, Gri95]. Dilution [FBS93, FHR^{+}99, MKM96, FSD91, FBL^{+}95]. Dimension [PSLD96, Bur90, FPPSK95, XCS^{+}95]. Dimensional [BG96, CB94b, Pol92b, AVM^{+}99, BP99a, DSP92, FST99, FBH91, Spe90, SS94c, Tho95, ZH97]. Dimensionality [Mar93a]. dimensionally [TUHT97]. Dimensionless [LT98, ME93]. Dimensions [Has94a, KD99a, Mul94, PH93, ZBB97, Kaa95, MBC95]. dimenthylurea [LBND98]. Dimerization [FR99, WAJ91]. dinucleotide [AM95]. dioecious [Mau92]. dioecy [TM92a]. diol [Mac90b]. dioxygen [HA90]. Diploid [Sza93b, Tan99, Tyv93, YZ95, Yi98, WBS95, Bar92]. diploidy [Hur90]. dipole [THS^{+}95]. Dipterocarp [KKH97]. Dipterosiphonia [MWAS91]. Direct [Ham98, Kee96, KG93, Kir92, Mar90, RSB92, Di 95]. Directed [Sym94]. Direction [BF99c, Gaá93, ML97b, Ros98]. Directional [Bür93a]. Directions [Gaá93, May90]. directly [Gor95]. disadvantage [TM92a]. Disappear [Hut99]. Disappearing [Wic99]. disassembly [THS^{+}95]. Disc [TP94, Won99]. Disc-Sphere [Won99]. discharges [HI92]. discoideum [AW94, AKMS97, DO98, MPH99, SS96b, VSW97, VVKZ94, WA94, vOPH^{+}96]. Discounting [SNT95]. Discrete [DSM99, FS94, GR96, GR97c, GR97b, KD99b, McC94, MB93, ODA97, Sal96, Szá99, YQsZgZw97, AMT91, CSC90, CSS91, Hur95, Kee91, Lib91]. Discreteness [DSP96]. Discriminant [ZZ99]. Discriminate [Bed96]. Discrimination [AFS95, CLY99, For94b, Wal96, For92, Kat90]. discriminatory [Rat91]. Discussion [GLH97]. Disease [BOS98, BKA97, CGM^{+}98, CI97, PK98, RBBG98, Whi97, Gli91, Hoc91, JR92, SAS92]. Diseases [KS96a, WJ97, Lun90]. Disentangling [GH98b]. Disequilibria [DA98]. Disparity [FD92]. Dispersal [BBM^{+}97, BOG^{+}99, CH96, Ezo98, GM99, Gan99, JS97, Par97, RGAP97, SG98, BBM^{+}13a, BBM^{+}13b, BBM^{+}13c, SW95b]. Disperser [Har99a]. Dispersers [WJH97]. Dispersing [MNH97]. Dispersion [CB97b, CH94, Wei97, BBB91a]. Displacement [DM99, HL98b, OSM95, ZL97]. displaying [Wys92]. Displays [DS99, Kok97, PM96, AMG95, Liq91]. disproportionate [FD92]. disproveable [Shu91]. dissipation [Sch91b]. dissipative [Com90, LVL92]. Distance [LTZ96, Ste94, TJ93, Pen98, RKM92, WL91]. Distances [DCM94]. Distinct [Fle97]. distinctions [Res95]. distinctive [Kun92]. Distinguish [Koz96, ZLYZ98, Gin95]. Distinguishing [BR91, HA97]. Distortion [Sta96a, vBW98, Lib91]. Distributed [CH96, DL97, TMZ99, Bra90b, GGT92, HC91]. distributed-parameter [HC91]. Distribution [CSB98, CB94a, DSP96, DEK98, DBS97, FB97, KSF98, KIF96, Mar97a, Oex98, RM97, SW98, Sas97, SNK^{+}99, Sin96, Sto94, SA93, THY91, Ute93, WRP99, WG97a, WGH99, Yam93, ZZ94, tBázsáSH96, Bro91, Cam95, FTCM90, FS95b, HCS91, JV91, JH95, MS90a, SM90b, SC90, Yee91, ZW95]. Distributions [BFG96, CCM^{+}97, EL96, GV98, HK99, JJ98, Koz96, McN94a, Moc99, Mow97, DM92b, FAB92, Gin95, Wys92]. Disturbed [DGSK99, GG97b]. Disulfide [EKK98, Mar90]. dithiol [Mar90]. dithiol-disulfide [Mar90]. Diurnal [SHW94]. Dive [Mor98]. Divers [Mor98]. diverse [Sib95]. Diversification [KY99, Koc94a, MS98]. Diversity [BJS97, HS98b, Jud97, MN94, PH99, Sek95, Hur90, LA90a]. dividing [Mor92]. Division [NG99, NT93, TF93, FW90, KA91]. Divisions [Ver94, VpM95]. DNA [Alm99, Bai93, BKA97, BF99b, BM95, CCM^{+}97, CT95, CT96, CL98, CS94b, Col93a, DLC96, Dru97, FPPSK95, For99, FD92, GOM94, Han95, Hei94, How92, JV91, KOTL97, Kat96, Koz91, KDBK93, LJ92, LPBR96, LWHZ91, LBW^{+}95, LSB^{+}97, LJ97, Mac90b, Man92a, Man92b, Mat91, MO93, NC97, Nus91, Nus92, OBGGGRR93, Pro94, SW98, SH97b, SS95, SYK91, SST91, SB90a, SVRF94, Sto98, TH98b, TET91, TET97, TCLB90, Vin98, Vol90, WH92, Wol91, XCC94, ZS91, Zha97]. DNA-Protein [Sto98]. DNA. [For98]. DNAs [For96]. Do [DP94b, Kau98, Lor94b, Mar97a, MHN^{+}94, Myl99, BMS93, GS95b, Tho96b]. Docking [CS94a]. Does [CSSF93, DBH94, EL97, ISSI97, KS91, MHW93, QG92, RBH93, Sil98, SLR97, CB91a, DH92, DC99, LGSM97]. Doing [EL97]. Dollars [ME93]. Domain [SW91]. Domains [Con93, SB97, Nor92b]. Domestic [Emm97, MA95]. domestica [WH94a]. Dominance [Cam93, For94a, GMN96, Por96, Sav92, Van99, Arm90, McN95, Thi90]. Dominance-Recessivity [Van99]. dominulus [TBD95]. Don’t [Enr90]. dopa [SP90]. Dopamine [NT97]. Dormancy [LNF96, ML94a]. Dorso [SvHT97, Tak95b]. Dorso-ventral [SvHT97, Tak95b]. Dosage [For94b]. dose [FSD91]. Double [ABH94, Dem94, GCB96, MFG99, MO93, Nun98, Kan92]. Double-marker [ABH94]. Double-stranded [MO93]. Doublet [Dav98b, RWK93, Man92b]. Doublets [LH94]. Doubling [SLR^{+}98, SAGB95]. dove [Cus95]. Down [Bra96, GLO98, KQ96, Mar96, Abe98, Her95]. Dr [Fel96]. Drag [Tho96b]. Dragline [SFKD99]. Dragline-mediated [SFKD99]. Dragonfly [SZK98]. Drift [BH94, Sas94, Tyv93, Tyv94, ZL97, Pec92]. Drift/Switching [Sas94]. Drifting [EIU94]. Drinking [KB98b]. Drive [Gra93, GS93, JD99, LPB96, HG91]. driven [DMP99, Kle98, LAV^{+}97, MBC95, TM92b]. Driving [TF93]. Drosophila [LG99, Mar97b, Bod97, GK90, HKG90, KG90, Lac90, LWS97, Mar91, SvHT97, VpM95, Tak95b]. Druckery [TB91]. Drug [AWA98, ABK94, NBSM97, PG99, Dex95, JB95, LPS90, SC90]. Drugs [Liu93, Mac90a, PST91, WCPR91]. Drying [Ben97]. dT [Boe90]. Dual [FPB99, VC98, But92, Smi90c]. Dual-activity [FPB99]. Duck [HDF^{+}99]. Duct [Mul94, Mul99]. Due [LNF96, TWB^{+}93, BSGB91, HP97b, Kat90, Ste90b, Wie90]. Duets [Smi94]. Dung [RWB98]. Duplex [LSB^{+}97]. Duplexes [PG94b]. Duplication [LH91, MS98, Di 95, LH90]. Duration [NM93b, Nuy94]. Durations [Woo93]. During [Adl93, BVW97, DBH94, Hor93, MR98, Mor98, RGAP97, SvHT97, TF93, Vis93, vOPH^{+}96, AH90, BLB95, Boi95, EGDG^{+}98, JH95, KSG90, Kun95, MMH90, Mor90, MBS92, O’N91, SC92a, SK92, Wie96, ZT92]. DV [SST91]. DV-X [SST91]. Dynamic [BGS^{+}99, EL96, ISN91, JSR99, KYH99, KG93, KWD^{+}94, KQ96, LPP^{+}94, Leb96, Man90, MH97, MPAA97, Niw94, Pet99, RH96, Ric96, SEB99, VK98, Van91, VKH^{+}98, WNY98, WZN97, Whi98, Wil93a, Doe95, Her90, JFM90, LRJ92, Pet91]. Dynamical [BG96, DC96a, DC96b, DC96c, Gre97, Khr98, Niw96, ORG94, Sch99a, SQ98, Vol97, WLH94, FBH91, LOLM91, LOH92, Sza92a]. Dynamics [AOL98, AW93, And98, Ano97-28, AWA98, BB97a, BBIK97, BP99a, Boc98, BBM^{+}97, BB99b, BS98c, Bür93a, CCFS96, CB97a, CC96, CS99, CCD^{+}98, DSM99, DA98, DM91, Doe94, DDS96, EV94, Fra98, FCV99a, GWL94, GR97a, GLR96, HZ93, Här99b, HH98, JIDH97, KH98, KB99, KD99b, KIF96, LL97, LGM98, MS97a, MFG99, MMB99, MN94, McG93, MB97, MFHA^{+}95, MAB98, MS94, Min94, MD96, MA97b, Myl99, NKR97, Neu97, Niw98a, NBSM97, NS98, NH99, PH99, PK98, RWN98, RA96, RKM92, Rig94, RDHD98, Rux96, Sem96, Sin96, Smi90b, Smi98, SFGG^{+}99, SV90, SS94d, Sza93a, Sza93b, Tan94a, TL98, VS90, VW94a, VC98, WG97b, Wil99, WLJN99, WE98, Woo99, AS95, BSM90, BN92b, BBM^{+}13a, BBM^{+}13b, BBM^{+}13c, BDK91, BM90, CF92, CPLA95, DML95, FPPSK95, GB95]. dynamics [HA92, HS95, Hoc91, HE95, JP90a, JP90b, JM95, KB95, LG90a, Lib91, Llo95, McC92, PVT91, SV91, SK95, Tak95a, Tch91, Vol90, Wei90b]. dynein [Mur90, Mur91, Mur92, Spu91]. Dysfunction [CBS^{+}99]. Dyson [Ish90]. Dystrophy [KL98].

E. [ASE94, Kuz95a, SVRF94]. Early [Bal97, BSN94a, Coc98, Fed93, For91b, Hol96, HE95, MHC95, MP98a, MP99, TW99, TL98, GK90, KG90, NH95]. earth [Tor90, Coc98, Li94]. Easy [CPK96, HS91]. eating [TH98a]. Ebenman [Lor90]. ecdysial [HC93]. Echinocytosis [Won99]. Eco [PH99]. Eco-evolutionary [PH99]. Ecological [DM99, Gag94, GLO99, GE96, Hir93, LKS98, Lor94b, ML97a, Nak98, Par94, iT94, BDK91, Jør95, Ste91a]. Ecology [GLO98, Van99, WP93]. economica [Gra98]. Economics [SLR97]. Ecosystem [JC96, Tan94a, And98, FTD90, Han91]. Ecosystems [Bur94, Fle97, LKS98, Lor94b, SG98, TA99, Her90, HBP92]. Ectodomain [WRP99]. ectopic [CL92]. Ectotherms [vdHdJ96]. EDF [Gin95, Koz96]. Edge [BdS97, Wil94, Jør95, KJ91]. Editing [CC96]. Editor [Ano98d, Laz97]. Editorial [Ano90a, Ano90b, Ano90c, Ano90d, Ano90e, Ano90f, Ano91a, Ano91b, Ano91c, Ano91d, Ano91e, Ano91f, Ano92a, Ano92b, Ano92c, Ano92d, Ano92e, Ano92f]. EDRF [GE90]. EF [FR99]. EF-Hand [FR99]. Effect [ASA98b, ASA98a, Adl93, Anc99, BSW97, BB94, BH94, CMG^{+}96, CMB95, Dib98, ES95, HK98, KH99, KD93, KH94, LNF96, Nak96, PK92, PL93, Pet99, RWN98, RVMR99, Sch99a, Sme90, SC90, STI93, VS97, WS94, WPP90, WGH99, YZ97, vP97, BS91b, But92, Cam95, CH90a, CB91a, HT95, LSM^{+}92, Nak95, Nun90, Smi98, Val92]. Effective [KL99]. Effectiveness [BSTV93]. Effectors [SAH96]. Effects [Aok90, BS98b, CB97a, Che94, Coc98, CV97, FR93, GM94, GWL94, HS98a, HT97, Hin98, Hor93, JK98, Jon97, JE98, JM98, LHA94, McC92, MPAA97, PW91, PK93, PGL94, RMH96, Seg94, SJ94, Sto94, VB93, WJH97, YC99, ZL97, ZJ97, ZJ98, dLOG93, sDAB91, vOvO97, vdHdJ96, AP91b, DSD92, Kee91, LVL92, Llo95, LPS90, McN95, Sai95, Str95, Vol90, Won90, Zon97, Wan90c, Wan90d]. Efferent [DBH94]. efficacies [Mac90a]. Efficacy [ROL99, Trz99a, Trz99b, WMLK96c]. Efficiency [BSB90, Ben92, Doi97, Jon94, KD97, Rob90b, SAB97, BM95]. Efficient [PGBYP97]. Effort [Daw96, SH94]. Efforts [LS94, AC90]. Egg [KW98a, BBB90]. Eggs [BMS93, BT98, FL96, MvB93, Yos99]. egocentric [BSB90]. Ejaculate [Cor95]. Elastance [CK97b]. Elastic [BC94, Sch94, SW95a]. Elasticities [Gie94, GCB96]. Elasticity [Che94, SW95a]. Electric [BT99, Edm96, MHM94, Nov97, PK99, STI93, MR90, Meg90]. Electrical [BD94, OR97, PO91, Ste90b]. electrically [BE92]. electrogenic [LRJ92]. Electrolyte [HLT96, MB93, LW94, RB91]. Electromagnetic [MLS92]. Electron [CL93, TP93, AdG90, PT92a, PT92b]. electron-transfer [AdG90]. Electronic [SST91]. Electrophoretic [HLT96]. Electroreception [Pau95]. Electrostatic [LBDDG90, SL91a, RB91]. Element [WE98, DS92, LOB92]. Elementary [SV91, Wan90b]. Elements [BB97a, CV93, Dwy98, HS97, Kat96, QF94, Smi98]. Elevation [LD94]. Eliminate [Zon96]. Elimination [ABR97, ABR99, SP99, SSS93, WGH98]. Elitzur [Yoc95]. ellipsoidal [PO91]. elongation [BH92]. embedded [Van90]. embodying [JR92]. Embryo [Bod97, VKH^{+}98, VHSC92]. Embryonic [SvHT97, PAS91, ZCS^{+}90]. Embryos [BA97a]. Emergence [AP96c, Eli96, GL95, JIDH97, LKS98, Li94, MS98, MA97b, NSW97, PD92]. Emergent [CSB98, Thi93, KJ95, Mil92b, dBP91]. Emerging [ND99]. Emitted [Baj99]. Emphasis [PBJ^{+}92]. Empirical [Han91, McK95]. Enable [KDBK93]. Enabling [MS98, Mag95]. Encoded [Jai97, Say94, NAZG95, NH95]. encoding [Kel95b, LG90b, Wil92b]. encrusting [Grü95]. End [Dav98b, SGP98, Ste94]. End-atom [Dav98b]. End-capillary [SGP98]. End-to-end [Ste94]. endings [OHS95]. Endocrinologist [KV93]. Endocytosis [AKMS97, Koc95]. Endogenous [CCK98]. endothelial [SL91c]. Endplate [FT93]. Endplates [MKGB98]. Endurance [PM96]. Energetic [DRB97, Par94, SM97d]. Energetical [Kuz95b]. Energetics [KD93, Koo95, San99, Ste90a, Wes91, PAS91]. Energy [Bal97, BSS96, Bla94, BC94, Cha93, DY99, Hag93, HA97, Hir93, JC96, KD97, KL96, KKH99, Kre94, Lor94b, Sch91b, Sto98, TT93, TP96, YY94, Her90, Kre92, Mar90, PT92b, SOR91, SL91a, SH91a, Sol95, WW92, ZS91]. Engine [KH98, Koe95]. engulfment [Aga95]. Enhanced [Meg90]. Enhances [Tan96, FD92]. enhancing [JB95]. Enrichment [Com94a]. ensembles [BPHH90, Won90]. Enslave [SRH97]. Ensures [WJ97]. Ensuring [ORM99].

Enterococcus [Koc92]. enterokinase [VHV^{+}90]. entities [ZW95]. Entrainment [BBPD93, LWS97, TUHT97]. entrant [BHP92, CL92, ZWH98]. Entropic [OBGGGRR93]. Entropy [Aok91, Aok92, ASML97, Ber95c, Hum92, LPBR96, Rig94, SH97b, Aok90, CCRM90, RCCM92, Yee91]. Entropy-controlled [Hum92]. Entry [McN94b, HZ95, PH93, ZH97]. Envelope [Tho96a]. Environment [AH94, CH94, Ezo98, GAD97, GG97b, KK94b, KRG95, McN94b, MGR97, MTM^{+}96, ML97b, Nis99, RP98, Sak97, Tan96, Web99, Ben92, CG91, Fer92, For91a, GD90, Nis92, Sak92, Sak95, SS90b, Van92a, Van92b]. Environment-Dependent [MTM^{+}96]. Environmental [PC97, Won90, dLOG93, CC95, Gro92, Hai91, McA91]. Environments [BN96, Dav98a, Iwa93, KKN98, LJ96, MSN96, Niw98a, Pay97, RBP99, SH93a, Sak96, WNY98, WT94, AC90, CH90b]. Enzymatic [BMG98, HK96, Kal98, MR98, HH91a]. enzymatically [TB91]. enzymatically-altered [TB91]. Enzyme [Bri93, Dem98, EEG93, FK98, GMN96, KH94, KV93, Pet96a, Pet96b, Pet99, Rou98, SM97a, SS96a, YYYU97, ZW96, Bat91, BS91a, BS91b, Cár91, FO92, Has92, Hav91, LOH92, MS90a, Maz91, MK92, Pat95, Sal91, Sol95, Top90, VHG^{+}90, VH90, VHGV92, WK90, Wan90c, Wan90d, Wan90e, WKSH92, ZT92]. enzyme-catalysed [Top90]. enzyme-enzyme [Cár91]. Enzyme-substrate [Dem98, Sol95]. Enzymes [KV93, SL98, WBS97, ABW90, Bar90, Fri91, GC91, MHTS90, Pet91, RKM92, Sri91, TS91]. Enzymic [Rak97, HRW93]. Enzymological [SM97b]. enzymology [Wan90b, WK90, Wan90c, Wan90d]. Epidemic [CDG^{+}98, LM96, ONRH96, RA96, WJ97]. Epidemics [DSD93, Joh96]. Epidemiological [Dus96, KD99b, Ons92]. Epidemiology [BRT98]. epidermal [Hoa90]. Epidermis [SPS93]. Epigenetic [Hol99, Kel94, PM99, Kel95a]. epiphenomenon [Hem91]. Epistasis [SBRG99, Tak95b]. Epistatic [Sil98]. epithelia [Mor92]. Epithelial [CLW93, Cum96a, GMST99, LBP^{+}97, Mor99, BL91, DF91b]. epithelium [SRW95]. epitope [VK91]. epitopes [NMS95]. epoxide [Mac90b]. Equality [Kur99]. equalization [MH90]. Equation [HE93, MB93, RL94, TW97, VR94a, YYYU97, HA92, Hes90, SL91a]. Equations [CPK96, CB94b, Kab94, LR94c, Ste94, WZ96, Zha90, FF90, FTBC92, LH90, LL90, Van91, VHG^{+}90]. Equilibrations [Cze93]. Equilibria [MP93, Mil92a]. Equilibrium [BLGG^{+}93, BW98, Che94, Com94a, HMEW96, Kem94, RMH96, Seg94, SH93b, Bla95, CKR95, Cra90a, DD98]. Equivalent [Ham98]. Eradication [BRT98, WDP98]. Errata [Ano90g, Ano90h, Ano90i, Ano91g, Ano91h, Ano91i, Ano92g]. Erratum [Ano90j, Ano90k, Ano91j, Ano91k, Ano91l, Ano92h, Ano92i, Ano93a, Ano94a, Ano95a, Ano96b, Ano97c, Ano98c, Ano99d, Ano99e, Ano99f]. Error [BS93, FSPC98, Kok96, KK94e, MO93, SNT95, WBS95, FPPSK95, HH91b, JB95]. error-enhancing [JB95]. Errors [AB98, MWC^{+}97, Pet96a]. Erythrocyte [AKMV96, CMB95, HBJ^{+}97, Won94, Won99, Sax92]. Erythrocytes [AKV96, AKMV96, LL93]. Erythropoiesis [MBM98]. Escalation [HS96, PM96]. Escape [BN94, MK97b, Lim90]. Escherichia [AKL97, BDA99, GRRF93, HLG92, HB91, KKV95, Kun92, LL96, Mah93, MBSB95, Nor92b, Nor93, OWM96, RWK93, Ric96, SP92d, Say94, SG90, SLR97, VP93a, VP93b, Yar96, VK98]. ESS [BPHH90, Fra97c, FCV99a, HT95, Hin98, Lib91, Mot91, RWB98, VV96, vdHH90]. ESS-analysis [BPHH90]. Essays [BLLÅ94]. essentials [Hah91]. Establishement [SvHT97]. Establishment [Rod96]. Esterolytic [Hav99]. Estimate [CAPP99, Woo93, FSDT91, vdGvdBB92]. Estimated [SVRF94, Mac90a]. Estimates [ELS97, SC92a, Don90, HH91b, WKSH92]. Estimating [FBS93, GRUH99, Mar97b, Moc99, SSCvdM99, SH97b, CB91b, FTL^{+}90, TH90]. Estimation [CK97b, Rog94, SSCC93, Jol90, JD92, SWA91, Val92]. Estimator [TL96, CK95b]. Ethylene [GPL98, WAO^{+}99]. Eubacteria [HS97]. Euchromatin [SW96]. Eukaryotes [ONT97, ALM95]. Eukaryotic [AMO93, NAB^{+}94, SSM93, NS91, Nus92, RH92]. Euphotic [KKS97]. europaeum [KKA96]. European [SAM97]. Eusocial [Yam93, KSG90, RR92, TBD95]. Eusociality [Ano94b, Hur97b, Hai92]. Evading [Sas94]. Evaluate [WM98]. evaluating [FC95]. Evaluation [CBFR98, Coc98, KD97, LRZ^{+}96, NT99, WGH98, WW97, BC94, BS91b, LW94, MPRB91]. Even [Wes93]. Even-aged [Wes93]. Events [HP97a, Kor93, Nak99, IL95]. Evergreen [Wal99b]. Everybody [EL97]. Evidence [CS94a, GBD99, JLL92, Jam96, Jam99a, Jam99b, KM96a, KM96b, SLR^{+}98, Spi91, SAM97, Ute93, Yar96, Pet91, VpM95]. Evolution [ASE94, Ano94b, ASML97, AF99, AKMS97, Bal97, BBL^{+}97a, Bar94, BRV99, BK95, Bea93, BN97, BA94, BG98, BP99b, BS94, Bla98, Bon99, Bow99, BKD99, Bur93b, CCVH98, CST98, CMS93, CW98, DD98, DCM94, Dwy98, Eld91b, Eli94, EFB98, Fed93, Fra94, Fre96, FW97, FO92, FOO94, FO96, Gan99, GM98, GG97a, GR97b, GMN96, HS98b, Hur97b, Iwa93, IP99b, Jab94, JS97, JW95, Jon96, KS96b, Kel96, KD99b, Laz97, LM94, Lin97, Lun90, MI98, MN94, MH96, MHWM94, Mel93b, Mit96a, MP93, MP98a, MA98, Nak98, NMI97, NNI98, Ned97, OWM96, Pam91, Pau93, Pec93, PGBYP97, PB97a, Pro94, QAM97, Rat91, RH96, RWN98, RRG98, RN96, RBD97, Sas94, Sau94, SRH97, SR98, SR99, SC93, Sjö97, SMS98, SJ94, SS93c]. Evolution [Tak95a, Tak98, TH98b, TF99, TéS96, Vrb98, WW98, WN99c, WT94, WH96, Yam99, Yoc95, vBR98, vBW98, And95, BN95, Ber91, Ber95c, Bla95, BW91, BBB90, BR91, Bur92, Cha91, Dug90, DH95, FR90, Fin92a, FM91, For90, For91b, Fre90, FD92, GH90, Gos92, Hai92, Hur90, Hur91, HN91, JG90, Kau92, KM95, Kel95c, Lah91, LN95, LB95, Mag92, Man95, MHM91, NH95, Nor92a, Now91, NC92, Pec90, Pec92, Ric90, Rob91, SHB90, Ser91, Sri91, SHMH90, Tan91, Tre95, TET91, Won90, Wor95]. Evolution-the [ASML97, Man95]. Evolutionarily [BP93, Daw98, FC92, GM99, GV98, Har99a, HT97, HMW95, Lim90, Lor94a, MH90, MWC^{+}97, Mow97, TS95, Wit95, YQsZgZw97, YZ97, ZJ97, Zha98, ZJ98, dLOG93, BR90, Cra90b, Cre90, MS95b, Now90, RG95]. Evolutionary [Alm99, AFM97, BEE97, BSN^{+}94b, Bea90, BB99b, Bro97, CV97, Daw96, Dem98, DML95, Doe95, EMS97, GL95, GR97a, Här99b, Her94, HKH93, IWH^{+}99, JSB93, Kal97, Kal98, KD98, Kim93, Kis99, KH94, Koe97, LTZ96, Mar93c, MC93, MK97b, McN93, McN94a, NC97, NPK99, RHS^{+}95, Sea99, Sjö97, SBRG99, SNH99, Van99, Woj92, Yam93, YZ94, Bro91, Cra90a, FS95a, Fra95a, GD90, HH91a, LT92, Mat95, PH99, PV95, RdM91, Rob90b, Wie90]. Evolutive [TéS96]. Evolving [CCG^{+}98, LR94b, PB97b]. Exact [Tyv93]. Exactly [MK97a]. Examination [DF91a, KB98b, Leb96, WBC99, Lor90]. Examined [ES95]. Example [AF99, GLO98]. excess [ZT92]. Exchange [AVM^{+}99, ABK94, FR93, JLK92, SSSF95]. Exchanger [VS97]. Exchanges [ODA97, SR98]. Excision [AKL97]. Excitability [VR94b, LF95]. Excitable [BD94, Mur91, NH99, RAGO99, Cha90, CL92, CLF95, CJ90, CJMBM92, EMRS90, Mur90, Ste90b]. Excitation [AP96a, ZWH98]. Excitatory [MKZ97]. Excited [BT91]. Exciton [CC97c]. Exclusion [DP94a, SAM97, Pec90]. exclusive [Jär95]. Exemplified [OKAP98, Vrb98]. exercise [Aok90, Ho90, SC92a]. exerted [Mor91, TS91]. Exhaustive [Boc98]. exhibit [Rux95]. Existence [AW93, KK94c, HZ93]. Existing [KD97, FO92]. exists [Fel95]. Exocytosis [CB96a, GMU95]. exon [XCS^{+}95]. Exons [CL98, Dwy98]. Exonuclease [Say94]. Expands [Tho96a]. Expansion [Bje94, Bje96, FK97, KIF96, ORM99, WR93, Koc90, Sak92]. Expansions [BKA97, Fin92b]. expect [Enr90]. Experiment [Rot98, THF92]. Experimental [AB98, FLR93, HCB96, KQ96, KGFB97, MBSB95, Pet96a, RS97, SC99a, JLK92, KA91, RV90, Yee91]. Experimentally [TF94]. Experiments [CCD^{+}98, HP97b, Jon94, Mar97b, RP96, SK99b, FR90, Mar91, SH90a]. Explain [Hes94, HE93, Kok97, Lin94, Whi97, MGL^{+}95, UJ95]. Explained [KL96, LBP^{+}97, Tan99]. Explaining [Yam93, JH95]. Explanation [FK97, SSA94, Spe90, WK90]. Explicit [Neu98, UP96]. Exploitation [HMW95, DM91, Nis92]. Exploiter [Gav97, HL96a]. Exploiter-mediated [HL96a]. exploiting [Sib95]. Exploration [Nis99, SSA94]. Explore [VV96, Ing90]. explored [JR92]. Exploring [OP96, WT94]. exponent [Fel95, Spe90]. Exponential [SK99b]. Exponentially [CK96]. Export [KL98, FAF90]. Exposure [CCK98]. Expression [All93a, CK99, Ham93, IP99a, KDBK93, LL96, LNF96, Mah93, BLB95, BHVC90, GK90, HT95]. Expulsion [WS95]. Extant [FVB97]. Extended [Bas98, Her94, SBV92]. Extended-spectrum [Bas98]. Extending [GCB96, HS98c]. Extension [RBR90, Cus95, Han92]. extent [FTL^{+}90]. External [BP96, BP97a, MHM94, NMdlR93, SAH96]. Extinction [Bje94, Bur99, CDG^{+}98, Flo98, Kok96, LL99, New97, RN96, SG98, Tan96, GLO98]. Extinctions [GM99]. Extra [Wag93]. Extra-pair [Wag93]. Extracellular [Com90, DSM99, HE94, Leb96]. Extracellularly [FHB^{+}97]. Extracellularly-derived [FHB^{+}97]. extrachromosomal [How92]. Extraembryonic [VKH^{+}98]. extrapolation [TWW90]. Extreme [Koc93, JR92]. Eye [BM98, HC93, LSG97, TWB^{+}93].

Fabricius [THF92]. facet [Hav91]. Facilitated [VGV93, Wil95]. Facilitation [MD96, Kum91]. Fact [GC91, Mar97a]. Factor [KKS97, MPSK94, Hoa90]. Factors [Ale97b, BB97a, Kel94, MBS93, NVB99, Kel95b]. Factory [CLMH^{+}96]. Facts [KS99, OVH^{+}96]. Faculatative [SMS98]. Failure [MB92]. falciparum [MB97, MB98, MMB99]. Families [Hol93]. famine [DSD92]. Far [LL94, MS95b]. Fast [All93b, AP96c, Cze93, KPS97, QF94, dLA97]. Fat [KD93, Pen98]. Fatty [BDK^{+}97, KQ96, VGV93, FSD91]. Favoring [JW95, LR94b]. Favour [Ng96]. Favoured [Pet96b]. Fc [BLGG^{+}93]. Fc-receptors [BLGG^{+}93]. Feasibility [Ben99]. Feathers [DVJ^{+}99, Fed93]. features [LRJ92]. February [Ano90-30, Ano90-37, Ano91-36, Ano91-40, Ano92-30, Ano92-35, Ano93q, Ano93w, Ano94p, Ano94y, Ano95m, Ano95s, Ano96o, Ano96t, Ano97q, Ano97w, Ano98q, Ano98u, Ano99v, Ano99-27]. feed [FTL^{+}90, PBJ^{+}92]. Feedback [BS98a, BH94, BM98, CMML96, DvH93, DSD94a, Gre97, KPS97, LKS98, MPFHG96, OP94, Rob91, Sea99, Thi90, VB93, GM90, NLS91, NSL91, Tor96]. Feedback-controlled [BM98]. feeders [GS95a]. Feeding [TAA^{+}98, Grü95, HH95, Hou90]. Feeds [FDT^{+}93]. Feline [LRZ^{+}96, CPLA95]. Female [HW98b, JM98, Kok97, MDGB94, MH96, Nor92a, RWB98, Wag93, Cor95, Kir92, SM92, SG95, Wie90]. Females [BMS93, SH94]. Femina [Gra98]. Femoral [Mul93a]. Femoris [MPSK94]. Femur [Mul93b]. fermentation [YI95]. Ferroelectric [BLM94, THS^{+}95]. Fertility [Gra98, Jos97, NNI98]. Fertilization [BP96, BP97a]. Fertilize [BMS93]. fetoprotein [Miz95]. Fiber [DSM99, TGS93, WG92, HC91]. Fiber-composite [WG92]. Fibers [VW94a, Mor90]. fibre [Mor91]. fibrinogen [BSGB91]. fibroblast [TM92b]. fibroblast-driven [TM92b]. fibrocartilages [De 95]. Fibronectin [Wie96]. fiction [GC91]. Fictions [KS99]. Fidelity [BP99b]. Field [BP93, Hie97, Lib96, Lib97, LÅ94, LÅ96a, MHM94, MTM^{+}96, PK99, STI93, Sza93a, BT90b, BT90a, Cra90b, Cra90a, Meg90, Mot91]. Fields [BT99, MR90, PO91]. Fifteen [Por96]. Fight [Nuy94]. Fighting [Hur97a]. Fights [GG98]. Filament [CF94b, MC96, ST93, HKT91, Mor90, Smi90a, Smi90b]. Filamentous [MCG95]. Filaments [MBSM93, SC99b, MB90]. file [FH99]. Filling [Adl93]. Film [HE94]. Filopodia [RS97]. Filopodia-based [RS97]. Filtering [FKJ^{+}94]. filtration [Med95]. find [BS95]. Finding [Kok96]. Findings [MBSB95]. Fine [For95]. Fingerprinting [WW97]. Finishing [NTCN^{+}99]. Finite [AH98a, DS92, SC90, Tyv93, Tyv94, Cra90a]. finite-population [Cra90a]. Fire [CSB98, BS98c, PCN95]. Firing [BS98c, Gaá93, Mat96, OHS95, Smi92]. firm [Leo92]. First [Ste94, SS97, BW92, FC92, MS90a, KPS97]. first-order [BW92, MS90a]. Fish [AHB^{+}91, BEE97, CB94a, DK95, EWHB94, EPT97, Kat96, KS98, KIA93, MCGPP97, Niw94, Niw96, Niw98a, Niw98b, TMI98, Dug90, ECB91, Gal92, HW92]. fisher [Gra90b, GV99]. fishery [Hou92]. fission [NT95]. fit [Pau93]. Fitness [AH96, AH98a, AH98b, BS93, Che93, GV99, GG97a, HW98a, Jon96, Nak98, SBRG99, SG93, SJ94, Sza93b, WS93, WH96, dLOG93, BK90b, Bur92, KJ91, QG92]. Fitness-sets [dLOG93]. Fitnesses [Koe94]. Fittedness [Bur94]. Fittest [AH98a, LL99]. Fitting [YCM97]. FitzHugh [Hor93, RL94]. five [Smi90b]. five-state [Smi90b]. Fixation [Kel95a]. Fixed [WM98, Gra95b, LGBP92]. fixed-delay [LGBP92]. Flagella [Lin94, Sch94, Mur91, SH90b]. Flagellar [Nab96]. flammability [PCN95]. flanks [NS91]. Flap [Tho96b]. Flat [SEB^{+}93, Hor95]. Flawed [GS95b]. Flexibility [Dem98, Eas91, Hum92, Nus92]. Flexible [NG99, Arm90]. Flies [SP99, Wil94]. Flight [Hed92, KIA93, LHA94, Tho96a, YA97]. flock [KB92]. Flow [BFK^{+}93, Cha98, FOO94, HFS97, HE93, JC96, LPP^{+}94, Lor94b, LT98, PD93, SC96, DB95, SF90, Sol95]. Flower [BSTV93, Gou94, MAB98, Sak97, FS95b]. Flowering [RGMG98, Ing90]. flowers [PSE92]. flows [Her90]. Fluctuating [Che93, ELS97, LJ96, MP93, RP98, WNY98]. Fluctuation [Jon94]. Fluctuations [MH95, MFB98, NMdlR93]. Fluid [AKMS97, HE93, Woo99, SF90]. Fluid-phase [AKMS97, SF90]. Fluorescence [Dem94, PN98, SGSS98]. Flux [ACC96, Gie94, GMN96, KN93, MRL96, Pet96a, Sal96, TF96b, Boi95, WCW95]. fluxes [SP92c, SP92d]. Fly [RWB98, Hou90]. Flying [KIA93, LD94, KB92]. foals [MA95]. foci [TB91]. fold [Daw95, TM92a]. Folded [PG94a]. Folding [CF94b, CZ93, GRRF93, Nöl98, Nöl99, Rot95, Sun95b, Tho95]. Folding-Studies [Rot95]. foliage [FAF90]. Following [CSSF93, KE92, WT90]. Food [CSB98, DVV99, DEK98, EM95, Emm97, HH98, JE98, KK94d, NRD94, ND99, HBP92, MA95, vdGvdBB92]. Foods [JD99]. Footprint [Fer94]. footprints [AHB^{+}91]. Forager [BSTV93, Nis99, PH90]. foragers [Lim90]. Foraging [CH94, GHW99, HC94, Nis92, PSE92, WEK95, Wit95, CG91, LÅ90b, Mil92b, MS95b, TPAS95, YI95]. Force [LD94, Lib97, LÅ94, LÅ96a, MM95, SQ98, TF93, Mat95, Mor91]. Forced [Ves98]. Forces [BF99a, Got94, MM97, ST93]. Forcing [Smi94]. Forecast [RWK93]. Foreign [PA98, SL91b]. Forest [BBIK97, KIF96, KKH97, MS97a, Wes93, Tri92]. Forests [SKI98, ISN91]. Form [BJ96, Cum96b, HC93, MBSB95, SM97a, Spi93]. Formal [HD97, Kaa94, Ker96, Kor93, Gra95b]. Formalism [LR94c, SP92a]. Format [ZLYZ98]. Formation [AP99, BVW97, CMML96, CS96, Cum99b, DBB99, DC96b, KMJ98, KM97, Kle98, Koc95, Kor97, Lub97, LL93, MS97a, Myl99, NYI99, OS97, Rak97, Sac94, TMI98, TMI99, VSW97, Wag93, WEK95, WLM96, Wie96, vOPH^{+}96, AP91b, BBB91b, Bla92, BT90b, BT90a, CSJM90, Com90, Cum90, DSP92, Hoa90, Koc92, KCM90, KB92, LQW92, LG91, MM91, Sav95b, Tak95b]. Formed [KL98, NMdlR93, JH95, NS95]. Formica [Ryt90]. Forming [ONT97, Pet91]. Forms [MHC95, RP99, Sza94]. Formulas [NVB99]. Formulation [MAC99a, VS97, VHSC92]. formulism [Zhu91]. Fossil [FVB97]. Fossils [CF94a]. found [Cár91]. Founder [RB97b]. Four [Ber95b, Hag93, Mul93a, MdR98, RJ96, áSá96, Ser91, KG90, Shc94, Spe90]. Four-bar [Mul93a, MdR98]. Four-compartment [Hag93]. four-dimensional [Spe90]. Fourier [FS94, SC90]. foxes [DM92b, MS92b]. Fractal [Bur93b, FPPSK95, Has94a, Lon94, VKH^{+}98, WS93, XCS^{+}95, XCC94, Zam99, Ber91, Bur90, CY90, Sav95a]. Fraction [But90, GW95]. Fragile [BKA97]. fragment [JV91, WAJ91]. Fragmented [Bur99, GAD97]. Fragments [MAGO97]. Framework [FK99, MA98, MPO95]. Francisco [Phi99]. Free [BSS96, Bla94, DY99, FB97, KD97, KK94e, LWS97, NBSM97, RM97, SVRF94, Sto98, YY94, BB91, CKR95, CB91a, KCM90, Nev90]. Free-living [KD97, BB91]. Free-running [LWS97]. Frequencies [BSGB91, RWK93, WR94, SYK91, Zha90]. Frequency [BS98c, CMPR93, Dru93, HL96a, LG90b, LNF96, MR90, PK93, YZ95, Yi98, FD92, Nak98, PO91, QG92, Rat90, RG95, SGG95, SH90b]. Frequency- [YZ95]. Frequency-Dependent [HL96a, Yi98, QG92]. Freshwater [Kat96]. Friction [Col93b, EIU94, IET99, MHM94, Sud90, TS91]. Friendship [Pec93]. Frog [BT98, CA92]. Fröhlich [LVL92]. Front [GL93]. Frustrated [BC97]. Fuel [HA97, KL96, Hed92]. Fugitive [HZ93]. Fuji [AH96, AH98a]. Fuji-Type [AH96, AH98a]. full [SL91a]. Function [BDK^{+}97, BDDF97, Blo97, Cha93, CFC98a, Emm97, GSS98, HDF^{+}99, OES93, RM93, Sin96, SK99b, TéS96, TT93, WB99, BSGB91, But92, DP95a, For95, Gli91, LOB92, Spa92]. Functional [HD97, LKS98, LÅ90b, ME93, ONT97, SNN^{+}97, Woo97, DC90, RKM92]. Functioning [Yar96]. Functions [Ber93, IWH^{+}99, BK90b]. Fundamental [CJ90, GV99, Han92]. Fungal [Dav98a]. Fungi [Koc94b, FTD90]. Further [FCV97, Jam99a, Kal98, MMH90, MBS92, Sau94]. Fusion [Eas96, KG93, KV93, PSLB99, Rau94, GMU95, RB91].

G [For96, For99, O’N91, Jär95, LRB94, O’N91, ROL99, WBS97]. g-protein [Jär95, AOL98, ROL99]. G0S2 [For98]. gag [Gla91, Liq91]. Gaia [Mar95b]. gait [SJK90]. Gaits [MA97a, SJK90]. Galvanotropism [RS97]. Game [CV97, DT98, KD98, Koe97, Kok97, MI98, MHWM94, MA97b, Mot93, Myl99, NPK99, VRH93, Cus95, JG92, OBGGGRR93, PV95, Sak92]. Game-theoretic [VRH93]. Games [BP96, BP97a, BP98, Her94, Lei97, Pos99, Szá99, BK95, Dug90, Hur95, Kim95, Rob90b]. Games-Memory [Pos99]. Gamete [JM98, JM95]. Gametic [Var93b]. Gametocytes [MB98]. Gamma [EL96, FTCM90]. Gap [All93c, BSW97, KM97, KIF96, KKH97, Lac90, MS97a, CM95, Kee91]. Gap-Junction [KM97]. Gaps [Sax92]. garden [GV91]. Garfinkel [Koh91]. Gas [FDT^{+}93, LVKH94, SSSF95, WGH98]. Gaseous [AVM^{+}99, ODA97]. Gases [TWB^{+}93]. gastrointestinal [VFS92]. Gastrulation [Spi93]. gated [LC97]. gates [Sta90]. gathering [FS95b]. gating [MLS92, SM90a]. gave [Di 95]. GC [SST91]. Gel [HE93]. Gender [FAB92, KK94b]. Gene [All93a, ACC96, AP96b, ABK94, BN97, Big93, CK99, Dam99, For98, Ham93, HB91, KG93, KDBK93, LL96, LNF96, Mah93, Mil91b, SW98, Say94, WE98, Boe90, Bur95, Di 95, Fin92a, GK90, Iid90, Ko91, LWHZ91, MPO95, Sai95, TET91, UJ95]. gene-culture [Fin92a]. Gene-for-gene [Dam99]. genealogical [GW95]. General [BP98, FB97, Has97, KW98c, MWC^{+}97, NH99, Wat97, WH97, YYYU97, Zhu91, BW92, DP95a, FS95a, Mil92a, O’N91, ROMM90, Van92b]. generalist [CW92]. generalization [MHTS90]. Generalized [BJS97, Sch96, MPRB91, PVT91, Tak95b, Tch91, Top90]. Generate [TMI98]. Generated [KMM^{+}97, SNN^{+}96, SNK^{+}99, FO92, IET99]. Generates [DSD94a]. Generating [EL96, Nak98, Nak99, Rou98]. Generation [CCK98, Eug95, KD97, Mor99, RWDM98, LF95, LGBP92, Mor91, Sta92]. Generations [KD99b, Sib94]. Generative [BL91]. generator [KB98b]. Generators [MM95]. Generosity [SR98, Gri95]. Genes [ASE94, AMO93, AM96, AM97, CMS93, IP99b, Wed94, ZLYZ98, AM90, ALM95, Eld91b, FO92, Gra95a, Kyp90, Now91, SHB90]. Genet [Sak94]. Genetic [BN96, BN97, BA94, CZ93, CV93, Dav98b, DT98, DCM94, Di 97, Di 98, Doe97, Doi97, DM99, Duf97, Dwy98, EFB98, For94a, FK97, FOO94, GVL93, HG91, IWH^{+}99, JIDH97, KKN98, KKN99, Koc93, KL97, KV97, MH96, MAB98, Mit96a, NSW97, PM99, RH96, Rig94, Sch97a, SV97, Shc93a, Shc93b, Shc94, Smi98, SW95b, SHMH90, Sza93b, TL96, Tyv93, Tyv94, Var93a, Ver94, Wil93b, Yam93, AM92, CV91, Hol91, Kel95b, Mil92a, Pec92, PVT91, RV90, Tch91, VGP95, Won90]. Genetical [Ric95]. Genetically [WW94]. Genetics [BRT98, BA97b, Fra94, Fra99, Hai93, And95, BBB90, Fra95a, Has92, Vic91]. geniculate [McI95]. Genome [BKNR97, HS97, MJLW96, Kon95]. Genomes [All93a, HK99, MK97a, NS91]. Genomic [IP99a, KDBK93, LPRB96, MJLW96, Rig94, WE98, Koz91]. Genotype [FS98, KS96a]. Genotype-phenotype [FS98]. Genotypes [MK97a]. Genu [MPSK94]. genus [MWAS91]. Geographic [Gag94, NNM99]. Geometric [Koz96, Lin94, Gin95, Ric90, Spe90]. Geometric-Normal [Koz96, Gin95]. Geometrical [Cum96a, GS93, PP92, PS93, GE90]. Geometry [Bur93b, Gri94b, KKA96, Koc95, MdR98, NH99, Rot95, Tho96a, XCC94]. George [Fra95a]. Germ [Sza94]. Germinal [PGBYP97, RDHD98]. Germination [Hag93, Und90]. Get [Laz97]. GG [SST91]. Giant [KL98, NYI99]. gibberellin [WH92]. gibberellin-DNA [WH92]. gibbons [RD90]. gift [Leo92]. gill [MR91b]. Give [PDP97]. given [But90]. Gives [Gre98, Alm99]. Giving [Nis99]. Giving-up [Nis99]. Gland [FHB^{+}97, MFM97, FBL^{+}95]. glass [Won90]. gliding [Koc90]. Glimpses [Ace96b]. Global [BJ96, CFC98a, NRD94, PLB96, GR95, Van91, VS90]. Globin [Eld97, Boe90, Iid90]. Globular [CL93, MP91]. glomerular [Med95]. Glomerulus [CLW93]. Glucose [CG96, HRW93, Rog94, QC91b, RBR90]. Glucose-induced [CG96]. Glycation [HRW93]. Glycogen [CG96]. glycol [RB91]. glycol-induced [RB91]. Glycolysis [NSH97, AGG91, BS91a]. glycolytic [BS91b, GC91, Smo95]. Glycosidase [Nak96].

Glycosidase-Catalyzed [Nak96]. Glycosidases [FPB99]. Glyoxalase [Kal98]. Go [Gui94]. Go-slow [Gui94]. Gompertz [Eas99, SK99b]. Gompertzian [Abe98]. gonadal [BZGM91]. gonadoliberin [Gla91]. Good [IP99b, RWK93, GN95]. Goodness [Pau93]. Goodness-of-fit [Pau93]. Goodwin [RR96, RVMR99]. graded [TCE95]. Gradient [DPG^{+}94, WH94b]. Gradients [Hon98, Yos99]. Gradual [MP93]. Grafen [ES95, Kir92]. Gram [KW94]. Gram-negative [KW94]. Graph [BDA99, IWH^{+}99, KL91, Koh92, Pat95]. graph-theoretic [Pat95]. Graphic [Mog93, SRW95, YH90]. graphs [PSE92]. Grass [Kor97]. Gravitaxis [MNB97]. Gravity [SEB^{+}93]. Great [Jam99b, SB94, AP91a, SB92, ZT92]. greater [Gri95]. Green [KKS97]. ground [LJ92, Wes93]. Group [BF99a, BTD98, BR90, GL95, Pec92, WE93, BBB91a, BBB91b, Boy95, DM92b, Hem90]. Groups [Bla97, Koe90, PA90, WL91]. Growing [PGF94, YY99, Sak95]. grown [MS92a]. Growth [AP96c, AWA98, BA97a, BM98, CC97a, CC97b, Com94a, CB97b, Cum96b, DH98, Eas99, GPL98, HVKS99, Han93, HW94, Jon94, Kaa94, KD93, KG93, Koc94b, KD99a, LQL94, Mat96, MTS97, McN94b, ML94a, NKR97, Nar98, PS93, QZDA93, RLRV97, RLF93, SH93a, SS93b, TE96, Tab98, VP93b, Vrb98, WH94b, Wes93, WOQP93, ZBB97, vdHdJ96, Aok91, Ber91, FB95, FF90, FAF90, Gro92, HB91, Hoa90, Jør95, Kaa95, KH91, KA91, KDG92, KE92, LdQW95, LG91, MS92a, PAS91, SS95, Van92a, Van92b, WG92]. Growth-promoting [GPL98]. guanine [NS91]. Guarding [FCV99b]. Guild [HL98b]. Guilty [DBAP^{+}98]. Guinea [And98]. gyrus [AMT91].

H [BK97, Hav99, NK90, Rad91]. H-chain [Hav99]. H-Y [BK97]. H1 [Fli99]. Habitat [BS98b, GL95, Har99a, Neu98, SH93a, Sak94, SA93, BBB91b, Boi95]. Habitats [Bur99, Has94a]. Haemophilus [LPRB96, RDHD98]. Hair [SSGR^{+}98]. Hairpin [Di 94, Di 95, TCLB90]. Hairpins [KL98]. Half [AP96b]. Half-Life [AP96b]. Halobacterium [Nab96]. Hamadryas [GG98]. Hamilton [Sib94]. Hand [FR99]. Handedness [MB92]. Handicap [IP99b, Sil98, Szá99]. Handicaps [Wed94, Gra90a]. handshake [Rob90b]. Hanson [TW97]. Haplochromis [Gal92]. haploid [FR90]. Happen [ISSI97]. Hard [SEB^{+}93]. Hares [BSM97]. harmonics [GK90, KG90]. Harpoon [Kuz96]. Harvester [CH94, HC94]. Harvesting [ELS97, AHL95]. Hatchability [HDF^{+}99]. Hatching [MvB93]. having [Bar90]. Hawaiian [KM96a]. hawk [Cus95]. Hayflick [PMA97]. Head [MC96, MKC^{+}98, Mor91]. Heads [Mor96, MM97, MB90, MBSM92]. Healing [GMST99]. Health [PAG99]. Heart [AP96a, Wie96, WLH94, FBH91, SCG91, ZCS^{+}90]. Heat [DVJ^{+}99, For94a, KGFB97, RVMR99, WHL93, Boi95, LL90, MR92, PAS91]. heat-induced [MR92]. Heat-shock [For94a]. Heavy [HS96, KL96, OHS91]. Hebb [Ada98]. hedging [Hut99]. Height [MS92a]. Helianthus [Han93]. Helical [Kat96, WG92]. Helicid [AC96, Leo92]. helicobacter [KB95].

Helix [BKK96, KK96, WBS97, ZZ99, Kan92, Sun95b, TSC90]. Helix/Strand [ZZ99]. Helper [AL96, BGE94, MK90]. Helpers [Mot93]. helps [GE90]. Hematopoietic [MBM98]. Hemodynamic [Adl93]. Hemoglobin [CLY99]. Hemoglobin/O [CLY99]. hemolysis [Del95]. Henrik [Fel96, CB96b]. Hepatic [ABR97, ABR99, Rog94]. hepatitis [MPRB91, MRB91, PNB92]. hepatocellular [PNB92, Ste91b]. Hepatocyte [AVK94, SK92]. Heptapeptide [Rot95]. Herbaceous [Sek95, Sak91]. Herbivore [TAN94b, GH98b]. Herbivores [GL95, LÅ90b]. herbivorous [YI95]. Herbivory [Sto94]. Herbs [Sak94]. Herd [MHN^{+}94]. Herds [GL93, GLR96, Hut99]. Hermaphrodite [Leo90b]. Hermaphrodites [JM98]. Hermaphroditism [CCG^{+}98]. Heterochromatic [CMG^{+}96]. Heterochromatin [SW96]. heterochronic [GS95b]. Heteroclinic [Sta96a]. Heteroduplexes [DLC96]. Heterogeneity [CDG^{+}98, LM96, PC97, KJ95, Llo95]. Heterogeneous [BOG^{+}99, Dav98a, HW94, HMEW96, JJ98, Web99, AS92, BBB91b, DM92b, GD90]. Heterothermal [Niw98a]. Heterotrophic [KKH99]. Heterotrophy [EGDG^{+}98]. Heterotrophy-autotrophy [EGDG^{+}98]. heterozygosity [Cam95]. Heterozygous [ZJ94]. Heuristic [EKK98]. hexapod [Fra95c]. hexokinase [AS92]. hidden [CT95, Cze92]. Hierarchical [Bur94, Jen98, Ned97, Bur92]. hierarchies [TBD95]. Hierarchy [BTM93, Jen98, Bur95, HKG90, Pol92b]. High [BG96, Coc98, GMN96, LPBR96, PK93, SEB99, SS94c, SW96, SSA94, TP94, vBW98, BS91a, LVL92]. High-dimensional [SS94c]. Higher [Hol96, SGSS98]. Highlighted [RLF93]. Highly [NH96]. hippocampal [AMT91]. Histograms [HFS97]. Histone [Fli99]. histones [BSGB91]. Historicity [Toq97]. histories [Tay91]. History [Alm99, Kal98, ONRH96, Woo93, AVK94, Cou91, HP95, IL95, McN93, PNB92, SC93, WH98a, WT94]. Hit [MNH97]. hitchhiking [Ber95a]. HIV [BN94, DP98, EP94, Fin92b, Gla91, Ham93, HLS94, LN95, MN92, Mel93a, NM92a, Now92, RBH93, RBB96, TCZ95, TL98, WZN97, WDP98, Whi97, Wic99, WLJN99]. HIV-1 [Fin92b, Ham93, Mel93a, TCZ95, TL98, WDP98]. HIV/AIDS [LN95]. Hives [Wat97]. Hodgkin [LR94c, SH91b]. Holding [Kur99]. Homeostasis [DRB97, Eas91, Man95, OKAP98, PDP97, Vin98]. Homeostatic [KB98b]. Homing [BB93, CSSF93, Wal98, Wal99a, Wal99b]. Homogeneous [BS98a, DVV99, ZH97, RCCM92]. homologies [Liq91]. homologous [BSOB92, IMPR95]. homology [DKL95, Mil91b]. homooligomer [Nus91]. Homopolar [LH94]. Homopolymer [MP99]. Homosexual [HLS94]. Homosexuality [BK97]. Homotypic [AR99]. Honest [AC96, Cor95, Joh95, NS99]. Honey [BSTV93, Vis93, Wat97, CS91, LL90]. Honeybee [Mye93, WC95b, CSJM90]. Hopf [SH91b]. Horizontal [MS98, Now91]. hormonal [Jam90a, Jam92]. Hormone [BW98, HS97, HU97, Jam96, Jam99a, Jam99b, Ros94b, BZGM91, LA90a, SP90]. Hormone-Binding-Protein [BW98]. hormone-like [SP90]. Hormones [WZN93]. horse [MA95]. Host [AP95, BB97a, BB94, BB99b, Bow99, CDG^{+}98, CCVH98, CH96, Dam99, Gat95, GG97b, HS96, KYH99, KD99b, NBBH99, RRG98, SRH97, WJ97, WJH97, WG97b, Yam99, AWA98, BB91, HS95, Kun92, RdM91]. Host-Host-Pathogen [BB94, BB91]. Host-Parasite [HS96, Yam99, GG97b, RdM91]. Host-Parasitoid [CH96, KYH99, SRH97, RRG98, WJH97]. Host-Pathogen [Dam99]. house [Sav92]. Housefly [WH94a]. hox [JO96]. Hsu [CB91b]. HTLV [Liq91]. Hubbell [ZL97]. Huddles [SA97]. Human [AOL98, And98, AKV96, AKMV96, rvdBvdG96, Bje96, CZ93, CC99, For98, HE94, JK98, JR92, Kel96, KSM97, Khr98, KS99, MJLW96, NS96, Pen91, RBB96, SLR^{+}98, Tan99, Vrb98, Wil99, ZZ94, Aok90, Aok91, Gla91, HA92, Iid90, Jam90a, JP90a, JP90b, KB95, Koz91, MK90, Nor92a, SS90a, vdGvdBB92]. Humans [SSS93]. Hummingbirds [LD94]. Humoral [RDHD98, SV90]. Hundred [Por96]. Huntington [BKA97]. Huxley [LR94c, MC96, SH91b, TGS93]. Huxley-Simmons [MC96]. Huxley-type [TGS93]. Hybrid [JW95, Lei94]. hybridoma [SP92b, SP92d]. Hybrids [Rod96]. Hydraulic [FK99]. hydrocarbons [CH90b]. Hydrodynamic [Com94b, CMB95, VSW97]. Hydrodynamics [Com94b]. Hydrogen [Sta98, BW92, BT91, Mró95]. Hydrolysis [FPB99, RP99]. hydrophilic [RHS^{+}95]. Hydrophobic [DJZK96, PGK94, RHS^{+}95]. Hydrostatic [SWKS96]. Hydroxyaluminosilicates [EPT97]. hydroxylaminolysis [RP99]. Hymenoptera [RB97a, Rat90]. hymenopteran [RR92]. Hypercube [GG97a]. Hypercycle [NMdlR93]. Hypercycles [CB94b, BH95]. hyperlogistic [HA92]. Hypermutation [KP93b]. Hyperoscillations [Mur92]. Hyperoxidant [HA90]. Hyperparasitism [HH98]. Hyperthermia [SHW94]. Hyphal [Koc94b]. hypothalamo [BZGM91]. hypothalamo-pituitary-gonadal [BZGM91]. Hypotheses [AVK94, Jam98a, LM94, Her95, Tre95]. Hypothesis [Bea90, Bro97, Di 98, DSD93, GRRF93, Jam98b, LGSM97, LÅ94, LÅ96a, Lin94, Nor93, Plu94, PA98, Pol96, Wag93, WW94, BBB91a, FL95, Fre90, FW90, Gra95b, Jam90b, Kan92, Kra95a, MMH90, MBS92, Shu91, VpM95]. Hypothesis-testing [Bea90]. hypothesized [Jam90a, Jam92]. Hypothetical [OK94]. Hypoxia [KWD^{+}94]. Hysteresis [Kle98, PA98, SQ98]. Hysteresis-driven [Kle98].

I-Prototypical [CCFS96]. IDDM [RBR90]. idea [Cle91b]. Ideal [Cha93, FB97, RM97, SF90]. Ideas [Duk99, Dus96]. Identical [ZH98]. Identification [AMO93, ALM95, HH91b, KL91, QF94, Say94, WW97, FAB92, MRB91]. Identified [Kat96]. Identifying [GCB96, Sve94]. Identity [BSN94a, TL96]. Ideotype [CCI94]. idiotype [VS90]. Idiotypic [BvHS93, CS94a, DBSV94, DSP96, WZN93, WDP90]. If [Cár91, Kel91, Kau98]. Ignorance [Web99]. II [CCDD92, CH94, AW94, CC97b, CCS96, CL99b, Cho93, DRB97, DC96b, ER94, FCV99b, GS93, HH98, Kat94b, KG90, Mal94b, MRB91, MAC99b, MP98c, NSL91, PVT91, RBR90, SP92b, SP92d, Sch91b, Shc94, Smi90a, SV90, SS93c, Trz99b, VP93b, WN99b, WMLK96b, sDAB91]. II-Comparison [CL99b]. II-Hyperparasitism [HH98]. II-Nonbalanced [CC97b]. II-Post-copulatory [FCV99b]. II-Ternary [Trz99b]. III [BR94, DC96c, GRRB90, JP90a, Mar97b, Smi90b, SW95a, WMLK96c, ZJ97]. illustrated [BW92, CJ90]. Illustration [ON98]. Images [OBGGGRR93, VpM95]. Imaging [LPP^{+}94]. Imbalance [Jam99b]. Immature [BBM^{+}97, BBM^{+}13a, BBM^{+}13b, BBM^{+}13c]. Immediate [NM93b]. Immigration [RM95, McC92]. immortal [RGK91]. Immune [BvHS93, BdS97, BR94, CBSV95, CCFS96, CCS96, Cho93, DP98, FV93, FV94, FCV97, HW98a, LCP^{+}98, MTLK96, NMS95, PGBYP97, RWN98, RDHD98, TH98b, WR93, ALW91, BK90a, Car90, CSC90, CSS91, Cou91, DSP92, For92, MPRB91, MRB91, SC92b, SV90, SV91, VS90, VK91, Wei90b, dBP91]. Immunglobulin [FBHK98]. Immunity [AK94, MS98, YCM97, YC99, CSS91, MS92b, TF99]. Immunodeficiency [Kel96, CPLA95, MK90]. Immunogenicity [HW98a]. Immunoglobulin [KDBK93]. Immunological [Dus96, SFHP97, SS94c, Whi97]. Impact [BDDF97, DDS96, HH91b]. Impaired [WW94]. imperfect [Fia92]. Implications [Ber91, CC96, DRB97, Dwy98, FV94, Gat95, GMN96, Kra99, MP99, MS90b, RBH93, Sau94, Sav91b, TET91, THS^{+}95, VVC95, Gro92, Sav95a]. implies [RV90]. Importance [AP96b, HE98, KK93, Mac90b, WWC90]. important [Cár91]. imposed [Hou92]. Imprinting [IP99a, Var93b]. Improve [AKMV96, Fis99]. Improved [CMPR93, PG99, BZGM91]. Improvement [McN93, Vid91, sDA91]. Improving [Doi97]. Impulse [EMRS90, Mar95a]. Impurities [Seg94]. in-feed [PBJ^{+}92]. IN1 [Ano92-29]. inaccessible [Now90]. inactivating [LCR94, Ste90b]. Inactivation [Rou98, BS92]. Inacurate [BDK^{+}97]. Inbreeding [DG93, Gan99, LR94b, BW91]. incest [SW95b]. incidence [Ste91b]. incidence/age [Ste91b]. incisiform [GKMSS90]. include [RBR90]. Including [AVM^{+}99, MKC^{+}98]. incomes [MH90]. Incompatability [Fra98]. Incompatibility [Fra97a, Hur97b, NH96, WS94, Hur91]. incomplete [Arm90]. Inconsistency [Wil98]. Incorporating [Dus96, Kab94, MS98, Par94]. Incorporation [HE93]. Increase [Cam93, PN98]. increased [Lun90]. Increases [Sch99a]. Increasing [CSFC91]. Incubation [HDF^{+}99]. Indefinite [HS96]. Independence [SSCvdM99, GR97c]. Independent [DSP96, Fri98, Hol96, MM95, SS94b, Ute93, BC90, Tak95a]. Index [Ano90l, Ano90m, Ano90n, Ano90o, Ano90p, Ano90q, Ano91m, Ano91n, Ano91o, Ano91p, Ano91q, Ano91r, Ano92j, Ano92k, Ano92l, Ano92m, Ano92n, Ano92o, BJS97, KS98]. Indicator [FVB97, WM98]. indicators [BDK91, Her90]. Indices [Ber95b]. Indirect [KB98b, NS98, Van90, Van91]. Indistinguishable [NMdlR93]. Individual [BSM97, CY98, FPB99, KK98, Kur99, OES93, SA97, Sch97a, Vis93, Wes93, KDG92, Pec92, RG95].

Individual-based [FPB99]. Individual-Oriented [BSM97]. Individuality [Jab94]. Individuals [FGLO99, GLR96, Hur97a, MA97b, Rux96, SA93]. Induced [AP95, LCP^{+}98, MR98, MKC^{+}98, PN98, WOQP93, BB99a, CG96, DYL92, FV94, FLR93, HKT91, KD94, Kum91, LOLM91, LC97, MR92, MR90, Mor90, RB91, SKI98, UJ95]. Induction [BB99a, RBH93, SGSS98, Ko91, Kuz95a, Nun90]. Inequalities [Ano94b]. Inert [TWB^{+}93, WGH98]. Inertness [KW94]. Inevitability [Koe97]. Inexcitable [PK93]. Infant [WW94]. infanticide [Nor92a]. Infarct [Vol97]. Infected [GH97, NBSM97]. Infection [BR94, Boc98, BT97, BGE94, DP98, EP94, HLS94, MB97, Wic99, WLJN99, KB95, MK90, PBJ^{+}92, PNB92]. Infections [Fra98, MMB99, RBBG98, NM92a, Now92]. Infectious [AWA98, Hoc91, Lun90]. infective [BB91]. Inference [Bro97, RN95, GS95b]. Inferring [ONRH96]. Infertile [Rod96]. Infertility [BP96]. Inflammation [TM92b]. Inflammation-mediation [TM92b]. Influence [BMG98, BTD98, Com94b, CB97b, Get93, Gou94, KKH97, LVL92, NMdlR93, Par97, Ric96, SEB^{+}93, SG91, WRP99, YY94, MP91, OHS95, WS90]. Influences [MDGB94]. Influenza [BR94]. influenzae [LPRB96, RDHD98]. Information [CV93, Dru97, DGSK99, For99, FOO94, HWO90, Hir93, MSN96, MTM^{+}96, Sch97a, Sch99b, SH93b, Sto98, Tor90, WW98, Wil95, CV91, Ish90, KC91, Koh92, LB95, Man90, Mar95b, SB95, THS^{+}95]. infusion [SF90]. inherent [Gos92]. Inheritance [GR98, Jab94, LJ96, PM99, Tak98, JLL92, Smi90c]. inhibited [SH90b]. Inhibition [BS98a, CMML96, KPS97, McC93, MBSB95, UP96, vOvO97, Abe90, AMT91, THY91, WT90]. Inhibitor [Nöl99, Bro90]. inhibitors [Wan90c]. Inhibitory [Tor96]. Inhomogeneity [AP96a]. Inhomogeneous [Tyv94, ZWH98]. Initial [Bar94, Rak97, RP99, Seg94, Spi93, TéS96, Coh91, MS90a]. Initiated [ZWH98]. Initiation [SAF^{+}93, Boe90, HB91]. initio [NT97, NT99, Nov97, STI93]. Injection [TWB^{+}93, FSD91]. Inner [RS98, Bra90a]. Innervation [RBC98, vOW99]. Inorganic [Mar97a]. Inositol [LC97]. Input [BDDF97, Lac90, Smi92]. Insect [BBIK97, Chi96, Fra97d, Gou94, NG99, NM93b, CK95a, Cor95, GB95, Hor95, SSSF95]. Insects [AP99, BBM^{+}97, NM93a, Nun98, SH93b, TB95, WNY98, Yam93, BBM^{+}13a, BBM^{+}13b, BBM^{+}13c, Nun90, NLS91, NSL91]. Insertion [KSF98, PSLB99]. Insertions [SK96a, AM92]. Insight [Alm99, ROL99, Rig94]. insights [Kuz95a]. Insoluble [FPB99]. InsP [LR94c]. Inspection [Fis99, Dug90]. inspiratory [Eug95]. Instability [BG98, GH97, MC93, Rig94, SM99, Lib91]. Instantaneous [ODA97]. Instrument [AC96]. insufficiency [Tri92]. insulin [But92, GRRB90, HCS91, HH91b, QC91a, QC91b, RBR90]. insulin-dependent [RBR90]. Insurance [For90, FL96]. Intact [Mar93b]. Intact-body [Mar93b]. Intake [Emm97, PGF94]. Integral [SKW98]. Integrated [Com94a, Isr96, VF98]. Integrating [DS99]. Integration [Bjö94, NM93b, PGF94, SK96a, Sea99, BS95, MS95a]. integrator [Eug95]. integumental [Hoa90]. intelligence [GV91]. Intensity [MDGB94]. Inter [BP97a, Nöl98, Nöl99, Ste97]. Inter- [BP97a]. Inter-Cellular [Ste97]. Inter-residue [Nöl98, Nöl99]. interact [BSB95]. Interacting [BLGG^{+}93, PGK94, TMZ99, SL91a, Sta90, ZW95]. Interaction [ASA98a, Bec96, Che94, Gat95, HKT91, KYH99, KKN99, Mat96, MB93, MK93, NT97, SOR91, Spi93, Tor96, AdG90, Don90, Hem90, KJ95, MGL^{+}95, McA91, TS91, vdGvdBB92]. Interactionistic [LÅ94]. Interactions [AL96, BvHS93, Big93, BBM^{+}97, Cam93, CS94a, CH96, Doe97, Dwy98, FP93, FAF90, GG97b, HE98, Kem94, KO93, MSSG94, NG99, Neu98, OS97, Sto98, SK96b, VC98, Yam99, BH95, BBM^{+}13a, BBM^{+}13b, BBM^{+}13c, BS92, Com90, Dex95, DML95, MN92, RB91, RKM92, Van90, Van91, VS90]. interactive [Jen98]. Interbeat [LGSM97]. Intercalation [BT90b, Mac90b]. Intercellular [BHP92, CMML96, Mel93a, RBC98, WS98]. Interchange [Hem91, LTZ96]. Interconnected [ABR97, ABR99]. Interconnected-Tubes [ABR99, ABR97]. Interconversion [MS90b, Mar90]. Interconvertible [RP99]. Interdependence [EG98]. Interests [Vis93]. Interface [AVM^{+}99]. Interfacial [SM99]. interference [BSGB91, Grü95, PW91]. interferon [BLB95, BB99a, Mar96]. Interferon-induced [BB99a]. Intergrate [BS98c]. Intergrate-and-fire [BS98c]. intermediary [SP92a, SP92b]. Intermediate [TP94, Ujh96, HKT91]. Intermediates [CLY99, FK97, RP98]. Intermittency [YFG^{+}93]. Intermittent [CK95a, Doe94]. Intermolecular [ST93, Mar92a]. Intermonomeric [EKK98]. Internal [FLR93, LRB94, LT98]. Internalization [HL98a]. Interneuronal [Leb96]. Internodal [Kor93]. Interplay [BT97, FGLO99]. Interpret [FDT^{+}93, Mor96]. Interpretation [KS96b, KDBK93, Mac90a, OK94, PL93, SP92b, Wei97, SW95a]. interpretations [Sal91]. interpreting [FSD91]. Intersexuality [MH96]. Interspecies [TWW90]. Interspecific [BT97, Gou94, CPJL92]. Interspike [ML94b]. interstitial [Fie91]. Intertidal [CSSF93]. Interval [Jol90]. Intervals [BB97b, HS98a, LGSM97, ML94b]. Intestinal [LBWP93, LBP^{+}97, LW94, LG91, LSM^{+}92]. Intra [BP97a]. Intra-species [BP97a]. Intracellular [AMW95, AP96b, Bra90b, DP95b, For94b, JSR99, KB93a, MGR97, MFB98, RBC98, WSF99, BT90a, CLF95, For95, MH95, Mir90]. Intraclonal [Tan99]. Intracolony [Vis93]. Intragenomic [Hai92]. Intramitotic [Tan99]. intramolecular [Sol95]. Intramural [BP99a]. intraocular [Rei95]. intraruminal [FSDT91]. intravenous [SC90]. intravesicular [NK90]. Intrinsic [For99, Gaá93, Gri94b, iT94]. Intrinsically [GKM93]. Introduced [KM96a, Moc99]. Introduction [BBL^{+}97a, BLLÅ94]. Intron [FLR93, NC92, ZLYZ98, Dib91, XCS^{+}95]. Intron-containing [ZLYZ98]. Intronless [ZLYZ98]. Introns [CL98, Eld97]. Intruder [ES95]. Invariant [SK99b, WH98a, Doe95, Phi99]. invariants [DS90, Fel91b, FS95a, Fu95]. Invasion [CGM^{+}98, Get93, SS93b]. invasive [Str95]. inverse [SW95a]. invertebrates [Rat91]. investigate [SG95]. Investigating [CK99, EL97, RS97, SK92]. Investigation [ASA98b, BS92, PMA97, WB99, BN95, PST91, PVT91]. investigations [Hem90]. investment [GD90]. Inviability [JW95, Lei94]. Involved [Mah93, PSLB99, For95]. Involvement [Dum94, Yar96]. involves [ALW91, Sai95]. Involving [MR98, SH94, LH91]. Ion [BLM94, CAPP99, Che94, DCL96a, Got94, Gre98, LPB96, Leb96, Nov97, Sta98, Sta99, Ute93, UP96, DF91a, LT91, MLS92, SGG95, SM90a, SH91b]. Ion-Channel [Che94]. Ionic [CMB95, DLC98, Kab94, LJ95, STI93, VR94b, DL90, DF91b, LG90a]. Ionizable [tBázsáSH96]. Ionizing [CS94b]. ions [FG92, LOLM91]. IP [KD94]. Iron [OKAP98]. Irradiance [Zon97]. Irradiated [MFHA^{+}95]. Irradiation [AKL97]. Irregular [AVM^{+}99]. Irregularity [OS97]. Irrelevance [WHL93]. irreversibility [Ber95c]. Irreversible [HE93, Coh91, Top90, Wan90e, ZT92]. IS10 [Jai97]. ischemic [BHP92]. Isentropic [WZ96]. islet [AGG91]. Isoelectric [WW97]. isolated [FBH91, HKT91, Lim90, SH91b]. Isolation [GG97a, Mar97b, HA90, Mar91]. Isologous [KY99]. Isomerase [OVH^{+}96]. Isomeric [RP99]. isometric [Mor91]. Isotherms [BLGG^{+}93]. Isotope [FBS93, FHR^{+}99, FSD91, FBL^{+}95]. Isotopic [Rog94, Ber92]. isotopy [Han95]. isozymes [Ure91]. Issue [Ros94b]. Italy [Ano99a]. Iterated [HS98c, MI98, NS97, SNT95]. Iterations [DC96a]. Iterative [VR94a]. Iteroparity [Rod96]. iteroparous [Tak95a]. Itself [Gán97]. IV [JP90b]. IV. [Zha98].

J [BBM^{+}13a, BBM^{+}13b]. James [MB92]. January [Ano90r, Ano90x, Ano91v, Ano91-29, Ano92t, Ano92z, Ano93c, Ano93j, Ano94d, Ano94k, Ano95c, Ano95i, Ano96c, Ano96i, Ano97d, Ano97j, Ano98e, Ano98k, Ano99j, Ano99p]. jaws [Gal92]. Jet [MMH90, MBS92]. Joint [CZ93, Fra97d, Mul93a, MdR98, ZJ98]. Jointly [LR94b]. Juglans [EGDG^{+}98]. Juke [Fu95]. July [Ano90s, Ano90y, Ano91u, Ano91-30, Ano92q, Ano92w, Ano93g, Ano93m, Ano94h, Ano94i, Ano94f, Ano95f, Ano95l, Ano96g, Ano96y, Ano97i, Ano97o, Ano98j, Ano98p, Ano99h, Ano99n]. jump [NK90]. Junction [BSW97, KM97, Kee91, SWA91]. Junctions [BFG96, CM95]. June [Ano90-31, Ano90-36, Ano91-31, Ano91-37, Ano92-28, Ano92-36, Ano93n, Ano93t, Ano94r, Ano94w, Ano95q, Ano95w, Ano96k, Ano96q, Ano97s, Ano97y, Ano98s, Ano98y, Ano99x, Ano99-29, Ano99a].

K-ATPase [Ste90a]. K562 [Boe90]. Kacser [CB96b, Fel96]. Karyotypes [Ost96]. Kauffman [BP97b, SBRG99, SJ94]. keep [GE90]. Keller [Sai95]. Ketogenesis [KQ96]. key [YC91]. Killed [Eas99]. Killer [Kuz96]. Killing [TAN94b]. Kin [DT98, Fra97c, GM99, Gan99, Kel92, Lei94, Lew98, TF96a]. Kinase [EEG93, Kem94, MS94, CKR95, DC90, KH98]. kinds [Smi91]. Kinematics [FB95]. Kinetic [ARC97, AOU98, EWHB96, FH99, HCM^{+}92, HH91a, HBJ^{+}97, KH94, LPP^{+}94, LR94c, Mal94a, MKZ97, Pea96, Pet96b, PA98, Rak97, RP99, SMA95, Sta98, Sta99, YYYU97, ZW96, BW92, DL90, Eas91, SM90b, TCLB90, WKSH92]. Kinetically [NMdlR93, Cze92]. Kinetics [AP96b, AKMS97, BIV97, Ber93, DBS97, GMST99, GMN96, HV90, KV93, Mar93b, Mor99, NK90, Pet99, Rau94, RP96, SM97a, SLBT99, VHV^{+}90, VH90, VHGV92, Wan90a, YYYU97, ZT92, tBázsáSH96, Bar90, BSB95, GRRB90, JR92, KJ95, LT91, LPS90, MK92, RBR90, Sal91, Sav95a, SG95, Smo95, Sta90, SK92, Top90, Wan90e]. kinety [LH90]. Kinship [Hol93, Thi90]. Kleptoparasitism [RM97]. Klinokinesis [Wil94]. Knee [Mul93a, MdR98]. Knew [WK96]. knocks [CPJL92]. knots [LJ92]. know [Kum91]. Knowledge [KKS90]. Knowledge-based [KKS90]. Known [Nuy94]. Koinophilia [Koe90, KK94a, Koe94, KK94b]. Kurtosis [Wys92].

L [Hav99, RWB98, SP90, MWAS91]. L-chain [Hav99]. L-dopa [SP90]. L-system [MWAS91]. L-tyrosine [SP90]. L. [EGDG^{+}98, Mal94a, WWC90]. Labeled [Seg94]. Labor [NG99]. Labour [Gra98]. labyrinth [BS95]. lac [Rad91, Yar96]. lac-Permease [Yar96]. Lactating [BDK^{+}97, FHR^{+}99, MFM97]. Lactation [KD93]. Lag [Bar98]. lagging [FD92]. lambda [RV90]. lamination [McI95]. Land [FVB97, Ste91a]. land-sea [Ste91a]. Landscape [AH96, AH98a, AH98b, BOG^{+}99, WH96]. Landscapes [BSS96, BS93, Che93, Jon96, SG93, SJ94, WS93, KJ91, WBS95]. Language [BKNR97, NPK99, TET97, Ujh96, Aga95]. lapse [VpM95]. Large [BB97b, BC94, Cra90a, GL95, GL93, MA97b, SB94, TF96b, BSB90, But90, DS92, FC95, McK95, Mur91, Van90]. Large- [Cra90a]. large-amplitude [Mur91]. Large-scale [MA97b]. Larger [Woo99]. Larvae [CSB98]. Lasius [BDG92]. Last [Koc94a]. Late [MH97, McK95]. Latency [PK98]. Later [Mar96]. Lateral [CMML96, McI95, TH98a, THY91]. Lattice [Big93, Har99a, NMI97, NNI98, NZPB90, RA96, Sat96, TA99, TB95]. Lattice-Structured [Har99a, NMI97]. lattices [THS^{+}95]. Law [KB98a, PMdM^{+}99, TA99, MH90, Spa92]. Laws [KKH99, Tab98, WW98]. Layer [Ves98, LSM^{+}92]. Layers [MB93]. LCMV [Boc98]. Lead [DP94b, GMN96]. Leading [SS97, FD92]. Leaf [AVM^{+}99, Jon97, Kor97, LAV^{+}97, Sak94, Ves98, Sak92]. leak [Bra90a]. leakage [HKT91]. Leaky [BS98c]. learn [MA95]. Learning [And95, Fra97b, GHW99, KRG95, TS95, Orp90]. Leaves [Baj99, LBND98, Sek95, WBC99, SW92]. Leech [SWKS96]. Left [Cha98, CBS^{+}99]. Leg [Fra97d, Fra95c]. legs [Kuz95b]. legume [Str95]. Lekking [Kok97, DH95]. Leks [Hut99, MDGB94]. Lemna [Mal94a]. Lemuridae [Fed93]. Length [Doi97, FR93, MS97b, MM97, Oex98, Pos99, Reg93, MBSM92, WOQP93]. Lengths [Mul93b, JV91]. Lens [Bet99, HC93, MH97, Rei95]. Leonardo [Lon94]. Leptothorax [THF92]. less [KSH95]. Lessons [Vol97]. Let [Eli94, Yoc95]. Lethal [JD99]. lethality [For95]. Letter [Ano98d, Laz97].

Leucine [FHR^{+}99, GRUH99, FBL^{+}95]. Leukemia [LRZ^{+}96, Phi99]. Leukocyte [DS92]. Level [Hag93, LL96, BDK91, GR97a, Hem90, Her90, LBDDG90, NM93a, NRD94]. Levels [Cha91, DS99, For99, Jab94, Jam96, Jam99a, MO96, Pos99, DF91a, Pol92b, RV90, Sza92b, YC91]. Libet [LÅ96a]. Libraries [KOTL97, MCG95]. Life [AP96b, BBL^{+}97a, Blo97, Edw96, Eli94, KC91, LBP^{+}97, McN93, MA97b, PSLB99, PAG99, RBD97, Rus97, SC93, SS97, Var98, Wäc97, WH98a, WT94, Woo93, Yoc95, Cha91, FTD90, IL95, SAS92, Sza92b, Tay91, Eli96]. Life-history [AVK94, McN93, SC93, WH98a, WT94]. lifespan [RGK91]. Ligament [Mul93a, MdR98]. Ligaments [Mul93b]. Ligand [AOL98, FV94, Hon98, HMEW96, Mat96, ROL99, GWS90, Wil95, WAJ91]. Ligand-induced [FV94]. ligand-mediated [GWS90]. Ligand-Receptor [Mat96]. Ligand/Receptor/G [AOL98]. Ligand/Receptor/G-Protein [AOL98]. Ligands [BLGG^{+}93, CAPP99]. Light [AKL97, CCI94, DPG^{+}94, HVKS99, JK98, JW99, KKS97, LWS97, LAV^{+}97, RB97b, WH94b, ZvdBK97]. Light-driven [LAV^{+}97]. Light-limited [HVKS99, ZvdBK97]. Light-pulse [LWS97]. Like [LR94c, BE92, CLF95, Enr90, Fle97, Joh93, LBW^{+}95, RBH93, SP90]. Likelihood [TL96, SWA91]. likely [Bat91]. Limb [AP99, DO99]. Limit [DA96, JK98, PMA97, But90, Wor95]. Limitation [RWN98, ZJ97, BSB90, HP95, KE92]. Limitations [Mul99]. Limited [Com94a, DP98, Kau98, HVKS99, KJ95, RGK91, Van92a, Van92b, ZvdBK97]. Limiting [CS94a, KKN98, MKM96, Mal94a, Ste91b]. Limits [AM98, Sey91]. Lindahl [Bec96, Lib97]. Line [BP97b]. Lineage [Moc99, Gra95b]. Lineages [Che93, Gra95b, Sta92]. Linear [CMB95, Edm96, Fu95, JTC99, SM99, Ste94, WN99b, BT90b, BT90a, ER94, EV94, Hoc91, SP92a, SP92b, Sta90, WN99a]. linguistic [Ser91]. Link [JO96, WOQP93]. Link-length [WOQP93]. Linkage [SS93a]. Linked [Rux96]. links [LJ92]. Linnaeus [Gri94a]. lipases [HCM^{+}92]. Lipid [FH99, GET96, Nor93, HKT91, RB91, Sme90]. lipids [LOLM91]. Lipoproteins [TP94]. liposomes [SH91a]. Litter [KD93]. littermates [Jam91]. little [GN95]. lived [CLY99, Rou98]. Liver [CG96, Cze93, FHR^{+}99, Wei97, HCS91, Rog95]. Living [MA97b, BB91, HM95, KD97, MH95]. Lizards [GHW99]. Load [HS96, Atk91, Hed92]. loaded [BSB95]. Loads [HA97, KL96, WDP98]. Local [AS92, AP96a, Bai93, CH96, Hie97, Jos97, KV93, LKS^{+}93, ML97a, Neu98, OR97, PLB96, ZJ95, BBM^{+}97, BBM^{+}13a, BBM^{+}13b, BBM^{+}13c, PW91]. Localization [Fer94, Miz95]. Localized [PTBL98, WDP90]. Locations [Ost96]. locked [CLF95]. Locking [YFG^{+}93, ZCS^{+}90]. Locomotion [LQL94, MNB97, McG93]. Locomotor [LWS97]. Locus [SBRG99, Yi98, Gre92, Hin98, ZJ94]. Locus-specific [SBRG99]. Locust [NM92b]. Lodging [BBS^{+}98]. Loft [HP97b]. Logging [KKH97]. Logic [BNS97, Sib95]. Logical [SvHT97, Tho91]. logistic [Llo95]. Lognormal [Koz96, Gin95]. Long [EFB98, Har99a, Hol99, Man92b, MNH97, NM93b, Pen98, Rou98, SS99, VGV93, JH95, RKM92, VC95, Vol90]. Long-distance [Pen98]. Long-lived [Rou98]. Long-range [Har99a, Man92b, Vol90]. Long-term [EFB98, Hol99]. Longevity [Pro94, McK95]. look [Jac91]. Looks [Hop94]. Loop [DvH93, Thi90]. looped [GM90]. looping [Nus92]. loops [For98]. loose [Bat91]. lordosis [SG95]. Lose [Pos99]. Lose-Shift [Pos99]. Loss [KGFB97, MvB93, SA93, WHL93, AKW95]. lot [GN95]. Lotka [DM91, Sat96, Van91]. Lottery [ZJ97]. love [Leo92]. love-dart [Leo92]. Low [BH94, Koz96, WDP98, ZGZ94, Gin95]. Lower [BvLVR96, BVvL96, LIWE90]. Lowering [GPL98]. luciferase [AS92]. Lucilia [FWF97]. Luminescence [Baj99]. lung [Bla92]. Luria [Jon94]. Lyme [CGM^{+}98, SAS92]. Lymphocyte [AL96, FP95, MFHA^{+}95]. Lymphocyte/Antigen [AL96]. Lymphocyte/Antigen-presenting [AL96]. Lymphocytes [Ste97, CW92]. Lysis [RZF^{+}99, Abe90]. lysozyme [Nak95]. lysozyme- [Nak95].

M [NT93, RSB92]. M-phase [NT93]. Macaque [DvH93, Thi90]. machine [KC91, Sud90]. Machines [Kre94, Kre92, Sch91a, Sch91b]. Macroevolution [Sýk92]. macroion [Com90]. Macromolecules [Com94b, Meg90]. Macroparasite [WG97b]. Macrophage [OS97, WLJN99]. Macrophage-Tumour [OS97]. Macrorestriction [MAGO97]. Macroscopic [MPAA97, Aok92]. Made [CPK96]. Mafia [SMS98]. Magnetic [LPP^{+}94, Phi96, Wal98, Wal99a, Wal99b]. magneto [Jac91]. Magnitude [WOQP93]. Main [LH94]. Maintaining [PAG99]. Maintains [TH98a]. Maintenance [Gag94, Get93, Mau92, WDP98, WLM96]. Major [Bal97, Sav92, SS97, Jol90, JD92, Mat95]. Make [BMS93, áSá96, TF96a, DH92]. Making [SK96a, SM97e, MWC^{+}97, SFKD99]. Malaria [DG93, How92]. malariae [MMB99]. Malaysian [KKH97]. Male [GG98, JM95, JM98, Kok97, Mar97b, Wie90]. Male-choice [Mar97b]. male-male [Wie90]. Males [Bed96, BMS93, SH94, TM92a, Whi94, Whi90]. Malignant [Rig94]. Maltose [MBSB95]. Maltose-binding [MBSB95]. Mammal [CCK98, MH96]. Mammalian [Bai93, Dru93, Jam96, Jam98a, Wil99, Wol91, Fie91, Jam91, Jam92, KA91, RGK91, SB90a, Woj92]. Mammals [Dru93, Jam97, KWB93, KGFB97, Var93b, WHL93, WN99c, BBP95, Beu90, Hok90, KS91, KE92, McK95, SK92, YI95]. Mammary [FHB^{+}97, MFM97, FBL^{+}95]. man [BZGM91, SC92a, Tre95]. maneuvers [Mag95]. Manipulation [AC96, Kra95a]. Many [Bow99, KS96a, MO96, RR96]. Map [Big93, FK97, Wal99b, Llo95]. Maple [YLT98]. Mapped [BD94]. Mapping [FS98, Hon98, GKMSS90]. maps [LBDDG90]. March [Ano90w, Ano90-27, Ano91w, Ano91-28, Ano92u, Ano92-27, Ano93e, Ano93l, Ano94g, Ano94z, Ano95d, Ano95j, Ano96e, Ano96w, Ano97g, Ano97m, Ano98f, Ano98l, Ano99i, Ano99o]. marching [NM92b]. Marine [KGFB97, WHL93, Hok90, Rat91]. markedly [Smi91]. Marker [JC96, ABH94, SF90, SH90a]. Markers [Moc99, Vol97, FSDT91]. market [PSE92]. Markov [CS94b, CMG^{+}96, Tyv94, VK99]. Markovian [Bar92, Gil92, Kok96]. Marrow [LL93, SH90a]. Marsh [YLT98]. Mass [BTD98, Dru93, HSM94, KKH99, Lor94b, New97, PTBL98, WW97, Wit95, Fel95]. Mastication [rvdBvdG96, vdGvdBB92]. Masting [ISSI97]. matching [HM95, PSE92, TPAS95]. Mate [Hut99, NAZG95, OJ99, Smi94, WM98, ZJ95, DH95]. Mate-seeking [Hut99]. Material [CBFR98]. Materially [Tan94a]. Maternal [Jam99b, Smi98]. Maternal-effect [Smi98]. Mathematical [AOL98, AL96, AKL97, Ano99a, AKW95, AKV96, AKMS97, BB99a, BKK96, BR94, BSB95, BS91b, CI97, CFC98b, CMML96, DSM99, DRLB91, DO99, DCL96a, DCL96b, Flo98, FBHK98, GMST99, HS95, HL96b, KK96, LBND98, Lor94b, MSHN97, MCG95, MPRB91, MRB91, MvB93, MAGO97, MBSB95, MAC99a, MAC99b, PG99, RBM97, RDHD98, SC96, SSS93, SGP98, SDB90, SSA94, TS94, WSF99, WDP98, WGH98, YH90, tBázsáSH96, BSOB92, BM95, BZGM91, CSJM90, FSDT91, GRRB90, HKT91, JH95, KA91, MPO95, NM92b, Pol92a, PN90, QC91a, QC91b, RBR90, SP92a, SP92c, Zhe91]. Mathematics [Ano99a]. Mating [AR99, Daw96, Get93, Koe94, LR94b, NM93b, RWB98, SH94, Tyv93, Cam95, Fia92, Kir92, Rat90]. Matrices [EG98, SK99a]. Matricide [Bou94]. Matrix [CPK96, DSM99, HE94, TN96, TJ93, Com90, Hem90, SAS92]. Matter [BLLÅ94, Blo94, DC99, Tar94, vW94b]. Maturation [CC96, HC93, PGBYP97, WW94, CW92, Wei90a]. maturational [FTCM90]. Mature [TP94]. Maturity [Emm97, TC97]. Maximal [Cha98, HK96, HH91a]. Maximization [Fra97c, Sto94]. maximizing [MHW93]. Maximum [MS97b, SWA91, TL96, VV96, WJ97, CK95b, CCRM90, LB95, RCCM92]. Maximum-Likelihood [TL96]. Maxwell [ON98]. May [Ano90t, Ano90z, Ano91s, Ano91y, Ano92p, Ano92v, Ano93b, Ano93h, Ano94e, Ano94l, Ano95e, Ano95k, Ano96f, Ano96x, Ano97f, Ano97l, Ano98g, Ano98m, Ano99k, Ano99q, ÉT90, Now90, Hut99]. may-wigner [ÉT90]. MBNW [WGH98]. MBP [Liq91]. MCA [HB96]. Mean [Ber93, DY99, GV99, GL95, Hie97, SGP98, Tar94, LOB92, SC92a]. Mean-Variance [Tar94]. meander [HZ95]. Meaning [BP99b]. Means [Hon98, Mor96, RWK93, Seg94, YG94, Di 95, Sta92]. Measles [KB99]. Measure [DBH94, Mar93c, RP96, Vol97]. Measurement [KN93, Nun98, HH91b, Med95, SP92c, SP92d, Wil92a]. Measurements [Gie94, ODA97, SQ98, CSFC91]. Measures [LPBR96, LRZ^{+}96, Cam95, Doe95]. Measuring [KE96, LPP^{+}94, MIT96b, Sch99b, SB95, Wil98, Kre92]. Mechanical [Adl93, CMB95, Fer94, GM94, Gho93, KG93, MTS97, SM99, dLA97, MR91a, Ss90c, WK90]. Mechanics [CK97b, GM93, RLRV97, SFW^{+}96, SC99b, WS95, Dra90, VC95, Wan90b, WK90, Wan90c, Wan90d]. Mechanics-based [RLRV97]. Mechanism [AOU98, BM98, CCK98, DBB99, FA94, HL98a, HL96b, Jam97, JBSdlG97, KCL^{+}96, MKC^{+}98, Orp94, OK94, Rou98, RV99, SFKD99, Sea99, SL98, SAM97, Var93b, VGV93, Won94, BT91, Car90, ECB91, Gli91, GW90, Hoa90, Jac91, KCM90, Kuz95a, Man95, Miz95, Nak95, Pol92a, RD90, Sav95a, Und90, Vol90, WCW95]. Mechanism-Based [HL96b]. Mechanisms [BOS98, BC94, BB99b, Got94, KW98b, KO93, LQL94, Nak98, OP96, Orp90, PK97, RGK91, SS98, SS93c, TP96, Trz99a, Trz99b, Zhu96, vWP98, AMW95, BR91, JLL92, Rad91, VHG^{+}90]. Mechanistic [CK97a, CK99, CLY99, MFM97]. Mechano [RWDM98, Mur90]. Mechano-sensitive [RWDM98]. mechano-sensitivity [Mur90]. Mechanochemical [OSM95]. Media [RAGO99, EMRS90]. Mediated [AP96b, NG99, NBBH99, WS94, BB97d, CM95, CPJL92, GWS90, HL96a, LR94c, MK97b, RB91, SFKD99]. Mediating [ACC96]. Mediation [DBB99, TM92b]. Medicine [Ano97-28]. medium [NK90]. Meiosis [Hai93, Mir98, RH96, HN91, Mag92]. meiotic [DKL95, HG91]. Melanocytes [SPS93]. melanogaster [LWS97, SvHT97, VpM95]. melanoma [Ste91b]. Melding [AWA98]. Meles [SAM97]. mellifera [SDB90]. mellitus [RBR90]. Membrane [BG97, CMB95, FK98, FO96, Gho93, GM93, Gre98, HBJ^{+}97, Joh93, KN93, KSF98, KC97, Mar93a, MB93, Nov97, Rau94, VR94a, VR94b, WRP99, BSB95, Bra90a, CS90, Cla95, GMU95, KCM90, LRS90, MGL^{+}95, Nor92b, RKM92, Sax92, VR90]. membrane-bound [RKM92]. Membranes [HE99, KG93, KSF98, LJ95, PSLB99, RL92, AH90, CJ90, NZPB90, SL91a, Ss90c]. memories [WDP90]. Memory [BvHS93, Duk99, HC94, Hol99, Joh93, McL94, OR97, Pos99, ALW91, BSB90, FP95, Kra95b, Val92]. Men [BK97, HLS94]. Mendel [Por96]. Meningococcal [BGS^{+}99]. Mental [LÅ96a]. Menten [BIV97, CB91b, RP96, Sav95a, TH90]. Meristems [Bar94]. Mesophyll [AVM^{+}99]. Message [SB94, NAZG95]. Messenger [KL98, SMA95]. Meta [GLO99, SV91]. meta-dynamics [SV91]. Meta-stability [GLO99]. Metabolic [Ace93, Ace96a, AB98, Bra96, CLMH^{+}96, CPK96, Eas96, EG98, Fri91, GCB96, Gie97, GMN96, HR91, HH93, HCB96, JP90a, JP90b, Kei96, KW94, KQ96, MHTS90, MHWM94, MS97b, MA97a, OA98, PL93, RSW96, Sch96, Sen96, SSCC93, SAH96, TF94, TF96b, Tor96, VP93a, VP93b, WM91, BS91a, Bra90a, Bro91, FTCM90, FO92, HS91, Hof91, Irv90, KL91, Ová91a, Sav92, SP92c, SP92d, Spa92, Spe90, Sri91]. Metabolism [AW94, AVK94, AKV96, AKMV96, Beu90, BBL97b, HE99, MHWM94, NS96, WA94, WLR96, vRGTV98, FSD91, FTBC92, Ish90, MHM91, Sav91b, SP92a, SP92b, SP92d, WW92]. Metabolism/Life [AVK94]. Metabolism/Life-history [AVK94]. Metabolists [Kal97]. Metabolite [Edw96, Gie94, Hof91, Kay92, KV93, Mat91, Sav91b, WW92, Cle91b, CB91a, Fel91a, Kel91, Ová91b, Pet91, Ure91]. Metabolizable [KD97]. MetaNet [Koh92]. metaphor [KC91]. Metapopulation [BS98b, BOG^{+}99, Dam99, GM99, HZ93, KMJ98, MN94, RMH96, RGAP97]. Metapopulations [Par97, CC95]. Metazoa [KW98a]. meters [RL92]. Methanogenesis [FBS93]. Method [BBS^{+}98, CMPR93, CS96, Emm97, FKJ^{+}94, Fra97c, GGML93, GE96, MA99, RWK93, TW97, D.92, FC95, O’N91, OT90, PVT91, SK92, Tak95b, Tch91, WT90]. Methodological [Wäc97, vdS90]. Methods [AL96, MU92, Mar93c, ODA97, PGL94, SS96b, WD97, DM92a, FSDT91, MK92, SP92c]. methyl [PK92]. methylation [LWHZ91]. Methylglyoxalase [Kal97]. Meyer [Kat94a, Kat94b]. Meyer-Overton [Kat94a, Kat94b]. MgATP [IET99]. MgGTP [IET99]. MHC [For91b, Mal94b]. Mice [Boc98, JO96]. Michaelis [BIV97, CB91b, RP96, Sav95a, TH90]. Micro [HH97]. Micro-organisms [HH97]. Microbes [Kau98, Koc93]. Microbial [BB94, DVV99, KK94c, NKR97, Nar98, RLF93, BB91, HA90]. Microchip [SK99a]. Microcirculatory [KB94]. microclimate [McA91]. microcompartmentation [Her91]. microcomputer [YH90]. Microdensitometric [DH98]. microencapsulated [LBN95]. Microheterogeneity [MJLW96]. microheterogeneous [MS90a]. Microlocations [SK99a]. Micromechanical [SSGR^{+}98]. Micronucleus [LL93]. Microorganisms [HFS97, HV93, Mar95b]. Microparasites [BB99b, Bow99]. Microparasites-Avoidance [BB99b]. Microscale [DFG95]. Microscopic [HM95]. Microspora [Und90]. Microsporidia [Hai93]. microsporidian [Und90]. Microstructural [SFW^{+}96]. Microtine [JD99]. Microtubule [Sem96, SLBT99, THS^{+}95]. Microtubules [Yos99]. Microtus [AFS95]. microvascular [KH91]. microvessel [SL91c]. Midlines [KS98]. Migating [KL96]. MIGET [WGH98]. Might [KL98]. Migrants [Web99]. migrating [AP91a]. Migration [AP96c, HZ93, MPH99, Niw98a, Pen98, WH97, WH98a, WT94, Wie96, Jam90b]. Migrations [MPAA97]. Migratory [HA97]. Milk [FHB^{+}97, MFM97]. milleri [HC93]. mimesis [BBB90]. Mimicking [Kau98]. Mimicry [GH98a]. Mind [Bec96, BLLÅ94, Blo94, Lib96, Lib97, LÅ94, LA96b, Rol94, Sve94, vW94b]. Mind-Brain [Bec96]. Mind-Matter [Blo94]. Mineral [Kal98]. Mini [RBH98]. Mini-Review [RBH98]. minichromosome [SG90]. Minimal [BG97, CJMBM92, KB98b, vRGTV98]. minimising [WH97]. Minimization [BF99a, HA97, JW99]. minimizing [KL96]. Minimum [CAPP99, Pau93, PN98, Sib95, Yee91]. Mirror [HP97b]. misalignment [Str91]. Misleading [Sal91]. Mismatch [FK97]. Mismatch-stimulated [FK97]. misreading [FD92]. missing [RV90]. Mites [Nun98, Nun90, NLS91, NSL91]. Mitochondria [AM97, AS92]. mitochondria-luciferase-hexokinase [AS92]. Mitochondrial [All93a, All96, BA97b, BM95, GGML93, HS97, Kem94, Bra90a, Fer92]. Mitosis [KH98, Sem96, Meg90]. Mitotic [Abe98, BT98, NT95, TB91, UJ95]. Mixed [Ale97b, DVV99, LR94b, MMB99, Whi98, HVKS99]. Mixture [AM98, MA99, SEB^{+}93]. Mixtures [IET99, NKR97, Nar98, FAB92]. Mobile [Wit95, Now91]. Mobility [HLT96]. Mobilization [MKGB98]. Mode [LJ97, JM95, RH92, SC90]. Model [AVM^{+}99, ASA98b, ASA98a, AOL98, ABH94, AKL97, ARC97, Anc99, ABR97, ABR99, AK94, AFM97, AKV96, BTM93, BP97a, BEE97, BD94, BW98, BS98b, BvLVR96, BVvL96, BRV99, BB99a, BN97, Ben99, BB97b, BKK96, Bla94, BBL97b, BR94, BP93, Bon99, BT98, BGS^{+}99, BB93, Bow99, BGE94, BVW97, BVW99, Bri93, BCV97, BFK^{+}93, Bur94, BG97, CC97a, CC97b, CCFS96, CCS96, CK96, CI97, CLW93, Cho93, CL93, CMS93, CMG^{+}96, CMML96, CH94, CB94b, Cum96b, CS98, Dam99, DH98, DVV99, DST96, Di 94, DFG95, DO99, Doe97, DCL96a, DCL96b, DGSK99, DSD94a, DK95, ER94, Edm96, EM95, FV94, Fed93, FPB99, Flo98, FDT^{+}93, FHB^{+}97, FHR^{+}99, FR99, Fra97d, FCV99a, FBHK98, Gag94, GPL98, GR97a, GC98, Gou94, GR97b, GS93, Gre98]. Model [Gri94a, GLH97, HC94, HSM94, Hak94, HL96a, HDF^{+}99, HL98b, HV93, HH97, HW94, Hor93, HK98, IET99, JS97, Jea92, JK98, Joh96, Jos97, Jud97, KH99, KD97, KB98b, KD94, Kel95b, Ker96, KSM97, KMJ98, KC97, KM97, KK96, KK94e, KW98c, KCL^{+}96, KKH97, Kuz96, LL97, LLS94, LC97, LRMS98, LH94, LG99, LCP^{+}98, LR94c, LL99, Liu93, LBW^{+}95, LBP^{+}97, LW94, LL93, LVKH94, MFM97, MBM98, MCG95, MS98, MPH99, MKZ97, MH97, Mel93a, MAGO97, MS94, MAC99a, MAC99b, MC96, MKC^{+}98, MIT96b, MP98a, MCGPP97, MTLK96, ML94b, Mye93, NYI99, NNI98, NG99, Neu97, New97, NS96, NT99, Niw94, Nun98, NM93b, OSM95, OKAP98, Osb93, OS97, PL93, PG99, PLB96, PTBL98, QZDA93, RZF^{+}99, RAGO99, RLRV97, RL94, RBC98, Rau94, RBBG98, RGMG98]. Model [RWDM98, RS97, RN96, Rod96, RBM97, RLF93, RV99, RDHD98, RR96, RVMR99, Sak97, SP94, SP99, SEB99, Sav95b, SS98, SSS93, Sem96, SLBT99, SGP98, SSM93, SS96b, SWKS96, SB90b, SS93b, SAF^{+}93, SFGG^{+}99, SBRG99, Spi93, Sta98, Sta99, TN96, TH98a, TP94, Tan94a, TF96a, Toq97, TT93, TMZ99, Trz99a, Trz99b, TRVL96, TL98, TWB^{+}93, Tyv94, VK98, VVKZ94, VRH93, VR94a, VR94b, VC98, Vol97, WNY98, Wal98, Wat97, WH97, WZN93, WMLK96a, WMLK96b, WMLK96c, WOI99, Wes93, WLM96, Wie96, WLH94, Woo97, YCM97, YLT98, YC99, YFG^{+}93, YZ94, YQsZgZw97, Yi98, Yot93, ZL97, Zon98, dLA97, vOvP96, vOW99, vRGTV98, AGG91, ALW91, Arm90, AMT91, AHL95, BBB91a, BBB91b, BL95, Bak95, BSOB92, BBP95, BM95, Ber92, BE92, BZGM91, Bla92, BCQ92, BB91]. model [BT90b, BT90a, BBB90, BJZ^{+}91, Bur95, BH92, CD92, CCDD92, CSJM90, CS91, CSC90, CSS91, Cla95, Cle91a, CJ90, CJMBM92, CW92, CB91b, Cra90b, Cum90, DRLB91, Dib91, DM92b, Dug90, DSD92, DF91b, FTCM90, FC92, Fia92, Fie91, For92, FTL^{+}90, FBL^{+}95, Fu95, Gal92, GRRB90, GKMSS90, GGT92, Gre92, Grü95, HC91, HCS91, HS95, HB91, HZ95, HKT91, HH91b, Ish90, JFM90, JV91, Jär95, JB95, JG92, JP90b, JR92, JH95, KSG90, KG90, KA91, Ko91, Koc90, KB92, LOLM91, LOH92, LF95, LH91, LRS90, LL90, LG90a, LGBP92, LdQW95, LT91, Llo95, LG91, Lor90, LT92, MR92, MGL^{+}95, Man91, MPRB91, MRB91, Mar92a, MK90, Mil92b, MSR91, Mur90, Mur91, NH95, NZPB90, NS95, Nor92b, NT95, Nun90, NLS91, NSL91]. model [NM92b, OHS95, PW91, PNB92, PT92a, Pol92a, PN90, QC91a, QC91b, RBR90, RB91, RdM91, Ric90, ROMM90, Sak91, SAS92, SGG95, Sax92, SM92, SG90, SC92b, SM90b, SBM92, SH91b, SCG91, SDB90, Smi90b, Smo95, Sta92, SS90a, SV91, SH90a, Str95, Sýk92, Thi90, Tho95, TM92b, TH90, TCLB90, VHGV92, VSW97, VR90, VHSC92, Wei90a, WG92, Won90, Zhe91, sDA91, vVvP92, vdGvdBB92, vdLLH95, WMLK96a, WA94]. Model-Analysis [YC99]. Modeled [SA97, SNN^{+}97, TD96]. Modeling [AL96, AP96a, ABK94, CK97a, CK99, EEK93, EP94, FP93, Ham99, KMM^{+}97, Lac90, LRJ92, LCR94, LQW92, LQL94, MGP95, MP99, NT93, NH99, PMA97, ROL99, WRP99, ZvdBK97, Zon97, AKW95, BK90a, BS91b, CH90b, DS92, FPPSK95, GH91, ISN91, KL91, O’N91, SG95, WH92]. modelled [LSM^{+}92]. Modelling [AKMS97, BSW97, BA97a, Ben97, Boc98, CHM98, DSM99, Dav98a, EGDG^{+}98, FCV97, GMST99, Hah91, HBJ^{+}97, KWD^{+}94, KKA96, LBND98, LH90, LAV^{+}97, Mal94a, McL94, MvB93, PBJ^{+}92, PB97b, Rot95, SH97a, SV97, SB97, SGSS98, SK96b, TMI98, TB95, WEK95, WSF99, WS90, WBC99, ZH97, ZBB97, Eug95, GV91, MS95a]. Models [AP96c, BDK^{+}97, BB94, BLM94, CS94b, CFC98b, Com94a, CH96, CCD^{+}98, Dav91, DT98, DP98, DLC98, DM99, DL96, Dus96, EL96, FB97, FV93, FV94, FWF97, FH99, GLO99, Gat95, GH98b, Ham98, Hie97, HT97, Hin98, HL96b, IP99b, Jai97, Kei96, KD99b, Kok96, KKH99, KGFB97, LL94, LM96, LBWP93, Mar93b, MR98, MN92, MSSG94, Nab96, Neu98, OR97, PS93, Qui91, RA96, Rog95, Sat96, SNN^{+}96, SNK^{+}99, Smi90c, SH95, SSCC93, SS94c, SSGR^{+}98, Sza94, TGS93, TD92, VF98, Vrb98, WHL93, WT94, Wil94, WOQP93, WLR96, ZWH98, Ale91, AHA92, BS95, Bla95, Cha90, CLF95, Cra90a, DSP92, EKK98, Fad92, FS95a, Gro92, GMU95, JLL92, Jør95, Kir92, Koh92, Kra95b, Mar91, MS95a, NC92, PVT91, RM95, Rux95, SMA95, Smi90a, Smi90b, Sta90]. models [Stö95, Tch91, UJ95, WK90, WL91, WM95, sDA91, sDAB91]. Modern [Eld97, Tre95]. Modes [AW93, AP99, DC96a, HMUA93]. Modification [Ber91, MC96, ZW96, GW90, Wan90e, ZT92]. Modifications [HS96, Mor90]. Modified [GSS98, SK96a]. Modifier [ZW96, Top90, ZT92]. Modifying [AL96]. Modular [FW97]. Modulate [FK98]. Modulated [Bai93, GRRF93, TE96]. Modulating [HL98a]. Modulation [GCB96, MHM94, MR90]. Modulations [GCB96]. Modules [FW97]. modulo [AM92]. Moiety [CPK96]. Moiety-Conservations [CPK96]. Moisture [MvB93]. molds [MH90]. Molecular [ASML97, Bon99, BWK96, CF94a, DD98, DRB97, For94a, GH91, HWO90, Iwa93, Kre94, Li94, LBW^{+}95, MFG99, MMH90, MHM94, MHSR^{+}96, NT97, NT99, Plu94, RBD97, Sch99b, SS98, Sud90, SS94d, STI93, SS93c, Var98, WW97, BT91, CF92, DM92a, JH95, Kau92, KM95, Kre92, LT92, MBS92, NH95, NT95, PVT91, Sch91a, Sch91b, SHB90, Tch91, ZM95]. Molecule [Di 99, YY94, Di 92, Di 95, How92]. Molecules [Cum96b, KSF98, PAG99, PG94b, Ste94, Mar92a, SL91a]. Moles [JBSdlG97]. mollis [TLS94]. Mollusc [KM96b]. molt [Boi95]. Moment [CB94a, dLA97]. Moment-arms [dLA97]. monarch [WWC90]. Monitored [FR99]. mono [Dav91]. mono-oriented [Dav91]. monoclonal [OHS91]. Monoclonality [LBWP93]. Monoculture [Wes93]. monocyclic [Pat95, VH90]. Monofunctional [RSW96]. Monomeric [OES93]. Monomers [SS94d]. Mononeuronal [vOW99]. Mononucleotide [HK99]. Monophasic [KL99]. monophyletic [Di 99]. Monovalent [Gie97]. Monoxide [SSS93]. Monte [AB98, HH91b, KKV95]. Monte-Carlo [AB98]. Months [SLR^{+}98]. morphalactic [BT90b]. Morphogen [AP99, Ker96, KW98b, Lac90]. Morphogenesis [GKM93, MH97, MAB98, MJIB98, Osb93, SH97a, SV91, BL91, CS90, Cle91a, HE95, LdQW95, PP92, SH91a]. Morphogenetic [RGMG98, Cum90, LF95, VHSC92]. Morphogens [Osb93]. Morphological [KS96b, Par94]. Morphology [HW98b, Sek95, FAF90, JS92, VVC95]. Morphospace [ME93]. Mortality [Abe98, AP95, Rig94, RGAP97, JR92, McN95]. Mosaic [TMI98, TMI99]. moschata [HDF^{+}99]. Most [MRL96]. Motif [SW96]. Motifs [MJLW96, Mal94b, PGK94]. Motile [CFC98a, CFC98b]. Motility [Fra97d, MAC99a, MAC99b, Koc90, SL91c]. Motion [MC96, Ste94, SSGR^{+}98]. Motivated [TMI98]. Motor [Edm96, Nab96, Sem96, ST93, TS91]. Mould [AKMS97, vOPH^{+}96]. Mound [BVW97, VSW97, Ryt90]. mound-building [Ryt90]. Mountain [Has94a, PLB96, PTBL98]. Mouse [LL93, OHS91]. movelength [TCB91]. Movement [BVW97, BVW99, CH94, Gaá93, HC94, IET99, MB90, MBSM93, MHN^{+}94, Sem96, SSM93, ZGZ94, Dav91, HW92, Kum91]. Movements [Ale97a, Dru93, LSG97, McC93, OPRN93, dLA97, BSB90, DF91b]. Moving [MBM98]. mRNA [CK97a, CC96, GRRF93, Iid90, KO93, OK94, Pen91, Wil92b]. MS2 [GRRF93]. Mt. [AH96, AH98a]. much [CB91a, IMPR95]. Muller [For94b]. Multi [AH98b, Hin98, HMEW96, TD96, GM90, GWS90, VK91, WAJ91, vdLLH95]. Multi-Determinant [TD96]. multi-epitope [VK91]. Multi-locus [Hin98]. multi-looped [GM90]. multi-model [vdLLH95]. Multi-site [HMEW96]. multi-state [GWS90, WAJ91]. Multi-step [AH98b]. multicellular [KDG92]. Multicellularity [Sza94, Hur90]. Multidrug [AOU98, WZN97]. Multifactorial [WD97]. multifunctional [Boy95]. Multilocus [HT97, KKN98, KKN99, Sza93b]. Multiphasic [Vrb98]. Multiple [BLGG^{+}93, Fra98, Gie94, HKH93, Kau98, Mal94b, McN94a, NBBH99, WGH98, WC94a, BC90, Bar90, CY90, Joh95, Mar91, Mur91, NMS95, SAS92]. multiple-choice [Mar91]. Multiplicative [Sza93b, WH96]. Multiplicity [LJ95, Pat95]. Multiply [tBázsáSH96]. multiregional [Tre95]. Multiscale [GC98]. Multispecies [CHM98, CH96]. Multivalent [Hol93, CH90a]. multivalued [DS90]. Multivariate [Fra97c, JS92]. Murray [GE90]. Musca [WH94a]. Muscle [Ale97a, KS98, Kem94, MM95, MM97, Pen98, PK99, SAB97, San99, Tab98, TGS93, Zah96, Ale91, BT91, CKR95, MMH90, Mor90, Mor91, MBS92, OHS95, Smi90a, Smi90b, Smo95, WPP90]. Muscles [Min94, WB99, dLA97]. Muscovy [HDF^{+}99]. Muscular [MKC^{+}98]. musculus [LIWE90]. Must [Cam93]. Mutagenesis [OK94, Str91]. Mutant [AKL97, Bje96, LWS97]. Mutants [AH96, AH98a, BN94, HMUA93, vRGTV98]. Mutated [Bje94]. Mutation [DC99, Daw98, HS96, JTC99, LBWP93, Mal97, MR98, Sym94, Wol91, LWHZ91, SG91, Wei90a]. Mutational [GMR98]. Mutations [Bas98, HT96, Iid90, Isr96, Koe94, MR98, MK97b, AM95, BL95, FD92, Kon95, Kyp90, SHB90, SG91, Stö95]. Mutualism [Fra94]. Mx1 [BB99a]. Mycelia [Dav98a]. Myelin [Kau98, Liq91, PGK94]. myelinated [HC91]. Myeloperoxidase [Sta98]. Myeloperoxidase-Hydrogen [Sta98]. mykiss [EWHB96]. Myocardial [Reg93, Vol97, VR90]. Myocardium [BP99a, PH93, BHP92, Kee91]. Myocybernetic [BvLVR96, BVvL96]. Myopic [BM98]. Myosin [Edm96, IET99, MC96, MKC^{+}98, MBSM93, Mor96, MM97, CKR95, Hum92, MB90, Mor91, MBSM92]. Myotonic [KL98]. Mystery [KL99].

N [Gla91, GRUH99, HKT91]. N-termini [Gla91]. N-terminus [HKT91]. Na-Ca [VS97]. Na/K [Ste90a]. Na/K-ATPase [Ste90a]. NADH [FW97, Spi91]. Nagumo [Hor93, RL94]. Name [BS94]. Nasal [KSM97]. Nascent [TP94]. Nash [Cra90a, Rob90b, SH93b]. natal [KS99]. native [Kyp90, SHB90]. Natural [CBSV95, Chi96, Fle97, HDF^{+}99, Kuz96, LCP^{+}98, Nak98, Ng91, PAG99, Sau94, Sza92a, PNB92, YC91, ZS91]. Nature [Baj99, FSPC98, LT98, MB93, RV99, SSA94, HI92, Pri95, RHS^{+}95]. Navigating [McC94]. Navigation [Phi96, Wal98, BBP95, Qui91]. NCP [PDP97]. Neanderthal [Flo98]. Near [KKS97, MS95b, Gla91, Nus91]. Near-Far [MS95b]. Nearest [HL98b, LTZ96, ZW95]. Nearest-Neighbor [HL98b, ZW95]. Nearly [HT96]. Necessary [GCB96, YZ94, ZJ94]. Necessity [RBM97]. nectar [CS91, FS95b]. Need [NS99, PGBYP97]. Needed [CAPP99]. Negative [KPS97, LJ95, Sat96, BH95, GM90, KW94, MGP95]. Neighbor [HL98b, ZW95]. Neighborhood [PDP97, Arm90]. Neighbour [LTZ96]. Neighbourhood [Sas97]. neighbours [WS90]. Neocortical [Wil99]. Neophenogenesis [JG90]. Nephrotic [CLW93]. Nernst [Kat95]. Nerve [LPM95, RL94, Cla95, HC91, Kum91, SB90b]. Nest [HDF^{+}99, Mot93, ORM99, CH91, KSG90]. Nesting [KSG90]. Net [KS98, YYYU97, MH90, Sax92, SB90b]. NetNews [BP99b]. Network [BdS97, Bur94, CBSV95, CCFS96, CCS96, DBSV94, DSP96, FV93, FV94, FCV97, KK94e, LCP^{+}98, LSG97, MAB98, SP92a, SP92b, TD96, TCZ95, VKH^{+}98, WZN93, WP93, Whi98, vOvP96, vOvO97, Bur95, HBP92, Mar95b, SV91, VVC95, VK91, dBP91]. Networks [BP97b, Bea93, BC97, CMS93, Fra99, Fre94, HS98b, Kle98, OKAP98, SV97, TPN98, WEK95, vOvP94, Cou91, KH91, KL91, MPO95, SV90, Tho91, VS90, WDP90]. Neural [Fre94, LLS94, LSG97, Nak91, Orp94, TCZ95, vOvP94, vOvP96, vOvO97, Eug95, VK91]. Neurite [LQW92, RS97, vOvP96, vOvO97, VVC95]. neurites [vVvP92]. Neuroendocrine [MTLK96]. Neurogenesis [KS99]. neuromodulator [Kum91]. neuromodulator-induced [Kum91]. Neuromuscular [BFG96, vOW99, SWA91]. Neuron [SLR^{+}98, TRVL96, BJZ^{+}91, HI92]. Neuronal [CS99, Joh93, ML94b, RBC98, LdQW95, Smi92, VVC95]. neurone [Orp90].

Neurons [BKK96, BS98c, Gaá93, KK96, Orp94, Wil99, vOvP94]. Neuroscience [Hes94]. Neuroscientific [Rol94]. Neurospora [MR98, RVMR99]. Neurotransmitter [CB97a, KPS97, LPS90]. Neutral [BRV99, HT96, NH96, SG91]. neutralizing [FBHK98]. Neutron [Mor96]. Neutropenia [HHM98]. Neutrophils [AOL98, HP97a]. Newborns [CY98]. Newtonian [DB95, Niw96]. Niches [Ng96, HBP92, PCN95]. Nicholson [Doe97]. Nicholson-Bailey [Doe97]. Nicotinic [Got94]. niger [BDG92]. Nitric [Sal96]. Nitrifying [AIK99]. Nitrogen [Sta99, Sto94, WGH98, vRGTV98]. nitrogenase [Str95]. nitrous [KPK92]. NK [SBRG99]. No [Lor94a, Pet91, MHTS90]. Noble [GLH97]. nodal [SAGB95]. Node [DCL96a, PST91]. nodules [Str95]. Noise [Col93b, DGSK99, Hor93, Pos99, SKI98, VB93, WN99b, Vid91]. Noise-induced [SKI98]. Non [Alm99, AK94, AMO93, BBM^{+}97, BBM^{+}13a, BBM^{+}13b, BBM^{+}13c, CK96, DvH93, DB95, DD98, Di 99, Don90, Dru97, DSD94a, ER94, EV94, GR97c, HRW93, Hav99, Hin98, Hoc91, JD99, KPK92, LCR94, LKS^{+}93, MKM96, Mar93b, Mat96, NVB99, RAGO99, SW98, Whi98, Zah96, BC90, BT90b, BT90a, CLF95, FSD91, JV91, Jär95, Kat90, Man92a, SF90, SB90b, SWA91, Sta90, Ste90b, Str95, Whi91]. Non-adaptive [DvH93]. Non-additive [Hin98]. Non-axial [Zah96]. Non-catalytic [Hav99]. Non-coding [Alm99, AMO93, SW98, Man92a]. non-cognate [Kat90]. Non-compartmental [Mar93b]. non-coplanarity [KPK92]. non-covalent [BC90]. non-cycling [Whi91]. Non-decaying [DSD94a]. non-diffuse [SB90b]. Non-enzymic [HRW93]. Non-equilibrium [DD98]. non-excitable [CLF95]. non-exclusive [Jär95]. non-ideal [SF90]. Non-inactivating [LCR94, Ste90b]. Non-independence [GR97c]. non-invasive [Str95]. Non-linear [ER94, EV94, Hoc91, BT90b, BT90a, Sta90]. Non-local [BBM^{+}97, BBM^{+}13a, BBM^{+}13b, BBM^{+}13c, LKS^{+}93]. Non-monophyletic [Di 99]. Non-Newtonian [DB95]. non-overlapping [JV91]. Non-parametric [Don90, NVB99]. Non-planar [KPK92]. Non-preferred [JD99]. Non-random [AMO93]. Non-Saturating [MKM96]. Non-specific [AK94]. Non-Stationary [Mat96]. non-steady [FSD91]. Non-Transcribed [Dru97]. Non-trophic [Whi98]. non-uniform [SWA91]. Non-uniformity [CK96]. Non-Uniformly [RAGO99]. Nonbalanced [CC97b]. noncoding [AM90]. Nonlinear [Ano97-28, CCD^{+}98, GBD99, GMN96, Hem98, MD96, Vol97, HE95, Nak95]. Nonlinearities [BSW97]. nonspecific [Wan90c]. Nontrophic [WP93]. Normal [KWD^{+}94, Koz96, Gin95, Jol90, JD92]. Normality [Koz96, Gin95]. Normalization [All93c]. Norms [SP94]. Not-self [For94b, For92]. Notch [CMML96]. Note [Sza93b, Wei97, HN91, Smi92, SK95]. notice [BBM^{+}13c]. Notion [AO96]. Novel [CAPP99, KV93, RH96, ZLYZ98]. November [Ano90u, Ano90-29, Ano91x, Ano91-27, Ano92r, Ano92y, Ano93f, Ano93i, Ano94m, Ano95b, Ano95h, Ano96d, Ano96j, Ano97e, Ano97k, Ano98h, Ano98n, Ano99l, Ano99-30]. Nuclear [Bai93, Ker96, KL98, LPP^{+}94, Mau92, SB90a, VpM95]. nuclear-cytoplasmic [Mau92]. nuclease [Boe90]. nuclease-sensitive [Boe90]. Nucleation [GLO99]. Nucleic [NGP97]. nucleocytoplasmic [SMA95]. Nucleon [RJ96]. nucleoside [SOR91]. Nucleosides [GSS98]. Nucleosome [CF94b]. Nucleosomes [MGR97]. Nucleotide [DY99, HS97, ONT97, WZ96, AM92, LB95, ROMM90, SYK91]. nucleotides [Nus91, SYK91]. Nucleotidic [CMPR93, CCRM90, RCCM92]. Number [AF99, Cam93, CAPP99, Eas99, LT98, MAGO97, Mor98, RJ96, SK99a, SLR^{+}98, Cam95, CK95b, For91a, Rak98, SP92e]. Numbering [MP98b, MP98c]. Numbers [ME93, WS90, YYY91]. Numerical [BP97b, FV94, FK99, GKMSS90, HL96a, Kat95, KSM97, MAC99a, MAC99b, PL93, Ste94, SGSS98, LSM^{+}92, SL91a]. Numerically [Ham98]. Nutrient [Com94a, Hir93, Her90]. nutritional [KE92].

O [CLY99]. obesus [KD93]. object [May90]. Objective [Wil93a]. Objectives [Irv90]. Objects [Chi96]. Obligate [RP98]. Obligatory [Wed94]. Observation [Kra99]. Observations [AFS95, UJ95]. Obstacle [PK93]. Obstacles [Rol94]. Obtained [Mor96, SSCC93]. Occupancy [KMJ98, NNM99, Trz99a, Trz99b, WMLK96a, WMLK96b, WMLK96c]. Occurrence [CMPR93, MJLW96, MR98, WR94, SNG92]. ocean [Jam90b, Qui91]. Oceanic [KKS97]. October [Ano90-35, Ano90-41, Ano91-33, Ano91-39, Ano92-32, Ano92-38, Ano93r, Ano93u, Ano94q, Ano94v, Ano95t, Ano95y, Ano96m, Ano96v, Ano97r, Ano97x, Ano98v, Ano98-27, Ano99u, Ano99y]. Ocular [Bet99, Rei95]. Odorant [KSM97]. Off [Got94, Daw96, Eug95, NS97]. offs [HW98a]. Offspring [FM98, Jam99a, LS94, Møl97, PRG98, TS94, Zha98, For90, For91a, GD90, ZJ98]. Often [Rod98]. OH [NK90]. OH-permeabilities [NK90]. Oil [Jon97]. Old [BS94, Ng91]. Olfaction [HSM94, KSM97]. Olfactory [LPM95, TRVL96]. Oligomer [MJLW96]. oligomers [SNG92]. Oligonucleotide [AIK99, CCM^{+}97, HK99, PG94b, SK99a, FPPSK95]. Oligonucleotides [SK99a, NS91]. Oligopeptide [Ste94]. Oligoribotides [FC93]. omeprazole [Bro90]. On-Off [Got94]. Onchidium [HI92]. oncopelti [SH90b]. Oncorhynchus [EWHB96]. One [EEG93, FL96, Jam98b, KS96a, Por96, TUHT97, ZJ94, ZH97, BS95, DSP92, FBH91, Fu95, Tho95, Van91, VHGV92]. one- [DSP92]. One-dimensional [ZH97, FBH91, Tho95]. One-dimensionally [TUHT97]. One-locus [ZJ94]. one-parameter [Fu95]. Only [DBAP^{+}98, JBSdlG97, Nuy94, YZ95]. onto [BFG96]. Ontogenesis [CCS96]. Ontogenetic [McN95]. Ontogeny [DC96c, Mar95b]. oocysts [PBJ^{+}92]. oocytes [Jaf95]. Oparin [Laz97]. Open [Col93b, Ros94b, Bat91, CK95b, MS92a, Qui91]. open-grown [MS92a]. Opening [BP94, KN93, BP92]. Openness [GL95]. operation [BR91, CCFS96, CCS96, Dug90, Mot91, Pec93, Sza93a]. Operational [Sav91a, Trz99a, Trz99b]. operative [LOH92, PA90, PT92a, PT92b, TP93, Vid91]. operativity [CLW93, RV90].

Operons [OWM96]. Opponent [Chi96, Fra90]. Opportunity [CPJL92, She91]. Opsariichthys [Kat96]. Optical [BM98]. Optics [HC93]. Optimal [All93b, All93c, AC90, Bea93, Chi96, Ezo98, For91a, Fra97b, Fra95c, Han93, HH95, Hut99, KIA93, KP93b, KB91, KB93b, KB94, Kok96, LHA94, LP93, MB98, MR91b, McN94b, MAGO97, Mor98, Nis99, OA98, PH90, Rot95, SH93a, Sak94, Sak95, SP92c, SP92d, SH93b, SM97e, SNH99, Sto94, TN96, TC97, Tay91, Van92a, Van92b, VP93b, WNY98, Fad92, SS90b]. Optimality [KWB93, BK90a]. optimally [LÅ90b]. Optimising [dLA97]. Optimism [FM98]. Optimization [ALS93, DCM94, KH94, MHWM94, NM93b, San99, SNN^{+}96, SNN^{+}97, SNK^{+}99, MR91a, Pol92b, SP92a, SP92b]. Optimize [SLR97, QG92]. Optimum [Ale91, Ale97a, DM92b, HA97, TA99, WJ97, vL91, GE90, KB92]. optional [BK95]. oral [RBR90]. oranges [Enr90]. Orbital [NT97, NT99, RM93, STI93]. Order [HE93, Ste94, BW92, BSM90, CSFC91, CK95a, HH93, LVL92, MS90a]. Ordered [ORM99]. ordering [Nus91]. ordinary [Jol90]. ordinate [Gaá93]. Ordination [GE96]. organ [Mar95a]. organelle [Miz95]. Organic [SV97, Tar94]. Organisation [Blo97]. Organism [Sea99]. Organismal [Bjö94]. Organisms [Bur94, HK99, Isr96, JIDH97, KKH99, MNH97, Van92a, Wit95, ZJ97, Zha98, ZJ98, HH97, KDG92, WBS95, Zha90]. Organization [BJ96, Hir93, Jen98, Jon96, LBWP93, LBP^{+}97, MTM^{+}96, Ned97, WE98, BSM90, CS91, GL93, HM95, KP93a, Nor92b, TBD95, Wil93b]. Organizational [YC91]. Organized [KM96a, SA97, WC95b, CSJM90, KD98, Mil92b, ORM99, SM95]. Organizes [PMdM^{+}99]. Organizing [DC96a, DC96b, DC96c, MSN96, Niw94]. Orientation [AP91a, DSM99, HV93, HP97b, MNB97, Pol96, CD92, CCDD92, Qui91, TCB91]. Oriented [BSM97, Dav91]. Origin [Ano94b, Bal97, BBL^{+}97a, Bas98, BA94, Blo97, Di 94, Di 97, Di 98, Di 99, Edw96, Fra95b, FW97, GOM94, JSB93, KW98a, Key97, Lew98, Li94, RBD97, Rus97, SLS93, SS93a, TLS94, Wäc97, Cha91, Di 92, Dib91, Hol96, How92, Ish90, Ss90c, Sza92b, Tre95]. Originates [Abe98]. Origins [Nor93, PRG98]. Ornaments [Cor98]. Orthogonal [KB99]. Orthophosphate [KB93a]. oryzae [WLR96]. Oscillating [TP96]. Oscillation [DSD94a, TH98a, TUHT97]. Oscillations [BBL97b, BG97, DSD94b, FKJ^{+}94, Fri98, KD94, LC97, LG99, LR94c, Lin94, Mur90, NT93, Rou98, SLBT99, SMG93, AS95, CJMBM92, DSD92, Mur91, Smo95, VR90]. Oscillator [BD94, HK98, JK98, KH99, Nun98, TUHT97, NLS91, NSL91, SM92]. Oscillators [SS94b]. Oscillatory [Ale97a, DDS96, NU96, SV90]. Osmotic [Kab94, Koc95]. Osmotolerance [SM97d]. osmotropotactic [CD92, CCDD92]. osteoarthosis [PN90]. Osteoarthrosis [MPSK94]. other [Hem90, Jon97, LOB92, MN92, VHG^{+}90]. Outbreaks [SB94, GB95]. outdoors [McA91]. Outer [CMB95]. Outgrowth [DO99, vOvP94, vOvP96, vOvO97, VVC95, vVvP92]. Outlines [CG96].

Output [JM98, JM95, LOB92, vL91]. outsider [Pec90]. Ovarian [JSB93, SM92]. ovarian-cycle [SM92]. Over-stimulation [Wic99]. Overdispersed [Iwa93]. Overdispersion [BRV99]. Overlap [KOTL97]. Overlapping [Sib94, JV91]. Overshoot [vOvP94]. Overton [Kat94a, Kat94b]. overwinter [WWC90]. Ovule [Sak96]. Owner [ES95]. Owner-Intruder [ES95]. Oxford [DM92b]. oxidase [BW92]. Oxidation [KQ96, SS98]. Oxidative [SS94a]. Oxide [Sal96, Sta99]. Oxide-Superoxide [Sta99]. Oxidoreductase [FW97]. Oxygen [BIV97, DP95b, KWB93, WGH99, Man95, WPP90, PAS91]. oxygenation [LCD95]. oxytocin [WCPR91]. oyster [Han91].

P [Bec96, BP93, WBC99]. Pacemaker [ASA98b, ASA98a, BKK96, DCL96a, DLC98, End97, JK98, KK96, HI92]. Pacemakers [HE99]. Pacemaking [VW94a, SCG91]. Pacific [Jam90b, Qui91]. Pack [WLM96]. packing [MR91b, RRB90]. Pages [Ano90-33, Ano90-39, Ano90-32, Ano90-38, Ano90-34, Ano90-40, Ano90-30, Ano90-37, Ano90r, Ano90x, Ano90-31, Ano90-36, Ano90s, Ano90y, Ano90w, Ano90-27, Ano90t, Ano90z, Ano90u, Ano90-29, Ano90-35, Ano90-41, Ano90v, Ano90-28, Ano91-32, Ano91-38, Ano91-35, Ano91-41, Ano91-34, Ano91-42, Ano91-36, Ano91-40, Ano91v, Ano91-29, Ano91-31, Ano91-37, Ano91u, Ano91-30, Ano91w, Ano91-28, Ano91s, Ano91y, Ano91x, Ano91-27, Ano91-33, Ano91-39, Ano91t, Ano91z, Ano92-33, Ano92-39, Ano92-31, Ano92-37, Ano92-34, Ano92-30, Ano92-35, Ano92-29, Ano92t, Ano92z, Ano92-28, Ano92-36, Ano92q, Ano92w, Ano92u, Ano92-27, Ano92p, Ano92v, Ano92r, Ano92y, Ano92-32, Ano92-38, Ano92s, Ano92x, Ano93p, Ano93v, Ano93s, Ano93x, Ano93o, Ano93y, Ano93q, Ano93w, Ano93c, Ano93j, Ano93n, Ano93t, Ano93g, Ano93m, Ano93e, Ano93l]. Pages [Ano93b, Ano93h, Ano93f, Ano93i, Ano93r, Ano93u, Ano93d, Ano93k, Ano94n, Ano94t, Ano94o, Ano94u, Ano94s, Ano94x, Ano94p, Ano94y, Ano94d, Ano94k, Ano94r, Ano94w, Ano94h, Ano94i, Ano94f, Ano94g, Ano94z, Ano94e, Ano94l, Ano94m, Ano94q, Ano94v, Ano94c, Ano94j, Ano95p, Ano95v, Ano95r, Ano95x, Ano95n, Ano95u, Ano95m, Ano95s, Ano95c, Ano95i, Ano95q, Ano95w, Ano95f, Ano95l, Ano95d, Ano95j, Ano95e, Ano95k, Ano95b, Ano95h, Ano95t, Ano95y, Ano95g, Ano95o, Ano96l, Ano96r, Ano96n, Ano96s, Ano96p, Ano96u, Ano96o, Ano96t, Ano96c, Ano96i, Ano96k, Ano96q, Ano96g, Ano96y, Ano96e, Ano96w, Ano96f, Ano96x, Ano96d, Ano96j, Ano96m, Ano96v, Ano96h, Ano96z, Ano97t, Ano97z, Ano97u, Ano97-27, Ano97p, Ano97v, Ano97q, Ano97w]. Pages [Ano97d, Ano97j, Ano97s, Ano97y, Ano97i, Ano97o, Ano97g, Ano97m, Ano97f, Ano97l, Ano97e, Ano97k, Ano97r, Ano97x, Ano97h, Ano97n, Ano98r, Ano98w, Ano98x, Ano98-28, Ano98t, Ano98z, Ano98q, Ano98u, Ano98e, Ano98k, Ano98s, Ano98y, Ano98j, Ano98p, Ano98f, Ano98l, Ano98g, Ano98m, Ano98h, Ano98n, Ano98v, Ano98-27, Ano98i, Ano98o, Ano99w, Ano99-28, Ano99t, Ano99z, Ano99r, Ano99s, Ano99v, Ano99-27, Ano99j, Ano99p, Ano99x, Ano99-29, Ano99h, Ano99n, Ano99i, Ano99o, Ano99k, Ano99q, Ano99l, Ano99-30, Ano99u, Ano99y, Ano99g, Ano99m]. Pair [MFG99, Myl99, Lac90, Wag93, SW98]. pair-rule [Lac90]. Pairing [Jai97, DKL95]. Pairings [DST96]. Pairs [ONT97, KPK92]. Palm [Jon97]. pancreatic [AGG91]. Panels [HP97b]. paper [Por96]. Parable [Sau94]. Paradigm [Dum94]. Paradox [For94b, MH96, Beu90, GWS90, Ng91]. Paradoxical [PK98]. Parallel [KOTL97, LSB^{+}97, RP99, Bra90b, Fer92, GGT92, Nak95]. Paramecium [LH90, LH94, LF95, LH91]. Parameter [CAPP99, CB91b, HL96b, SSCC93, Fu95, HC91, HH91b, TH90, vRGTV98]. Parameters [BDDF97, Fra97b, KH94, MRB91, San99, SFHP97, WBC99, FTD90, HH91a, MPRB91, WKSH92]. Parametric [SGP98, Don90, NVB99]. parametrization [MU92]. parasagittal [Kuz95b]. Parasegment [SB97]. Parasegment-specific [SB97]. Parasite [DG93, HS96, Hur90, NBBH99, WS94, Yam99, Atk91, CPJL92, GG97b, JFM90, RdM91]. Parasite-Mediated [WS94, CPJL92]. Parasites [CB94b, HH98, LM94]. Parasitic [TF99]. Parasitism [NM93a, RdM91, HS95]. Parasitoid [CH96, KYH99, SRH97, RRG98, WJH97]. Parent [Hou90, IP99b, LS94, Mot93, PRG98, TS94, Zha98, ZJ98]. Parent-Offspring [LS94, PRG98, TS94, Zha98, ZJ98]. Parental [BEE97, Daw96, FM98, Get93, Jam96, Jam99a, SH94, BK90b, GD90]. parents [MB92]. Parish [DSD94b]. Parity [BF99c, TA99]. parsimony [Bea90]. Part [CC97b, CL99b, DC96a, DC96b, DC96c, Gra93, GS93, HH98, KG90, Shc94, Trz99a, Trz99b, Zhu91, CC97a, CL99a, GK90, JP90a, JP90b, Shc93b, ZJ98]. Parthenogenesis [Gag94]. Parthenogenetic [SLS93, TLS94]. Parthenogens [RP98]. Partial [CB94b, Zhu96, Hem90, Rad91]. Partially [Jam96]. participates [Hav91]. Participating [AH98b]. Particle [rvdBvdG96, CB97b]. Particles [Ber93, MNH97, TP94, vdGvdBB92]. Partitioning [FHR^{+}99, KD93, Wes93, FBL^{+}95, She91, Wal91]. Partnerships [CW98]. parvoviral [TCLB90]. parvoviruses [FM91]. passage [Fad92]. Passive [CBS^{+}99, FH99, SAM97, Vin98]. Passive-homeostasis [Vin98]. Patch [KMJ98, LGM98, MHW93, Mor98, Nis92, Val92, PH90]. Patches [BSTV93]. Patchiness [LPRB96]. Patchy [MPAA97, Ben92, SS90b]. Path [KIA93, Key97, Rot95, Wal96, Whi98, BDG92, BS95, MS95a, SS90b, ZM95]. Path-based [Whi98]. Pathogen [AP95, AWA98, BB94, CGM^{+}98, Dam99, Lew98, BB91]. Pathogenesis [MPSK94, RBM97, Whi97]. Pathogens [LM94, Sas94, GB95, MN92]. Paths [McN93, MH96]. Pathway [Kal97, MFB98, Nöl98, Nöl99, SSCC93, Sre91]. Pathways [Ben99, Cum99b, GCB96, Ker96, MHWM94, BS91a, FO92, Sri91]. Pattern [AF99, BSTV93, CMML96, Cum99b, DSP92, FST99, Hem98, KMJ98, KM97, Kor97, Lub97, MJIB98, MCGPP97, NYI99, Orp94, OS97, PC97, RGMG98, RWB98, Sac94, SvHT97, Sen96, Spi93, TSC90, WD97, vOPH^{+}96, AP91b, BT90b, BT90a, CSJM90, Cum90, GK90, Hoa90, KG90, Lac90, LQW92, MM91, Sav95b]. Pattern-form [Spi93]. patterned [VpM95]. Patterning [Cum99a, DO99, KCL^{+}96, SAF^{+}93, EMRS90]. Patterns [AP96a, BSM97, BFG96, Con93, DC96b, GL93, KMM^{+}97, MAGO97, MAC99a, MAC99b, ND99, NM93b, SKI98, SRH97, TMI98, VP93b, WR94, BH95, GR95, Hem90, Hor95, KJ95, McK95, McN95, Mor92, Nus91, OHS95, Res95, Vic91]. Pauses [GRRF93]. Pay [NS97, Kur99]. Pay-off [NS97]. PCO [SC92a]. PCR [SM97b, DA96, KDBK93]. PCR-derived [KDBK93]. Peak [WH96]. peaked [WBS95]. pelagic [GS95a]. pelagicus [Gri94a]. Penguin [DVJ^{+}99]. Pentose [MHWM94, Rog95]. peptidases [LBDDG90]. Peptide [DBB99, FC93, FR99, MCG95, Rot95, WW97, WCPR91]. Peptide-assisted [FC93]. Peptides [FHB^{+}97, Kau98, PSLB99, RBH98, IMPR95]. Peptidyl [DBB99]. Percolation [GG97a]. Perennial [TN96, Zha98, ZJ98, Tak95a]. perennials [Ing90]. Perfect [Doi97]. Performance [Tho96a, Hed92]. Performed [Got94]. perfused [HCS91]. Perfusion [WGH99]. Period [AH94, Hor93, KE92, LWS97, SAGB95]. period-doubling [SAGB95]. Periodic [AH94, Ber95a, DC96a, YQsZgZw97, vOvP96, LCD95, SH91b]. Periodicities [AM90, AM92, ALM95, CT95]. Periodicity [Bai93, DC96b, DSD93, GOM94, LRB94, TET91, WM95, KS91]. Periodograms [DR91, Enr90]. Peripheral [HHM98]. Permanence [OSM95]. Permanent [Kor93]. permeabilities [NK90]. Permeability [KC97]. permeable [DB95, PD93]. Permease [Yar96, Rad91]. permeation [DL90]. Peroxidase [Sta98]. Peroxidase-Chloride [Sta98]. peroxide [BW92]. Persistence [AHA92, Che93, CH96, Doe97, GG97b, NBBH99, RA96, RM95]. Persistence-time [AHA92]. person [Dug90, MI98]. Personal [Ace96b]. Perspective [Dem98, Par94, SQ98, Sym94, Mil91a]. Perspectives [KD93, Rus97, HBP92]. pertinent [Res95]. Perturbation [CS94b, TA99, WG97a, Man95]. Perturbations [Kat94a, Kat94b]. Pest [BB94, BB91, CK95a]. pestilence [DSD92]. Peter [BEE97]. pH [AKMS97, Fri98, NK90, Wan90d, WSF99, sDAB91]. Phacops [HC93]. Phage [MCG95, RZF^{+}99, SK96a, RV90]. Phagolysosomes [Sta98]. Phagotrophic [HFS97]. Phantom [Wal99b]. Pharmacodynamic [Liu93]. Pharmacokinetics [AWA98, TWW90]. Pharmacological [Ben99, MAC99b]. pharyngeal [Gal92]. Phase [AVM^{+}99, ASA98a, LGBP92, LWS97, NU96, SS94b, YFG^{+}93, AKMS97, CLF95, LOLM91, Mar92a, NT93, SCG91, SF90, Sme90, VHG^{+}90, VH90, Whi91, ZCS^{+}90]. phase-locked [CLF95]. phase-sensitivity [SCG91]. Phases [RBBG98]. Phasic [UP96]. Phenology [Sak94, FAF90]. Phenomena [GLO98, Her94, vOvP94, Mil91a]. Phenomenological [FH99]. Phenomenon [TS94]. Phenomes [MK97a]. Phenotype [YZ94, YQsZgZw97, Yi98, FS98]. Phenotypes [KS96a, LJ96, MK97a, RGK91]. Phenotypic [BN96, Fin92a, GVL93, BN95, JG90]. Pheromone [Wat97]. Philip [JG92]. Philosophical [Rus97]. Philosophy [BLLÅ94]. phocid [Boi95]. phosholipid [SL91a]. Phosphatase [EEG93]. Phosphatases [Cum99a].

Phosphate [ASML97, MHWM94, VK98]. Phosphate-Starvation [VK98]. phosphatidylinositol [HKT91]. Phosphoenolpyruvate [Ric96]. Phosphofructokinase [Ric96, Smo95]. phosphogluconolactone [RP99]. Phospholipid [HBJ^{+}97, Nor92b, Mar92a]. Phosphorus [Rob90a]. Phosphorylase [CG96]. Phosphorylation [EEG93, SS94a]. Phosphorylation-Dephosphorylation [EEG93]. Photo [TP96]. Photoacclimation [Zon98]. Photoadaption [Zon97]. Photoinhibition [Zon98]. Photoneuroendocrine [WW94]. Photoperiodic [Dum94, Nun98, Nun90, NLS91, NSL91]. Photophosphorylation [SS94a, SS98]. photorespiration [Hah91]. Photosynthesis [Fri98, Hah91, LAV^{+}97, Sto94, WBC99, Zon97, Sti91, Tor90]. Photosynthesis-Irradiance [Zon97]. Photosynthetic [BBL97b, Sak94, SP94, SS98, WAO^{+}99, Bea90, LOB92]. Phototactic [HV93]. Phototranslocations [Mal94a]. Phototrophic [AIK99]. Photovoltage [DPG^{+}94]. Phycomyces [EV94, WG92]. Phyletic [PSLD96]. Phyllotactic [BJ96]. Phyllotaxis [CS98, DC96a, DC96b, DC96c, Yot93, Jea92]. Phylogenetic [FS95a, KS96b, ONRH96, PGL94, TW99, WN99c, Wil98, Fel91b, Gra92]. phylogenies [FC95]. Phylogeny [ONT97]. Physarum [MTM^{+}96, NU96, NYI99, TUHT97]. Physical [JW99, NSW97, SM97d, Del95, Doe95, McC91, Mil91a, Ss90c]. Physicist [Blo94]. Physicochemical [DCM94]. Physics [BLLÅ94, FGLO99, Hop94]. Physiological [DGSK99, Fel91a, Med95, Ová91a, Ová91b, Sri91, Bat91, Cle91b, Kel91, Man95]. Physiologically [KKH99, Cár91]. Physiology [Ano99a, SKW98, Aok92, Cor95, WM91]. phytochrome [GH91]. Phytoplankton [HVKS99, KKS97, ZvdBK97, Zon98]. piceatus [Gal92]. picture [MPRB91]. Pigeons [Wal98, Wal99a, Wal99b]. Piggyback [RBH93]. Pigmentation [KM96b, MM91]. Pillow [Gie91]. Pinaceae [FF90, FAF90]. Pine [PLB96, PTBL98]. pineal [Car90]. Pinwheel [Rot98]. Pioneer [Pea96]. pipits [FAB92]. pituitary [BZGM91]. place [BBP95]. Placenta [BU98]. Placentation [Bla98, QAM97, Bla95]. plague [DSD92]. Planar [dLA97, KPK92]. Planck [Kat95]. planetary [Mar95b]. planktivores [GS95a]. Planning [BHVC90]. Plant [AO96, BA97b, BG97, Cha93, Com94a, Dam99, GPL98, GC98, GH98b, JD99, KG93, LGM98, LBND98, Mel93a, NVB99, RB97b, SC96, SW95a, TAN94b, BN92b, FB95, PW91, PCN95, Sak95, SW92]. Plant-herbivore [GH98b]. Plant-Soil [Com94a]. Plants [Bar94, Gou94, RGMG98, Rod96, SH93a, Sek95, SM97d, SGSS98, TN96, WBC99, Zha98, ZJ98, AC90, Bak95, HP95, JS92, Sak91, Tak95a, VHSC92]. Plaque [YY99]. Plasma [KSF98]. Plasmid [Bas98, SP92e, RSB92, Sýk92]. Plasmid-borne [Bas98]. Plasmids [ACC96, BA97b, Sýk92]. Plasmodium [How92, MB97, MB98, NU96, NYI99, MMB99, MTM^{+}96]. Plasmolysis [Koc95]. Plasticity [BN96, GVL93, BN95]. Plastid [How92]. Plating [Jon94]. Plausibility [WR93]. played [NS91]. player [Bla97]. Playing [BP93, Mot91, Sza93a, Cra90b, Cra90a]. Pleasure [Cab92]. Pleistocene [McK95]. plexippus [WWC90]. plexus [SB90b]. Pliocene [McK95]. plot [Wan91]. Plots [Seg94]. pN. [Mor91]. Pocket [Nor93]. Pogonomyrmex [CH94, HC94]. Point [JD92, MR98, WD97, WW97, FL95, HT95, Kyp90, SHB90, SG91]. Points [SFGG^{+}99]. poised [KJ91]. Poisson [Kat95, SL91a]. Pol [YFG^{+}93]. polar [Meg90]. Polarity [DPG^{+}94, Ss90c]. Polarization [BGE94, Dem94, HP97b, Hor95]. Polaron [CL93]. Pole [KW94]. Poles [KD99a, Koc92]. Polistes [TBD95]. Pollen [RGMG98, ZJ97]. Pollinated [Gou94]. Pollination [FS95b, PSE92]. Pollinator [Sak96]. Poly [Mat93, vOW99, Boe90]. Poly- [vOW99]. polyadenylate [Coh91]. polycephalum [NU96, TUHT97]. Polydactyly [Jam98b]. Polydrusus [TLS94]. Polyethism [WNY98]. polyethylene [RB91]. Polygenic [Bür93a, Gav97]. Polymer [RB91, SW95a]. Polymer-mediated [RB91]. Polymerase [Col93a, Mat93, SM97b, SM97c, VK99, CF92, Coh91]. Polymerization [AOL98, Coh91, Kuz95a, Meg90]. Polymers [MP98a, MP98b, MP98c, Ost96, Won90]. Polymorphic [NH96, JM95]. Polymorphism [CTV93, ZJ94, vBW98, For91b, Mau92]. Polymorphisms [KKN99, TH98a]. polynomial [GKMSS90]. Polynuclear [EWHB94, FG92]. polypeptide [Mró95]. Polyploid [TLS94, TL96]. polyploidization [SK92]. Polyploidy [AVK94, Rod96, SLS93]. polyproteins [Gla91]. Polyribosome [Sin96]. Polysaccharide [FPB99]. polysaccharides [Com90]. Pomatia [BKK96, KK96]. Pool [WJH97]. Pools [ML97a]. Population [AP96c, AWA98, BB97a, BSM97, BA97b, CY98, CB97b, CPLA95, CCD^{+}98, Doe94, DSD94a, EL96, ELS97, EFB98, Fra97a, Fra99, GAD97, HS98a, HMUA93, JS97, KD99b, Kok96, KKH99, KS91, McN94a, NMI97, NM93a, Neu97, ONRH96, Rux95, áSá96, Tan96, TL98, WS94, WT94, Woo93, WH96, YZ95, AS95, Ber95a, BBB90, CK95a, Cra90a, DML95, DSD92, FTD90, GB95, Gro92, HA92, HS95, Has92, Hou92, Kel95a, KDG92, Llo95, McC92, RM95, SK95, Tan91, Van90, Vic91, WM95]. Population-level [NM93a]. population/genetics [BBB90]. Populations [AMO93, BTM93, Bje96, BKD99, CCK98, DA98, DVV99, DL97, DG93, ELS97, FWF97, Gat95, HMEW96, HLS94, JJ98, Koe94, LGM98, NBSM97, Pay97, PB97b, Rux96, Sch99a, SK99b, TC97, Tyv93, Tyv94, vBR98, Bar92, BN92b, CPLA95, Fin92a, Han92, JM95, Kel92, Lor90, Lun90, McN95, O’N91, Whi91]. Populus [Kra99]. pore [BE92]. Pores [KL98]. Porous [Com94b]. Portunid [Gri94a]. Portunus [Gri94a]. Position [CMG^{+}96, McC94, Wal99a]. Positional [Dru97, KW98b, MTM^{+}96]. Positive [KS98, MGP95]. Possibilities [Mul93b, WR93]. Possibility [CC97c, FSPC98, Jam91]. Possible [ACC96, AKV96, AKMV96, BOS98, BSN94a, BT99, DBB99, FS98, GVL93, GL95, HS97, HW98b, HE99, Kal97, KKS97, Li94, LRB94, MPSK94, Orp94, San99, Vin98, Won94, Zhu96, Ber92, Jac91, Man95, Mró95, Ss90c]. Possibly [SM97e]. Post [FCV99b, HC93, KS99, Orp90, WCW95, vdGvdBB92]. post-albers [WCW95]. post-canine [vdGvdBB92]. Post-ecdysial [HC93]. Post-natal [KS99]. post-synaptic [Orp90]. Postcopulatory [HW98b]. Postnatal [SLR^{+}98, Bla92, SK92]. Postural [MSN96]. Potassium [GWL94, Leb96, OHS95, Sta90]. potencies [Mac90a, Kat94a]. Potential [All93a, FPB99, Fis99, Gre98, Jam97, Kab94, Kur99, MSHN97, MKC^{+}98, Nov97, Say94, Tan99, Wil92b, vBW98, CJ90, Jac91, LBDDG90, LG90a, McI95, SL91a, Sun95a, Wie90]. Potentials [LRMS98, ML94b, RAGO99, Mar92a, TCE95]. Power [Koz96, SAB97, CSFC91, Gin95, Mat95, vL91]. practical [vdS90]. PRC [ASA98b]. Pre [HE94, FO92, Pen91, Sme90]. pre- [Sme90]. Pre-corneal [HE94]. pre-existing [FO92]. pre-mRNA [Pen91]. Prebiotic [Lah91, LL99, ML97b, ON98, Won90]. Precedence [NM93b]. precision [HH91b]. Preconceptual [Pol96]. Precursor [Sch97b, Gla91]. Precursors [Rak98, VP93a, Iid90]. Predation [MHN^{+}94, vdLLH95]. Predator [BS98b, BF99a, Doe97, FB97, Fis99, JE98, KK94c, MC93, MPAA97, PC97, SK96b, VRH93, AHL95, CA92, DML95, Dug90, Van91]. predator-avoidance [CA92]. predatory [KJ95]. predatory-prey [KJ95]. Predict [Emm97, Gou94, HDF^{+}99, MHW93]. Predicted [Bur99, SGP98, CW92]. Predicting [Fer94, MFB98, SA93]. Prediction [BDK^{+}97, CZ93, Com94a, Sch94, TCZ95, WAO^{+}99, WR94, vRGTV98, CPJL92, JP90b, KKS90, Sun95a, VGP95]. Predictions [Bür93a, Duk99, DSD94b, OVH^{+}96, RB97a, Doe95]. Predictive [MGL^{+}95, CSFC91]. Predictors [HVKS99]. Predicts [DBH94]. Preference [ASE94, Kat96]. Preferences [AR99, CY98, Kir92, Kun92, Kun95, MA95]. Preferential [GWS90, LJ97, WAJ91]. Preferred [Gaá93, JD99]. Preliminary [PG99]. Prenatal [CCK98]. Preparations [BIV97, GGML93]. prepattern [HKG90]. Prerequisites [Sjö97]. Presence [NBBH99, JB95, Kat90, Top90, Vid91]. Present [HS97]. Presentation [AL96, BOS98, SL91b]. Presenting [FP93, AL96]. Presents [RP99]. Press [TA99]. Pressing [KKS97]. Pressure [Bet99, HE93, KB93b, LD94, TWB^{+}93, KB91, Rei95, Sme90, SW95a]. pressure-volume [SW95a]. Prestress [SC99b]. Prevail [Lor94b]. prevalence [Koz91]. Prey [BS98b, CA92, Hut99, KK94c, MHW93, SK96b, SA93, VRH93, AHL95, DML95, Doe97, JE98, KJ95, MC93, MPAA97, PC97, Van91]. Prey-catching [CA92, Hut99]. Prey-Predator [BS98b, KK94c]. Price [Fra95a]. Primary [OP96, BN92a, OHS95, PNB92, Wil95]. Primates [Fed93]. Prime [YYY91]. primeval [KSH95]. priming [BLB95]. Primitive [Laz97, Di 95]. primitively [KSG90, TBD95]. Primordia [KCL^{+}96, SAF^{+}93]. primordial [Fin92b]. Principle [ES95, PDP97, Shc93a, Sil98, Szá99, VS97, VV96, Aok91, LB95, Wan90b]. Principles [BH94, SM97d, JR92, Ste90a, vdS90]. Prion [KS96a, PK98]. priori [Sai95]. prisoner [Gri95, BPHH90, HS98c, Koe97, Lor94a, MI98, Mow97, NS97, NS94, SNT95, WN99a, WN99b]. Probabilistic [ABK94, BB93, Maz91, MK92]. Probabilities [Nak99, SWA91]. Probability [rvdBvdG96, Bje94, CC97a, CC97b, CFC98a, JJ98, Kok96, MJLW96, MKGB98, Nak99, SG91, SNG92, AM95]. Probable [HC93]. Probands [Jam99b]. Probe [SQ98]. probes [AS92]. Problem [Blo94, FCV97, Gui94, Jon96, Koc94b, Mat96, Sch97b, Sve94, WP93, Whi98, WC94b, Cze92, DH92, HS91, TCB91]. Problems [MBSM93, NAB^{+}94, Wan90e]. Procedure [EL97, Yee91]. procedures [CB91b, TH90]. Process [DH98, DC96a, DC96b, DC96c, DGSK99, EKK98, GKM93, KN93, SFKD99, SLR97, VK99, BSB90, Bra90b, BHVC90, CH91, Gra90b, JS92, PST91, Tak95b]. Process-based [DH98]. Processes [BWK96, Bur99, Nak98, Nak99, PSLB99, Tor96, WP93, Whi98, AM92, Bar92, BN92a, Doe95, GS95b, Han95, HR91, HH95, LT91, MGP95, RR92, YC91]. Processing [AL96, Orp94, SL91b, THS^{+}95]. produce [SB90b]. Produces [KS98]. Producing [CG91, SNH99]. Product [Rak97]. Production [HHM98, MB98, Sak96, Sak97, WAO^{+}99, Aok90, BDK91, LL90, Mar95a, PAS91]. Productivity [CLMH^{+}96, SH93a, FAF90]. Products [DVV99]. Profile [Ben97, OSM95]. Profiles [DH98, FDT^{+}93, OBGGGRR93]. profit [Kuz95b]. Progenitor [MFHA^{+}95]. Progenitors [MSSG94]. Progeny [FM98, Cam95, MB92]. programmed [VGP95]. Programming [Bod97, Pol96, Wil93a, Aga95, Sib95]. Progress [vWP98]. Progression [Isr96, SS95]. Progressive [CI97, HMEW96]. Projection [Hon98, McI95]. Projections [PLB96]. prokaryote [RH92]. prokaryotes [ALM95]. Prokaryotic [CK97a, NAB^{+}94, NS91]. Proliferation [Ano99a, HU97, PMA97, vWP98, DP95a, Zhe91]. Proliferative [Tan99]. Prolonged [Vrb98]. Promiscuous [OWM96]. Promoter [WR94]. Promoters [RWK93]. Promotes [ZJ95, Koe90]. promoting [GPL98]. Promotion [FD92]. propagated [LG90a]. Propagation [Hor93, Koh92, RAGO99, WLH94, Coh91, EMRS90, JB95, Kee91, LF95, Mir90, Mur91, SB90b, WBS95]. propensity [Rux95]. Properties [AMO93, BJS97, BSS96, BU98, CSB98, Com94b, GP93, Hav99, HKH93, Joh93, KG93, LW94, OR97, PA98, RR96, Sch94, TS95, Zam99, CJ90, GD90, KL91, LF95, YYY91, dBP91]. Property [WH98b]. Propithecus [Fed93]. proportionality [MHTS90]. Proposal [FC93, Ros98, SK96a]. Proposed [Gli91, MC96, Rad91, Miz95, Und90]. Protandry [Zon96]. Protease [TCZ95]. Proteases [NT99]. Protect [NC97]. Protein [AOL98, AM96, AM97, BW98, BB99a, Bla94, BB97d, CL99a, CL99b, Di 94, Dwy98, EEG93, FHB^{+}97, FSPC98, FOO94, GRUH99, GP93, GRRF93, GOM94, Hei94, IET99, KK94e, MFM97, MBSB95, MS97b, ML97b, Nöl98, PGK94, PAG99, RJ96, Rak98, ROL99, Sto98, ST93, TS91, VGV93, WBS97, WH98b, WW97, YG94, ZZ94, ZZ99, ALM95, BSGB91, BSB95, Bro91, Cle91a, Eld91a, For95, Gor95, HKT91, Jär95, KKS90, KCM90, Kuz95a, Kyp90, Liq91, MGL^{+}95, OT90, PK92, SG91, Sib95, Sun95a, Tho95, sDA91, sDAB91]. protein-free [KCM90]. protein-induced [HKT91]. Protein-mediated [BB97d]. Protein-Protein [Dwy98]. Proteinase [Mel93a]. Proteins [BSN^{+}94b, BRV99, BWK96, CZ93, CL93, CC97c, Duf97, FST99, FSPC98, FO96, HRW93, LG99, Mah93, Mel93a, OWM96, PGK94, PG94a, RBH93, SVRF94, XCC94, AdG90, BSB95, Boy95, FO92, GW90, GMU95, Liq91, MP91, SL91b, TSC90]. prothrombin [WAJ91]. Proto [Dib91]. Proto-splice [Dib91]. protocols [WCPR91]. Proton [MFG99, TP96, Bro90]. protons [Bra90a]. Protoplast [KG93]. Prototype [DVV99]. Prototypical [CCFS96]. Protozoa [OPRN93, Und90]. Provided [Fed93]. Provides [ROL99]. Provisioners [DEK98]. Proximity [ACC96, Nov97]. PrP [Wil92b]. PrP-encoding [Wil92b]. Pruning [Jon97, vP97]. Psammomys [KD93]. pseudoadiabatic [Nöl95]. pseudoknots [Wil92b]. Psoralen [Str91]. Psychiatric [PMdM^{+}99]. Psychomimesis [Ros94a]. PTP [Cum99a]. Pulmonary [KB93b]. pulsatile [LG90b]. pulse [LWS97]. Pump [LR94a, LK91, LRJ92, WCW95]. Pumping [CK97b]. Pumps [Kab94]. Punctuated [MP93, Sea99, Bla95]. Pups [SA97]. Pure [YZ95]. purification [BHVC90]. purine [Nus91]. Purkinje [MKZ97, VW94a, WOQP93]. Pursuit [Wag93]. Putting [FL96]. pylori [KB95]. Pyramidal [Orp94]. pyrene [Mac90b]. pyrimidine [Nus91]. pyrimidine-purine [Nus91]. pyruvate [DC90].

Q [Bea90]. Q-type [Bea90]. QC [SM97b]. QC-PCR [SM97b]. QSAR [DBS97, tBázsáSH96]. quadrants [McI95]. Quadratic [DS90, ZZ99]. Quadriceps [MPSK94]. quadrupedal [SJK90]. Qualitative [Dav98a]. Qualitatively [Neu97]. qualities [Joh95]. Quality [FM98, GGML93, Wal96, WM98]. Quantal [BFG96, GT97, MKGB98]. Quantification [Alm99, AO96, FL95, Iid90, SNN^{+}96]. Quantifying [BSM90, FHB^{+}97]. Quantitative [Anc99, BB97d, DMP99, DM99, Fra99, FT93, Ham93, Hol91, JO96, KV93, Leb96, NT95, Nun98, RRB90, SM97b, SC99b, And95, ALW91, BS91b, FPPSK95, Her95, PST91, Sav91b]. Quantitatively [SR99]. Quantization [Ben99]. Quantum [PN98, LBDDG90, LWHZ91]. Quantumchemical [BMG98]. Quasi [FSPC98, SP92e]. Quasi-species [FSPC98]. quasi-steady-state [SP92e]. Quasispecies [SFGG^{+}99]. Quaternary [Ros94b]. queen [DML95, Rat90, RR92, Jud97, OJ99, SFGG^{+}99]. quenching [Fri91]. quest [Ing90]. question [vdHH90]. quiescence [Mur90]. Quota [ZvdBK97]. Quotas [Duf97].

Rabbit [CBFR98, HZ95]. rabies [MS92b]. Race [HS96]. Races [Här99b, TLS94]. Radial [MM97]. Radiate [Kaa94, Kaa95]. Radiation [Baj99, BB97b, CS94b, Coc98, Sek95, BS92, LOB92]. Radical [KK94e, Sta99, SS94a]. Radical-Nitrogen [Sta99]. Radio [SSCvdM99, Don90]. Radio-telemetry [SSCvdM99]. radio-tracking [Don90]. Rain [KKH97]. Rainbow [EWHB96]. rainforests [SM95]. rana [HC93]. Random [AH96, AH98a, EIU94, HH97, Isr96, Koe94, KRG95, KW98c, SS96a, Tyv93, AMO93, AM95, Jam90b, JV91, Kyp90, May90, SBM92, SL91c, Wil92a, ZM95]. Randomized [HL98b]. Randomly [Koe94, Moc99, AC90]. Range [Ezo98, NNM99, SAM97, Har99a, Man92b, SS90a, Vol90]. Rapid [MS98, OT90, Str95]. Rat [CBS^{+}99, KD93, SA97, WOQP93, Bla92, HCS91]. Rate [CB97a, GM99, Jon94, Kel96, Mal94a, MHW93, MP93, Rig94, Ste91b, WS94, WAO^{+}99, YYYU97, YG94, ZL97, Bra90a, BS91b, BH92, Coh91, FTCM90, FTL^{+}90, JLK92, Jør95, Med95, Spa92, TB91, WC95a, WT90, Wol91]. Rate-limiting [Ste91b]. Rate-maximizing [MHW93]. Rates [BSN^{+}94b, BTD98, DC99, Daw98, Gie94, JTC99, Mal97, Hoc91, JP90a, KDG92]. rather [LT91]. Ratio [CY98, DP94b, DG93, Jam97, Jam98a, RB97a, WG97a, WBC99, Jam90a, Jam92, Kra95a, MH90, Mil92a, Nor92a, WCW95]. Rational [Ng96, Zhu96, Ho90, Ng91]. Rationale [WE93]. rationalization [BS91a]. Ratios [Jam96, Jam99a, Jam99b, JE98, HT95]. rats [SM92, SG95]. Ray [Eas99, Mor96]. RC [BN92a]. Re [BC94, ES95, LW94, PH93, Whi97, ZH97, ZWH98, BHP92, CL92, HZ95, Lor90]. Re-assessing [Whi97]. Re-entrant [ZWH98, BHP92, CL92]. Re-entry [PH93, ZH97, HZ95]. Re-evaluation [BC94, LW94]. re-examination [Lor90]. Re-Examined [ES95]. Reactant [RP99]. Reaction [AH98b, Col93a, FT93, MCGPP97, NYI99, Rak97, SP94, SM97b, SM97c, VK99, YYYU97, Bar90, Bea90, Cze92, JP90a, Lac90, LH90, Nak95, SG95, Top90, VHGV92, WT90, Wan90c, ZT92]. Reaction-Diffusion [MCGPP97, FT93, Lac90, LH90]. Reaction-Diffusion-Advection [NYI99]. Reactions [BS98a, Bla94, Nak96, Pet96b, Pet99, Rak97, RP99, ZW96, AH90, HH91a, Maz91, Nak95, Sav91b]. reactivated [SH90b]. Reactive [WN99a, WN99b]. Reactors [NKR97]. Reading [Ker96]. Realistic [GH98b, KD94, Myl99, Kan92]. Reality [Sjö97]. Realization [SC99a]. Really [JBSdlG97, TAN94b]. Reappraisal [CPJL92]. Rearrangement [GS93, Gra93]. reassessment [Jam90b]. RecA [Mat93]. Receptor [AOL98, BW98, BFG96, BGS^{+}99, CLW93, CS96, FV94, Got94, HL98a, Hon98, JJ98, LR94c, LKS^{+}93, Mat96, NT97, TRVL96, WMLK96a, WMLK96b, WMLK96c, Zhu96, But92, FL95, HH91b, LG90b, QC91a]. Receptor-mediated [LR94c]. Receptor-Occupancy [WMLK96a, WMLK96b, WMLK96c]. Receptors [Chi96, Ker96, ROL99, Ros94b, SL98, TMZ99, WBS97, WRP99, WZN93, BLGG^{+}93, Jär95]. Recessives [Ano94b]. Recessivity [Van99]. Reciprocal [McC93, Sey91]. Reciprocity [NS98, PD92, Hem90]. Recirculation [Ste97]. Recognition [CS94a, FST99, KK93, Kuz96, PA98, Kan92, Man91, SC92b, TSC90, Wil95]. Recognize [RSW96]. recognized [SL91b]. Recombinant [WAO^{+}99]. Recombination [FK97, SK96a, TH98b, TF99, BN92a, BSOB92, Han95, MS90b, UJ95]. Reconciliation [LM94, Tre95]. reconciling [HP95]. reconsideration [NAZG95]. Reconsidered [KM96a, Sav95a]. Reconstructing [Bla98].

Reconstruction [HC93, Kor93, Yee91, Tak95b, YH90]. recordings [CK95b]. Recovery [BB99b]. Recruitment [BTD98, ND99, SFKD99, HP95]. Rectangularly [LL93]. Rectifier [LCR94]. rectilinear [SB90b]. Recurrence [ML94a]. Recurrent [Joh96, AMT91]. Red [CC95, Hai93, Jud97, NS96, OJ99, SFGG^{+}99, DML95, DM92b, JP90a, JP90b, KCM90, MWAS91]. Redox [All93a]. Reduce [MHN^{+}94, RR92]. Reduced [LRMS98, Sun95a, Rux95]. Reducing [SK99a, Tho96b, HS91]. Reduction [Sza93b, GH90, NS95]. Reductive [Kal97]. Redundant [Vin98]. Reef [SB94, Han91, SB92]. reference [FTBC92]. Refinement [JK98, NS96]. Reflection [Hor95, HP97b, ML97b]. Reflection-polarization [Hor95]. Reflections [Di 98, Eli94, Sve94, Yoc95]. Refractory [Hor93]. regain [LA90a]. Regarding [KS99, Stö95]. Regeneration [Col93a, DSM99, LAV^{+}97, MFHA^{+}95, Gos92, ISN91]. regia [EGDG^{+}98]. Regime [CCI94]. Regimes [DC96c]. Region [ASE94, KDBK93, OHS91]. Regional [LPP^{+}94, ML97a]. Regions [CMG^{+}96, McC94, ONT97, Sal96, AM90, Koz91, KCM90, Man92a, Man92b]. Register [DSD94b]. Regression [ER94, Mal94b, Mar93c, ML94a, GKMSS90, Gra92]. Regrowth [WOQP93]. Regular [TMI98, Han95]. Regularities [Shc93b, Shc94]. Regularity [SKI98]. Regulate [SW98, SS99]. regulated [Hum92, Kel92]. Regulation [AKL97, BOS98, BB99a, BB94, BN97, FP94, GH97, Gri94a, KB93a, KQ96, LC97, MPFHG96, MTLK96, Mye93, NU96, OWM96, PGF94, WSF99, WG97b, BLB95, BT90b, CGM^{+}98, FW90, HB91, Hof91, Hok90, KA91, Kuz95a, QC91a, QC91b, Rei95, Sav91b]. Regulatory [Fra99, MAB98, OKAP98, OWM96, SB97, SPS93, Tho91, WE98, KW91, KL91, MPO95, SYK91]. Rein [SKW98]. Reinforcing [BW91, SRH97]. reject [vdHH90]. Related [MBSM93, Boe90, Fin92b, Gla91]. Relating [Jam99a, BK90b]. Relation [HS97, LL96, LA96b, Mar92a, NU96, Nöl95, San99, Sek95, Hed92, Lib91, Mil91a, Mró95]. Relations [Hol93, Koc95, Rak98, Reg93, LWHZ91]. Relationship [BT99, CCVH98, For94b, For99, GL95, GP93, GW95, Mar93b, Mul93b, SGP98, SZK98, WE98, Won94, Jol90, KDG92, WKSH92, Wil95]. Relationships [HLG92, Bea90, MU92, Mil91b, RdM91]. Relative [GV99, KK93, MHN^{+}94, Mac90a]. Relatives [Jam99b]. Relaxation [GLO99, Gho93, SS94b, EMRS90, HR91]. relay [LF95]. Release [BFG96, CB97a, KPS97, LPS90, Sin96, SWA91]. Relevance [Mor99, Hor95]. Relevant [CBFR98, KDBK93]. Reliability [CK95b, Kol97]. Rely [CSSF93]. remarks [BN92a]. Remodeling [BM98]. removal [WPP90]. Renal [CI97]. Renewing [HMW95, Dib98]. Repair [AKL97, Pro94]. Repeat [MR98]. Repeat-Induced [MR98]. Repeated [DBAP^{+}98, FR93, Lei97, Lor94a, Mow97, Pos99]. Repeats [BKA97, KL98, MS90b]. Repertoire [DMP99, TD96, TH98b, Man91, Wei90a, WS90]. Repertoires [DP94a]. Repetition [SB90a]. Repetitive [Kat96, QF94, CJ90, OHS95]. Replaces [Koe94]. replicate [MB92]. Replicating [PB97a, FC93, MP98a, MP98b, MP98c]. Replication [Kel96, LL99, Nor93, Gor95, HB91, Kan92, Lah91, MP99, MS90b, SS95, TCLB90, Wol91]. Replicator [HS98b]. Replicators [GR97a, SS97]. Reply [CB91b, Eli96, Lib97, LÅ96a, LA96b, Sai95]. Repolarization [VR94b, LRS90]. repopulation [SH90a]. Representation [Jon97, CV91, MWAS91, Sun95a, YH90]. representations [Sav91a]. Representing [Con93, SAS92]. Reproducers [SS97]. reproducing [FO96, ON98]. Reproduction [Daw95, KK94b, McN94b, RBP99, TN96, Yam99, ZH98, AC90, FC92, KC91, Mag95, Rux95, Sak95, Van92a, Van92b, WBS95]. Reproductive [For99, GG97a, HW98b, LS94, ZJ97, Zha98, ZJ98, AC90, BK90b, Cor95, JM95, Mau92, WWC90]. Reptilian [JSB93, Lub97]. reputation [PD92]. Required [ACC96, Koz96, Gin95]. Requirement [All93a, FAF90]. Requirements [KD97, MA99]. Reservoirs [NBBH99]. Resetting [SS94b, LGBP92]. Resident [MFHA^{+}95]. Residual [CBFR98]. Residue [For99, MP91, Nöl98, Nöl99]. residue-residue [MP91]. residues [RHS^{+}95]. Resistance [ABK94, Bas98, BB99a, BB99b, Bow99, CK97b, Dam99, JIDH97, Kyp90, Lew98, NBSM97, Ves98, Wed94]. Resolution [Beu90, For94b, Nöl98, Nak91]. Resolving [Cze93]. Resonance [Dib98, Gre98, LPP^{+}94, BJZ^{+}91]. Resonant [BH94]. Resource [GH98b, HMW95, KD93, Kur99, LP93, Sak96, Sak97, WC94b, YS92, BBB91a, MBC95, Sak95, Wal91]. Resource-dependent [YS92]. resource-driven [MBC95]. resources [DM92b]. Respect [Mul93a]. Respiration [ODA97]. Respiratory [ASA98b, BBPD93, DBH94, FKJ^{+}94, FW97, ODA97, DRLB91, Eug95, LGBP92, Str95]. Respiring [Yar96]. Response [ASA98a, AKL97, Bet99, BR94, Cho93, Com94a, Dav98a, For94a, FBHK98, Gre98, HD97, Isr96, Kau98, Kei96, KWD^{+}94, Koc93, LPM95, LWS97, Mat96, MTLK96, RDHD98, SSCC93, Szá99, VK98, WOI99, BK90a, CSC90, Gie97, HCB96, Hur95, JH95, Kuz95a, LOB92, LÅ90b, Mac90a, MPRB91, MRB91, Mar95a, Ová91b, PO91, RBR90, SGG95, SC92b]. Responses [Boc98, Gro92, OP96, RBC98, RWN98, LOB92, NMS95, NSL91]. Responsible [KN93, RGK91]. Responsive [HS97, SR99, BE92]. Resting [Kab94]. Restricted [Sol95, SW95b, Wal96]. restriction [JV91]. Result [PH99, Sak92]. Resulting [BF99a, DC96a, GRRF93]. Results [BP97a, CS96, PG99, Irv90, MB92]. Resurrecting [WMLK96c]. Rethinking [CS94a]. reticulum [AH90]. Retina [Con93, TMI98, TMI99]. Retinotectal [Hon98]. RETRACTED [BBM^{+}97, BBM^{+}13c]. Retraction [BBM^{+}13c]. Retrovirus [AP96b]. Retroviruses [AP96b, Liq91]. Reveal [Ver94]. Revealed [DSD93]. revealing [Fer92]. reveals [BSB95]. Reversal [ML94b, BT90a]. Reversals [GAP95]. Reverse [OK94]. reversible [BW92, Kuz95a]. Reverting [Rou98]. Review [DLC98, RBH98, MS95a]. Revisited [Ham99, JM98, KB99, MHN^{+}94, QAM97, Rot98, May90]. revisited-but [May90]. Rhizopus [WLR96]. Rhodomelaceae [MWAS91].

RHP [KK98]. Rhythm [BBPD93, DCL96b, LWS97, RVMR99, Eug95, LGBP92]. Rhythmic [BD94, NYI99]. Rhythms [LWS97, ORG94, SS99, CLF95, Ing90, NLS91]. Ribonucleic [YY94]. ribose [Mat93]. Ribosomal [Fli99, ONT97]. Ribosome [DBB99, Woo97, ZGZ94]. Ribosomes [Kre94, Sin96]. ribosyl [Mat93]. ridiculous [Sti91]. Right [Laz97]. rigid [DB95]. RIP [MR98]. ripple [Mar92a]. Rise [PDP97, Di 95]. Risk [BBS^{+}98, BP98, MHN^{+}94]. Risks [FL96, JR92]. Risky [KB98a, RBP99, Kor91]. RNA [BSS96, Cha91, CF94a, DST96, Di 92, Di 94, Di 95, Eld91a, Fer92, Fer94, Fli99, FS98, GSS98, Gor95, HW98a, HKH93, KL98, Lah91, NYY96, SMA95, Sch97b, SFGG^{+}99, VGP95]. RNAs [BDA99, Jai97, KL98, Kun92, Mah93, ONT97]. Road [Blo94, Rol94]. Robust [GKM93, KY99, Sac94, ZL97]. Robustness [Pay97, RB97a]. Rod [KW94]. Rodent [BB93]. Role [ACC96, AKV96, AKMV96, BB97a, BOS98, CB96a, CS99, Col93b, Dem98, GVL93, Kal97, KG93, KW98a, Koc95, Li94, MS90a, MBSM93, ORM99, RWDM98, RS97, SC99b, Ste97, SA93, Vin98, VC98, BN92a, BN95, Car90, Fer92, Hoa90, Kor91, Man95, Mar95b, McI95, PNB92, Rei95, Ure91]. Roles [CH94, HC94, Mat93, Nuy94, NS91, SL91c]. Romanes [For99]. Root [Enn93, MTS97, VV96, ZBB97, But90, FB95]. Root-shoot [VV96]. Rooted [Ber95b]. Rooting [Ben97]. roots [Rob90a]. Rotating [PK99, PO91]. Rotation [Mul93b]. Rotational [BP99a]. Route [Kal97]. Routes [CTV93]. Roving [Whi94, Whi90]. RRE [Fin92b]. rRNAs [AIK99]. RuBP [LAV^{+}97]. Rugged [SJ94, TMII98]. Rule [BF99c, Kat94a, Kat94b, KV97, Lon94, Ost96, Sib94, WM98, Lac90, THY91]. Rule-based [KV97]. Rules [Dav98b, GCB96, Gie97, MHN^{+}94, SA97, TMI98, TS95, Web99, Whi90, BL91, CS91, PSE92]. Rumen [BDK^{+}97, BDDF97, FTL^{+}90, FTD90, FSD91, FSDT91]. Ruminant [FDT^{+}93, FHB^{+}97, FTBC92, VFS92]. Ruminants [FBS93, PGF94]. Runaway [BW91]. running [LWS97]. rupture [KCM90]. RX [CS96].

S [SSCC93, Whi91]. S-system [SSCC93]. Sabre [DM95]. Sabre-shaped [DM95]. Saccadic [LSG97]. Saccharomyces [LPRB96, vRGTV98, GW95]. saccharum [YLT98]. sacculus [Koc90]. Safety [Ale97b, MBS93, NVB99]. salinarium [Nab96]. salmon [Jam90b, Qui91]. Salmonid [BA97a]. Saltationist [Bla95]. Samaras [YA97]. Same [SS96b]. Sample [Koz96, MA99, WM98, Gin95]. samples [MS90a]. sampling [Fad92, FSDT91]. San [Phi99]. Sand [KD93, ME93]. SAR [SW96]. SAR-Associated [SW96]. sarcin [MGL^{+}95]. sarcina [JH95]. sarcomeres [vL91]. sarcoplasmic [AH90]. Satiety [TAA^{+}98]. Saturating [MKM96]. Save [NC97]. Say [HB96]. Scale [Phi99, SB94, TH98a, Wal96, BSB90, KS91, MA97b]. Scale-eating [TH98a]. Scale-invariant [Phi99]. scaled [CY90]. Scales [Dru93, HS98a, Mar93a, Ons92]. Scaling [BSS96, BU98, Enn93, FVB97, WW98, WBC99, Beu90]. Scatchard [Seg94]. Scathophaga [RWB98]. scattered [CM95]. Scattering [JW99, Mor96]. Schedule [SH93a, Fad92]. Scheie [Koc95]. schemas [CA92]. Schematic [MK92]. scheme [Ber91]. Schemes [FBS93, FSD91]. Schistosomiasis [YCM97, YC99]. Schizochroal [HC93]. School [Ano97-28, Ano99a, Niw98b]. Schooling [Niw94, Niw96]. Schools [Niw98a, HW92]. Sciarid [SP99]. Scleral [BM98]. Sclerosis [Kau98]. Scope [AM98]. Score [NNI98]. Score-dependent [NNI98]. Scotian [DK95]. scrambling [HG91]. Screening [FM98]. Script [KP93a]. Scroll [BP99a]. Scrotum [WN99c, Fre90]. scrounging [CG91]. Sea [Spi93, Sug98, Yos99, SL91b, SB90b, Ste91a]. Sea-anchor [Sug98]. seabird [Hou92]. seals [Boi95]. Search [CT95, KM95, Wal96, WM98, MS95b, SS90b]. Searching [KRG95, Ben92, DKL95]. Seasonal [SH93a, SK96b, Sak92]. Second [Ano97-28, BF99c, HH93, SSA94, BW92, CSFC91, Jam92]. Second-order [HH93, BW92, CSFC91]. Secondary [BDA99, DY99, FST99, GRRF93, HKH93, NYY96, Cle91a, Mró95, TSC90, VGP95, Woj92]. secret [Rob90b]. Secreted [YG94]. Secretion [GET96, MKGB98, BZGM91]. Seed [CH94, Ezo98, Hag93, LNF96, TN96, WS95, Zha98, RRB90]. Seedlings [EGDG^{+}98]. Seeking [Mar93c, Wil94, Hut99]. Seen [DO98, SB90b, Tho91, ZM95]. Segment [AF99, Kuz95b]. segmentation [HKG90, Pol92a]. Segmented [CL98]. Segments [CCM^{+}97, Sza92a]. Segregation [Sta96a, vBW98, Aga95, HLG92, Lib91, SG90]. Selected [SBRG99]. Selection [Abe90, AH94, BL95, Bjö94, Bür93a, DT98, DMP99, Fle97, Fra97b, Fra97c, GM98, GR97a, Han92, HL96a, HT96, Joh97, JW95, Jos97, Kei96, LKS98, Lei94, LNF96, MCG95, MDGB94, Møl97, Nak98, OJ99, OK94, PAG99, PB97a, Ric95, Sau94, Tan96, TF96a, Wed94, WE93, YZ95, Yi98, Zha98, vBR98, BDG92, Ber95a, Boi95, BR90, BM90, BW91, CS91, Cha91, CPJL92, FP95, Gra90b, HH91b, Kel92, Kir92, Man91, MGP95, NAZG95, Ng91, Pec92, Pri95, QG92, SP92c, SP92d, She91, Sza92a, Sza92b, YS92, vdGvdBB92, SS93a]. Selective [LRB94]. Selectively [NH96, Pet96b]. Selectivity [STI93, Vid91]. Self [Bai93, BB94, CS91, CGM^{+}98, Dib98, DC96a, DC96b, DC96c, ES95, FC93, FO96, GL93, Jon96, Kan92, KP93a, KM96a, KD98, MTM^{+}96, MP98a, MP98b, MP98c, MP99, Niw94, ORM99, ON98, SRH97, SA97, SM97e, TUHT97, TD96, TBD95, WC95b, WR93, Wil93b, CSJM90, For92, For94b, GWS90, HM95, KC91, Lah91, Mil92b, SL91b, SM95]. Self-assessment [ES95]. self-association [GWS90]. Self-Organization [Jon96, MTM^{+}96, CS91, GL93, KP93a, TBD95, Wil93b, HM95]. Self-Organized [KM96a, SA97, WC95b, KD98, ORM99, CSJM90, Mil92b, SM95]. Self-organizing [Niw94]. Self-regulation [CGM^{+}98]. Self-Reinforcing [SRH97]. Self-renewing [Dib98]. Self-replicating [FC93, MP98a, MP98b, MP98c]. Self-replication [Kan92, MP99, Lah91]. Self-reproducing [FO96, ON98]. self-reproduction [KC91].

Self-similarity [Bai93]. Self-sufficiency [SM97e]. Self-sustained [TUHT97]. Self-tolerance [WR93]. Self/Not [For94b, For92]. Self/Not-self [For94b, For92]. Selfish [BB97a, GR97a, MHN^{+}94, Smi98, PA90]. SEM [GMR98]. Semantics [Her91]. semelparous [Tak95a]. Semen [Cor98]. Semi [PD93]. Semi-permeable [PD93]. Semicircular [Mul94, Mul99]. Semiphenomenological [LVKH94]. senescence [Rob90a]. sense [FO92, Pau95]. sensing [YG94]. Sensitive [Cha98, Boe90, QC91b, RWDM98]. Sensitivity [BCQ92, FWF97, GRRB90, Mul94, PL93, TF94, Mur90, SCG91, Vid91]. Sensory [LSG97, MSN96, SPS93]. Separable [EG98]. Separate [All96, NK90]. Separated [AVM^{+}99]. September [Ano90v, Ano90-28, Ano91t, Ano91z, Ano92s, Ano92x, Ano93d, Ano93k, Ano94c, Ano94j, Ano95g, Ano95o, Ano96h, Ano96z, Ano97h, Ano97n, Ano98i, Ano98o, Ano99g, Ano99m]. Sequence [All93c, Bai93, BRV99, FOO94, FO96, LJ97, OWM96, YY94, FPPSK95, Gla91, Iid90, JV91, Wil95]. Sequences [Alm99, CT96, CL98, CMPR93, DY99, DLC96, FR93, GOM94, HWO90, HS97, KDBK93, LGSM97, LPBR96, OBGGGRR93, QF94, SH97b, Sch97a, SS96a, WZ96, WH98b, ZZ94, Zha97, AKW95, Boe90, CT95, CCRM90, DM92a, Fin92b, FPPSK95, GK90, Iid90, LT92, LB95, Miz95, RCCM92, SYK91, TET91, XCS^{+}95, ZS91, ZM95]. Sequencing [MO93]. Sequential [BCV97, BP92, BP94, vdLLH95]. Sequential-Arrivals [BCV97]. Sequestration [Nor93]. sericata [FWF97]. Series [DSD93, DSD94b, Rak97, DSD92]. Serine [NT99, LBDDG90]. Serum [Vol97]. Set [McC94]. Sets [Chi96, RCCM92, dLOG93]. Settling [CB97b]. Seven [Joh97, WBS97]. Sex [Bed96, CY98, DP94b, DG93, Gag94, IP99a, Jam96, Jam97, Jam98a, Jam99a, Jam99b, JBSdlG97, Jon97, JM98, Koe97, MH96, OJ99, RB97a, SP99, WG97a, Cha91, FR90, For95, GN95, Hai91, HT95, Jam90a, Jam92, Kra95a, MB92, Mil92a, Nor92a, Sza92b]. sex-determination [For95]. Sexes [All96, Myl99, Jam91]. Sexual [Atk91, BB97c, Gra90b, JIDH97, KK94b, Mag95, Mar91, Mar97b, MDGB94, Møl97, Pol96, RBP99, Tan96, Wed94, Yam99, ZJ97, Zha98, ZJ98, ZH98, CPJL92, Koe90, Rux95, Sak95, She91, Woj92]. sexually [LN95]. Shannon [HWO90]. Shape [BT99, CS94a, DBSV94, JO96, MBSM92, MM97, SFHP97, TMII98, Won94, dLOG93, DSP92, FC95, HLG92, MR91a, SV91, Wei90b]. shape-space [DSP92, SV91]. Shaped [HU97, TP94, DM95]. shapes [Kyp90, Mar92a]. Shaping [FS98, Fer92]. Sharing [EEG93]. sharks [Pau95]. Shear [SNK^{+}99]. shedding [Sak92]. Sheep [FWF97]. Sheet [Wie96, THY91]. Sheets [Cum96a]. Shelf [DK95]. Shell [KM96b]. Shift [Pos99]. shifted [ALM95]. Shifting [Bri93]. Shimagare [ISN91]. Shock [KL99, RVMR99, For94a]. Shoot [Sek95, VV96]. Shooting [AC96]. Short [CLY99, EFB98, Har99a, HK99, Kum91, FPPSK95]. Short- [Har99a]. Short-lived [CLY99]. Short-term [EFB98, Kum91]. Shortening [RV99, Tan99, Mor90]. Should [PA90, GS95a]. Show [BB99b, Rux95]. showiness [Atk91]. shows [Grü95]. shuttle [Bro90]. Sib [LNF96]. siblicide [GH90]. Sibling [ZJ98]. sibs [Jam91]. Sickle [Ben99]. Side [TPN98, MP91]. sided [PSE92]. Sidewinds [LHA94]. Sidney [JG92]. Sieve [PD93]. Sign [Sen96]. Signal [AC96, Cum99b, DO98, HL98a, LC97, LRB94, Mat96, RBC98, SC99a, SSA94, TP96, WLH94, For92, Iid90, Miz95]. Signal-induced [LC97]. Signaling [CS99, Ham98, Has97, NS99, VRH93, Kim95, Smi91, THS^{+}95]. Signalling [CMML96, Gui94, HP97a, Hur97a, HE98, JSR99, KW98b, LKS^{+}93, MM98, MFB98, PRG98, Bra90b, JG92]. Signals [BSN94a, DS99, Has94b, LSG97, SH95, BSGB91, Gra90a, Joh95, LG90b, NAZG95, Pen91]. significance [Bat91, Cle91b, Fel91a, Kel91, Ová91a, Ová91b, Sav91a]. Significant [PL93, Liq91]. Similar [HS97, ZJ95, Rad91]. Similarities [OWM96, PGK94]. Similarity [DST96, FOO94, FO96, HS97, Bai93]. Simmons [MC96]. Simple [AKV96, Bje94, CBSV95, Del95, DH98, DBS97, FV94, FPB99, GHW99, HV93, HK98, Joh96, Kra95b, MHN^{+}94, Mye93, Neu97, RWK93, SA97, SS94c, Sza94, Toq97, YQsZgZw97, dLA97, Ale91, Ber91, CS91, Cla95, CJ90, Fel91b, GKMSS90, Hes90, JG92, LGBP92, Llo95, RdM91, Tho95]. Simplicity [MHWM94]. Simplification [KD94]. Simplified [Wil99]. Simpson [Anc99]. Simpson-Baldwin [Anc99]. Simulated [BDDF97, CS94a, GWL94, GRUH99, Hem98, AH90, McK95, Sun95a]. Simulates [RR96]. Simulating [Osb93, RVMR99, FTBC92, SC92b]. Simulation [AB98, AMO93, BDK^{+}97, BSM97, BKA97, Big93, CCS96, CSSF93, DKL95, DC96c, Fad92, FWF97, FC93, FAF90, FT93, FK99, HA92, Hon98, Jon97, KKV95, LK91, LQL94, MFG99, MTS97, MAC99a, MAC99b, MK93, Mur90, NG99, OPRN93, PH93, Pen98, PN98, RP96, Rot95, SS93b, SS94c, SGSS98, Yos99, AM92, ALM95, BJZ^{+}91, CH91, Fia92, GKMSS90, HKG90, HW92, KH91, LQW92, LSM^{+}92, QC91a, QC91b, RSB92, SS95, SRW95]. Simulations [DBS94, Hie97, KV97, KM96b, LWS97, NSL91, PG99, SS94d, Aga95, BB93, Ing90, NZPB90, Ste90b, Sun95b]. simulator [FF90]. Simultaneous [CCG^{+}98]. Simultaneously [Bje94, Jon94, CK95b]. Singing [Smi91]. Single [But92, CK97b, DCL96a, FH99, LH94, Leb96, SH97a, SS96a, WH96, Yi98, YY94, BJZ^{+}91, CK95b, DF91a, FPPSK95, FSD91, HZ95, Kyp90, RR92, SM90a, Tre95, WBS95]. Single-beat [CK97b]. single-channel [CK95b]. single-dose [FSD91]. Single-file [FH99]. Single-Locus [Yi98]. Single-Peak [WH96]. single-queen [RR92]. Single-spike [Leb96]. single-stranded [FPPSK95]. Singular [CS94b]. Sinoatrial [DCL96a, DCL96b, PST91]. sinuosity [Wil92a]. Sinus [ASA98b, DBH94, DRLB91, DYL92]. Sir [JG92]. SIS [CDG^{+}98]. Sister [ABK94, FR93]. Sit [Nis99]. Sit-and-wait [Nis99]. Site [FLR93, RWB98, SK96a, Dib91, FK98, Han95, HMEW96, KV93]. Site-specific [SK96a, Han95]. Sites [KSF98, TCZ95, Bar90, Boe90, JV91, LBDDG90, Mur91]. Sitka [SEB99]. Situ [KV93, SSCC93, Bat91]. SIV [Gla91, RBB96]. Sixty [Shc94].

Sixty-four [Shc94]. Size [rvdBvdG96, Cou91, DSP96, DK95, Ezo98, GL95, HS98a, JE98, KD93, Koz96, KIF96, LGM98, LP93, Mel93b, Mul99, MA99, Niw98b, NT93, Sin96, SJ94, TC97, Vol97, WJ97, Zha98, dBP91, vdHdJ96, BN92b, BDK91, DM92b, For90, For91a, Gin95, GS95b, KA91, Kor91, KS91, KB92, TD92, WS90, Wys92]. size-based [GS95b]. Size-Distribution [Sin96]. Size-Structured [TC97, BN92b]. Sizes [SH93b]. Skeletal [Kem94, WB99, Smo95]. Skeleton [SWKS96, KCM90, Sax92]. Skew [Fra97d]. skewness [Wys92]. Skin [CBFR98, MJIB98, TMII98, SB90b]. slab [SC90]. Sliding [IET99, Smi90a, Smi90b]. Slight [YY94]. slightly [Kon95]. Slime [AKMS97, vOPH^{+}96, MH90]. slope [JD92, TB91]. slopes [Jol90]. Slow [AP96c, HK98, KH99, KN93, TMZ99, Cha90, Gui94]. Slowing [GLO98]. Slowly [Rou98, OHS95]. Slug [BVW99]. Slugs [MPH99, SH97a]. Small [CCK98, MRL96, MAC99a, MAC99b, TCE95]. Smallpox [DSD93, DDS96]. smooth [CKR95]. Snail [BKK96, KK96]. Snails [AC96, Leo92]. snakes [MM91]. Snowshoe [BSM97]. Soaring [Sug98]. Social [Bou94, CG91, DvH93, KP93a, MD96, NG99, RB97a, SFKD99, TB95, WNY98, BW91, Hem90, Koe90, Rat90, Tan91, WL91]. sociality [Kel95c]. Societies [SMG93, RR92, TBD95]. Society [KY99, MA97b]. sociology [MU92]. Sodium [JLK92, STI93, SH91b, Ste90b, WCW95]. Sodium-calcium [JLK92]. Soft [JW95]. Soil [Ben97, Com94a]. Solar [Sek95]. Soldier [Yam93]. solid [CH90a]. solitary [KSG90]. Soliton [CC97c]. Soliton/Exciton [CC97c]. Solution [Kat94b, Mat96, MGR97, Mor96, SM97a, Ste94, SS94d, UP96, WP93, Whi98, HS91, SL91a]. Solutions [HLT96, NAB^{+}94, TGS93, TRVL96, AM95, RB91, SH91b, TD92]. Solving [CCG^{+}98, DH92, FTBC92]. Soma [Sza94]. Somatic [BFG96, KP93b, LBWP93, PGBYP97, TH98b, Wei90a, Van92a, Van92b]. Some [BN92a, GD90, GE96, Nak96, Sjö97, TS94, Wan90b, WK90, Wan90c, Wan90d, WS93, Fel91b, JP90b, Res95]. somitic [Pol92a]. Song [Spe94, Wil93a, WS90]. Sorting [MIT96b, Gra93]. SOS [AKL97, Kuz95a]. Sosigonic [Joh97]. Soup [Laz97]. source [CS91]. Sources [PH93]. Southern [Ano99c, Oex98, McK95]. Space [BRV99, CS94a, DBSV94, FS98, Lei97, McC94, McN93, RS98, SFHP97, Ste90b, Arm90, BVvL96, DSP92, SV91, Wei90b]. Space- [Ste90b]. spacer [Nus92]. Spaces [Koc95]. spacing [Ryt90]. Spanish [JBSdlG97]. Spatial [BSM97, BSB90, CDG^{+}98, CV97, CH94, Cum99a, DO99, DEK98, DL96, GLO99, GK90, HC94, Hie97, KG90, KJ95, KD98, KIF96, KCL^{+}96, LGM98, LM96, MS97a, Mor92, Phi99, RMH96, SKI98, SRH97, SB94, SAF^{+}93, Vic91, WD97, WRP99, vOPH^{+}96, BH95, Kra95b, Llo95, MU92, MS92b, Nor92b, Ons92, SB95, Spu91]. spatialized [Gri95]. Spatially [BKD99, CH96, DL97, Her94, JS97, Neu98, TMZ99, Web99, Sak95, SBV92]. Spatio [BS98a, KMM^{+}97]. Spatio-Temporal [KMM^{+}97, BS98a]. Spatiotemporal [OS97]. Special [SP94]. Specialization [NC97, SW91]. Speciation [Che93, Koe95, Kor93, PM99, Var93b, Fia92, For95, Gre92, MWAS91, RD90, Sýk92]. Species [AFS95, Bjö94, Cam93, DL96, EM95, EL96, For96, For99, Fra94, Key97, KKN99, Kor93, Mar97a, ML97a, NNM99, Ng96, NBBH99, Nov97, NMdlR93, Ric95, TL96, WW97, Yi98, AHA92, BP97a, DH95, FSPC98, FS95b, GH98a, GR96, GR97c, HBP92, JFM90, Koe90, Man95, McK95, RR92, SBV92, Wie90]. Species-Area [DL96, NNM99]. Species-Dependent [For99]. Specific [BDDF97, IP99a, LNF96, MS98, Rig94, AIK99, AK94, FD92, Han95, Hav97, MR90, SK96a, SB97, SBRG99, SNG92]. Specification [McA91, GK90, KG90]. Specificity [Bri93, KO93, LA90a]. Specifying [Kel94]. Spectra [HS98a, BDK91, CC95, TD92]. spectral [DR91, Enr90]. Spectrum [AH96, DK95, GMR98, HMUA93, Bas98]. Speculation [Sch97b, Bat91]. Speculations [Mat93]. Speculative [Di 94]. Speed [LHA94, Wor95]. Sperm [BP96, BP97a, BP98, FCV99a, FCV99b, NM93b, RBH93, Sch94, Zon96]. Sphere [Won99]. Spheres [SEB^{+}93]. Spherical [TP94]. Spider [KV97, GV91]. Spiders [SFKD99]. Spike [Hor93, Leb96]. spikes [CLF95, Nak91]. spin [Won90]. spin-glass [Won90]. spindle [OHS95]. spinoletta [FAB92]. Spiral [DC96a, DC96b, WS98]. spite [Hur91]. splice [Dib91, Iid90]. splicing [Pen91]. Split [RB97a, Eld91b]. Splits [Kor93, TAA^{+}98]. Sponge [LR94a]. Spontaneous [CGB98, DC96b, GL95, Nab96]. Sporangiophores [EV94]. spore [MH90]. spores [Und90]. sporocarps [MR91b]. Sporulation [RVMR99, Kun95]. Spread [GVL93, HLS94, RA96, WS94, YY99, MS92b]. Spruce [SEB99]. squamous [SRW95]. St. [BEE97]. Stability [AHL95, BG98, CMB95, DLC96, EM95, EMS97, FBH91, GR98, Hut99, JE98, Kim93, KK94c, KK93, McN94a, MM98, NRD94, OK94, PG94a, PG94b, RMH96, RGAP97, Sch99a, SBV92, TW99, AMG95, Cra90a, Cre90, GLO99, Lor90, McN95, SW92, SV92, VS90, VR90]. Stabilization [AKMV96, Han91]. Stabilizes [KK94b, Koe90]. Stabilizing [ME93, Par97, WB99, YC99]. Stable [BP93, Daw98, Ezo98, GM99, GV98, Har99a, HT97, HMW95, LR94b, Lor94a, McN94a, MWC^{+}97, Mow97, ONT97, TS95, WNY98, WG97a, YZ94, YQsZgZw97, YZ97, ZJ97, Zha98, ZJ98, ZH98, dLOG93, BR90, Cra90b, Cre90, FC92, Lim90, MH90, MS95b, Now90, RG95]. Stackelberg [SH93b]. stacking [NZPB90, ZS91]. stacking/unstacking [NZPB90]. Stage [FC93, MMB99, Ujh96, Woo93, Lor90, MB97, NH95, SS90a]. stage-structured [Lor90]. Stages [CC96, Sjö97, BB91]. staircase [CLF95]. stalk [MH90]. Standard [Alm99, Bra90a, McA91]. standardized [Jol90]. Starfish [SB94, SB92]. Starvation [VK98]. Starved [YG94]. Stasis [MK97b, Par94, Sea99, Koe90]. State [BVvL96, Ber93, CPK96, Lei97, LBP^{+}97, MBM98, RBC98, San99, UP96, YQsZgZw97, YYYU97, AH90, ABR99, Bar90, FSD91, Gie94, GWS90, JLK92, KSG90, LA90a, MS90a, Maz91, Pet99, SC92a, SP92e, Smi90b, SM95, VFS92, VH90, WC95a, WAJ91]. State-dependent [UP96]. State-space [BVvL96]. stated [ÉT90]. States [Bon99, HK96, Kel94, MKC^{+}98, Nar98, HA90, HH91a, KJ91, Kel95a, Pat95, sDA91]. static [BE92]. Stationary [LL93, Mat96]. Statistical [AMO93, Dex95, GM94, Gho93, GM93, Gin95, GR96, GR97c, Koz96, LGSM97, NVB99, ORG94, TW97, NC92, Sun95a, Wan90b, WK90, Wan90c, Wan90d, WK90]. Statistics [LPBR96, Niw98b]. Status [Kim95]. Stay [Pos99]. Steady [ABR99, Bar90, CPK96, Gie94, Nar98, RBC98, San99, WC95a, YYYU97, FSD91, Maz91, Pat95, SC92a, SP92e, VFS92, VH90]. Steady-State [CPK96, YYYU97, ABR99, Gie94, WC95a, Maz91, SC92a, VFS92, VH90]. Steering [RS97]. Stein [ML94b]. Stem [Bje94, Bje96, For98, LBWP93, LBP^{+}97, ONT97]. Stem-loops [For98]. Stems [NVB99, SW92]. Step [KPS97, AH98b, Coh91]. Steps [Key97, Coh91, Cze92, Ste91b]. stercoraria [RWB98]. Stereochemical [RJ96]. Stereocilia [SSGR^{+}98]. Steric [CS94a]. Sterically [KL98]. sterile [Pam91]. Sterility [Ano94b, Kel95c]. Steroid [Ros94b]. Stewart [HC93]. stiff [FTBC92]. Stiffening [WOI99]. Stiffness [Zah96, SW92]. Stigmergic [KP93a]. Stimulated [PT92b, FK97]. Stimulation [ASA98a, PK93, LGBP92, Wic99]. Stimulus [LPM95, Trz99a, Trz99b]. Stochastic [Bar98, Bje94, BFK^{+}93, BJZ^{+}91, Bur99, Col93b, FWF97, HP97a, Hie97, LBWP93, LBP^{+}97, MS98, MDGB94, McN94b, MIT96b, MA97b, RV99, SHB90, SS93b, TMZ99, TL98, CG91, Hav91, Ko91, LT91, LG91, LT92, ROMM90, SBM92]. Stochasticity [KE96, Rod96, Gil92]. Stoichiometric [LJ95, SNH99]. Stoichiometry [Koo95]. stomach [KB95]. Stomatal [AVM^{+}99, LVKH94, YLT98]. Stomatocytosis [Won99]. Stop [ASE94]. Storage [BC94]. stores [Mir90]. Strain [Cha93, FLR93, KS98]. Strains [Bow99, TF99]. Strand [TUHT97, ZZ99, FD92]. strand-specific [FD92]. Stranded [YY94, FPPSK95, MO93]. strands [FO92, FD92]. Strategic [AR99]. Strategies [CCG^{+}98, ELS97, HW98a, HV93, MI98, McN93, MWC^{+}97, NS96, Pos99, WN99a, WN99b, WG97a, YZ95, ZJ97, Zha98, ZJ98, dLOG93, BPHH90, BR90, BHVC90, Cre90, KSG90, KM95, MS92a, YI95]. Strategy [BP93, GR98, HT97, HCB96, KD99a, Lor94a, McN93, Mow97, SvH99, TA99, WC94a, YZ94, YQsZgZw97, YZ97, Boy95, CW91, Cra90b, Fra90, Gie91, MBC95, MS95b, NH95, Now90, RG95]. Stratification [CB97b]. stratified [SH90a]. Strength [Ben97, CMB95, KE96, STI93, WR94, BC90, SW92]. streptococci [Boy95]. Streptomycetes [Koc94b]. Stress [CBFR98, Fer94, Isr96, Par94, Reg93, SNK^{+}99, TE96, Zah96, Man95]. Stress-Length [Reg93]. Stress-Modulated [TE96]. Stresses [Jon97]. stretched [Sax92]. Striatal [AW93]. striated [WPP90]. string [GMU95]. Strings [All93b]. Strong [Cre90, HS97, Nus91]. Structural [Bar94, EM95, Jen98, Mel93a, Nöl98, PGK94, SNN^{+}96, FAF90, Jol90, MGL^{+}95, SV92, vL91]. Structural-interactive [Jen98]. Structure [BDA99, BT97, CL98, Dav98b, DY99, DJZK96, DVV99, FLR93, FST99, FF90, Fra97a, GRRF93, Has94a, Hie97, HE98, HKH93, Kat96, KB94, Kle98, LKS98, LBW^{+}95, ME93, Ned97, RMH96, SZK98, TéS96, ZZ99, Cam95, Com90, Fer92, Gli91, HBP92, KPK92, Mil91b, OT90, PK92, RR92, SG91, SST91, SB90a, Sun95a, TD92, TSC90, VGP95, Wil95]. Structure-induced [FLR93]. Structured [BKD99, FT93, Har99a, JS97, KKH99, NMI97, áSá96, TC97, WT94, Woo93, vRGTV98, BN92b, Fin92a, HLS94, Lor90, Lun90, WM95]. Structures [CCFS96, CL99a, CL99b, Di 94, GC98, Koc95, NYY96, NGP97, Ros94b, TP93, Cle91a, KKS90, MS90b, Mró95, MP91, Rad91, Sib95, Spa92]. Studies [CL92, FV93, FCV97, FC93, LSB^{+}97, RWDM98, Rot95, SOR91]. Study [BMG98, BP97b, BN96, Ben99, CCM^{+}97, CT96, CL98, For98, GRUH99, GE96, Ham93, Hon98, JO96, KW98b, KKH97, LSG97, MPH99, ND99, NT97, ONT97, RP96, STI93, TE96, VKH^{+}98, WD97, WKSH92, WZ96, XCC94, AS92, HM95, HH91b, Jaf95, LRS90, MGL^{+}95, MU92, OHS91, OHS95, QC91a, QC91b, SC92a, SP92e, VR90, WBS95, ZS91, vdLLH95]. Studying [DST96, HP97b, Var98]. Subalpine [SKI98, ISN91]. Subcellular [DBS97, KV93, tBázsáSH96, Miz95]. Subcellular-site [KV93]. Subclass [GMR98]. subconductance [DF91a]. Subconscious [Khr98]. Subcoxal [Fra97d]. subcycles [SS95]. Subdivided [DA98, Tan91]. Subdivision [WS94]. Subject [Koe94]. Subjects [KWD^{+}94]. sublime [Sti91]. Subordinate [Cam93]. Subpopulations [EP94]. Subset [BGE94]. substances [Cor95]. Substituent [Nak96, Nak95]. Substitutable [Nar98]. Substitute [GV99]. substitution [PK92, ROMM90]. Substitutions [BMG98, SYK91]. Substrate [DFG95, GC91, SS96a, BSB95, Dem98, Sol95, WT90, ZT92]. Substrates [EEG93, KV93, NKR97, Nar98, ABW90, Bar90, VHGV92, Wan90a]. subsurface [CH90b]. Subthreshold [CS99]. subtilis [KMM^{+}97, Kun95]. Subtilisin [BMG98, LBDDG90]. subtransition [Sme90]. subunit [BS92]. subunits [LG91]. Success [Get93, WWC90]. successful [Hur91, Lim90]. succession [Dra90]. such [BS91a]. Sudden [WW94]. sufficiency [SM97e]. Sufficient [YZ94]. Sugar [YLT98]. Suggested [FA94]. Suggestion [OWM96]. Suggests [CS94a, KS98, Mal94b, ALW91]. suicide [Wan90a]. Sulfation [Com90]. Summer [Ano99a, Ano97-28]. sun [AP91a]. sunflower [RRB90]. Supercomputer [HKG90]. Superconduction [TP93]. Superinfection [MN94, MA98]. superior [Kac91]. Superoxide [Sta99]. Superparasitism [vdHH90]. Supply [Gra98]. Support [Kal98, LGSM97, FAF90]. Suppression [RL94]. Surface [BLGG^{+}93, Bla94, CMB95, Gra93, GS93, Kal97, KSF98, RBH93, SEB^{+}93, Spa92, Yos99, Boy95, Hoa90, May90, Mir90]. Surfaces [HMEW96, PA98, CH90a, Hor95, SM90a]. Surfactant [PGK94]. surveillance [CSS91]. survey [vdS90]. Survival [Eas99, Isr96, LL99, MD96, WJ97]. susceptibility [Lun90]. Suspended [MNH97]. Sustained [Adl93, TUHT97]. Swarming [Vis93]. Swarms [Mye93, TB95, Mil92b]. Swelling [HE93]. Swimmers [BC94]. Swimming [CB94a, DFG95, HH97, KS98]. swine [Aok92]. Switch [BB96, CG96, Joh93, Mit96a, Eug95, RV90]. Switch-like [Joh93]. Switching [NU96, PTBL98, Sas94]. Switchlike [OKAP98]. Switiching [Nab96].

Syllables [Wil93a]. Symbiosis [Fra95b, PV95]. Symmetric [Jos97]. Symmetrical [Shc93a, Zha97]. Symmetry [KL97, SNN^{+}96, Spi93]. Sympathetic [ASA98a]. Sympatric [Fia92]. Symplocarpus [BJ96]. Synapses [TMZ99]. Synaptic [BS98c, CB96a, CB97a, LLS94, MKZ97, BP92, FTCM90, GMU95, Orp90]. Synaptotagmin [CB96a]. Synchronic [RJ96]. Synchronicity [SB94]. Synchronisation [Rux96]. synchronizable [Nev90]. Synchronization [EL97]. Synchronize [ISSI97]. Synchrony [HS98c, SM92]. Syncytial [RBC98, VpM95]. Syndrome [BKA97, Mar96, WW94, Tri92]. Synecological [GE96]. Syneretic [Bet99]. Synergetic [TP96, SJK90]. Synergetics [Hak94]. Synergies [McC93]. Synergistic [Fra95b, KV93]. Synergy [BvHS93]. Synonym [Duf97]. Synonymous [Kun92, Kun95]. syntactic [CV91]. syntax [CV91]. Synthase [CG96, Sal96, Kay92]. syntheses [FPPSK95]. Synthesis [Di 94, Fli99, FHB^{+}97, Kem94, MFM97, Sal96, Kaa95, VP93a]. Synthetic [FR99]. System [AW94, BvHS93, CPK96, Dam99, Dum94, FV93, For99, FT93, FK99, FO96, GBD99, Gre97, Her90, JSR99, Kal98, KWB93, Khr98, LKS98, Mul93a, Mul94, MdR98, OWM96, ON98, PC97, PGBYP97, RH96, RRG98, RSW96, SRH97, SH94, Sta98, Sta99, TUHT97, UP96, Var93a, VB93, WW94, WR93, AS92, BDK91, DSP92, Ish90, Kat95, Lac90, MWAS91, NLS91, NSL91, SC92b, Smi90c, SSSF95, SSCC93, VK91, Wei90b]. System-dependent [LKS98]. System-level [Her90, BDK91]. Systematic [MA99]. Systematization [Shc93a]. Systems [Ale97b, BG96, BSW97, BW98, BSN94a, BB97d, CK99, CPK96, CK94, Chi96, Com94b, CLY99, DvH93, Eas96, FK98, Fra97b, Gaá93, Gav97, GH98a, HW94, Hir93, Jab94, KKN98, KK94c, KQ96, LR94b, MS98, MC93, MPAA97, Mog93, Mul99, Nak99, Ned97, NS96, NSW97, OR97, SK96a, Sch96, SQ98, SNH99, Sza93b, SAH96, TéS96, VF98, VC98, WJH97, Yam93, ZJ94, BSM90, BSB95, Bur90, CSFC91, DM91, FBH91, GM90, GAP95, GWS90, HH91b, Irv90, Kan92, LVL92, Mau92, Mur92, PVT91, Sav91a, SV92, Tan91, Tch91, VHG^{+}90, Vid91, WAJ91, YC91]. Systolic [CK97b].

T [CK96, DP94a, EP94, RDHD98, Ste97, AL96, BGE94, CCFS96, CCS96, DP95a, DMP99, FP93, Kau98, Man91, MR98, MK90, McL94, MSSG94, MPFHG96, OP96, PMA97, Wic99, WLJN99]. T-cell [EP94, PMA97]. T-tube [CK96]. T-zone [RDHD98]. T4 [Kun92]. tactic [RG95]. Tail [BC94, Tho96a, Tho96b]. Tails [Tho96b]. Take [LL94]. Tandem [Gra95a]. Tapered [CK96]. Target [DP98, Kuz96, RWN98]. targeting [Miz95]. Targets [Hut99]. taste [Enr90]. TATA [Nus92]. Taxa [Hol96]. Taxon [AIK99]. Taxon-specific [AIK99]. taxonomic [Bur90]. taxonomy [Ho90]. Taylor [Rak97]. TCR [BOS98]. TCR-Presentation [BOS98]. Tear [HE94]. Technique [MWC^{+}97, WGH98]. Techniques [GE96, NYY96, CCRM90, RCCM92]. tectonics [Nor92b]. Teeth [KCL^{+}96, SAF^{+}93, vdGvdBB92]. telemetry [SSCvdM99]. Telomere [Oex98, RV99, Tan99, AKW95]. Temperate [Mit96a]. Temperature [GM94, Hok90, MvB93, Mye93, PN98, RR96, vdHdJ96, HLG92, McA91, Nun90, Sun95b, Tor90]. Temperature-Compensated [RR96]. Temperatures [RLF93, LIWE90]. templates [Com90, GMU95]. Temporal [Eas96, HE98, KMM^{+}97, MFB98, Ons92, SSCvdM99, BS98a]. temporo [Spu91]. temporo-spatial [Spu91]. Tendency [WL91]. Tendency-distance [WL91]. tendon [De 95]. Tensegrity [SFW^{+}96, WOI99]. Tension [SM99, BH92, Hoa90, Tri92]. tentative [RB91]. Term [NM93b, SS99, EFB98, Hol99, Kum91]. Terminal [ASE94, Kum91, MS90b]. Terminator [MO93]. termini [Gla91]. Terminology [SH95]. terminus [HKT91]. Termite [ORM99, CH91]. Termites [MD96]. Termoli [Ano99a]. Terms [CK97b, Hak94, Sch96, WR93]. Ternary [CS96, Cze93, Trz99b, WMLK96a, WMLK96b, WMLK96c]. Territorial [Bla97, PH90, Wie90]. territoriality [BBB91a]. Territory [BCV97, WLM96, BBB91b]. Tertiary [Fer94, PK92, Sun95a]. Test [Dum94, Hem90]. testability [THF92]. Testicular [WN99c]. Testing [AVK94, Bro97, Nuy94, SS90a, TW99, Bea90, Jea92, WCPR91]. Tests [FLR93, GR96, GR97c, Koz96, Gin95, GS95b]. Tetrahymena [BT90a]. tetraplex [UJ95]. Tetrapods [FVB97, TW99]. Text [PB97a]. TGTTCT [MK93]. Th [ALW91]. Th1 [FP94]. Th1/Th2 [FP94]. Th2 [FP94]. thalassemia [Iid90]. thaliana [MAB98]. Their [For96, Gaá93, HS97, JBSdlG97, KG93, KV93, MSHN97, Rak98, Sak94, SB97, WZ96, ABW90, Alm99, BMS93, CZ93, GW95, Hor95, Jon97, JM98, Kau98, KSF98, Koc95, Kyp90, Mar93b, MA95, Sav91a, Ste91b, TD92]. them [NAZG95]. Theme [LS94]. Theor [BBM^{+}13a, BBM^{+}13b]. Theorem [GV99, ÉT90, Han92]. theorems [Leó90a]. Theoretic [DT98, Mot93, WZN97, Pat95, VRH93]. Theoretical [Adl93, BLB95, Blo94, Boi95, CG96, Cha98, CF94b, Coc98, CH90b, DBH94, DST96, DP95b, FK99, HU97, KW98b, KB92, Mir90, Mró95, NT97, NT99, ONT97, ON98, PN98, PTBL98, RZF^{+}99, RV90, Ryt90, SM97b, SM97c, Sug98, TE96, TF99, TWB^{+}93, Vis93, Wan91, WW97, ZCS^{+}90, Bak95, GW95, Jaf95, KL91, SC92a, Top90, Wan90e, Mar91]. Theories [KD97, PGF94]. Theory [Ale97b, All96, BDA99, Bje94, BJZ^{+}91, CPK96, CS94b, CV97, DS99, DMP99, EFB98, FLR93, Gie94, Ham99, Has97, Joh97, Kat94b, KD98, KK94e, Mat96, ML94a, MA97a, Mit96a, NS96, Ost96, Ric95, Rol94, RBD97, Sch91a, Sch91b, SS94a, THF92, Tyv93, WW98, WR93, Zha97, Zhu96, AMW95, BH92, CSFC91, CS90, Cus95, DC90, HCM^{+}92, Irv90, JG90, JR92, Kac91, KW91, PO91, Rob91, Sav91a, SJK90, Smi90a, Smi90b, Ste91a, Whi91, Wil95, Zhu91]. Therapeutic [MSHN97]. Therapies [WZN97]. Therapy [WDP98, vWP98, Jac91]. There [CCI94, Eli94, JO96, MO96, Ujh96, Hol99, Ing90, Yoc95]. Thermal [McA91]. Thermodynamic [Che94, Com94b, Eli94, Got94, HBJ^{+}97, HE93, JJ98, SAB97, San99, TW97, Yoc95, Leó90a, Zhu91]. Thermodynamics [AH90, DD98, Mar93a, WW98]. Thermoregulation [WC95b, WHL93, LL90]. Thermotaxis [MPH99]. thiamin [Bro90]. Thinking [DGSK99, For98]. Third [LRB94, PSLD96, Kuz95b]. thorns [SB92, SB94]. those [VHG^{+}90]. though [Bat91]. Thoughts [GE96]. threat [AMG95]. Threats [Fis99]. Three [AVM^{+}99, All93b, BW98, BP99a, Bon99, BB99b, EM95, FST99, GOM94, KD99a, LW94, MH96, PH93, UP96, FL95, Kaa95, MI98]. Three-compartment [LW94]. Three-dimensional [AVM^{+}99, BP99a, FST99]. three-person [MI98]. Three-Species [EM95]. Three-state [UP96]. Threshold [DM92a, HLS94, YG94, GW90, PVT91, Tch91, Vid91]. Thresholds [BRT98, BS93, FK98, GHW99]. Through-bond [AdG90]. Throwing [CC99, Ale91]. thylakoid [NZPB90]. thymic [Man91]. Thymine [MFG99]. Thymus [MSSG94, MPFHG96, Car90, MGP95]. thymus-pineal [Car90]. Thymuses [MFHA^{+}95]. Tibia [Mul93b]. Tibial [Mul93a]. Tick [NBBH99, AS95, SAS92]. Tick-borne [NBBH99]. tiger [MR91a]. Tilted [PSLB99]. TIM [LG99]. Time [AMT91, BB97b, DC96a, Dru93, DSD92, DSD93, DSD94b, Emm97, GLO98, HA97, Jam96, Jam99a, KL96, Mar93a, Mor98, Nis99, Nun98, PM96, Rak97, Rod98, SM97a, SAH96, TRVL96, WH97, YQsZgZw97, YZ97, AHA92, BP92, CH90a, Cle91b, HA92, HM95, Med95, Nak91, Ste90b, VpM95, Val92]. time-delay [HA92]. Time-Dependent [SAH96, TRVL96, SM97a, Ste90b]. time-lapse [VpM95]. Time-minimising [WH97]. Time-minimizing [KL96]. Time-Periodic [DC96a]. time-resolution [Nak91]. Time-to-extinction [GLO98]. times [GW95, HS91, Kon95]. Timing [CC99, McN94b, Ale91, AC90, IL95]. tip [CS90]. Tissue [BIV97, DSM99, HE93, JW99, Kee96, OSM95, PK93, PDP97, PK97, RBC98, SGP98, Tab98, ZH97, ZWH98, Com90, KWB93]. Tissues [Dib98, SC96, CL92]. toad [CA92]. Tolerance [BB99b, CBSV95, LCP^{+}98, PG99, SSA94, WZN93, WR93]. Tolerance-Show [BB99b]. tool [Kum91, LPS90]. Tooth [Osb93, GKMSS90]. Top [Bra96, KQ96, Her95]. Top-Down [KQ96, Her95]. Topographic [Hon98]. Topological [DBS94, GC98, Kra99, Ber95b, Han95]. Topology [Ber95b, DY99, vP97]. Torsional [SZK98]. Total [Bro91, KIF96]. Toxicity [EWHB94, EWHB96, EPT97, MTLK96, ECB91, EB92, FG92]. Toxins [JD99]. Toy [Sza94]. Tracer [GRUH99, LPP^{+}94, Wei97]. Tracer-Kinetic [LPP^{+}94]. tracheal [SSSF95]. Track [MNB97]. tracking [Don90]. tracks [MB90]. Tract [ABH94, BvLVR96, BVvL96, HW98b, VFS92]. Traction [Edm96]. tracts [Nus91]. Trade [Daw96, HW98a, SR99]. Trade-off [Daw96]. Trade-offs [HW98a]. tradeoff [Tay91]. traditional [Fad92, Mor91]. traffic [MH95]. Trail [CSSF93, WEK95]. Trails [ALS93, BDG92]. Trait [KKN99]. Trait-for-Trait [KKN99]. Traits [AVK94, DT98, But90, Cor95, SHB90]. Trajectories [HH97, Hut99]. trans [BN97]. Transcribed [Dru97]. transcript [Fer92]. Transcription [BSN94a, BF99c, Boe90, Kel94, SW98, BSGB91, Kel95b, LWHZ91]. Transcriptional [OWM96, Koz91]. Transduction [HL98a, Ker96, RBC98, SC99a, TP96, Mar90]. Transfer [AP96b, BDA99, CL93, CC97c, DVJ^{+}99, KG93, MFG99, MS98, TP93, AdG90, Di 92, Kun92, PT92a, PT92b, TCLB90]. Transferase [DBB99, Mat93]. Transferrin [BGS^{+}99]. Transform [FS94, SC90]. Transformation [KB99, LSG97, TP94]. Transformations [Won99]. Transient [DC96c, KK94d, LL93, Pet99, SSCC93, AH90, HS91, Str91, VHG^{+}90, VH90]. transient-phase [VH90]. Transient-state [Pet99, AH90]. Transients [CGB98, Leb96]. Transition [CC97a, CC97b, EGDG^{+}98, FC93, KH98, MR98, TN96, CK95a, Vol90]. Transitions [HK98, KH99, SS97, LOLM91, O’N91, SJK90, Sme90]. transitory [Cár91]. Translation [BSN94a, Gor95, Lah91, RH92, SLR97]. Translational [GRRF93]. Translocation [FH99, HBJ^{+}97, Sin96]. Transmembrane [BD94, RHS^{+}95]. Transmission [BP99b, LLS94, MKZ97, Tak98, Wat97, YCM97, YC99, Fin92a, Hoc91, Now91, RBH93]. transmits [SAS92]. transmitted [LN95]. Transmitter [BFG96, TMZ99, BP92, SWA91]. Transparency [JW99]. Transport [Ben99, BB97d, CFC98b, DFG95, HSM94, HRW93, JC96, KWB93, KW98b, KSM97, LVKH94, MNH97, Mel93a, MBSB95, MS97b, Spa92, TP96, VGV93, AMW95, AH90, BSB95, LRS90, Rad91, SMA95, WC95a]. Transporter [AOU98, QC91b]. Transporters [TMZ99, Wil95]. Transposable [BB97a, Dwy98]. Transposition [Jam99b]. Transposon [MK97b]. Transposon-mediated [MK97b]. Trap [Wil94]. travel [Val92]. travelling [Vic91]. Treat [Ros98, SvH99]. Treatment [Gat95, KP93b, MBSB95, NBSM97, RWB98, BSB95, Top90]. treatments [Str95]. Tree [BU98, CCI94, ONRH96, SNN^{+}96, SNK^{+}99, vP97, Ber91, CY90, DS90, YH90]. Trees [Ber95b, DM95, FK99, KW98c, LTZ96, Lon94, MBS93, PSLD96, SNN^{+}97, Wes93, Wil98, Zam99, But90, MS92a]. Trend [Kol97]. Trends [AS95, KS96b, NS91]. Triad [BMG98]. Triads [NT99]. Trial [WC94a, CW91]. Tricarboxylic [SS96b]. Trigger [NT93, PA98]. Triggered [GWL94]. Trilobite [HC93]. Trim [Tho96b]. Triosephosphate [OVH^{+}96]. triple [Kan92]. Triplet [BKA97, DLC96]. Triplets [AIK99, DBAP^{+}98, Shc94]. Triplex [LSB^{+}97]. Trisomy [Mar96]. trisphosphate [LC97]. trisulca [Mal94a]. tRNA [Di 95, Di 99, Kat90, KK93, PG94b]. Trochilidae [LD94]. Trophic [HBP92, MO96, WP93, Whi98, Her95]. Tropism [RBH93]. Trout [EWHB96]. Trunk [ALS93]. Truth [PGL94]. trypsin [VHV^{+}90]. trypsinogen [VHV^{+}90]. tryptophan [Kay92]. Tsao [TW97]. Tsembaga [And98]. Tseng [CB91b]. Tsetse [Wil94]. Tsou [Wan91]. tube [CK96]. Tubes [ABR99, CFC98b, PD93, ABR97, BL91]. Tubular [Koc95, SW92]. tubulins [BSGB91]. Tumor [Gat95, ML94a, Rig94, TMII98, WSF99, CSS91, SC90]. Tumor-Host [Gat95]. Tumorigenesis [Ned97]. Tumour [Isr96, OS97, SS93b]. tuning [For95]. tunneling [Ber92]. Turbidostat [Fle97]. Turbidostat-like [Fle97]. Turbulent [MNH97]. turn [TCB91]. Turning [CFC98a]. Turnover [GRUH99, McK95, VFS92]. turns [BDG92]. Twenty [Shc93b]. Two [AFS95, AW93, CB94b, Daw95, EEG93, Fle97, For92, For99, GH98a, GR97a, KB98b, KD99a, LM94, LL93, Mal94a, MKC^{+}98, Nab96, NRD94, NMdlR93, PH93, PK97, PGL94, áSá96, SEB^{+}93, SS96b, SH94, SSA94, TH98a, Wan90e, WC94a, YZ94, YZ95, YQsZgZw97, Yi98, ZJ94, ZBB97, CW91, DSP92, FTL^{+}90, FW90, Gre92, MBC95, NC92, PSE92, RP99, Rau94, Sai95, Smi91, SBV92, SS90a, TM92a, Van91, sDA91]. Two-Allele [Yi98, ZJ94]. two-channel [FW90]. Two-compartment [LL93, FTL^{+}90]. Two-component [SEB^{+}93]. two-dimensional [DSP92]. Two-fold [Daw95, TM92a]. Two-generator [KB98b]. Two-level [GR97a, NRD94]. two-locus [Gre92]. Two-Phenotype [YZ94, YQsZgZw97, Yi98]. two-sided [PSE92]. Two-Species [Yi98, GH98a, SBV92]. two-stage [SS90a]. two-strategy [CW91]. two-trial [CW91]. Type [AH96, EL96, RDHD98, SJ94, AH98a, Bea90, Joh93, Mil92b, NH95, SGG95, TGS93]. Types [CZ93, Fle97, MdR98, RJ96, Mró95, Sib95]. Typogenetics [Var93a]. tyrosine [SP90]. Tyrosyl [Sta99].

U [BDG92]. U-turns [BDG92]. Ubiquinone [FW97]. Ultimate [Thi93]. Ultrabithorax [SB97]. ultrasonic [Nöl95]. Ultrastructure [JW99]. Ultraviolet [AKL97, Coc98]. Unbranched [HK96]. Uncertain [ELS97, Neu97, Nis99, Man90, Nis92]. Uncertainty [iT94]. Unchanging [KK94b]. uncirostris [Kat96]. Unclassified [AFS95]. uncoupled [SS95]. Undecidability [KM96b]. undergoing [O’N91]. Underlie [HP97a]. Underlying [DCL96a, Orp94]. Understanding [De 95, WMLK96b, YC91]. Unequal [ABK94, FR93, KA91, TF93]. uneven [Mil92a]. Unexpected [RLF93]. unexplored [GWS90]. Unfittest [LL99]. Unfolding [WP93, Whi98, HBP92]. unhandicapped [Gra90b]. Uniarticular [dLA97]. Unicellular [Isr96, PS93, PP92]. Unification [Gie97]. Unified [Cho93, MA98, CSC90, Her95, Man91]. uniform [SWA91]. uniformity [CK96]. Uniformly [RAGO99]. Unifying [ARC97, DT98, ML94a, WW94, FSDT91]. Unique [Kau98]. Uniqueness [Pat95, Gra92]. Unit [vBR98]. Unitary [Dwy98]. Units [BF99b, CV93, GM98, Gie97, GE96, OES93, PB97a, RSW96, CV91, LH90, LH91]. Universal [Koc94a, Ver94]. universality [CLF95]. Unlabeled [Seg94]. Unliganded [MBSB95]. unloaded [BSB95]. Unpinning [PK99]. Unpredictable [Sak96, Sak97]. unselective [GS95a]. Unstable [Møl97, Top90]. unstacking [NZPB90]. Unsteady [LT98]. unstirred [LSM^{+}92]. Untransformed [Ros94b]. Unusual [CMPR93, Hai93]. update [For92, Jam90a, Jam92]. upon [LS94, PA98]. upstream [Boe90]. Uptake [FHR^{+}99, FBL^{+}95, HCS91, Mar95a]. Urchin [Spi93, Yos99]. urchins [SB90b]. urea [LBN95]. urease [LBN95]. Urinary [BvLVR96, BVvL96]. urine [Beu90]. Usage [ASE94, BSN^{+}94b, GRRF93, LL96, ZGZ94, BL95]. Use [BP99b, DBS97, FTD90, HDF^{+}99, HW98b, Jon96, Kok96, KDBK93, LT98, MHW93, Mor98, Rak97, Vol97, WM98, WBC99, tBázsáSH96, AHA92, Han95, Kun92, PH90, Rad91, SG95]. Use-What [MHW93]. Used [All93c, HP97b, Hem90]. Useful [Mil91a]. Usefulness [Ish90]. Using [BB93, CLY99, GRUH99, KKH97, LCR94, MCG95, Mar97b, NYY96, NS96, SK96a, SS99, TW97, VV96, WT94, YYYU97, Aga95, AFS95, AP91a, Bea90, CSFC91, CCRM90, DSM99, FAB92, FSDT91, HH91b, Joh95, LH90, LL90, LT91, SM90a, SP92a, SP92b, Sib95, WAO^{+}99, WCPR91, WW97]. usurpation [KSG90]. Utilization [KD97].

V [BP93, WBC99, ZJ98]. Vaccination [DDS96, RBBG98, PBJ^{+}92]. Vaccine [RBB96]. Vagal [ASA98b, ASA98a, DBH94, DCL96b, DLC98]. Vagal-Cardiac [DLC98]. vagally [DYL92]. Validation [KGFB97, MAC99b]. Validity [HWO90, Szá99, Mor91]. Value [PL93, BK90b, JR92]. Values [Ber93, HL96b, Nöl98, Nöl99]. Valves [MCGPP97]. vanadate [SH90b]. vanishing [AHA92]. Variability [For99, HW98a, Koz96, Now92, SNN^{+}97, SS99, BHP92, Gin95, Gro92, Smi92]. Variable [Hol96, KDBK93, RBP99, Sac94, Tho96a, WT94, For91a, Gro92, OHS91, Sta92]. variable-yield [Gro92]. Variables [Adl93, Sch96, TF94]. Variance [NS97, Tar94, CC95]. variant [Sav91a]. Variation [AF99, Boy95, Doe97, Fra97b, Hem98, Jam97, Jam98a, Mah93, MP93, Tan99, Ber91, Cam95, Hol91]. Variations [Adl93, HK99, LS94, SBRG99]. Varicosities [CGB98]. Variegation [CMG^{+}96]. Various [CC96, HK99, ODA97]. varum [MPSK94]. Varying [BN96, DY99, Doi97, GAD97, Koe94, AC90, Sak95, SH91b]. Vascular [KWB93, Kra99, SNN^{+}97, VKH^{+}98, CKR95]. Vascular-tissue [KWB93]. Vector [DSD94a, Wal98]. Vectoral [DRB97]. Vegetative [NH96, TN96, Sak95]. velocimetry [Nöl95]. Velocity [Mor90, Rak97, Seg94, BHP92]. Velocity-induced [Mor90]. Venous [GK91]. Ventilation [WGH99]. Ventilation-Perfusion [WGH99]. Ventilatory [Gri94a, KWD^{+}94]. ventral [SvHT97, Tak95b]. Ventricle [Cha98, CBS^{+}99]. Ventricular [Adl93, CK97b, Cha98]. ventriculi [JH95]. Verification [BK90a, KB98b]. Versus [SH93b, Cam95, CB94b, LL99]. Vertebrae [JO96]. Vertebral [JO96]. Vertebrate [DO99, MS98, MH97, Mul94, Pol92a]. Vertebrates [AP99]. vertical [Now91]. very [JH95, Kon95]. Vesicle [CB96a, GMU95, HKT91]. Vesicles [CB97a, KN93, LPB96, MKGB98, HKT91]. vessel [DB95]. Vestibular [Mul94]. via [Cum99a, Fri91, HH95, Leb96, SL91a, Val92]. Viability [BOG^{+}99, AHA92]. Viable [Rod96]. Vibrational [Hav99]. Victim [Gav97]. View [Kat94b, SC93, WN99c, Maz91]. viewed [SH91a]. Viewpoint [FOO94, Aok92]. vigilant [PA90]. vimentin [HKT91]. Vinci [Lon94]. Viral [HMUA93, Sza92b, WDP98, MRB91, NBSM97]. Viroids [CF94a].

Virological [Whi97]. Virulence [CCVH98, Dam99, KD99b, LM94, MA98, SMS98, Boy95, LN95]. Virus [BR94, FBHK98, Kel96, LRZ^{+}96, NBBH99, NBSM97, RWN98, SFGG^{+}99, Sza93a, CPLA95, MK90, NH95, PNB92, Sza92a]. Virus-neutralizing [FBHK98]. virus-type [NH95]. Viruses [AH94, HW98a, Mel93a, RBH93, YY99, ALM95, Cha91, Now91]. viscoelastic [DS92]. Viscous [vBR98]. Vision [Chi96, Hak94, Hor95, McI95]. Visual [VB93]. Vital [BSN94a]. Vitro [FDT^{+}93, MBSM93, BW98, Cár91, IET99]. Viviparity [Bla98, QAM97, Bla95]. Vivo [HRW93, BS91b, Cár91, Fer92, Jai97, Kuz95a, Rog94, SK96a, Sal96, Sav91b, SP92c, SP92d]. Voids [RS98]. Volant [JSB93]. Volatile [BDK^{+}97, FSD91]. Voltage [BH94, BG97, LCR94, LK91, Mar92b, SM90a, SW91]. voltage-dependent [SW91]. Volterra [DM91, Sat96, Van91]. Volume [Adl93, AKMV96, Ben99, Nak96, Ano90l, Ano90m, Ano90n, Ano90o, Ano90p, Ano90q, Ano91m, Ano91n, Ano91o, Ano91p, Ano91q, Ano91r, Ano92j, Ano92k, Ano92l, Ano92m, Ano92n, Ano92o, FSDT91, SW95a]. Voronoi [MU92]. vs [Har99a, Ng91, NS91]. Vulnerability [ZH97]. Vulpes [DM92b]. Vulture [HL98b].

Waals [Nak96]. wait [Nis99]. Walk [EIU94, HH97, Hou90]. walker [Fra95c]. Walking [MHC95]. Walks [AH98a, Wil92a]. Wall [Cha93, KW94, SW95a, WG92]. Walls [PD93, DB95]. Walnut [EGDG^{+}98]. War [Bla97, Kim93, KK98, Nuy94]. Ward [PMdM^{+}99]. Wars [MGMD96]. Was [Di 95]. Washout [WGH98]. wasp [TBD95]. Wasps [Bou94, KP93a]. Water [GM94, HVKS99, SS98, SC96, Boi95, FAB92, Hor95, Tri92]. water-cohesion-tension [Tri92]. waterfalls [GK91]. waters [Kno91]. Watson [KPK92]. Watson-Crick [KPK92]. Wave [PK99, SKI98, FTCM90, ISN91, LF95, LH91, Mir90]. waveform [Nak91]. Wavelengths [For96]. Wavelet [LPM95]. Waves [BP99a, Cum99b, SB94, WS98, Jaf95, Vic91]. Way [Fis99, For98, BS95, LA90a]. Weak [LW94, VC98, TS91]. weak-binding [TS91]. Weak-electrolyte [LW94]. Weaker [Hur97a]. Weakly [End97]. Weakly-coupled [End97]. web [GV91, KV97]. web-building [GV91, KV97]. Weber [KB98a]. Weevil [TLS94]. Weevils [SLS93]. Weibull [HDF^{+}99]. weights [JH95]. Well [HVKS99]. Well-mixed [HVKS99]. wenckebach [CLF95]. wenckebach-like [CLF95]. Were [Gor95]. whales [Boi95, LIWE90]. Wheat [BBS^{+}98]. wheels [KG90]. Where [Sak97, Coh91, EEG93, Kel91, Sak96, ZT92]. which [Rad91, RP99, TCE95, VHGV92]. White [HHM98, CC95]. Who [MHSR^{+}96]. Whole [Rak98, Mar95a]. whole-body [Mar95a]. Whole-number [Rak98]. Whorled [DC96b]. whose [Cle91b]. wide [SS90a]. Width [WGH99]. wigner [ÉT90]. Wildebeest [GL93]. Will [Hes94, Ing94]. Willful [Ing94]. Win [Pos99]. Win-Stay [Pos99]. Winding [Abe98]. Winding-down [Abe98]. windows [Val92]. Winds [SEB99]. windthrow [Bak95]. Wine [BS94]. Wing [SZK98, Wal98]. Wings [LD94, Tho96a]. winter [LL90]. Wishing [DH92]. Within [AWA98, Wat97, BWK96, CPLA95, Grü95, Kan92]. Within-host [AWA98]. Without [GH97, MGMD96, áSá96, AHB^{+}91, Atk91, Dam99, DBB99, HT96, HS98c, Osb93, Sau94, TH98b]. Wobbling [Plu94]. Wolbachia [Fra98, Hur97b]. Wolf [WLM96]. Wolf-Pack [WLM96]. Womersley [LT98]. Wood [DH98]. Words [CMPR93]. Work [KS98, Sil98, GS95b]. Worker [Bou94, Yam93]. workers [Rat90]. World [CF94a, GSS98, Gor95, Lah91]. worlds [Man90]. Worth [TAN94b]. Wound [GMST99, OSM95, TM92b]. wren [SW95b].

X [SP99, BKA97, Eas99, For94b, IP99a, Mor96, SST91]. X-chromosome [SP99]. X-Ray [Eas99, Mor96]. Xenopus [Con93, Jaf95]. Xylem [DH98, Kra99].

Years [JBSdlG97, Mar96, SLR^{+}98, Por96]. Yeast [FLR93, LL96, NAZG95, NT95, RSB92]. Yellow [RWB98]. Yield [PN98, Gro92]. Yockey [Eli96]. young [Hou90, Ng91].

Z [ZLYZ98]. Zebra [Kor97]. Zebrafish [TMI99]. Zee [FL95]. Zipf [PMdM^{+}99]. Zone [KKS97, SSA94, WHL93, RDHD98]. Zones [MKGB98]. zooids [Grü95]. zooplankton [Jør95]. zoospore [FTD90].

References

Ainscow:1998:EAM

[AB98] Edward K. Ainscow and Martin D. Brand. Errors associated with metabolic control analysis. application of Monte-Carlo simulation of experimental data. Journal of Theoretical Biology, 194(2):223–233, September 21, 1998. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

              tronic). URL http://www.sciencedirect.com/science/
              article/pii/S0022519398907603.

Abedon:1990:SLI

[Abe90] Stephen T. Abedon. Selection for lysis inhibition in bacteriophage. Journal of Theoretical Biology, 146(4):501–511, October 21, 1990. CODEN JTBIAP. ISSN 0022-5193 (print),

              1095-8541 (electronic). URL http://www.sciencedirect.
              com/science/article/pii/S0022519305803753.

Abernethy:1998:GMO

[Abe98] John Abernethy. Gompertzian mortality originates in the winding-down of the mitotic clock. Journal of Theoretical Biology, 192(4):419–435, June 21, 1998. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519398906579.

Aharoni:1994:DMD

[ABH94] Y. Aharoni, A. Brosh, and Z. Holzer. Double-marker digestive tract model. Journal of Theoretical Biology, 166(4):407–417, February 21, 1994. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519384710368.

Axelrod:1994:GAU

[ABK94] David E. Axelrod, Keith A. Baggerly, and Marek Kimmel. Gene amplification by unequal sister chromatid exchange: Probabilistic modeling and analysis of drug resistance data. Journal of Theoretical Biology, 168(2):151–159, May 21, 1994. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519384710952.

Anissimov:1997:ITM

[ABR97] Yuri G. Anissimov, Anthony J. Bracken, and Michael S. Roberts. Interconnected-tubes model of hepatic elimination. Journal of Theoretical Biology, 188(1):89–101, September 7, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519397904550.

Anissimov:1999:ITM

[ABR99] Yuri G. Anissimov, Anthony J. Bracken, and Michael S. Roberts. Interconnected-tubes model of hepatic elimination: Steady-state considerations. Journal of Theoretical Biology, 199(4):435–447, August 21, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519399909700.

Albe:1990:CCE

[ABW90] Kathy R. Albe, Margaret H. Butler, and Barbara E. Wright. Cellular concentrations of enzymes and their substrates. Journal of Theoretical Biology, 143(2):163–195, March 22, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305802668.

Amir:1990:ORE

[AC90] Shmuel Amir and Dan Cohen. Optimal reproductive efforts and the timing of reproduction of annual plants in randomly varying environments. Journal of Theoretical Biology, 147(1):17–42, November 7, 1990. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305802504.

Adamo:1996:DSH

[AC96] Shelley A. Adamo and R. Chase. Dart shooting in helicid snails: An ‘honest’ signal or an instrument of manipulation? Journal of Theoretical Biology, 180(1):77–80, May 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519396900806.

Amabile-Cuevas:1996:PRP

[ACC96] Carlos F. Amábile-Cuevas and M. E. Chicurel. A possible role for plasmids in mediating the cell-cell proximity required for gene flux. Journal of Theoretical Biology, 181(3):237–243, August 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396901289.

Acerenza:1993:MCD

[Ace93] Luis Acerenza. Metabolic control design. Journal of Theoretical Biology, 165(1):63–85, November 7, 1993. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251938371177X.

Acerenza:1996:HCM

[Ace96a] Luis Acerenza. How constrained is metabolic control? Journal of Theoretical Biology, 182(3):277–283, October 7, 1996.

CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396901654.

Acerenza:1996:PG

[Ace96b] Luis Acerenza. Personal glimpses. Journal of Theoretical Biology, 182(3):201–207, October 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519396901563.

Adams:1998:HD

[Ada98] Paul Adams. Hebb and Darwin. Journal of Theoretical Biology, 195(4):419–438, December 21, 1998. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519397906202.

Arnobio:1990:TBE

[AdG90] A. Arnóbio and S. da Gama. Through-bond electrontransfer interaction in proteins. Journal of Theoretical Biology, 142(2):251–260, January 23, 1990. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305802267.

Adler:1993:TAH

[Adl93] David Adler. Theoretical analysis of the hemodynamic variables during sustained mechanical alternans: Effect of variations in ventricular filling volume. Journal of Theoretical Biology, 163(1):1–14, July 1, 1993. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519383711033.

Arthur:1999:PVC

[AF99] Wallace Arthur and Malcolm Farrow. The pattern of variation in centipede segment number as an example of developmental constraint in evolution. Journal of Theoretical Biology, 200(2):183–191, September 21, 1999. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519399909864.

Arques:1997:EMC

[AFM97] Didier G. Arquès, Jean-Paul Fallot, and Christian J. Michel. An evolutionary model of a complementary circular code. Journal of Theoretical Biology, 185(2):241–253, March 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519396903057.

Airoldi:1995:DBT

[AFS95] J.-P. Airoldi, B. D. Flury, and M. Salvioni. Discrimination between two species of microtus using both classified and unclassified observations. Journal of Theoretical Biology, 177(3):247–262, December 7, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519385702421.

Agarwal:1995:CSE

[Aga95] Pankaj Agarwal. Cellular segregation and engulfment simulations using the cell programming language. Journal of Theoretical Biology, 176(1):79–89, September 7, 1995. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385701786.

Achs:1991:CMP

[AGG91] Murray J. Achs, Lillian Garfinkel, and David Garfinkel. A computer model of pancreatic islet glycolysis. Journal of Theoretical Biology, 150(1):109–135, May 7, 1991. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305804783.

Alonso:1990:TCT

[AH90] Guillermo L. Alonso and Juan P. Hecht. Thermodynamics of Ca^{2+} transport through sarcoplasmic reticulum membranes during the transient-state of simulated reactions. Journal of Theoretical Biology, 147(2):161–176, November 21, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305800505.

Aita:1994:PDS

[AH94] Takuyo Aita and Yuzuru Husimi. Period dependent selection in continuous culture of viruses in a periodic environment. Journal of Theoretical Biology, 168(3):281–289, June 7, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251938471109X.

Aita:1996:FSA

[AH96] Takuyo Aita and Yuzuru Husimi. Fitness spectrum among random mutants on Mt. Fuji-type fitness landscape. Journal of Theoretical Biology, 182(4):469–485, October 21, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396901897.

Aita:1998:AWF

[AH98a] Takuyo Aita and Yuzuru Husimi. Adaptive walks by the fittest among finite random mutants on a Mt. Fuji-type fitness landscape. Journal of Theoretical Biology, 193(3):383–405, August 7, 1998. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519398907093.

Aita:1998:FLB

[AH98b] Takuyo Aita and Yuzuru Husimi. Fitness landscape of a biopolymer participating in a multi-step reaction. Journal of Theoretical Biology, 191(4):377–390, April 21, 1998. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519397905956.

Allen:1992:PTM

[AHA92] E. J. Allen, J. M. Harris, and L. J. S. Allen. Persistencetime models for use in viability analyses of vanishing species. Journal of Theoretical Biology, 155(1):33–53, March 7, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305805478.

Ahlborn:1991:FF

[AHB^{+}91] Boye Ahlborn, David G. Harper, Robert W. Blake, Dorit Ahlborn, and Michael Cam. Fish without footprints. Journal of Theoretical Biology, 148(4):521–533, February 21, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305802346.

Azar:1995:SAH

[AHL95] Christian Azar, John Holmberg, and Kristian Lindgren. Stability analysis of harvesting in a predator–prey model. Journal of Theoretical Biology, 174(1):13–19, May 7, 1995. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385700768.

Abella:1999:TSC

[AIK99] Charles A. Abella, Volodymyr N. Ivanov, and In S. Kim. Taxon-specific content of oligonucleotide triplets in 16s rRNAs of anoxygenic phototrophic and nitrifying bacteria. Journal of Theoretical Biology, 196(3):289–296, February 7, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519398908396.

Antia:1994:MNS

[AK94] Rustom Antia and Jacob C. Koella. A model of non-specific immunity. Journal of Theoretical Biology, 168(2):141–150, May 21, 1994. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519384710940.

Aksenov:1997:MMS

[AKL97] S. V. Aksenov, E. A. Krasavin, and A. A. Litvin. Mathematical model of the SOS response regulation of an excision repair deficient mutant of Escherichia coli after ultraviolet light irradiation. Journal of Theoretical Biology, 186(2):251–260, May 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396903537.

Aubrey:1997:FPE

[AKMS97] Laurence Aubrey, Gérard Klein, Jean-Louis Martiel, and Michel Satre. Fluid-phase endocytosis in the amoebae of the cellular slime mould Dictyostelium discoideum: Mathematical modelling of kinetics and pH evolution. Journal of Theoretical Biology, 184(1):89–98, January 7, 1997. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396902441.

Ataullakhanov:1996:PRAb

[AKMV96] Fazoil I. Ataullakhanov, Svetlana V. Komarova, Mikhail V. Martynov, and Victor M. Vitvitsky. A possible role of adenylate metabolism in human erythrocytes. 2. adenylate metabolism is able to improve the erythrocyte volume stabilization. Journal of Theoretical Biology, 183(3):307–316, December 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396902222.

Ataullakhanov:1996:PRAa

[AKV96] Fazoil I. Ataullakhanov, Svetlana V. Komarova, and Victor M. Vitvitsky. A possible role of adenylate metabolism in human erythrocytes: Simple mathematical model. Journal of Theoretical Biology, 179(1):75–86, March 7, 1996. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396900508.

Arino:1995:MML

[AKW95] Ovide Arino, Marek Kimmel, and Glenn F. Webb. Mathematical modeling of the loss of telomere sequences. Journal of Theoretical Biology, 177(1):45–57, November 7, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385702238.

Agrawal:1996:MMH

[AL96] Nancy G. B. Agrawal and Jennifer J. Linderman. Mathematical modeling of helper T lymphocyte/antigen-presenting cell interactions: Analysis of methods for modifying antigen processing and presentation. Journal of Theoretical Biology, 182(4):487–504, October 21, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519396901903.

Aarhem:1997:CCC

[ÅL97] Peter Århem and Hans Liljenström. On the coevolution of cognition and consciousness. Journal of Theoretical Biology, 187(4):601–612, August 21, 1997. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396903938.

Alexander:1991:OTM

[Ale91] R. McN. Alexander. Optimum timing of muscle activation for simple models of throwing. Journal of Theoretical Biology, 150(3):349–372, June 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305804345.

Alexander:1997:OMD

[Ale97a] R. McN. Alexander. Optimum muscle design for oscillatory movements. Journal of Theoretical Biology, 184(3):253–259, February 7, 1997. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396902714.

Alexander:1997:TMC

[Ale97b] R. McN. Alexander. A theory of mixed chains applied to safety factors in biological systems. Journal of Theoretical Biology, 184(3):247–252, February 7, 1997. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396902702.

Allen:1993:CGE

[All93a] John F. Allen. Control of gene expression by redox potential and the requirement for chloroplast and mitochondrial genomes. Journal of Theoretical Biology, 165(4):609–631, December 21, 1993. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519383712105.

Allison:1993:FAO

[All93b] Lloyd Allison. A fast algorithm for the optimal alignment of three strings. Journal of Theoretical Biology, 164(2):

261–269, September 21, 1993. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519383711537.

Allison:1993:NAG

[All93c] Lloyd Allison. Normalization of affine gap costs used in optimal sequence alignment. Journal of Theoretical Biology, 161(2):263–269, March 21, 1993. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519383710544.

Allen:1996:SSM

[All96] John F. Allen. Separate sexes and the mitochondrial theory of ageing. Journal of Theoretical Biology, 180(2):135–140, May 21, 1996. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396900892.

Arques:1995:ISS

[ALM95] Didier G. Arquès, Jean-Christophe Lapayre, and Christian J. Michel. Identification and simulation of shifted periodicities common to protein coding genes of eukaryotes, prokaryotes and viruses. Journal of Theoretical Biology, 172(3):279–291, February 7, 1995. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519385700240.

Almirantis:1999:SDB

[Alm99] Yannis Almirantis. A standard deviation based quantification differentiates coding from non-coding DNA sequences and gives insight to their evolutionary history. Journal of Theoretical Biology, 196(3):297–308, February 7, 1999. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519398908402.

Acosta:1993:BAA

[ALS93] Francisco J. Acosta, Francisco López, and Jose M. Serrano. Branching angles of ant trunk trails as an optimization cue. Journal of Theoretical Biology, 160(3):297–310, February 7, 1993. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519383710209.

Antia:1991:QMS

[ALW91] Rustom Antia, Bruce Levin, and Patrick Williamson. A quantitative model suggests immune memory involves the colocalization of B and Th cells. Journal of Theoretical Biology, 153(3):371–384, December 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305805764.

Arques:1990:PCN

[AM90] Didier G. Arques and Christian J. Michel. Periodicities in coding and noncoding regions of the genes. Journal of Theoretical Biology, 143(3):307–318, April 5, 1990. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305800323.

Arques:1992:SGP

[AM92] Didier G. Arquès and Christian J. Michel. A simulation of the genetic periodicities modulo 2 and 3 with processes of nucleotide insertions and deletions. Journal of Theoretical Biology, 156(1):113–127, May 7, 1992. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305806599.

Arques:1995:ASD

[AM95] Didier G. Arquès and Christian J. Michel. Analytical solutions of the dinucleotide probability after and before random mutations. Journal of Theoretical Biology, 175(4):533–544, August 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519385701610.

Arques:1996:CCC

[AM96] Didier G. Arquès and Christian J. Michel. A complementary circular code in the protein coding genes. Journal of Theoretical Biology, 182(1):45–58, September 7, 1996. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396901423.

Arques:1997:CCP

[AM97] Didier G. Arquès and Christian J. Michel. A circular code in the protein coding genes of mitochondria. Journal of Theoretical Biology, 189(3):273–290, December 7, 1997. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519397905130.

Acuna:1998:MAB

[AM98] Jose D. Acuña and Miguel A. Muñoz. Mixture analysis in biology: Scope and limits. Journal of Theoretical Biology, 191(3):341–344, April 7, 1998. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519397905944.

Adams:1995:CTD

[AMG95] Eldridge S. Adams and Michael Mesterton-Gibbons. The cost of threat displays and the stability of deceptive communication. Journal of Theoretical Biology, 175(4):405–421, August 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519385701518.

Arques:1993:ISN

[AMO93] Didier G. Arquès, Christian J. Michel, and Karine Orieux. Identification and simulation of new non-random statistical properties common to different populations of eukaryotic noncoding genes. Journal of Theoretical Biology, 161(3):329–342, April 7, 1993. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519383710593.

Ashida:1991:TDM

[AMT91] Hiroshi Ashida, Ehchi Maru, and Jiro Tatsuno. Time discrete model of recurrent inhibition in hippocampal dentate gyrus. Journal of Theoretical Biology, 148(1):1–16, January 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251930580462X.

Agutter:1995:ITM

[AMW95] P. S. Agutter, P. C. Malone, and D. N. Wheatley. Intracellular transport mechanisms: a critique of diffusion theory. Journal of Theoretical Biology, 176(2):261–272, September 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519385701968.

Ancel:1999:QMS

[Anc99] Lauren W. Ancel. A quantitative model of the Simpson-Baldwin effect. Journal of Theoretical Biology, 196(2):197–209, January 21, 1999. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519398908335.

Anderson:1995:LEQ

[And95] Russell W. Anderson. Learning and evolution: a quantitative genetics approach. Journal of Theoretical Biology, 175(1):89– 101, July 7, 1995. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519385701233.

Anderies:1998:CHA

[And98] John M. Anderies. Culture and human agro-ecosystem dynamics: the tsembaga of new guinea. Journal of Theoretical Biology, 192(4):515–530, June 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519398906816.

Anonymous:1990:EBa

[Ano90a] Anonymous. Editorial Board. Journal of Theoretical Biology, 142(1):i, January 9, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519305800086.

Anonymous:1990:EBb

[Ano90b] Anonymous. Editorial Board. Journal of Theoretical Biology, 143(1):i, March 8, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519305802838.

Anonymous:1990:EBc

[Ano90c] Anonymous. Editorial Board. Journal of Theoretical Biology, 144(1):i, May 10, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519305802942.

Anonymous:1990:EBd

[Ano90d] Anonymous. Editorial Board. Journal of Theoretical Biology, 145(1):i, July 9, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519305805296.

Anonymous:1990:EBe

[Ano90e] Anonymous. Editorial Board. Journal of Theoretical Biology, 146(1):i, September 7, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519305800396.

Anonymous:1990:EBf

[Ano90f] Anonymous. Editorial Board. Journal of Theoretical Biology, 147(1):i, November 7, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519305802474.

Anonymous:1990:Ed

[Ano90g] Anonymous. Errata. Journal of Theoretical Biology, 144(1): 143, May 10, 1990. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305803066.

Anonymous:1990:Ee

[Ano90h] Anonymous. Errata. Journal of Theoretical Biology, 145(1): 141, July 9, 1990. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305805429.

Anonymous:1990:Ef

[Ano90i] Anonymous. Errata. Journal of Theoretical Biology, 147 (1):143, November 7, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519305802577.

Anonymous:1990:Ea

[Ano90j] Anonymous. Erratum. Journal of Theoretical Biology, 142 (1):139, January 9, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519305800190.

Anonymous:1990:Ec

[Ano90k] Anonymous. Erratum. Journal of Theoretical Biology, 143 (2):285, March 22, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519305802735.

Anonymous:1990:IVa

[Ano90l] Anonymous. Index to volume 142. Journal of Theoretical Biology, 142(1):vi–x, January 9, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305800098.

Anonymous:1990:IVb

[Ano90m] Anonymous. Index to volume 143. Journal of Theoretical Biology, 143(1):vi–x, March 8, 1990. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251930580284X.

Anonymous:1990:IVc

[Ano90n] Anonymous. Index to volume 144. Journal of Theoretical Biology, 144(1):vi–xi, May 10, 1990. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305802954.

Anonymous:1990:IVd

[Ano90o] Anonymous. Index to volume 145. Journal of Theoretical Biology, 145(1):vi–xi, July 9, 1990. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305805302.

Anonymous:1990:IVe

[Ano90p] Anonymous. Index to volume 146. Journal of Theoretical Biology, 146(1):vi–xi, September 7, 1990. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305800402.

Anonymous:1990:IVf

[Ano90q] Anonymous. Index to volume 147. Journal of Theoretical Biology, 147(1):vi–ix, November 7, 1990. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305802486.

Anonymous:1990:PJa

[Ano90r] Anonymous. Pages 1–139 (9 January 1990). Journal of Theoretical Biology, 142(1):??, January 9, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1990:PJe

[Ano90s] Anonymous. Pages 1–141 (9 July 1990). Journal of Theoretical Biology, 145(1):??, July 9, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1990:PMc

[Ano90t] Anonymous. Pages 1–143 (10 May 1990). Journal of Theoretical Biology, 144(1):??, May 10, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1990:PNa

[Ano90u] Anonymous. Pages 1–143 (7 November 1990). Journal of Theoretical Biology, 147(1):??, November 7, 1990. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1990:PSa

[Ano90v] Anonymous. Pages 1–143 (7 September 1990). Journal of Theoretical Biology, 146(1):??, September 7, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1990:PMa

[Ano90w] Anonymous. Pages 1–143 (8 March 1990). Journal of Theoretical Biology, 143(1):??, March 8, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1990:PJb

[Ano90x] Anonymous. Pages 141–284 (23 January 1990). Journal of Theoretical Biology, 142(2):??, January 23, 1990. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1990:PJf

[Ano90y] Anonymous. Pages 143–286 (24 July 1990). Journal of Theoretical Biology, 145(2):??, July 24, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1990:PMd

[Ano90z] Anonymous. Pages 145–285 (14 May 1990). Journal of Theoretical Biology, 144(2):??, May 14, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1990:PMb

[Ano90-27] Anonymous. Pages 145–285 (22 March 1990). Journal of Theoretical Biology, 143(2):??, March 22, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1990:PSb

[Ano90-28] Anonymous. Pages 145–287 (21 September 1990). Journal of Theoretical Biology, 146(2):??, September 21, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1990:PNb

[Ano90-29] Anonymous. Pages 145–288 (21 November 1990). Journal of Theoretical Biology, 147(2):??, November 21, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1990:PFa

[Ano90-30] Anonymous. Pages 285–427 (9 February 1990). Journal of Theoretical Biology, 142(3):??, February 9, 1990. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1990:PJc

[Ano90-31] Anonymous. Pages 287–427 (7 June 1990). Journal of Theoretical Biology, 144(3):??, June 7, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1990:PAc

[Ano90-32] Anonymous. Pages 287–428 (9 August 1990). Journal of Theoretical Biology, 145(3):??, August 9, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1990:PAa

[Ano90-33] Anonymous. Pages 287–430 (5 April 1990). Journal of Theoretical Biology, 143(3):??, April 5, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1990:PDa

[Ano90-34] Anonymous. Pages 289–432 (7 December 1990). Journal of Theoretical Biology, 147(3):??, December 7, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1990:POa

[Ano90-35] Anonymous. Pages 289–432 (7 October 1990). Journal of Theoretical Biology, 146(3):??, October 7, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1990:PJd

[Ano90-36] Anonymous. Pages 429–571 (21 June 1990). Journal of Theoretical Biology, 144(4):??, June 21, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1990:PFb

[Ano90-37] Anonymous. Pages 429–571 (22 February 1990). Journal of Theoretical Biology, 142(4):??, February 22, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1990:PAd

[Ano90-38] Anonymous. Pages 429–572 (23 August 1990). Journal of Theoretical Biology, 145(4):??, August 23, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1990:PAb

[Ano90-39] Anonymous. Pages 431–573 (19 April 1990). Journal of Theoretical Biology, 143(4):??, April 19, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1990:PDb

[Ano90-40] Anonymous. Pages 433–571 (21 December 1990). Journal of Theoretical Biology, 147(4):??, December 21, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1990:POb

[Ano90-41] Anonymous. Pages 433–573 (21 October 1990). Journal of Theoretical Biology, 146(4):??, October 21, 1990. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1991:EBa

[Ano91a] Anonymous. Editorial Board. Journal of Theoretical Biology, 148(1):i, January 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519305804606.

Anonymous:1991:EBb

[Ano91b] Anonymous. Editorial Board. Journal of Theoretical Biology, 149(1):i, March 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519305800669.

Anonymous:1991:EBc

[Ano91c] Anonymous. Editorial Board. Journal of Theoretical Biology, 150(1):i, May 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519305804692.

Anonymous:1991:EBd

[Ano91d] Anonymous. Editorial Board. Journal of Theoretical Biology, 151(1):i, July 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519305801407.

Anonymous:1991:EBe

[Ano91e] Anonymous. Editorial Board. Journal of Theoretical Biology, 152(1):i, September 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519305804989.

Anonymous:1991:EBf

[Ano91f] Anonymous. Editorial Board. Journal of Theoretical Biology, 153(1):i, November 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519305803480.

Anonymous:1991:Ec

[Ano91g] Anonymous. Errata. Journal of Theoretical Biology, 148 (4):563, February 21, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519305802371.

Anonymous:1991:Ee

[Ano91h] Anonymous. Errata. Journal of Theoretical Biology, 152 (2):285–286, September 21, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002251930580459X.

Anonymous:1991:Ef

[Ano91i] Anonymous. Errata. Journal of Theoretical Biology, 152 (4):569–571, October 21, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519305803996.

Anonymous:1991:Ea

[Ano91j] Anonymous. Erratum. Journal of Theoretical Biology, 148 (3):431, February 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519305802462.

Anonymous:1991:Ed

[Ano91k] Anonymous. Erratum. Journal of Theoretical Biology, 151 (4):573, August 21, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519305803716.

Anonymous:1991:Eg

[Ano91l] Anonymous. Erratum. Journal of Theoretical Biology, 153 (4):603, December 21, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519305801584.

Anonymous:1991:IVa

[Ano91m] Anonymous. Index to volume 148. Journal of Theoretical Biology, 148(1):vi–x, January 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305804618.

Anonymous:1991:IVb

[Ano91n] Anonymous. Index to volume 149. Journal of Theoretical Biology, 149(1):vi–ix, March 7, 1991. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305800670.

Anonymous:1991:IVc

[Ano91o] Anonymous. Index to volume 150. Journal of Theoretical Biology, 150(1):vi–xi, May 7, 1991. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305804709.

Anonymous:1991:IVd

[Ano91p] Anonymous. Index to volume 151. Journal of Theoretical Biology, 151(1):vi–x, July 7, 1991. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305801419.

Anonymous:1991:IVe

[Ano91q] Anonymous. Index to volume 152. Journal of Theoretical Biology, 152(1):vii–xiii, September 7, 1991. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305804990.

Anonymous:1991:IVf

[Ano91r] Anonymous. Index to volume 153. Journal of Theoretical Biology, 153(1):vii–xi, November 7, 1991. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305803492.

Anonymous:1991:PMc

[Ano91s] Anonymous. Pages 1–135 (7 May 1991). Journal of Theoretical Biology, 150(1):??, May 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1991:PSa

[Ano91t] Anonymous. Pages 1–141 (7 September 1991). Journal of Theoretical Biology, 152(1):??, September 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1991:PJe

[Ano91u] Anonymous. Pages 1–143 (7 July 1991). Journal of Theoretical Biology, 151(1):??, July 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1991:PJa

[Ano91v] Anonymous. Pages 1–144 (7 January 1991). Journal of Theoretical Biology, 148(1):??, January 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1991:PMa

[Ano91w] Anonymous. Pages 1–144 (7 March 1991). Journal of Theoretical Biology, 149(1):??, March 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1991:PNa

[Ano91x] Anonymous. Pages 1–144 (7 November 1991). Journal of Theoretical Biology, 153(1):??, November 7, 1991. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1991:PMd

[Ano91y] Anonymous. Pages 137–276 (21 May 1991). Journal of Theoretical Biology, 150(2):??, May 21, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1991:PSb

[Ano91z] Anonymous. Pages 143–286 (21 September 1991). Journal of Theoretical Biology, 152(2):??, September 21, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1991:PNb

[Ano91-27] Anonymous. Pages 145–285 (21 November 1991). Journal of Theoretical Biology, 153(2):??, November 21, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1991:PMb

[Ano91-28] Anonymous. Pages 145–286 (21 March 1991). Journal of Theoretical Biology, 149(2):??, March 21, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1991:PJb

[Ano91-29] Anonymous. Pages 145–288 (21 January 1991). Journal of Theoretical Biology, 148(2):??, January 21, 1991. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1991:PJf

[Ano91-30] Anonymous. Pages 145–288 (21 July 1991). Journal of Theoretical Biology, 151(2):??, July 21, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1991:PJc

[Ano91-31] Anonymous. Pages 277–420 (7 June 1991). Journal of Theoretical Biology, 150(3):??, June 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1991:PAa

[Ano91-32] Anonymous. Pages 287–427 (7 April 1991). Journal of Theoretical Biology, 149(3):??, April 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1991:POa

[Ano91-33] Anonymous. Pages 287–428 (7 October 1991). Journal of Theoretical Biology, 152(3):??, October 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1991:PDa

[Ano91-34] Anonymous. Pages 287–444 (7 December 1991). Journal of Theoretical Biology, 153(3):??, December 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1991:PAc

[Ano91-35] Anonymous. Pages 289–431 (7 August 1991). Journal of Theoretical Biology, 151(3):??, August 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1991:PFa

[Ano91-36] Anonymous. Pages 289–431 (7 February 1991). Journal of Theoretical Biology, 148(3):??, February 7, 1991. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1991:PJd

[Ano91-37] Anonymous. Pages 421–563 (21 June 1991). Journal of Theoretical Biology, 150(4):??, June 21, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1991:PAb

[Ano91-38] Anonymous. Pages 429–571 (21 April 1991). Journal of Theoretical Biology, 149(4):??, April 21, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1991:POb

[Ano91-39] Anonymous. Pages 429–571 (21 October 1991). Journal of Theoretical Biology, 152(4):??, October 21, 1991. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1991:PFb

[Ano91-40] Anonymous. Pages 433–563 (21 February 1991). Journal of Theoretical Biology, 148(4):??, February 21, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1991:PAd

[Ano91-41] Anonymous. Pages 433–573 (21 August 1991). Journal of Theoretical Biology, 151(4):??, August 21, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1991:PDb

[Ano91-42] Anonymous. Pages 445–603 (21 December 1991). Journal of Theoretical Biology, 153(4):??, December 21, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1992:EBa

[Ano92a] Anonymous. Editorial Board. Journal of Theoretical Biology, 154(1):i, January 7, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002251930580181X.

Anonymous:1992:EBb

[Ano92b] Anonymous. Editorial Board. Journal of Theoretical Biology, 155(1):i, March 7, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519305805430.

Anonymous:1992:EBc

[Ano92c] Anonymous. Editorial Board. Journal of Theoretical Biology, 156(1):i, May 7, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519305806526.

Anonymous:1992:EBd

[Ano92d] Anonymous. Editorial Board. Journal of Theoretical Biology, 157(1):i, July 7, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519305807519.

Anonymous:1992:EBe

[Ano92e] Anonymous. Editorial Board. Journal of Theoretical Biology, 158(1):i, September 7, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519305806435.

Anonymous:1992:EBf

[Ano92f] Anonymous. Editorial Board. Journal of Theoretical Biology, 159(1):i, November 7, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519305807623.

Anonymous:1992:Ed

[Ano92g] Anonymous. Errata. Journal of Theoretical Biology, 158 (2):269, September 21, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519305807234.

Anonymous:1992:Ea

[Ano92h] Anonymous. Erratum. Journal of Theoretical Biology, 154 (4):539, February 21, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519305801687.

Anonymous:1992:Ec

[Ano92i] Anonymous. Erratum. Journal of Theoretical Biology, 157(2): 269, July 21, 1992. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305806253.

Anonymous:1992:IVa

[Ano92j] Anonymous. Index to volume 154. Journal of Theoretical Biology, 154(1):vii–xii, January 7, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305801821.

Anonymous:1992:IVb

[Ano92k] Anonymous. Index to volume 155. Journal of Theoretical Biology, 155(1):vi–viii, x, March 7, 1992. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305805442.

Anonymous:1992:IVc

[Ano92l] Anonymous. Index to volume 156. Journal of Theoretical Biology, 156(1):vi–ix, May 7, 1992. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305806538.

Anonymous:1992:IVd

[Ano92m] Anonymous. Index to volume 157. Journal of Theoretical Biology, 157(1):vi–x, July 7, 1992. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305807520.

Anonymous:1992:IVe

[Ano92n] Anonymous. Index to volume 158. Journal of Theoretical Biology, 158(1):vi–x, September 7, 1992. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305806447.

Anonymous:1992:IVf

[Ano92o] Anonymous. Index to volume 158. Journal of Theoretical Biology, 159(1):vi–x, November 7, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305807635.

Anonymous:1992:PMc

[Ano92p] Anonymous. Pages 1–132 (7 May 1992). Journal of Theoretical Biology, 156(1):??, May 7, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1992:PJe

[Ano92q] Anonymous. Pages 1–133 (7 July 1992). Journal of Theoretical Biology, 157(1):??, July 7, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1992:PNa

[Ano92r] Anonymous. Pages 1–133 (7 November 1992). Journal of Theoretical Biology, 159(1):??, November 7, 1992. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1992:PSa

[Ano92s] Anonymous. Pages 1–134 (7 September 1992). Journal of Theoretical Biology, 158(1):??, September 7, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1992:PJa

[Ano92t] Anonymous. Pages 1–136 (7 January 1992). Journal of Theoretical Biology, 154(1):??, January 7, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1992:PMa

[Ano92u] Anonymous. Pages 1–136 (7 March 1992). Journal of Theoretical Biology, 155(1):??, March 7, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1992:PMd

[Ano92v] Anonymous. Pages 133–267 (21 May 1992). Journal of Theoretical Biology, 156(2):??, May 21, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1992:PJf

[Ano92w] Anonymous. Pages 135–269 (21 July 1992). Journal of Theoretical Biology, 157(2):??, July 21, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1992:PSb

[Ano92x] Anonymous. Pages 135–269 (21 September 1992). Journal of Theoretical Biology, 158(2):??, September 21, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1992:PNb

[Ano92y] Anonymous. Pages 135–270 (21 November 1992). Journal of Theoretical Biology, 159(2):??, November 21, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1992:PJb

[Ano92z] Anonymous. Pages 137–270 (21 January 1992). Journal of Theoretical Biology, 154(2):??, January 21, 1992. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1992:PMb

[Ano92-27] Anonymous. Pages 137–272 (21 March 1992). Journal of Theoretical Biology, 155(2):??, March 21, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1992:PJc

[Ano92-28] Anonymous. Pages 269–403 (7 June 1992). Journal of Theoretical Biology, 156(3):??, June 7, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1992:PID

[Ano92-29] Anonymous. Pages 271–352, IN1, 353-379 (7 December 1992). Journal of Theoretical Biology, 159(3):??, December 7, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1992:PFa

[Ano92-30] Anonymous. Pages 271–404 (7 February 1992). Journal of Theoretical Biology, 154(3):??, February 7, 1992. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1992:PAc

[Ano92-31] Anonymous. Pages 271–406 (7 August 1992). Journal of Theoretical Biology, 157(3):??, August 7, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1992:POa

[Ano92-32] Anonymous. Pages 271–407 (7 October 1992). Journal of Theoretical Biology, 158(3):??, October 7, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1992:PAa

[Ano92-33] Anonymous. Pages 273–407 (7 April 1992). Journal of Theoretical Biology, 155(3):??, April 7, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1992:PD

[Ano92-34] Anonymous. Pages 381–552 (21 December 1992). Journal of Theoretical Biology, 159(4):??, December 21, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1992:PFb

[Ano92-35] Anonymous. Pages 405–539 (21 February 1992). Journal of Theoretical Biology, 154(4):??, February 21, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1992:PJd

[Ano92-36] Anonymous. Pages 405–539 (21 June 1992). Journal of Theoretical Biology, 156(4):??, June 21, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1992:PAd

[Ano92-37] Anonymous. Pages 407–542 (21 August 1992). Journal of Theoretical Biology, 157(4):??, August 21, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1992:POb

[Ano92-38] Anonymous. Pages 409–537 (21 October 1992). Journal of Theoretical Biology, 158(4):??, October 21, 1992. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1992:PAb

[Ano92-39] Anonymous. Pages 409–544 (21 April 1992). Journal of Theoretical Biology, 155(4):??, April 21, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1993:E

[Ano93a] Anonymous. Erratum. Journal of Theoretical Biology, 165 (4):673, December 21, 1993. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519383712130.

Anonymous:1993:PMc

[Ano93b] Anonymous. Pages 1–130 (7 May 1993). Journal of Theoretical Biology, 162(1):??, May 7, 1993. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1993:PJa

[Ano93c] Anonymous. Pages 1–133 (7 January 1993). Journal of Theoretical Biology, 160(1):??, January 7, 1993. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1993:PSa

[Ano93d] Anonymous. Pages 1–133 (7 September 1993). Journal of Theoretical Biology, 164(1):??, September 7, 1993. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1993:PMa

[Ano93e] Anonymous. Pages 1–134 (7 March 1993). Journal of Theoretical Biology, 161(1):??, March 7, 1993. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1993:PNa

[Ano93f] Anonymous. Pages 1–134 (7 November 1993). Journal of Theoretical Biology, 165(1):??, November 7, 1993. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1993:PJe

[Ano93g] Anonymous. Pages 1–144 (1 July 1993). Journal of Theoretical Biology, 163(1):??, July 1, 1993. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1993:PMd

[Ano93h] Anonymous. Pages 131–266 (21 May 1993). Journal of Theoretical Biology, 162(2):??, May 21, 1993. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1993:PNb

[Ano93i] Anonymous. Pages 135–267 (21 November 1993). Journal of Theoretical Biology, 165(2):??, November 21, 1993. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1993:PJb

[Ano93j] Anonymous. Pages 135–269 (21 January 1993). Journal of Theoretical Biology, 160(2):??, January 21, 1993. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1993:PSb

[Ano93k] Anonymous. Pages 135–269 (21 September 1993). Journal of Theoretical Biology, 164(2):??, September 21, 1993. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1993:PMb

[Ano93l] Anonymous. Pages 137–269 (21 March 1993). Journal of Theoretical Biology, 161(2):??, March 21, 1993. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1993:PJf

[Ano93m] Anonymous. Pages 145–275 (21 July 1993). Journal of Theoretical Biology, 163(2):??, July 21, 1993. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1993:PJc

[Ano93n] Anonymous. Pages 267–401 (7 June 1993). Journal of Theoretical Biology, 162(3):??, June 7, 1993. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1993:PDa

[Ano93o] Anonymous. Pages 269–453 (7 December 1993). Journal of Theoretical Biology, 165(3):??, December 7, 1993. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1993:PAa

[Ano93p] Anonymous. Pages 271–405 (7 April 1993). Journal of Theoretical Biology, 161(3):??, April 7, 1993. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1993:PFa

[Ano93q] Anonymous. Pages 271–405 (7 February 1993). Journal of Theoretical Biology, 160(3):??, February 7, 1993. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1993:POa

[Ano93r] Anonymous. Pages 271–406 (7 October 1993). Journal of Theoretical Biology, 164(3):??, October 7, 1993. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1993:PAc

[Ano93s] Anonymous. Pages 277–411 (7 August 1993). Journal of Theoretical Biology, 163(3):??, August 7, 1993. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1993:PJd

[Ano93t] Anonymous. Pages 403–537 (21 June 1993). Journal of Theoretical Biology, 162(4):??, June 21, 1993. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1993:POb

[Ano93u] Anonymous. Pages 407–541 (21 October 1993). Journal of Theoretical Biology, 164(4):??, October 21, 1993. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1993:PAb

[Ano93v] Anonymous. Pages 407–542 (21 April 1993). Journal of Theoretical Biology, 161(4):??, April 21, 1993. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1993:PFb

[Ano93w] Anonymous. Pages 407–546 (21 February 1993). Journal of Theoretical Biology, 160(4):??, February 21, 1993. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1993:PAd

[Ano93x] Anonymous. Pages 413–548 (21 August 1993). Journal of Theoretical Biology, 163(4):??, August 21, 1993. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1993:PDb

[Ano93y] Anonymous. Pages 455–673 (21 December 1993). Journal of Theoretical Biology, 165(4):??, December 21, 1993. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1994:E

[Ano94a] Anonymous. Erratum. Journal of Theoretical Biology, 166 (4):479, February 21, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519384710435.

Anonymous:1994:SDR

[Ano94b] Anonymous. Is sterility by deleterious recessives an origin of inequalities in the evolution of eusociality? Journal of Theoretical Biology, 166(1):113–115, January 7, 1994. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384710095.

Anonymous:1994:PSa

[Ano94c] Anonymous. Pages 1–114 (7 September 1994). Journal of Theoretical Biology, 170(1):??, September 7, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1994:PJa

[Ano94d] Anonymous. Pages 1–115 (7 January 1994). Journal of Theoretical Biology, 166(1):??, January 7, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1994:PMc

[Ano94e] Anonymous. Pages 1–115 (7 May 1994). Journal of Theoretical Biology, 168(1):??, May 7, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1994:PJg

[Ano94f] Anonymous. Pages 1–122 (July 1994). Journal of Theoretical Biology, 171(1):??, July 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1994:PMa

[Ano94g] Anonymous. Pages 1–92 (7 March 1994). Journal of Theoretical Biology, 167(1):??, March 7, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1994:PJe

[Ano94h] Anonymous. Pages 1–99 (7 July 1994). Journal of Theoretical Biology, 169(1):??, July 7, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1994:PJf

[Ano94i] Anonymous. Pages 101–207 (21 July 1994). Journal of Theoretical Biology, 169(2):??, July 21, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1994:PSb

[Ano94j] Anonymous. Pages 115–230 (21 September 1994). Journal of Theoretical Biology, 170(2):??, September 21, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1994:PJb

[Ano94k] Anonymous. Pages 117–223 (21 January 1994). Journal of Theoretical Biology, 166(2):??, January 21, 1994. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1994:PMd

[Ano94l] Anonymous. Pages 117–231 (21 May 1994). Journal of Theoretical Biology, 168(2):??, May 21, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1994:PN

[Ano94m] Anonymous. Pages 123–337 (21 November 1994). Journal of Theoretical Biology, 171(2):??, November 21, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1994:PAa

[Ano94n] Anonymous. Pages 201–300 (7 April 1994). Journal of Theoretical Biology, 167(3):??, April 7, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1994:PAc

[Ano94o] Anonymous. Pages 209–315 (7 August 1994). Journal of Theoretical Biology, 169(3):??, August 7, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1994:PFa

[Ano94p] Anonymous. Pages 225–363 (7 February 1994). Journal of Theoretical Biology, 166(3):??, February 7, 1994. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1994:POa

[Ano94q] Anonymous. Pages 231–330 (7 October 1994). Journal of Theoretical Biology, 170(3):??, October 7, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1994:PJc

[Ano94r] Anonymous. Pages 233–348 (7 June 1994). Journal of Theoretical Biology, 168(3):??, June 7, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1994:PDa

[Ano94s] Anonymous. Pages 239–353 (7 December 1994). Journal of Theoretical Biology, 171(3):??, December 7, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1994:PAb

[Ano94t] Anonymous. Pages 301–416 (21 April 1994). Journal of Theoretical Biology, 167(4):??, April 21, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1994:PAd

[Ano94u] Anonymous. Pages 317–432 (21 August 1994). Journal of Theoretical Biology, 169(4):??, August 21, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1994:POb

[Ano94v] Anonymous. Pages 331–430 (21 October 1994). Journal of Theoretical Biology, 170(4):??, October 21, 1994. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1994:PJd

[Ano94w] Anonymous. Pages 349–463 (21 June 1994). Journal of Theoretical Biology, 168(4):??, June 21, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1994:PDb

[Ano94x] Anonymous. Pages 355–437 (21 December 1994). Journal of Theoretical Biology, 171(4):??, December 21, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1994:PFb

[Ano94y] Anonymous. Pages 365–479 (21 February 1994). Journal of Theoretical Biology, 166(4):??, February 21, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1994:PMb

[Ano94z] Anonymous. Pages 93–200 (21 March 1994). Journal of Theoretical Biology, 167(2):??, March 21, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1995:E

[Ano95a] Anonymous. Erratum. Journal of Theoretical Biology, 177 (3):313, December 7, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519385702494.

Anonymous:1995:PNa

[Ano95b] Anonymous. Pages 1–102 (7 November 1995). Journal of Theoretical Biology, 177(1):??, November 7, 1995. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1995:PJa

[Ano95c] Anonymous. Pages 1–105 (7 January 1995). Journal of Theoretical Biology, 172(1):??, January 7, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1995:PMa

[Ano95d] Anonymous. Pages 1–108 (7 March 1995). Journal of Theoretical Biology, 173(1):??, March 7, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1995:PMc

[Ano95e] Anonymous. Pages 1–117 (7 May 1995). Journal of Theoretical Biology, 174(1):??, May 7, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1995:PJe

[Ano95f] Anonymous. Pages 1–126 (7 July 1995). Journal of Theoretical Biology, 175(1):??, July 7, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1995:PSa

[Ano95g] Anonymous. Pages 1–197 (7 September 1995). Journal of Theoretical Biology, 176(1):??, September 7, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1995:PNb

[Ano95h] Anonymous. Pages 103–198 (21 November 1995). Journal of Theoretical Biology, 177(2):??, November 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1995:PJb

[Ano95i] Anonymous. Pages 107–197 (21 January 1995). Journal of Theoretical Biology, 172(2):??, January 21, 1995. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1995:PMb

[Ano95j] Anonymous. Pages 109–215 (21 March 1995). Journal of Theoretical Biology, 173(2):??, March 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1995:PMd

[Ano95k] Anonymous. Pages 119–236 (21 May 1995). Journal of Theoretical Biology, 174(2):??, May 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1995:PJf

[Ano95l] Anonymous. Pages 127–266 (21 July 1995). Journal of Theoretical Biology, 175(2):??, July 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1995:PFa

[Ano95m] Anonymous. Pages 199–297 (7 February 1995). Journal of Theoretical Biology, 172(3):??, February 7, 1995. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1995:PDa

[Ano95n] Anonymous. Pages 199–313 (7 December 1995). Journal of Theoretical Biology, 177(3):??, December 7, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1995:PSb

[Ano95o] Anonymous. Pages 199–316 (21 September 1995). Journal of Theoretical Biology, 176(2):??, September 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1995:PAa

[Ano95p] Anonymous. Pages 217–327 (7 April 1995). Journal of Theoretical Biology, 173(3):??, April 7, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1995:PJc

[Ano95q] Anonymous. Pages 237–353 (7 June 1995). Journal of Theoretical Biology, 174(3):??, June 7, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1995:PAc

[Ano95r] Anonymous. Pages 267–396 (7 August 1995). Journal of Theoretical Biology, 175(3):??, August 7, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1995:PFb

[Ano95s] Anonymous. Pages 299–390 (21 February 1995). Journal of Theoretical Biology, 172(4):??, February 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1995:POa

[Ano95t] Anonymous. Pages 317–432 (7 October 1995). Journal of Theoretical Biology, 176(3):??, October 7, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1995:PDb

[Ano95u] Anonymous. Pages 325–429 (21 December 1995). Journal of Theoretical Biology, 177(4):??, December 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1995:PAb

[Ano95v] Anonymous. Pages 329–440 (21 April 1995). Journal of Theoretical Biology, 173(4):??, April 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1995:PJd

[Ano95w] Anonymous. Pages 355–472 (21 June 1995). Journal of Theoretical Biology, 174(4):??, June 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1995:PAd

[Ano95x] Anonymous. Pages 405–603 (21 August 1995). Journal of Theoretical Biology, 175(4):??, August 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1995:POb

[Ano95y] Anonymous. Pages 433–542 (21 October 1995). Journal of Theoretical Biology, 176(4):??, October 21, 1995. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1996:A

[Ano96a] Anonymous. Announcement. Journal of Theoretical Biology, 178(2):227–228, January 21, 1996. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519396900211.

Anonymous:1996:E

[Ano96b] Anonymous. Erratum. Journal of Theoretical Biology, 183 (1):119, November 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519396902076.

Anonymous:1996:PJa

[Ano96c] Anonymous. Pages 1–111 (7 January 1996). Journal of Theoretical Biology, 178(1):??, January 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1996:PNa

[Ano96d] Anonymous. Pages 1–119 (7 November 1996). Journal of Theoretical Biology, 183(1):??, November 7, 1996. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1996:PMa

[Ano96e] Anonymous. Pages 1–90 (7 March 1996). Journal of Theoretical Biology, 179(1):??, March 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1996:PMc

[Ano96f] Anonymous. Pages 1–92 (7 May 1996). Journal of Theoretical Biology, 180(1):??, May 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1996:PJe

[Ano96g] Anonymous. Pages 1–96 (7 July 1996). Journal of Theoretical Biology, 181(1):??, July 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1996:PSa

[Ano96h] Anonymous. Pages 1–98 (7 September 1996). Journal of Theoretical Biology, 182(1):??, September 7, 1996. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1996:PJb

[Ano96i] Anonymous. Pages 113–228 (21 January 1996). Journal of Theoretical Biology, 178(2):??, January 21, 1996. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1996:PNb

[Ano96j] Anonymous. Pages 121–236 (21 November 1996). Journal of Theoretical Biology, 183(2):??, November 21, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1996:PJc

[Ano96k] Anonymous. Pages 181–270 (7 June 1996). Journal of Theoretical Biology, 180(3):??, June 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1996:PAa

[Ano96l] Anonymous. Pages 187–273 (7 April 1996). Journal of Theoretical Biology, 179(3):??, April 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1996:POa

[Ano96m] Anonymous. Pages 193–457 (7 October 1996). Journal of Theoretical Biology, 182(3):??, October 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1996:PAc

[Ano96n] Anonymous. Pages 197–298 (7 August 1996). Journal of Theoretical Biology, 181(3):??, August 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1996:PFa

[Ano96o] Anonymous. Pages 229–340 (7 February 1996). Journal of Theoretical Biology, 178(3):??, February 7, 1996. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1996:PDa

[Ano96p] Anonymous. Pages 237–349 (7 December 1996). Journal of Theoretical Biology, 183(3):??, December 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1996:PJd

[Ano96q] Anonymous. Pages 271–364 (21 June 1996). Journal of Theoretical Biology, 180(4):??, June 21, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1996:PAb

[Ano96r] Anonymous. Pages 275–368 (21 April 1996). Journal of Theoretical Biology, 179(4):??, April 21, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1996:PAd

[Ano96s] Anonymous. Pages 299–397 (21 August 1996). Journal of Theoretical Biology, 181(4):??, August 21, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1996:PFb

[Ano96t] Anonymous. Pages 341–426 (21 February 1996). Journal of Theoretical Biology, 178(4):??, February 21, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1996:PDb

[Ano96u] Anonymous. Pages 351–469 (21 December 1996). Journal of Theoretical Biology, 183(4):??, December 21, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1996:POb

[Ano96v] Anonymous. Pages 459–559 (21 October 1996). Journal of Theoretical Biology, 182(4):??, October 21, 1996. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1996:PMb

[Ano96w] Anonymous. Pages 91–186 (14 March 1996). Journal of Theoretical Biology, 179(2):??, March 14, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1996:PMd

[Ano96x] Anonymous. Pages 93–180 (21 May 1996). Journal of Theoretical Biology, 180(2):??, May 21, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1996:PJf

[Ano96y] Anonymous. Pages 97–196 (21 July 1996). Journal of Theoretical Biology, 181(2):??, July 21, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1996:PSb

[Ano96z] Anonymous. Pages 99–191 (21 September 1996). Journal of Theoretical Biology, 182(2):??, September 21, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1997:Ab

[Ano97a] Anonymous. Announcement. Journal of Theoretical Biology, 188(2):259–260, September 21, 1997. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519397905397.

Anonymous:1997:Aa

[Ano97b] Anonymous. Announcements. Journal of Theoretical Biology, 184(4):499–500, February 21, 1997. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519396903744.

Anonymous:1997:E

[Ano97c]  Anonymous.  Erratum.  Journal of Theoretical Biol-

ogy, 188(4):525–526, October 21, 1997. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519397905567.

Anonymous:1997:PJa

[Ano97d] Anonymous. Pages 1–100 (7 January 1997). Journal of Theoretical Biology, 184(1):??, January 7, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1997:PNa

[Ano97e] Anonymous. Pages 1–119 (7 November 1997). Journal of Theoretical Biology, 189(1):??, November 7, 1997. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1997:PMc

[Ano97f] Anonymous. Pages 1–129 (7 May 1997). Journal of Theoretical Biology, 186(1):??, May 7, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1997:PMa

[Ano97g] Anonymous. Pages 1–138 (7 March 1997). Journal of Theoretical Biology, 185(1):??, March 7, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1997:PSa

[Ano97h] Anonymous. Pages 1–142 (7 September 1997). Journal of Theoretical Biology, 188(1):??, September 7, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1997:PJe

[Ano97i] Anonymous. Pages 1–146 (7 July 1997). Journal of Theoretical Biology, 187(1):??, July 7, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1997:PJb

[Ano97j] Anonymous. Pages 101–226 (21 January 1997). Journal of Theoretical Biology, 184(2):??, January 21, 1997. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1997:PNb

[Ano97k] Anonymous. Pages 121–225 (21 November 1997). Journal of Theoretical Biology, 189(2):??, November 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1997:PMd

[Ano97l] Anonymous. Pages 131–266 (21 May 1997). Journal of Theoretical Biology, 186(2):??, May 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1997:PMb

[Ano97m] Anonymous. Pages 139–280 (21 March 1997). Journal of Theoretical Biology, 185(2):??, March 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1997:PSb

[Ano97n] Anonymous. Pages 143–260 (21 September 1997). Journal of Theoretical Biology, 188(2):??, September 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1997:PJf

[Ano97o] Anonymous. Pages 147–296 (21 July 1997). Journal of Theoretical Biology, 187(2):??, July 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1997:PDa

[Ano97p] Anonymous. Pages 227–331 (7 December 1997). Journal of Theoretical Biology, 189(3):??, December 7, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1997:PFa

[Ano97q] Anonymous. Pages 229–358 (7 February 1997). Journal of Theoretical Biology, 184(3):??, February 7, 1997. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1997:POa

[Ano97r] Anonymous. Pages 261–389 (7 October 1997). Journal of Theoretical Biology, 188(3):??, October 7, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1997:PJc

[Ano97s] Anonymous. Pages 267–396 (7 June 1997). Journal of Theoretical Biology, 186(3):??, June 7, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1997:PAa

[Ano97t] Anonymous. Pages 281–413 (7 April 1997). Journal of Theoretical Biology, 185(3):??, April 7, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1997:PAc

[Ano97u] Anonymous. Pages 297–451 (7 August 1997). Journal of Theoretical Biology, 187(3):??, August 7, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1997:PDb

[Ano97v] Anonymous. Pages 333–452 (21 December 1997). Journal of Theoretical Biology, 189(4):??, December 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1997:PFb

[Ano97w] Anonymous. Pages 359–500 (21 February 1997). Journal of Theoretical Biology, 184(4):??, February 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1997:POb

[Ano97x] Anonymous. Pages 391–526 (21 October 1997). Journal of Theoretical Biology, 188(4):??, October 21, 1997. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1997:PJd

[Ano97y] Anonymous. Pages 397–534 (21 June 1997). Journal of Theoretical Biology, 186(4):??, June 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1997:PAb

[Ano97z] Anonymous. Pages 415–555 (21 April 1997). Journal of Theoretical Biology, 185(4):??, April 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1997:PAd

[Ano97-27] Anonymous. Pages 453–629 (21 August 1997). Journal of Theoretical Biology, 187(4):??, August 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1997:SAS

[Ano97-28] Anonymous. Second annual Summer school on nonlinear dynamics in biology and medicine. Journal of Theoretical Biology, 184(2):225–226, January 21, 1997. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396903227.

Anonymous:1998:Aa

[Ano98a]  Anonymous.  Announcement.  Journal of Theoretical

Biology, 190(4):395, February 21, 1998. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519397906251.

Anonymous:1998:Ab

[Ano98b] Anonymous. Announcement. Journal of Theoretical Biology, 193(2):369, July 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519398907354.

Anonymous:1998:E

[Ano98c] Anonymous. Erratum. Journal of Theoretical Biology, 194 (4):605, October 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002251939890830X.

Anonymous:1998:LE

[Ano98d] Anonymous. Letter to the Editor. Journal of Theoretical Biology, 195(3):413–416, December 7, 1998. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519398906932.

Anonymous:1998:PJa

[Ano98e] Anonymous. Pages 1–106 (7 January 1998). Journal of Theoretical Biology, 190(1):??, January 7, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1998:PMa

[Ano98f] Anonymous. Pages 1–114 (7 March 1998). Journal of Theoretical Biology, 191(1):??, March 7, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1998:PMc

[Ano98g] Anonymous. Pages 1–130 (7 May 1998). Journal of Theoretical Biology, 192(1):??, May 7, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1998:PNa

[Ano98h] Anonymous. Pages 1–137 (7 November 1998). Journal of Theoretical Biology, 195(1):??, November 7, 1998. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1998:PSa

[Ano98i] Anonymous. Pages 1–161 (7 September 1998). Journal of Theoretical Biology, 194(1):??, September 7, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1998:PJe

[Ano98j] Anonymous. Pages 1–177 (7 July 1998). Journal of Theoretical Biology, 193(1):??, July 7, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1998:PJb

[Ano98k] Anonymous. Pages 107–199 (21 January 1998). Journal of Theoretical Biology, 190(2):??, January 21, 1998. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1998:PMb

[Ano98l] Anonymous. Pages 115–236 (21 March 1998). Journal of Theoretical Biology, 191(2):??, March 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1998:PMd

[Ano98m] Anonymous. Pages 131–263 (21 May 1998). Journal of Theoretical Biology, 192(2):??, May 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1998:PNb

[Ano98n] Anonymous. Pages 139–280 (21 November 1998). Journal of Theoretical Biology, 195(2):??, November 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1998:PSb

[Ano98o] Anonymous. Pages 163–311 (21 September 1998). Journal of Theoretical Biology, 194(2):??, September 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1998:PJf

[Ano98p] Anonymous. Pages 179–369 (21 July 1998). Journal of Theoretical Biology, 193(2):??, July 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1998:PFa

[Ano98q] Anonymous. Pages 201–293 (7 February 1998). Journal of Theoretical Biology, 190(3):??, February 7, 1998. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1998:PAa

[Ano98r] Anonymous. Pages 237–344 (7 April 1998). Journal of Theoretical Biology, 191(3):??, April 7, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1998:PJc

[Ano98s] Anonymous. Pages 265–413 (7 June 1998). Journal of Theoretical Biology, 192(3):??, June 7, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1998:PDa

[Ano98t] Anonymous. Pages 281–416 (7 December 1998). Journal of Theoretical Biology, 195(3):??, December 7, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1998:PFb

[Ano98u] Anonymous. Pages 295–395 (21 February 1998). Journal of Theoretical Biology, 190(4):??, February 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1998:POa

[Ano98v] Anonymous. Pages 313–460 (7 October 1998). Journal of Theoretical Biology, 194(3):??, October 7, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1998:PAb

[Ano98w] Anonymous. Pages 345–465 (21 April 1998). Journal of Theoretical Biology, 191(4):??, April 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1998:PAc

[Ano98x] Anonymous. Pages 371–541 (7 August 1998). Journal of Theoretical Biology, 193(3):??, August 7, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1998:PJd

[Ano98y] Anonymous. Pages 415–584 (21 June 1998). Journal of Theoretical Biology, 192(4):??, June 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1998:PDb

[Ano98z] Anonymous. Pages 419–542 (21 December 1998). Journal of Theoretical Biology, 195(4):??, December 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1998:POb

[Ano98-27] Anonymous. Pages 461–605 (21 October 1998). Journal of Theoretical Biology, 194(4):??, October 21, 1998. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1998:PAd

[Ano98-28] Anonymous. Pages 543–741 (21 August 1998). Journal of Theoretical Biology, 193(4):??, August 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1999:SSM

[Ano99a] Anonymous. 2nd Summer School of Mathematical Biology on Mathematics in Cell Physiology and Proliferation 6–19 June 1999 Termoli, CB, Italy. Journal of Theoretical Biology, 196(1):139, January 7, 1999. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519398908542.

Anonymous:1999:A

[Ano99b]  Anonymous.  Announcement.  Journal of Theoretical

Biology, 196(2):267, January 21, 1999. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519398908736.

Anonymous:1999:ASB

[Ano99c] Anonymous. Announcement — 19th Southern Biomedical Conference. Journal of Theoretical Biology, 201(2):157, November 21, 1999. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519399910330.

Anonymous:1999:Ea

[Ano99d] Anonymous. Erratum. Journal of Theoretical Biology, 198(1): 143, May 7, 1999. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251939999999X.

Anonymous:1999:Eb

[Ano99e] Anonymous. Erratum. Journal of Theoretical Biology, 198(3): 477, June 7, 1999. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519399909517.

Anonymous:1999:Ec

[Ano99f] Anonymous. Erratum. Journal of Theoretical Biology, 200 (1):119, September 7, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519399909980.

Anonymous:1999:PSa

[Ano99g] Anonymous. Pages 1–119 (7 September 1999). Journal of Theoretical Biology, 200(1):??, September 7, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1999:PJe

[Ano99h] Anonymous. Pages 1–123 (7 July 1999). Journal of Theoretical Biology, 199(1):??, July 7, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1999:PMa

[Ano99i] Anonymous. Pages 1–133 (7 March 1999). Journal of Theoretical Biology, 197(1):??, March 7, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1999:PJa

[Ano99j] Anonymous. Pages 1–139 (7 January 1999). Journal of Theoretical Biology, 196(1):??, January 7, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1999:PMc

[Ano99k] Anonymous. Pages 1–143 (7 May 1999). Journal of Theoretical Biology, 198(1):??, May 7, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1999:PNa

[Ano99l] Anonymous. Pages 1–92 (7 November 1999). Journal of Theoretical Biology, 201(1):??, November 7, 1999. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1999:PSb

[Ano99m] Anonymous. Pages 121–247 (21 September 1999). Journal of Theoretical Biology, 200(2):??, September 21, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1999:PJf

[Ano99n] Anonymous. Pages 125–233 (21 July 1999). Journal of Theoretical Biology, 199(2):??, July 21, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1999:PMb

[Ano99o] Anonymous. Pages 135–280 (21 March 1999). Journal of Theoretical Biology, 197(2):??, March 21, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1999:PJb

[Ano99p] Anonymous. Pages 141–267 (21 January 1999). Journal of Theoretical Biology, 196(2):??, January 21, 1999. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1999:PMd

[Ano99q] Anonymous. Pages 145–285 (21 May 1999). Journal of Theoretical Biology, 198(2):??, May 21, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1999:PDa

[Ano99r] Anonymous. Pages 159–218 (7 December 1999). Journal of Theoretical Biology, 201(3):??, December 7, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1999:PDb

[Ano99s] Anonymous. Pages 219–291 (21 December 1999). Journal of Theoretical Biology, 201(4):??, December 21, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1999:PAc

[Ano99t] Anonymous. Pages 235–358 (7 August 1999). Journal of Theoretical Biology, 199(3):??, August 7, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1999:POa

[Ano99u] Anonymous. Pages 249–344 (7 October 1999). Journal of Theoretical Biology, 200(3):??, October 7, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1999:PFa

[Ano99v] Anonymous. Pages 269–395 (7 February 1999). Journal of Theoretical Biology, 196(3):??, February 7, 1999. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1999:PAa

[Ano99w] Anonymous. Pages 281–423 (7 April 1999). Journal of Theoretical Biology, 197(3):??, April 7, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1999:PJc

[Ano99x] Anonymous. Pages 287–477 (7 June 1999). Journal of Theoretical Biology, 198(3):??, June 7, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1999:POb

[Ano99y] Anonymous. Pages 345–434 (21 October 1999). Journal of Theoretical Biology, 200(4):??, October 21, 1999. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1999:PAd

[Ano99z] Anonymous. Pages 359–471 (21 August 1999). Journal of Theoretical Biology, 199(4):??, August 21, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1999:PFb

[Ano99-27] Anonymous. Pages 397–525 (21 February 1999). Journal of Theoretical Biology, 196(4):??, February 21, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1999:PAb

[Ano99-28] Anonymous. Pages 425–567 (21 April 1999). Journal of Theoretical Biology, 197(4):??, April 21, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1999:PJd

[Ano99-29] Anonymous. Pages 479–602 (21 June 1999). Journal of Theoretical Biology, 198(4):??, June 21, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1999:PNb

[Ano99-30] Anonymous. Pages 93–157 (21 November 1999). Journal of Theoretical Biology, 201(2):??, November 21, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anand:1996:CPC

[AO96] Madhur Anand and Lászó Orlóci. Complexity in plant communities: the notion and quantification. Journal of Theoretical Biology, 179(2):179–186, March 14, 1996. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396900582.

Aoki:1990:EEC

[Aok90] Ichiro Aoki. Effects of exercise and chills on entropy production in human body. Journal of Theoretical Biology, 145(3):421–428, August 9, 1990. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305801201.

Aoki:1991:EPH

[Aok91] Ichiro Aoki. Entropy principle for human development, growth and aging. Journal of Theoretical Biology, 150(2):215–223, May 21, 1991. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305803339.

Aoki:1992:EPS

[Aok92] Ichiro Aoki. Entropy physiology of swine — a macroscopic viewpoint. Journal of Theoretical Biology, 157(3):363–371, August 7, 1992. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305806150.

Adams:1998:MML

[AOL98] James A. Adams, Geneva M. Omann, and Jennifer J. Linderman. A mathematical model for ligand/receptor/G-protein dynamics and actin polymerization in human neutrophils. Journal of Theoretical Biology, 193(4):543–560, August 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519398907214.

Ashida:1998:KAM

[AOU98] Hiroshi Ashida, Tadahiro Oonishi, and Nobuhiro Uyesaka. Kinetic analysis of the mechanism of action of the multidrug transporter. Journal of Theoretical Biology, 195(2):219–232, November 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519398907871.

Alerstam:1991:OAG

[AP91a] Thomas Alerstam and Sven-Göran Pettersson. Orientation along great circles by migrating birds using a sun compass. Journal of Theoretical Biology, 152(2):191–202, September 21, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305804527.

Almirantis:1991:CDE

[AP91b] Y. Almirantis and S. Papageorgiou. Cross-diffusion effects on chemical and biological pattern formation. Journal of Theoretical Biology, 151(3):289–311, August 7, 1991. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305803790.

Andreasen:1995:PCI

[AP95] Viggo Andreasen and Andrea Pugliese. Pathogen coexistence induced by density-dependent host mortality. Journal of Theoretical Biology, 177(2):159–165, November 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519319301729.

Aliev:1996:MHE

[AP96a] Rubin R. Aliev and Alexander V. Panfilov. Modeling of heart excitation patterns caused by a local inhomogeneity. Journal of Theoretical Biology, 181(1):33–40, July 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519396901125.

Andreadis:1996:KRM

[AP96b] Stylianos T. Andreadis and Bernhard O. Palsson. Kinetics of retrovirus mediated gene transfer: The importance of intracellular half-life of retroviruses. Journal of Theoretical Biology, 182(1):1–20, September 7, 1996. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251939690140X.

Auger:1996:EPG

[AP96c] Pierre Auger and Jean-Christophe Poggiale. Emergence of population growth models: Fast migration and slow growth. Journal of Theoretical Biology, 182(2):99–108, September 21, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519396901459.

Almirantis:1999:MMC

[AP99] Yannis Almirantis and Spyros Papageorgiou. Modes of morphogen cooperation for limb formation in vertebrates and insects. Journal of Theoretical Biology, 199(3):235–242, August 7, 1999. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519399909542.

Alpern:1999:SMH

[AR99] Steve Alpern and Diane Reyniers. Strategic mating with homotypic preferences. Journal of Theoretical Biology, 198(1):71– 88, May 7, 1999. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519399909037.

Almeida:1997:UKM

[ARC97] J. S. Almeida, M. A. M. Reis, and M. J. T. Carrondo. A unifying kinetic model of denitrification. Journal of Theoretical Biology, 186(2):241–249, May 21, 1997. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396903525.

Armstrong:1990:FMI

[Arm90] Robert A. Armstrong. A flexible model of incomplete dominance in neighborhood competition for space. Journal of Theoretical Biology, 144(3):287–302, June 7, 1990. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305800773.

Aflalo:1992:LPH

[AS92] Claude Aflalo and Lee A. Segel. Local probes and heterogeneous catalysis: a case study of a mitochondrialuciferase-hexokinase coupled system. Journal of Theoretical Biology, 158(1):67–108, September 7, 1992. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305806484.

Awerbuch:1995:TOT

[AS95] Tamara E. Awerbuch and Sonja Sandberg. Trends and oscillations in tick population dynamics. Journal of Theoretical Biology, 175(4):511–516, August 21, 1995. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385701580.

Scheuring:1996:WTT

[áSá96] Istv án Scheuring and Imre M. J. ánosi. When two and two make four: a structured population without chaos. Journal of Theoretical Biology, 178(1):89–97, January 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396900089.

Abramovich-Sivan:1998:PRC

[ASA98a] Sarit Abramovich-Sivan and Solange Akselrod. A phase response curve based model: Effect of vagal and sympathetic stimulation and interaction on a pacemaker cell. Journal of Theoretical Biology, 192(4):567–579, June 21, 1998. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519398906841.

Abramovich-Sivan:1998:PBM

[ASA98b] Sarit Abramovich-Sivan and Solange Akselrod. A PRC based model of a pacemaker cell: Effect of vagal activity and investigation of the respiratory sinus arrhythmia. Journal of Theoretical Biology, 192(2):219–234, May 21, 1998. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519398906580.

Alff-Steinberger:1994:CPT

[ASE94] C. Alff-Steinberger and R. Epstein. Codon preference in the terminal region of E. coli genes and evolution of stop codon usage. Journal of Theoretical Biology, 168(4):461–463, June 21, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519384711246.

Arrhenius:1997:ECM

[ASML97] G. Arrhenius, B. Sales, S. Mojzsis, and T. Lee. Entropy and charge in molecular evolution-the case of phosphate. Journal of Theoretical Biology, 187(4):503–522, August 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519396903859.

Atkinson:1991:SSP

[Atk91] David Atkinson. Sexual showiness and parasite load: Correlations without parasite coevolutionary cycles. Journal of Theoretical Biology, 150(2):251–260, May 21, 1991. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305803364.

Anatskaya:1994:HPM

[AVK94] Olga V. Anatskaya, Alexander E. Vinogradov, and Boris N. Kudryavtsev. Hepatocyte polyploidy and metabolism/lifehistory traits: Hypotheses testing. Journal of Theoretical Biology, 168(2):191–199, May 21, 1994. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519384710988.

Aalto:1999:TDS

[AVM^{+}99] Tuula Aalto, Timo Vesala, Timo Mattila, Pawel Simbierowicz, and Pertti Hari. A three-dimensional stomatal CO 2 exchange model including gaseous phase and leaf mesophyll separated by irregular interface. Journal of Theoretical Biology, 196(1):115–128, January 7, 1999. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519398908207.

Alexander:1993:ASD

[AW93] M. E. Alexander and J. R. Wickens. Analysis of striatal dynamics: The existence of two modes of behaviour. Journal of Theoretical Biology, 163(4):413–438, August 21, 1993. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519383711288.

Albe:1994:CMD

[AW94] Kathy R. Albe and Barbara E. Wright. Carbohydrate metabolism in Dictyostelium discoideum: II. systems’ analysis. Journal of Theoretical Biology, 169(3):243–251, August 7, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519384711453.

Austin:1998:DDA

[AWA98] D. J. Austin, N. J. White, and R. M. Anderson. The dynamics of drug action on the within-host population growth of infectious agents: Melding pharmacokinetics with pathogen population dynamics. Journal of Theoretical Biology, 194(3):313–339, October 7, 1998. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519397904380.

Beland:1994:OEG

[BA94] Pierre Béland and T. F. H. Allen. The origin and evolution of the genetic code. Journal of Theoretical Biology, 170(4):359–365, October 21, 1994. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519384711982.

Beer:1997:MGS

[BA97a] W. Nicholas Beer and James J. Anderson. Modelling the growth of salmonid embryos. Journal of Theoretical Biology, 189(3):297–306, December 7, 1997. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519397905154.

Bengtsson:1997:PGP

[BA97b] Bengt O. Bengtsson and Helene Andersson. The population genetics of plant mitochondrial plasmids. Journal of Theoretical Biology, 188(2):163–176, September 21, 1997. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251939790463X.

Bains:1993:LSS

[Bai93] William Bains. Local self-similarity of sequence in mammalian nuclear DNA is modulated by a 180 bp periodicity. Journal of Theoretical Biology, 161(2):137–143, March 21, 1993. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519383710465.

Bajpai:1999:CNR

[Baj99] R. P. Bajpai. Coherent nature of the radiation emitted in delayed luminescence of leaves. Journal of Theoretical Biology, 198(3):287–299, June 7, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519399908998.

Baker:1995:DTM

[Bak95] C. J. Baker. The development of a theoretical model for the windthrow of plants. Journal of Theoretical Biology, 175(3):355–372, August 7, 1995. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519385701476.

Baltscheffsky:1997:MAO

[Bal97] Herrick Baltscheffsky. Major “anastrophes” in the origin and early evolution of biological energy conversion. Journal of Theoretical Biology, 187(4):495–501, August 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396903847.

Barnsley:1990:SSK

[Bar90] Eric A. Barnsley. Steady state kinetics of enzymes with substrates having multiple reaction sites. Journal of Theoretical Biology, 142(2):275–279, January 23, 1990. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305802280.

Bartoloni:1992:MPA

[Bar92] N. Bartoloni. Markovian processes in aneuploids: I. Diploid populations. Journal of Theoretical Biology, 155(3):273–280, April 7, 1992. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305805995.

Barlow:1994:ESI

[Bar94] P. W. Barlow. Evolution of structural initial cells in apical meristems of plants. Journal of Theoretical Biology, 169(2):163–177, July 21, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519384711386.

Baranyi:1998:CSD

[Bar98]  József Baranyi.  Comparison of stochastic and deter-

ministic concepts of bacterial lag. Journal of Theoretical Biology, 192(3):403–408, June 7, 1998. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519398906737.

Bastarrachea:1998:OPB

[Bas98] Fernando Bastarrachea. On the origin of plasmid-borne, extended-spectrum, antibiotic resistance mutations in bacteria. Journal of Theoretical Biology, 190(4):379–387, February 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519397905634.

Batke:1991:CLE

[Bat91] József Batke. Channelling by loose enzyme complexes in situ is likely, though physiological significance is open for speculation. Journal of Theoretical Biology, 152(1):41–46, September 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305805065.

Bowers:1991:HHP

[BB91] Roger G. Bowers and Michael Begon. A host-host-pathogen model with free-living infective stages, applicable to microbial pest control. Journal of Theoretical Biology, 148(3):305–329, February 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305802401.

Bovet:1993:CSR

[BB93] Jacques Bovet and Pierre Bovet. Computer-simulations of rodent homing behavior, using a probabilistic model. Journal of Theoretical Biology, 161(2):145–156, March 21, 1993. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519383710477.

Begon:1994:HHP

[BB94] Michael Begon and Roger G. Bowers. Host-host-pathogen models and microbial pest control: The effect of host self regulation. Journal of Theoretical Biology, 169(3):275–287, August 7, 1994. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519384711489.

Bodnar:1996:CS

[BB96] John W. Bodnar and Margaret K. Bradley. A chromatin switch. Journal of Theoretical Biology, 183(1):1–7, November 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396901952.

Badge:1997:RHF

[BB97a] Richard M. Badge and John F. Y. Brookfield. The role of host factors in the population dynamics of selfish transposable elements. Journal of Theoretical Biology, 187(2):261–271, July 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251939790432X.

Beregovskaya:1997:RDA

[BB97b] Neonila Beregovskaya and Yuri Beregovski. Radiation damage accumulation over large time intervals: a descriptive model. Journal of Theoretical Biology, 189(1):81–88, November 7, 1997. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519397905002.

Bernstein:1997:SC

[BB97c] Carol Bernstein and Harris Bernstein. Sexual communication. Journal of Theoretical Biology, 188(1):69–78, September 7, 1997. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519397904598.

Bohl:1997:QAB

[BB97d] Erich Bohl and Winfried Boos. Quantitative analysis of binding protein-mediated ABC transport systems. Journal of Theoretical Biology, 186(1):65–74, May 7, 1997. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396903422.

Bazhan:1999:IIA

[BB99a] S. I. Bazhan and O. E. Belova. Interferon-induced antiviral resistance. a mathematical model of regulation of mx1 protein induction and action. Journal of Theoretical Biology, 198(3):375– 393, June 7, 1999. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519399909219.

Boots:1999:TMH

[BB99b] Michael Boots and Roger G. Bowers. Three mechanisms of host resistance to microparasites-avoidance, recovery and tolerance-show different evolutionary dynamics. Journal of Theoretical Biology, 201(1):13–23, November 7, 1999. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519399910093.

Brooker:1990:APG

[BBB90] L. C. Brooker, M. G. Brooker, and A. M. H. Brooker. An alternative population/genetics model for the evolution of egg mimesis and egg crypsis in cuckoos. Journal of Theoretical Biology, 146(1):123–143, September 7, 1990. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305800487.

Bacon:1991:AMG

[BBB91a] P. J. Bacon, Frank Ball, and Paul Blackwell. Analysis of a model of group territoriality based on the resource dispersion hypothesis. Journal of Theoretical Biology, 148(4):433–444, February 21, 1991. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305802309.

Bacon:1991:MTG

[BBB91b] P. J. Bacon, Frank Ball, and Paul Blackwell. A model for territory and group formation in a heterogeneous habitat. Journal of Theoretical Biology, 148(4):445–468, February 21, 1991. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305802310.

Bazykin:1997:DFI

[BBIK97] A. D. Bazykin, F. S. Berezovskaya, A. S. Isaev, and R. G. Khlebopros. Dynamics of forest insect density: Bifurcation approach. Journal of Theoretical Biology, 186(3):267–278, June 7, 1997. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251939690363X.

Baltscheffsky:1997:OEL

[BBL^{+}97a] H. Baltscheffsky, C. Blomberg, H. Liljenström, B. I. B. Lindahl, and P. Århem. On the origin and evolution of life: An introduction. Journal of Theoretical Biology, 187(4):453–459, August 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251939690380X.

Blasius:1997:MPO

[BBL97b] B. Blasius, F. Beck, and U. Lüttge. A model for photosynthetic oscillations in crassulacean acid metabolism (CAM). Journal of Theoretical Biology, 184(3):345–351, February 7, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519396902878.

Boldrini:1997:RNL

[BBM^{+}97] J. L. Boldrini, R. C. Bassanezi, A. C. Moretti, F. J. Von Zuben, W. A. C. Godoy, C. J. Von Zuben, and S. F. dos Reis. RETRACTED: Non-local interactions and the dynamics of dispersal in immature insects. Journal of Theoretical Biology, 185(4):523–531, April 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S0022519396903203. See [BBM^{+}13c, BBM^{+}13a, BBM^{+}13b].

Boldrini:2013:CNLa

[BBM^{+}13a] J. L. Boldrini, R. C. Bassanezi, A. C. Moretti, F. J. Von Zuben, W. A. C. Godoy, C. J. Von Zuben, and S. F. dos Reis. Corrigendum to “Non-local interactions and the dynamics of dispersal in immature insects” [J. Theor. Biol. 185 (4) (1997) 523– 531]. Journal of Theoretical Biology, 336(??):251, November 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519313003445. See [BBM^{+}97,

BBM^{+}13c].

Boldrini:2013:CNLb

[BBM^{+}13b] J. L. Boldrini, R. C. Bassanezi, A. C. Moretti, F. J. Von Zuben, W. A. C. Godoy, C. J. Von Zuben, and S. F. dos Reis. Corrigendum to “Non-local interactions and the dynamics of dispersal in immature insects” [J. Theor. Biol. 185 (4) (1997) 523– 531]. Journal of Theoretical Biology, 336(??):252–258, November 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519313003433. See [BBM^{+}97].

Boldrini:2013:RRN

[BBM^{+}13c] J. L. Boldrini, R. C. Bassanezi, A. C. Moretti, F. J. Von Zuben, W. A. C. Godoy, C. J. Von Zuben, and S. F. dos Reis. RETRACTED: Retraction notice to “non-local interactions and the dynamics of dispersal in immature insects”. Journal of Theoretical Biology, 336(??):250, November 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519313000441. See [BBM^{+}97].

Benhamou:1995:MPN

[BBP95] Simon Benhamou, Pierre Bovet, and Bruno Poucet. A model for place navigation in mammals. Journal of Theoretical Biology, 173(2):163–178, March 21, 1995. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385700537.

Baconnier:1993:ERR

[BBPD93] P. F. Baconnier, G. Benchetrit, P. Pachot, and J. Demongeot. Entrainment of the respiratory rhythm: a new approach. Journal of Theoretical Biology, 164(2):149–162, September 21, 1993. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519383711458.

Baker:1998:MAR

[BBS^{+}98] C. J. Baker, P. M. Berry, J. H. Spink, R. Sylvester-Bradley, J. M. Griffin, R. K. Scott, and R. W. Clare. A method for the assessment of the risk of wheat lodging. Journal of Theoretical Biology, 194(4):587–603, October 21, 1998. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519398907780.

Baltz:1990:SNC

[BC90] Jay M. Baltz and Richard A. Cone. The strength of noncovalent biological bonds and adhesions by multiple independent bonds. Journal of Theoretical Biology, 142(2):163–178, January 23, 1990. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251930580219X.

Blickhan:1994:ESE

[BC94] Reinhard Blickhan and Jian-Yu Cheng. Energy storage by elastic mechanisms in the tail of large swimmers — a reevaluation. Journal of Theoretical Biology, 168(3):315–321, June 7, 1994. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251938471112X.

Bersini:1997:FCB

[BC97] Hugues Bersini and Vera Calenbuhr. Frustrated chaos in biological networks. Journal of Theoretical Biology, 188(2): 187–200, September 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519397904689.

Bois:1992:SAN

[BCQ92] Frédéric Y. Bois and Penelope J. E. Compton-Quintana. Sensitivity analysis of a new model of carcinogenesis. Journal of Theoretical Biology, 159(3):361–375, December 7, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305807301.

Broom:1997:SAM

[BCV97] M. Broom, C. Cannings, and G. T. Vickers. A sequentialarrivals model of territory acquisition. Journal of Theoretical Biology, 189(3):257–272, December 7, 1997. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519397905191.

Bardakjian:1994:MCO

[BD94] Berj L. Bardakjian and Nicholas E. Diamant. A mapped clock oscillator model for transmembrane electrical rhythmic activity in excitable cells. Journal of Theoretical Biology, 166(3):225–235, February 7, 1994. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519384710204.

Bermudez:1999:CCE

[BDA99] Clara I. Bermúdez, Edgar E. Daza, and Eugenio Andrade. Characterization and comparison of Escherichia coli transfer RNAs by graph theory based on secondary structure. Journal of Theoretical Biology, 197(2):193–205, March 21, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519398908669.

Bannink:1997:IDS

[BDDF97] A. Bannink, H. De Visser, J. Dijkstra, and J. France. Impact of diet-specific input parameters on simulated rumen function. Journal of Theoretical Biology, 184(4):371–384, February 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396902507.

Beckers:1992:TUT

[BDG92] R. Beckers, J. L. Deneubourg, and S. Goss. Trails and uturns in the selection of a path by the ant lasius niger. Journal of Theoretical Biology, 159(4):397–415, December 21, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305806861.

Boudreau:1991:BSS

[BDK91] P. R. Boudreau, L. M. Dickie, and S. R. Kerr. Body-size spectra of production and biomass as system-level indicators of ecological dynamics. Journal of Theoretical Biology, 152(3):329–339, October 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305801985.

Bannink:1997:CIP

[BDK^{+}97] A. Bannink, H. De Visser, A. Klop, J. Dijkstra, and J. France. Causes of inacurate prediction of volatile fatty acids by simulation models of rumen function in lactating cows. Journal of Theoretical Biology, 189(4):353–366, December 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519397905105.

Bernardes:1997:INE

[BdS97] Américo T. Bernardes and Rita M. Zorzenon dos Santos. Immune network at the edge of chaos. Journal of Theoretical Biology, 186(2):173–187, May 21, 1997. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396903161.

Berry:1992:CLB

[BE92] R. Berry and D. T. Edmonds. Carrier-like behaviour from a static but electrically responsive model pore. Journal of Theoretical Biology, 154(2):249–260, January 21, 1992. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305804060.

Beanland:1990:ERB

[Bea90] Timothy J. Beanland. Evolutionary relationships between “qtype” photosynthetic reaction centres: Hypothesis-testing using parsimony. Journal of Theoretical Biology, 145(4):535–545, August 23, 1990. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305804874.

Beaumont:1993:EOB

[Bea93] Mark A. Beaumont. Evolution of optimal behaviour in networks of Boolean automata. Journal of Theoretical Biology, 165(4):455–476, December 21, 1993. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519383712014.

Beck:1996:MBI

[Bec96] Friedrich Beck. Mind-brain interaction: Comments on an article by B. I. B. Lindahl and P. Århem. Journal of Theoretical Biology, 180(1):87–89, May 7, 1996. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251939690082X.

Bednekoff:1996:CSW

[Bed96] Peter A. Bednekoff. The cost of sex when males care, but can’t discriminate. Journal of Theoretical Biology, 179(3):187– 189, April 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396900594.

Balshine-Earn:1997:EMP

[BEE97] Sigal Balshine-Earn and David J. D. Earn. An evolutionary model of parental care in St. Peter’s fish. Journal of Theoretical Biology, 184(4):423–431, February 21, 1997. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396902544.

Benhamou:1992:EAC

[Ben92] Simon Benhamou. Efficiency of area-concentrated searching behaviour in a continuous patchy environment. Journal of Theoretical Biology, 159(1):67–81, November 7, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305807684.

Bengough:1997:MRD

[Ben97] A. G. Bengough. Modelling rooting depth and soil strength in a drying soil profile. Journal of Theoretical Biology, 186(3):327– 338, June 7, 1997. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396903677.

Bennekou:1999:FPV

[Ben99] Poul Bennekou. The feasibility of pharmacological volume control of sickle cells is dependent on the quantization of the transport pathways. a model study. Journal of Theoretical Biology, 196(1):129–137, January 7, 1999. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519398908190.

Berger:1991:MSF

[Ber91] David S. Berger. Modification of a simple fractal tree growth scheme: Implications on growth, variation, and evolution. Journal of Theoretical Biology, 152(4):513–529, October 21, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305803959.

Berezin:1992:CIT

[Ber92] Alexander A. Berezin. Correlated isotopic tunneling as a possible model for consciousness. Journal of Theoretical Biology, 154(4):415–420, February 21, 1992. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305801602.

Bergner:1993:KCP

[Ber93] Per-Erik E. Bergner. On kinetics of classic particles: Mean values of state functions. Journal of Theoretical Biology, 165(3):357–372, December 7, 1993. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251938371194X.

Berg:1995:PSH

[Ber95a] Otto G. Berg. Periodic selection and hitchhiking in a bacterial population. Journal of Theoretical Biology, 173(3):307–320, April 7, 1995. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519385700641.

Berntson:1995:CTC

[Ber95b] G. M. Berntson. The characterization of topology: a comparison of four topological indices for rooted binary trees. Journal of Theoretical Biology, 177(3):271–281, December 7, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519385702445.

Berry:1995:EIE

[Ber95c] Stephan Berry. Entropy, irreversibility and evolution. Journal of Theoretical Biology, 175(2):197–202, July 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385701324.

Bettelheim:1999:SRP

[Bet99] Frederick A. Bettelheim. Syneretic response to pressure in ocular lens. Journal of Theoretical Biology, 197(2):277–280, March 21, 1999. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519398908554.

Beuchat:1990:MSU

[Beu90] Carol A. Beuchat. Metabolism and the scaling of urine concentrating ability in mammals: Resolution of a paradox? Journal of Theoretical Biology, 143(1):113–122, March 8, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305802917.

Beecham:1999:AGF

[BF99a] J. A. Beecham and K. D. Farnsworth. Animal group forces resulting from predator avoidance and competition minimization. Journal of Theoretical Biology, 198(4):533–548, June 21, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251939990930X.

Bell:1999:AUD

[BF99b] S. J. Bell and D. R. Forsdyke. Accounting units in DNA. Journal of Theoretical Biology, 197(1):51–61, March 7, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519398908578.

Bell:1999:DCS

[BF99c] S. J. Bell and D. R. Forsdyke. Deviations from Chargaff’s second parity rule correlate with direction of transcription. Journal of Theoretical Biology, 197(1):63–76, March 7, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251939890858X.

Bennett:1996:QTR

[BFG96] M. R. Bennett, L. Farnell, and W. G. Gibson. Quantal transmitter release onto different patterns of receptor distributions at somatic neuromuscular junctions. Journal of Theoretical Biology, 183(2):231–236, November 21, 1996. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396902167.

Bruni:1993:SMC

[BFK^{+}93] C. Bruni, L. Ferrante, G. Koch, C. Scoglio, and G. Starace. A stochastic model for cell debris in flow cytometry. Journal of Theoretical Biology, 161(2):157–174, March 21, 1993. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519383710489.

Bagley:1996:CCA

[BG96] R. J. Bagley and Leon Glass. Counting and classifying attractors in high dimensional dynamical systems. Journal of Theoretical Biology, 183(3):269–284, December 7, 1996. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396902209.

Buschmann:1997:MMO

[BG97] P. Buschmann and D. Gradmann. Minimal model for oscillations of membrane voltage in plant cells. Journal of Theoretical Biology, 188(3):323–332, October 7, 1997. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519397904835.

Benci:1998:SIE

[BG98] Vieri Benci and Lodovico Galleni. Stability and instability in evolution. Journal of Theoretical Biology, 194(4):541–549, October 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519398907767.

Brass:1994:CAM

[BGE94] Andy Brass, Richard K. Grencis, and Kathryn J. Else. A cellular automata model for helper T cell subset polarization in chronic and acute infection. Journal of Theoretical Biology, 166(2):189–200, January 21, 1994. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384710174.

Boulton:1999:DMM

[BGS^{+}99] I. C. Boulton, A. R. Gorringe, J. K. Shergill, C. L. Joannou, and R. W. Evans. A dynamic model of the meningococcal transferrin receptor. Journal of Theoretical Biology, 198(4):497–505, June 21, 1999. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519399909281.

Buxbaum:1992:ART

[BH92] R. E. Buxbaum and S. R. Heidemann. An absolute rate theory model for tension control of axonal elongation. Journal of Theoretical Biology, 155(4):409–426, April 21, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305806265.

Biktashev:1994:DPL

[BH94] V. N. Biktashev and A. V. Holden. Design principles of a low voltage cardiac defibrillator based on the effect of feedback resonant drift. Journal of Theoretical Biology, 169(2):101–112, July 21, 1994. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519384711325.

Boerlijst:1995:ASP

[BH95] Maarten C. Boerlijst and Pauline Hogeweg. Attractors and spatial patterns in hypercycles with negative interactions. Journal of Theoretical Biology, 176(2):199–210, September 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519385701919.

Ben-Haim:1992:ICV

[BHP92] Shlomo A. Ben-Haim and Yoram Palti. Intercellular conduction velocity variability as the basis for re-entrant arrhythmias in the ischemic myocardium. Journal of Theoretical Biology, 154(3):317–330, February 7, 1992. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305801730.

Brunden:1990:PPP

[BHVC90] Marshall N. Brunden, Rita M. Huff, Thomas J. Vidmar, and Murray M. Cooper. Planning the purification process of active cDNA in expression cloning strategies. Journal of Theoretical Biology, 144(2):145–154, May 14, 1990. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305803157.

Bignone:1993:CGI

[Big93] Franco A. Bignone. Cells-gene interactions simulation on a coupled map lattice. Journal of Theoretical Biology, 161(2):231–249, March 21, 1993. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519383710520.

Bassom:1997:ODT

[BIV97] Andrew P. Bassom, Achim Ilchmann, and Heinrich Voß. Oxygen diffusion in tissue preparations with Michaelis–Menten kinetics. Journal of Theoretical Biology, 185(1):119–127, March 7, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519396902982.

Barabe:1996:CGF

[BJ96] Denis Barabé and Roger V. Jean. The constraints of global form on phyllotactic organization: the case of symplocarpus (araceae). Journal of Theoretical Biology, 178(4):393–397, February 21, 1996. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396900363.

Bjerknes:1994:SST

[Bje94] Matthew Bjerknes. Simple stochastic theory of stem cell differentiation is not simultaneously consistent with crypt extinction probability and the expansion of mutated clones. Journal of Theoretical Biology, 168(4):349–365, June 21, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384711155.

Bjerknes:1996:EMS

[Bje96] Matthew Bjerknes. Expansion of mutant stem cell populations in the human colon. Journal of Theoretical Biology, 178(4):381–385, February 21, 1996. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251939690034X.

Bjorklund:1994:SSO

[Bjö94] Mats Björklund. Species selection on organismal integration. Journal of Theoretical Biology, 171(4):427–430, December 21, 1994. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251938471246X.

Baczkowski:1997:PGD

[BJS97] A. J. Baczkowski, D. N. Joanes, and G. M. Shamia. Properties of a generalized diversity index. Journal of Theoretical Biology, 188(2):207–213, September 21, 1997. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519397904677.

Bulsara:1991:SRS

[BJZ^{+}91] Adi Bulsara, E. W. Jacobs, Ting Zhou, Frank Moss, and Laszlo Kiss. Stochastic resonance in a single neuron model: Theory and analog simulation. Journal of Theoretical Biology, 152(4):531–555, October 21, 1991. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305803960.

Batt:1990:VIR

[BK90a] Brian C. Batt and Dhinakar S. Kompala. Verification of immune response optimality through cybernetic modeling. Journal of Theoretical Biology, 142(3):317–340, February 9, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305805569.

Beauchamp:1990:FFR

[BK90b] Guy Beauchamp and Alejandro Kacelnik. On the fitness functions relating parental care to reproductive value. Journal of Theoretical Biology, 146(4):513–522, October 21, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305803765.

Batali:1995:EAO

[BK95] John Batali and Philip Kitcher. Evolution of altriusm in optional and compulsory games. Journal of Theoretical Biology, 175(2):161–171, July 21, 1995. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519385701282.

Blanchard:1997:HAH

[BK97] Ray Blanchard and Philip Klassen. H-y antigen and homosexuality in men. Journal of Theoretical Biology, 185(3):373–378, April 7, 1997. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251939690315X.

Bat:1997:CSE

[BKA97] O. Bat, M. Kimmel, and D. E. Axelrod. Computer simulation of expansions of DNA triplet repeats in the fragile X syndrome and Huntington’s disease. Journal of Theoretical Biology, 188(1):53–67, September 7, 1997. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519397904513.

Brauchli:1999:ECS

[BKD99] Kurt Brauchli, Timothy Killingback, and Michael Doebeli. Evolution of cooperation in spatially structured populations. Journal of Theoretical Biology, 200(4):405–417, October 21, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519399910007.

Berezetskaya:1996:MMP

[BKK96] Natalia M. Berezetskaya, Valery N. Kharkyanen, and Nikolai I. Kononenko. Mathematical model of pacemaker activity in bursting neurons of snail, helix pomatia. Journal of Theoretical Biology, 183(2):207–218, November 21, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519396902143.

Bodnar:1997:DLG

[BKNR97] John W. Bodnar, Jeffrey Killian, Michael Nagle, and Suneil Ramchandani. Deciphering the language of the genome. Journal of Theoretical Biology, 189(2):183–193, November 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519397905075.

Beloussov:1991:GRM

[BL91] L. V. Beloussov and A. V. Lakirev. Generative rules for the morphogenesis of epithelial tubes. Journal of Theoretical Biology, 152(4):455–468, October 21, 1991. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305803923.

Bagnoli:1995:SMC

[BL95] Franco Bagnoli and Pietro Liò. Selection, mutations and codon usage in a bacterial model. Journal of Theoretical Biology, 173(3):271–281, April 7, 1995. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519385700628.

Blanco:1992:MPF

[Bla92] Luciano N. Blanco. A model of postnatal formation of alveoli in rat lung. Journal of Theoretical Biology, 157(4):427–446, August 21, 1992. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305806629.

Blank:1994:PAR

[Bla94] Martin Blank. Protein aggregation reactions: Surface free energy model. Journal of Theoretical Biology, 169(4):323–326, August 21, 1994. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519384711544.

Blackburn:1995:SPE

[Bla95] Daniel G. Blackburn. Saltationist and punctuated equilibrium models for the evolution of viviparity and placentation. Journal of Theoretical Biology, 174(2):199–216, May 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519385700926.

Blackwell:1997:PWA

[Bla97] P. G. Blackwell. The n-player war of attrition and territorial groups. Journal of Theoretical Biology, 189(2):175–181, November 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251939790504X.

Blackburn:1998:REV

[Bla98] Daniel G. Blackburn. Reconstructing the evolution of viviparity and placentation. Journal of Theoretical Biology, 192 (2):183–190, May 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519397905920.

Bazhan:1995:TAR

[BLB95] S. I. Bazhan, V. A. Likhoshvai, and O. E. Belova. Theoretical analysis of the regulation of interferon expression during priming and blocking. Journal of Theoretical Biology, 175(2):149– 160, July 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519385701270.

Barbet:1993:CCM

[BLGG^{+}93] Jacques Barbet, Jean-Marc Le Doussal, Anne Gruaz-Guyon, Marie Martin, Emmanuel Gautherot, and Michel Delaage. Computer calculation of multiple binding equilibrium isotherms: Application to the binding of bivalent ligands to antibodies interacting with cell surface fc-receptors. Journal of Theoretical Biology, 165(3):321–340, December 7, 1993. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519383711926.

Blomberg:1994:MME

[BLLÅ94] C. Blomberg, H. Liljenström, B. I. B. Lindahl, and P. Århem. Mind and matter: Essays from biology, physics and philosophy: An introduction. Journal of Theoretical Biology, 171(1): 1–5, July 1994. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519384712057.

Bystrov:1994:FAM

[BLM94] Vladimir S. Bystrov, Victor D. Lakhno, and M. Molchanov. Ferroelectric active models of ion channels in biomembranes. Journal of Theoretical Biology, 168(4):383–393, June 21, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384711180.

Blomberg:1994:PRT

[Blo94] Clas Blomberg. The Physicist’s road to theoretical biology and the mind-matter problem. Journal of Theoretical Biology, 171(1):41–52, July 1994. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519384712100.

Blomberg:1997:AFO

[Blo97] Clas Blomberg. On the appearance of function and organisation in the origin of life. Journal of Theoretical Biology, 187(4):541–554, August 21, 1997. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519396903884.

Boyer:1990:DCS

[BM90] John F. Boyer and David B. Mertz. The dynamics of cohort selection. Journal of Theoretical Biology, 142(3):393–414, February 9, 1990. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305805594.

Beregovskaya:1995:MDD

[BM95] Neonila Beregovskaya and Rostislav Maiboroda. Mitochondrial DNA damage and efficiency of ATP biosynthesis: mathematical model. Journal of Theoretical Biology, 172(2):161–168, January 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519385700136.

Bryant:1998:OFC

[BM98] Michael R. Bryant and Peter J. McDonnell. Optical feedbackcontrolled scleral remodeling as a mechanism for myopic eye growth. Journal of Theoretical Biology, 193(4):613–622, August 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519398907275.

Baeten:1998:QSC

[BMG98] Anja Baeten, Dominique Maes, and Paul Geerlings. Quantumchemical study of the catalytic triad in subtilisin: the influence of amino acid substitutions on enzymatic activity. Journal of Theoretical Biology, 195(1):27–40, November 7, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251939890769X.

Birkhead:1993:WDF

[BMS93] T. R. Birkhead, A. P. Møller, and W. J. Sutherland. Why do females make it so difficult for males to fertilize their eggs? Journal of Theoretical Biology, 161(1):51–60, March 7, 1993. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519383710398.

Barvik:1992:SRA

[BN92a] Ivan Barvik and Ladislav Nedbal. Some remarks about the role of a back recombination in RC primary processes. Journal of Theoretical Biology, 154(3):303–315, February 7, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305801729.

Bishir:1992:DDD

[BN92b] John Bishir and Gene Namkoong. Density-dependent dynamics in size-structured plant populations. Journal of Theoretical Biology, 154(2):163–188, January 21, 1992. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305804011.

Berry:1994:DEM

[BN94] Richard M. Berry and Martin A. Nowak. Defective escape mutants of HIV. Journal of Theoretical Biology, 171(4):387–395, December 21, 1994. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519384712422.

Behera:1995:IRP

[BN95] Narayan Behera and Vidyanand Nanjundiah. An investigation into the role of phenotypic plasticity in evolution. Journal of Theoretical Biology, 172(3):225–234, February 7, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519385700197.

Behera:1996:CPP

[BN96] Narayan Behera and Vidyanand Nanjundiah. The consequences of phenotypic plasticity in cyclically varying environments: a genetic algorithm study. Journal of Theoretical Biology, 178(2):135–144, January 21, 1996. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396900120.

Behera:1997:TGR

[BN97] Narayan Behera and Vidyanand Nanjundiah. trans gene regulation in adaptive evolution: a genetic algorithm model. Journal of Theoretical Biology, 188(2):153–162, September 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519397904616.

Boerlijst:1997:LC

[BNS97] Maarten C. Boerlijst, Martin A. Nowak, and Karl Sigmund. The logic of contrition. Journal of Theoretical Biology, 185(3):281–293, April 7, 1997. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519396903264.

Bocharov:1998:MDL

[Boc98] G. A. Bocharov. Modelling the dynamics of LCMV infection in mice: Conventional and exhaustive CTL responses. Journal of Theoretical Biology, 192(3):283–308, June 7, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519397906123.

Bodnar:1997:PDE

[Bod97] John W. Bodnar. Programming the Drosophila embryo. Journal of Theoretical Biology, 188(4):391–445, October 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519396903288.

Boettcher:1990:TIN

[Boe90] Barry Boettcher. Transcription initiation and nucleasesensitive sites upstream of the ϵ-globin gene in k562 cells are related to poly (dA). poly (dT) sequences. Journal of Theoretical Biology, 146(3):333–339, October 7, 1990. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930580743X.

Brachet:1999:DMV

[BOG^{+}99] Stephanie Brachet, Isabelle Olivieri, Bernard Godelle, Etienne Klein, Nathalie Frascaria-Lacoste, and Pierre-Henri Gouyon. Dispersal and metapopulation viability in a heterogeneous landscape. Journal of Theoretical Biology, 198(4):479–495, June 21, 1999. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519399909268.

Boily:1995:THF

[Boi95] Patrice Boily. Theoretical heat flux in water and habitat selection of phocid seals and beluga whales during the annual molt. Journal of Theoretical Biology, 172(3):235–244, February 7, 1995. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519385700203.

Bonnaz:1999:ATS

[Bon99] D. Bonnaz. About a three states molecular evolution model. Journal of Theoretical Biology, 197(3):415–423, April 7, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519398908839.

Bar-Or:1998:RPD

[BOS98] Ruth Lev Bar-Or and Lee A. Segel. On the role of a possible dialogue between cytokine and TCR-presentation mechanisms in the regulation of autoimmune disease. Journal of Theoretical Biology, 190(2):161–178, January 21, 1998. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519397905452.

Bourke:1994:WMS

[Bou94] Andrew F. G. Bourke. Worker matricide in social bees and wasps. Journal of Theoretical Biology, 167(3):283–292, April 7, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519384710708.

Bowers:1999:BMA

[Bow99] Roger G. Bowers. A baseline model for the apparent competition between many host strains: the evolution of host resistance to microparasites. Journal of Theoretical Biology, 200(1):65–75, September 7, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519399909761.

Boyle:1995:VMS

[Boy95] Michael D. P. Boyle. Variation of multifunctional surface binding proteins — a virulence strategy for group A streptococci? Journal of Theoretical Biology, 173(4):415–426, April 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519385700732.

Buchman:1992:SAD

[BP92] E. Buchman and H. Parnas. Sequential approach to describe the time course of synaptic channel opening under constant transmitter concentration. Journal of Theoretical Biology, 158(4):517–534, October 21, 1992. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305807131.

Bomze:1993:DES

[BP93] I. M. Bomze and B. M. Pötscher. “On the Definition of an Evolutionarily Stable Strategy in the Playing the Field Model” by V. P. Crawford. Journal of Theoretical Biology, 161(3):405, April 7, 1993. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519383710647.

Buchman:1994:SAD

[BP94] E. Buchman and H. Parnas. Sequential approach for describing channel opening and desensitization. Journal of Theoretical Biology, 167(4):381–395, April 21, 1994. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384710770.

Ball:1996:SCG

[BP96] M. A. Ball and G. A. Parker. Sperm competition games: External fertilization and “adapative” infertility. Journal of Theoretical Biology, 180(2):141–150, May 21, 1996. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396900909.

Ball:1997:SCG

[BP97a] M. A. Ball and G. A. Parker. Sperm competition games: Inter- and intra-species results of a continuous external fertilization model. Journal of Theoretical Biology, 186(4):459–466, June 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519397904069.

Bastolla:1997:NSC

[BP97b] U. Bastolla and G. Parisi. A numerical study of the critical line of Kauffman networks. Journal of Theoretical Biology, 187(1):117–133, July 7, 1997. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519397904239.

Ball:1998:SCG

[BP98] M. A. Ball and G. A. Parker. Sperm competition games: a general approach to risk assessment. Journal of Theoretical Biology, 194(2):251–262, September 21, 1998. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519398907561.

Berenfeld:1999:DIS

[BP99a] Omer Berenfeld and Arkady M. Pertsov. Dynamics of intramural scroll waves in three-dimensional continuous myocardium with rotational anisotropy. Journal of Theoretical Biology, 199(4):383–394, August 21, 1999. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519399909657.

Best:1999:MUT

[BP99b] Michael L. Best and Richard Pocklington. Meaning as use: Transmission fidelity and evolution in NetNews. Journal of Theoretical Biology, 196(3):389–395, February 7, 1999. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519398908505.

Borstnik:1990:EAE

[BPHH90] B. Borstnik, D. Pumpernik, I. L. Hofacker, and G. L. Hofacker. An ESS-analysis for ensembles of Prisoner’s dilemma strategies. Journal of Theoretical Biology, 142(2):189–200, January 23, 1990. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305802218.

Boyd:1990:GSA

[BR90] Robert Boyd and Peter J. Richerson. Group selection among alternative evolutionarily stable strategies. Journal of Theoretical Biology, 145(3):331–342, August 9, 1990. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305801134.

Bull:1991:DME

[BR91] J. J. Bull and W. R. Rice. Distinguishing mechanisms for the evolution of co-operation. Journal of Theoretical Biology, 149(1):63–74, March 7, 1991. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305800724.

Bocharov:1994:MMA

[BR94] G. A. Bocharov and A. A. Romanyukha. Mathematical model of antiviral immune response III. influenza a virus infection. Journal of Theoretical Biology, 167(4):323–360, April 21, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519384710745.

Brand:1990:CLP

[Bra90a] Martin D. Brand. The contribution of the leak of protons across the mitochondrial inner membrane to standard metabolic rate. Journal of Theoretical Biology, 145(2):267– 286, July 24, 1990. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305801316.

Bray:1990:ISP

[Bra90b] D. Bray. Intracellular signalling as a parallel distributed process. Journal of Theoretical Biology, 143(2):215–231, March 22, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305802681.

Brand:1996:TMC

[Bra96] Martin D. Brand. Top down metabolic control analysis. Journal of Theoretical Biology, 182(3):351–360, October 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519396901745.

Britt:1993:SSM

[Bri93] Billy Mark Britt. A shifting specificity model for enzyme catalysis. Journal of Theoretical Biology, 164(2):181–190, September 21, 1993. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519383711471.

Brown:1990:PCB

[Bro90] Robert D. Brown. The proton channel blocking agent omeprazole is an inhibitor of the thiamin shuttle. Journal of Theoretical Biology, 143(4):565–573, April 19, 1990. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930580030X.

Brown:1991:TCP

[Bro91] Guy C. Brown. Total cell protein concentration as an evolutionary constraint on the metabolic control distribution in cells. Journal of Theoretical Biology, 153(2):195–203, November 21, 1991. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305804229.

Brookfield:1997:HTE

[Bro97] J. F. Y. Brookfield. Hypothesis testing in evolutionary inference. Journal of Theoretical Biology, 185(4):533–538, April 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519396903367.

Bascompte:1998:ETE

[BRT98] Jordi Bascompte and Francisco Rodríguez-Trelles. Eradication thresholds in epidemiology, conservation biology and genetics. Journal of Theoretical Biology, 192(4):415–418, June 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519397905889.

Bastolla:1999:NEM

[BRV99] Ugo Bastolla, H. Eduardo Roman, and Michele Vendruscolo. Neutral evolution of model proteins: Diffusion in sequence space and overdispersion. Journal of Theoretical Biology, 200(1):49–64, September 7, 1999. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251939990975X.

Betts:1991:RHE

[BS91a] G. F. Betts and D. K. Srivastava. The rationalization of high enzyme concentration in metabolic pathways such as glycolysis. Journal of Theoretical Biology, 151(2):155–167, July 21, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305803595.

Brooks:1991:QEE

[BS91b] Stephen P. J. Brooks and Kenneth B. Storey. A quantitative evaluation of the effect of enzyme complexes on the glycolytic rate in vivo: Mathematical modeling of the glycolytic complex. Journal of Theoretical Biology, 149(3):361– 375, April 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251930580311X.

Buxbaum:1992:ISI

[BS92] Engelbert Buxbaum and Wilhelm Schoner. Investigation of subunit interactions by radiation inactivation: The case of Na^{+} /K^{+}-ATPase. Journal of Theoretical Biology, 155(1):21–31, March 7, 1992. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305805466.

Bonhoeffer:1993:ETC

[BS93] Sebastian Bonhoeffer and Peter F. Stadler. Error thresholds on correlated fitness landscapes. Journal of Theoretical Biology, 164(3):359–372, October 7, 1993. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519383711604.

Blackburn:1994:OWN

[BS94] D. G. Blackburn and C. W. Schneider. Old wine in a new bottle: Evolution by another name. Journal of Theoretical Biology, 171(2):233–337, November 21, 1994. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384712276.

Benhamou:1995:HFO

[BS95] Simon Benhamou and Valérie Séguinot. How to find one’s way in the labyrinth of path integration models. Journal of Theoretical Biology, 174(4):463–466, June 21, 1995. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385701129.

Baier:1998:HST

[BS98a] Gerold Baier and Sven Sahle. Homogeneous and spatiotemporal chaos in biochemical reactions with feedback inhibition. Journal of Theoretical Biology, 193(2):233–242, July 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519398906956.

Bascompte:1998:EHD

[BS98b] Jordi Bascompte and Ricard V. Solé. Effects of habitat destruction in a prey-predator metapopulation model. Journal of Theoretical Biology, 195(3):383–393, December 7, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519398908037.

Brunel:1998:FFL

[BS98c] Nicolas Brunel and Simone Sergi. Firing frequency of leaky intergrate-and-fire neurons with synaptic current dynamics. Journal of Theoretical Biology, 195(1):87–95, November 7, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519398907822.

Benhamou:1990:SML

[BSB90] Simon Benhamou, Jean-Pierre Sauvé, and Pierre Bovet. Spatial memory in large scale movements: Efficiency and limitation of the egocentric coding process. Journal of Theoretical Biology, 145(1):1–12, July 9, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305805314.

Bohl:1995:MTK

[BSB95] Erich Bohl, Howard A. Shuman, and Winfried Boos. Mathematical treatment of the kinetics of binding protein dependent transport systems reveals that both the substrate loaded and unloaded binding proteins interact with the membrane components. Journal of Theoretical Biology, 172(1):83–94, January 7, 1995. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519385700069.

Barrai:1991:FCH

[BSGB91] I. Barrai, C. Scapoli, R. Gambari, and F. Brungnoli. Frequencies of codons in histones, tubulins and fibrinogen: Bias due to interference between transcription signals and protein function. Journal of Theoretical Biology, 152(3):405–426, October 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305802024.

Banerjee:1990:QDO

[BSM90] Satindranath Banerjee, Peter R. Sibbald, and Jack Maze. Quantifying the dynamics of order and organization in biological systems. Journal of Theoretical Biology, 143(1):91–112, March 8, 1990. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305802905.

Bascompte:1997:PCS

[BSM97] Jordi Bascompte, Ricard V. Solé, and Norbert Martínez. Population cycles and spatial patterns in snowshoe hares: an individual-oriented simulation. Journal of Theoretical Biology, 187(2):213–222, July 21, 1997. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519397904306.

Barrai:1994:PIT

[BSN94a] I. Barrai, C. Scapoli, and C. Nesti. Possible identity of transcription and translation signals in early vital systems. Journal of Theoretical Biology, 169(3):289–294, August 7, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519384711490.

Barrai:1994:CUE

[BSN^{+}94b] I. Barrai, C. Scapoli, C. Nesti, G. Poli, R. Gambari, and M. Beretta. Codon usage and evolutionary rates of proteins. Journal of Theoretical Biology, 166(3):331–337, February 7, 1994. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519384710307.

Bastin:1992:MMH

[BSOB92] Marcel Bastin, Luc St-Onge, and Louise Bouchard. A mathematical model of homologous recombination in cultured cells. Journal of Theoretical Biology, 156(4):513–524, June 21, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305806411.

Baskaran:1996:ASP

[BSS96] Subbiah Baskaran, Peter F. Stadler, and Peter Schuster. Approximate scaling properties of RNA free energy landscapes. Journal of Theoretical Biology, 181(4):299–310, August 21, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519396901320.

Bartholdi:1993:PEF

[BSTV93] John J. Bartholdi, Thomas D. Seeley, Craig A. Tovey, and John H. Vande Vate. The pattern and effectiveness of forager allocation among flower patches by honey bee colonies. Journal of Theoretical Biology, 160(1):23–40, January 7, 1993. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519383710027.

Baigent:1997:MEG

[BSW97] Stephen Baigent, Jaroslav Stark, and Anne Warner. Modelling the effect of gap junction nonlinearities in systems of coupled cells. Journal of Theoretical Biology, 186(2):223–239, May 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396903513.

Brandts:1990:NLFb

[BT90a] W. A. M. Brandts and L. E. H. Trainor. A non-linear field model of pattern formation: Application to intracellular pattern reversal in tetrahymena. Journal of Theoretical Biology, 146(1):57–85, September 7, 1990. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930580044X.

Brandts:1990:NLFa

[BT90b] W. A. M. Brandts and L. E. H. Trainor. A non-linear field model of pattern formation: Intercalation in morphalactic regulation. Journal of Theoretical Biology, 146(1):37–56, September 7, 1990. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305800438.

Bespalova:1991:EHB

[BT91] S. V. Bespalova and K. B. Tolpygo. Excited hydrogen bonds in the molecular mechanism of muscle contraction. Journal of Theoretical Biology, 153(2):145–155, November 21, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305804199.

Bowers:1997:CSI

[BT97] Roger G. Bowers and Joanne Turner. Community structure and the interplay between interspecific infection and competition. Journal of Theoretical Biology, 187(1):95–109, July 7, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519397904185.

Borisuk:1998:BAM

[BT98] Mark T. Borisuk and John. J. Tyson. Bifurcation analysis of a model of mitotic control in frog eggs. Journal of Theoretical Biology, 195(1):69–85, November 7, 1998. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519398907810.

Brown:1999:PRB

[BT99] J. A. Brown and J. A. Tuszyński. The possible relationship between cell shape and electric fields. Journal of Theoretical Biology, 200(2):245–247, September 21, 1999. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519399909840.

Bonabeau:1998:GMR

[BTD98] Eric Bonabeau, Guy Theraulaz, and Jean-Louis Deneubourg. Group and mass recruitment in ant colonies: the influence of contact rates. Journal of Theoretical Biology, 195(2):157–166, November 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519398907895.

Baier:1993:CHM

[BTM93] Gerold Baier, Jesper S. Thomsen, and Erik Mosekilde. Chaotic hierarchy in a model of competing populations. Journal of Theoretical Biology, 165(4):593–607, December 21, 1993. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519383712099.

Bergman:1998:SPP

[BU98] Dan L. Bergman and Ulla Ullberg. Scaling properties of the Placenta’s arterial tree. Journal of Theoretical Biology, 193(4):731–738, August 21, 1998. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251939890741X.

Burlando:1990:FDT

[Bur90] Bruno Burlando. The fractal dimension of taxonomic systems. Journal of Theoretical Biology, 146(1):99–114, September 7, 1990. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305800463.

Burns:1992:AEH

[Bur92] Thomas P. Burns. Adaptedness, evolution and a hierarchical concept of fitness. Journal of Theoretical Biology, 154(2):219–237, January 21, 1992. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305804047.

Burger:1993:PDP

[Bür93a] Reinhard Bürger. Predictions of the dynamics of a polygenic character under directional selection. Journal of Theoretical Biology, 162(4):487–513, June 21, 1993. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251938371101X.

Burlando:1993:FGE

[Bur93b] Bruno Burlando. The fractal geometry of evolution. Journal of Theoretical Biology, 163(2):161–172, July 21, 1993. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519383711148.

Burns:1994:FOA

[Bur94] Thomas P. Burns. On the fittedness of organisms and the ascendency of ecosystems: Toward a hierarchical model of network development. Journal of Theoretical Biology, 170(2):115–127, September 21, 1994. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519384711726.

Burstein:1995:NMD

[Bur95] Zippora Burstein. A network model of developmental gene hierarchy. Journal of Theoretical Biology, 174(1):1–11, May 7, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519385700756.

Burkey:1999:EFH

[Bur99] Tormod Vaaland Burkey. Extinction in fragmented habitats predicted from stochastic birth-death processes with density dependence. Journal of Theoretical Biology, 199(4):395–406, August 21, 1999. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519399909670.

Butler:1990:FTG

[But90] James P. Butler. Fraction of trees with given root traits; the limit of large trees. Journal of Theoretical Biology, 147(2):265–274, November 21, 1990. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305800566.

Butte:1992:IRS

[But92] John C. Butte. The insulin receptor — single function and dual effect. Journal of Theoretical Biology, 155(4):427–436, April 21, 1992. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305806277.

Behn:1993:MAC

[BvHS93] Ulrich Behn, J. Leo van Hemmen, and Bernhard Sulzer. Memory to antigenic challenge of the immune system: Synergy of idiotypic interactions and memory B cells. Journal of Theoretical Biology, 165(1):1–25, November 7, 1993. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519383711732.

Bastiaanssen:1996:MML

[BvLVR96] E. H. C. Bastiaanssen, J. L. van Leeuwen, J. Vanderschoot, and P. A. Redert. A myocybernetic model of the lower urinary tract. Journal of Theoretical Biology, 178(2):113–133, January 21, 1996. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396900119.

Bastiaanssen:1996:SSA

[BVvL96] E. H. C. Bastiaanssen, J. Vanderschoot, and J. L. van Leeuwen. State-space analysis of a myocybernetic model of the lower urinary tract. Journal of Theoretical Biology, 180(3):215–227, June 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396900983.

Bretschneider:1997:MCM

[BVW97] Till Bretschneider, Bakhtier Vasiev, and Cornelis J. Weijer. A model for cell movement during Dictyostelium mound formation. Journal of Theoretical Biology, 189(1):41–51, November 7, 1997. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519397904902.

Bretschneider:1999:MDS

[BVW99] Till Bretschneider, Bakhtier Vasiev, and Cornelis J. Weijer. A model for Dictyostelium slug movement. Journal of Theoretical Biology, 199(2):125–136, July 21, 1999. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251939990944X.

Breden:1991:RSE

[BW91] Felix Breden and Michael J. Wade. “Runaway” social evolution: Reinforcing selection for inbreeding and altruism. Journal of Theoretical Biology, 153(3):323–337, December 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305805739.

Baker:1992:AFO

[BW92] Gary M. Baker and Lichun Weng. Apparent first-order behavior under second-order kinetic conditions: a general concept illustrated by the reversible binding of hydrogen peroxide to cytochrome c oxidase. Journal of Theoretical Biology, 158(2):221–229, September 21, 1992. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305807209.

Barnard:1998:AAE

[BW98] Ross Barnard and Rodney C. Wolff. Analysis and application of an equilibrium model for in vitro bioassay systems with three components: Receptor, hormone and hormone-binding-protein. Journal of Theoretical Biology, 190(4):333–339, February 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519397905555.

Brown:1996:MCA

[BWK96] Guy C. Brown, Hans V. Westerhoff, and Boris N. Kholodenko. Molecular control analysis: Control within proteins and molecular processes. Journal of Theoretical Biology, 182(3):389–396, October 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396901782.

Bing-Zheng:1991:IMM

[BZGM91] Liu Bing-Zheng and Deng Gou-Min. An improved mathematical model of hormone secretion in the hypothalamo-pituitary-gonadal axis in man. Journal of Theoretical Biology, 150(1): 51–58, May 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305804746.

Cobas:1992:PCP

[CA92] Alberto Cobas and Michael Arbib. Prey-catching and predatoravoidance in frog and toad: Defining the schemas. Journal of Theoretical Biology, 157(3):271–304, August 7, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305806125.

Cabanac:1992:PCC

[Cab92] Michel Cabanac. Pleasure: the common currency. Journal of Theoretical Biology, 155(2):173–200, March 21, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305805946.

Camargo:1993:MDI

[Cam93] J. A. Camargo. Must dominance increase with the number of subordinate species in competitive interactions? Journal of Theoretical Biology, 161(4):537–542, April 21, 1993. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519383710726.

Campbell:1995:EMS

[Cam95] R. B. Campbell. The effect of mating structure and progeny distribution on heterozygosity versus the number of alleles as measures of variation. Journal of Theoretical Biology, 175(4):503–509, August 21, 1995. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385701579.

Castrignano:1999:NPE

[CAPP99] Tiziana Castrignanò, Filippo Aluffi-Pentini, and Valerio Parisi. A novel parameter to estimate the minimum number of bound ligands needed to activate an ion channel. Journal of Theoretical Biology, 199(1):97–103, July 7, 1999. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519399909451.

Cardarelli:1990:RTP

[Car90] Nate F. Cardarelli. The role of a thymus-pineal axis in an immune mechanism of aging. Journal of Theoretical Biology, 145(3):397–405, August 9, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305801183.

Cardenas:1991:TEE

[Cár91] María Luz Cárdenas. Are the transitory enzyme-enzyme complexes found in vitro also transitory in vivo? if so, are they physiologically important? Journal of Theoretical Biology, 152(1):111–113, September 7, 1991. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305805223.

Cornish-Bowden:1991:HME

[CB91a] Athel Cornish-Bowden. How much effect on free metabolite concentrations does channelling have? Journal of Theoretical Biology, 152(1):39–40, September 7, 1991. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305805053.

Cornish-Bowden:1991:PEP

[CB91b] Athel Cornish-Bowden. Parameter estimating procedures for the Michaelis–Menten model: Reply to Tseng and Hsu. Journal of Theoretical Biology, 153(3):437–440, December 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305805806.

Cheng:1994:BMD

[CB94a] Jian-Yu Cheng and Reinhard Blickhan. Bending moment distribution along swimming fish. Journal of Theoretical Biology, 168(3):337–348, June 7, 1994. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519384711143.

Cronhjort:1994:HVP

[CB94b] Mikael B. Cronhjort and Clas Blomberg. Hypercycles versus parasites in a two dimensional partial differential equations model. Journal of Theoretical Biology, 169(1):31–49, July 7, 1994. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519384711283.

Chaudhuri:1996:RSS

[CB96a] Surasri Chaudhuri and Kamales Bhaumik. On the role of synaptotagmin in synaptic vesicle exocytosis. Journal of Theoretical Biology, 178(4):419–422, February 21, 1996. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396900399.

Cornish-Bowden:1996:HKA

[CB96b] Athel Cornish-Bowden. Henrik Kacser (1918–1995): an annotated bibliography. Journal of Theoretical Biology, 182 (3):195–199, October 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519396901551.

Chaudhuri:1997:DSV

[CB97a] Surasri Chaudhuri and Kamales Bhaumik. On the dynamics of synaptic vesicles: Effects on the rate of neurotransmitter release. Journal of Theoretical Biology, 187(3):351–362, August 7, 1997. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519397904446.

Condie:1997:IDS

[CB97b] Scott A. Condie and Myriam Bormans. The influence of density stratification on particle settling, dispersion and population growth. Journal of Theoretical Biology, 187(1):65–75, July 7, 1997. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519397904173.

Chaudhry:1998:ERS

[CBFR98] H. R. Chaudhry, B. Bukiet, T. Findley, and A. B. Ritter. Evaluation of residual stress in rabbit skin and the relevant material constants. Journal of Theoretical Biology, 192(2):191–195, May 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519397906160.

Chaudhry:1999:APR

[CBS^{+}99] H. R. Chaudhry, B. Bukiet, M. Siegel, T. Findley, A. B. Ritter, and N. Guzelsu. Adaptation of passive rat left ventricle in diastolic dysfunction. Journal of Theoretical Biology, 201(1):37–46, November 7, 1999. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519399910111.

Calenbuhr:1995:NTS

[CBSV95] Vera Calenbuhr, Hugues Bersini, John Stewart, and Francisco J. Varela. Natural tolerance in a simple immune network. Journal of Theoretical Biology, 177(3):199–213, December 7, 1995. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519385702391.

Caswell:1995:RWB

[CC95] Hal Caswell and Joel E. Cohen. Red, white and blue: environmental variance spectra and coexistence in metapopulations. Journal of Theoretical Biology, 176(2):301–316, September 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519385702007.

Chen:1996:CAB

[CC96] Lin Chen and Lawrence Chan. Control of apolipoprotein B mRNA editing: Implications of mRNA dynamics at various maturation stages. Journal of Theoretical Biology, 183(4):391–407, December 21, 1996. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396902301.

Cain:1997:TPCa

[CC97a] Scott J. Cain and Pao C. Chau. Transition probability cell cycle model. Part I — balanced growth. Journal of Theoretical Biology, 185(1):55–67, March 7, 1997. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396902891.

Cain:1997:TPCb

[CC97b] Scott J. Cain and Pao C. Chau. Transition probability cell cycle model. part II-nonbalanced growth. Journal of Theoretical Biology, 185(1):69–79, March 7, 1997. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396902908.

Ciblis:1997:PSE

[CC97c] Peter Ciblis and Irena Cosic. The possibility of soliton/ exciton transfer in proteins. Journal of Theoretical Biology, 184(3):331–338, February 7, 1997. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519396902817.

Chowdhary:1999:TAH

[CC99] Aasim G. Chowdhary and John H. Challis. Timing accuracy in human throwing. Journal of Theoretical Biology, 201(4):219–229, December 21, 1999. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251939991024X.

Cushing:1998:NPD

[CCD^{+}98] J. M. Cushing, R. F. Costantino, Brian Dennis, R. A. Desharnais, and Shandelle M. Henson. Nonlinear population dynamics: Models, experiments and data. Journal of Theoretical Biology, 194(1):1–9, September 7, 1998. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519398907366.

Calenbuhr:1992:MOOb

[CCDD92] V. Calenbuhr, L. Chrétien, J.-L. Deneubourg, and C. Detrain. A model for osmotropotactic orientation (II). Journal of Theoretical Biology, 158(3):395–407, October 7, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305807398.

Carneiro:1996:MINa

[CCFS96] Jorge Carneiro, António Coutinho, Jose Faro, and John Stewart. A model of the immune network with B–T cell cooperation. i-prototypical structures and dynamics. Journal of Theoretical Biology, 182(4):513–529, October 21, 1996. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396901927.

Crowley:1998:SCD

[CCG^{+}98] Philip H. Crowley, Ted Cottrell, Tiffany Garcia, Margret Hatch, R. Craig Sargent, Betty Jo Stokes, and James M. White. Solving the complementarity dilemma: Evolving strategies for simultaneous hermaphroditism. Journal of Theoretical Biology, 195(1):13–26, November 7, 1998. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519398907494.

Chen:1994:TLR

[CCI94] S. G. Chen, R. Ceulemans, and I. Impens. Is there a light regime determined tree ideotype? Journal of Theoretical Biology, 169(2):153–161, July 21, 1994. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384711374.

Cowell:1998:DDP

[CCK98] Lindsay G. Cowell, Larry B. Crowder, and Thomas B. Kepler. Density-dependent prenatal androgen exposure as an endogenous mechanism for the generation of cycles in small mammal populations. Journal of Theoretical Biology, 190(1):93–106, January 7, 1998. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519397905439.

Castrigano:1997:SOD

[CCM^{+}97] T. Castriganò, A. Colosimo, S. Morante, V. Paris, and G. C. Rossi. A study of oligonucleotide distributions in DNA coding segments. Journal of Theoretical Biology, 184(4):451–469, February 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396902842.

Cosmi:1990:CNS

[CCRM90] Carmelina Cosmi, Vincenzo Cuomo, Maria Ragosta, and Maria F. Macchiato. Characterization of nucleotidic sequences using maximum entropy techniques. Journal of Theoretical Biology, 147(3):423–432, December 7, 1990. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305804977.

Carneiro:1996:MINb

[CCS96] Jorge Carneiro, António Coutinho, and John Stewart. A model of the immune network with B–T cell co-operation. II — the simulation of ontogenesis. Journal of Theoretical Biology, 182(4):531–547, October 21, 1996. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396901939.

Castillo-Chavez:1998:RBE

[CCVH98] Carlos Castillo-Chavez and Jorge X. Velasco-Hernández. On the relationship between evolution of virulence and host demography. Journal of Theoretical Biology, 192(4):437–444, June 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519398906610.

Calenbuhr:1992:MOOa

[CD92] V. Calenbuhr and J.-L. Deneubourg. A model for osmotropotactic orientation (i). Journal of Theoretical Biology, 158(3):359–393, October 7, 1992. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305807386.

Caraco:1998:HSH

[CDG^{+}98] Thomas Caraco, Maria Duryea, Geoffrey Gardner, William Maniatty, and Boleslaw K. Szymanski. Host spatial heterogeneity and extinction of an SIS epidemic. Journal of Theoretical Biology, 192(3):351–361, June 7, 1998. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519398906634.

Chela-Flores:1992:TMB

[CF92] Julian Chela-Flores. Towards the molecular bases of polymerase dynamics. Journal of Theoretical Biology, 154(4): 519–539, February 21, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519305801675.

Chela-Flores:1994:VMF

[CF94a] Julian Chela-Flores. Are viroids molecular fossils of the RNA world? Journal of Theoretical Biology, 166(2):163–166, January 21, 1994. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519384710149.

Chela-Flores:1994:TTB

[CF94b] Julian Chela-Flores. Towards the theoretical bases of the folding of the 100-Å nucleosome filament. Journal of Theoretical Biology, 168(1):65–73, May 7, 1994. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519384710885.

Chen:1998:GTP

[CFC98a] Kevin C. Chen, Roseanne M. Ford, and Peter T. Cummings. The global turning probability density function for motile bacteria and its applications. Journal of Theoretical Biology, 195(2):139–155, November 21, 1998. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519398907688.

Chen:1998:MMM

[CFC98b] Kevin C. Chen, Roseanne M. Ford, and Peter T. Cummings. Mathematical models for motile bacterial transport in cylindrical tubes. Journal of Theoretical Biology, 195(4):481–504, December 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519398908086.

Caraco:1991:SFP

[CG91] Thomas Caraco and Luc-Alain Giraldea. Social foraging: Producing and scrounging in a stochastic environment. Journal of Theoretical Biology, 153(4):559–583, December 21, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305801560.

Cardenas:1996:GIS

[CG96] María Luz Cárdenas and Albert Goldbeter. The glucoseinduced switch between glycogen phosphorylase and glycogen synthase in the liver: Outlines of a theoretical approach. Journal of Theoretical Biology, 182(3):421–426, October 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519396901824.

Coop:1998:SCT

[CGB98] A. D. Coop, W. G. Gibson, and M. R. Bennett. Spontaneous calcium transients in automatic boutons and varicosities. Journal of Theoretical Biology, 195(3):395–411, December 7, 1998. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519398908049.

Caraco:1998:LDS

[CGM^{+}98] Thomas Caraco, Geoffrey Gardner, William Maniatty, Ewa Deelman, and Boleslaw K. Szymanski. Lyme disease: Selfregulation and pathogen invasion. Journal of Theoretical Biology, 193(4):561–575, August 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519398907226.

Chang:1990:EMC

[CH90a] You-Im Chang and Jyh-Ping Hsu. The effect of multivalent cations on adhesion time for cellular adhesion to solid surfaces. Journal of Theoretical Biology, 147(4):509–516, December 21, 1990. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305802620.

Corapcioglu:1990:TMB

[CH90b] M. Yavuz Corapcioglu and M. Akhter Hossain. Theoretical modeling of biodegradation and biotransformation of hydrocarbons in subsurface environments. Journal of Theoretical Biology, 142(4):503–516, February 22, 1990. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305801055.

Courtois:1991:SCP

[CH91] P.-J. Courtois and F. Heymans. A simulation of the construction process of a termite nest. Journal of Theoretical Biology, 153(4):469–475, December 21, 1991. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305801511.

Crist:1994:SMM

[CH94] Thomas O. Crist and James W. Haefner. Spatial model of movement and foraging in harvester ants (Pogonomyrmex) (II): The roles of environment and seed dispersion. Journal of Theoretical Biology, 166(3):315–323, February 7, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384710289.

Comins:1996:PMH

[CH96] H. N. Comins and M. P. Hassell. Persistence of multispecies host-parasitoid interactions in spatially distributed models with local dispersal. Journal of Theoretical Biology, 183(1):19–28, November 7, 1996. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519396901976.

Chay:1990:BEC

[Cha90] Teresa Ree Chay. Bursting excitable cell models by a slow Ca^{2+} current. Journal of Theoretical Biology, 142(3):305–315, February 9, 1990. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305805557.

Chao:1991:LSE

[Cha91] Lin Chao. Levels of selection, evolution of sex in RNA viruses, and the origin of life. Journal of Theoretical Biology, 153(2):229–246, November 21, 1991. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305804242.

Chaplain:1993:SEF

[Cha93] M. A. J. Chaplain. The strain energy function of an ideal plant cell wall. Journal of Theoretical Biology, 163(1):77–97, July 1, 1993. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519383711082.

Chang:1998:TMF

[Cha98] Kuo Chu Chang. Theoretical maximal flow of the left ventricle is sensitive to change in ventricular afterload. Journal of Theoretical Biology, 194(3):409–417, October 7, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519398907731.

Cheetham:1993:AFF

[Che93] Tom Cheetham. Adaptation on fluctuating fitness landscapes: Speciation and the persistence of lineages. Journal of Theoretical Biology, 161(3):287–297, April 7, 1993. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519383710568.

Chernyak:1994:ECE

[Che94] Yuri B. Chernyak. Effects of conformational elasticity and ionchannel interaction on thermodynamic equilibrium of biomembranes. Journal of Theoretical Biology, 166(4):375–392, February 21, 1994. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519384710344.

Chittka:1996:OSC

[Chi96] Lars Chittka. Optimal sets of color receptors and color opponent systems for coding of natural objects in insect vision. Journal of Theoretical Biology, 181(2):179–196, July 21, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519396901241.

Caldarelli:1998:MCM

[CHM98] Guido Caldarelli, Paul G. Higgs, and Alan J. McKane. Modelling coevolution in multispecies communities. Journal of Theoretical Biology, 193(2):345–358, July 21, 1998. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519398907068.

Chowdhury:1993:UMI

[Cho93] Debashish Chowdhury. A unified model of immune response II: Continuum approach. Journal of Theoretical Biology, 165(2):135–159, November 21, 1993. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251938371180X.

Chaturvedi:1997:MMP

[CI97] Pawan Chaturvedi and Michael F. Insana. Mathematical model of progressive renal disease. Journal of Theoretical Biology, 185(3):357–365, April 7, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519396903112.

Colding-Jorgensen:1990:FPA

[CJ90] Morten Colding-Jørgensen. Fundamental properties of the action potential and repetitive activity in excitable membranes illustrated by a simple model. Journal of Theoretical Biology, 144(1):37–67, May 10, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251930580298X.

Colding-Jorgensen:1992:MMC

[CJMBM92] Morten Colding-Jørgensen, Henrik Østergaard Madsen, Brian Bodholdt, and Erik Mosekilde. Minimal model for Ca^{2+}- dependent oscillations in excitable cells. Journal of Theoretical Biology, 156(3):309–326, June 7, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305806782.

Cavalieri:1994:CBC

[CK94] Liebe F. Cavalieri and Huseyin Koçak. Chaos in biological control systems. Journal of Theoretical Biology, 169(2):179–187, July 21, 1994. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519384711398.

Cavalieri:1995:ITB

[CK95a] Liebe F. Cavalieri and Huseyin Koçak. Intermittent transition between order and chaos in an insect pest population. Journal of Theoretical Biology, 175(2):231–234, July 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251938570135X.

Chang:1995:RMN

[CK95b] Hangil Chang and Kiyoshi Kurokawa. Reliability of maximum number of simultaneously open channels as an estimator for the number of channels in single-channel recordings. Journal of Theoretical Biology, 173(1):61–65, March 7, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519385700434.

Chang:1996:ETT

[CK96] Kuo-Chu Chang and Te-Son Kuo. Exponentially tapered ttube model in the characterization of arterial non-uniformity. Journal of Theoretical Biology, 183(1):35–46, November 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251939690199X.

Carrier:1997:MMP

[CK97a] Trent A. Carrier and J. D. Keasling. Mechanistic modeling of prokaryotic mRNA decay. Journal of Theoretical Biology, 189(2):195–209, November 21, 1997. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519397905099.

Chang:1997:SBE

[CK97b] Kuo-Chu Chang and Te-Son Kuo. Single-beat estimation of the ventricular pumping mechanics in terms of the systolic elastance and resistance. Journal of Theoretical Biology, 189(1):89–95, November 7, 1997. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519397904987.

Carrier:1999:IAG

[CK99] Trent A. Carrier and J. D. Keasling. Investigating autocatalytic gene expression systems through mechanistic modeling. Journal of Theoretical Biology, 201(1):25–36, November 7, 1999. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251939991010X.

Clark:1995:CKE

[CKR95] Joseph F. Clark, Graham J. Kemp, and George K. Radda. The creatine kinase equilibrium, free [ADP] and myosin AT-Pase in vascular smooth muscle cross-bridges. Journal of Theoretical Biology, 173(2):207–211, March 21, 1995. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385700562.

Chay:1992:SRE

[CL92] Teresa Ree Chay and Young Seek Lee. Studies on re-entrant arrhythmias and ectopic beats in excitable tissues by bifurcation analyses. Journal of Theoretical Biology, 155(2):137–171, March 21, 1992. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305805934.

Chuev:1993:PME

[CL93] G. N. Chuev and V. D. Lakhno. A polaron model for electron transfer in globular proteins. Journal of Theoretical Biology, 163(1):51–60, July 1, 1993. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519383711069.

Chechetkin:1998:SCS

[CL98] V. R. Chechetkin and V. V. Lobzin. Study of correlations in segmented DNA sequences: Application to structure coupling between exons and introns. Journal of Theoretical Biology, 190(1):69–83, January 7, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251939790535X.

Chechetkin:1999:CCPa

[CL99a] V. R. Chechetkin and V. V. Lobzin. Characterization and comparison of protein structures. Part I — characterization. Journal of Theoretical Biology, 198(2):197–218, May 21, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519399909104.

Chechetkin:1999:CCPb

[CL99b] V. R. Chechetkin and V. V. Lobzin. Characterization and comparison of protein structures. part II-comparison. Journal of Theoretical Biology, 198(2):219–237, May 21, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519399909116.

Clay:1995:SMK

[Cla95] John R. Clay. A simple model of K^{+} channel activation in nerve membrane. Journal of Theoretical Biology, 175(2):257– 262, July 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519385701385.

Clearwater:1991:MMP

[Cle91a] Scott H. Clearwater. A model of morphogenesis for protein secondary structures. Journal of Theoretical Biology, 153(3):353–370, December 7, 1991. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305805752.

Clegg:1991:PSM

[Cle91b] James S. Clegg. The physiological significance of metabolite channeling: an idea whose time has come. Journal of Theoretical Biology, 152(1):63–64, September 7, 1991. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305805119.

Chay:1995:APL

[CLF95] Teresa Ree Chay, Young Seek Lee, and Yin Shui Fan. Appearance of phase-locked wenckebach-like rhythms, devil’s staircase and universality in intracellular calcium spikes in nonexcitable cell models. Journal of Theoretical Biology, 174(1): 21–44, May 7, 1995. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251938570077X.

Cascante:1996:MPC

[CLMH^{+}96] Marta Cascante, Mónica Lloréns, Enrique Meléndez-Hevia, Joaquim Puigjaner, Francisco Montero, and Enric Martín. The metabolic productivity of the cell factory. Journal of Theoretical Biology, 182(3):317–325, October 7, 1996. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396901708.

Cho:1993:ECD

[CLW93] Carolyn R. Cho, Charles J. Lumsden, and Catharine I. Whiteside. Epithelial cell detachment in the nephrotic glomerulus: a receptor co-operativity model. Journal of Theoretical Biology, 160(4):407–426, February 21, 1993. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519383710271.

Czerlinski:1999:HSU

[CLY99] George Czerlinski, Richard Levin, and Tjalling Ypma. Hemoglobin/O_{2} systems: Using short-lived intermediates for mechanistic discrimination. Journal of Theoretical Biology, 199(1):25–44, July 7, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519399909396.

Chen:1995:CSG

[CM95] Lin Chen and Medea Q. Meng. Compact and scattered gap junctions in diffusion mediated cell-cell communication. Journal of Theoretical Biology, 176(1):39–45, September 7, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519385701749.

Cortez-Maghelly:1995:EIS

[CMB95] Célia Cortez-Maghelly and Paulo M. Bisch. Effect of ionic strength and outer surface charge on the mechanical stability of the erythrocyte membrane: a linear hydrodynamic analysis. Journal of Theoretical Biology, 176(3):325–339, October 7, 1995. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519385702020.

Clarke:1996:MMA

[CMG^{+}96] B. Clarke, I. McKay, T. Grigliatti, V. Lloyd, and A. Yuan. A Markov model for the assembly of heterochromatic regions in position effect variegation. Journal of Theoretical Biology, 181(2):137–155, July 21, 1996. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519396901216.

Collier:1996:PFL

[CMML96] Joanne R. Collier, Nicholas A. M. Monk, Philip K. Maini, and Julian H. Lewis. Pattern formation by lateral inhibition with feedback: a mathematical model of delta-notch intercellular signalling. Journal of Theoretical Biology, 183(4):429–446, December 21, 1996. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396902337.

Colosimo:1993:IMD

[CMPR93] A. Colosimo, S. Morante, V. Parisi, and G. C. Rossi. An improved method for detection of words with unusual occurrence frequency in nucleotidic sequences. Journal of Theoretical Biology, 165(4):659–672, December 21, 1993. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519383712129.

Clarke:1993:MEN

[CMS93] Bertrand Clarke, Jay E. Mittenthal, and Mark Senn. A model for the evolution of networks of genes. Journal of Theoretical Biology, 165(3):269–289, December 7, 1993. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519383711896.

Cockell:1998:BEH

[Coc98] Charles S. Cockell. Biological effects of high ultraviolet radiation on early Earth — a theoretical evaluation. Journal of Theoretical Biology, 193(4):717–729, August 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251939890738X.

Cohen:1991:DRC

[Coh91] Robert J. Cohen. Determining rate constants for irreversible polymerization where the initial step and propagation steps have different rate constants: Consideration of polyadenylate polymerase. Journal of Theoretical Biology, 150(4):529–537, June 21, 1991. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305804448.

Colgan:1993:DRP

[Col93a] D. J. Colgan. DNA regeneration in the polymerase chain reaction. Journal of Theoretical Biology, 162(3):289–307, June 7, 1993. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251938371088X.

Colmenares:1993:RSF

[Col93b] Pedro J. Colmenares. Role of a stochastic friction coefficient in open channel noise. Journal of Theoretical Biology, 161(2):175–198, March 21, 1993. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519383710490.

Comper:1990:EMI

[Com90] Wayne D. Comper. Extracellular matrix interactions: Sulfation of connective tissue polysaccharides creates macroion binding templates and conditions for dissipative structure formation. Journal of Theoretical Biology, 145(4):497–509, August 23, 1990. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305804849.

Comins:1994:EAI

[Com94a] Hugh N. Comins. Equilibrium analysis of integrated plantsoil models for prediction of the nutrient limited growth response to CO 2 enrichment. Journal of Theoretical Biology, 171(4):369–385, December 21, 1994. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519384712410.

Comper:1994:THP

[Com94b] Wayne D. Comper. The thermodynamic and hydrodynamic properties of macromolecules that influence the hydrodynamics of porous systems. Journal of Theoretical Biology, 168(4):421– 427, June 21, 1994. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519384711222.

Conway:1993:DPD

[Con93] Kevin M. Conway. Diffusion patterns on domains representing developing xenopus retina. Journal of Theoretical Biology, 163(2):181–197, July 21, 1993. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519383711161.

Cordero:1995:ESA

[Cor95] Carlos Cordero. Ejaculate substances that affect female insect reproductive physiology and behavior: Honest or arbitrary traits? Journal of Theoretical Biology, 174(4):453–461, June 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519385701117.

Cordero:1998:COS

[Cor98] Carlos Cordero. Chemical ornaments of semen. Journal of Theoretical Biology, 192(4):581–584, June 21, 1998. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519397906214.

Coutinho:1991:SCB

[Cou91] Antonio Coutinho. Size and connectivity: a bit of the history of immune networks. Journal of Theoretical Biology, 149(3):425– 427, April 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305803145.

Clayton:1992:RIP

[CPJL92] Dale H. Clayton, S. G. Pruett-Jones, and Russell Lande. Reappraisal of the interspecific prediction of parasite-mediated sexual selection: Opportunity knocks. Journal of Theoretical Biology, 157(1):95–108, July 7, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305807593.

Cascante:1996:SSC

[CPK96] Marta Cascante, Joaquim Puigjaner, and Boris Kholodenko. Steady-state characterization of systems with moietyconservations made easy: Matrix equations of metabolic control analysis and biochemical system theory. Journal of Theoretical Biology, 178(1):1–6, January 7, 1996. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396900016.

Courchamp:1995:PDF

[CPLA95] Franck Courchamp, Dominique Pontier, Michel Langlais, and Marc Artois. Population dynamics of feline immunodeficiency virus within cat populations. Journal of Theoretical Biology, 175(4):553–560, August 21, 1995. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385701634.

Crawford:1990:NEE

[Cra90a] Vincent P. Crawford. Nash equilibrium and evolutionary stability in large- and finite-population “playing the field” models. Journal of Theoretical Biology, 145(1):83–94, July 9, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305805363.

Crawford:1990:DES

[Cra90b] Vincent P. Crawford. On the definition of an evolutionarily stable strategy in the “playing the field” model. Journal of Theoretical Biology, 143(2):269–273, March 22, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305802711.

Cressman:1990:SSD

[Cre90] R. Cressman. Strong stability and density-dependent evolutionarily stable strategies. Journal of Theoretical Biology, 145(3):319–330, August 9, 1990. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305801122.

Chaplain:1990:AMT

[CS90] M. A. J. Chaplain and B. D. Sleeman. An application of membrane theory to tip morphogenesis in acetabularia. Journal of Theoretical Biology, 146(2):177–200, September 21, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305801341.

Camazine:1991:MCN

[CS91] Scott Camazine and James Sneyd. A model of collective nectar source selection by honey bees: Self-organization through simple rules. Journal of Theoretical Biology, 149(4):547–571, April 21, 1991. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305800980.

Carneiro:1994:RSS

[CS94a] Jorge Carneiro and John Stewart. Rethinking “shape space”: Evidence from simulated docking suggests that steric shape complementarity is not limiting for antibody– antigen recognition and idiotypic interactions. Journal of Theoretical Biology, 169(4):391–402, August 21, 1994. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384711611.

Chen:1994:SPT

[CS94b] Pei-Li Chen and Rainer Sachs. Singular perturbation theory applied to Markov models for DNA damage caused by ionizing radiation. Journal of Theoretical Biology, 166(2):117–126, January 21, 1994. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519384710101.

Coret:1996:CRC

[CS96] I. A. Coret and D. A. Schwalbe. On the contributions of RX complex and agonist-receptor complex to formation of the ternary complex: Results of a new method of analysis. Journal of Theoretical Biology, 179(4):287–301, April 21, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519396900685.

Cummings:1998:MP

[CS98] F. W. Cummings and J. C. Strickland. A model of phyllotaxis. Journal of Theoretical Biology, 192(4):531–544, June 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519398906828.

Clay:1999:RSD

[CS99] John R. Clay and Alvin Shrier. On the role of subthreshold dynamics in neuronal signaling. Journal of Theoretical Biology, 197(2):207–216, March 21, 1999. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519398908670.

Cassill:1998:EPF

[CSB98] Deby Lee Cassill, Alexander Stuy, and Ronald G. Buck. Emergent properties of food distribution among fire ant larvae. Journal of Theoretical Biology, 195(3):371–381, December 7, 1998. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519398908025.

Chowdhury:1990:UDM

[CSC90] Debashish Chowdhury, Dietrich Stauffer, and Prabhakara V. Choudary. A unified discrete model of immune response. Journal of Theoretical Biology, 145(2):207–215, July 24, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305801262.

Cascante:1991:BST

[CSFC91] Marta Cascante, Albert Sorribas, Rafael Franco, and Enric I. Canela. Biochemical systems theory: Increasing predictive power by using second-order derivatives measurements. Journal of Theoretical Biology, 149(4):521–535, April 21, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305800967.

Camazine:1990:MMS

[CSJM90] Scott Camazine, James Sneyd, Michael J. Jenkins, and J. D. Murray. A mathematical model of self-organized pattern formation on the combs of honeybee colonies. Journal of Theoretical Biology, 147(4):553–571, December 21, 1990. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305802644.

Chowdhury:1991:DMI

[CSS91] Debashish Chowdhury, Muhammad Sahimi, and Dietrich Stauffer. A discrete model for immune surveillance, tumor immunity and cancer. Journal of Theoretical Biology, 152(2):263–270, September 21, 1991. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305804576.

Chelazzi:1993:DHA

[CSSF93] G. Chelazzi, G. Santini, P. Della Santina, and S. Focardi. Does the homing accuracy of intertidal chitons rely on active trail following? a simulation approach. Journal of Theoretical Biology, 160(2):165–178, January 21, 1993. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519383710118.

Castro:1998:CCC

[CST98] Laureano Castro, José M. Serrano, and Miguel A. Toro. Conceptual capacity to categorize and the evolution of altruism. Journal of Theoretical Biology, 192(4):561–565, June 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519398906853.

Chechetkin:1995:SHP

[CT95] V. R. Chechetkin and A. Yu. Turygin. Search of hidden periodicities in DNA sequences. Journal of Theoretical Biology, 175(4):477–494, August 21, 1995. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385701555.

Chechetkin:1996:SCD

[CT96] V. R. Chechetkin and A. Y. Turygin. Study of correlations in DNA sequences. Journal of Theoretical Biology, 178(2):205–217, January 21, 1996. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251939690017X.

Cannings:1993:RP

[CTV93] C. Cannings, J. P. Tyrer, and G. T. Vickers. Routes to polymorphism. Journal of Theoretical Biology, 165(2):213–223, November 21, 1993. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519383711860.

Cummings:1990:MMP

[Cum90] F. W. Cummings. A model of morphogenetic pattern formation. Journal of Theoretical Biology, 144(4):547–566, June 21, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251930580089X.

Cummings:1996:GCE

[Cum96a] F. W. Cummings. Geometrical concepts in epithelial sheets. Journal of Theoretical Biology, 179(1):41–49, March 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396900466.

Cummings:1996:MGF

[Cum96b] Frederick W. Cummings. A model of growth and form based on adhesion molecules. Journal of Theoretical Biology, 178(3):229–238, February 7, 1996. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519396900223.

Cummings:1999:SPP

[Cum99a] F. W. Cummings. Spatial patterning via PTP adhesive phosphatases. Journal of Theoretical Biology, 196(1):19–26, January 7, 1999. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519398908141.

Cummings:1999:WPF

[Cum99b] F. W. Cummings. Waves of pattern formation and signal pathways. Journal of Theoretical Biology, 196(1):27–31, January 7, 1999. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519398908153.

Cushing:1995:WHC

[Cus95] Bruce S. Cushing. When a hawk can damage a dove: an extension of game theory. Journal of Theoretical Biology, 175(2):173–176, July 21, 1995. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519385701294.

Collado-Vides:1991:SRU

[CV91] Julio Collado-Vides. A syntactic representation of units of genetic information — a syntax of units of genetic information. Journal of Theoretical Biology, 148(3):401–429, February 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305802450.

Collado-Vides:1993:ECU

[CV93] Julio Collado-Vides. The elements for a classification of units of genetic information with a combinatorial component. Journal of Theoretical Biology, 163(4):527–548, August 21, 1993. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519383711367.

Cressman:1997:SDE

[CV97] R. Cressman and G. T. Vickers. Spatial and density effects in evolutionary game theory. Journal of Theoretical Biology, 184(4):359–369, February 21, 1997. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396902519.

Cannings:1991:TTT

[CW91] C. Cannings and J. C. Whittaker. A two-trial two-strategy conflict. Journal of Theoretical Biology, 149(2):281–286, March 21, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305802826.

Conrad:1992:CBL

[CW92] Michael Conrad and Richard G. Weinand. CD5 + B lymphocytes correspond to generalist lymphocytes predicted by a computational model of affinity maturation. Journal of Theoretical Biology, 154(3):399–401, February 7, 1992. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305801791.

Cooper:1998:EPC

[CW98] Ben Cooper and Chris Wallace. Evolution, partnerships and cooperation. Journal of Theoretical Biology, 195(3):315–328, December 7, 1998. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519398907949.

Crawford:1990:MSF

[CY90] John W. Crawford and Iain M. Young. A multiple scaled fractal tree. Journal of Theoretical Biology, 145(2):199–206, July 24, 1990. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305801250.

Chu:1998:ISP

[CY98] C. Y. Cyrus Chu and R. Yu. Individual sex preferences and the population sex ratio of newborns. Journal of Theoretical Biology, 194(3):383–390, October 7, 1998. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519398907640.

Chou:1993:JPF

[CZ93] James Jeiwen Chou and Chun-Ting Zhang. A joint prediction of the folding types of 1490 human proteins from their genetic codons. Journal of Theoretical Biology, 161(2):251–262, March 21, 1993. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519383710532.

Czerlinski:1992:PKH

[Cze92] George H. Czerlinski. On the problem of “kinetically hidden” reaction steps. Journal of Theoretical Biology, 154(1):119–126, January 7, 1992. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305801924.

Czerlinski:1993:RFE

[Cze93] George H. Czerlinski. Resolving the fast equilibrations of ternary complexes in liver alcohol dehydrogenase. Journal of Theoretical Biology, 165(3):313–320, December 7, 1993. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519383711914.

D:1992:MAC

[D.92] Mark D. A method for the analysis of comparative data. Journal of Theoretical Biology, 156(4):431–442, June 21, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930580637X.

Dimitrov:1996:LPA

[DA96] Dimiter S. Dimitrov and Mariana A. Apostolova. The limit of PCR amplification. Journal of Theoretical Biology, 178(4):425–426, February 21, 1996. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519396900417.

Datta:1998:DCD

[DA98] Susmita Datta and Jonathan Arnold. Dynamics of cytonuclear disequilibria in subdivided populations. Journal of Theoretical Biology, 192(1):99–111, May 7, 1998. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519397906238.

Damgaard:1999:CPH

[Dam99] Christian Damgaard. Coevolution of a plant host-pathogen gene-for-gene system in a metapopulation model without cost of resistance or cost of virulence. Journal of Theoretical Biology, 201(1):1–12, November 7, 1999. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251939991007X.

Davis:1991:MMM

[Dav91] Danny N. Davis. Models for the movement of mono-oriented chromosomes. Journal of Theoretical Biology, 150(3):305–322, June 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305804308.

Davidson:1998:MQR

[Dav98a] F. A. Davidson. Modelling the qualitative response of fungal mycelia to heterogeneous environments. Journal of Theoretical Biology, 195(3):281–292, December 7, 1998. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519398907391.

Davydov:1998:AAC

[Dav98b] Oleg V. Davydov. Amino acid contribution to the genetic code structure: End-atom chemical rules of doublet composition. Journal of Theoretical Biology, 193(4):679–690, August 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519398907330.

Dawson:1995:AAR

[Daw95] Kevin J. Dawson. The advantage of asexual reproduction: When is it two-fold? Journal of Theoretical Biology, 176(3):341–347, October 7, 1995. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519385702032.

Dawson:1996:ECT

[Daw96] Kevin J. Dawson. Evolutionary consequences of a trade-off between parental effort and mating effort. Journal of Theoretical Biology, 183(2):139–158, November 21, 1996. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396902088.

Dawson:1998:ESM

[Daw98] Kevin J. Dawson. Evolutionarily stable mutation rates. Journal of Theoretical Biology, 194(1):143–157, September 7, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519398907524.

Das:1995:NNF

[DB95] Bigyani Das and R. L. Batra. Non-Newtonian flow of blood in an arteriosclerotic blood vessel with rigid permeable walls. Journal of Theoretical Biology, 175(1):1–11, July 7, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519385701154.

DeFonzo:1998:ORT

[DBAP^{+}98] Valeria De Fonzo, Enrico Bersani, Filippo Aluffi-Pentini, Tiziana Castrignanò, and Valerio Parisi. Are only repeated triplets guilty? Journal of Theoretical Biology, 194(1):125–142, September 7, 1998. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519398907536.

Das:1999:PMP

[DBB99] Gourab Kanti Das, Dhananjay Bhattacharyya, and Debi Prosad Burma. A possible mechanism of peptide bond formation on ribosome without mediation of peptidyl transferase. Journal of Theoretical Biology, 200(2):193–205, September 21, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519399909876.

Dexter:1994:TAP

[DBH94] Franklin Dexter and Shlomo Ben-Haim. Theoretical analysis predicts that respiratory sinus arrhythmia does not accurately measure efferent vagal activity during anesthesia. Journal of Theoretical Biology, 169(2):133–141, July 21, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384711350.

deBoer:1991:SCE

[dBP91] Rob J. de Boer and Alan S. Perelson. Size and connectivity as emergent properties of a developing immune network. Journal of Theoretical Biology, 149(3):381–424, April 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305803133.

Duvdevani-Bar:1994:TSD

[DBS94] Sharon Duvdevani-Bar and Lee Segel. On topological simulations in developmental biology. Journal of Theoretical Biology, 166(1):33–50, January 7, 1994. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519384710034.

Dvorsky:1997:KSD

[DBS97] R. Dvorsky, S. Balaz, and R. J. Sawchuk. Kinetics of subcellular distribution of compounds in simple biosystems and its use in QSAR. Journal of Theoretical Biology, 185(2):213–222, March 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396903082.

Detours:1994:DIN

[DBSV94] Vincent Detours, Hugues Bersini, John Stewart, and Francisco Varela. Development of an idiotypic network in shape space. Journal of Theoretical Biology, 170(4):401–414, October 21, 1994. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251938471201X.

Dillon:1990:TDF

[DC90] Patrick F. Dillon and Joseph F. Clark. The theory of diazymes and functional coupling of pyruvate kinase and creatine kinase. Journal of Theoretical Biology, 143(2):275–284, March 22, 1990. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305802723.

Douady:1996:PDSa

[DC96a] S. Douady and Y. Couder. Phyllotaxis as a dynamical self organizing process part i: The spiral modes resulting from time-periodic iterations. Journal of Theoretical Biology, 178(3):255–273, February 7, 1996. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396900247.

Douady:1996:PDSb

[DC96b] S. Douady and Y. Couder. Phyllotaxis as a dynamical self organizing process part II: The spontaneous formation of a periodicity and the coexistence of spiral and whorled patterns. Journal of Theoretical Biology, 178(3):275–294, February 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396900259.

Douady:1996:PDSc

[DC96c] S. Douady and Y. Couder. Phyllotaxis as a dynamical self organizing process part III: The simulation of the transient regimes of ontogeny. Journal of Theoretical Biology, 178(3):295–312, February 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396900260.

Dall:1999:MRD

[DC99] Sasha R. X. Dall and Innes C. Cuthill. Mutation rates: does complexity matter? Journal of Theoretical Biology, 198(2):283–285, May 21, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519399909153.

Dokos:1996:ICU

[DCL96a] Socrates Dokos, Branko Celler, and Nigel Lovell. Ion currents underlying sinoatrial node pacemaker activity: a new single cell mathematical model. Journal of Theoretical Biology, 181(3):245–272, August 7, 1996. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396901290.

Dokos:1996:VCS

[DCL96b] Socrates Dokos, Branko Celler, and Nigel Lovell. Vagal control of sinoatrial rhythm: a mathematical model. Journal of Theoretical Biology, 182(1):21–44, September 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396901411.

DiGiulio:1994:OPD

[DCM94] Massimo Di Giulio, M. Rosaria Capobianco, and Mario Medugno. On the optimization of the physicochemical distances between amino acids in the evolution of the genetic code. Journal of Theoretical Biology, 168(1):43–51, May 7, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519384710861.

Dewey:1998:NET

[DD98] T. Gregory Dewey and Mariano Delle Donne. Non-equilibrium thermodynamics of molecular evolution. Journal of Theoretical Biology, 193(4):593–599, August 21, 1998. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251939890724X.

Duncan:1996:ODS

[DDS96] C. J. Duncan, S. R. Duncan, and Susan Scott. Oscillatory dynamics of smallpox and the impact of vaccination. Journal of Theoretical Biology, 183(4):447–454, December 21, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519396902349.

DeCarvalho:1995:UBT

[De 95] Hernandes Faustino De Carvalho. Understanding the biomechanics of tendon fibrocartilages. Journal of Theoretical Biology, 172(3):293–297, February 7, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519385700252.

Dukas:1998:SDC

[DEK98] Reuven Dukas and Leah Edelstein-Keshet. The spatial distribution of colonial food provisioners. Journal of Theoretical Biology, 190(2):121–134, January 21, 1998. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519397905300.

Delano:1995:SPC

[Del95] Michael D. Delano. Simple physical constraints in hemolysis. Journal of Theoretical Biology, 175(4):517–524, August 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519385701592.

Demidov:1994:DFP

[Dem94] Andrey A. Demidov. Determination of fluorescence polarization of double chromophore complexes. Journal of Theoretical Biology, 170(4):355–358, October 21, 1994. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384711970.

Demetrius:1998:RES

[Dem98] Lloyd Demetrius. Role of enzyme-substrate flexibility in catalytic activity: an evolutionary perspective. Journal of Theoretical Biology, 194(2):175–194, September 21, 1998. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519398907482.

Dexter:1995:SAD

[Dex95] Franklin Dexter. Statistical analysis of drug interactions in anesthesia. Journal of Theoretical Biology, 172(4):305–314, February 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519385700276.

Dani:1991:ESL

[DF91a] John A. Dani and James A. Fox. Examination of subconductance levels arising from a single ion channel. Journal of Theoretical Biology, 153(3):401–423, December 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305805788.

Duszyk:1991:AMI

[DF91b] Marek Duszyk and Andrew S. French. An analytical model of ionic movements in airway epithelial cells. Journal of Theoretical Biology, 151(2):231–247, July 21, 1991. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305803625.

Dillon:1995:MMB

[DFG95] Robert Dillon, Lisa Fauci, and Donald Gaver III. A microscale model of bacterial swimming, chemotaxis and substrate transport. Journal of Theoretical Biology, 177(4):325–340, December 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519385702512.

Dye:1993:SRI

[DG93] C. Dye and H. C. F. Godfray. On sex ratio and inbreeding in malaria parasite populations. Journal of Theoretical Biology, 161(1):131–134, March 7, 1993. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519383710453.

Dubischar:1999:AMP

[DGSK99] Daniel Dubischar, Volker Matthias Gundlach, Oliver Steinkamp, and Andrei Khrennikov. A p-adic model for the process of thinking disturbed by physiological and information noise. Journal of Theoretical Biology, 197(4):451–467, April 21, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519398908876.

D:1992:SCP

[DH92] Mark D. and Paul H. Harvey. On solving the correct problem: Wishing does not make it so. Journal of Theoretical Biology, 156(4):425–430, June 21, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305806368.

Dugatkin:1995:DBE

[DH95] Lee Alan Dugatkin and Jacob Höglund. Delayed breeding and the evolution of mate copying in lekking species. Journal of Theoretical Biology, 174(3):261–267, June 7, 1995. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385700975.

Deleuze:1998:SPB

[DH98] Christine Deleuze and François Houllier. A simple processbased xylem growth model for describing wood microdensitometric profiles. Journal of Theoretical Biology, 193(1):99– 113, July 7, 1998. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519398906890.

DiGiulio:1992:OTR

[Di 92] Massimo Di Giulio. On the origin of the transfer RNA molecule. Journal of Theoretical Biology, 159(2):199–214, November 21, 1992. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305807027.

DiGiulio:1994:OPS

[Di 94] Massimo Di Giulio. On the origin of protein synthesis: a speculative model based on hairpin RNA structures. Journal of Theoretical Biology, 171(3):303–308, December 7, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384712331.

DiGiulio:1995:WIA

[Di 95] Massimo Di Giulio. Was it an ancient gene codifying for a hairpin RNA that, by means of direct duplication, gave rise to the primitive tRNA molecule? Journal of Theoretical Biology, 177(1):95–101, November 7, 1995. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305800074.

DiGiulio:1997:OGC

[Di 97] Massimo Di Giulio. On the origin of the genetic code. Journal of Theoretical Biology, 187(4):573–581, August 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519396903902.

DiGiulio:1998:ROG

[Di 98] Massimo Di Giulio. Reflections on the origin of the genetic code: a hypothesis. Journal of Theoretical Biology, 191(2):191–196, March 21, 1998. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519397905804.

DiGiulio:1999:NMO

[Di 99] Massimo Di Giulio. The non-monophyletic origin of the tRNA molecule. Journal of Theoretical Biology, 197(3):403–414, April 7, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519398908827.

Dibb:1991:PSS

[Dib91] N. J. Dibb. Proto-splice site model of intron origin. Journal of Theoretical Biology, 151(3):405–416, August 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305803881.

Dibrov:1998:RES

[Dib98] B. F. Dibrov. Resonance effect in self-renewing tissues. Journal of Theoretical Biology, 192(1):15–33, May 7, 1998. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519397906135.

Denesyuk:1996:SCH

[DJZK96] Alexander I. Denesyuk, Mark S. Johnson, Vladimir P. Zav’yalov, and Timo Korpela. Structure of cytokine hydrophobic cores. Journal of Theoretical Biology, 180(4):297–307, June 21, 1996. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396901046.

Duplisea:1995:ABS

[DK95] Daniel E. Duplisea and Stephen R. Kerr. Application of a biomass size spectrum model to demersal fish data from the scotian shelf. Journal of Theoretical Biology, 177(3):263–269, December 7, 1995. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519385702433.

Dorninger:1995:SCH

[DKL95] Dietmar Dorninger, Günther Karigl, and Josef Loidl. Simulation of chromosomal homology searching in meiotic pairing. Journal of Theoretical Biology, 176(2):247–260, September 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519385701956.

Dani:1990:DKA

[DL90] John A. Dani and David G. Levitt. Diffusion and kinetic approaches to describe permeation in ionic channels. Journal of Theoretical Biology, 146(3):289–301, October 7, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305807404.

Durrett:1996:SMS

[DL96] Rick Durrett and Simon Levin. Spatial models for speciesarea curves. Journal of Theoretical Biology, 179(2):119–127, March 14, 1996. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396900533.

Durrett:1997:ASD

[DL97]  Rick Durrett and Simon Levin.  Allelopathy in spa-

tially distributed populations. Journal of Theoretical Biology, 185(2):165–171, March 21, 1997. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519396902921.

deLussanet:1997:SMF

[dLA97] Marc H. E. de Lussanet and R. McNeill Alexander. A simple model for fast planar arm movements; optimising mechanical activation and moment-arms of uniarticular and biarticular arm muscles. Journal of Theoretical Biology, 184(2):187–201, January 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396902787.

Dodin:1996:TCD

[DLC96] G. Dodin, P. Levoir, and C. Cordier. Triplet correlation in DNA sequences and stability of heteroduplexes. Journal of Theoretical Biology, 183(3):341–343, December 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396902258.

Dokos:1998:RIM

[DLC98] Socrates Dokos, Nigel H. Lovell, and Branko G. Celler. Review of ionic models of vagal-cardiac pacemaker control. Journal of Theoretical Biology, 192(3):265–274, June 7, 1998. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519397906196.

deLaguerie:1993:EEF

[dLOG93] Patrick de Laguérie, Isabelle Olivieri, and Pierre-Henri Gouyon. Environmental effects on fitness-sets shape and evolutionarily stable strategies. Journal of Theoretical Biology, 163(1):113–125, July 1, 1993. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519383711100.

Derrick:1991:DLV

[DM91] William Derrick and Lee Metzgar. Dynamics of Lotka– Volterra systems with exploitation. Journal of Theoretical Biology, 153(4):455–468, December 21, 1991. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930580150X.

Day:1992:TCM

[DM92a] William H. E. Day and F. R. McMorris. Threshold consensus methods for molecular sequences. Journal of Theoretical Biology, 159(4):481–489, December 21, 1992. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305806927.

Doncaster:1992:OGS

[DM92b] C. Patrick Doncaster and David W. Macdonald. Optimum group size for defending heterogeneous distributions of resources: a model applied to red foxes, Vulpes vulpes, in Oxford city. Journal of Theoretical Biology, 159(2):189–198, November 21, 1992. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305807015.

Dietrich:1995:CSS

[DM95] F. Dietrich and C. Mattheck. Cracks in sabre-shaped trees and branches. Journal of Theoretical Biology, 173(3):321–327, April 7, 1995. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519395900012.

Drossel:1999:ECD

[DM99] Barbara Drossel and Alan McKane. Ecological character displacement in quantitative genetic models. Journal of Theoretical Biology, 196(3):363–376, February 7, 1999. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519398908475.

Dieckmann:1995:ECP

[DML95] Ulf Dieckmann, Paul Marrow, and Richard Law. Evolutionary cycling in predator–prey interactions: population dynamics and the red queen. Journal of Theoretical Biology, 176(1):91–102, September 7, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519385701798.

Detours:1999:QTA

[DMP99] Vincent Detours, Ramit Mehr, and Alan S. Perelson. A quantitative theory of affinity-driven T cell repertoire selection. Journal of Theoretical Biology, 200(4):389–403, October 21, 1999. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519399910032.

Dallon:1998:CAC

[DO98] John C. Dallon and Hans G. Othmer. A continuum analysis of the chemotactic signal seen by Dictyostelium discoideum. Journal of Theoretical Biology, 194(4):461–483, October 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519398907664.

Dillon:1999:MMO

[DO99] Robert Dillon and Hans G. Othmer. A mathematical model for outgrowth and spatial patterning of the vertebrate limb bud. Journal of Theoretical Biology, 197(3):295–330, April 7, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519398908761.

Doebeli:1994:ICP

[Doe94] Michael Doebeli. Intermittent chaos in population dynamics. Journal of Theoretical Biology, 166(3):325–330, February 7, 1994. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519384710290.

Doebeli:1995:EPI

[Doe95] Michael Doebeli. Evolutionary predictions from invariant physical measures of dynamic processes. Journal of Theoretical Biology, 173(4):377–387, April 21, 1995. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385700707.

Doebeli:1997:GVP

[Doe97] Michael Doebeli. Genetic variation and persistence of predator–prey interactions in the Nicholson-Bailey model. Journal of Theoretical Biology, 188(1):109–120, September 7, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519397904549.

Doig:1997:IEG

[Doi97] Andrew J. Doig. Improving the efficiency of the genetic code by varying the codon length — the perfect genetic code. Journal of Theoretical Biology, 188(3):355–360, October 7, 1997. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519397904896.

Doncaster:1990:NPE

[Don90] C. Patrick Doncaster. Non-parametric estimates of interaction from radio-tracking data. Journal of Theoretical Biology, 143(4):431–443, April 19, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305800207.

DeBoer:1994:CRC

[DP94a] Rob J. De Boer and Alan S. Perelson. T cell repertoires and competitive exclusion. Journal of Theoretical Biology, 169(4):375–390, August 21, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251938471160X.

Dickinson:1994:DAL

[DP94b] Heather Dickinson and Louise Parker. Do alcohol and lead change the sex ratio? Journal of Theoretical Biology, 169(3):313–315, August 7, 1994. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519384711520.

DeBoer:1995:TGF

[DP95a] Rob J. De Boer and A. S. Perelson. Towards a general function describing T cell proliferation. Journal of Theoretical Biology, 175(4):567–576, August 21, 1995. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385701658.

Dutta:1995:TAI

[DP95b] A. Dutta and A. S. Popel. A theoretical analysis of intracellular oxygen diffusion. Journal of Theoretical Biology, 176(4):433–445, October 21, 1995. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519385702111.

DeBoer:1998:TCL

[DP98] R. J. De Boer and A. S. Perelson. Target cell limited and immune control models of HIV infection: a comparison. Journal of Theoretical Biology, 190(3):201–214, February 7, 1998. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519397905488.

Dobek:1994:APL

[DPG^{+}94] A. Dobek, G. Paillotin, J. Gapiński, J. Breton, W. Leibl, and H.-W. Trissl. Amplitude and polarity of the light gradient photovoltage from chloroplasts. Journal of Theoretical Biology, 170(2):129–143, September 21, 1994. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384711738.

Dowse:1991:CBP

[DR91] Harold B. Dowse and John M. Ringo. Comparisons between “periodograms” and spectral analysis: Apples are apples after all. Journal of Theoretical Biology, 148(1):139–144, January 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305804680.

Drake:1990:MCA

[Dra90] James A. Drake. The mechanics of community assembly and succession. Journal of Theoretical Biology, 147(2):213–233, November 21, 1990. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305800530.

Dillon:1997:MCI

[DRB97] Patrick F. Dillon and Robert S. Root-Bernstein. Molecular complementarity II: Energetic and vectoral basis of biological homeostasis and its implications for death. Journal of Theoretical Biology, 188(4):481–493, October 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251939790477X.

Dexter:1991:MMC

[DRLB91] Franklin Dexter, Yoram Rudy, Matthew N. Levy, and Eugene N. Bruce. Mathematical model of cellular basis for the respiratory sinus arrhythmia. Journal of Theoretical Biology, 150(2):157–173, May 21, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305803297.

Druzinsky:1993:TAM

[Dru93] Robert E. Druzinsky. The time allometry of mammalian chewing movements: Chewing frequency scales with body mass in mammals. Journal of Theoretical Biology, 160(4):427–440, February 21, 1993. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519383710283.

Drummond:1997:PIN

[Dru97] R. Drummond. Positional information and non-transcribed DNA. Journal of Theoretical Biology, 188(1):103–108, September 7, 1997. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519397904562.

Drolet:1990:QTI

[DS90] Suzanne Drolet and David Sankoff. Quadratic tree invariants for multivalued characters. Journal of Theoretical Biology, 144(1):117–129, May 10, 1990. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305803030.

Dong:1992:LDF

[DS92] Cheng Dong and Richard Skalak. Leukocyte deformability: Finite element modeling of large viscoelastic deformation. Journal of Theoretical Biology, 158(2):173–174, IN1–IN2, 175–193, September 21, 1992. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305807167.

Deag:1999:CSA

[DS99] John M. Deag and Graham W. Scott. “Conventional” signals in avian agonistic displays: Integrating theory, data and different levels of analysis. Journal of Theoretical Biology, 196(2):155–162, January 21, 1999. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519398908256.

Duncan:1992:TSA

[DSD92] S. R. Duncan, Susan Scott, and C. J. Duncan. Time series analysis of oscillations in a model population: The effects of plague, pestilence and famine. Journal of Theoretical Biology, 158(3):293–311, October 7, 1992. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305807350.

Duncan:1993:HPS

[DSD93] S. R. Duncan, Susan Scott, and C. J. Duncan. An hypothesis for the periodicity of smallpox epidemics as revealed by time series analysis. Journal of Theoretical Biology, 160(2):231–248, January 21, 1993. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519383710167.

Duncan:1994:DFV

[DSD94a] S. R. Duncan, Susan Scott, and C. J. Duncan. Determination of a feedback vector that generates a non-decaying oscillation in a model population. Journal of Theoretical Biology, 167(1):67– 71, March 7, 1994. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519384710502.

Duncan:1994:PTS

[DSD94b] S. R. Duncan, Susan Scott, and C. J. Duncan. Predictions from time series analysis of the oscillations in Parish register series. Journal of Theoretical Biology, 168(1):95–103, May 7, 1994. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519384710903.

Dallon:1999:MME

[DSM99] John C. Dallon, Jonathan A. Sherratt, and Philip K. Maini. Mathematical modelling of extracellular matrix dynamics using discrete cells: Fiber orientation and tissue regeneration. Journal of Theoretical Biology, 199(4):449–471, August 21, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519399909712.

DeBoer:1992:PFO

[DSP92] Rob J. De Boer, Lee A. Segel, and Alan S. Perelson. Pattern formation in one- and two-dimensional shape-space models of the immune system. Journal of Theoretical Biology, 155(3):295–333, April 7, 1992. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305806010.

Detours:1996:SCI

[DSP96] Vincent Detours, Bernhard Sulzer, and Alan S. Perelson. Size and connectivity of the idiotypic network are independent of the discreteness of the affinity distribution. Journal of Theoretical Biology, 183(4):409–416, December 21, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519396902313.

DiGennaro:1996:ATM

[DST96] Patrizia Di Gennaro, Guido Sello, and Manuela Termini. Analysis of a theoretical model based on similarity for studying RNA base pairings. Journal of Theoretical Biology, 181(4):359–371, August 21, 1996. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251939690137X.

Day:1998:UGG

[DT98] Troy Day and Peter D. Taylor. Unifying genetic and game theoretic models of kin selection for continuous traits. Journal of Theoretical Biology, 194(3):391–407, October 7, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519398907627.

Dufton:1997:GCS

[Duf97] Mark J. Dufton. Genetic code synonym quotas and amino acid complexity: Cutting the cost of proteins? Journal of Theoretical Biology, 187(2):165–173, July 21, 1997. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519397904434.

Dugatkin:1990:PGE

[Dug90] Lee Alan Dugatkin. N-person games and the evolution of co-operation: a model based on predator inspection in fish. Journal of Theoretical Biology, 142(1):123–135, January 9, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305800177.

Dukas:1999:CMI

[Duk99] Reuven Dukas. Costs of memory: Ideas and predictions. Journal of Theoretical Biology, 197(1):41–50, March 7, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519398908566.

Dumortier:1994:CPT

[Dum94] B. Dumortier. The “circadian paradigm”: a test of involvement of the circadian system in the photoperiodic clock. Journal of Theoretical Biology, 166(1):101–112, January 7, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519384710083.

Dushoff:1996:III

[Dus96] Jonathan Dushoff. Incorporating immunological ideas in epidemiological models. Journal of Theoretical Biology, 180 (3):181–187, June 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519396900946.

Das:1993:NAF

[DvH93] M. Das and J. A. R. A. M. van Hooff. A non-adaptive feedback loop to account for the differences between the macaques’ social systems — is it convincing? Journal of Theoretical Biology, 160(3):399–402, February 7, 1993. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519383710258.

Dawson:1999:HTT

[DVJ^{+}99] Colin Dawson, Julian F. V. Vincent, George Jeronimidis, Graham Rice, and Paula Forshaw. Heat transfer through penguin feathers. Journal of Theoretical Biology, 199(3):291–295, August 7, 1999. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519399909591.

Dens:1999:PMS

[DVV99] E. J. Dens, K. M. Vereecken, and J. F. Van Impe. A prototype model structure for mixed microbial populations in homogeneous food products. Journal of Theoretical Biology, 201(3):159–170, December 7, 1999. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519399910214.

Dwyer:1998:AEU

[Dwy98] Donard S. Dwyer. Assembly of exons from unitary transposable genetic elements: Implications for the evolution of protein-protein interactions. Journal of Theoretical Biology, 194(1):11–27, September 7, 1998. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519398906762.

Dawson:1999:MFE

[DY99] Wayne K. Dawson and Kenji Yamamoto. Mean free energy topology for nucleotide sequences of varying composition based on secondary structure calculations. Journal of Theoretical Biology, 201(2):113–140, November 21, 1999. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519399910184.

Dexter:1992:AVI

[DYL92] Franklin Dexter, Tianen Yang, and Matthew N. Levy. Analysis of vagally induced sinus arrhythmias. Journal of Theoretical Biology, 159(4):507–511, December 21, 1992. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305806940.

Easterby:1991:HFC

[Eas91] J. S. Easterby. Homeostasis, flexibility and conflict in the kinetic advantage of channelling. Journal of Theoretical Biology, 152(1):47–48, September 7, 1991. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305805077.

Easterby:1996:FCA

[Eas96] John S. Easterby. The fusion of control analysis and temporal analysis of metabolic systems. Journal of Theoretical Biology, 182(3):327–331, October 7, 1996. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251939690171X.

Easton:1999:XRS

[Eas99] Dexter M. Easton. X-ray survival as Gompertz growth in number killed. Journal of Theoretical Biology, 196(1):1–8, January 7, 1999. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519398908128.

Exley:1992:CTA

[EB92] Christopher Exley and J. Derek Birchall. The cellular toxicity of aluminium. Journal of Theoretical Biology, 159(1):83–98, November 7, 1992. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305807696.

Exley:1991:MAA

[ECB91] C. Exley, J. S. Chappell, and J. D. Birchall. A mechanism for acute aluminium toxicity in fish. Journal of Theoretical Biology, 151(3):417–428, August 7, 1991. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305803893.

Edmonds:1996:LET

[Edm96] Donald T. Edmonds. Linear electric traction as an alternative model of the actin/myosin motor. Journal of Theoretical Biology, 179(1):55–60, March 7, 1996. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251939690048X.

Edwards:1996:MCO

[Edw96] Matthew R. Edwards. Metabolite channeling in the origin of life. Journal of Theoretical Biology, 179(4):313–322, April 21, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519396900703.

Erneux:1993:ESP

[EEG93] Thomas Erneux, Ronald D. Edstrom, and Albert Goldbeter. Enzyme sharing in phosphorylation-dephosphorylation cascades: The case where one protein kinase (or phosphatase) acts on two different substrates. Journal of Theoretical Biology, 165(1):43–61, November 7, 1993. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519383711768.

Ermentrout:1993:CAA

[EEK93] G. Bard Ermentrout and Leah Edelstein-Keshet. Cellular automata approaches to biological modeling. Journal of Theoretical Biology, 160(1):97–133, January 7, 1993. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519383710076.

Eshel:1998:LTE

[EFB98] Ilan Eshel, Marcus W. Feldman, and Aviv Bergman. Longterm evolution, short-term evolution, and population genetic theory. Journal of Theoretical Biology, 191(4):391–396, April 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251939790597X.

Elsner:1998:MCA

[EG98] Ludwig Elsner and Christoph Giersch. Metabolic control analysis: Separable matrices and interdependence of control coefficients. Journal of Theoretical Biology, 193(4):649–661, August 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519398907287.

Escobar-Gutierrez:1998:MAB

[EGDG^{+}98] A. J. Escobar-Gutiérrez, F.-A. Daudet, J.-P. Guadillère, P. Maillard, and J.-S. Frossard. Modelling of allocation and balance of carbon in walnut (Juglans regia L.) seedlings during heterotrophy-autotrophy transition. Journal of Theoretical Biology, 194(1):29–47, September 7, 1998. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519398907408.

Ezoe:1994:ACR

[EIU94] Hideo Ezoe, Yoh Iwasa, and Tamiki Umeda. Aggregation by chemotactic random walk: Drifting clusters and chemotactic friction. Journal of Theoretical Biology, 168(3):259–267, June 7, 1994. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519384711076.

Evans:1998:HMI

[EKK98] David A. Evans, John Klemer, and Stephen L. Kaattari. Heuristic models of the intermonomeric disulfide bonding process. Journal of Theoretical Biology, 195(4):505–524, December 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519398908098.

Engen:1996:PDM

[EL96] Steinar Engen and Russell Lande. Population dynamic models generating species abundance distributions of the gamma type. Journal of Theoretical Biology, 178(3):325–331, February 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396900284.

Engel:1997:DWE

[EL97] Jürgen Engel and Jürg Lamprecht. Doing what everybody does? a procedure for investigating behavioural synchronization. Journal of Theoretical Biology, 185(2):255–262, March 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519396903598.

Elder:1991:WCA

[Eld91a] D. Elder. Why is Crick’s adaptor an RNA and not a protein? Journal of Theoretical Biology, 148(2):287–288, January 21, 1991. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305803479.

Elder:1991:ESG

[Eld91b] David Elder. Evolution of split genes. Journal of Theoretical Biology, 152(3):427–428, October 7, 1991. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305802036.

Elder:1997:GIA

[Eld97] David Elder. Globin introns: Ancient or modern? Journal of Theoretical Biology, 187(1):77–80, July 7, 1997. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251939790415X.

Elitzur:1994:LTL

[Eli94] Avshalom C. Elitzur. Let there be life: Thermodynamic reflections on biogenesis and evolution. Journal of Theoretical Biology, 168(4):429–459, June 21, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519384711234.

Elitzur:1996:LEA

[Eli96] Avshalom C. Elitzur. Life’s emergence is not an axiom: a reply to yockey. Journal of Theoretical Biology, 180(2):175–180, May 21, 1996. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396900934.

Engen:1997:HSF

[ELS97] Steinar Engen, Russell Lande, and Bernt-Erik Sæther. Harvesting strategies for fluctuating populations based on uncertain population estimates. Journal of Theoretical Biology, 186(2):201–212, May 21, 1997. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519396903562.

Eisenberg:1995:SST

[EM95] Joseph N. Eisenberg and Don R. Maszle. The structural stability of a three-species food chain model. Journal of Theoretical Biology, 176(4):501–510, October 21, 1995. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385702160.

Emmans:1997:MPF

[Emm97] G. C. Emmans. A method to predict the food intake of domestic animals from birth to maturity as a function of time. Journal of Theoretical Biology, 186(2):189–199, May 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519396903574.

Elphick:1990:IPR

[EMRS90] C. Elphick, E. Meron, J. Rinzel, and E. A. Spiegel. Impulse patterning and relaxation propagation in excitable media. Journal of Theoretical Biology, 146(2):249–268, September 21, 1990. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305801389.

Eshel:1997:CSE

[EMS97] Ilan Eshel, Uzi Motro, and Emilia Sansone. Continuous stability and evolutionary convergence. Journal of Theoretical Biology, 185(3):333–343, April 7, 1997. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396903124.

Endresen:1997:CWC

[End97] Lars Petter Endresen. Chaos in weakly-coupled pacemaker cells. Journal of Theoretical Biology, 184(1):41–50, January 7, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519396902404.

Ennos:1993:SRA

[Enn93] A. R. Ennos. The scaling of root anchorage. Journal of Theoretical Biology, 161(1):61–75, March 7, 1993. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519383710404.

Enright:1990:CBP

[Enr90] J. T. Enright. Comparisons between periodograms and spectral analysis: Don’t expect apples to taste like oranges. Journal of Theoretical Biology, 143(3):425–430, April 5, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305800384.

Essunger:1994:MHI

[EP94] Paulina Essunger and Alan S. Perelson. Modeling HIV infection of CD4 + t-cell subpopulations. Journal of Theoretical Biology, 170(4):367–391, October 21, 1994. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384711994.

Exley:1997:HAA

[EPT97] Christopher Exley, John K. Pinnegar, and Helen Taylor. Hydroxyaluminosilicates and acute aluminium toxicity in fish. Journal of Theoretical Biology, 189(2):133–139, November 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519397905014.

Ebert:1994:AMI

[ER94] Thomas A. Ebert and Michael P. Russell. Allometry and model II non-linear regression. Journal of Theoretical Biology, 168(4):367–372, June 21, 1994. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519384711167.

Eshel:1995:OIC

[ES95] Ilan Eshel and Emilia Sansone. Owner-intruder conflict, Grafen effect and self-assessment. The bourgeois principle reexamined. Journal of Theoretical Biology, 177(4):341–356, December 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519385702524.

Erdi:1990:WWS

[ÉT90] Péter Érdi and János Tóth. What is and what is not stated by the may-wigner theorem? Journal of Theoretical Biology, 145(1):137–140, July 9, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305805417.

Eugenin:1995:GRR

[Eug95] J. Eugenín. Generation of the respiratory rhythm: modelling the inspiratory off switch as a neural integrator. Journal of Theoretical Biology, 172(2):107–120, January 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385700082.

Ensminger:1994:DNL

[EV94] Peter A. Ensminger and Michael Vinson. Deterministic nonlinear dynamics in phycomyces sporangiophores. Journal of Theoretical Biology, 170(3):259–266, October 7, 1994. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384711866.

Exley:1994:PAA

[EWHB94] C. Exley, A. J. Wicks, R. B. Hubert, and J. D. Birchall. Polynuclear aluminium and acute aluminium toxicity in the fish. Journal of Theoretical Biology, 167(4):415–416, April 21, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519384710812.

Exley:1996:KCA

[EWHB96] Christopher Exley, Alan J. Wicks, Rupert B. Hubert, and Derek J. Birchall. Kinetic constraints in acute aluminium toxicity in the rainbow trout (Oncorhynchus mykiss). Journal of Theoretical Biology, 179(1):25–31, March 7, 1996. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396900442.

Ezoe:1998:ODR

[Ezo98] Hideo Ezoe. Optimal dispersal range and seed size in a stable environment. Journal of Theoretical Biology, 190(3):287–293, February 7, 1998. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519397905531.

Forbes:1994:SMA

[FA94] W. F. Forbes and N. Agwani. A suggested mechanism for aluminum biotoxicity. Journal of Theoretical Biology, 171(2):207–214, November 21, 1994. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519384712240.

Flury:1992:GIW

[FAB92] Bernard D. Flury, Jean-Pierre Airoldi, and Jean-Pierre Biber. Gender identification of water pipits (Anthus spinoletta) using mixtures of distributions. Journal of Theoretical Biology, 158(4):465–480, October 21, 1992. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305807106.

Fadel:1992:STO

[Fad92] James G. Fadel. Simulation of traditional and optimal sampling schedule designs for digestion and passage models. Journal of Theoretical Biology, 157(1):109–126, July 7, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251930580760X.

Ford:1990:SBG

[FAF90] E. David Ford, Anne Avery, and R. Ford. Simulation of branch growth in the pinaceae: Interactions of morphology, phenology, foliage productivity, and the requirement for structural support, on the export of carbon. Journal of Theoretical Biology, 146(1):15–36, September 7, 1990. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305800426.

Feng:1995:KAP

[FB95] Y. Feng and L. Boersma. Kinematics of axial plant root growth. Journal of Theoretical Biology, 174(1):109–117, May 7, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519385700835.

Farnsworth:1997:BIF

[FB97] K. D. Farnsworth and J. A. Beecham. Beyond the ideal free distribution: More general models of predator distribution. Journal of Theoretical Biology, 187(3):389–396, August 7, 1997. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519397904367.

Fruchter:1991:SAO

[FBH91] Gila Fruchter and Shlomo Ben-Haim. Stability analysis of onedimensional dynamical systems applied to an isolated beating heart. Journal of Theoretical Biology, 148(2):175–192, January 21, 1991. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305803406.

Funk:1998:MMV

[FBHK98] G. A. Funk, A. D. Barbour, H. Hengartner, and U. Kalinke. Mathematical model of a virus-neutralizing immunglobulin response. Journal of Theoretical Biology, 195(1):41–52, November 7, 1998. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519398907792.

France:1995:IDM

[FBL^{+}95] J. France, B. J. Bequette, G. E. Lobley, J. A. Metcalf, D. Wray-Cahen, M. S. Dhanoa, F. R. C. Backwell, M. D. Hanigan, J. C. Macrae, and D. E. Beever. An isotope dilution model for partitioning leucine uptake by the bovine mammary gland. Journal of Theoretical Biology, 172(4):369–377, February 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519385700343.

France:1993:CSE

[FBS93] J. France, D. E. Beever, and R. C. Siddons. Compartmental schemes for estimating methanogenesis in ruminants from isotope dilution data. Journal of Theoretical Biology, 164(2):207–218, September 21, 1993. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519383711495.

Ferriere:1992:ESA

[FC92] Régis Ferrière and Jean Clobert. Evolutionarily stable age at first reproduction in a density-dependent model. Journal of Theoretical Biology, 157(2):253–267, July 21, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305806241.

Ferreira:1993:SSS

[FC93] Ricardo Ferreira and Kaline R. Coutinho. Simulation studies of self-replicating oligoribotides, with a proposal for the transition to a peptide-assisted stage. Journal of Theoretical Biology, 164(3):291–305, October 7, 1993. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519383711550.

Fusco:1995:NME

[FC95] Giuseppe Fusco and Quentin C. B. Cronk. A new method for evaluating the shape of large phylogenies. Journal of Theoretical Biology, 175(2):235–243, July 21, 1995. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385701361.

Faro:1997:FSP

[FCV97] Jose Faro, Jorge Carneiro, and Santiago Velasco. Further studies on the problem of immune network modelling. Journal of Theoretical Biology, 184(4):405–421, February 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519396902520.

Fryer:1999:SCB

[FCV99a] T. Fryer, C. Cannings, and G. T. Vickers. Sperm competition i: Basic model, ESS and dynamics. Journal of Theoretical Biology, 196(1):81–100, January 7, 1999. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519398908220.

Fryer:1999:SCI

[FCV99b] T. Fryer, C. Cannings, and G. T. Vickers. Sperm competition. II-post-copulatory guarding. Journal of Theoretical Biology, 197(3):343–360, April 7, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519398908797.

Furusawa:1992:PED

[FD92] Mitsuru Furusawa and Hirofumi Doi. Promotion of evolution: Disparity in the frequency of strand-specific misreading between the lagging and leading DNA strands enhances disproportionate accumulation of mutations. Journal of Theoretical Biology, 157(1):127–133, July 7, 1992. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305807611.

France:1993:MIG

[FDT^{+}93] J. France, M. S. Dhanoa, M. K. Theodorou, S. J. Lister, D. R. Davies, and D. Isac. A model to interpret gas accumulation profiles associated with in vitro degradation of ruminant feeds. Journal of Theoretical Biology, 163(1):99–111, July 1, 1993. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519383711094.

Feduccia:1993:AME

[Fed93] Alan Feduccia. Aerodynamic model for the early evolution of feathers provided by propithecus (primates, lemuridae). Journal of Theoretical Biology, 160(2):159–164, January 21, 1993. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519383710106.

Fell:1991:PSM

[Fel91a] David A. Fell. Physiological significance of metabolite channelling. Journal of Theoretical Biology, 152(1):109–110, September 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305805211.

Felsenstein:1991:CPI

[Fel91b] Joseph Felsenstein. Counting phylogenetic invariants in some simple cases. Journal of Theoretical Biology, 152(3):357–376, October 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305802000.

Feldman:1995:AME

[Fel95] Henry A. Feldman. On the allometric mass exponent, when it exists. Journal of Theoretical Biology, 172(2):187–197, January 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251938570015X.

Fell:1996:DHK

[Fel96] David Fell. Dr Henrik Kacser (1918–1995). Journal of Theoretical Biology, 182(3):193–194, October 7, 1996. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251939690154X.

Fernandez:1992:PCR

[Fer92] Ariel Fernández. A parallel computation revealing the role of the in vivo environment in shaping the catalytic structure of a mitochondrial RNA transcript. Journal of Theoretical Biology, 157(4):487–503, August 21, 1992. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305806654.

Fernandez:1994:SLR

[Fer94] Ariel Fernández. Stress localization in the RNA backbone: a mechanical footprint for predicting base-backbone tertiary contacts. Journal of Theoretical Biology, 166(4):443–452, February 21, 1994. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519384710393.

Ford:1990:SBE

[FF90] R. Ford and E. David Ford. Structure and basic equations of a simulator for branch growth in the pinaceae. Journal of Theoretical Biology, 146(1):1–13, September 7, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305800414.

Flaten:1992:PIA

[FG92] Trond Peder Flaten and Ralph M. Garruto. Polynuclear ions in aluminum toxicity. Journal of Theoretical Biology, 156(1):129– 132, May 7, 1992. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305806605.

Flierl:1999:IAI

[FGLO99] G. Flierl, D. Grünbaum, S. Levins, and D. Olson. From individuals to aggregations: the interplay between behavior and physics. Journal of Theoretical Biology, 196(4):397–454, February 21, 1999. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519398908426.

Frickenhaus:1999:KMP

[FH99] Stephan Frickenhaus and Reinhart Heinrich. Kinetic models and phenomenological analysis of passive lipid translocation in single-file. Journal of Theoretical Biology, 197(2):175–191, March 21, 1999. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519398908712.

France:1997:MQC

[FHB^{+}97] J. France, M. D. Hanigan, B. J. Bequette, J. A. Metcalf, D. E. Beevers, G. E. Lobley, J. C. MacRae, and F. R. C. Backwell. A model for quantifying the contribution of extracellularlyderived peptides to milk protein synthesis in the ruminant mammary gland. Journal of Theoretical Biology, 185(3):379– 388, April 7, 1997. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396903173.

France:1999:IDM

[FHR^{+}99] J. France, M. D. Hanigan, C. K. Reynolds, J. Dijkstra, L. A. Crompton, J. A. Maas, B. J. Bequette, J. A. Metcalf, G. E. Lobley, J. C. MacRae, and D. E. Beever. An isotope dilution model for partitioning leucine uptake by the liver of the lactating dairy cow. Journal of Theoretical Biology, 198(1):121– 133, May 7, 1999. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519399908974.

Fialkowski:1992:SSS

[Fia92] Konrad R. Fialkowski. Sympatric speciation: a simulation model of imperfect assortative mating. Journal of Theoretical Biology, 157(1):9–30, July 7, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305807544.

Fields:1991:MIC

[Fie91] Gregg B. Fields. A model for interstitial collagen catabolism by mammalian collagenases. Journal of Theoretical Biology, 153(4):585–602, December 21, 1991. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305801572.

Findlay:1992:PEU

[Fin92a] C. Scott Findlay. Phenotypic evolution under gene-culture transmission in structured populations. Journal of Theoretical Biology, 156(3):387–400, June 7, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305806824.

Finkel:1992:HAP

[Fin92b] David L. Finkel. HIV-1 ancestry primordial expansions of RRE and RRE related sequences. Journal of Theoretical Biology, 154(3):285–302, February 7, 1992. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305801717.

Fishman:1999:PIC

[Fis99] Michael A. Fishman. Predator inspection: Closer approach as a way to improve assessment of potential threats. Journal of Theoretical Biology, 196(2):225–235, January 21, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519398908347.

Fujitani:1997:MSD

[FK97] Y. Fujitani and I. Kobayashi. Mismatch-stimulated destruction of intermediates as an explanation for map expansion in genetic recombination. Journal of Theoretical Biology, 189(4):443–447, December 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519397905762.

Fogelson:1998:MBS

[FK98] Aaron L. Fogelson and Andrew L. Kuharsky. Membrane binding-site density can modulate activation thresholds in enzyme systems. Journal of Theoretical Biology, 193(1):1–18, July 7, 1998. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519398906701.

Fruh:1999:HST

[FK99] Thomas Früh and Winfried Kurth. The hydraulic system of trees: Theoretical framework and numerical simulation. Journal of Theoretical Biology, 201(4):251–270, December 21, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519399910287.

Fowler:1994:MFR

[FKJ^{+}94] A. C. Fowler, G. Kember, P. Johnson, S. J. Walter, P. Fleming, and M. Clements. A method for filtering respiratory oscillations. Journal of Theoretical Biology, 170(3):273–281, October 7, 1994. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251938471188X.

Fountain:1995:QTP

[FL95] K. R. Fountain and Charles M. Leland. Quantification of the three point receptor hypothesis of Cheng and Zee– Cheng. Journal of Theoretical Biology, 173(4):329–337, April 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519385700665.

Forbes:1996:IDA

[FL96] Scott L. Forbes and Timothy C. Lamey. Insurance, developmental accidents, and the risks of putting all your eggs in one basket. Journal of Theoretical Biology, 180(3):247–256, June 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396901009.

Flegr:1997:TDT

[Fle97] Jaroslav Flegr. Two distinct types of natural selection in turbidostat-like and chemostat-like ecosystems. Journal of Theoretical Biology, 188(1):121–126, September 7, 1997. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519397904586.

Flickinger:1999:HHR

[Fli99] R. A. Flickinger. Histone H1^{∘} ribosomal RNA synthesis and cell differentiation. Journal of Theoretical Biology, 196(4):521–525, February 21, 1999. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519398908530.

Flores:1998:MMN

[Flo98] J. C. Flores. A mathematical model for Neanderthal extinction. Journal of Theoretical Biology, 191(3):295–298, April 7, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519397905816.

Fernandez:1993:SIS

[FLR93] Ariel Fernández, Alfred S. Lewin, and Herschel Rabitz. Structure-induced strain determining the internal cyclization site in the yeast cobI5 autocatalytic intron: Theory and experimental tests. Journal of Theoretical Biology, 164(1):121–133, September 7, 1993. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519383711434.

Fisher:1991:EDA

[FM91] R. E. Fisher and H. D. Mayor. The evolution of defective and autonomous parvoviruses. Journal of Theoretical Biology, 149(4):429–439, April 21, 1991. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305800918.

Forbes:1998:POP

[FM98] L. Scott Forbes and Douglas W. Mock. Parental optimism and progeny choice: When is screening for offspring quality affordable. Journal of Theoretical Biology, 192(1):3–14, May 7, 1998. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519397905968.

Fukuchi:1992:EMP

[FO92] Satoshi Fukuchi and Jinya Otsuka. Evolution of metabolic pathways by chance assembly of enzyme proteins generated from sense and antisense strands of pre-existing genes. Journal of Theoretical Biology, 158(3):271–291, October 7, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305807349.

Fukuchi:1996:ESR

[FO96] Satoshi Fukuchi and Jinya Oysuka. Evolution of the selfreproducing system to the biosynthesis of the membrane: An approach from the amino acid sequence similarity in proteins. Journal of Theoretical Biology, 182(2):117–136, September 21, 1996. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396901472.

Fukuchi:1994:EGI

[FOO94] Satoshi Fukuchi, Tetsuya Okayama, and Jinya Otsuka. Evolution of genetic information flow from the viewpoint of protein sequence similarity. Journal of Theoretical Biology, 171(2):179–195, November 21, 1994. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519384712227.

Forbes:1990:IOE

[For90] L. Scott Forbes. Insurance offspring and the evolution of avian clutch size. Journal of Theoretical Biology, 147(3):345–359, December 7, 1990. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305804928.

Forbes:1991:OSN

[For91a] L. Scott Forbes. Optimal size and number of offspring in a variable environment. Journal of Theoretical Biology, 150(3):299– 304, June 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305804291.

Forsdyke:1991:EEM

[For91b] D. R. Forsdyke. Early evolution of MHC polymorphism. Journal of Theoretical Biology, 150(4):451–456, June 21, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305804394.

Forsdyke:1992:TSM

[For92] D. R. Forsdyke. Two signal model of self/not-self immune discrimination: An update. Journal of Theoretical Biology, 154(1):109–118, January 7, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305801912.

Forsdyke:1994:HSR

[For94a] D. R. Forsdyke. The heat-shock response and the molecular basis of genetic dominance. Journal of Theoretical Biology, 167(1):1–5, March 7, 1994. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519384710447.

Forsdyke:1994:RXC

[For94b] D. R. Forsdyke. Relationship of X chromosome dosage compensation to intracellular self/not-self discrimination: a resolution of Muller’s paradox? Journal of Theoretical Biology, 167(1):7– 12, March 7, 1994. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519384710459.

Forsdyke:1995:FTI

[For95] D. R. Forsdyke. Fine tuning of intracellular protein concentrations, a collective protein function involved in aneuploid lethality, sex-determination and speciation? Journal of Theoretical Biology, 172(4):335–345, February 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385700318.

Forsdyke:1996:DBS

[For96] D. R. Forsdyke. Different biological species “broadcast” their DNAs at different (g+c)% “wavelengths”. Journal of Theoretical Biology, 178(4):405–417, February 21, 1996. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396900387.

Forsdyke:1998:AWT

[For98] D. R. Forsdyke. An alternative way of thinking about stemloops in DNA. a case study of the human g0s2 gene. Journal of Theoretical Biology, 192(4):489–504, June 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519398906749.

Forsdyke:1999:TLI

[For99] D. R. Forsdyke. Two levels of information in DNA: Relationship of Romanes’ “intrinsic” variability of the reproductive system, and Bateson’s “residue” to the species-dependent component of the base composition, (c+g)%. Journal of Theoretical Biology, 201(1):47–61, November 7, 1999. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519399910135.

Fishman:1993:MCA

[FP93] Michael A. Fishman and Alan S. Perelson. Modeling T cell-antigen presenting cell interactions. Journal of Theoretical Biology, 160(3):311–342, February 7, 1993. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519383710210.

Fishman:1994:TTC

[FP94] Michael A. Fishman and Alan S. Perelson. Th1/Th2 cross regulation. Journal of Theoretical Biology, 170(1):25–56, September 7, 1994. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519384711660.

Fishman:1995:LMA

[FP95] Michael A. Fishman and Alan S. Perelson. Lymphocyte memory and affinity selection. Journal of Theoretical Biology, 173(3):241–262, April 7, 1995. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519395800034.

Fenske:1999:SIB

[FPB99] John J. Fenske, Michael H. Penner, and John P. Bolte. A simple individual-based model of insoluble polysaccharide hydrolysis: the potential for autosynergism with dual-activity glycosidases. Journal of Theoretical Biology, 199(1):113–118, July 7, 1999. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519399909384.

Foldes-Papp:1995:FDE

[FPPSK95] Zeno Földes-Papp, Wei-Guo Peng, Hartmut Seliger, and Albrecht K. Kleinschmidt. Fractal dimension of error sequence dynamics in quantitative modeling of syntheses of short oligonucleotide and single-stranded DNA sequences. Journal of Theoretical Biology, 174(4):391–408, June 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385701075.

Findlay:1990:CEE

[FR90] Scott Findlay and Glenn Rowe. Computer experiments on the evolution of sex: the haploid case. Journal of Theoretical Biology, 146(3):379–393, October 7, 1990. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305807489.

Fletcher:1993:EUS

[FR93] H. L. Fletcher and J. A. Rafferty. The effects of unequal sister chromatid exchange on length of arrays of repeated sequences. Journal of Theoretical Biology, 164(4):507–514, October 21, 1993. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519383711690.

Franchini:1999:MCD

[FR99] Patrick L. A. Franchini and Ronald E. Reid. A model for circular dichroism monitored dimerization and calcium binding in an EF-hand synthetic peptide. Journal of Theoretical Biology, 199(2):199–211, July 21, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519399909475.

Frank:1990:WCA

[Fra90] Steven A. Frank. When to copy or avoid an opponent’s strategy. Journal of Theoretical Biology, 145(1):41–46, July 9, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305805338.

Frank:1994:GME

[Fra94] Steven A. Frank. Genetics of mutualism: The evolution of altruism between species. Journal of Theoretical Biology, 170(4):393–400, October 21, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519384712008.

Frank:1995:GPC

[Fra95a] Steven A. Frank. George Price’s contributions to evolutionary genetics. Journal of Theoretical Biology, 175(3):373–388, August 7, 1995. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519385701488.

Frank:1995:OSS

[Fra95b] Steven A. Frank. The origin of synergistic symbiosis. Journal of Theoretical Biology, 176(3):403–410, October 7, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385702081.

Frantsevich:1995:OLD

[Fra95c] Leonid Frantsevich. Optimal leg design in a hexapod walker. Journal of Theoretical Biology, 175(4):561–566, August 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519385701646.

Frank:1997:CIP

[Fra97a] Steven A. Frank. Cytoplasmic incompatibility and population structure. Journal of Theoretical Biology, 184(3):327–330, February 7, 1997. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396902763.

Frank:1997:DAS

[Fra97b] Steven A. Frank. The design of adaptive systems: Optimal parameters for variation and selection in learning and development. Journal of Theoretical Biology, 184(1):31–39, January 7, 1997. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396902416.

Frank:1997:MAC

[Fra97c] Steven A. Frank. Multivariate analysis of correlated selection and kin selection, with an ESS maximization method. Journal of Theoretical Biology, 189(3):307–316, December 7, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519397905166.

Frantsevich:1997:AMS

[Fra97d] L. I. Frantsevich. Artifactual motility of the subcoxal axis in a model insect leg with skew joint axes. Journal of Theoretical Biology, 184(3):271–277, February 7, 1997. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396902611.

Frank:1998:DCI

[Fra98] Steven A. Frank. Dynamics of cytoplasmic incompatability with multiple wolbachia infections. Journal of Theoretical Biology, 192(2):213–218, May 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251939890652X.

Frank:1999:PQG

[Fra99] Steven A. Frank. Population and quantitative genetics of regulatory networks. Journal of Theoretical Biology, 197(3):281– 294, April 7, 1999. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519398908724.

Freeman:1990:ESN

[Fre90] Scott Freeman. The evolution of the scrotum: a new hypothesis. Journal of Theoretical Biology, 145(4):429–445, August 23, 1990. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305804795.

Freeman:1994:NNC

[Fre94] Walter J. Freeman. Neural networks and chaos. Journal of Theoretical Biology, 171(1):13–18, July 1994. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384712070.

Frean:1996:EDC

[Fre96] Marcus Frean. The evolution of degrees of cooperation. Journal of Theoretical Biology, 182(4):549–559, October 21, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519396901940.

Friedrich:1991:MCQ

[Fri91] Peter Friedrich. Metabolic compartmentation via “quenching enzymes”. Journal of Theoretical Biology, 152(1):115–116, September 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305805235.

Fridlyand:1998:ICA

[Fri98] Leonid E. Fridlyand. Independent changes of ATP/ADP or Δ pH could cause oscillations in photosynthesis. Journal of Theoretical Biology, 193(4):739–741, August 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519398907329.

Forrest:1994:DFT

[FS94] T. G. Forrest and Robert B. Suter. The discrete Fourier transform (DFT) in behavioural analysis. Journal of Theoretical Biology, 166(4):419–429, February 21, 1994. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251938471037X.

Ferretti:1995:PIM

[FS95a] V. Ferretti and D. Sankoff. Phylogenetic invariants for more general evolutionary models. Journal of Theoretical Biology, 173(2):147–162, March 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519385700525.

Friedman:1995:PGN

[FS95b] James W. Friedman and Avi Shmida. Pollination, gathering nectar and the distribution of flower species. Journal of Theoretical Biology, 175(2):127–138, July 21, 1995. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385701257.

Fontana:1998:SSP

[FS98] Walter Fontana and Peter Schuster. Shaping space: the possible and the attainable in RNA genotype-phenotype mapping. Journal of Theoretical Biology, 194(4):491–515, October 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519398907718.

France:1991:ACS

[FSD91] J. France, R. C. Siddons, and M. S. Dhanoa. Adaptation of compartmental schemes for interpreting isotope dilution data on volatile fatty acid metabolism in the rumen to the non-steady state and for single-dose injection. Journal of Theoretical Biology, 153(2):247–254, November 21, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305804254.

France:1991:UMA

[FSDT91] J. France, R. C. Siddons, M. S. Dhanoa, and J. H. M. Thornley. A unifying mathematical analysis of methods to estimate rumen volume using digesta markers and intraruminal sampling. Journal of Theoretical Biology, 150(2):145–155, May 21, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305803285.

Freist:1998:APB

[FSPC98] Wolfgang Freist, Hans Sternbach, Iancu Pardowitz, and Friedrich Cramer. Accuracy of protein biosynthesis: Quasispecies nature of proteins and possibility of error catastrophes. Journal of Theoretical Biology, 193(1):19–38, July 7, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519398906725.

Figureau:1999:SSP

[FST99] Alain Figureau, M. Angélica Soto, and José Tohá. Secondary structure of proteins and three-dimensional pattern recognition. Journal of Theoretical Biology, 201(2):103–111, November 21, 1999. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519399910160.

Friboulet:1993:RDC

[FT93] Alain Friboulet and Daniel Thomas. Reaction-diffusion coupling in a structured system: Application to the quantitative simulation of endplate currents. Journal of Theoretical Biology, 160(4):441–455, February 21, 1993. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519383710295.

France:1992:SSE

[FTBC92] J. France, J. H. M. Thornley, R. L. Baldwin, and K. A. Crist. On solving stiff equations with reference to simulating ruminant metabolism. Journal of Theoretical Biology, 156(4):525– 539, June 21, 1992. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305806423.

Feinberg:1990:GDM

[FTCM90] I. Feinberg, H. C. Thode, H. T. Chugani, and J. D. March. Gamma distribution model describes maturational curves for delta wave amplitude, cortical metabolic rate and synaptic density. Journal of Theoretical Biology, 142(2):149–161, January 23, 1990. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305802188.

France:1990:UZC

[FTD90] J. France, M. K. Theodorou, and D. Davies. Use of zoospore concentrations and life cycle parameters in determining the population of anaerobic fungi in the rumen ecosystem. Journal of Theoretical Biology, 147(3):413–422, December 7, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305804965.

France:1990:TCM

[FTL^{+}90] J. France, J. H. M. Thornley, S. Lopez, R. C. Siddons, M. S. Dhanoa, P. J. Van Soest, and M. Gill. On the two-compartment model for estimating the rate and extent of feed degradation in the rumen. Journal of Theoretical Biology, 146(2):269–287, September 21, 1990. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305801390.

Fu:1995:LIU

[Fu95] Yun-Xin Fu. Linear invariants under jukes’ and Cantor’s oneparameter model. Journal of Theoretical Biology, 173(4):339– 352, April 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519385700677.

Faro:1993:SRC

[FV93] Jose Faro and Santiago Velasco. Studies on a recent class of network models of the immune system. Journal of Theoretical Biology, 164(3):271–290, October 7, 1993. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519383711549.

Faro:1994:NAM

[FV94] Jose Faro and Santiago Velasco. Numerical analysis of a model of ligand-induced b-cell antigen-receptor clustering. implications for simple models of b-cell activation in an immune network. Journal of Theoretical Biology, 167(1):45–53, March 7, 1994. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519384710484.

Farina:1997:SIA

[FVB97] Richard A. Fariña, Sergio F. Vizcaíno, and R. Ernesto Blanco. Scaling of the indicator of athletic capability in fossil and extant land tetrapods. Journal of Theoretical Biology, 185(4):441–446, April 21, 1997. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396903239.

Freeman:1990:TCH

[FW90] Raymond A. Freeman and James D. Wilson. A two-channel hypothesis for regulation of cell division and differentiation. Journal of Theoretical Biology, 146(3):303–315, October 7, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305807416.

Friedrich:1997:MER

[FW97] Thorsten Friedrich and Hanns Weiss. Modular evolution of the respiratory NADH:Ubiquinone oxidoreductase and the origin of its modules. Journal of Theoretical Biology, 187(4):529–540, August 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396903872.

Fenton:1997:SAD

[FWF97] Andrew Fenton, Richard Wall, and Nigel French. Sensitivity analysis of deterministic and stochastic simulation models of populations of the sheep blowfly, lucilia sericata. Journal of Theoretical Biology, 184(2):139–148, January 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519396902556.

Gaal:1993:CMD

[Gaá93] Gyöngyi Gaál. Calculation of movement direction from firing activities of neurons in intrinsic co-ordinate systems defined by their preferred directions. Journal of Theoretical Biology, 162(1):103–130, May 7, 1993. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519383710799.

Gruntfest:1997:FPV

[GAD97] Yana Gruntfest, Roger Arditi, and Yuri Dombrovsky. A fragmented population in a varying environment. Journal of Theoretical Biology, 185(4):539–547, April 21, 1997. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396903586.

Gaggiotti:1994:EMM

[Gag94] Oscar E. Gaggiotti. An ecological model for the maintenance of sex and geographic parthenogenesis. Journal of Theoretical Biology, 167(3):201–221, April 7, 1994. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384710642.

Galis:1992:MBP

[Gal92] Frietson Galis. A model for biting in the pharyngeal jaws of a cichlid fish: Haplochromis piceatus. Journal of Theoretical Biology, 155(3):343–368, April 7, 1992. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305806034.

Ganti:1997:BI

[Gán97] Tibor Gánti. Biogenesis itself. Journal of Theoretical Biology, 187(4):583–593, August 21, 1997. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519396903914.

Gandon:1999:KCC

[Gan99] Sylvain Gandon. Kin competition, the cost of inbreeding and the evolution of dispersal. Journal of Theoretical Biology, 200(4):345–364, October 21, 1999. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519399909943.

Gonzalez-Andujar:1995:RCB

[GAP95] J. L. González-Andújar and J. N. Perry. Reversals of chaos in biological control systems. Journal of Theoretical Biology, 175(4):603, August 21, 1995. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519385701695.

Gatenby:1995:MTH

[Gat95] Robert A. Gatenby. Models of tumor-host interaction as competing populations: Implications for tumor biology and treatment. Journal of Theoretical Biology, 176(4):447–455, October 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519385702123.

Gavrilets:1997:CCE

[Gav97] Sergey Gavrilets. Coevolutionary chase in exploiter–victim systems with polygenic characters. Journal of Theoretical Biology, 186(4):527–534, June 21, 1997. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519397904264.

Godfray:1995:PDP

[GB95] H. C. J. Godfray and C. J. Briggs. The population dynamics of pathogens that control insect outbreaks. Journal of Theoretical Biology, 176(1):125–136, September 7, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385701828.

Ghosh:1999:ENC

[GBD99] Subhendu Ghosh, Amal K. Bera, and Sudipto Das. Evidence for nonlinear capacitance in biomembrane channel system. Journal of Theoretical Biology, 200(3):299–305, October 7, 1999. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519399909931.

Gutfreund:1991:SCA

[GC91] H. Gutfreund and P. B. Chock. Substrate channeling among glycolytic enzymes: Fact or fiction. Journal of Theoretical Biology, 152(1):117–121, September 7, 1991. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305805247.

Godin:1998:MMP

[GC98] C. Godin and Y. Caraglio. A multiscale model of plant topological structures. Journal of Theoretical Biology, 191(1):1–46, March 7, 1998. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519397905610.

Giersch:1996:EDM

[GCB96] Christoph Giersch and Athel Cornish-Bowden. Extending double modulation: Combinatorial rules for identifying the modulations necessary for determining elasticities in metabolic pathways. Journal of Theoretical Biology, 182(3):361–369, October 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396901757.

Garcia-Dorado:1990:SEP

[GD90] Aurora García-Dorado. Some evolutionary properties of parental investment per offspring in a heterogeneous environment. Journal of Theoretical Biology, 147(1):101–114, November 7, 1990. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305802541.

Griffith:1990:BEA

[GE90] T. M. Griffith and D. H. Edwards. Basal EDRF activity helps to keep the geometrical configuration of arterial bifurcations close to the Murray optimum. Journal of Theoretical Biology, 146(4):545–573, October 21, 1990. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305803789.

Gutierrez-Espeleta:1996:STO

[GE96] Edgar E. Gutiérrez-Espeleta. Some thoughts on ordination techniques, ecological study units, and the method of synecological coordinates. Journal of Theoretical Biology, 179(1):13– 23, March 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396900430.

Getz:1993:IMA

[Get93] Wayne M. Getz. Invasion and maintenance of alleles that influence mating and parental success. Journal of Theoretical Biology, 162(4):515–537, June 21, 1993. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519383711021.

Groen:1996:CAB

[GET96] Albert K. Groen, Ronald P. J. Oude Elferink, and Joseph M. Tager. Control analysis of biliary lipid secretion. Journal of Theoretical Biology, 182(3):427–436, October 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396901836.

Gavrilets:1997:PFH

[GG97a] Sergey Gavrilets and Janko Gravner. Percolation on the fitness hypercube and the evolution of reproductive isolation. Journal of Theoretical Biology, 184(1):51–64, January 7, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519396902428.

Gubbins:1997:PHP

[GG97b] Simon Gubbins and Christopher A. Gilligan. Persistence of host-parasite interactions in a disturbed environment. Journal of Theoretical Biology, 188(2):241–258, September 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519397904665.

Guth:1998:MFH

[GG98] Sandra Güth and Werner Güth. Male fights of hamadryas baboons. Journal of Theoretical Biology, 190(1):1–14, January 7, 1998. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519397905208.

Gallinet:1993:MAQ

[GGML93] J. P. Gallinet, J. J. Girardot, J. Ph. Massonie, and C. Lance. A method for the assessment of the quality of mitochondrial preparations. Journal of Theoretical Biology, 161(4):431–460, April 21, 1993. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519383710660.

Gordon:1992:PDM

[GGT92] Deborah M. Gordon, Brian C. Goodwin, and L. E. H. Trainor. A parallel distributed model of the behaviour of ant colonies. Journal of Theoretical Biology, 156(3):293–307, June 7, 1992. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305806770.

Godfray:1990:EBR

[GH90] H. C. J. Godfray and A. B. Harper. The evolution of brood reduction by siblicide in birds. Journal of Theoretical Biology, 145(2):163–175, July 24, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305801225.

Gabriel:1991:MMP

[GH91] Jerome L. Gabriel and J. Kenneth Hoober. Molecular modeling of phytochrome. Journal of Theoretical Biology, 151(4):541–556, August 21, 1991. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305803698.

Greenman:1997:ICI

[GH97] J. V. Greenman and P. J. Hudson. Infected coexistence instability with and without density-dependent regulation. Journal of Theoretical Biology, 185(3):345–356, April 7, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396903094.

Gavrilets:1998:CCT

[GH98a] Sergey Gavrilets and Alan Hastings. Coevolutionary chase in two-species systems with applications to mimicry. Journal of Theoretical Biology, 191(4):415–427, April 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519397906159.

Grover:1998:DRA

[GH98b] James P. Grover and Robert D. Holt. Disentangling resource and apparent competition: Realistic models for plant-herbivore communities. Journal of Theoretical Biology, 191(4):353–376, April 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519397905622.

Ghosh:1993:RMC

[Gho93] Subhendu Ghosh. Relaxation of membrane channels: a statistical mechanical approach. Journal of Theoretical Biology, 165(2):171–176, November 21, 1993. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519383711823.

Goldberg:1999:LFT

[GHW99] Leslie Ann Goldberg, William E. Hart, and David Bruce Wilson. Learning foraging thresholds for lizards: an analysis of a simple learning algorithm. Journal of Theoretical Biology, 197(3):361–369, April 7, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519398908773.

Giersch:1991:PS

[Gie91] Christoph Giersch. Pillow strategy. Journal of Theoretical Biology, 152(1):71, September 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305805132.

Giersch:1994:DEM

[Gie94] Christoph Giersch. Determining elasticities from multiple measurements of steady-state flux rates and metabolite concentrations: Theory. Journal of Theoretical Biology, 169(1):89–99, July 7, 1994. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519384711313.

Giersch:1997:CRC

[Gie97] Christoph Giersch. Co-response coefficients, monovalent units, and combinatorial rules: Unification of concepts in metabolic control analysis. Journal of Theoretical Biology, 189(1):1–9, November 7, 1997. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519397905427.

Gilpin:1992:DSM

[Gil92] Michael Gilpin. Demographic stochasticity: a Markovian approach. Journal of Theoretical Biology, 154(1):1–8, January 7, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305801833.

Gingerich:1995:SPE

[Gin95] Philip D. Gingerich. Statistical power of EDF tests of normality and the sample size required to distinguish geometricnormal (lognormal) from arithmetic-normal distributions of low variability. Journal of Theoretical Biology, 173(2):125–136, March 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519385700501.

Goodwin:1990:SHP

[GK90] Brian C. Goodwin and Stuart A. Kauffman. Spatial harmonics and pattern specification in early Drosophila development. Part I. Bifurcation sequences and gene expression. Journal of Theoretical Biology, 144(3):303–319, June 7, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305800785.

Gosselin:1991:VWC

[GK91] R. E. Gosselin and S. M. Kaplow. Venous waterfalls in coronary circulation. Journal of Theoretical Biology, 149(2):265–279, March 21, 1991. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305802814.

Goodwin:1993:MIR

[GKM93] Brian C. Goodwin, Stuart Kauffman, and J. D. Murray. Is morphogenesis an intrinsically robust process? Journal of Theoretical Biology, 163(1):135–144, July 1, 1993. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519383711124.

Goodall:1990:NSC

[GKMSS90] C. Goodall, S. H. Kim, L. Moss-Salentijn, and M. Shinozuka. Numerical simulation of the crown of an incisiform tooth by conformal and polynomial regression mapping of a simple model. Journal of Theoretical Biology, 147(4):433–448, December 21, 1990. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305802589.

Gueron:1993:SOF

[GL93] Shay Gueron and Simon A. Levin. Self-organization of front patterns in large wildebeest herds. Journal of Theoretical Biology, 165(4):541–552, December 21, 1993. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519383712063.

Gerard:1995:SER

[GL95] Jean-François Gerard and Patrice Loisel. Spontaneous emergence of a relationship between habitat openness and mean group size and its possible evolutionary consequences in large herbivores. Journal of Theoretical Biology, 176(4):511–522, October 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519385702172.

Glass:1991:SRH

[Gla91] John D. Glass. A sequence related to the human gonadoliberin precursor near the n-termini of HIV and SIV gag polyproteins. Journal of Theoretical Biology, 150(4):489–496, June 21, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305804424.

Guan:1997:DAD

[GLH97] Shan Guan, Qishao Lu, and Kelei Huang. A discussion about the DiFrancesco-noble model. Journal of Theoretical Biology, 189(1):27–32, November 7, 1997. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519397904860.

Glick:1991:PMA

[Gli91] J. Leslie Glick. Proposed mechanism for alteration of albumin structure and function in Alzheimer’s disease. Journal of Theoretical Biology, 148(2):283–286, January 21, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305803467.

Gandhi:1998:CST

[GLO98] Amar Gandhi, Simon Levin, and Steven Orszag. “Critical slowing down” in time-to-extinction: an example of critical phenomena in ecology. Journal of Theoretical Biology, 192(3):363– 376, June 7, 1998. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519398906609.

Gandhi:1999:NRM

[GLO99] Amar Gandhi, Simon Levin, and Steven Orszag. Nucleation and relaxation from meta-stability in spatial ecological models. Journal of Theoretical Biology, 200(2):121–146, September 21, 1999. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519399909785.

Gueron:1996:DHI

[GLR96] Shay Gueron, Simon A. Levin, and Daniel I. Rubenstein. The dynamics of herds: From individuals to aggregations. Journal of Theoretical Biology, 182(1):85–98, September 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519396901447.

Glass:1990:CML

[GM90] Leon Glass and C. P. Malta. Chaos in multi-looped negative feedback systems. Journal of Theoretical Biology, 145(2):217– 223, July 24, 1990. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305801274.

Ghosh:1993:SMM

[GM93] Subhendu Ghosh and Amitabha Mukherjee. Statistical mechanics of membrane channels. Journal of Theoretical Biology, 160(2):151–157, January 21, 1993. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251938371009X.

Gelman:1994:DTW

[GM94] N. Gelman and J. Machin. Diffusion through the water barrier of arthropod cuticles: a statistical mechanical approach to the analysis of temperature effects. Journal of Theoretical Biology, 167(4):361–372, April 21, 1994. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384710757.

Gatherer:1998:USC

[GM98] Derek Gatherer and Neil R. McEwan. On units of selection in cultural evolution. Journal of Theoretical Biology, 192(3):409– 413, June 7, 1998. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251939890666X.

Gandon:1999:ESD

[GM99] Sylvain Gandon and Yannis Michalakis. Evolutionarily stable dispersal rate in a metapopulation with extinctions and kin competition. Journal of Theoretical Biology, 199(3):275–290, August 7, 1999. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519399909608.

Grossniklaus:1996:NEK

[GMN96] Ueli Grossniklaus, M. S. Madhusudhan, and Vidyanand Nanjundiah. Nonlinear enzyme kinetics can lead to high metabolic flux control coefficients: Implications for the evolution of dominance. Journal of Theoretical Biology, 182(3):299–302, October 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396901678.

Glazko:1998:SAM

[GMR98] Galina B. Glazko, Luciano Milanesi, and Igor B. Rogozin. The subclass approach for mutational spectrum analysis: Application of the SEM algorithm. Journal of Theoretical Biology, 192(4):475–487, June 21, 1998. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519398906683.

Gaffney:1999:MMC

[GMST99] E. A. Gaffney, P. K. Maini, J. A. Sherratt, and S. Tuft. The mathematical modelling of cell kinetics in corneal epithelial wound healing. Journal of Theoretical Biology, 197(1):15–40, March 7, 1999. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519398908529.

Gundersen:1995:CSP

[GMU95] Cameron B. Gundersen, Alessandro Mastrogiacomo, and Joy A. Umbach. Cysteine-string proteins as templates for membrane fusion: models of synaptic vesicle exocytosis. Journal of Theoretical Biology, 172(3):269–277, February 7, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385700239.

Green:1995:LBS

[GN95] Richard F. Green and David L. G. Noakes. Is a little bit of sex as good as a lot? Journal of Theoretical Biology, 174(1):87– 96, May 7, 1995. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519385700811.

Gutierrez:1994:OPT

[GOM94] G. Gutiérrez, J. L. Oliver, and A. Marín. On the origin of the periodicity of three in protein coding DNA sequences. Journal of Theoretical Biology, 167(4):413–414, April 21, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519384710800.

Gordon:1995:WRR

[Gor95] Karl H. J. Gordon. Were RNA replication and translation directly coupled in the RNA (+protein?) world? Journal of Theoretical Biology, 173(2):179–193, March 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385700549.

Goss:1992:ERA

[Gos92] Richard J. Goss. The evolution of regeneration: Adaptive or inherent? Journal of Theoretical Biology, 159(2):241–260, November 21, 1992. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305807040.

Goto:1994:MNA

[Got94] Kunihiko Goto. The on-off mechanisms of the nicotinic acetylcholine receptor ion channel are performed by thermodynamic forces. Journal of Theoretical Biology, 170(3):267–272, October 7, 1994. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519384711878.

Goulson:1994:MPI

[Gou94] D. Goulson. A model to predict the influence of insect flower constancy on interspecific competition between insect pollinated plants. Journal of Theoretical Biology, 168(3):309–314, June 7, 1994. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519384711118.

Gromiha:1993:RBA

[GP93] M. Michael Gromiha and P. K. Ponnuswamy. Relationship between amino acid properties and protein compressibility. Journal of Theoretical Biology, 165(1):87–100, November 7, 1993. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519383711781.

Glick:1998:MLP

[GPL98] Bernard R. Glick, Donna M. Penrose, and Jiping Li. A model for the lowering of plant ethylene concentrations by plant growth-promoting bacteria. Journal of Theoretical Biology, 190(1):63–68, January 7, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519397905324.

Guerreiro:1995:GAB

[GR95] J. Guerreiro and F. Rothen. A global approach to botanic patterns. Journal of Theoretical Biology, 176(2):233–245, September 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519385701944.

Grafen:1996:STD

[GR96] Alan Grafen and Mark Ridley. Statistical tests for discrete cross-species data. Journal of Theoretical Biology, 183(3):255–267, December 7, 1996. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519396902192.

Godelle:1997:EDS

[GR97a] B. Godelle and X. Reboud. The evolutionary dynamics of selfish replicators: a two-level selection model. Journal of Theoretical Biology, 185(3):401–413, April 7, 1997. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396903604.

Grafen:1997:NMD

[GR97b] Alan Grafen and Mark Ridley. A new model for discrete character evolution. Journal of Theoretical Biology, 184(1):7–14, January 7, 1997. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396902696.

Grafen:1997:NIS

[GR97c] Alan Grafen and Mark Ridley. Non-independence in statistical tests for discrete cross-species data. Journal of Theoretical Biology, 188(4):507–514, October 21, 1997. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519397904847.

Gross:1998:SIC

[GR98] Mart R. Gross and Joe Repka. Stability with inheritance in the conditional strategy. Journal of Theoretical Biology, 192(4):445–453, June 21, 1998. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519398906658.

Grafen:1990:BSH

[Gra90a] Alan Grafen. Biological signals as handicaps. Journal of Theoretical Biology, 144(4):517–546, June 21, 1990. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305800888.

Grafen:1990:SSU

[Gra90b] Alan Grafen. Sexual selection unhandicapped by the fisher process. Journal of Theoretical Biology, 144(4):473–516, June 21, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305800876.

Grafen:1992:UPR

[Gra92] Alan Grafen. The uniqueness of the phylogenetic regression. Journal of Theoretical Biology, 156(4):405–423, June 21, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305806356.

Graner:1993:CSAa

[Gra93] François Graner. Can surface adhesion drive cell-rearrangement? part i: Biological cell-sorting. Journal of Theoretical Biology, 164(4):455–476, October 21, 1993. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519383711677.

Graham:1995:TGC

[Gra95a] Geoffrey J. Graham. Tandem genes and clustered genes. Journal of Theoretical Biology, 175(1):71–87, July 7, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385701221.

Gratch:1995:FDC

[Gra95b] A. V. Gratch. A formal description of cell lineage and a hypothesis: all cell lineages are fixed. Journal of Theoretical Biology, 174(3):237–260, June 7, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519385700963.

Grafen:1998:FLS

[Gra98] Alan Grafen. Fertility and labour supply in femina economica. Journal of Theoretical Biology, 194(3):429–455, October 7, 1998. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519398907962.

Gregorius:1992:TLM

[Gre92] Hans-Rolf Gregorius. A two-locus model of speciation. Journal of Theoretical Biology, 154(3):391–398, February 7, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930580178X.

Gregorius:1997:ASD

[Gre97] Hans-Rolf Gregorius. The adaptational system as a dynamical feedback system. Journal of Theoretical Biology, 189(1):97–105, November 7, 1997. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519397904999.

Green:1998:RMG

[Gre98] Michael E. Green. A resonance model gives the response to membrane potential for an ion channel. Journal of Theoretical Biology, 193(3):475–483, August 7, 1998. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519398907159.

Gribble:1994:CMR

[Gri94a] N. Gribble. A conceptual model for the regulation of cardiac and ventilatory activity of the portunid crab, portunus pelagicus (linnaeus). Journal of Theoretical Biology, 168(1):19–29, May 7, 1994. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519384710848.

Griffin:1994:IGC

[Gri94b] Lewis D. Griffin. The intrinsic geometry of the cerebral cortex. Journal of Theoretical Biology, 166(3):261–273, February 7, 1994. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519384710241.

Grim:1995:GGS

[Gri95] Patrick Grim. The greater generosity of the spatialized prisoner’s dilemma. Journal of Theoretical Biology, 173(4):353– 359, April 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519385700689.

Grover:1992:CVY

[Gro92] James P. Grover. Constant- and variable-yield models of population growth: Responses to environmental variability and implications for competition. Journal of Theoretical Biology, 158(4):409–428, October 21, 1992. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305807076.

Geevan:1990:MMI

[GRRB90] C. P. Geevan, J. Subba Rao, Gita Subba Rao, and J. S. Bajaj. A mathematical model for insulin kinetics III. sensitivity analysis of the model. Journal of Theoretical Biology, 147(2):255–263, November 21, 1990. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305800554.

Guisez:1993:FMC

[GRRF93] Yves Guisez, Johan Robbens, Erik Remaut, and Walter Fiers. Folding of the MS2 coat protein in Escherichia coli is modulated by translational pauses resulting from mRNA secondary structure and codon usage: a hypothesis. Journal of Theoretical Biology, 162(2):243–252, May 21, 1993. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519383710854.

Grunbaum:1995:MFC

[Grü95] Daniel Grünbaum. A model of feeding currents in encrusting bryozoans shows interference between zooids within a colony. Journal of Theoretical Biology, 174(4):409–410, IN1, 411–425, June 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519385701087.

Gowrie:1999:EPT

[GRUH99] Ian J. Gowrie, Abdul V. Roudsari, A. Margot Umpleby, and Roman Hovorka. Estimating protein turnover with a [ 13 n, 15 c]leucine tracer: a study using simulated data. Journal of Theoretical Biology, 198(2):165–172, May 21, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519399909074.

Graner:1993:CSAb

[GS93] François Graner and Yasuji Sawada. Can surface adhesion drive cell rearrangement? part II: a geometrical model. Journal of Theoretical Biology, 164(4):477–506, October 21, 1993. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519383711689.

Giske:1995:WPP

[GS95a] Jarl Giske and Anne Gro Vea Salvanes. Why pelagic planktivores should be unselective feeders. Journal of Theoretical Biology, 173(1):41–50, March 7, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305800037.

Godfrey:1995:FIW

[GS95b] Laurie R. Godfrey and Michael R. Sutherland. Flawed inference: why size-based tests of heterochronic processes do not work. Journal of Theoretical Biology, 172(1):43–61, January 7, 1995. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519385700045.

Gimenez:1998:FMN

[GSS98] Juan A. Martínez Giménez, Guillermo T. Sáez, and Rafael Tabarés Seisdedos. On the function of modified nucleosides in the RNA world. Journal of Theoretical Biology, 194(4):485–490, October 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519398907706.

Green:1997:ABQ

[GT97] H. S. Green and T. Triffet. The animal brain as a quantal computer. Journal of Theoretical Biology, 184(4):385–403, February 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396902490.

Guilford:1994:GSS

[Gui94] Tim Guilford. “Go-slow” signalling and the problem of automimicry. Journal of Theoretical Biology, 170(3):311–316, October 7, 1994. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519384711921.

Gotts:1991:AIM

[GV91] N. M. Gotts and F. Vollrath. Artificial intelligence modelling of web-building in the garden cross spider. Journal of Theoretical Biology, 152(4):485–511, October 21, 1991. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305803947.

Garay:1998:ESA

[GV98] József Garay and Zoltán Varga. Evolutionarily stable allele distributions. Journal of Theoretical Biology, 191(2):163–172, March 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519397905774.

Garay:1999:RAS

[GV99] József Garay and Zoltán Varga. Relative advantage: a substitute for mean fitness in Fisher’s fundamental theorem? Journal of Theoretical Biology, 201(3):215–218, December 7, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519399910202.

Gerard:1993:SPP

[GVL93] Jean-François Gerard, Michel Vancassel, and Brigitte Laffort. Spread of phenotypic plasticity or genetic assimilation: The possible role of genetic constraints. Journal of Theoretical Biology, 164(3):341–349, October 7, 1993. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519383711586.

Goldbeter:1990:CMP

[GW90] A. Goldbeter and L. Wolpert. Covalent modification of proteins as a threshold mechanism in development. Journal of Theoretical Biology, 142(2):243–250, January 23, 1990. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305802255.

Grover:1995:RBF

[GW95] N. B. Grover and C. L. Woldringh. Relationship between the fraction of cells of different genealogical ages and their cycle times in saccharomyces cerevisiae: a theoretical analysis. Journal of Theoretical Biology, 174(2):223–225, May 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251938570094X.

Gibb:1994:ESP

[GWL94] William J. Gibb, Mary B. Wagner, and Michael D. Lesh. Effects of simulated potassium blockade on the dynamics of triggered cardiac activity. Journal of Theoretical Biology, 168(3):245–257, June 7, 1994. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519384711064.

Gow:1990:PLB

[GWS90] Alexander Gow, Donald J. Winzor, and Ross Smith. Preferential ligand binding to multi-state acceptor systems: the unexplored paradox of acceptor self-association that is ligandmediated but detrimental to ligand binding. Journal of Theoretical Biology, 145(3):407–420, August 9, 1990. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305801195.

Hansberg:1990:HSC

[HA90] Wilhelm Hansberg and Jesus Aguirre. Hyperoxidant states cause microbial cell differentiation by cell isolation from dioxygen. Journal of Theoretical Biology, 142(2):201–221, January 23, 1990. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251930580222X.

Haberl:1992:SHP

[HA92] Helmut Haberl and Hans Peter Aubauer. Simulation of human population dynamics by a hyperlogistic time-delay equation. Journal of Theoretical Biology, 156(4):499–511, June 21, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251930580640X.

Hedenstrom:1997:OFL

[HA97] Anders Hedenström and Thomas Alerstam. Optimum fuel loads in migratory birds: Distinguishing between time and energy minimization. Journal of Theoretical Biology, 189(3):227–234, December 7, 1997. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519397905051.

Hageseth:1993:FCE

[Hag93]  G. T. Hageseth.  The four-compartment energy level

diagram for seed germination. Journal of Theoretical Biology, 160(1):73–84, January 7, 1993. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519383710052.

Hahn:1991:PPM

[Hah91] B. D. Hahn. Photosynthesis and photorespiration: Modelling the essentials. Journal of Theoretical Biology, 151(1):123–139, July 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251930580147X.

Haig:1991:DAE

[Hai91] David Haig. Developmental asynchrony and environmental sex determination in alligators. Journal of Theoretical Biology, 150(3):373–383, June 7, 1991. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305804357.

Haig:1992:ICE

[Hai92] David Haig. Intragenomic conflict and the evolution of eusociality. Journal of Theoretical Biology, 156(3):401–403, June 7, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305806836.

Haig:1993:AMU

[Hai93] David Haig. Alternatives to meiosis: The unusual genetics of red algae, microsporidia, and others. Journal of Theoretical Biology, 163(1):15–31, July 1, 1993. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519383711045.

Haken:1994:BMV

[Hak94] Hermann Haken. A brain model for vision in terms of synergetics. Journal of Theoretical Biology, 171(1):75–85, July 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519384712136.

Hammond:1993:QSC

[Ham93] B. J. Hammond. Quantitative study of the control of HIV-1 gene expression. Journal of Theoretical Biology, 163(2):199– 221, July 21, 1993. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519383711173.

Hammer:1998:DDS

[Ham98] Øyvind Hammer. Diffusion and direct signaling models are numerically equivalent. Journal of Theoretical Biology, 192(1):129–130, May 7, 1998. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519398906543.

Hamrick:1999:CMT

[Ham99] Mark W. Hamrick. A chondral modeling theory revisited. Journal of Theoretical Biology, 201(3):201–208, December 7, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519399910251.

Hannon:1991:ECS

[Han91] Bruce Hannon. Empirical cyclic stabilization of an oyster reef ecosystem. Journal of Theoretical Biology, 149(4):507–519, April 21, 1991. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305800955.

Hansen:1992:SAP

[Han92] Thomas Fredrik Hansen. Selection in asexual populations: An extension of the fundamental theorem. Journal of Theoretical Biology, 155(4):537–544, April 21, 1992. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305806344.

Hannon:1993:OGH

[Han93] Bruce Hannon. The optimal growth of Helianthus Annuus. Journal of Theoretical Biology, 165(4):523–531, December 21, 1993. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251938371204X.

Hand:1995:URI

[Han95] Brian J. Hand. The use of regular isotopy in the topological description of site-specific DNA recombination processes. Journal of Theoretical Biology, 176(1):161–174, September 7, 1995. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519385701853.

Harada:1999:SVL

[Har99a] Yuko Harada. Short- vs. long-range disperser: the evolutionarily stable allocation in a lattice-structured habitat. Journal of Theoretical Biology, 201(3):171–187, December 7, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519399910226.

Hardling:1999:ARC

[Här99b] Roger Härdling. Arms races, conflict costs and evolutionary dynamics. Journal of Theoretical Biology, 196(2):163–167, January 21, 1999. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519398908281.

Hastings:1992:PGA

[Has92] Ian M. Hastings. The population genetics of alleles affecting enzyme activity. Journal of Theoretical Biology, 157(3):305–316, August 7, 1992. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305806137.

Haslett:1994:CSF

[Has94a] John R. Haslett. Community structure and the fractal dimensions of mountain habitats. Journal of Theoretical Biology, 167(4):407–411, April 21, 1994. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384710794.

Hasson:1994:CS

[Has94b] Oren Hasson. Cheating signals. Journal of Theoretical Biology, 167(3):223–238, April 7, 1994. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519384710654.

Hasson:1997:TGT

[Has97] Oren Hasson. Towards a general theory of biological signaling. Journal of Theoretical Biology, 185(2):139–156, March 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519396902581.

Havsteen:1991:SAP

[Hav91] Bent H. Havsteen. A stochastic attractor participates in chymotrypsin catalysis. a new facet of enzyme catalysis. Journal of Theoretical Biology, 151(4):557–571, August 21, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305803704.

Havsteen:1997:ACB

[Hav97] B. Havsteen. Attractor control of the binding of digoxin to a specific antibody. Journal of Theoretical Biology, 189 (4):367–376, December 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002251939790518X.

Havsteen:1999:EAV

[Hav99] Bent H. Havsteen. An esterolytic antibody with vibrational properties of a catalytic h-chain and a non-catalytic l-chain. Journal of Theoretical Biology, 196(3):377–387, February 7, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519398908487.

Herrick:1991:GRU

[HB91] John Herrick and David Bensimon. Gene regulation under growth conditions. a model for the regulation of initiation of replication in Escherichia coli. Journal of Theoretical Biology, 151(3):359–366, August 7, 1991. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305803856.

Hatzimanikatis:1996:MMS

[HB96] Vassily Hatzimanikatis and James E. Bailey. MCA has more to say. Journal of Theoretical Biology, 182(3):233–242, October 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396901605.

Heinrich:1997:MPT

[HBJ^{+}97] Reinhart Heinrich, Milan Brumen, Annekathrin Jaeger, Peter Müller, and Andreas Herrmann. Modelling of phospholipid translocation in the erythrocyte membrane: a combined kinetic and thermodynamic approach. Journal of Theoretical Biology, 185(3):295–312, April 7, 1997. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396903252.

Higashi:1992:TNS

[HBP92] Masahiko Higashi, Thomas P. Burns, and Bernard C. Patten. Trophic niches of species and trophic structure of ecosystems: Complementary perspectives through food network unfolding. Journal of Theoretical Biology, 154(1):57–76, January 7, 1992. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305801882.

Halter:1991:DPM

[HC91] J. A. Halter and J. W. Clark. A distributed-parameter model of the myelinated nerve fiber. Journal of Theoretical Biology, 148(3):345–382, February 7, 1991. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305802425.

Horvath:1993:CRP

[HC93] Gábor Horváth and Euan N. K. Clarkson. Computational reconstruction of the probable change of form of the corneal lens and maturation of optics in the post-ecdysial development of the schizochroal eye of the Devonian trilobite Phacops rana milleri Stewart 1927. Journal of Theoretical Biology, 160(3):343–373, February 7, 1993. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519383710222.

Haefner:1994:SMM

[HC94] James W. Haefner and Thomas O. Crist. Spatial model of movement and foraging in harvester ants (Pogonomyrmex) (I): The roles of memory and communication. Journal of Theoretical Biology, 166(3):299–313, February 7, 1994. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384710277.

Hofmeyr:1996:CRA

[HCB96] Jan-Hendrik S. Hofmeyr and Athel Cornish-Bowden. Coresponse analysis: a new experimental strategy for metabolic control analysis. Journal of Theoretical Biology, 182(3): 371–380, October 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519396901769.

Havsteen:1992:KTA

[HCM^{+}92] B. H. Havsteen, R. Varón Castellanos, M. Molina, M. J. García Meseguer, E. Valero, and M. García-Moreno. Kinetic theory of the action of lipases. Journal of Theoretical Biology, 157(4):523–533, August 21, 1992. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305806678.

Hammond:1991:MID

[HCS91] Bruce J. Hammond, Jillian R. Christensen, and Geoffrey D. Smith. A model of insulin distribution and uptake in the perfused rat liver. Journal of Theoretical Biology, 149(1):121–139, March 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251930580075X.

Huisman:1997:FDB

[HD97] Gert Huisman and Rob J. De Boer. A formal derivation of the “beddington” functional response. Journal of Theoretical Biology, 185(3):389–400, April 7, 1997. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396903185.

Harun:1999:BBM

[HDF^{+}99] M. A. S. Harun, G. Draisma, K. Frankena, R. J. Veeneklaas, and M. Van Kampen. Breeding biology of muscovy duck (Cairina moschata) under natural incubation: the use of the Weibull function and a beta-binomial model to predict nest hatchability. Journal of Theoretical Biology, 198(1):89–99, May 7, 1999. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519399909025.

Hodson:1993:IGP

[HE93] Stuart Hodson and Richard Earlam. The incorporation of gel pressure into the irreversible thermodynamic equation of fluid flow in order to explain biological tissue swelling. Journal of Theoretical Biology, 163(2):173–180, July 21, 1993. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251938371115X.

Hodson:1994:EMH

[HE94] Stuart Hodson and Richard Earlam. Of an extracellular matrix in human pre-corneal tear film. Journal of Theoretical Biology, 168(4):395–398, June 21, 1994. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384711192.

Hunding:1995:EBM

[HE95] A. Hunding and R. Engelhardt. Early biological morphogenesis and nonlinear dynamics. Journal of Theoretical Biology, 173(4):401–413, April 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519385700720.

Hurd:1998:CSA

[HE98] Peter L. Hurd and Magnus Enquist. Conventional signalling in aggressive interactions: the importance of temporal structure. Journal of Theoretical Biology, 192(2):197–211, May 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519397906111.

Hulbert:1999:MPP

[HE99] A. J. Hulbert and Paul Lewis Else. Membranes as possible pacemakers of metabolism. Journal of Theoretical Biology, 199(3):257–274, August 7, 1999. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519399909554.

Hedenstrom:1992:FPR

[Hed92] Anders Hedenström. Flight performance in relation to fuel load in birds. Journal of Theoretical Biology, 158(4):535–537, October 21, 1992. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305807143.

Hein:1994:ACD

[Hei94] Jotun Hein. An algorithm combining DNA and protein alignment. Journal of Theoretical Biology, 167(2):169–174, March 21, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519384710629.

Hemelrijk:1990:MPC

[Hem90] C. K. Hemelrijk. A matrix partial correlation test used in investigations of reciprocity and other social interaction patterns at group level. Journal of Theoretical Biology, 143(3):405–420, April 5, 1990. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305800360.

Hemelrijk:1991:IAA

[Hem91] C. K. Hemelrijk. Interchange of “altruistic” acts as an epiphenomenon. Journal of Theoretical Biology, 153(1):137–139, November 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251930580356X.

Hemsley:1998:NVS

[Hem98] Alan R. Hemsley. Nonlinear variation in simulated complex pattern development. Journal of Theoretical Biology, 192(1): 73–79, May 7, 1998. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251939790610X.

Herendeen:1990:SLI

[Her90] Robert Herendeen. System-level indicators in dynamic ecosystems: Comparison based on energy and nutrient flows. Journal of Theoretical Biology, 143(4):523–553, April 19, 1990.

CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305800281.

Herve:1991:SCM

[Her91] Guy Hervé. Semantics of channeling and microcompartmentation. Journal of Theoretical Biology, 152(1):131–132, September 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305805260.

Herz:1994:CPS

[Her94]  Andreas V. M. Herz.  Collective phenomena in spa-

tially extended evolutionary games. Journal of Theoretical Biology, 169(1):65–87, July 7, 1994. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519384711301.

Herendeen:1995:UQA

[Her95] R. A. Herendeen. A unified quantitative approach to trophic cascade and bottom-up: top-down hypotheses. Journal of Theoretical Biology, 176(1):13–26, September 7, 1995. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385701713.

Heschl:1990:CSE

[Hes90] Adolf Heschl. L = C a simple equation with astonishing consequences. Journal of Theoretical Biology, 145(1):13–40, July 9, 1990. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305805326.

Hesslow:1994:WNE

[Hes94] Germund Hesslow. Will neuroscience explain consciousness? Journal of Theoretical Biology, 171(1):29–39, July 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519384712094.

Hatzis:1997:CCA

[HFS97] Christos Hatzis, A. G. Frederickson, and Friedrich Srienc. Cellcycle analysis in phagotrophic microorganisms from flow cytometric histograms. Journal of Theoretical Biology, 186(2):131– 144, May 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396903483.

Haig:1991:GSD

[HG91] David Haig and Alan Grafen. Genetic scrambling as a defence against meiotic drive. Journal of Theoretical Biology, 153(4):531–558, December 21, 1991. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305801559.

Heinrich:1991:KPE

[HH91a] Reinhart Heinrich and Edda Hoffmann. Kinetic parameters of enzymatic reactions in states of maximal activity; an evolutionary approach. Journal of Theoretical Biology, 151(2):249–283, July 21, 1991. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305803637.

Hovorka:1991:IIR

[HH91b] R. Hovorka and J. Hilgertova. Identification of insulin receptor systems: Assessing the impact of model selection and measurement error on precision of parameter estimates using Monte Carlo study. Journal of Theoretical Biology, 151(3):367–383, August 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305803868.

Hofer:1993:SOA

[HH93] Thomas Höfer and Reinhart Heinrich. A second-order approach to metabolic control analysis. Journal of Theoretical Biology, 164(1):85–102, September 7, 1993. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519383711410.

Henson:1995:OFC

[HH95] Shandelle M. Henson and Thomas G. Hallam. Optimal feeding via constrained processes. Journal of Theoretical Biology, 176(1):33–37, September 7, 1995. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385701737.

Hill:1997:BRW

[HH97] N. A. Hill and D.-P. Häder. A biased random walk model for the trajectories of swimming micro-organisms. Journal of Theoretical Biology, 186(4):503–526, June 21, 1997. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519397904215.

Holt:1998:CCP

[HH98] Robert D. Holt and Michael E. Hochberg. The coexistence of competing parasites. part II-hyperparasitism and food chain dynamics. Journal of Theoretical Biology, 193(3):485–495, August 7, 1998. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519398907172.

Hearn:1998:CNP

[HHM98] Taryn Hearn, Caroline Haurie, and Michael C. Mackey. Cyclical neutropenia and the peripheral control of white blood cell production. Journal of Theoretical Biology, 192(2):167–181, May 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519397905890.

Hayashi:1992:CNB

[HI92] Hatsuo Hayashi and Satoru Ishizuka. Chaotic nature of bursting discharges in the onchidium pacemaker neuron. Journal of Theoretical Biology, 156(3):269–291, June 7, 1992. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305806769.

Hiebeler:1997:SSM

[Hie97] David Hiebeler. Stochastic spatial models: From simulations to mean field and local structure approximations. Journal of Theoretical Biology, 187(3):307–319, August 7, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519397904227.

Hines:1998:MLE

[Hin98] W. G. S. Hines. Multi-locus ESS models: Non-additive effects. Journal of Theoretical Biology, 191(2):181–189, March 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519397905786.

Hirata:1993:IOE

[Hir93] Hironori Hirata. Information of organization in ecological systems: Nutrient > energy > carbon. Journal of Theoretical Biology, 162(2):187–194, May 21, 1993. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519383710829.

Heinrich:1996:CAU

[HK96] Reinhart Heinrich and Edda Klipp. Control analysis of unbranched enzymatic chains in states of maximal activity. Journal of Theoretical Biology, 182(3):243–252, October 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519396901617.

Hunding:1998:ESA

[HK98] A. Hunding and M. Kærn. The effect of slow allosteric transitions in a simple biochemical oscillator model. Journal of Theoretical Biology, 191(3):309–322, April 7, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519397905877.

Haring:1999:VMS

[HK99]  David Häring and Jaroslav Kypr.  Variations of the

mononucleotide and short oligonucleotide distributions in the genomes of various organisms. Journal of Theoretical Biology, 201(2):141–156, November 21, 1999. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519399910196.

Hunding:1990:DSS

[HKG90] Axel Hunding, Stuart A. Kauffman, and Brian C. Goodwin. Drosophila segmentation: Supercomputer simulation of prepattern hierarchy. Journal of Theoretical Biology, 145(3):369–384, August 9, 1990. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251930580116X.

Huynen:1993:MCE

[HKH93] Martijn A. Huynen, Danielle A. M. Konings, and Paulien Hogeweg. Multiple coding and the evolutionary properties of RNA secondary structure. Journal of Theoretical Biology, 165(2):251–267, November 21, 1993. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519383711884.

Horkovics-Kovats:1991:MMP

[HKT91] Stefan Horkovics-Kovats and Peter Traub. A mathematical model for protein-induced lipid vesicle leakage: Interaction of the intermediate filament protein vimentin and its isolated n-terminus with phosphatidylinositol vesicles. Journal of Theoretical Biology, 153(1):89–110, November 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305803546.

Harding:1996:EMC

[HL96a] Stephan P. Harding and James E. Lovelock. Exploitermediated coexistence and frequency-dependent selection in a numerical model of biodiversity. Journal of Theoretical Biology, 182(2):109–116, September 21, 1996. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396901460.

Hopkins:1996:DAP

[HL96b] Joy C. Hopkins and Robert J. Leipold. On the dangers of adjusting the parameter values of mechanism-based mathematical models. Journal of Theoretical Biology, 183(4):417–427, December 21, 1996. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396902325.

Haugh:1998:ARI

[HL98a] Jason M. Haugh and Douglas A. Lauffenburger. Analysis of receptor internalization as a mechanism for modulating signal transduction. Journal of Theoretical Biology, 195(2):187–218, November 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519398907913.

Hertel:1998:RNN

[HL98b] Fritz Hertel and Niles Lehman. A randomized nearest-neighbor approach for assessment of character displacement: the vulture guild as a model. Journal of Theoretical Biology, 190(1):51–61, January 7, 1998. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519397905312.

Helmstetter:1992:RBC

[HLG92] Charles E. Helmstetter, Alan C. Leonard, and Julia E. Grimwade. Relationships between chromosome segregation, cell shape and temperature in Escherichia coli. Journal of Theoretical Biology, 159(2):261–266, November 21, 1992. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305807052.

Hyman:1994:TCS

[HLS94] James M. Hyman, Jia Li, and E. Ann Stanley. Threshold conditions for the spread of the HIV infection in agestructured populations of homosexual men. Journal of Theoretical Biology, 166(1):9–31, January 7, 1994. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384710022.

Hsu:1996:EMB

[HLT96] Jyh-Ping Hsu, Sung-Hwa Lin, and Shiojenn Tseng. Electrophoretic mobility of biological cells in asymmetric electrolyte solutions. Journal of Theoretical Biology, 182(2):

137–145, September 21, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519396901484.

Hess:1995:MSO

[HM95] Benno Hess and Alexander Mikhailov. Microscopic selforganization in living cells: a study of time matching. Journal of Theoretical Biology, 176(1):181–184, September 7, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385701877.

Hubble:1996:PDC

[HMEW96] John Hubble, Fang Ming, Robert Eisenthal, and William Whish. Progressive detachment of cells from surfaces: a consequence of heterogeneous ligand populations or a multisite binding equilibrium? Journal of Theoretical Biology, 182(2):169–171, September 21, 1996. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519396901514.

Husimi:1993:MPC

[HMUA93] Yuzuru Husimi, Takuya Mori, Shinobu Ushijima, and Takuyo Aita. Modes of population changeovers in continuous culture of bacterial and viral mutants spectrum. Journal of Theoretical Biology, 165(4):577–591, December 21, 1993. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519383712087.

Houston:1995:ESE

[HMW95] Alasdair I. Houston, John M. McNamara, and James N. Webb. The evolutionarily stable exploitation of a renewing resource. Journal of Theoretical Biology, 177(2):151–158, November 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519385702330.

Hurst:1991:NEM

[HN91] Laurence D. Hurst and Paul Nurse. A note on the evolution of meiosis. Journal of Theoretical Biology, 150(4):561–563, June 21, 1991. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305804473.

Ho:1990:ERT

[Ho90] Mae-Wan Ho. An exercise in rational taxonomy. Journal of Theoretical Biology, 147(1):43–57, November 7, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305802516.

Hoath:1990:CRS

[Hoa90] Steven B. Hoath. Considerations on the role of surface tension and epidermal growth factor in the mechanism of integumental pattern formation. Journal of Theoretical Biology, 143(1):1–14, March 8, 1990. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305802851.

Hochberg:1991:NLT

[Hoc91] Michael E. Hochberg. Non-linear transmission rates and the dynamics of infectious disease. Journal of Theoretical Biology, 153(3):301–321, December 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305805727.

Hofmeyr:1991:MCM

[Hof91]  Jan-Hendrik S. Hofmeyr.  Metabolite channelling and

metabolic regulation. Journal of Theoretical Biology, 152 (1):101, September 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519305805193.

Hokkanen:1990:TRM

[Hok90] J. E. I. Hokkanen. Temperature regulation of marine mammals. Journal of Theoretical Biology, 145(4):465–485, August 23, 1990. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305804825.

Holliday:1991:QGV

[Hol91] Robin Holliday. Quantitative genetic variation and developmental clocks. Journal of Theoretical Biology, 151(3):351–358, August 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305803844.

Holgate:1993:FMK

[Hol93] P. Holgate. Families of multivalent kinship relations. Journal of Theoretical Biology, 164(4):537–541, October 21, 1993. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519383711720.

Holman:1996:IVE

[Hol96] Eric W. Holman. The independent variable in the early origin of higher taxa. Journal of Theoretical Biology, 181(1):85– 94, July 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396901162.

Holliday:1999:TEC

[Hol99] Robin Holliday. Is there an epigenetic component in long-term memory? Journal of Theoretical Biology, 200(3):339–341, October 7, 1999. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519399909955.

Honda:1998:TMR

[Hon98] Hisao Honda. Topographic mapping in the retinotectal projection by means of complementary ligand and receptor gradients: a computer simulation study. Journal of Theoretical Biology, 192(2):235–246, May 21, 1998. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519398906622.

Hopfield:1994:PCW

[Hop94] J. J. Hopfield. Physics, computation, and why biology looks so different. Journal of Theoretical Biology, 171(1):53–60, July 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519384712112.

Horikawa:1993:NES

[Hor93] Yo Horikawa. Noise effects on spike propagation during the refractory period in the FitzHugh–Nagumo model. Journal of Theoretical Biology, 162(1):41–59, May 7, 1993. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519383710751.

Horvath:1995:RPP

[Hor95] Gábor Horváth. Reflection-polarization patterns at flat water surfaces and their relevance for insect polarization vision. Journal of Theoretical Biology, 175(1):27–37, July 7, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251938570118X.

Houston:1990:PBF

[Hou90] Alasdair I. Houston. Parent birds feeding young — to walk or to fly? Journal of Theoretical Biology, 142(2):141–147, January 23, 1990. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305802176.

Houston:1992:CIF

[Hou92] Alasdair I. Houston. The cost imposed on a fishery by a seabird population. Journal of Theoretical Biology, 158(3):313–327, October 7, 1992. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305807362.

Howe:1992:POE

[How92] Christopher J. Howe. Plastid origin of an extrachromosomal DNA molecule from plasmodium, the causative agent of malaria. Journal of Theoretical Biology, 158(2):199– 205, September 21, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519305807180.

Hurtt:1995:CRL

[HP95] George C. Hurtt and Stephen W. Pacala. The consequences of recruitment limitation: reconciling chance, history and competitive differences between plants. Journal of Theoretical Biology, 176(1):1–12, September 7, 1995. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385701701.

Hallett:1997:SEU

[HP97a] M. B. Hallett and E. J. Pettit. Stochastic events underlie Ca^{2+} signalling in neutrophils. Journal of Theoretical Biology, 186(1):1–6, May 7, 1997. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519396903458.

Horvath:1997:HCP

[HP97b] Gábor Horváth and István Pomozi. How celestial polarization changes due to reflection from the deflector panels used in deflector loft and mirror experiments studying avian orientation. Journal of Theoretical Biology, 184(3):291–300, February 7, 1997. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396902453.

Heinrich:1991:MCA

[HR91] Reinhart Heinrich and Christine Reder. Metabolic control analysis of relaxation processes. Journal of Theoretical Biology, 151(3):343–350, August 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305803832.

Hatton:1993:GTN

[HRW93] Mark W. C. Hatton, Mary Richardson, and Peter D. Winocour. On glucose transport and non-enzymic glycation of proteins in vivo. Journal of Theoretical Biology, 161(4):481–490, April 21, 1993. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519383710684.

Heinrich:1991:MCC

[HS91] R. Heinrich and S. Schuster. Is metabolic channelling the complicated solution to the easy problem of reducing transient times? Journal of Theoretical Biology, 152(1):57–61, September 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305805107.

Haraguchi:1995:MMP

[HS95] Yoshihiro Haraguchi and Hiromi Seno. Mathematical model of the population dynamics of brood parasitism: what is the benefit for the host? Journal of Theoretical Biology, 174(3):281– 297, June 7, 1995. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519385700999.

Haraguchi:1996:HPA

[HS96] Yoshihiro Haraguchi and Akira Sasaki. Host-parasite arms race in mutation modifications: Indefinite escalation despite a heavy load? Journal of Theoretical Biology, 183(2):121–137, November 21, 1996. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396999993.

Hatzoglou:1997:DNS

[HS97] Evi Hatzoglou and Constantine E. Sekeris. The detection of nucleotide sequences with strong similarity to hormone responsive elements in the genome of eubacteria and archaebacteria and their possible relation to similar sequences present in the mitochondrial genome. Journal of Theoretical Biology, 184(3):339–344, February 7, 1997. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396902854.

Han:1998:SDB

[HS98a] Bo-Ping Han and Milan Straskraba. Size dependence of biomass spectra and population density I. The effects of size scales and size intervals. Journal of Theoretical Biology, 191(3):259–265, April 7, 1998. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519397905749.

Happel:1998:EDR

[HS98b] Robert Happel and Peter F. Stadler. The evolution of diversity in replicator networks. Journal of Theoretical Biology, 195(3):329–338, December 7, 1998. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519398907974.

Hauert:1998:EIP

[HS98c] Ch. Hauert and H. G. Schuster. Extending the iterated Prisoner’s dilemma without synchrony. Journal of Theoretical Biology, 192(2):155–166, May 21, 1998. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519397905907.

Hahn:1994:MTM

[HSM94] Intaek Hahn, Peter W. Scherer, and Maxwell M. Mozell. A mass transport model of olfaction. Journal of Theoretical Biology, 167(2):115–128, March 21, 1994. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384710575.

Hatcher:1995:EPE

[HT95] Melanie J. Hatcher and Chris Tofts. The effect of point of expression on ESS sex ratios. Journal of Theoretical Biology, 175(2):263–266, July 21, 1995. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519385701397.

Hartl:1996:CNN

[HT96] Daniel L. Hartl and Clifford H. Taubes. Compensatory nearly neutral mutations: Selection without adaptation. Journal of Theoretical Biology, 182(3):303–309, October 7, 1996. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251939690168X.

Hines:1997:MES

[HT97] W. G. S. Hines and Michael Turelli. Multilocus evolutionarily stable strategy models: Additive effects. Journal of Theoretical Biology, 187(3):379–388, August 7, 1997. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519397904409.

Higo:1997:TAB

[HU97] Junichi Higo and Hideaki Umeyama. A theoretical approach to the Bell-shaped dependency of cell proliferation on the hormone concentration. Journal of Theoretical Biology, 186(4):477–490, June 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519397904124.

Hummel:1992:BMA

[Hum92] Z. Hummel. The binding of myosin to actin regulated by flexibility. entropy-controlled association. Journal of Theoretical Biology, 154(3):331–334, February 7, 1992. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305801742.

Hurst:1990:PDE

[Hur90] Laurence D. Hurst. Parasite diversity and the evolution of diploidy, multicellularity and anisogamy. Journal of Theoretical Biology, 144(4):429–443, June 21, 1990. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305800852.

Hurst:1991:ECI

[Hur91] Laurence D. Hurst. The evolution of cytoplasmic incompatibility or when spite can be successful. Journal of Theoretical Biology, 148(2):269–277, January 21, 1991. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305803443.

Hurd:1995:CDA

[Hur95] Peter L. Hurd. Communication in discrete action-response games. Journal of Theoretical Biology, 174(2):217–222, May 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519385700938.

Hurd:1997:SFA

[Hur97a] Peter L. Hurd. Is signalling of fighting ability costlier for weaker individuals? Journal of Theoretical Biology, 184(1):83–88, January 7, 1997. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396902465.

Hurst:1997:WCI

[Hur97b] Gregory D. D. Hurst. Wolbachia, cytoplasmic incompatibility, and the evolution of eusociality. Journal of Theoretical Biology, 184(1):99–100, January 7, 1997. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396902647.

Hutchinson:1999:BHW

[Hut99] John M. C. Hutchinson. Bet-hedging when targets May disappear: Optimal mate-seeking or prey-catching trajectories and the stability of leks and herds. Journal of Theoretical Biology, 196(1):33–49, January 7, 1999. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519398908177.

Havsteen:1990:KCC

[HV90] B. H. Havsteen and R. Varón. Kinetics of the classical complement activation cascade. Journal of Theoretical Biology, 145(1):47–64, July 9, 1990. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251930580534X.

Hill:1993:SMS

[HV93] N. A. Hill and R. V. Vincent. A simple model and strategies for orientation in phototactic microorganisms. Journal of Theoretical Biology, 163(2):223–235, July 21, 1993. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519383711185.

Han:1999:PLL

[HVKS99] Bo-Ping Han, Markku Virtanen, Jorma Koponen, and Milan Straskraba. Predictors of light-limited growth and competition of phytoplankton in a well-mixed water column. Journal of Theoretical Biology, 197(4):439–450, April 21, 1999.

CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519398908852.

Huth:1992:SMF

[HW92] Andreas Huth and Christian Wissel. The simulation of the movement of fish schools. Journal of Theoretical Biology, 156(3):365–385, June 7, 1992. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305806812.

Hills:1994:NMB

[HW94] B. P. Hills and K. M. Wright. A new model for bacterial growth in heterogeneous systems. Journal of Theoretical Biology, 168(1):31–41, May 7, 1994. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251938471085X.

Haydon:1998:IAS

[HW98a] Daniel T. Haydon and Mark E. J. Woolhouse. Immune avoidance strategies in RNA viruses: Fitness continuums arising from trade-offs between immunogenicity and antigenic variability. Journal of Theoretical Biology, 193(4):601–612, August 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519398907263.

Hellriegel:1998:CFR

[HW98b] Barbara Hellriegel and Paul I. Ward. Complex female reproductive tract morphology: Its possible use in postcopulatory female choice. Journal of Theoretical Biology, 190(2):179–186, January 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519397905464.

Hariri:1990:VSI

[HWO90] Ali Hariri, Bruce Weber, and John Olmsted III. On the validity of Shannon information calculations for molecular biological sequences. Journal of Theoretical Biology, 147(2):235–254, November 21, 1990. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305800542.

Hanski:1993:MMD

[HZ93] Ilkka Hanski and Da-Yong Zhang. Migration, metapopulation dynamics and fugitive co-existence. Journal of Theoretical Biology, 163(4):491–504, August 21, 1993. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519383711343.

Holden:1995:CAR

[HZ95] A. V. Holden and H. Zhang. Characteristics of atrial reentry and meander computed from a model of a rabbit single atrial cell. Journal of Theoretical Biology, 175(4):545–551, August 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519385701622.

Imafuku:1999:PFM

[IET99] Yasuhiro Imafuku, Yumiko Emoto, and Katsuhisa Tawada. A protein friction model of the actin sliding movement generated by myosin in mixtures of MgATP and MgGTP in vitro. Journal of Theoretical Biology, 199(4):359–370, August 21, 1999. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519399909633.

Iida:1990:QAS

[Iid90] Yôichi Iida. Quantification analysis of 5’-splice signal sequences in mRNA precursors. mutations in 5’-splice signal sequence of human β-globin gene and β-thalassemia. Journal of Theoretical Biology, 145(4):523–533, August 23, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305804862.

Iwasa:1995:TLH

[IL95] Yoh Iwasa and Simon A. Levin. The timing of life history events. Journal of Theoretical Biology, 172(1):33–42, January 7, 1995. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519385700033.

Incardona:1995:HMH

[IMPR95] F. Incardona, S. Morante, V. Parisi, and G. C. Rossi. How much are homologous peptides homologous? Journal of Theoretical Biology, 175(4):437–455, August 21, 1995. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385701531.

Inghe:1990:CSF

[Ing90] Ola Inghe. Computer simulations of flowering rhythms in perennials — is there a new area to explore in the quest for chaos? Journal of Theoretical Biology, 147(4):449–469, December 21, 1990. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305802590.

Ingvar:1994:WBC

[Ing94] David H. Ingvar. The will of the brain: Cerebral correlates of willful acts. Journal of Theoretical Biology, 171(1): 7–12, July 1994. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519384712069.

Iwasa:1999:SSX

[IP99a] Y. Iwasa and A. Pomiankowski. Sex specific X chromosome expression caused by genomic imprinting. Journal of Theoretical Biology, 197(4):487–495, April 21, 1999. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519398908888.

Iwasa:1999:GPG

[IP99b] Yoh Iwasa and Andrew Pomiankowski. Good parent and good genes models of handicap evolution. Journal of Theoretical Biology, 200(1):97–109, September 7, 1999. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519399909797.

Irvine:1990:OAR

[Irv90] Douglas H. Irvine. Objectives, assumptions and results of metabolic control theory and biochemical systems theory. Journal of Theoretical Biology, 143(1):139–143, March 8, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305802930.

Ishida:1990:UIB

[Ish90] K. Ishida. Usefulness of information in bistable system and Dyson model for the origin of metabolism. Journal of Theoretical Biology, 146(2):233–248, September 21, 1990. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305801377.

Iwasa:1991:DMW

[ISN91] Yoh Iwasa, Kazunori Sato, and Seigo Nakashima. Dynamic modeling of wave regeneration (shimagare) in subalpine abies forests. Journal of Theoretical Biology, 152(2): 143–158, September 21, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519305804485.

Israel:1996:TPR

[Isr96] Lucien Israel. Tumour progression: Random mutations or an integrated survival response to cellular stress conserved from unicellular organisms? Journal of Theoretical Biology, 178(4):375–380, February 21, 1996. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396900338.

Isagi:1997:HDM

[ISSI97] Y. Isagi, K. Sugimura, A. Sumida, and H. Ito. How does masting happen and synchronize? Journal of Theoretical Biology, 187(2):231–239, July 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519397904422.

Tainaka:1994:IUE

[iT94] Kei ichi Tainaka. Intrinsic uncertainty in ecological catastrophe. Journal of Theoretical Biology, 166(1):91–99, January 7, 1994. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519384710071.

Iwasa:1993:OME

[Iwa93] Yoh Iwasa. Overdispersed molecular evolution in constant environments. Journal of Theoretical Biology, 164(3):373–393, October 7, 1993. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519383711616.

Imai:1999:CGE

[IWH^{+}99] Hirotami T. Imai, Masayasu Y. Wada, Hirohisa Hirai, Yoichi Matsuda, and Kimiyuki Tsuchiya. Cytological, genetic and evolutionary functions of chiasmata based on chiasma graph analysis. Journal of Theoretical Biology, 198(2):239–257, May 21, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251939990913X.

Jablonka:1994:ISE

[Jab94] Eva Jablonka. Inheritance systems and the evolution of new levels of individuality. Journal of Theoretical Biology, 170(3):301–309, October 7, 1994. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251938471191X.

Jacobson:1991:LPM

[Jac91] Jerry I. Jacobson. A look at the possible mechanism and potential of magneto therapy. Journal of Theoretical Biology, 149(1):97–119, March 7, 1991. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305800748.

Jafri:1995:TSC

[Jaf95] Saleet M. Jafri. A theoretical study of cytosolic calcium waves in xenopus oocytes. Journal of Theoretical Biology, 172(3):209–216, February 7, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519385700173.

Jain:1997:MPI

[Jai97] Chaitanya Jain. Models for pairing of IS10 encoded antisense RNAs in vivo. Journal of Theoretical Biology, 186(4):431–439, June 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396903719.

James:1990:HHC

[Jam90a] William H. James. The hypothesized hormonal control of human sex ratio at birth — an update. Journal of Theoretical Biology, 143(4):555–564, April 19, 1990. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305800293.

Jamon:1990:RRH

[Jam90b] Marc Jamon. A reassessment of the random hypothesis in the ocean migration of Pacific salmon. Journal of Theoretical Biology, 143(2):197–214, March 22, 1990. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930580267X.

James:1991:CSM

[Jam91] William H. James. The combinations of the sexes of mammalian sibs and littermates: and the possibility of adaptive control. Journal of Theoretical Biology, 152(4):567–568, October 21, 1991. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305803984.

James:1992:HHC

[Jam92] William H. James. The hypothesized hormonal control of mammalian sex ratio at birth — a second update. Journal of Theoretical Biology, 155(1):121–128, March 7, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305805521.

James:1996:EMS

[Jam96] William H. James. Evidence that mammalian sex ratios at birth are partially controlled by parental hormone levels at the time of conception. Journal of Theoretical Biology, 180(4):271– 286, June 21, 1996. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396901022.

James:1997:PMS

[Jam97] William H. James. A potential mechanism for sex ratio variation in mammals. Journal of Theoretical Biology, 189(3):253–255, December 7, 1997. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251939790521X.

James:1998:HMS

[Jam98a] William H. James. Hypotheses on mammalian sex ratio variation at birth. Journal of Theoretical Biology, 192(1):113–116, May 7, 1998. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251939890649X.

James:1998:HOC

[Jam98b] William H. James. Hypothesis: One cause of polydactyly. Journal of Theoretical Biology, 192(1):1–2, May 7, 1998. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519397905919.

James:1999:FER

[Jam99a] William H. James. Further evidence relating offspring sex ratios to parental hormone levels around the time of conception. Journal of Theoretical Biology, 197(2):261–263, March 21, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519398908840.

James:1999:TGA

[Jam99b] William H. James. Is transposition of the great arteries a consequence of maternal hormone imbalance? evidence from the sex ratios of relatives of probands. Journal of Theoretical Biology, 198(3):301–303, June 7, 1999. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519399909165.

Jarv:1995:MNE

[Jär95] Jaak Järv. A model of non-exclusive binding of agonist and antagonist on g-protein coupled receptors. Journal of Theoretical Biology, 175(4):577–582, August 21, 1995. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251938570166X.

Johansson:1995:MEP

[JB95] Jan Johansson and Clas Blomberg. A model of error propagation in the presence of an error-enhancing drug. Journal of Theoretical Biology, 173(1):1–13, March 7, 1995. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385700379.

Jimenez:1997:DSM

[JBSdlG97] Rafael Jiménez, Miguel Burgos, Antonio Sánchez, and Rafael Díaz de la Guardia. Did Spanish moles really change their mechanism of sex determination in only 5 years? Journal of Theoretical Biology, 188(1):141–142, September 7, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519397904574.

Jeffries:1996:MTT

[JC96] Clark Jeffries and Joel E. Cohen. Marker transport through ecosystem energy flow. Journal of Theoretical Biology, 179(4):323–328, April 21, 1996. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519396900715.

Jolicoeur:1992:BAP

[JD92] Pierre Jolicoeur and Gilles Ducharme. Bivariate allometry: Point estimation of the slope of the normal major axis. Journal of Theoretical Biology, 154(1):35–41, January 7, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305801869.

Jensen:1999:LTN

[JD99] Susanne Plesner Jensen and C. Patrick Doncaster. Lethal toxins in non-preferred foods: How plant chemical defences can drive microtine cycles. Journal of Theoretical Biology, 199(1): 63–85, July 7, 1999. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519399909414.

Jonsson:1998:EPP

[JE98] Tomas Jonsson and Bo Ebenman. Effects of predator–prey body size ratios on the stability of food chains. Journal of Theoretical Biology, 193(3):407–417, August 7, 1998. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519398907081.

Jean:1992:MTP

[Jea92] Roger V. Jean. Model testing in phyllotaxis. Journal of Theoretical Biology, 156(1):41–62, May 7, 1992. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305806551.

Jenner:1998:HOS

[Jen98] Ronald A. Jenner. Hierarchical organization and the structural-interactive hierarchy. Journal of Theoretical Biology, 194(3):457–460, October 7, 1998. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519398907652.

Janovy:1990:MDB

[JFM90] J. Janovy, M. T. Ferdig, and Mary Ann McDowell. A model of dynamic behavior of a parasite species assemblage. Journal of Theoretical Biology, 142(4):517–529, February 22, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305801067.

Johnston:1990:NDT

[JG90] Timothy D. Johnston and Gilbert Gottlieb. Neophenogenesis: a developmental theory of phenotypic evolution. Journal of Theoretical Biology, 147(4):471–495, December 21, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305802607.

Johnstone:1992:CSP

[JG92] Rufus A. Johnstone and Alan Grafen. The continuous Sir Philip Sidney game: a simple model of biological signalling. Journal of Theoretical Biology, 156(2):215–234, May 21, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305806745.

Jung:1995:MME

[JH95] Seunho Jung and Rawle I. Hollingsworth. A mathematical model explaining the molecular weights and distribution of very long chain dicarboxylic acids formed during the adaptive response of sarcina ventriculi. Journal of Theoretical Biology, 172(2):121–126, January 21, 1995. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385700094.

Jaffe:1997:DEG

[JIDH97] Klaus Jaffe, Solange Issa, Eric Daniels, and Dan Haile. Dynamics of the emergence of genetic resistance to biocides among asexual and sexual organisms. Journal of Theoretical Biology, 188(3):289–299, October 7, 1997. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519397904720.

Jose:1998:PDT

[JJ98] Marco V. José and Jorge V. José. Probability distributions of thermodynamic affinities for heterogeneous receptor populations. Journal of Theoretical Biology, 190(1):85–92, January 7, 1998. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519397905361.

Jewett:1998:RLC

[JK98] Megan E. Jewett and Richard E. Kronauer. Refinement of limit cycle oscillator model of the effects of light on the human circadian pacemaker. Journal of Theoretical Biology, 192(4):455– 465, June 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519398906671.

Johnson:1992:SCE

[JLK92] Edward A. Johnson, D. Renald Lemieux, and J. Mailen Kootsey. Sodium-calcium exchange: Derivation of a state diagram and rate constants from experimental data. Journal of Theoretical Biology, 156(4):443–483, June 21, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305806381.

Jablonka:1992:EMM

[JLL92] Eva Jablonka, Michael Lachmann, and Marion J. Lamb. Evidence, mechanisms and models for the inheritance of acquired characters. Journal of Theoretical Biology, 158(2): 245–268, September 21, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519305807222.

Joshi:1995:MGO

[JM95] Amitabh Joshi and Michael E. Moody. Male gamete output of asexuals and the dynamics of populations polymorphic for reproductive mode. Journal of Theoretical Biology, 174(2):189– 197, May 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519385700914.

Joshi:1998:CSR

[JM98] Amitabh Joshi and Michael E. Moody. The cost of sex revisited: Effects of male gamete output of hermaphrodites that are asexual in their female capacity. Journal of Theoretical Biology, 195(4):533–542, December 21, 1998. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519398908116.

Johnson:1996:TLB

[JO96] D. R. Johnson and P. O’Higgins. Is there a link between changes in the vertebral “hox code” and the shape of vertebrae? A quantitative study of shape change in the cervical vertebral column of mice. Journal of Theoretical Biology, 183(1):89–93, November 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396902040.

Johansson:1993:SLM

[Joh93] Staffan Johansson. Switch-like membrane properties computed from classical-type Na^{+} and K^{+} currents: a neuronal memory device. Journal of Theoretical Biology, 164(4):515–529, October 21, 1993. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519383711707.

Johnstone:1995:HAM

[Joh95] Rufus A. Johnstone. Honest advertisement of multiple qualities using multiple signals. Journal of Theoretical Biology, 177(1):87–94, November 7, 1995. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305800062.

Johansen:1996:SMR

[Joh96] Anders Johansen. A simple model of recurrent epidemics. Journal of Theoretical Biology, 178(1):45–51, January 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396900053.

John:1997:SCT

[Joh97] Jeremy Leighton John. Seven comments on the theory of sosigonic selection. Journal of Theoretical Biology, 187(3):333–349, August 7, 1997. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519397904410.

Jolicoeur:1990:BAI

[Jol90] Pierre Jolicoeur. Bivariate allometry: Interval estimation of the slopes of the ordinary and standardized normal major axes and structural relationship. Journal of Theoretical Biology, 144(2):275–285, May 14, 1990. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305803261.

Jones:1994:LDF

[Jon94] M. E. Jones. Luria–Delbrück fluctuation experiments; accounting simultaneously for plating efficiency and differential growth rate. Journal of Theoretical Biology, 166(3):355–363, February 7, 1994. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519384710320.

Jongeling:1996:SOC

[Jon96] Tjeerd B. Jongeling. Self-organization and competition in evolution: a conceptual problem in the use of fitness landscapes. Journal of Theoretical Biology, 178(4):369–373, February 21, 1996. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396900326.

Jones:1997:ELP

[Jon97] L. H. Jones. The effects of leaf pruning and other stresses on sex determination in the oil Palm and their representation by a computer simulation. Journal of Theoretical Biology, 187(2):241–260, July 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519397904331.

Jorgensen:1995:GRZ

[Jør95] Sven Erik Jørgensen. The growth rate of zooplankton at the edge of chaos: ecological models. Journal of Theoretical Biology, 175(1):13–21, July 7, 1995. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519385701166.

Josic:1997:LBS

[Jos97] Kresimir Josić. Local bifurcations in the symmetric model of selection with fertility differences. Journal of Theoretical Biology, 189(3):291–295, December 7, 1997. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519397905142.

Joshi:1990:MDHa

[JP90a] Abhay Joshi and Bernhard O. Palsson. Metabolic dynamics in the human red cell. Part III — metabolic reaction rates. Journal of Theoretical Biology, 142(1):41–68, January 9, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305800128.

Joshi:1990:MDHb

[JP90b] Abhay Joshi and Bernhard O. Palsson. Metabolic dynamics in the human red cell. Part IV — data prediction and some model computations. Journal of Theoretical Biology, 142(1):69–85, January 9, 1990. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251930580013X.

Juckett:1992:HDM

[JR92] David A. Juckett and Barnett Rosenberg. Human disease mortality kinetics are explored through a chain model embodying principles of extreme value theory and competing risks. Journal of Theoretical Biology, 155(4):463–483, April 21, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305806307.

Jeune:1992:MAP

[JS92] Bernard Jeune and Rolf Sattler. Multivariate analysis in process morphology of plants. Journal of Theoretical Biology, 156(2):147–167, May 21, 1992. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305806708.

Janosi:1997:EDD

[JS97] Imre M. Jánosi and István Scheuring. On the evolution of density dependent dispersal in a spatially structured population model. Journal of Theoretical Biology, 187(3):397–408, August 7, 1997. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519397904343.

Jones:1993:RAO

[JSB93] Richard E. Jones, Hobart M. Smith, and Carl E. Bock. Reptilian and avian ovarian cycles and the evolutionary origin of volant birds. Journal of Theoretical Biology, 161(4):527–536, April 21, 1993. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519383710714.

James:1999:CBD

[JSR99] Adrienne James, Karl Swann, and Michael Recce. Cell behaviour as a dynamic attractor in the intracellular signalling system. Journal of Theoretical Biology, 196(3):269–288, February 7, 1999. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251939890827X.

Jones:1999:ALA

[JTC99] M. E. Jones, S. M. Thomas, and K. Clarke. The application of a linear algebra to the analysis of mutation rates. Journal of Theoretical Biology, 199(1):11–23, July 7, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519399909335.

Judson:1997:MAC

[Jud97] Olivia P. Judson. A model of asexuality and clonal diversity: Cloning the red Queen. Journal of Theoretical Biology, 186(1): 33–40, May 7, 1997. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396903392.

Jarchau:1991:DRF

[JV91] Th. Jarchau and H. Vogt. The distribution of restriction fragment lengths for non-overlapping restriction sites in a random DNA sequence model. Journal of Theoretical Biology, 153(4):445–453, December 21, 1991. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305801493.

Johnson:1995:CSS

[JW95] N. A. Johnson and M. J. Wade. Conditions for soft selection favoring the evolution of hybrid inviability. Journal of Theoretical Biology, 176(4):493–499, October 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385702159.

Johnsen:1999:PBT

[JW99] Sönke Johnsen and Edith A. Widder. The physical basis of transparency in biological tissue: Ultrastructure and the minimization of light scattering. Journal of Theoretical Biology, 199(2):181–198, July 21, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519399909487.

Kimmel:1991:UCD

[KA91] Marek Kimmel and David E. Axelrod. Unequal cell division, growth regulation and colony size of mammalian cells: a mathematical model and analysis of experimental data. Journal of Theoretical Biology, 153(2):157–180, November 21, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305804205.

Kaandorp:1994:FDR

[Kaa94] Jaap A. Kaandorp. A formal description of radiate accretive growth. Journal of Theoretical Biology, 166(2):149–161, January 21, 1994. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519384710137.

Kaandorp:1995:ASR

[Kaa95] Jaap A. Kaandorp. Analysis and synthesis of radiate accretive growth in three dimensions. Journal of Theoretical Biology, 175(1):39–55, July 7, 1995. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519385701191.

Kabakov:1994:RPE

[Kab94] Anatolii Yu. Kabakov. The resting potential equations incorporating ionic pumps and osmotic concentration. Journal of Theoretical Biology, 169(1):51–64, July 7, 1994. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384711295.

Kacser:1991:ST

[Kac91]  H. Kacser. A superior theory?  Journal of Theoret-

ical Biology, 149(1):141–144, March 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305800761.

Kalapos:1997:PER

[Kal97] Miklós Péter Kalapos. Possible evolutionary role of methylglyoxalase pathway: Anaplerotic route for reductive citric acid cycle of surface metabolists. Journal of Theoretical Biology, 188(2):201–206, September 21, 1997. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519397904690.

Kalapos:1998:MSE

[Kal98] Milkós Péter Kalapos. From mineral support to enzymatic catalysis — further assumptions for the evolutionary history of glyoxalase system. Journal of Theoretical Biology, 193(1): 91–98, July 7, 1998. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519398906907.

Kanavarioti:1992:SRC

[Kan92] Anastassia Kanavarioti. Self-replication of chemical systems based on recognition within a double or a triple helix: a realistic hypothesis. Journal of Theoretical Biology, 158(2):207–219, September 21, 1992. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305807192.

Kato:1990:CDD

[Kat90] Mikio Kato. Codon discrimination due to presence of abundant non-cognate competitive tRNA. Journal of Theoretical Biology, 142(1):35–39, January 9, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305800116.

Katz:1994:AMOa

[Kat94a] Yehuda Katz. Anesthesia and the Meyer-Overton rule I. Potencies and perturbations. Journal of Theoretical Biology, 167(2):93–97, March 21, 1994. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251938471054X.

Katz:1994:AMOb

[Kat94b] Yehuda Katz. Anesthesia and the Meyer-Overton rule II. a solution theory view of anesthesia and perturbations. Journal of Theoretical Biology, 167(2):99–105, March 21, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384710551.

Kato:1995:NAN

[Kat95] Masayuki Kato. Numerical analysis of the Nernst–Planck– Poisson system. Journal of Theoretical Biology, 177(3):299– 304, December 7, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519385702470.

Kato:1996:CPH

[Kat96] Mikio Kato. Conformational preference in the helical structure of repetitive DNA elements identified in a freshwater fish opsariichthys uncirostris. Journal of Theoretical Biology, 180(2):121–124, May 21, 1996. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519396900879.

Kauffman:1992:AME

[Kau92] Stuart A. Kauffman. Applied molecular evolution. Journal of Theoretical Biology, 157(1):1–7, July 7, 1992. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305807532.

Kaufman:1998:DMP

[Kau98] Michael D. Kaufman. Do microbes with peptides mimicking myelin cause multiple sclerosis if the T cell response to their unique peptides is limited? Journal of Theoretical Biology, 193(4):691–708, August 21, 1998. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519398907342.

Kayastha:1992:MCT

[Kay92] Arvind M. Kayastha. Metabolite channelling in tryptophan synthase. Journal of Theoretical Biology, 158(1):133–134, September 7, 1992. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305806514.

Khanin:1991:OAB

[KB91] M. A. Khanin and I. B. Bukharov. Optimal arterial blood pressure. Journal of Theoretical Biology, 148(3):289–294, February 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305802383.

Kshatriya:1992:TMO

[KB92] M. Kshatriya and R. W. Blake. Theoretical model of the optimum flock size of birds flying in formation. Journal of Theoretical Biology, 157(2):135–174, July 21, 1992. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305806186.

Kemp:1993:RIO

[KB93a] G. J. Kemp and A. Bevington. The regulation of intracellular orthophosphate concentration. Journal of Theoretical Biology, 161(1):77–94, March 7, 1993. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519383710416.

Khanin:1993:OPA

[KB93b] M. A. Khanin and I. B. Bukharov. Optimal pulmonary arterial blood pressure. Journal of Theoretical Biology, 162(4):431–445, June 21, 1993. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519383710982.

Khanin:1994:OSM

[KB94] M. A. Khanin and I. B. Bukharov. Optimal structure of the microcirculatory bed. Journal of Theoretical Biology, 169(3):267–273, August 7, 1994. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519384711477.

Kirschner:1995:DHP

[KB95] Denise E. Kirschner and Martin J. Blaser. The dynamics of helicobacter pylori infection of the human stomach. Journal of Theoretical Biology, 176(2):281–290, September 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519385701981.

Kacelnik:1998:RCW

[KB98a] Alex Kacelnik and Fausto Brito e Abreu. Risky choice and Weber’s law. Journal of Theoretical Biology, 194(2): 289–298, September 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519398907639.

Karadi:1998:IVM

[KB98b] Kázmér Karádi and István Bende. The indirect verification of minimal two-generator model of homeostatic drinking by complexity examination. Journal of Theoretical Biology, 192(1): 55–60, May 7, 1998. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519397906093.

Kanjilal:1999:RMC

[KB99] P. P. Kanjilal and J. Bhattacharya. Revisited measles and chickenpox dynamics through orthogonal transformation. Journal of Theoretical Biology, 197(2):163–174, March 21, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519398908657.

Kampis:1991:LSR

[KC91] G. Kampis and V. Csányi. Life, self-reproduction and information: Beyond the machine metaphor. Journal of Theoretical Biology, 148(1):17–32, January 7, 1991. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305804631.

Kier:1997:CAM

[KC97] Lemont B. Kier and Chao-Kun Cheng. Cellular automata model of membrane permeability. Journal of Theoretical Biology, 186(1):75–80, May 7, 1997. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519396903331.

Kulesa:1996:MMS

[KCL^{+}96] P. M. Kulesa, G. C. Cruywagen, S. R. Lubkin, P. K. Main, J. Sneyd, M. W. J. Ferguson, and J. D. Murray. On a model mechanism for the spatial patterning of teeth primordia in the alligator. Journal of Theoretical Biology, 180(4):287–296, June 21, 1996. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396901034.

Kozlov:1990:MFP

[KCM90] M. M. Kozlov, L. V. Chernomordik, and V. S. Markin. A mechanism of formation of protein-free regions in the red cell membrane: the rupture of the membrane skeleton. Journal of Theoretical Biology, 144(3):347–365, June 7, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305800803.

Kam:1993:ELG

[KD93] Michael Kam and A. Allan Degen. Energetics of lactation and growth in the fat sand rat, psammomys obesus: New perspectives of resource partitioning and the effect of litter size. Journal of Theoretical Biology, 162(3):353–369, June 7, 1993. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519383710921.

Keizer:1994:SRM

[KD94] Joel Keizer and Gary De Young. Simplification of a realistic model of IP_{3}-induced Ca^{2+} oscillations. Journal of Theoretical Biology, 166(4):431–442, February 21, 1994. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384710381.

Kam:1997:ERE

[KD97] Michael Kam and A. Allan Degen. Energy requirements and the efficiency of utilization of metabolizable energy in freeliving animals: Evaluation of existing theories and generation of a new model. Journal of Theoretical Biology, 184(2):101–104, January 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396902799.

Killingback:1998:SOC

[KD98] Timothy Killingback and Michael Doebeli. Self-organized criticality in spatial evolutionary game theory. Journal of Theoretical Biology, 191(3):335–340, April 7, 1998. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519397906020.

Koch:1999:ACC

[KD99a] Arthur L. Koch and Ronald J. Doyle. Attachment of the chromosome to the cell poles: the strategy for the growth of bacteria in two and three dimensions. Journal of Theoretical Biology, 199(2):213–221, July 21, 1999. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519399909505.

Koella:1999:PDE

[KD99b] Jacob C. Koella and Michael Doebeli. Population dynamics and the evolution of virulence in epidemiological models with discrete host generations. Journal of Theoretical Biology, 198(3):461–475, June 7, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519399909256.

Koziol:1993:UBA

[KDBK93] James A. Koziol, Debra J. Decker, Nora E. Boyle, and Norman R. Klinman. The use of Bayesian analysis of PCR-derived genomic DNA sequences to enable a biologically relevant interpretation of immunoglobulin variable region gene expression. Journal of Theoretical Biology, 160(3):387–398, February 7, 1993. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519383710246.

Kindlmann:1992:RBI

[KDG92] P. Kindlmann, A. F. G. Dixon, and L. J. Gross. The relationship between individual and population growth rates in multicellular organisms. Journal of Theoretical Biology, 157(4):535–542, August 21, 1992. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251930580668X.

Kyriazakis:1992:GMF

[KE92] I. Kyriazakis and G. C. Emmans. The growth of mammals following a period of nutritional limitation. Journal of Theoretical Biology, 156(4):485–498, June 21, 1992. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305806393.

Kokko:1996:MSD

[KE96] Hanna Kokko and Torbjörn Ebenhard. Measuring the strength of demographic stochasticity. Journal of Theoretical Biology, 183(2):169–178, November 21, 1996. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396902106.

Keener:1991:EDG

[Kee91] James P. Keener. The effects of discrete gap junction coupling on propagation in myocardium. Journal of Theoretical Biology, 148(1):49–82, January 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305804655.

Keener:1996:DAD

[Kee96] James P. Keener. Direct activation and defibrillation of cardiac tissue. Journal of Theoretical Biology, 178(3):313–324, February 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396900272.

Keightley:1996:MMS

[Kei96] Peter D. Keightley. Metabolic models of selection response. Journal of Theoretical Biology, 182(3):311–316, October 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519396901691.

Kell:1991:PSM

[Kel91] Douglas B. Kell. On the physiological significance of metabolite channelling: If, how, and where, but not why. Journal of Theoretical Biology, 152(1):49–51, September 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305805089.

Kelly:1992:KSD

[Kel92] John K. Kelly. Kin selection in density regulated populations. Journal of Theoretical Biology, 157(4):447–461, August 21, 1992. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305806630.

Keller:1994:SES

[Kel94] Andrew D. Keller. Specifying epigenetic states with autoregulatory transcription factors. Journal of Theoretical Biology, 170(2):175–181, September 21, 1994. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384711775.

Keller:1995:FES

[Kel95a] Andrew D. Keller. Fixation of epigenetic states in a population. Journal of Theoretical Biology, 176(2):211–219, September 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519385701920.

Keller:1995:MGC

[Kel95b] Andrew D. Keller. Model genetic circuits encoding autoregulatory transcription factors. Journal of Theoretical Biology, 172(2):169–185, January 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519385700148.

Keller:1995:SDA

[Kel95c] Laurent Keller. Sterility by deleterious alleles and the evolution of sociality. Journal of Theoretical Biology, 174(4):467–469, June 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519385701130.

Kelly:1996:RRE

[Kel96] John K. Kelly. Replication rate and evolution in the human immunodeficiency virus. Journal of Theoretical Biology, 180(4):359–364, June 21, 1996. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519396901083.

Kemp:1994:IMA

[Kem94] G. J. Kemp. Interactions of mitochondrial ATP synthesis and the creatine kinase equilibrium in skeletal muscle. Journal of Theoretical Biology, 170(3):239–246, October 7, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519384711842.

Kerszberg:1996:ARM

[Ker96] Michel Kerszberg. Accurate reading of morphogen concentrations by nuclear receptors: a formal model of complex transduction pathways. Journal of Theoretical Biology, 183(1):95–104, November 7, 1996. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519396902052.

Keynes:1997:SPO

[Key97] Richard D. Keynes. Steps to the path to the origin of species. Journal of Theoretical Biology, 187(4):461–471, August 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396903811.

Kauffman:1990:SHP

[KG90] Stuart A. Kauffman and Brian C. Goodwin. Spatial harmonics and pattern specification in early Drosophila development. part II. the four colour wheels model. Journal of Theoretical Biology, 144(3):321–345, June 7, 1990. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305800797.

Kell:1993:MDP

[KG93] Antje Kell and Ralf W. Glaser. On the mechanical and dynamic properties of plant cell membranes: Their role in growth, direct gene transfer and protoplast fusion. Journal of Theoretical Biology, 160(1):41–62, January 7, 1993. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519383710039.

Kvadsheim:1997:EVH

[KGFB97] Petter H. Kvadsheim, Anne R. L. Gotaas, Lars P. Folkow, and Arnoldus S. Blix. An experimental validation of heat loss models for marine mammals. Journal of Theoretical Biology, 184(1):15–23, January 7, 1997. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519396902568.

Kiani:1991:CSG

[KH91] Mohammad F. Kiani and Antal G. Hudetz. Computer simulation of growth of anastomosing microvascular networks. Journal of Theoretical Biology, 150(4):547–560, June 21, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305804461.

Klipp:1994:EOE

[KH94] Edda Klipp and Reinhart Heinrich. Evolutionary optimization of enzyme kinetic parameters; effect of constraints. Journal of Theoretical Biology, 171(3):309–323, December 7, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519384712343.

Kaern:1998:DCC

[KH98] Mads Kærn and Axel Hunding. Dynamics of the cell cycle engine: Cdk2-kinase and the transition into mitosis. Journal of Theoretical Biology, 193(1):47–57, July 7, 1998. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519398906774.

Kaern:1999:ESA

[KH99] Mads Kærn and Axel Hunding. The effect of slow allosteric transitions in a coupled biochemical oscillator model. Journal of Theoretical Biology, 198(2):269–281, May 21, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519399909128.

Khrennikov:1998:HSA

[Khr98] Andrei Khrennikov. Human subconscious as a p-adic dynamical system. Journal of Theoretical Biology, 193(2):179–196, July 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519397906044.

Kawachi:1993:OFP

[KIA93] Keiji Kawachi, Yoshinobu Inada, and Akira Azuma. Optimal flight path of flying fish. Journal of Theoretical Biology, 163(2):145–159, July 21, 1993. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519383711136.

Kubo:1996:FSD

[KIF96] Takuya Kubo, Yoh Iwasa, and Naoki Furumoto. Forest spatial dynamics with gap expansion: Total gap area and gap size distribution. Journal of Theoretical Biology, 180(3):229– 246, June 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396900995.

Kim:1993:ESA

[Kim93] Yong-Gwan Kim. Evolutionary stability in the asymmetric war of attrition. Journal of Theoretical Biology, 161(1):13–21, March 7, 1993. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519383710362.

Kim:1995:SSG

[Kim95] Yong-Gwan Kim. Status signaling games in animal contests. Journal of Theoretical Biology, 176(2):221–231, September 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519385701932.

Kirkpatrick:1992:DSF

[Kir92] Mark Kirkpatrick. Direct selection of female mating preferences: Comments on Grafen’s models. Journal of Theoretical Biology, 154(1):127–129, January 7, 1992. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305801936.

Kisdi:1999:EBU

[Kis99] Éva Kisdi. Evolutionary branching under asymmetric competition. Journal of Theoretical Biology, 197(2):149–162, March 21, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519398908645.

Kauffman:1991:CEC

[KJ91] Stuart A. Kauffman and Sonke Johnsen. Coevolution to the edge of chaos: Coupled fitness landscapes, poised states, and coevolutionary avalanches. Journal of Theoretical Biology, 149(4):467–505, April 21, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305800943.

Keitt:1995:SHA

[KJ95] Timothy H. Keitt and Alan R. Johnson. Spatial heterogeneity and anomalous kinetics: emergent patterns in diffusion-limited predatory-prey interaction. Journal of Theoretical Biology, 172(2):127–139, January 21, 1995. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385700100.

Kowald:1993:ATC

[KK93] A. Kowald and T. B. L. Kirkwood. Accuracy of tRNA charging and codon: Anticodon recognition; relative importance for cellular stability. Journal of Theoretical Biology, 160(4):493–508, February 21, 1993. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519383710325.

Koeslag:1994:K

[KK94a] Johan H. Koeslag and Peter D. Koeslag. Koinophilia. Journal of Theoretical Biology, 167(1):55–65, March 7, 1994. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384710496.

Koeslag:1994:KSB

[KK94b] Peter D. Koeslag and Johan H. Koeslag. Koinophilia stabilizes Bi-gender sexual reproduction against asex in an unchanging environment. Journal of Theoretical Biology, 166(3):251–260, February 7, 1994. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251938471023X.

Kooi:1994:ESM

[KK94c] B. W. Kooi and S. A. L. M. Kooijman. Existence and stability of microbial prey-predator systems. Journal of Theoretical Biology, 170(1):75–85, September 7, 1994. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384711696.

Kooi:1994:TBF

[KK94d] B. W. Kooi and S. A. L. M. Kooijman. The transient behaviour of food chains in chemostats. Journal of Theoretical Biology, 170(1):87–94, September 7, 1994. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384711702.

Kowald:1994:TNT

[KK94e] A. Kowald and T. B. L. Kirkwood. Towards a network theory of ageing: a model combining the free radical theory and the protein error theory. Journal of Theoretical Biology, 168(1):75– 94, May 7, 1994. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519384710897.

Komendantov:1996:DCM

[KK96] Alexander O. Komendantov and Nikolai I. Kononenko. Deterministic chaos in mathematical model of pacemaker activity in bursting neurons of snail, helix pomatia. Journal of Theoretical Biology, 183(2):219–230, November 21, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519396902155.

Kura:1998:WAI

[KK98] Takuya Kura and Kenya Kura. War of attrition with individual differences on RHP. Journal of Theoretical Biology, 193(2):335–344, July 21, 1998. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519398907056.

Kisljuk:1996:MCG

[KKA96] O. S. Kisljuk, T. V. Kusnetzova, and A. A. Agafonova. On the modelling of clone geometry in asarum europaeum. Journal of Theoretical Biology, 178(4):399–404, February 21, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519396900375.

Kurpick:1997:ILM

[KKH97] P. Kürpick, U. Kürpick, and A. Huth. The influence of logging on a Malaysian dipterocarp rain forest: a study using a forest gap model. Journal of Theoretical Biology, 185(1):47–54, March 7, 1997. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251939690288X.

Kooijman:1999:AME

[KKH99] S. A. L. M. Kooijman, B. W. Kooi, and T. G. Hallam. The application of mass and energy conservation laws in physiologically structured population models of heterotrophic organisms. Journal of Theoretical Biology, 197(3):371–392, April 7, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519398908815.

Kirzhner:1998:CLB

[KKN98] V. M. Kirzhner, A. B. Korol, and E. Nevo. Complex limiting behaviour of multilocus genetic systems in cyclical environments. Journal of Theoretical Biology, 190(3):215–225, February 7, 1998. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519397905476.

Kirzhner:1999:AMP

[KKN99] V. M. Kirzhner, A. B. Korol, and E. Nevo. Abundant multilocus polymorphisms caused by genetic interaction between species on trait-for-trait basis. Journal of Theoretical Biology, 198(1):61–70, May 7, 1999. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519399909001.

Kaden:1990:KBP

[KKS90] F. Kaden, I. Koch, and J. Selbig. Knowledge-based prediction of protein structures. Journal of Theoretical Biology, 147(1):85–100, November 7, 1990. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251930580253X.

Kukhtin:1997:GLP

[KKS97] V. V. Kukhtin, N. V. Kuzmenko, and O. V. Shramko. Green light as a possible pressing factor for oceanic phytoplankton near the base of the euphotic zone. Journal of Theoretical Biology, 188(3):319–322, October 7, 1997. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519397904823.

Keasling:1995:MCS

[KKV95] J. D. Keasling, Henry Kuo, and Gizette Vahanian. A Monte Carlo simulation of the Escherichia coli cell cycle. Journal of Theoretical Biology, 176(3):411–430, October 7, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519385702093.

Kohn:1991:IRP

[KL91] Michael C. Kohn and D. Renald Lemieux. Identification of regulatory properties of metabolic networks by graph theoretical modeling. Journal of Theoretical Biology, 150(1):3–25, May 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305804722.

Klaassen:1996:DFL

[KL96] Marcel Klaassen and Åke Lindström. Departure fuel loads in time-minimizing migating birds can be explained by the energy costs of being heavy. Journal of Theoretical Biology, 183(1):29–34, November 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519396901988.

Koch:1997:ASG

[KL97] A. J. Koch and J. Lehmann. About a symmetry of the genetic code. Journal of Theoretical Biology, 189(2):171–174, November 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519397905038.

Koch:1998:GHF

[KL98] K. S. Koch and H. L. Leffert. Giant hairpins formed by CUG repeats in myotonic dystrophy messenger RNAs might sterically block RNA export through nuclear pores. Journal of Theoretical Biology, 192(4):505–514, June 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519398906798.

Keener:1999:BMW

[KL99] J. P. Keener and T. J. Lewis. The biphasic mystery: Why a biphasic shock is more effective than a monophasic shock for defibrillation. Journal of Theoretical Biology, 200(1):1–17, September 7, 1999. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519399909724.

Klein:1998:HDS

[Kle98] Christian Th. Klein. Hysteresis-driven structure formation in biochemical networks. Journal of Theoretical Biology, 194(2):263–274, September 21, 1998. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519398907573.

Kauffman:1995:SSA

[KM95] Stuart A. Kauffman and William G. Macready. Search strategies for applied molecular evolution. Journal of Theoretical Biology, 173(4):427–440, April 21, 1995. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385700744.

Keitt:1996:IHA

[KM96a]  Timothy H. Keitt and Pablo A. Marquet.  The in-

troduced Hawaiian avifauna reconsidered: Evidence for self-organized criticality? Journal of Theoretical Biology, 182(2):161–167, September 21, 1996. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519396901502.

Kusch:1996:MSP

[KM96b] Ingo Kusch and Mario Markus. Mollusc Shell pigmentation: Cellular automaton simulations and evidence for undecidability. Journal of Theoretical Biology, 178(3):333–340, February 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396900296.

Klein:1997:MPF

[KM97] Christian Th. Klein and Bernd Mayer. A model for pattern formation in gap-junction coupled cells. Journal of Theoretical Biology, 186(1):107–115, May 7, 1997. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396903379.

Keymer:1998:PFP

[KMJ98] Juan E. Keymer, Pablo A. Marquet, and Alan R. Johnson. Pattern formation in a patch occupancy metapopulation model: a cellular automata approach. Journal of Theoretical Biology, 194(1):79–90, September 7, 1998. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519398907457.

Kawasaki:1997:MST

[KMM^{+}97] K. Kawasaki, A. Mochizuki, M. Matsushita, T. Umeda, and N. Shigesada. Modeling spatio-temporal patterns generated by bacillus subtilis. Journal of Theoretical Biology, 188(2):177–185, September 21, 1997. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519397904628.

Kasai:1993:COP

[KN93] Michiki Kasai and Kazuki Nunogaki. Channel opening process is responsible for the slow component in the flux measurement of membrane vesicles. Journal of Theoretical Biology, 161(4):461–480, April 21, 1993. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519383710672.

Knowles:1991:CWC

[Kno91] Jeremy R. Knowles. Calmer waters in the channel? Journal of Theoretical Biology, 152(1):53–55, September 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305805090.

Ko:1991:SMG

[Ko91] Minoru S. H. Ko. A stochastic model for gene induction. Journal of Theoretical Biology, 153(2):181–194, November 21, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305804217.

Kyrpides:1993:MSM

[KO93] Nikos C. Kyrpides and Christos A. Ouzounis. Mechanisms of specificity in mRNA degradation: Autoregulation and cognate interactions. Journal of Theoretical Biology, 163(3):373–392, August 7, 1993. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519383711264.

Koch:1990:SCE

[Koc90] Arthur L. Koch. The sacculus contraction/expansion model for gliding motility. Journal of Theoretical Biology, 142(1):95–112, January 9, 1990. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305800153.

Koch:1992:DFP

[Koc92] Arthur L. Koch. Differences in the formation of poles of enterococcus and bacillus. Journal of Theoretical Biology, 154(2):205–217, January 21, 1992. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305804035.

Koch:1993:GRM

[Koc93] Arthur L. Koch. Genetic response of microbes to extreme challenges. Journal of Theoretical Biology, 160(1):1–21, January 7, 1993. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519383710015.

Koch:1994:DDL

[Koc94a] Arthur L. Koch. Development and diversification of the last universal ancestor. Journal of Theoretical Biology, 168(3):269– 280, June 7, 1994. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519384711088.

Koch:1994:PHG

[Koc94b] Arthur L. Koch. The problem of hyphal growth in streptomycetes and fungi. Journal of Theoretical Biology, 171(2):137–150, November 21, 1994. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519384712197.

Koch:1995:GOR

[Koc95] Arthur L. Koch. The geometry and osmotic relations of plasmolysis spaces in bacteria and the role of endocytosis, tubular structures and scheie structures in their formation. Journal of Theoretical Biology, 176(4):471–492, October 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519385903005.

Koeslag:1990:KGS

[Koe90] Johan H. Koeslag. Koinophilia groups sexual creatures into species, promotes stasis, and stabilizes social behaviour. Journal of Theoretical Biology, 144(1):15–35, May 10, 1990. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305802978.

Koeslag:1994:KRR

[Koe94] J. H. Koeslag. Koinophilia replaces random mating in populations subject to mutations with randomly varying fitnesses. Journal of Theoretical Biology, 171(3):341–345, December 7, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519384712367.

Koeslag:1995:ES

[Koe95] Johan H. Koeslag. On the engine of speciation. Journal of Theoretical Biology, 177(4):401–409, December 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519385702561.

Koeslag:1997:SPD

[Koe97] Johan H. Koeslag. Sex, the Prisoner’s dilemma game, and the evolutionary inevitability of cooperation. Journal of Theoretical Biology, 189(1):53–61, November 7, 1997. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519397904963.

Kohn:1991:DG

[Koh91] Michael C. Kohn. David garfinkel. Journal of Theoretical Biology, 150(1):1–2, May 7, 1991. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305804710.

Kohn:1992:PIM

[Koh92] Michael C. Kohn. Propagation of information in MetaNet graph models. Journal of Theoretical Biology, 154(4):505–517, February 21, 1992. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305801663.

Kokko:1996:FPE

[Kok96] Hanna Kokko. Finding the probability of extinction: Error analysis and optimal use of Markovian population models. Journal of Theoretical Biology, 183(1):77–87, November 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396902039.

Kokko:1997:LGC

[Kok97] Hanna Kokko. The lekking game: Can female choice explain aggregated male displays? Journal of Theoretical Biology, 187(1):57–64, July 7, 1997. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519397904148.

Koltover:1997:RCT

[Kol97] V. K. Koltover. Reliability concept as a trend in biophysics of aging. Journal of Theoretical Biology, 184(2):157–163, January 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396902477.

Kondrashov:1995:CGV

[Kon95] Alexey S. Kondrashov. Contamination of the genome by very slightly deleterious mutations: why have we not died 100 times over? Journal of Theoretical Biology, 175(4):583–594, August 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519385701671.

Koojiman:1995:SAE

[Koo95] S. A. L. M. Koojiman. The stoichiometry of animal energetics. Journal of Theoretical Biology, 177(2):139–149, November 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519385702329.

Korona:1991:RAB

[Kor91] Ryszard Korona. On the role of age and body size in risky animal contests. Journal of Theoretical Biology, 152(2):165–176, September 21, 1991. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305804503.

Kornet:1993:PSS

[Kor93] D. J. Kornet. Permanent splits as speciation events: a formal reconstruction of the internodal species concept. Journal of Theoretical Biology, 164(4):407–435, October 21, 1993. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519383711641.

Korn:1997:PFL

[Kor97] Robert W. Korn. Pattern formation in the leaf of zebra grass. Journal of Theoretical Biology, 187(3):449–451, August 7, 1997. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519397904355.

Kaplan:1997:POA

[KOTL97] Peter D. Kaplan, Qi Ouyang, David S. Thaler, and Albert Libchaber. Parallel overlap assembly for the construction of computational DNA libraries. Journal of Theoretical Biology, 188(3):333–341, October 7, 1997. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519397904756.

Koziol:1991:PTR

[Koz91] James A. Koziol. On the prevalence of transcriptional regions in human genomic DNA. Journal of Theoretical Biology, 149(3):377–380, April 7, 1991. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305803121.

Koziol:1996:CSP

[Koz96] James A. Koziol. Comment on the statistical power of EDF tests of normality and the sample size required to distinguish geometric-normal (lognormal) from arithmeticnormal distributions of low variability. Journal of Theoretical Biology, 180(1):81–85, May 7, 1996. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519396900818.

Karsai:1993:CBS

[KP93a] István Karsai and Zsolt Pénzes. Comb building in social wasps: Self-organization and stigmergic script. Journal of Theoretical Biology, 161(4):505–525, April 21, 1993. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519383710702.

Kepler:1993:SHB

[KP93b] Thomas B. Kepler and Alan S. Perelson. Somatic hypermutation in B cells: An optimal control treatment. Journal of Theoretical Biology, 164(1):37–64, September 7, 1993. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519383711392.

Komarov:1992:NPS

[KPK92] V. M. Komarov, R. V. Polozov, and G. G. Konoplev. Nonplanar structure of nitrous bases and non-coplanarity of Watson-Crick pairs. Journal of Theoretical Biology, 155(3):

281–294, April 7, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519305806009.

Khanin:1997:FSN

[KPS97] Raya Khanin, Hanna Parnas, and Lee Segel. “First step” negative feedback accounts for inhibition of fast neurotransmitter release. Journal of Theoretical Biology, 188(3):261–276, October 7, 1997. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251939790480X.

Krauss:1996:RCC

[KQ96] Stefan Krauss and Patti A. Quant. Regulation and control in complex, dynamic metabolic systems: Experimental application of the top-down approaches of metabolic control analysis to fatty acid oxidation and ketogenesis. Journal of Theoretical Biology, 182(3):381–388, October 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396901770.

Krackow:1995:DAH

[Kra95a] Sven Krackow. The developmental asynchrony hypothesis for sex ratio manipulation. Journal of Theoretical Biology, 176(2):273–280, September 21, 1995. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251938570197X.

Krakauer:1995:SCM

[Kra95b] David C. Krakauer. Simple connectionist models of spatial memory in bees. Journal of Theoretical Biology, 172 (2):149–160, January 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519385700124.

Kramer:1999:OTD

[Kra99] Eric M. Kramer. Observation of topological defects in the xylem of populus deltoides and implications for the vascular cambium. Journal of Theoretical Biology, 200(2):223– 230, September 21, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002251939990989X.

Kremen:1992:BME

[Kre92] Alexandr Kremen. Biological molecular energy machines as measuring devices. Journal of Theoretical Biology, 154 (4):405–413, February 21, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519305801596.

Kremen:1994:RME

[Kre94] Alexandr Kremen. Ribosomes as molecular energy machines. Journal of Theoretical Biology, 170(3):231–238, October 7, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519384711830.

Krakauer:1995:SLR

[KRG95] David C. Krakauer and Miguel A. Rodríguez-Gironés. Searching and learning in a random environment. Journal of Theoretical Biology, 177(4):417–429, December 21, 1995. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385702585.

Krukonis:1991:PCM

[KS91] G. Krukonis and W. M. Schaffer. Population cycles in mammals and birds: Does periodicity scale with body size? Journal of Theoretical Biology, 148(4):469–493, February 21, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305802322.

Kacser:1996:HMP

[KS96a] Henrik Kacser and J. Rankin Small. How many phenotypes from one genotype? the case of prion diseases. Journal of Theoretical Biology, 182(3):209–218, October 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396901575.

Kalinin:1996:AMT

[KS96b] Vladimir I. Kalinin and Valentin A. Stonik. Application of morphological trends of evolution to phylogenetic interpretation of chemotaxonomic data. Journal of Theoretical Biology, 180(1): 1–10, May 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396900739.

Katz:1998:CSF

[KS98] Stephen L. Katz and Robert E. Shadwick. Curvature of swimming fish midlines as an index of muscle strain suggests swimming muscle produces net positive work. Journal of Theoretical Biology, 193(2):243–256, July 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519398906968.

Korr:1999:FFR

[KS99] Hubert Korr and Christoph Schmitz. Facts and fictions regarding post-natal neurogenesis in the developing human cerebral cortex. Journal of Theoretical Biology, 200(3):291–297, October 7, 1999. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251939990992X.

Khanin:1998:DMA

[KSF98] Raya Khanin, Lee Segel, and Anthony H. Futerman. The diffusion of molecules in axonal plasma membranes: the sites of insertion of new membrane molecules and their distribution along the axon surface. Journal of Theoretical Biology, 193(3):371–382, August 7, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251939890707X.

Kaitala:1990:NSP

[KSG90] Veijo Kaitala, Brian H. Smith, and Wayne M. Getz. Nesting strategies of primitively eusocial bees: a model of nest usurpation during the solitary state of the nesting cycle. Journal of Theoretical Biology, 144(4):445–471, June 21, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305800864.

Konecny:1995:CCC

[KSH95] Jaromir Konecny, Michael Schöniger, and Ludwig G. Hofacker. Complementary coding conforms to the primeval comma-less code. Journal of Theoretical Biology, 173(3):263–270, April 7, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519385700616.

Keyhani:1997:NMN

[KSM97] Keyvan Keyhani, Peter W. Scherer, and Maxwell M. Mozell. A numerical model of nasal odorant transport for the analysis of human olfaction. Journal of Theoretical Biology, 186(3):279– 301, June 7, 1997. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396903471.

Kumamoto:1991:STF

[Kum91] Eiichi Kumamoto. Short-term facilitation as a tool to know neuromodulator-induced change in Ca^{2+} movement in the nerve terminal. Journal of Theoretical Biology, 149(3):317– 323, April 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251930580308X.

Kunisawa:1992:SCP

[Kun92] Takashi Kunisawa. Synonymous codon preferences in bacteriophage T4: a distinctive use of transfer RNAs from T4 and from its host Escherichia coli. Journal of Theoretical Biology, 159(3):287–298, December 7, 1992. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305807258.

Kunisawa:1995:ASA

[Kun95] Takashi Kunisawa. Alteration in synonymous arginine codon preferences of bacillus subtilis during sporulation. Journal of Theoretical Biology, 172(4):387–390, February 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385700367.

Kura:1999:DEA

[Kur99] Takuya Kura. Dilemma of the equality: an all-pay contest with individual differences in resource holding potential. Journal of Theoretical Biology, 198(3):395–404, June 7, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519399909189.

Kuzminov:1995:MIS

[Kuz95a] Andrei Kuzminov. A mechanism for induction of the SOS response in E. coli: insights into the regulation of reversible protein polymerization in vivo. Journal of Theoretical Biology, 177(1):29–43, November 7, 1995. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385702226.

Kuznetsov:1995:EPT

[Kuz95b] Alexander N. Kuznetsov. Energetical profit of the third segment in parasagittal legs. Journal of Theoretical Biology, 172(1):95–105, January 7, 1995. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519385700070.

Kuznetsov:1996:HMC

[Kuz96] Vladimir A. Kuznetsov. “Harpoon” model for cell–cell adhesion and recognition of target cells by the natural killer cells. Journal of Theoretical Biology, 180(4):321–342, June 21, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251939690106X.

Kuhn-Velten:1993:FEK

[KV93] W. K. Kühn-Velten. Fusion of enzyme kinetics with quantitative cell biology constitutes a novel, synergistic approach to calculate local enzyme activities in situ from subcellularsite concentrations of enzymes and their substrates: An Endocrinologist’s contribution to the “metabolite channelling” debate. Journal of Theoretical Biology, 165(3):447–453, December 7, 1993. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519383712002.

Krink:1997:ASW

[KV97] Thiemo Krink and Fritz Vollrath. Analysing spider Webbuilding behaviour with rule-based simulations and genetic algorithms. Journal of Theoretical Biology, 185(3):321–331, April 7, 1997. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396903069.

Kahn:1991:CTR

[KW91] Daniel Kahn and Hans V. Westerhoff. Control theory of regulatory cascades. Journal of Theoretical Biology, 153(2):255–285, November 21, 1991. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305804266.

Koch:1994:MIP

[KW94] Arthur L. Koch and Conrad L. Woldringh. The metabolic inertness of the pole wall of a Gram-negative rod. Journal of Theoretical Biology, 171(4):415–425, December 21, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519384712458.

Kerszberg:1998:OME

[KW98a] M. Kerszberg and L. Wolpert. The origin of metazoa and the egg: a role for cell death. Journal of Theoretical Biology, 193(3):535–537, August 7, 1998. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519398907147.

Kerszberg:1998:MPS

[KW98b] Michel Kerszberg and Lewis Wolpert. Mechanisms for positional signalling by morphogen transport: a theoretical study. Journal of Theoretical Biology, 191(1):103–114, March 7, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519397905750.

Kruszewski:1998:GRC

[KW98c] Paul Kruszewski and Sue Whitesides. A general random combinatorial model of botanical trees. Journal of Theoretical Biology, 191(2):221–236, March 21, 1998. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519397905865.

Kamiya:1993:OAV

[KWB93] Akira Kamiya, Hiroshi Wakayama, and Keiichi Baba. Optimality analysis of vascular-tissue system in mammals for oxygen transport. Journal of Theoretical Biology, 162(2):229–242, May 21, 1993. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519383710842.

Kirby:1994:MDV

[KWD^{+}94] T. P. Kirby, P. K. Wraith, S. C. De Cort, M. A. A. Airlie, J. E. Hill, E. R. Carson, D. C. Flenley, and P. M. Warren. Modelling the dynamic ventilatory response to hypoxia in normal subjects. Journal of Theoretical Biology, 166(2):135–147, January 21, 1994. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519384710125.

Kaneko:1999:IDR

[KY99] Kunihiko Kaneko and Tetsuya Yomo. Isologous diversification for robust development of cell society. Journal of Theoretical Biology, 199(3):243–256, August 7, 1999. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519399909529.

Kaitala:1999:DCH

[KYH99] Veijo Kaitala, Janica Ylikarjula, and Mikko Heino. Dynamic complexities in host-parasitoid interaction. Journal of Theoretical Biology, 197(3):331–341, April 7, 1999. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519398908785.

Kypr:1990:RNP

[Kyp90] Jaroslav Kypr. Resistance of native protein shapes to random single point mutations in their genes. Journal of Theoretical Biology, 145(1):133–134, July 9, 1990. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305805399.

Lichtstein:1990:CSW

[LA90a] D. Lichtstein and H. Atlan. The “cellular state”: The way to regain specificity and diversity in hormone action. Journal of Theoretical Biology, 145(3):287–294, August 9, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305801109.

Lundberg:1990:FRO

[LÅ90b] Per Lundberg and Mårten Åström. Functional response of optimally foraging herbivores. Journal of Theoretical Biology, 144(3):367–377, June 7, 1990. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305800815.

Lindahl:1994:MFF

[LÅ94] B. I. B. Lindahl and P. Århem. Mind as a force field: Comments on a new interactionistic hypothesis. Journal of Theoretical Biology, 171(1):111–122, July 1994. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384712173.

Lindahl:1996:MFF

[LÅ96a] B. I. B. Lindahl and Peter Århem. The mental force field hypothesis: a reply to libet. Journal of Theoretical Biology, 178(2):225–226, January 21, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251939690020X.

Lindahl:1996:RBC

[LA96b] B. I. B. Lindahl and Peter Arhem. The relation between the conscious mind and the brain: a reply to Beck. Journal of Theoretical Biology, 181(1):95–96, July 7, 1996. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396901174.

Lacalli:1990:MDP

[Lac90] Thurston C. Lacalli. Modeling the Drosophila pair-rule pattern by reaction-diffusion: Gap input and pattern control in a 4morphogen system. Journal of Theoretical Biology, 144(2):171– 194, May 14, 1990. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305803170.

Lahav:1991:PCE

[Lah91] Noam Lahav. Prebiotic co-evolution of self-replication and translation or RNA world? Journal of Theoretical Biology, 151(4):531–539, August 21, 1991. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305803686.

Lushnikov:1997:MLD

[LAV^{+}97] A. A. Lushnikov, T. Ahonen, T. Vesala, E. Juurola, E. Nikinmaa, and P. Hari. Modelling of light-driven RuBP regeneration, carboxylation and CO 2 diffusion for leaf photosynthesis. Journal of Theoretical Biology, 188(2):143–151, September 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519397904653.

Lazcano:1997:LEC

[Laz97] Antonio Lazcano. Letter to the Editor: Chemical evolution and the primitive soup: Did oparin get it all right? Journal of Theoretical Biology, 184(2):219–223, January 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519396902738.

Luo:1995:MIP

[LB95] Liaofu Luo and Guoyi Bai. The maximum information principle and the evolution of nucleotide sequences. Journal of Theoretical Biology, 174(2):131–136, May 21, 1995. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385700860.

Lamotte-Brasseur:1990:EPM

[LBDDG90] Josette Lamotte-Brasseur, Georges Dive, Dominique Dehareng, and Jean-Marie Ghuysen. Electrostatic potential maps at the quantum chemistry level of the active sites of the serine peptidases, α-chymotrypsin and subtilisin. Journal of Theoretical Biology, 145(2):183–190, 191–198, July 24, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305801249.

Lee:1995:BUC

[LBN95] K. B. Lee, D. K. Boadi, and R. J. Neufeld. Blood urea clearance with microencapsulated urease. Journal of Theoretical Biology, 175(3):295–303, August 7, 1995. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385701439.

Lazar:1998:MMD

[LBND98] Dusan Lazár, Miroslav Brokes, Jan Naus, and Lubomír Dvorák. Mathematical modelling of 3-(3’,4’-dichlorophenyl)- 1,1-dimenthylurea action in plant leaves. Journal of Theoretical Biology, 191(1):79–86, March 7, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251939790566X.

Loeffler:1997:CLC

[LBP^{+}97] M. Loeffler, T. Bratke, U. Paulus, Y. Q. Li, and C. S. Potten. Clonality and life cycles of intestinal crypts explained by a state dependent stochastic model of epithelial stem cell organization. Journal of Theoretical Biology, 186(1):41–54, May 7, 1997. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396903409.

Liu:1995:MMB

[LBW^{+}95] C.-Q. Liu, C.-L. Bai, Y. Wang, J.-F. Huang, S.-S. Wang, X.-Q. Zhu, E.-H. Cao, and J.-S. Zhao. A molecular model of braid-like DNA structure. Journal of Theoretical Biology, 177(4):411–416, December 21, 1995. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385702573.

Loeffler:1993:SMM

[LBWP93] Markus Loeffler, Andreas Birke, Douglas Winton, and Christopher Potten. Somatic mutation, monoclonality and stochastic models of stem cell organization in the intestinal crypt. Journal of Theoretical Biology, 160(4):471–491, February 21, 1993. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519383710313.

Laurent:1997:SIC

[LC97] Michel Laurent and Michel Claret. Signal-induced Ca^{2+} oscillations through the regulation of the inositol 1,4,5trisphosphate-gated Ca^{2+} channel: an allosteric model. Journal of Theoretical Biology, 186(3):307–326, June 7, 1997. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396903653.

Levine:1995:CPB

[LCD95] M. Levine, J. P. Cleave, and C. Dodds. Can periodic breathing have advantages for oxygenation? Journal of Theoretical Biology, 172(4):355–368, February 21, 1995. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385700331.

Leon:1998:NIT

[LCP^{+}98] Kalet León, Jorge Carneiro, Rolando Peréz, Enrique Montero, and Agustin Lage. Natural and induced tolerance in an immune network model. Journal of Theoretical Biology, 193(3):519–534, August 7, 1998. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519398907202.

Lemieux:1994:MNI

[LCR94] D. R. Lemieux, H. Chicoine, and F. A. Roberge. Modeling a non-inactivating delayed rectifier cardiac current using voltage clamp data. Journal of Theoretical Biology, 169(4):363–373, August 21, 1994. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519384711593.

Larimer:1994:CFB

[LD94] James L. Larimer and Robert Dudley. Centrifugal force and blood pressure elevation in the wings of flying hummingbirds (trochilidae). Journal of Theoretical Biology, 168(3):233–236, June 7, 1994. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519384711040.

Li:1995:CMG

[LdQW95] Guo-Hua Li, Cheng de Qin, and Li-Wen Wang. Computer model of growth cone behavior and neuronal morphogenesis. Journal of Theoretical Biology, 174(4):381–389, June 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519385701063.

Lebovitz:1996:QED

[Leb96] Robert M. Lebovitz. Quantitative examination of dynamic interneuronal coupling via single-spike extracellular potassium ion transients. Journal of Theoretical Biology, 180(1):11–25, May 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396900740.

Leibowitz:1994:SHI

[Lei94] A. H. Leibowitz. Selection for hybrid inviability through kin selection. Journal of Theoretical Biology, 170(2):163– 174, September 21, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519384711763.

Leimar:1997:RGS

[Lei97] Olof Leimar. Repeated games: a state space approach. Journal of Theoretical Biology, 184(4):471–498, February 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519396902866.

Leon:1990:CNT

[Leó90a] Antonio León. Coevolution: New thermodynamic theorems. Journal of Theoretical Biology, 147(2):205–212, November 21, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305800529.

Leonard:1990:HD

[Leo90b] Janet L. Leonard. The Hermaphrodite’s dilemma. Journal of Theoretical Biology, 147(3):361–371, December 7, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930580493X.

Leonard:1992:LDH

[Leo92] Janet L. Leonard. The “love-dart” in helicid snails: a gift of calcium or a firm commitment? Journal of Theoretical Biology, 159(4):513–521, December 21, 1992. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305806952.

Lewis:1998:PRO

[Lew98] Kim Lewis. Pathogen resistance as the origin of kin altruism. Journal of Theoretical Biology, 193(2):359–363, July 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519398907251.

Laurent:1995:MER

[LF95] Michel Laurent and Anne Fleury. A model with excitability and relay properties for the generation and the propagation of a Ca^{2+} morphogenetic wave in paramecium. Journal of Theoretical Biology, 174(2):227–236, May 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385700951.

Lewis:1990:CDI

[LG90a] Timothy J. Lewis and Michael R. Guevara. Chaotic dynamics in an ionic model of the propagated cardiac action potential. Journal of Theoretical Biology, 146(3):407–432, October 7, 1990. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305807507.

Li:1990:FEP

[LG90b] Yue-Xian Li and Albert Goldbeter. Frequency encoding of pulsatile signals of cAMP based on receptor desensitization in Dictyostelium cells. Journal of Theoretical Biology, 146(3):355–367, October 7, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305807465.

Loeffler:1991:SBM

[LG91] Markus Loeffler and Bernd Grossmann. A stochastic branching model with formation of subunits applied to the growth of intestinal crypts. Journal of Theoretical Biology, 150(2):175– 191, May 21, 1991. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305803303.

Leloup:1999:CBM

[LG99] Jean-Christophe Leloup and Albert Goldbeter. Chaos and birhythmicity in a model for circadian oscillations of the PER and TIM proteins in drosophila. Journal of Theoretical Biology, 198(3):445–459, June 7, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519399909244.

Lewis:1992:PRF

[LGBP92] John E. Lewis, Leon Glass, Manjit Bachoo, and Canio Polosa. Phase resetting and fixed-delay stimulation of a simple model of respiratory rhythm generation. Journal of Theoretical Biology, 159(4):491–506, December 21, 1992. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305806939.

Latore:1998:SDC

[LGM98] J. Latore, P. Gould, and A. M. Mortimer. Spatial dynamics and critical patch size of annual plant populations. Journal of Theoretical Biology, 190(3):277–285, February 7, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519397905580.

LePape:1997:SAS

[LGSM97] Gilles Le Pape, Hector Giacomini, Bernard Swynghedauw, and Pascale Mansier. A statistical analysis of sequences of cardiac interbeat intervals does not support the chaos hypothesis. Journal of Theoretical Biology, 184(2):123–131, January 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251939690257X.

LeGuyader:1990:MDC

[LH90] Hervé Le Guyader and Claude Hyver. Modelling of the duplication of cortical units along a kinety of paramecium using reaction-diffusion equations. Journal of Theoretical Biology, 143(2):233–250, March 22, 1990. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305802693.

LeGuyader:1991:DCU

[LH91] Hervé Le Guyader and Claude Hyver. Duplication of cortical units on the cortex of paramecium: a model involving a Ca^{2+} wave. Journal of Theoretical Biology, 150(2):261–276, May 21, 1991. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305803376.

LeGuyader:1994:SMC

[LH94] Hervé Le Guyader and Claude Hyver. A single model can account for the main asymmetries of the cortex of paramecium: Application to homopolar doublets. Journal of Theoretical Biology, 170(3):317–326, October 7, 1994. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384711933.

Liechti:1994:ESO

[LHA94] Felix Liechti, Anders Hedenström, and Thomas Alerstam. Effects of sidewinds on optimal flight speed of birds. Journal of Theoretical Biology, 170(2):219–225, September 21, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519384711817.

Li:1994:PRM

[Li94] T. Li. A possible role of molecular decay in the origin of asymmetric emergence of biomolecules on Earth. Journal of Theoretical Biology, 170(2):227–230, September 21, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519384711829.

Liberman:1991:RBI

[Lib91] Uri Liberman. On the relation between the instability of ESS in discrete dynamics and segregation distortion. Journal of Theoretical Biology, 150(4):421–436, June 21, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305804370.

Libet:1996:CMF

[Lib96] Benjamin Libet. Conscious mind as a field. Journal of Theoretical Biology, 178(2):223–224, January 21, 1996. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396900193.

Libet:1997:CMF

[Lib97] Benjamin Libet. Conscious mind as a force field: a reply to lindahl and Århem. Journal of Theoretical Biology, 185(1):137– 138, March 7, 1997. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396903136.

Lima:1990:ESA

[Lim90] Steven L. Lima. Evolutionarily stable antipredator behavior among isolated foragers: On the consequences of successful escape. Journal of Theoretical Biology, 143(1):77–89, March 8, 1990. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305802899.

Lindemann:1994:GCH

[Lin94] Charles B. Lindemann. A “geometric clutch” hypothesis to explain oscillations of the axoneme of cilia and flagella. Journal of Theoretical Biology, 168(2):175–189, May 21, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519384710976.

Lindahl:1997:CBE

[Lin97] B. I. B. Lindahl. Consciousness and biological evolution. Journal of Theoretical Biology, 187(4):613–629, August 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251939690394X.

Liquori:1991:MBP

[Liq91] Alfonso M. Liquori. Myelin basic protein (MBP) displays significant homologies with gag core proteins of HTLV retroviruses. Journal of Theoretical Biology, 148(2):279–281, January 21, 1991. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305803455.

Liu:1993:NPM

[Liu93] Wei-Min Liu. A new pharmacodynamic model of antiarrhythmic drugs. Journal of Theoretical Biology, 165(4):533–539, December 21, 1993. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519383712051.

Lavigne:1990:LCT

[LIWE90] D. M. Lavigne, S. Innes, G. A. J. Worthy, and Elizabeth F. Edwards. Lower critical temperatures of blue whales, balaenoptera musculus. Journal of Theoretical Biology, 144(2):249–257, May 14, 1990. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305803236.

Liang:1992:CGD

[LJ92] Chengzhi Liang and Yuansheng Jiang. The chirality of ground DNA knots and links. Journal of Theoretical Biology, 158(2):231–243, September 21, 1992. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305807210.

Lee:1995:MIM

[LJ95] Michael G. Lee and Jacob Jorné. Multiplicity in ionic membranes with negative stoichiometric diffusivity. Journal of Theoretical Biology, 177(2):103–111, November 21, 1995. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385702299.

Lachmann:1996:IPA

[LJ96] Michael Lachmann and Eva Jablonka. The inheritance of phenotypes: an adaptation to fluctuating environments. Journal of Theoretical Biology, 181(1):1–9, July 7, 1996. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396901095.

Luo:1997:PMA

[LJ97] Liaofu Luo and Fengmin Ji. The preferential mode analysis of DNA sequence. Journal of Theoretical Biology, 188(3):343–353, October 7, 1997. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519397904859.

Lemieux:1991:SVD

[LK91] D. Renald Lemieux and J. Mailen Kootsey. Simulation of the voltage dependence of the Na, K pump applied to cardiac cells. Journal of Theoretical Biology, 150(1):73–91, May 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251930580476X.

Lundstrom:1993:LNL

[LKS^{+}93] K. I. Lundström, J. O. G. Karlsson, S. P. S. Svensson, L. G. E. Mårtensson, H. Elwing, S. Ödman, and R. G. G. Andersson. Local and non-local receptor signalling. Journal of Theoretical Biology, 164(2):135–148, September 21, 1993. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519383711446.

Lansing:1998:SDS

[LKS98] J. Stephen Lansing, James N. Kremer, and Barbara B. Smuts. System-dependent selection, ecological feedback and the emergence of functional structure in ecosystems. Journal of Theoretical Biology, 192(3):377–391, June 7, 1998. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519398906646.

Lemke:1990:MHP

[LL90] M. Lemke and I. Lamprecht. A model for heat production and thermoregulation in winter clusters of honey bees using differential heat conduction equations. Journal of Theoretical Biology, 142(2):261–273, January 23, 1990. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305802279.

Ludwikow:1993:TCM

[LL93] F. Ludwików and G. Ludwików. A two-compartment model of micronucleus formation in erythrocytes and its application to mouse bone marrow. I. Rectangularly transient action of clastogen in stationary conditions. Journal of Theoretical Biology, 165(3):417–428, December 7, 1993. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519383711987.

Lansner:1994:CMB

[LL94] A. Lansner and H. Liljenström. Computer models of the brain-how far can they take us? Journal of Theoretical Biology, 171(1):61–73, July 1994. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519384712124.

Li:1996:RBC

[LL96] Hong Li and Liaofu Luo. The relation between codon usage, base correlation and gene expression level in Escherichia coli and yeast. Journal of Theoretical Biology, 181(2):111–124, July 21, 1996. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396901198.

Landau:1997:CBC

[LL97] Mayer Landau and Paco Lorente. Conduction block and chaotic dynamics in an asymmetrical model of coupled cardiac cells. Journal of Theoretical Biology, 186(1):93–105, May 7, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519396903434.

Lifson:1999:MPR

[LL99] Shneior Lifson and Hanna Lifson. A model of prebiotic replication: Survival of the fittest versus extinction of the unfittest. Journal of Theoretical Biology, 199(4):425–433, August 21, 1999. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519399909694.

Lloyd:1995:CLM

[Llo95] Alun L. Lloyd. The coupled logistic map: a simple model for the effects of spatial heterogeneity on population dynamics. Journal of Theoretical Biology, 173(3):217–230, April 7, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519385700586.

Lanska:1994:STD

[LLS94] Vera Lánská, Petr Lánský, and Charles E. Smith. Synaptic transmission in a diffusion model for neural activity. Journal of Theoretical Biology, 166(4):393–406, February 21, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519384710356.

Lenski:1994:EVP

[LM94] Richard E. Lenski and Robert M. May. The evolution of virulence in parasites and pathogens: Reconciliation between two competing hypotheses. Journal of Theoretical Biology, 169(3):253–265, August 7, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519384711465.

Lloyd:1996:SHE

[LM96] Alun L. Lloyd and Robert M. May. Spatial heterogeneity in epidemic models. Journal of Theoretical Biology, 179(1):1–11, March 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396900429.

Lipsitch:1995:EVS

[LN95] Marc Lipsitch and Martin A. Nowak. The evolution of virulence in sexually transmitted HIV/AIDS. Journal of Theoretical Biology, 174(4):427–440, June 21, 1995. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385701099.

Lundberg:1996:SDF

[LNF96] Sigfrid Lundberg, Patric Nilsson, and Torbjörn Fagerström. Seed dormancy and frequency dependent selection due to sib competition: the effect of age specific gene expression. Journal of Theoretical Biology, 183(1):9–17, November 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519396901964.

Lappi:1992:CPR

[LOB92] Juha Lappi and Pauline Oker-Blom. Characterizing photosynthetic radiation response or other output function as a mean of element responses. Journal of Theoretical Biology, 154(3):371–389, February 7, 1992. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305801778.

Lara-Ochoa:1992:DMC

[LOH92] Francisco Lara-Ochoa and A. Herrera. A dynamical model for a co-operative enzyme. Journal of Theoretical Biology, 154(2):189–204, January 21, 1992. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305804023.

Lara-Ochoa:1991:DMP

[LOLM91] Francisco Lara-Ochoa and Esther Lara-Moreno. A dynamical model for phase transitions of lipids induced by calcium ions. Journal of Theoretical Biology, 148(3):295–304, February 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305802395.

Long:1994:LVR

[Lon94] Charles A. Long. Leonardo da Vinci’s rule and fractal complexity in dichotomous trees. Journal of Theoretical Biology, 167(2):107–113, March 21, 1994. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384710563.

Loreau:1990:CBA

[Lor90] M. Loreau. Competition between age classes, and the stability of stage-structured populations: a re-examination of Ebenman’s model. Journal of Theoretical Biology, 144(4):567–571, June 21, 1990. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305800906.

Lorberbaum:1994:NSE

[Lor94a] Jeffrey Lorberbaum. No strategy is evolutionarily stable in the repeated Prisoner’s dilemma. Journal of Theoretical Biology, 168(2):117–130, May 21, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519384710927.

Loreau:1994:MEF

[Lor94b] Michel Loreau. Mass and energy flow in closed ecosystems: Do ecological or mathematical constraints prevail? Journal of Theoretical Biology, 168(3):237–243, June 7, 1994. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384711052.

Lundberg:1993:OBS

[LP93] Sigfrid Lundberg and Lennart Persson. Optimal body size and resource density. Journal of Theoretical Biology, 164(2):163–180, September 21, 1993. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251938371146X.

Lammert:1996:IDV

[LPB96] P. E. Lammert, J. Prost, and R. Bruinsma. Ion drive for vesicles and cells. Journal of Theoretical Biology, 178(4):387–391, February 21, 1996. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396900351.

Lio:1996:HSB

[LPBR96] Pietro Liò, Antonio Politi, Marcello Buiatti, and Stefano Ruffo. High statistics block entropy measures of DNA sequences. Journal of Theoretical Biology, 180(2):151–160, May 21, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519396900910.

Lewalle:1995:WAO

[LPM95] Jacques Lewalle, Frank W. Peek, and Stephen J. Murphy. Wavelet analysis of olfactory nerve response to stimulus. Journal of Theoretical Biology, 177(3):215–236, December 7, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519385702408.

Larson:1994:TKA

[LPP^{+}94] Kenneth B. Larson, William H. Perman, Joel S. Perlmutter, Mokhtar H. Gado, John M. Ollinger, and Kenneth Zierler. Tracer-kinetic analysis for measuring regional cerebral blood flow by dynamic nuclear magnetic resonance imaging. Journal of Theoretical Biology, 170(1):1–14, September 7, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519384711647.

Lio:1996:AGP

[LPRB96] Pietro Liò, Antonio Politi, Stefano Ruffo, and Marcello Buiatti. Analysis of genomic patchiness of haemophilus influenzae and saccharomyces cerevisiae chromosomes. Journal of Theoretical Biology, 183(4):455–469, December 21, 1996. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396902350.

Lustig:1990:RKT

[LPS90] Cornel Lustig, Hanna Parnas, and Lee Segel. Release kinetics as a tool to describe drug effects on neurotransmitter release. Journal of Theoretical Biology, 144(2):225–248, May 14, 1990. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305803224.

Li:1994:MGC

[LQL94] Guo-Hua Li, Cheng-De Qin, and Mao-Hui Li. On the mechanisms of growth cone locomotion: Modeling and computer simulation. Journal of Theoretical Biology, 169(4):355–362, August 21, 1994. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519384711581.

Li:1992:NBP

[LQW92] Guo-Hua Li, Cheng-De Qin, and Zun-Sheng Wang. Neurite branching pattern formation: Modeling and computer simulation. Journal of Theoretical Biology, 157(4):463–486, August 21, 1992. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305806642.

Larsen:1994:SP

[LR94a] Poul Scheel Larsen and Hans Ulrik Riisgåd. The sponge pump. Journal of Theoretical Biology, 168(1):53–63, May 7, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519384710873.

Latta:1994:CFS

[LR94b] Robert Latta and Kermit Ritland. Conditions favoring stable mixed mating systems with jointly evolving inbreeding depression. Journal of Theoretical Biology, 170(1):15–23, September 7, 1994. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519384711659.

Li:1994:EIR

[LR94c] Yue-Xian Li and John Rinzel. Equations for InsP_{3} receptormediated [Ca^{2+}]_{i} oscillations derived from a detailed kinetic model: a Hodgkin–Huxley like formalism. Journal of Theoretical Biology, 166(4):461–473, February 21, 1994. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384710411.

Lio:1994:TCG

[LRB94] Pietro Lió, Stefano Ruffo, and Marcello Buiatti. Third codon G + C periodicity as a possible signal for an “internal” selective constraint. Journal of Theoretical Biology, 171(2):215–223, November 21, 1994. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519384712252.

Lemieux:1992:MDF

[LRJ92] D. R. Lemieux, F. A. Roberge, and D. Joly. Modeling the dynamic features of the electrogenic na,k pump of cardiac cells. Journal of Theoretical Biology, 154(3):335–358, February 7, 1992. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305801754.

LeBeau:1998:ARM

[LRMS98] Andrew P. LeBeau, A. Bruce Robson, Alan E. McKinnon, and James Sneyd. Analysis of a reduced model of cortiocotroph action potentials. Journal of Theoretical Biology, 192(3):319– 339, June 7, 1998. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519398906567.

Lemieux:1990:MSC

[LRS90] D. R. Lemieux, F. A. Roberge, and P. Savard. A model study of the contribution of active Na–K transport to membrane repolarization in cardiac cells. Journal of Theoretical Biology, 142(1):1–34, January 9, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305800104.

Lubkin:1996:EFL

[LRZ^{+}96] Sharon R. Lubkin, Joseph Romatowski, Mei Zhu, Paul M. Kulesa, and K. A. J. White. Evaluation of feline leukemia virus control measures. Journal of Theoretical Biology, 178(1):53–60, January 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396900065.

Lundberg:1994:POC

[LS94] Sigfrid Lundberg and Henrik G. Smith. Parent-offspring conflicts over reproductive efforts: Variations upon a theme by charnov. Journal of Theoretical Biology, 170(2):215– 218, September 21, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519384711805.

Liu:1997:CSP

[LSB^{+}97] C.-Q Liu, X.-F. Shi, C.-L. Bai, Jinshi Zhao, and Y. Wang. Comparative studies on parallel and antiparallel duplex and triplex DNA. Journal of Theoretical Biology, 184(3):319–326, February 7, 1997. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251939690274X.

Liu:1997:TSS

[LSG97] Lin Liu, Haijian Sun, and Aike Guo. Transformation of sensory signals into commands for saccadic eye movements: a neural network study. Journal of Theoretical Biology, 189(2):121–131, November 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519397904811.

Lucas:1992:EAC

[LSM^{+}92] M. L. Lucas, L. Sood, M. McGregor, N. Sattar, A. Watt, and J. C. Taylor. The effect of auxiliary conditions on intestinal unstirred layer diffusion modelled by numerical simulation. Journal of Theoretical Biology, 159(3):271–286, December 7, 1992. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305807246.

Liebovitch:1991:MIC

[LT91] Larry S. Liebovitch and Tibor I. Toth. A model of ion channel kinetics using deterministic chaotic rather than stochastic processes. Journal of Theoretical Biology, 148(2):243–267, January 21, 1991. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305803431.

Luo:1992:SEM

[LT92] Liaofu Luo and L. E. H. Trainor. A stochastic evolutionary model of molecular sequences. Journal of Theoretical Biology, 157(1):83–94, July 7, 1992. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305807581.

Loudon:1998:UDW

[LT98] Catherine Loudon and Antoinette Tordesillas. The use of the dimensionless Womersley number to characterize the unsteady nature of internal flow. Journal of Theoretical Biology, 191(1):63–78, March 7, 1998. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519397905646.

Li:1996:NNI

[LTZ96] Ming Li, John Tromp, and Louxin Zhang. On the nearest neighbour interchange distance between evolutionary trees. Journal of Theoretical Biology, 182(4):463–467, October 21, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519396901885.

Lubkin:1997:PFR

[Lub97] S. R. Lubkin. On pattern formation in reptilian dentition. Journal of Theoretical Biology, 186(2):145–157, May 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519396903495.

Lundberg:1990:EIS

[Lun90] Sigfrid Lundberg. Evolution of increased susceptibility to infectious diseases in age structured populations. Journal of Theoretical Biology, 147(1):133–142, November 7, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305802565.

Lushnikov:1994:SMS

[LVKH94] A. A. Lushnikov, T. Vesala, M. Kulmala, and P. Hari. A semiphenomenological model for stomatal gas transport. Journal of Theoretical Biology, 171(3):291–301, December 7, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251938471232X.

Lauck:1992:FCE

[LVL92] Leonardo Lauck, Aurea R. Vasconcellos, and Roberto Luzzi. On Fröhlich’s coherent effects in biological systems: Influence of carriers and high order dissipative effects. Journal of Theoretical Biology, 158(1):1–13, September 7, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305806459.

Lucas:1994:REP

[LW94] M. L. Lucas and R. R. Whitehead. A re-evaluation of the properties of the three-compartment model of intestinal weak-electrolyte absorption. Journal of Theoretical Biology, 167(2):147–159, March 21, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519384710599.

Liu:1991:QBA

[LWHZ91] C. Q. Liu, Y. Wang, J. F. Huang, and H. Zhang. A quantum biological approach to the relations of DNA methylation with gene transcription and mutation. Journal of Theoretical Biology, 148(2):145–155, January 21, 1991. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305803388.

Lewis:1997:SFR

[LWS97] R. D. Lewis, G. R. Warman, and D. S. Saunders. Simulations of free-running rhythms, light entrainment and the light-pulse phase response curves for the locomotor activity rhythm in period mutant of Drosophila melanogaster. Journal of Theoretical Biology, 185(4):503–510, April 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396903306.

Marinier:1995:CAF

[MA95] S. L. Marinier and A. J. Alexander. Coprophagy as an avenue for foals of the domestic horse to learn food preferences from their dams. Journal of Theoretical Biology, 173(2):121–124, March 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519385700495.

Minetti:1997:TMC

[MA97a] A. E. Minetti and R. McN. Alexander. A theory of metabolic costs for bipedal gaits. Journal of Theoretical Biology, 186(4):467–476, June 21, 1997. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519397904070.

Monetti:1997:ELS

[MA97b] Roberto A. Monetti and Ezequiel V. Albano. On the emergence of large-scale complex behavior in the dynamics of a society of living individuals: the stochastic game of life. Journal of Theoretical Biology, 187(2):183–194, July 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519397904240.

Mosquera:1998:EVU

[MA98] J. Mosquera and Frederick R. Adler. Evolution of virulence: a unified framework for coinfection and superinfection. Journal of Theoretical Biology, 195(3):293–313, December 7, 1998. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519398907937.

Munoz:1999:SSR

[MA99] Miguel A. Muñoz and Jose D. Acuña. Sample size requirements of a mixture analysis method with applications in systematic biology. Journal of Theoretical Biology, 196(2):263–265, January 21, 1999. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519398908268.

Mendoza:1998:DGR

[MAB98] Luis Mendoza and Elena R. Alvarez-Buylla. Dynamics of the genetic regulatory network for Arabidopsis thaliana flower morphogenesis. Journal of Theoretical Biology, 193(2):307–319, July 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519398907019.

Mackay:1990:IRP

[Mac90a] Dennis Mackay. Interpretation of relative potencies, relative efficacies and apparent affinity constants of agonist drugs estimated from concentration-response curves. Journal of Theoretical Biology, 142(3):415–427, February 9, 1990. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305805600.

MacLeod:1990:IIC

[Mac90b] Michael C. MacLeod. The importance of intercalation in the covalent binding of benzo(a)pyrene diol epoxide to DNA. Journal of Theoretical Biology, 142(1):113–122, January 9, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305800165.

Miftakhov:1999:NSMa

[MAC99a] R. N. Miftakhov, G. R. Abdusheva, and J. Christensen. Numerical simulation of motility patterns of the small bowel.

1. formulation of a mathematical model. Journal of Theoretical Biology, 197(1):89–112, March 7, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519398908591.

Miftakhov:1999:NSMb

[MAC99b] R. N. Miftakhov, G. R. Abdusheva, and J. Christensen. Numerical simulation of motility patterns of the small bowel. II. comparative pharmacological validation of a mathematical model. Journal of Theoretical Biology, 200(3):261–290, October 7, 1999. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519399909918.

Maguire:1992:EM

[Mag92] Marjorie P. Maguire. The evolution of meiosis. Journal of Theoretical Biology, 154(1):43–55, January 7, 1992. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305801870.

Maguire:1995:SRD

[Mag95] M. P. Maguire. Sexual reproduction: the dawn of enabling chromosome maneuvers. Journal of Theoretical Biology, 173(2):213–215, March 21, 1995. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519385700574.

Mendez-Alvarez:1997:MMD

[MAGO97] S. Méndez-Alvarez, N. Gaju, and B. Oliva. A mathematical model to determine the optimal number of fragments for comparison of bacterial chromosomic macrorestriction patterns. Journal of Theoretical Biology, 185(3):367–372, April 7, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519396903148.

Mahaffy:1993:VCR

[Mah93] Joseph M. Mahaffy. Variation in concentrations of RNAs and proteins involved in gene expression of Escherichia coli. Journal of Theoretical Biology, 162(2):153–186, May 21, 1993. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519383710817.

Malec:1994:KMC

[Mal94a] Przemysław Malec. Kinetic modelling of chloroplast phototranslocations in Lemna trisulca L.: Two rate limiting components? Journal of Theoretical Biology, 169(2):189–195, July 21, 1994. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519384711404.

Mallios:1994:MRA

[Mal94b] Ronna R. Mallios. Multiple regression analysis suggests motifs for class II MHC binding. Journal of Theoretical Biology, 166(2):167–172, January 21, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519384710150.

Maley:1997:MRA

[Mal97] Carlo Maley. Mutation rates as adaptations. Journal of Theoretical Biology, 186(3):339–348, June 7, 1997. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396903628.

Mangel:1990:DIU

[Man90] Marc Mangel. Dynamic information in uncertain and changing worlds. Journal of Theoretical Biology, 146(3):317–332, October 7, 1990. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305807428.

Mannie:1991:UMC

[Man91] Mark D. Mannie. A unified model for T cell antigen recognition and thymic selection of the T cell repertoire. Journal of Theoretical Biology, 151(2):169–192, July 21, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305803601.

Mani:1992:CBC

[Man92a] G. S. Mani. Correlations between the coding and noncoding regions in DNA. Journal of Theoretical Biology, 158(4):429–445, October 21, 1992. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305807088.

Mani:1992:LRD

[Man92b] G. S. Mani. Long-range doublet correlations in DNA and the coding regions. Journal of Theoretical Biology, 158(4):447–464, October 21, 1992. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251930580709X.

Mantovani:1995:HPP

[Man95] Bernardo Mantovani. Homeostasis perturbation (physiological stress) and the mechanism of biological evolution-the possible role of active oxygen species. Journal of Theoretical Biology, 176(1):193–194, September 7, 1995. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385701890.

Martirosov:1990:DTE

[Mar90] S. M. Martirosov. Direct transduction of energy by dithioldisulfide interconversion. Journal of Theoretical Biology, 144 (1):69–74, May 10, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519305802991.

Marin:1991:SID

[Mar91] Ignacio Marin. Sexual isolation in Drosophila I. Theoretical models for multiple-choice experiments. Journal of Theoretical Biology, 152(2):271–284, September 21, 1991. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305804588.

Marek:1992:DSI

[Mar92a] J. Marek. A difference in the shapes of intermolecular potentials between phospholipid model molecules. relation to the ripple phase. Journal of Theoretical Biology, 159(4):417–429, December 21, 1992. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305806873.

Marom:1992:DAV

[Mar92b] Shimon Marom. A 3-D approach to voltage clamp data. Journal of Theoretical Biology, 154(4):475–484, February 21, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305801638.

Margineanu:1993:DTS

[Mar93a] D.-G. Margineanu. Dimensionality, time scales and membrane thermodynamics. Journal of Theoretical Biology, 164 (3):403–406, October 7, 1993. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002251938371163X.

Mari:1993:CMI

[Mar93b] Andrea Mari. Circulatory models of intact-body kinetics and their relationship with compartmental and non-compartmental analysis. Journal of Theoretical Biology, 160(4):509–531, February 21, 1993. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519383710337.

Mark:1993:SER

[Mar93c] Mark. Seeking the evolutionary regression coefficient: An analysis of what comparative methods measure. Journal of Theoretical Biology, 164(2):191–205, September 21, 1993. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519383711483.

Mari:1995:COW

[Mar95a] Andrea Mari. Calculation of organ and whole-body uptake and production with the impulse response approach. Journal of Theoretical Biology, 174(3):341–353, June 7, 1995. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385701038.

Markos:1995:OGR

[Mar95b] Anton Markos. The ontogeny of gaia:: The role of microorganisms in planetary information network. Journal of Theoretical Biology, 176(1):175–180, September 7, 1995. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385701865.

Maroun:1996:IAC

[Mar96] Leonard E. Maroun. Interferon action and chromosome 21 trisomy (down syndrome): 15 years later. Journal of Theoretical Biology, 181(1):41–46, July 7, 1996. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396901137.

Marcus:1997:DIC

[Mar97a] Yehouda Marcus. Distribution of inorganic carbon among its component species in cyanobacteria: Do cyanobacteria in fact actively accumulate inorganic carbon? Journal of Theoretical Biology, 185(1):31–45, March 7, 1997. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251939690291X.

Marin:1997:SID

[Mar97b] Ignacio Marín. Sexual isolation in drosophila. III. estimating isolation using male-choice experiments. Journal of Theoretical Biology, 188(4):515–524, October 21, 1997. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519397904926.

Mathews:1991:MCD

[Mat91] Christopher K. Mathews. Metabolite channelling in deoxyribonucleotide and DNA biosynthesis. Journal of Theoretical Biology, 152(1):25–28, September 7, 1991. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305805028.

Matsumoto:1993:SRA

[Mat93] Joji Matsumoto. Speculations on the roles of ADPribosyl transferase based on analogies between RecA and poly(ADP-ribose) polymerase. Journal of Theoretical Biology, 164(3):307–319, October 7, 1993. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519383711562.

Matsuno:1995:CPM

[Mat95] Koichiro Matsuno. Consumer power as the major evolutionary force. Journal of Theoretical Biology, 173(2):137–145, March 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519385700513.

Matsson:1996:RTN

[Mat96] Leif Matsson. Response theory for non-stationary ligandreceptor interaction and a solution to the growth signal firing problem. Journal of Theoretical Biology, 180(2):93–104, May 21, 1996. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396900843.

Maurice:1992:MNC

[Mau92] Sandrine Maurice. Maintenance of nuclear-cytoplasmic polymorphism under dioecious reproductive systems. Journal of Theoretical Biology, 154(2):239–247, January 21, 1992. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305804059.

Mayhew:1990:SAO

[May90] Terry M. Mayhew. The surface area of an object revisitedbut from random directions. Journal of Theoretical Biology, 144(2):259–265, May 14, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519305803248.

Mazur:1991:PVS

[Maz91] Alexey K. Mazur. A probabilistic view on steady-state enzyme reactions. Journal of Theoretical Biology, 148(2):229–242, January 21, 1991. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251930580342X.

Morel:1990:MAF

[MB90] Jean E. Morel and Nacima Bachouchi. Movement of actin filaments along tracks of myosin heads. Journal of Theoretical Biology, 145(1):135–136, July 9, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305805405.

McManus:1992:HPS

[MB92] I. C. McManus and M. P. Bryden. Handedness of parents and sex of progeny: Failure to replicate the results of James (1986). Journal of Theoretical Biology, 159(4):439–442, December 21, 1992. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305806897.

Mozrzymas:1993:DNB

[MB93] Jerzy W. Mozrzymas and Michal Bartoszkiewicz. The discrete nature of biological membrane conductance, channel interaction through electrolyte layers and the cable equation. Journal of Theoretical Biology, 162(3):371–380, June 7, 1993. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519383710933.

McKenzie:1997:DPF

[MB97] F. Ellis McKenzie and William H. Bossert. The dynamics of plasmodium falciparum blood-stage infection. Journal of Theoretical Biology, 188(1):127–140, September 7, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519397904781.

McKenzie:1998:OPG

[MB98] F. Ellis McKenzie and William H. Bossert. The optimal production of gametocytes by plasmodium falciparum. Journal of Theoretical Biology, 193(3):419–428, August 7, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251939890710X.

Mander:1995:RDS

[MBC95] A. P. Mander, P. G. Blackwell, and C. Cannings. A resource-driven strategy conflict in two dimensions. Journal of Theoretical Biology, 177(1):17–28, November 7, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385702214.

Mahaffy:1998:HMM

[MBM98] Joseph M. Mahaffy, Jacques Bélair, and Michael C. Mackey. Hematopoietic model with moving boundary condition and state dependent delay: Applications in erythropoiesis. Journal of Theoretical Biology, 190(2):135–146, January 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519397905373.

Morel:1992:CFA

[MBS92] Jean E. Morel and Nacima Bachouchi-Salhi. Comments on “Further analysis of the molecular jet hypothesis during muscle contraction”. Journal of Theoretical Biology, 154(3):403–404, February 7, 1992. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305801808.

Mattheck:1993:SFT

[MBS93] C. Mattheck, K. Bethge, and J. Schäfer. Safety factors in trees. Journal of Theoretical Biology, 165(2):185–189, November 21, 1993. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519383711847.

Merino:1995:IMT

[MBSB95] Gonzalo Merino, Winfried Boos, Howard A. Shuman, and Erich Bohl. The inhibition of maltose transport by the unliganded form of the maltose-binding protein of Escherichia coli: Experimental findings and mathematical treatment. Journal of Theoretical Biology, 177(2):171–179, November 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519385702366.

Morel:1992:SLM

[MBSM92] Jean E. Morel, Nacima Bachouchi-Salhi, and Zalika Merah. Shape and length of myosin heads. Journal of Theoretical Biology, 156(1):73–76, IN1–IN2, 77–90, May 7, 1992. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305806575.

Morel:1993:VMA

[MBSM93] J. E. Morel, N. Bachouchi-Salhi, and Z. Merah. In vitro movement of actin filaments over myosin: Role of actin and related problems. Journal of Theoretical Biology, 160(2):265–269, January 21, 1993. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519383710180.

Marrow:1993:EIP

[MC93] Paul Marrow and Chris Cannings. Evolutionary instability in predator–prey systems. Journal of Theoretical Biology, 160(2):135–150, January 21, 1993. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519383710088.

Mitsui:1996:PMH

[MC96] Toshio Mitsui and Hisahiro Chiba. Proposed modification of the Huxley-Simmons model for myosin head motion along an actin filament. Journal of Theoretical Biology, 182(2):147–159, September 21, 1996. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519396901496.

McArthur:1991:TIB

[McA91] A. J. McArthur. Thermal interaction between animal and microclimate: Specification of a “standard environmental temperature” for animals outdoors. Journal of Theoretical Biology, 148(3):331–343, February 7, 1991. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305802413.

McCorkle:1991:PBB

[McC91] Richard A. McCorkle. A physical basis for biochemistry. Journal of Theoretical Biology, 148(3):393–400, February 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305802449.

McCallum:1992:EIC

[McC92] H. I. McCallum. Effects of immigration on chaotic population dynamics. Journal of Theoretical Biology, 154(3):277–284, February 7, 1992. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305801705.

McCollum:1993:RIS

[McC93] Gin McCollum. Reciprocal inhibition, synergies, and movements. Journal of Theoretical Biology, 165(3):291–311, December 7, 1993. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519383711902.

McCollum:1994:NSD

[McC94] Gin McCollum. Navigating a set of discrete regions in body position space. Journal of Theoretical Biology, 167(3):263–271, April 7, 1994. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251938471068X.

McGeer:1993:DCB

[McG93] Tad McGeer. Dynamics and control of bipedal locomotion. Journal of Theoretical Biology, 163(3):277–314, August 7, 1993. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519383711215.

Mandecki:1995:MMB

[MCG95] Wlodek Mandecki, Jack Y.-C. Chen, and Nelson Grihalde. A mathematical model for biopanning (affinity selection) using peptide libraries on filamentous phage. Journal of Theoretical Biology, 176(4):523–530, October 21, 1995. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385702184.

Munoz-Chapuli:1997:RDM

[MCGPP97] Ramón Muñoz-Chápuli, Alejandro Gallego, and José María Pérez-Pomares. A reaction-diffusion model can account for the anatomical pattern of the cardiac conal valves in fish. Journal of Theoretical Biology, 185(2):233–240, March 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519396903033.

McIlwain:1995:LGL

[McI95] James T. McIlwain. Lateral geniculate lamination and the corticogeniculate projection: a potential role in binocular vision in the quadrants. Journal of Theoretical Biology, 172(4):329–333, February 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519385700306.

McKee:1995:TPS

[McK95] Jeffrey K. McKee. Turnover patterns and species longevity of large mammals from the Late Pliocene and Pleistocene of Southern Africa: a comparison of simulated and empirical data. Journal of Theoretical Biology, 172(2):141–147, January 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519385700112.

McLean:1994:MCM

[McL94] Angela R. McLean. Modelling T cell memory. Journal of Theoretical Biology, 170(1):63–74, September 7, 1994. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384711684.

McNamara:1993:EPS

[McN93] John M. McNamara. Evolutionary paths in strategy space: An improvement algorithm for life-history strategies. Journal of Theoretical Biology, 161(1):23–37, March 7, 1993. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519383710374.

McNamara:1994:MSA

[McN94a] John M. McNamara. Multiple stable age distributions in a population at evolutionary stability. Journal of Theoretical Biology, 169(4):349–354, August 21, 1994. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251938471157X.

McNamara:1994:TED

[McN94b] John M. McNamara. Timing of entry into diapause: Optimal allocation to “growth” and “reproduction” in a stochastic environment. Journal of Theoretical Biology, 168(2):201–209, May 21, 1994. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251938471099X.

McNair:1995:OPD

[McN95] James N. McNair. Ontogenetic patterns of density-dependent mortality: contrasting stability effects in populations with adult dominance. Journal of Theoretical Biology, 175(2):207– 230, July 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519385701348.

Miramontes:1996:NDS

[MD96] Octavio Miramontes and Og DeSouza. The nonlinear dynamics of survival and social facilitation in termites. Journal of Theoretical Biology, 181(4):373–380, August 21, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396901381.

Marks:1994:SIF

[MDGB94] Arnost Marks, James C. Deutsch, and T. H. Glutton-Brock. Stochastic influences, female copying and the intensity of sexual selection on leks. Journal of Theoretical Biology, 170(2):159–162, September 21, 1994. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384711751.

Muller:1998:DKJ

[MdR98] M. Muller and M. de Ruijter. The derivation of knee joint types from the geometry of the cruciate ligament four-bar system. Journal of Theoretical Biology, 193(3):507–518, August 7, 1998. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519398907196.

Moore:1993:FMB

[ME93] Anne M. F. Moore and Olaf Ellers. A functional morphospace, based on dimensionless numbers, for a circumferential, calcite, stabilizing structure in sand dollars. Journal of Theoretical Biology, 162(2):253–266, May 21, 1993. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519383710866.

Medler:1995:PTM

[Med95] S. Medler. Physiological time and the measurement of glomerular filtration rate. Journal of Theoretical Biology, 176(1):59–65, September 7, 1995. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519385701762.

Meggs:1990:EPP

[Meg90] William J. Meggs. Enhanced polymerization of polar macromolecules by an applied electric field with application to mitosis. Journal of Theoretical Biology, 145(2):245–255, July 24, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305801298.

Melcher:1993:HPS

[Mel93a] Ulrich Melcher. HIV-1 proteinase as structural model of intercellular transport proteins of plant viruses. Journal of Theoretical Biology, 162(1):61–74, May 7, 1993. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519383710763.

Melton:1993:BSE

[Mel93b] Robert H. Melton. Body size evolution and the concept of alloenergetic change. Journal of Theoretical Biology, 164(3):321–339, October 7, 1993. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519383711574.

Morton-Firth:1998:PTF

[MFB98] Carl Jason Morton-Firth and Dennis Bray. Predicting temporal fluctuations in an intracellular signalling pathway. Journal of Theoretical Biology, 192(1):117–128, May 7, 1998. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519397906512.

Maranon:1999:MDS

[MFG99] J. Marañón, A. Fantoni, and J. R. Grigera. Molecular dynamics simulation of double proton transfer: Adenine-thymine base pair. Journal of Theoretical Biology, 201(2):93–102, November 21, 1999. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519399910081.

Mehr:1995:LDI

[MFHA^{+}95] Ramit Mehr, Masha Fridkis-Hareli, Loya Abel, Lee Segel, and Amiela Globerson. Lymphocyte development in irradiated thymuses: Dynamics of colonization by progenitor cells and regeneration of resident cells. Journal of Theoretical Biology, 177(2):181–192, November 21, 1995. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385702378.

Maas:1997:MMP

[MFM97] J. A. Maas, J. France, and B. W. McBride. Model of milk protein synthesis. a mechanistic model of milk protein synthesis in the lactating bovine mammary gland. Journal of Theoretical Biology, 187(3):363–378, August 7, 1997. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519397904379.

Mancheno:1995:PSC

[MGL^{+}95] Jose M. Mancheño, Maria Gasset, Javier Lacadena, Alvaro Martínez del Pozo, Mercedes Oñaderra, and Jose G. Gavilanes. Predictive study of the conformation of the cytotoxic protein α-sarcin: a structural model to explain αsarcin-membrane interaction. Journal of Theoretical Biology, 172(3):259–267, February 7, 1995. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519385700227.

Mesterton-Gibbons:1996:WAA

[MGMD96] Michael Mesterton-Gibbons, James H. Marden, and Lee Alan Dugatkin. On wars of attrition without assessment. Journal of Theoretical Biology, 181(1):65–83, July 7, 1996. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396901150.

Mehr:1995:MPN

[MGP95] Ramit Mehr, Amiela Globerson, and Alan S. Perelson. Modeling positive and negative selection and differentiation processes in the thymus. Journal of Theoretical Biology, 175(1):103– 126, July 7, 1995. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519385701245.

Minsky:1997:NSC

[MGR97] Abraham Minsky, Rodolfo Ghirlando, and Ziv Reich. Nucleosomes: a solution to a crowded intracellular environment? Journal of Theoretical Biology, 188(3):379–385, October 7, 1997. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519397905257.

Matsuda:1990:ESS

[MH90] Hiroyuki Matsuda and Yasushi Harada. Evolutionarily stable stalk to spore ratio in cellular slime molds and the law of equalization in net incomes. Journal of Theoretical Biology, 147(3):329–344, December 7, 1990. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305804916.

Mikhailov:1995:FLC

[MH95] Alexander Mikhailov and Benno Hess. Fluctuations in living cells and intracellular traffic. Journal of Theoretical Biology, 176(1):185–192, September 7, 1995. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385701889.

McVean:1996:GCP

[MH96] Gilean McVean and Laurence D. Hurst. Genetic conflicts and the paradox of sex determination: Three paths to the evolution of female intersexuality in a mammal. Journal of Theoretical Biology, 179(3):199–211, April 7, 1996. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396900612.

Marzec:1997:DMM

[MH97] Christopher J. Marzec and Richard W. Hendrix. A dynamic model for the morphogenesis of the late vertebrate lens. Journal of Theoretical Biology, 186(3):349–372, June 7, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519396903549.

McCollum:1995:FEW

[MHC95] Gin McCollum, Clay Holroyd, and Ann M. Castelfranco. Forms of early walking. Journal of Theoretical Biology, 176(3):373–390, October 7, 1995. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519385702068.

Melendez-Hevia:1991:CEM

[MHM91] Enrigue Meléndez-Hevia and Francisco Montero. Channelling and evolution of metabolism. Journal of Theoretical Biology, 152(1):77–79, September 7, 1991. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305805156.

Marin:1994:MMF

[MHM94] J. L. Marin, M. Huerta, and J. Muñz. Modulation of molecular friction in actomyosin complex by an external electric field. Journal of Theoretical Biology, 167(2):167–168, March 21, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519384710617.

Morton:1994:SHR

[MHN^{+}94] Thomas L. Morton, James W. Haefner, Vasudevarao Nugala, Robert D. Decino, and Lloyd Mendes. The selfish herd revisited: Do simple movement rules reduce relative predation risk? Journal of Theoretical Biology, 167(1):73–79, March 7, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519384710514.

Melendez-Hevia:1996:MBW

[MHSR^{+}96] Enrique Meléndez-Hevia, Joaquín Sicilia, María Teresa Ramos, Enric I. Canela, and Marta Cascante. Molecular bureaucracy: Who controls the delays? Journal of Theoretical Biology, 182(3):333–339, October 7, 1996. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396901721.

Melendez-Hevia:1990:GMC

[MHTS90] Enrique Meléndez-Hevia, Néstor V. Torres, and Joaquín Sicilia. A generalization of metabolic control analysis to conditions of no proportionality between activity and concentration of enzymes. Journal of Theoretical Biology, 142(4):443–451, February 22, 1990. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305801006.

McNamara:1993:CPC

[MHW93] John M. McNamara, Alasdair I. Houston, and Wolfgang W. Weisser. Combining prey choice and patch use-what does rate-maximizing predict? Journal of Theoretical Biology, 164(2):219–238, September 21, 1993. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519383711501.

Melendez-Hevia:1994:OME

[MHWM94] Enrique Meléndez-Hevia, Thomas G. Waddell, and Francisco Montero. Optimization of metabolism: The evolution of metabolic pathways toward simplicity through the game of the pentose phosphate cycle. Journal of Theoretical Biology, 166(2):201–220, January 21, 1994. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384710186.

Matsushima:1998:EST

[MI98] Masanao Matsushima and Takashi Ikegami. Evolution of strategies in the three-person iterated Prisoner’s dilemma game. Journal of Theoretical Biology, 195(1):53–67, November 7, 1998. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519398907809.

Miller:1991:UPR

[Mil91a] Dale Armin Miller. Useful perspective on the relation between biological and physical descriptions of phenomena. Journal of Theoretical Biology, 152(3):341–355, October 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305801997.

Mills:1991:CCG

[Mil91b] Gordon C. Mills. Cytochrome c: Gene structure, homology and ancestral relationships. Journal of Theoretical Biology, 152(2):177–190, September 21, 1991. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305804515.

Milchtaich:1992:GAG

[Mil92a] Igal Milchtaich. A general approach to genetic equilibria with an uneven sex ratio. Journal of Theoretical Biology, 157(3):373–381, August 7, 1992. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305806162.

Millonas:1992:CTM

[Mil92b] Mark M. Millonas. A connectionist type model of selforganized foraging and emergent behavior in ant swarms. Journal of Theoretical Biology, 159(4):529–552, December 21, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305806976.

Minetti:1994:CDA

[Min94] Alberto E. Minetti. Contraction dynamics in antagonist muscles. Journal of Theoretical Biology, 169(3):295–304, August 7, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519384711507.

Mironov:1990:TAC

[Mir90] Sergej L. Mironov. Theoretical analysis of Ca wave propagation along the surface of intracellular stores. Journal of Theoretical Biology, 146(1):87–97, September 7, 1990. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305800451.

Mira:1998:WMA

[Mir98] Alex Mira. Why is meiosis arrested? Journal of Theoretical Biology, 194(2):275–287, September 21, 1998. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519398907615.

Mittler:1996:EGS

[Mit96a] John E. Mittler. Evolution of the genetic switch in temperate bacteriophage. I. Basic theory. Journal of Theoretical Biology, 179(2):161–172, March 14, 1996. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396900569.

Mochizuki:1996:SMC

[MIT96b] Atsushi Mochizuki, Yoh Iwasa, and Yasuhiko Takeda. A stochastic model for cell sorting and measuring cell–cell adhesion. Journal of Theoretical Biology, 179(2):129–146, March 14, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519396900545.

Mizejewski:1995:AFS

[Miz95] G. J. Mizejewski. Alpha-fetoprotein signal sequences: a proposed mechanism for subcellular localization and organelle targeting. Journal of Theoretical Biology, 176(1):103–113, September 7, 1995. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519385701804.

Moore:1998:PMS

[MJIB98] G. P. M. Moore, N. Jackson, K. Isaacs, and G. Brown. Pattern and morphogenesis in skin. Journal of Theoretical Biology, 191(1):87–94, March 7, 1998. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519397905671.

MacLeod:1996:POO

[MJLW96] Michael C. MacLeod, Dennis A. Johnston, Michael LaBate, and R. Allen White. The probability of occurrence of oligomer motifs in the human genome and genomic microheterogeneity. Journal of Theoretical Biology, 181(4):311–318, August 21, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519396901332.

McLean:1990:MHI

[MK90] A. R. McLean and T. B. L. Kirkwood. A model of human immunodeficiency virus infection in T helper cell clones. Journal of Theoretical Biology, 147(2):177–203, November 21, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305800517.

Mazur:1992:SMP

[MK92] Alexey K. Mazur and Alexander V. Kuchinski. Schematic methods for probabilistic enzyme kinetics. Journal of Theoretical Biology, 155(3):387–407, April 7, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305806058.

Mrigank:1993:CSI

[MK93] Mrigank and V. Kothekar. Computer simulation of interaction of deacylcortivazol with d(TGTTCT)2. Journal of Theoretical Biology, 163(1):33–50, July 1, 1993. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519383711057.

Mahner:1997:WEG

[MK97a] Martin Mahner and Michael Kary. What exactly are genomes, genotypes and phenotypes? and what about phenomes? Journal of Theoretical Biology, 186(1):55–63, May 7, 1997. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396903355.

McFadden:1997:EES

[MK97b] Johnjoe McFadden and Greg Knowles. Escape from evolutionary stasis by transposon-mediated deleterious mutations. Journal of Theoretical Biology, 186(4):441–447, June 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519397904033.

Mitsui:1998:IPM

[MKC^{+}98] T. Mitsui, S. Kumagai, H. Chiba, H. Yoshimura, and H. Ohshima. Induced potential model for muscular contraction mechanism, including two attached states of myosin head. Journal of Theoretical Biology, 192(1):35–41, May 7, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519397906147.

Macleod:1998:PQS

[MKGB98] G. T. Macleod, V. Khurana, W. G. Gibson, and M. R. Bennett. Probability of quantal secretion and the mobilization of vesicles at the active zones of endplates. Journal of Theoretical Biology, 191(3):323–334, April 7, 1998. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519397905932.

Maier:1996:ALD

[MKM96] Stefan Maier, Stephen Kaattari, and Mario E. Magaña. The analysis of limiting dilution assays conducted under non-saturating conditions. Journal of Theoretical Biology, 179(2):173–177, March 14, 1996. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519396900570.

Marienhagen:1997:KME

[MKZ97] Jörg Marienhagen, Bernhard U. Keller, and Annette Zippelius. Kinetic model of excitatory synaptic transmission to cerebellar Purkinje cells. Journal of Theoretical Biology, 188(2):227–240, September 21, 1997. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519397904641.

Michelson:1994:DRR

[ML94a] Seth Michelson and John T. Leith. Dormancy, regression, and recurrence: Towards a unifying theory of tumor growth control. Journal of Theoretical Biology, 169(4):327–338, August 21, 1994. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519384711556.

Musila:1994:IIC

[ML94b] Miroslav Musila and Petr Lánský. On the interspike intervals calculated from diffusion approximations of Stein’s neuronal model with reversal potentials. Journal of Theoretical Biology, 171(2):225–232, November 21, 1994. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384712264.

Morton:1997:RSP

[ML97a] R. Daniel Morton and Richard Law. Regional species pools and the assembly of local ecological communities. Journal of Theoretical Biology, 187(3):321–331, August 7, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519397904197.

Mroz:1997:DPB

[ML97b] Iwona Mróz and Zdzislaw Latajka. Direction of protein biosynthesis as a reflection of the prebiotic environment. Journal of Theoretical Biology, 189(2):151–158, November 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519397905026.

McLeod:1992:EGI

[MLS92] Bruce R. McLeod, Abraham R. Liboff, and Stephen D. Smith. Electromagnetic gating in ion channels. Journal of Theoretical Biology, 158(1):15–16, IN1, 17–31, September 7, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305806460.

Murray:1991:PPF

[MM91] J. D. Murray and M. R. Myerscough. Pigmentation pattern formation on snakes. Journal of Theoretical Biology, 149(3):339– 360, April 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305803108.

Morel:1995:MCC

[MM95] Jean-Emile Morel and Zalika Merah. Muscle contraction: Are the crossbridges independent force generators? Journal of Theoretical Biology, 176(3):431–432, October 7, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251938570210X.

Morel:1997:SLM

[MM97] Jean-Emile Morel and Zalika Merah. Shape and length of myosin heads and radial compressive forces in muscle. Journal of Theoretical Biology, 184(2):133–138, January 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519396902684.

Moller:1998:DSS

[MM98] Anders Pape Møller and Mark. Developmental stability and signalling among cells. Journal of Theoretical Biology, 193(3):497–506, August 7, 1998. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519398907184.

Mason:1999:BSD

[MMB99] Daniel P. Mason, F. Ellis McKenzie, and William H. Bossert. The blood-stage dynamics of mixed Plasmodium malariae–Plasmodium falciparum infections. Journal of Theoretical Biology, 198(4):549–566, June 21, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519399909323.

Marin:1990:FAM

[MMH90] J. L. Marin, J. Muñiz, and M. Huerta. Further analysis of the molecular jet hypothesis during muscle contraction. Journal of Theoretical Biology, 147(3):373–376, December 7, 1990. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305804941.

McLean:1992:MIB

[MN92] Angela R. McLean and Martin A. Nowak. Models of interactions between HIV and other pathogens. Journal of Theoretical Biology, 155(1):69–86, March 7, 1992. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305805491.

May:1994:SMD

[MN94] Robert M. May and Martin A. Nowak. Superinfection, metapopulation dynamics, and the evolution of diversity. Journal of Theoretical Biology, 170(1):95–114, September 7, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519384711714.

Machemer:1997:ADC

[MNB97] Hans Machemer, Ute Nagel, and Richard Bräucker. Assessment of g-dependent cellular gravitaxis: Determination of cell orientation from locomotion track. Journal of Theoretical Biology, 185(2):201–211, March 21, 1997. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251939690301X.

McNair:1997:TTS

[MNH97] James N. McNair, J. Denis Newbold, and David D. Hart. Turbulent transport of suspended particles and dispersing benthic organisms: How long to hit bottom? Journal of Theoretical Biology, 188(1):29–52, September 7, 1997. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519397904537.

Miller:1993:CTS

[MO93] R. L. Miller and T. Ohkawa. Chain terminator sequencing of double-stranded DNA with built-in error correction. Journal of Theoretical Biology, 161(4):407–429, April 21, 1993. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519383710659.

Matsuno:1996:HMT

[MO96] Koichiro Matsuno and Nobuaki Ono. How many trophic levels are there? Journal of Theoretical Biology, 180(2):105–109, May 21, 1996. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396900855.

Mochizuki:1999:ECL

[Moc99] Atsushi Mochizuki. Estimating cell lineage from distributions of randomly introduced markers. Journal of Theoretical Biology, 197(2):227–245, March 21, 1999. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519398908700.

Mogi:1993:GAC

[Mog93] Kenichiro Mogi. Graphic analysis of coupling in biological systems. Journal of Theoretical Biology, 162(3):337–352, June 7, 1993. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251938371091X.

Moller:1997:DSA

[Møl97] Anders Pape Møller. Developmental selection against developmentally unstable offspring and sexual selection. Journal of Theoretical Biology, 185(4):415–422, April 21, 1997.

CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251939690332X.

Morel:1990:VIM

[Mor90] Jean E. Morel. Velocity-induced modifications in the crossbridge and/or the actin filament behavior during shortening of muscle fibers. Journal of Theoretical Biology, 146(3):347–354, October 7, 1990. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305807453.

Morel:1991:IFE

[Mor91] Jean E. Morel. The isometric force exerted per myosin head in a muscle fibre is 8 pN. consequence on the validity of the traditional concepts of force generation. Journal of Theoretical Biology, 151(2):285–288, July 21, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305803649.

Morris:1992:SPC

[Mor92] Valerie B. Morris. Spatial patterns of cells in dividing epithelia. Journal of Theoretical Biology, 159(3):343–352, IN1, 353–360, December 7, 1992. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305807295.

Morel:1996:MHS

[Mor96] Jean-Emile Morel. Myosin heads in solution: How to interpret the data obtained by means of X-ray and neutron scattering? Journal of Theoretical Biology, 179(1):87–90, March 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251939690051X.

Mori:1998:OPU

[Mor98] Yoshihisa Mori. The optimal patch use in divers: Optimal time budget and the number of dive cycles during bout. Journal of Theoretical Biology, 190(2):187–199, January 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519397905506.

Morris:1999:KEC

[Mor99] J. A. Morris. The kinetics of epithelial cell generation: Its relevance to cancer and ageing. Journal of Theoretical Biology, 199(1):87–95, July 7, 1999. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519399909438.

Motro:1991:COD

[Mot91] Uzi Motro. Co-operation and defection: Playing the field and the ESS. Journal of Theoretical Biology, 151(2):145–154, July 21, 1991. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305803583.

Motro:1993:HPN

[Mot93] Uzi Motro. Helpers at parents’ nest: a game theoretic approach. Journal of Theoretical Biology, 163(1):127–134, July 1, 1993. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519383711112.

Mowbray:1997:ESS

[Mow97] Miranda Mowbray. Evolutionarily stable strategy distributions for the repeated Prisoner’s dilemma. Journal of Theoretical Biology, 187(2):223–229, July 21, 1997. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519397904318.

Muthusamy:1991:ISC

[MP91] R. Muthusamy and P. K. Ponnuswamy. The influence of side chain structures on the residue-residue associations in globular proteins. Journal of Theoretical Biology, 153(1):25–40, November 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305803510.

Moller:1993:PEG

[MP93] Anders Pape Møller and Andrew Pomiankowski. Punctuated equilibria or gradual evolution: Fluctuating asymmetry and variation in the rate of evolution. Journal of Theoretical Biology, 161(3):359–367, April 7, 1993. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519383710611.

Monteiro:1998:MEE

[MP98a] L. H. A. Monteiro and J. R. C. Piqueira. A model for the early evolution of self-replicating polymers. Journal of Theoretical Biology, 191(3):237–248, April 7, 1998. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519397905683.

Monteiro:1998:NSRa

[MP98b] L. H. A. Monteiro and J. R. C. Piqueira. Numbering selfreplicating polymers. Journal of Theoretical Biology, 193 (2):365–367, July 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519398906993.

Monteiro:1998:NSRb

[MP98c] L. H. A. Monteiro and J. R. C. Piqueira. Numbering self-replicating polymers II. Journal of Theoretical Biology, 194(1):159–161, September 7, 1998. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519398907585.

Monteiro:1999:MHS

[MP99] L. H. A. Monteiro and J. R. C. Piqueira. Modeling homopolymer self-replication: Implications for early competition. Journal of Theoretical Biology, 196(1):51–60, January 7, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519398908232.

Michalski:1997:MDE

[MPAA97] J. Michalski, J.-Ch. Poggiale, R. Arditi, and P. M. Auger. Macroscopic dynamic effects of migrations in patchy predator– prey systems. Journal of Theoretical Biology, 185(4):459–474, April 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396903276.

Mehr:1996:FRC

[MPFHG96] Ramit Mehr, Alan S. Perelson, Masha Fridkis-Hareli, and Amiela Globerson. Feedback regulation of T cell development in the thymus. Journal of Theoretical Biology, 181(2):157–167, July 21, 1996. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396901228.

Maree:1999:MTD

[MPH99] Athanasius F. M. Marée, Alexander V. Panfilov, and Paulien Hogeweg. Migration and thermotaxis of Dictyostelium discoideum slugs, a model study. Journal of Theoretical Biology, 199(3):297–309, August 7, 1999. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251939990958X.

Mestl:1995:MFD

[MPO95] Thomas Mestl, Erik Plahte, and Stig W. Omholt. A mathematical framework for describing and analysing gene regulatory networks. Journal of Theoretical Biology, 176(2): 291–300, September 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519385701993.

Marchuk:1991:MMAa

[MPRB91] G. I. Marchuk, R. V. Petrov, A. A. Romanyukha, and G. A. Bocharov. Mathematical model of antiviral immune response. I. Data analysis, generalized picture construction and parameters evaluation for hepatitis B. Journal of Theoretical Biology, 151(1):1–40, July 7, 1991. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305801420.

Marks:1994:QFA

[MPSK94] R. Marks, J. S. Percy, J. Semple, and S. Kumar. Quadriceps femoris activation changes in genu varum: a possible biomechanical factor in the pathogenesis of osteoarthrosis. Journal of Theoretical Biology, 170(3):283–289, October 7, 1994. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519384711891.

McLeod:1990:FSM

[MR90] Kenneth J. McLeod and Clinton T. Rubin. Frequency specific modulation of bone adaptation by induced electric fields. Journal of Theoretical Biology, 145(3):385–396, August 9, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305801171.

Mattheck:1991:CTA

[MR91a] C. Mattheck and S. Reuss. The claw of the tiger: An assessment of its mechanical shape optimization. Journal of Theoretical Biology, 150(3):323–328, June 7, 1991. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930580431X.

McKnight:1991:OGP

[MR91b]  Karl B. McKnight and Robin O. Roundy.  Optimal

gill packing in agaric sporocarps. Journal of Theoretical Biology, 150(4):497–528, June 21, 1991. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305804436.

Mackey:1992:MHI

[MR92] Michael A. Mackey and Joseph L. Roti Roti. A model of heat-induced clonogenic cell death. Journal of Theoretical Biology, 156(2):133–146, May 21, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305806691.

Mautino:1998:AMI

[MR98] Mario R. Mautino and Alberto L. Rosa. Analysis of models involving enzymatic activities for the occurrence of C → T transition mutations during repeat-induced point mutation (RIP) in Neurospora crassa. Journal of Theoretical Biology, 192(1): 61–71, May 7, 1998. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519397906081.

Marchuk:1991:MMAb

[MRB91] G. I. Marchuk, A. A. Romanyukha, and G. A. Bocharov. Mathematical model of antiviral immune response. II. parameters identification for acute viral hepatitis B. Journal of Theoretical Biology, 151(1):41–69, July 7, 1991. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305801432.

Mazat:1996:WMF

[MRL96] Jean-Pierre Mazat, Christine Reder, and Thierry Letellier. Why are most flux control coefficients so small? Journal of Theoretical Biology, 182(3):253–258, October 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396901629.

Mroz:1995:TCR

[Mró95] Iwona Mróz. Theoretical consideration of the relation between hydrogen bond types and possible secondary structures of a polypeptide chain. Journal of Theoretical Biology, 175(2):245– 255, July 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519385701373.

Malhotra:1990:RIS

[MS90a] Arum Malhotra and Ajit Sadana. Role of the initial state distribution on first-order deactivation of microheterogeneous enzyme samples. Journal of Theoretical Biology, 145(2):143–162, July 24, 1990. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305801213.

Morgan:1990:IRR

[MS90b] A. Richard Morgan and Alberto Severini. Interconversion of replication and recombination structures: Implications for terminal repeats and concatemers. Journal of Theoretical Biology, 144(2):195–202, May 14, 1990. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305803182.

Makela:1992:HGS

[MS92a] Annikki Mäkelä and Risto Sievänen. Height growth strategies in open-grown trees. Journal of Theoretical Biology, 159(4):443–467, December 21, 1992. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305806903.

Murray:1992:SSR

[MS92b] J. D. Murray and W. L. Seward. On the spatial spread of rabies among foxes with immunity. Journal of Theoretical Biology, 156(3):327–348, June 7, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305806794.

Michelson:1994:CKM

[MS94] Seth Michelson and Howard Schulman. CaM kinase: a model for its activation and dynamics. Journal of Theoretical Biology, 171(3):281–290, December 7, 1994. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384712318.

Maurer:1995:WMC

[MS95a] Roland Maurer and Valérie Séguinot. What is modelling for? a critical review of the models of path integration. Journal of Theoretical Biology, 175(4):457–475, August 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519385701543.

Motro:1995:NFS

[MS95b] Uzi Motro and Avi Shmida. Near-far search: An evolutionarily stable foraging strategy. Journal of Theoretical Biology, 173(1):15–22, March 7, 1995. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519385700380.

Manrubia:1997:FSD

[MS97a] Susanna C. Manrubia and Ricard V. Solé. On forest spatial dynamics with gap formation. Journal of Theoretical Biology, 187(2):159–164, July 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519397904094.

Miller:1997:AMC

[MS97b] Kyle E. Miller and David C. Samuels. The axon as a metabolic compartment: Protein degradation, transport, and maximum length of an axon. Journal of Theoretical Biology, 186(3):373– 379, June 7, 1997. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396903550.

Marchalonis:1998:SMR

[MS98] J. J. Marchalonis and S. F. Schluter. A stochastic model for the rapid emergence of specific vertebrate immunity incorporating horizontal transfer of systems enabling duplication and combinatorial diversification. Journal of Theoretical Biology, 193(3):429–444, August 7, 1998. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519398907111.

Madsen:1997:ATA

[MSHN97] Mark S. Madsen, David F. Snelling, Shaun Heaphy, and Vic Norris. Antiviruses as therapeutic agents: a mathematical analysis of their potential. Journal of Theoretical Biology, 184(2):111–116, January 21, 1997. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396902386.

McCollum:1996:OSI

[MSN96] Gin McCollum, Charlotte L. Shupert, and Lewis M. Nashner. Organizing sensory information for postural control in altered sensory environments. Journal of Theoretical Biology, 180(3):257–270, June 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519396901010.

Mjolsness:1991:CMD

[MSR91] Eric Mjolsness, David H. Sharp, and John Reinitz. A connectionist model of development. Journal of Theoretical Biology, 152(4):429–453, October 21, 1991. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305803911.

Mehr:1994:CTC

[MSSG94] Ramit Mehr, Lee Segel, Ayala Sharp, and Amiela Globerson. Colonization of the thymus by T cell progenitors: Models for cell–cell interactions. Journal of Theoretical Biology, 170(3):247–257, October 7, 1994. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519384711854.

Muraille:1996:TNR

[MTLK96] E. Muraille, D. Thieffry, O. Leo, and M. Kaufman. Toxicity and neuroendocrine regulation of the immune response: a model analysis. Journal of Theoretical Biology, 183(3):285–305, December 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396902210.

Miyake:1996:EDS

[MTM^{+}96] Yoshihiro Miyake, Sunao Tabata, Hirofumi Murakami, Masafumi Yano, and Hiroshi Shimizu. Environment-dependent self-organization of positional information field in chemotaxis of Physarum Plasmodium. Journal of Theoretical Biology, 178(4):341–353, February 21, 1996. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519396900302.

Mattheck:1997:MCR

[MTS97] C. Mattheck, M. Teschner, and J. Schäfer. Mechanical control of root growth: a computer simulation. Journal of Theoretical Biology, 184(3):261–269, February 7, 1997. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396902635.

Marcelpoil:1992:MSC

[MU92] R. Marcelpoil and Y. Usson. Methods for the study of cellular sociology: Voronoi diagrams and parametrization of the spatial relationships. Journal of Theoretical Biology, 154(3):359–369, February 7, 1992. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305801766.

Muller:1993:AFT

[Mul93a] M. Muller. The angles of femoral and tibial axes with respect to the cruciate ligament four-bar system in the knee joint. Journal of Theoretical Biology, 161(2):221–230, March 21, 1993. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519383710519.

Muller:1993:RBR

[Mul93b] M. Muller. The relationship between the rotation possibilities between femur and tibia and the lengths of the cruciate ligaments. Journal of Theoretical Biology, 161(2):199–220, March 21, 1993. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519383710507.

Muller:1994:SDD

[Mul94] M. Muller. Semicircular Duct dimensions and sensitivity of the vertebrate vestibular system. Journal of Theoretical Biology, 167(3):239–256, April 7, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519384710666.

Muller:1999:SLS

[Mul99] M. Muller. Size limitations in semicircular Duct systems. Journal of Theoretical Biology, 198(3):405–437, June 7, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519399909220.

Murase:1990:SCB

[Mur90] Masatoshi Murase. Simulation of ciliary beating by an excitable dynein model: Oscillations, quiescence and mechanosensitivity. Journal of Theoretical Biology, 146(2):209– 231, September 21, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519305801365.

Murase:1991:EDM

[Mur91] Masatoshi Murase. Excitable dynein model with multiple active sites for large-amplitude oscillations and bend propagation in flagella. Journal of Theoretical Biology, 149(2):181–202, March 21, 1991. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305802760.

Murase:1992:HDC

[Mur92] Masatoshi Murase. Hyperoscillations of dynein cross-bridge systems. Journal of Theoretical Biology, 154(1):27–33, January 7, 1992. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305801857.

Meijerhof:1993:MMT

[MvB93] R. Meijerhof and G. van Beek. Mathematical modelling of temperature and moisture loss of hatching eggs. Journal of Theoretical Biology, 165(1):27–41, November 7, 1993. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519383711756.

Morelli:1991:SRS

[MWAS91] R. A. Morelli, R. E. Walde, E. Akstin, and C. W. Schneider. L-system representation of speciation in the red algal genus Dipterosiphonia (Ceramiales, Rhodomelaceae). Journal of Theoretical Biology, 149(4):453–465, April 21, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305800931.

McNamara:1997:GTC

[MWC^{+}97] John M. McNamara, James N. Webb, E. J. Collins, Tamás Székely, and Alasdair I. Houston. A general technique for computing evolutionarily stable strategies based on errors in decision-making. Journal of Theoretical Biology, 189(2):211–225, November 21, 1997. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519397905117.

Myerscough:1993:SMT

[Mye93] M. R. Myerscough. A simple model for temperature regulation in honeybee swarms. Journal of Theoretical Biology, 162(3):381–393, June 7, 1993. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519383710945.

Mylius:1999:WPF

[Myl99] Sido D. Mylius. What pair formation can do to the battle of the sexes: Towards more realistic game dynamics. Journal of Theoretical Biology, 197(4):469–485, April 21, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251939890889X.

Norris:1994:CCC

[NAB^{+}94] V. Norris, J. A. Ayala, K. Begg, J.-P. Bouché, P. Bouloc, E. Boye, J. Canvin, S. Casaregola, A. J. Cozzone, E. Crooke, R. D’Ari, M. A. de Pedro, W. D. Donachie, R. J. Doyle, G. R. Drapeau, R. Fontana, S. Foster, J. A. Fralick, P. Freestone, R. C. Gayda, M. Goldberg, S. Grant, E. Guzman, J. H. Hageman, C. F. Higgins, M. Hofnung, I. B. Holland, J.-V. Holtje, P. Hughes, M. Inouye, S. Inouye, A. Jaffé, A. Jimenez-Sanchez, D. Joseleau-Petit, W. Keck, F. Kepes, A. Kornberg, P. Kuempel, H. Labischinski, A. Lobner-Olesen, J. Lutkenhaus, P. E. March, M. Matsuhashi, G. McGurk, W. Messer, J. Meury, Y. Milner, K. Modha, K. Nagai, T. Nagata, Y. Nishimura, S. Normark, E. Orr, A. Ottolenghi, L. Paolozzi, P. Poulsen, J. E. Rebollo, E. Z. Ron, J. Rouviere-Yaniv, K. Rudd, G. P. C. Salmond, G. Satta, U. Schwarz, S. Seror, A. J. Simon, B. G. Spratt, K. Sreekumar, S. Sweeney, I. Toth, R. Utsumi, D. Vinella, M. Wachi, B. M. Wilkins, P. H. Williams, and C. Yanofsky. Cell cycle control: Prokaryotic solutions to eukaryotic problems? Journal of Theoretical Biology, 168(2):227– 230, May 21, 1994. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519384711027.

Naber:1996:TAM

[Nab96] Holger Naber. Two alternative models for spontaneous flagellar motor switiching in halobacterium salinarium. Journal of Theoretical Biology, 181(4):343–358, August 21, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396901368.

Nakamura:1991:NTR

[Nak91] Kiyohiko Nakamura. Neural time-resolution depending on waveform of spikes. Journal of Theoretical Biology, 152(2):255–261, September 21, 1991. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305804564.

Nakatani:1995:ANS

[Nak95] Hiroshi Nakatani. Analysis of nonlinear substituent effect on lysozyme- and α-amylase-catalyzed reactions by parallel reaction mechanism. Journal of Theoretical Biology, 176(1):195–197, September 7, 1995. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519385701907.

Nakatani:1996:EVW

[Nak96] Hiroshi Nakatani. Effect of van der Waals volume of substituent on some glycosidase-catalyzed reactions. Journal of Theoretical Biology, 178(4):423–424, February 21, 1996. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396900405.

Nakajima:1998:EME

[Nak98] Toshiyuki Nakajima. Ecological mechanisms of evolution by natural selection: Causal processes generating densityand-frequency dependent fitness. Journal of Theoretical Biology, 190(4):313–331, February 21, 1998. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519397905543.

Nakajima:1999:BPC

[Nak99] Toshiyuki Nakajima. Biological probability: Cognitive processes of generating probabilities of events in biological systems. Journal of Theoretical Biology, 200(1):77–95, September 7, 1999. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519399909773.

Narang:1998:SSM

[Nar98] A. Narang. The steady states of microbial growth on mixtures of substitutable substrates in a chemostat. Journal of Theoretical Biology, 190(3):241–261, February 7, 1998. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251939790552X.

Nathon:1995:MSY

[NAZG95] Edna Nathon, Daniella Atzmony, Amotz Zahavi, and David Granot. Mate selection in yeast: a reconsideration of the signals and the message encoded by them. Journal of Theoretical Biology, 172(4):315–322, February 21, 1995. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519395900004.

Norman:1999:PTB

[NBBH99] R. Norman, R. G. Bowers, M. Begon, and P. J. Hudson. Persistence of tick-borne virus in the presence of multiple host species: Tick reservoirs and parasite mediated competition. Journal of Theoretical Biology, 200(1):111–118, September 7, 1999. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519399909827.

Nowak:1997:AVD

[NBSM97] Martin A. Nowak, Sebastian Bonhoeffer, George M. Shaw, and Robert M. May. Anti-viral drug treatment: Dynamics of resistance in free virus and infected cell populations. Journal of Theoretical Biology, 184(2):203–217, January 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519396903070.

Nyberg:1992:IES

[NC92] Anna M. Nyberg and Mikael B. Cronhjort. Intron evolution: a statistical comparison of two models. Journal of Theoretical Biology, 157(2):175–190, July 21, 1992. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305806198.

Niu:1997:EAC

[NC97] Deng K. Niu and Jia-Kuan Chen. Evolutionary advantages of cell specialization: Save and protect DNA. Journal of Theoretical Biology, 187(1):39–43, July 7, 1997. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519397904100.

Nicolis:1999:EPF

[ND99] S. C. Nicolis and J. L. Deneubourg. Emerging patterns and food recruitment in ants: an analytical study. Journal of Theoretical Biology, 198(4):575–592, June 21, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519399909347.

Nederbragt:1997:HOB

[Ned97] H. Nederbragt. Hierarchical organization of biological systems and the structure of adaptation in evolution and tumorigenesis. Journal of Theoretical Biology, 184(2):149–156, January 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396902660.

Neubert:1997:SPM

[Neu97] M. G. Neubert. A simple population model with qualitatively uncertain dynamics. Journal of Theoretical Biology, 189(4):399–411, December 21, 1997. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519397902587.

Neuhauser:1998:HDC

[Neu98] Claudia Neuhauser. Habitat destruction and competitive coexistence in spatially explicit models with local interactions. Journal of Theoretical Biology, 193(3):445–463, August 7, 1998. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519398907135.

Neveln:1990:CFS

[Nev90] Bob Neveln. Comma-free and synchronizable codes. Journal of Theoretical Biology, 144(2):209–212, May 14, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305803200.

Newman:1997:MME

[New97] M. E. J. Newman. A model of mass extinction. Journal of Theoretical Biology, 189(3):235–252, December 7, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519397905087.

Ng:1991:PAY

[Ng91] Yew-Kwang Ng. The paradox of the adventurous young and the cautious old: Natural selection vs. rational calculation. Journal of Theoretical Biology, 153(3):339–352, December 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305805740.

Ng:1996:CNF

[Ng96] Yew-Kwang Ng. Complex niches favour rational species. Journal of Theoretical Biology, 179(4):303–311, April 21, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396900697.

Naug:1999:FDL

[NG99] Dhruba Naug and Raghavendra Gadagkar. Flexible division of labor mediated by social interactions in an insect colony — a simulation model. Journal of Theoretical Biology, 197(1):123– 133, March 7, 1999. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519398908621.

Nilsen:1997:DNA

[NGP97] Thor W. Nilsen, Joseph Grayzel, and Wolf Prensky. Dendritic nucleic acid structures. Journal of Theoretical Biology, 187(2):273–284, July 21, 1997. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251939790446X.

Nemoto:1995:MVT

[NH95] Naoto Nemoto and Yuzuru Husimi. A model of the virustype strategy in the early stage of encoded molecular evolution. Journal of Theoretical Biology, 176(1):67–77, September 7, 1995. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519385701774.

Nauta:1996:VIA

[NH96] Maarten J. Nauta and Rolf F. Hoekstra. Vegetative incompatibility in ascomycetes: Highly polymorphic but selectively neutral? Journal of Theoretical Biology, 183(1):67–76, November 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396902027.

Nygren:1999:GAM

[NH99] A. Nygren and J. A. Halter. A general approach to modeling conduction and concentration dynamics in excitable cells of concentric cylindrical geometry. Journal of Theoretical Biology, 199(3):329–358, August 7, 1999. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519399909621.

Nishimura:1992:FUE

[Nis92] Kinya Nishimura. Foraging in an uncertain environment: Patch exploitation. Journal of Theoretical Biology, 156(1):91– 111, May 7, 1992. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305806587.

Nishimura:1999:EOG

[Nis99] Kinya Nishimura. Exploration of optimal giving-up time in uncertain environment: a sit-and-wait forager. Journal of Theoretical Biology, 199(3):321–327, August 7, 1999. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251939990961X.

Niwa:1994:SOD

[Niw94] Hiro-Sato Niwa. Self-organizing dynamic model of fish schooling. Journal of Theoretical Biology, 171(2):123–136, November 21, 1994. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519384712185.

Niwa:1996:NDA

[Niw96] Hiro-Sato Niwa. Newtonian dynamical approach to fish schooling. Journal of Theoretical Biology, 181(1):47–63, July 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519396901149.

Niwa:1998:MDF

[Niw98a] Hiro-Sato Niwa. Migration dynamics of fish schools in heterothermal environments. Journal of Theoretical Biology, 193(2):215–231, July 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519398906750.

Niwa:1998:SSS

[Niw98b] Hiro-Sato Niwa. School size statistics of fish. Journal of Theoretical Biology, 195(3):351–361, December 7, 1998. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519398908013.

Nunogaki:1990:HSH

[NK90] Kazuki Nunogaki and Michiki Kasai. How to separate the H^{+} and OH-permeabilities: Kinetics of intravesicular pH change after medium pH jump. Journal of Theoretical Biology, 144(3):415–427, June 7, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305800840.

Narang:1997:DMG

[NKR97] Atul Narang, Allan Konopka, and Doraiswami Ramkrishna. The dynamics of microbial growth on mixtures of substrates in batch reactors. Journal of Theoretical Biology, 184(3):301–317, February 7, 1997. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396902751.

Nunes:1991:COFa

[NLS91] M. Vaz Nunes, R. D. Lewis, and D. S. Saunders. A coupled oscillator feedback system as a model for the photoperiodic clock in insects and mites. I. The basic control system as a model for circadian rhythms. Journal of Theoretical Biology, 152(3):287–298, October 7, 1991. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930580195X.

Nowak:1992:CCH

[NM92a] Martin A. Nowak and Robert M. May. Coexistence and competition in HIV infections. Journal of Theoretical Biology, 159(3):329–342, December 7, 1992. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305807283.

Nunes:1992:LMM

[NM92b] Marlies Vaz Nunes and J. E. Moorhouse. Locust marching: a mathematical model. Journal of Theoretical Biology, 154(2):137–161, January 21, 1992. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251930580400X.

Nee:1993:PLC

[NM93a] S. Nee and R. M. May. Population-level consequences of conspecific brood parasitism in birds and insects. Journal of Theoretical Biology, 161(1):95–109, March 7, 1993. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519383710428.

Nuyts:1993:IIL

[NM93b] Erik Nuyts and Nico K. Michiels. Integration of immediate and long term sperm precedence patterns and mating costs in an optimization model of insect copulation duration. Journal of Theoretical Biology, 160(3):271–295, February 7, 1993. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519383710192.

Nuno:1993:IEF

[NMdlR93] J. C. Nuño, F. Montero, and F. J. de la Rubia. Influence of external fluctuations on a hypercycle formed by two kinetically indistinguishable species. Journal of Theoretical Biology, 165(4):553–575, December 21, 1993. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519383712075.

Nakamaru:1997:ECL

[NMI97] M. Nakamaru, H. Matsuda, and Y. Iwasa. The evolution of cooperation in a lattice-structured population. Journal of Theoretical Biology, 184(1):65–81, January 7, 1997. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251939690243X.

Nowak:1995:IRA

[NMS95] Martin A. Nowak, Robert M. May, and Karl Sigmund. Immune responses against multiple epitopes. Journal of Theoretical Biology, 175(3):325–353, August 7, 1995. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385701464.

Nakamaru:1998:SDF

[NNI98] Mayuko Nakamaru, Hajime Nogami, and Yoh Iwasa. Scoredependent fertility model for the evolution of cooperation in a lattice. Journal of Theoretical Biology, 194(1):101–124, September 7, 1998. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519398907500.

Ney-Nifle:1999:SAC

[NNM99] Muriel Ney-Nifle and Marc Mangel. Species-area curves based on geographic range and occupancy. Journal of Theoretical Biology, 196(3):327–342, February 7, 1999. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251939890844X.

Nolting:1995:RBA

[Nöl95] Bengt Nölting. Relation between adiabatic and pseudoadiabatic compressibility in ultrasonic velocimetry. Journal of Theoretical Biology, 175(2):191–196, July 21, 1995. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385701312.

Nolting:1998:SRF

[Nöl98] Bengt Nölting. Structural resolution of the folding pathway of a protein by correlation of Φ-values with inter-residue contacts. Journal of Theoretical Biology, 194(3):419–428, October 7, 1998. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519398907834.

Nolting:1999:AFP

[Nöl99] Bengt Nölting. Analysis of the folding pathway of chymotrypsin inhibitor by correlation of Φ-values with interresidue contacts. Journal of Theoretical Biology, 197(1):113– 121, March 7, 1999. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519398908608.

Nordborg:1992:FIH

[Nor92a] Magnus Nordborg. Female infanticide and human sex ratio evolution. Journal of Theoretical Biology, 158(2):195– 198, September 21, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519305807179.

Norris:1992:PDD

[Nor92b] Vic Norris. Phospholipid domains determine the spatial organization of the Escherichia coli cell cycle: the membrane tectonics model. Journal of Theoretical Biology, 154(1):91–107, January 7, 1992. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305801900.

Norris:1993:SOC

[Nor93] Vic Norris. Sequestration of origins of chromosome replication in Escherichia coli by lipid compartments: The pocket hypothesis. Journal of Theoretical Biology, 164(2):239–244, September 21, 1993. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519383711513.

Novak:1997:EPC

[Nov97] J. P. Novak. Electric potential and concentration of ion species in the proximity of a cell membrane: ab initio calculations. Journal of Theoretical Biology, 185(2):173–187, March 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396903008.

Nowak:1990:ESS

[Now90] Martin Nowak. An evolutionarily stable strategy may be inaccessible. Journal of Theoretical Biology, 142(2):237–241, January 23, 1990. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305802243.

Nowak:1991:EVC

[Now91] Martin Nowak. The evolution of viruses. competition between horizontal and vertical transmission of mobile genes. Journal of Theoretical Biology, 150(3):339–347, June 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305804333.

Nowak:1992:VHI

[Now92] Martin A. Nowak. Variability of HIV infections. Journal of Theoretical Biology, 155(1):1–20, March 7, 1992. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305805454.

Nowak:1999:ELG

[NPK99] Martin A. Nowak, Joshua B. Plotkin, and David C. Krakauer. The evolutionary language game. Journal of Theoretical Biology, 200(2):147–162, September 21, 1999. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519399909815.

Neutel:1994:GST

[NRD94] A. M. Neutel, J. B. T. M. Roerdink, and P. C. De Ruiter. Global stability of two-level detritus decomposer food chains. Journal of Theoretical Biology, 171(3):351–353, December 7, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519384712380.

Nussinov:1991:TVF

[NS91] Ruth Nussinov and Gary Smythers. Trends in the 5’ vs. 3’ flanks of oligonucleotides in eukaryotic and prokaryotic genomes: The asymmetric roles played by cytosine and guanine. Journal of Theoretical Biology, 153(1):111–135, November 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305803558.

Nowak:1994:APD

[NS94] Martin A. Nowak and Karl Sigmund. The alternating Prisoner’s dilemma. Journal of Theoretical Biology, 168(2):219– 226, May 21, 1994. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519384711015.

Nilsson:1995:MCR

[NS95] Nils-Otto Nilsson and Torbjörn Säll. A model of chiasma reduction of closely formed crossovers. Journal of Theoretical Biology, 173(1):93–98, March 7, 1995. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251938570046X.

Ni:1996:MAR

[NS96] Ta-Chen Ni and Michael A. Savageau. Model assessment and refinement using strategies from biochemical systems theory: Application to metabolism in human red blood cells. Journal of Theoretical Biology, 179(4):329–368, April 21, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519396900727.

Nishimura:1997:IPD

[NS97] K. Nishimura and D. W. Stephens. Iterated Prisoner’s dilemma: Pay-off variance. Journal of Theoretical Biology, 188(1):1–10, September 7, 1997. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519397904392.

Nowak:1998:DIR

[NS98] Martin A. Nowak and Karl Sigmund. The dynamics of indirect reciprocity. Journal of Theoretical Biology, 194(4):561–574, October 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519398907755.

Noldeke:1999:HCH

[NS99] Georg Nöldeke and Larry Samuelson. How costly is the honest signaling of need? Journal of Theoretical Biology, 197(4):527– 539, April 21, 1999. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519398908931.

Nielsen:1997:CG

[NSH97] Keld Nielsen, Preben Graae Sørensen, and Finn Hynne. Chaos in glycolysis. Journal of Theoretical Biology, 186(3):303–306, June 7, 1997. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396903665.

Nunes:1991:COFb

[NSL91] M. Vaz Nunes, D. S. Saunders, and R. D. Lewis. A coupled oscillator feedback system as a model for the photoperiodic clock in insects and mites. II. simulations of photoperiodic responses. Journal of Theoretical Biology, 152(3):299–317, October 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305801961.

Nieselt-Struwe:1997:EGC

[NSW97] Kay Nieselt-Struwe and Peter R. Wills. The emergence of genetic coding in physical systems. Journal of Theoretical Biology, 187(1):1–14, July 7, 1997. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519397904045.

Novak:1993:MCD

[NT93] Bela Novak and John J. Tyson. Modeling the cell division cycle: M-phase trigger, oscillations, and size control. Journal of Theoretical Biology, 165(1):101–134, November 7, 1993. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519383711793.

Novak:1995:QAM

[NT95] Bela Novak and John J. Tyson. Quantitative analysis of a molecular model of mitotic control in fission yeast. Journal of Theoretical Biology, 173(3):283–305, April 7, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251938570063X.

Nishihira:1997:TSI

[NT97] Jun Nishihira and Hiroto Tachikawa. Theoretical study on the interaction between dopamine and its receptor by ab initio molecular orbital calculation. Journal of Theoretical Biology, 185(2):157–163, March 21, 1997. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396902623.

Nishihira:1999:TEM

[NT99] Jun Nishihira and Hiroto Tachikawa. Theoretical evaluation of a model of the catalytic triads of serine and cysteine proteases by ab initio molecular orbital calculation. Journal of Theoretical Biology, 196(4):513–519, February 21, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519398908517.

Novak:1999:FCC

[NTCN^{+}99] Béla Novák, Attila Tóth, Attila Csikász-Nagy, Béla Györffy, John J. Tyson, and Kim Nasmyth. Finishing the cell cycle. Journal of Theoretical Biology, 199(2):223–233, July 21, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519399909566.

Nakagaki:1996:PSO

[NU96] Toshiyuki Nakagaki and Tetsuo Ueda. Phase switching of oscillatory contraction in relation to the regulation of amoeboid behavior by the plasmodium of Physarum polycephalum. Journal of Theoretical Biology, 179(3):261–267, April 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251939690065X.

Nunes:1990:ETP

[Nun90] Marlies Vaz Nunes. The effect of temperature on photoperiodic induction of diapause in insects and mites: a model for the photoperiodic “counter”. Journal of Theoretical Biology, 146(3):369–378, October 7, 1990. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305807477.

Nunes:1998:DCO

[Nun98] Marlies Vaz Nunes. A double circadian oscillator model for quantitative photoperiodic time measurement in insects and mites. Journal of Theoretical Biology, 194(2):299–311, September 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519398907676.

Nussinov:1991:OND

[Nus91] Ruth Nussinov. The ordering of nucleotides in the DNA: Strong pyrimidine-purine patterns near homooligomer tracts. Journal of Theoretical Biology, 149(1):21–42, March 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305800694.

Nussinov:1992:ECT

[Nus92] Ruth Nussinov. The eukaryotic CCAAT and TATA boxes, DNA spacer flexibility and looping. Journal of Theoretical Biology, 155(2):243–270, March 21, 1992. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305805971.

Nuyts:1994:TAW

[Nuy94] Erik Nuyts. Testing for the asymmetric war of attrition when only roles and fight duration are known. Journal of Theoretical Biology, 169(1):1–13, July 7, 1994. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384711258.

Niklas:1999:NPS

[NVB99] Karl J. Niklas, Jans Varna, and Lars A. Berglund. Nonparametric statistical formulas for factors of safety of plant stems. Journal of Theoretical Biology, 197(2):135–147, March 21, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519398908633.

Nakagaki:1999:RDA

[NYI99] Toshiyuki Nakagaki, Hiroyasu Yamada, and Masami Ito. Reaction-diffusion-advection model for pattern formation of rhythmic contraction in a giant amoeboid cell of the Physarum plasmodium. Journal of Theoretical Biology, 197(4):497–506, April 21, 1999. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519398908906.

Nakaya:1996:CRS

[NYY96] Akihiro Nakaya, Akinori Yonezawa, and Kenji Yamamoto. Classification of RNA secondary structures using the techniques of cluster analysis. Journal of Theoretical Biology, 183(1):105–117, November 7, 1996. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519396902064.

Nicholson:1990:LMS

[NZPB90] D. Nicholson, S. J. Zara, N. G. Parsonage, and J. Barber. Lattice model simulations of the stacking/unstacking behaviour in thylakoid membranes. Journal of Theoretical Biology, 145(1): 65–81, July 9, 1990. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305805351.

Ortega:1998:OMC

[OA98] Fernando Ortega and Luis Acerenza. Optimal metabolic control design. Journal of Theoretical Biology, 191(4):439–449, April 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251939790607X.

Oliver:1993:EPD

[OBGGGRR93] J. L. Oliver, P. Bernaola-Galván, J. Guerrero-García, and R. Román-Roldán. Entropic profiles of DNA sequences through chaos-game-derived images. Journal of Theoretical Biology, 160(4):457–470, February 21, 1993. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519383710301.

Ortigues:1997:CVM

[ODA97] I. Ortigues, C.-G. Dussap, and Y. Anglaret. Comparison of various methods of calculating the instantaneous respiratory gaseous exchanges from discrete measurements in respiration chambers. Journal of Theoretical Biology, 185(4):489–501, April 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396903380.

Ogurtsov:1993:CCI

[OES93] A. Yu. Ogurtsov, Yu. E. Elkin, and S. A. Shabalina. Calculation of contributions of individual monomeric units to biopolymer function. Journal of Theoretical Biology, 164(3):395–401, October 7, 1993. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519383711628.

Oexle:1998:TLD

[Oex98] Konrad Oexle. Telomere length distribution and southern blot analysis. Journal of Theoretical Biology, 190(4):369–377, February 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519397905592.

Olsson:1991:AMM

[OHS91] Per G. Olsson, Lennart Hammarström, and C. I. Edvard Smith. Antigenicity of mouse monoclonal antibodies. a study on the variable region of the heavy chain. Journal of Theoretical Biology, 151(1):111–122, July 7, 1991. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305801468.

Otten:1995:MSI

[OHS95] E. Otten, M. Hulliger, and K. A. Scheepstra. A model study on the influence of a slowly activating potassium conductance on repetitive firing patterns of muscle spindle primary endings. Journal of Theoretical Biology, 173(1):67–78, March 7, 1995. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519385700446.

Ochoa:1999:SMS

[OJ99] Gabriela Ochoa and Klaus Jaffé. On sex, mate selection and the red Queen. Journal of Theoretical Biology, 199(1):1–9, July 7, 1999. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519399909311.

Ouzounis:1994:RIH

[OK94] Christos A. Ouzounis and Nikos C. Kyrpides. Reverse interpretation: a hypothetical selection mechanism for adaptive mutagenesis based on autoregulated mRNA stability. Journal of Theoretical Biology, 167(4):373–380, April 21, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384710769.

Omholt:1998:DAS

[OKAP98] Stig W. Omholt, Xiang Kefang, Øivind Andersen, and Erik Plahte. Description and analysis of switchlike regulatory networks exemplified by a model of cellular iron homeostasis. Journal of Theoretical Biology, 195(3):339–350, December 7, 1998. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519398908001.

ONeill:1991:GMM

[O’N91] Michael C. O’Neill. A general method for modeling cell populations undergoing G_{1} → G_{0} transitions during development. Journal of Theoretical Biology, 153(4):499–518, December 21, 1991. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305801535.

Otsuka:1998:SRS

[ON98] Jinya Otsuka and Yuji Nozawa. Self-reproducing system can behave as Maxwell’s demon: Theoretical illustration under prebiotic conditions. Journal of Theoretical Biology, 194(2):205–221, September 21, 1998. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251939890755X.

Ong:1996:IPH

[ONRH96] Chi-Keong Ong, Sean Nee, Andrew Rambaut, and Paul H. Harvey. Inferring the population history of an epidemic from a phylogenetic tree. Journal of Theoretical Biology, 182(2):173–178, September 21, 1996. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519396901526.

Onstad:1992:TSS

[Ons92] David W. Onstad. Temporal and spatial scales in epidemiological concepts. Journal of Theoretical Biology, 158(4):495–515, October 21, 1992. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251930580712X.

Otsuka:1997:TSN

[ONT97] Jinya Otsuka, Takao Nakano, and Goro Terai. A theoretical study on the nucleotide changes under a definite functional constraint of forming stable base-pairs in the stem regions of ribosomal RNAs; its application to the phylogeny of eukaryotes. Journal of Theoretical Biology, 184(2):171–186, January 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396902775.

Omholt:1994:FC

[OP94] S. W. Omholt and E. Plahte. Feedback confusions. Journal of Theoretical Biology, 168(2):231, May 21, 1994. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384711039.

Oprea:1996:EMP

[OP96] Mihaela Oprea and Alan S. Perelson. Exploring the mechanisms of primary antibody responses to T cell-dependent antigens. Journal of Theoretical Biology, 181(3):215–235, August 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396901277.

Oliveira-Pinto:1993:SCM

[OPRN93] F. Oliveira-Pinto, N. Ricci, and L. Nieri. Simulation of creeping movements of ciliated protozoa. Journal of Theoretical Biology, 163(3):365–372, August 7, 1993. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519383711252.

Otani:1997:MME

[OR97] Niels F. Otani and Robert F.Gilmour, Jr. Memory models for the electrical properties of local cardiac systems. Journal of Theoretical Biology, 187(3):409–436, August 7, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519397904471.

Ortega:1994:SDA

[ORG94] Guillermo J. Ortega, Lilia Romanelli, and Diego A. Golombek. Statistical and dynamical analysis of circadian rhythms. Journal of Theoretical Biology, 169(1):15–21, July 7, 1994. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251938471126X.

OToole:1999:SOC

[ORM99] D. V. O’Toole, P. A. Robinson, and M. R. Myerscough. Self-organized criticality in termite architecture: a role for crowding in ensuring ordered nest expansion. Journal of Theoretical Biology, 198(3):305–327, June 7, 1999. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519399909177.

Orpwood:1990:MAL

[Orp90] Roger D. Orpwood. Mechanisms of association learning in the post-synaptic neurone. Journal of Theoretical Biology, 143(2):145–162, March 22, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305802656.

Orpwood:1994:PNM

[Orp94] R. D. Orpwood. A possible neural mechanism underlying consciousness based on the pattern processing capabilities of pyramidal neurons in the cerebral cortex. Journal of Theoretical Biology, 169(4):403–418, August 21, 1994. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384711623.

Owen:1997:PFS

[OS97] Markus R. Owen and Jonathan A. Sherratt. Pattern formation and spatiotemporal irregularity in a model for macrophage-tumour interactions. Journal of Theoretical Biology, 189(1):63–80, November 7, 1997. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251939790494X.

Osborn:1993:MST

[Osb93] Jeffrey W. Osborn. A model simulating tooth morphogenesis without morphogens. Journal of Theoretical Biology, 165(3):429–445, December 7, 1993. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519383711999.

Olsen:1995:MMA

[OSM95] Luke Olsen, Jonathan A. Sherratt, and Philip K. Maini. A mechanochemical model for adult dermal wound contraction and the permanence of the contracted tissue displacement profile. Journal of Theoretical Biology, 177(2):113–128, November 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519385702305.

Ostashevsky:1996:CLK

[Ost96] Joseph Y. Ostashevsky. Centromeric locations in karyotypes: a rule derived from the theory of brahnched polymers. Journal of Theoretical Biology, 181(3):293–298, August 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396901319.

Orengo:1990:RMP

[OT90] C. A. Orengo and W. R. Taylor. A rapid method of protein structure alignment. Journal of Theoretical Biology, 147(4):517–551, December 21, 1990. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305802632.

Ovadi:1991:PSMa

[Ová91a] Judit Ovádi. Physiological significance of metabolic channelling. Journal of Theoretical Biology, 152(1):1–22, September 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305805004.

Ovadi:1991:PSMb

[Ová91b] Judit Ovádi. Physiological significance of metabolite channelling: Author’s response to commentaries. Journal of Theoretical Biology, 152(1):135–141, September 7, 1991. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305805284.

Orosz:1996:TID

[OVH^{+}96] Ferenc Orosz, Beata G. Vértessy, Susan Hollán, Margit Horányi, and Judit Ovádi. Triosephosphate isomerase deficiency: Predictions and facts. Journal of Theoretical Biology, 182(3):437–447, October 7, 1996. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519396901848.

Otsuka:1996:ETR

[OWM96] Jinya Otsuka, Hidemi Watanabe, and Kentaro T. Mori. Evolution of transcriptional regulation system through promiscuous coupling of regulatory proteins with operons; suggestion from proteins sequence similarities in Escherichia coli. Journal of Theoretical Biology, 178(2):183–204, January 21, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519396900168.

Packer:1990:SCO

[PA90] Craig Packer and Peter Abrams. Should co-operative groups be more vigilant than selfish groups? Journal of Theoretical Biology, 142(3):341–357, February 9, 1990. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305805570.

Pokhilko:1998:CAB

[PA98] A. V. Pokhilko and F. I. Ataullakhanov. Contact activation of blood coagulation: Trigger properties and hysteresis hypothesis: Kinetic recognition of foreign surfaces upon contact activation of blood coagulation: a hypothesis. Journal of Theoretical Biology, 191(2):213–219, March 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519397905841.

Pirlet:1999:MLH

[PAG99] Karl Pirlet and Avril Arthur-Goettig. Maintaining life and health by natural selection of protein molecules. Journal of Theoretical Biology, 201(1):75–85, November 7, 1999. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519399910159.

Pamilo:1991:ESC

[Pam91] Pekka Pamilo. Evolution of the sterile caste. Journal of Theoretical Biology, 149(1):75–95, March 7, 1991. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305800736.

Parsons:1994:MSE

[Par94] P. A. Parsons. Morphological stasis: an energetic and ecological perspective incorporating stress. Journal of Theoretical Biology, 171(4):409–414, December 21, 1994. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384712446.

Paradis:1997:MCS

[Par97] Emmanuel Paradis. Metapopulations and chaos: On the stabilizing influence of dispersal. Journal of Theoretical Biology, 186(2):261–266, May 21, 1997. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519396903641.

Pearson:1991:EEG

[PAS91] S. D. Pearson, R. A. Ackerman, and R. C. Seagrave. The energetics of embryonic growth and development. I. Oxygen consumption, biomass growth, and heat production. Journal of Theoretical Biology, 152(2):233–240, September 21, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305804540.

Patnaik:1995:UMS

[Pat95] Pratap R. Patnaik. Uniqueness and multiplicity of steady states in monocyclic enzyme cascades: a graph-theoretic analysis. Journal of Theoretical Biology, 177(1):67–72, November 7, 1995. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519385702251.

Pauplin:1993:MEG

[Pau93] Yves Pauplin. Minimum evolution and goodness-of-fit. Journal of Theoretical Biology, 165(2):161–170, November 21, 1993. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519383711811.

Paulin:1995:ECS

[Pau95] Michael G. Paulin. Electroreception and the compass sense of sharks. Journal of Theoretical Biology, 174(3):325–339, June 7, 1995. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519385701026.

Payne:1997:RPC

[Pay97] Robert J. H. Payne. On the robustness of populations in changing environments. Journal of Theoretical Biology, 185(4):553– 555, April 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396903343.

Pocklington:1997:CEU

[PB97a] Richard Pocklington and Michael L. Best. Cultural evolution and units of selection in replicating text. Journal of Theoretical Biology, 188(1):79–87, September 7, 1997. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519397904604.

Prugel-Bennett:1997:MEP

[PB97b] Adam Prügel-Bennett. Modelling evolving populations. Journal of Theoretical Biology, 185(1):81–95, March 7, 1997. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396902957.

Parry:1992:MCI

[PBJ^{+}92] S. Parry, M. E. J. Barratt, S. Jones, S. McKee, and J. D. Murray. Modelling coccidial infection in chickens: Emphasis on vaccination by in-feed delivery of oocysts. Journal of Theoretical Biology, 157(4):407–425, August 21, 1992. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305806617.

Pascual:1997:EHB

[PC97] Mercedes Pascual and Hal Caswell. Environmental heterogeneity and biological pattern in a chaotic predator–prey system. Journal of Theoretical Biology, 185(1):1–13, March 7, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519396902726.

Possingham:1995:FFN

[PCN95] Hugh P. Possingham, Hugh N. Comins, and Ian R. Noble. The fire and flammability niches in plant communities. Journal of Theoretical Biology, 174(1):97–108, May 7, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519385700823.

Pollock:1992:RER

[PD92] Gregory Pollock and Lee Alan Dugatkin. Reciprocity and the emergence of reputation. Journal of Theoretical Biology, 159(1):25–37, November 7, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305807659.

Phillips:1993:AAF

[PD93] Ronald J. Phillips and Stephanie R. Dungan. Asymptotic analysis of flow in sieve tubes with semi-permeable walls. Journal of Theoretical Biology, 162(4):465–485, June 21, 1993. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519383711008.

Phipps:1997:HNC

[PDP97] Michel Phipps, Janusz Darozewski, and Jenny Phipps. How the neighborhood coherence principle (NCP) can give rise to tissue homeostasis: a cellular automaton approach. Journal of Theoretical Biology, 185(4):475–487, April 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396903021.

Peacocke:1996:PKA

[Pea96] Arthur R. Peacocke. A pioneer of the kinetic approach. Journal of Theoretical Biology, 182(3):219–222, October 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519396901587.

Peck:1990:EOE

[Pec90] Joel R. Peck. The evolution of outsider exclusion. Journal of Theoretical Biology, 142(4):565–571, February 22, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305801092.

Peck:1992:GSI

[Pec92] Joel R. Peck. Group selection, individual selection, and the evolution of genetic drift. Journal of Theoretical Biology, 159(2):163–187, November 21, 1992. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305807003.

Peck:1993:FEC

[Pec93] Joel R. Peck. Friendship and the evolution of co-operation. Journal of Theoretical Biology, 162(2):195–228, May 21, 1993. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519383710830.

Penotti:1991:HPM

[Pen91] Fabio E. Penotti. Human pre-mRNA splicing signals. Journal of Theoretical Biology, 150(3):385–420, June 7, 1991. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305804369.

Pennycuick:1998:CSF

[Pen98] C. J. Pennycuick. Computer simulation of fat and muscle burn in long-distance bird migration. Journal of Theoretical Biology, 191(1):47–61, March 7, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519397905725.

Pettersson:1991:NCE

[Pet91] Gösta Pettersson. No convincing evidence is available for metabolite channelling between enzymes forming dynamic complexes. Journal of Theoretical Biology, 152(1):65–69, September 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305805120.

Pettersson:1996:EAE

[Pet96a] Gösta Pettersson. Errors associated with experimental determinations of enzyme flux control coefficients. Journal of Theoretical Biology, 179(3):191–197, April 7, 1996. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396900600.

Pettersson:1996:NAD

[Pet96b] Gösta Pettersson. A new approach for determination of the selectively favoured kinetic design of enzyme reactions. Journal of Theoretical Biology, 183(2):179–183, November 21, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519396902118.

Pettersson:1999:EDC

[Pet99] G. Pettersson. Effect of dynamic channelling on the transientstate kinetics of coupled enzyme reactions. Journal of Theoretical Biology, 198(1):135–141, May 7, 1999. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519399909049.

Ponnuswamy:1994:CSF

[PG94a] P. K. Ponnuswamy and M. Michael Gromiha. On the conformational stability of folded proteins. Journal of Theoretical Biology, 166(1):63–74, January 7, 1994. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384710058.

Ponnuswamy:1994:CSO

[PG94b] P. K. Ponnuswamy and M. Michael Gromiha. On the conformational stability of oligonucleotide duplexes and tRNA molecules. Journal of Theoretical Biology, 169(4):419–432, August 21, 1994. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519384711635.

Peper:1999:PRS

[PG99] Abraham Peper and Cees A. Grimbergen. Preliminary results of simulations with an improved mathematical model of drug tolerance. Journal of Theoretical Biology, 199(1):119– 123, July 7, 1999. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519399909402.

Pierre:1997:SEI

[PGBYP97] Darren M. Pierre, Daniel Goldman, Yaneer Bar-Yam, and Alan S. Perelson. Somatic evolution in the immune system: The need for germinal centers for efficient affinity maturation. Journal of Theoretical Biology, 186(2):159–171, May 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519396903501.

Poppi:1994:ITI

[PGF94] D. P. Poppi, M. Gill, and J. France. Integration of theories of intake regulation in growing ruminants. Journal of Theoretical Biology, 167(2):129–145, March 21, 1994. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384710587.

Perez-Gil:1994:SSB

[PGK94] J. Pérez-Gil and K. M. W. Keogh. Structural similarities between myelin and hydrophobic surfactant associated proteins: Protein motifs for interacting with bilayers. Journal of Theoretical Biology, 169(3):221–229, August 7, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251938471143X.

Purvis:1994:TCE

[PGL94] Andy Purvis, John L. Gittleman, and Hang-Kwang Luh. Truth or consequences: Effects of phylogenetic accuracy on two comparative methods. Journal of Theoretical Biology, 167(3):293– 300, April 7, 1994. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251938471071X.

Possingham:1990:OPU

[PH90] Hugh P. Possingham and Alasdair I. Houston. Optimal patch use by a territorial forager. Journal of Theoretical Biology, 145(3):343–353, August 9, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305801146.

Panfilov:1993:CSR

[PH93] A. V. Panfilov and A. V. Holden. Computer simulation of re-entry sources in myocardium in two and three dimensions. Journal of Theoretical Biology, 161(3):271–285, April 7, 1993. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519383710556.

Pagie:1999:CDR

[PH99] Ludo Pagie and Paulien Hogeweg. Colicin diversity: a result of eco-evolutionary dynamics. Journal of Theoretical Biology, 196(2):251–261, January 21, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519398908384.

Phillips:1996:MN

[Phi96] John B. Phillips. Magnetic navigation. Journal of Theoretical Biology, 180(4):309–319, June 21, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519396901058.

Philippe:1999:SIS

[Phi99]  P. Philippe.  The scale-invariant spatial clustering of

leukemia in San Francisco. Journal of Theoretical Biology, 199(4):371–381, August 21, 1999. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519399909645.

Paik:1992:EMS

[PK92] Woon Ki Paik and Sangduk Kim. Effect of methyl substitution on protein tertiary structure. Journal of Theoretical Biology, 155(3):335–342, April 7, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305806022.

Panfilov:1993:EHF

[PK93] Alexander V. Panfilov and James P. Keener. Effects of high frequency stimulation on cardiac tissue with an inexcitable obstacle. Journal of Theoretical Biology, 163(4):439–448, August 21, 1993. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251938371129X.

Pumir:1997:TBM

[PK97] Alain Pumir and Valentin I. Krinsky. Two biophysical mechanisms of defibrillation of cardiac tissue. Journal of Theoretical Biology, 185(2):189–199, March 21, 1997. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396902969.

Payne:1998:PDP

[PK98] Robert J. H. Payne and David C. Krakauer. The paradoxical dynamics of prion disease latency. Journal of Theoretical Biology, 191(4):345–352, April 21, 1998. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519397906275.

Pumir:1999:URW

[PK99] Alain Pumir and Valentin Krinsky. Unpinning of a rotating wave in cardiac muscle by an electric field. Journal of Theoretical Biology, 199(3):311–319, August 7, 1999. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519399909578.

Palsson:1993:MCS

[PL93] Bernhard O. Palsson and I-Der Lee. Model complexity has a significant effect on the numerical value and interpretation of metabolic sensitivity coefficients. Journal of Theoretical Biology, 161(3):299–315, April 7, 1993. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251938371057X.

Powell:1996:LPG

[PLB96] James A. Powell, Jesse A. Logan, and Barbara J. Bentz. Local projections for a global model of mountain pine beetle attacks. Journal of Theoretical Biology, 179(3):243–260, April 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396900648.

Pluhar:1994:MBW

[Plu94] Wolfgang Pluhar. The molecular basis of wobbling: An alternative hypothesis. Journal of Theoretical Biology, 169 (3):305–312, August 7, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519384711519.

Payne:1996:ETC

[PM96] Robert J. H. Payne and Mark. Escalation and time costs in displays of endurance. Journal of Theoretical Biology, 183(2):185–193, November 21, 1996. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251939690212X.

Pal:1999:EIG

[PM99] Csaba Pál and István Miklós. Epigenetic inheritance, genetic assimilation and speciation. Journal of Theoretical Biology, 200(1):19–37, September 7, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519399909748.

Pilyugin:1997:MCP

[PMA97] Sergei Pilyugin, John Mittler, and Rustom Antia. Modeling T-cell proliferation: an investigation of the consequences of the Hayflick limit. Journal of Theoretical Biology, 186(1):117– 129, May 7, 1997. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396903197.

Piqueira:1999:ZLO

[PMdM^{+}99] J. R. C. Piqueira, L. H. A. Monteiro, T. M. C. de Magalhães, R. T. Ramos, R. B. Sassi, and E. G. Cruz. Zipf’s law organizes a psychiatric Ward. Journal of Theoretical Biology, 198(3):439– 443, June 7, 1999. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519399909232.

Pollatschek:1990:MMO

[PN90] M. A. Pollatschek and A. M. Nahir. A mathematical model of osteoarthosis. Journal of Theoretical Biology, 143(4):497–505, April 19, 1990. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305800268.

Pospisil:1998:TST

[PN98] Pavel Pospísil and Jan Naus. Theoretical simulation of temperature induced increase of quantum yield of minimum chlorophyll fluorescence Φ_{F (0)}. Journal of Theoretical Biology, 193(1):125–130, July 7, 1998. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519398906919.

Payne:1992:ACM

[PNB92] Robert J. H. Payne, Martin A. Nowak, and Baruch S. Blumberg. Analysis of a cellular model to account for the natural history of infection by the hepatitis B virus and its role in the development of primary hepatocellular carcinoma. Journal of Theoretical Biology, 159(2):215–240, November 21, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305807039.

Paul:1991:TFR

[PO91] R. Paul and M. Otwinowski. The theory of the frequency response of ellipsoidal biological cells in rotating electrical fields. Journal of Theoretical Biology, 148(4):495–519, February 21, 1991. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305802334.

Polezhaev:1992:MMM

[Pol92a] A. A. Polezhaev. A mathematical model of the mechanism of vertebrate somitic segmentation. Journal of Theoretical Biology, 156(2):169–181, May 21, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251930580671X.

Pollanen:1992:DOD

[Pol92b] M. S. Pollanen. Dimensional optimization at different levels of the arterial hierarchy. Journal of Theoretical Biology, 159(2):267–270, November 21, 1992. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305807064.

Pollard:1996:PPS

[Pol96] Irina Pollard. Preconceptual programming and sexual orientation: a hypothesis. Journal of Theoretical Biology, 179(3):269– 273, April 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396900661.

Porteous:1996:DOH

[Por96] John W. Porteous. Dominance — one hundred and fifteen years after Mendel’s paper. Journal of Theoretical Biology, 182(3):223–232, October 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519396901599.

Posch:1999:WSL

[Pos99] Martin Posch. Win-stay, lose-shift strategies for repeated games-memory length, aspiration levels and noise. Journal of Theoretical Biology, 198(2):183–195, May 21, 1999. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519399909098.

Pelce:1992:GAM

[PP92] Pierre Pelce and Alain Pocheau. Geometrical approach to the morphogenesis of unicellular algae. Journal of Theoretical Biology, 156(2):197–214, May 21, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305806733.

Payne:1998:OPO

[PRG98] Robert J. H. Payne and Miguel A. Rodríguez-Gironés. The origins of parent-offspring signalling. Journal of Theoretical Biology, 195(2):273–279, November 21, 1998. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519398907986.

Price:1995:NS

[Pri95] George R. Price. The nature of selection. Journal of Theoretical Biology, 175(3):389–396, August 7, 1995. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251938570149X.

Promislow:1994:DRE

[Pro94] Daniel E. L. Promislow. DNA repair and the evolution of longevity: a critical analysis. Journal of Theoretical Biology, 170(3):291–300, October 7, 1994. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384711908.

Pelce:1993:GMG

[PS93] Pierre Pelce and Jiong Sun. Geometrical models for the growth of unicellular algae. Journal of Theoretical Biology, 160(3):375–386, February 7, 1993. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519383710234.

Peleg:1992:FGC

[PSE92] B. Peleg, A. Shmida, and S. Ellner. Foraging graphs: Constraint rules on matching between bees and flowers in a two-sided pollination market. Journal of Theoretical Biology, 157(2):191–201, July 21, 1992. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305806204.

Peuvot:1999:FPI

[PSLB99] Jacques Peuvot, André Schanck, Laurence Lins, and Robert Brasseur. Are the fusion processes involved in birth, life and death of the cell depending on tilted insertion of peptides into membranes? Journal of Theoretical Biology, 198(2):173–181, May 21, 1999. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519399909086.

Pinou:1996:ADT

[PSLD96] Theodora Pinou, Tamar Schlick, Bin Li, and H. G. Dowling. Addition of Darwin’s third dimension to phyletic trees. Journal of Theoretical Biology, 182(4):505–512, October 21, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519396901915.

Plonsey:1991:QIB

[PST91] R. Plonsey, C. F. Starmer, and I. Tipans. The quantitative investigation of the binding process of calcium blocking drugs in a sinoatrial node. Journal of Theoretical Biology, 149(1):55–61, March 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305800712.

Pohlmann:1992:MCO

[PT92a] L. Pohlmann and H. Tributsch. A model for co-operative electron transfer. Journal of Theoretical Biology, 156(1):63–72, May 7, 1992. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305806563.

Pohlmann:1992:SCO

[PT92b] L. Pohlmann and H. Tributsch. Stimulated and co-operative electron transfer in energy conversion and catalysis. Journal of Theoretical Biology, 155(4):443–462, April 21, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305806290.

Powell:1998:TAS

[PTBL98] James Powell, Jessica Tams, Barbara Bentz, and Jesse Logan. Theoretical analysis of “switching” in a localized model for mountain pine beetle mass attack. Journal of Theoretical Biology, 194(1):49–63, September 7, 1998. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519398907421.

Poulin:1995:CSE

[PV95] Robert Poulin and William L. Vickery. Cleaning symbiosis as an evolutionary game: To cheat or not to cheat? Journal of Theoretical Biology, 175(1):63–70, July 7, 1995. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251938570121X.

Prokudina:1991:NMA

[PVT91] Elena I. Prokudina, Rustem Yu. Valeev, and Rustem N. Tchuraev. A new method for the analysis of the dynamics of the molecular genetic control systems. II. application of the method of generalized threshold models in the investigation of concrete genetic systems. Journal of Theoretical Biology, 151(1):89– 110, July 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305801456.

Pacala:1991:ECA

[PW91] S. W. Pacala and J. Weiner. Effects of competitive asymmetry on a local density model of plant interference. Journal of Theoretical Biology, 149(2):165–179, March 21, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305802759.

Qualls:1997:EVP

[QAM97] Carl P. Qualls, Robin M. Andrews, and Tom Mathies. The evolution of viviparity and placentation revisited. Journal of Theoretical Biology, 185(1):129–135, March 7, 1997. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396902945.

Quon:1991:MMCa

[QC91a] Michael J. Quon and L. Arthur Campfield. A mathematical model and computer simulation study of insulin receptor regulation. Journal of Theoretical Biology, 150(1):59–72, May 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305804758.

Quon:1991:MMCb

[QC91b] Michael J. Quon and L. Arthur Campfield. A mathematical model and computer simulation study of insulin sensitive glucose transporter regulation. Journal of Theoretical Biology, 150(1):93–107, May 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305804771.

Quentin:1994:FIR

[QF94] Yves Quentin and Gwennaele A. Fichant. Fast identification of repetitive elements in biological sequences. Journal of Theoretical Biology, 166(1):51–61, January 7, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384710046.

Quenette:1992:DFD

[QG92] Pierre-Yves Quenette and Jean-François Gerard. Does frequency-dependent selection optimize fitness? Journal of Theoretical Biology, 159(4):381–385, December 21, 1992. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305806848.

Quinn:1991:MPS

[Qui91] Thomas P. Quinn. Models of Pacific salmon orientation and navigation on the open ocean. Journal of Theoretical Biology, 150(4):539–545, June 21, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251930580445X.

Qi:1993:CAM

[QZDA93] An-Shen Qi, Xiang Zheng, Chan-Ying Du, and Bao-Sheng An. A cellular automaton model of cancerous growth. Journal of Theoretical Biology, 161(1):1–12, March 7, 1993. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519383710350.

Rhodes:1996:PDL

[RA96] C. J. Rhodes and R. M. Anderson. Persistence and dynamics in lattice models of epidemic spread. Journal of Theoretical Biology, 180(2):125–133, May 21, 1996. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396900880.

Radding:1991:PPS

[Rad91] Wilson Radding. Proposed partial β-structures for lac permease and the Na + H + antiporter which use similar transport and H + coupling mechanisms. Journal of Theoretical Biology, 150(2):239–249, May 21, 1991. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305803352.

Rabinovitch:1999:MPA

[RAGO99] A. Rabinovitch, I. Aviram, N. Gulko, and E. Ovsyshcher. A model for the propagation of action potentials in nonuniformly excitable media. Journal of Theoretical Biology, 196(2):141–154, January 21, 1999. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519398908244.

Rakitzis:1997:KAC

[Rak97] Emmanuel T. Rakitzis. Kinetic analysis of chemical or enzymic reactions: an algorithm for the determination of the initial velocity of product formation by the use of a Taylor series in reaction time. Journal of Theoretical Biology, 188(3):387–389, October 7, 1997. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519397905129.

Rakocevic:1998:WNR

[Rak98] Miloje M. Rakocević. Whole-number relations between protein amino acids and their biosynthetic precursors. Journal of Theoretical Biology, 191(4):463–465, April 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519397906068.

Ratnieks:1990:AQM

[Rat90] Francis L. W. Ratnieks. Assessment of queen mating frequency by workers in social hymenoptera. Journal of Theoretical Biology, 142(1):87–93, January 9, 1990. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305800141.

Ratnieks:1991:EDA

[Rat91] Francis L. W. Ratnieks. Evolution of discriminatory aggression in marine invertebrates. Journal of Theoretical Biology, 152(4):557–565, October 21, 1991. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305803972.

Raudino:1994:MCK

[Rau94] Antonio Raudino. A model of coagulation kinetics between two deformable bilayers. application to the membrane fusion. Journal of Theoretical Biology, 167(4):307–321, April 21, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519384710733.

Raudino:1991:PME

[RB91] Antonio Raudino and Patrizia Bianciardi. Polymer-mediated electrostatic interactions between charged lipid assemblies and electrolyte solutions: a tentative model of the polyethylene glycol-induced cell fusion. Journal of Theoretical Biology, 149(1):1–20, March 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519305800682.

Ratnieks:1997:RSS

[RB97a] Francis L. W. Ratnieks and Jacobus J. Boomsma. On the robustness of split sex ratio predictions in social hymenoptera. Journal of Theoretical Biology, 185(4):423–439, April 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396903318.

Rees:1997:ALC

[RB97b] Mark Rees and Joy Bergelson. Asymmetric light competition and founder control in plant communities. Journal of Theoretical Biology, 184(3):353–358, February 7, 1997. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396902970.

Root-Bernstein:1996:SHV

[RBB96] Robert S. Root-Bernstein and Maurine Bernstein. SIV as a human vaccine against HIV? Journal of Theoretical Biology, 180(1):91–92, May 7, 1996. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519396900831.

Reade:1998:MDV

[RBBG98] Brian Reade, Roger G. Bowers, Michael Begon, and Rosalind Gaskell. A model of disease and vaccination for infections with acute and chronic phases. Journal of Theoretical Biology, 190(4):355–367, February 21, 1998. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519397905579.

Ramanan:1998:NII

[RBC98] S. V. Ramanan, P. R. Brink, and G. J. Christ. Neuronal innervation, intracellular signal transduction and intercellular coupling: a model for syncytial tissue responses in the steady state. Journal of Theoretical Biology, 193(1):69–84, July 7, 1998. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519398906877.

Root-Bernstein:1997:MCC

[RBD97] Robert S. Root-Bernstein and Patrick F. Dillon. Molecular complementarity i: the complementarity theory of the origin and evolution of life. Journal of Theoretical Biology, 188(4):447–479, October 21, 1997. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519397904768.

Root-Bernstein:1993:DHP

[RBH93] Robert S. Root-Bernstein and Sheila H. Hobbs. Does HIV “piggyback” on CD4-like surface proteins of sperm, viruses, and bacteria? implications for co-transmission, cellular tropism and the induction of autoimmunity in AIDS. Journal of Theoretical Biology, 160(2):249–264, January 21, 1993. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519383710179.

Root-Bernstein:1998:APC

[RBH98] Robert S. Root-Bernstein and Daniel D. Holsworth. Antisense peptides: a critical mini-review. Journal of Theoretical Biology, 190(2):107–119, January 21, 1998. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519397905440.

Root-Bernstein:1997:NCP

[RBM97] Robert S. Root-Bernstein and Stephen J. Merrill. The necessity of cofactors in the pathogenesis of AIDS: a mathematical model. Journal of Theoretical Biology, 187(1):135–146, July 7, 1997. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519397904495.

Robson:1999:RBS

[RBP99] Arthur J. Robson, Carl T. Bergstrom, and Jonathan K. Pritchard. Risky business: Sexual and asexual reproduction in variable environments. Journal of Theoretical Biology, 197(4):541–556, April 21, 1999. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519398908943.

Rao:1990:MMI

[RBR90] Gita Subba Rao, J. S. Bajaj, and J. Subba Rao. A mathematical model for insulin kinetics II. extension of the model to include response to oral glucose administration and application to insulin-dependent diabetes mellitus (IDDM). Journal of Theoretical Biology, 142(4):473–483, February 22, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305801031.

Ragosta:1992:AME

[RCCM92] Maria Ragosta, Carmelina Cosmi, Vincenzo Cuomo, and Maria Macchiato. An application of maximum entropy techniques to determine homogeneous sets of nucleotidic sequences. Journal of Theoretical Biology, 155(1):129–136, March 7, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305805533.

Ramadevan:1990:GSM

[RD90] Sanjay Ramadevan and Michael A. B. Deakin. The gibbons speciation mechanism. Journal of Theoretical Biology, 145(4):447–456, August 23, 1990. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305804801.

Rundell:1998:HIR

[RDHD98] A. Rundell, R. DeCarlo, H. HogenEsch, and P. Doerschuk. The humoral immune response to haemophilus influenzae type b: a mathematical model based on t-zone and germinal center bcell dynamics. Journal of Theoretical Biology, 194(3):341–381, October 7, 1998. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519398907512.

Renaud:1991:SMH

[RdM91] F. Renaud and T. de Meeüs. A simple model of host-parasite evolutionary relationships. parasitism: Compromise or conflict? Journal of Theoretical Biology, 152(3):319–327, October 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305801973.

Regen:1993:CMS

[Reg93] David M. Regen. Characterization of myocardial stresslength relations. Journal of Theoretical Biology, 164(2): 245–259, September 21, 1993. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519383711525.

Reiff:1995:ROL

[Rei95] Theodore R. Reiff. The role of the ocular lens in the regulation of intraocular pressure. Journal of Theoretical Biology, 175(3):267–269, August 7, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519385701403.

Resnik:1995:DCP

[Res95] David Resnik. Developmental constraints and patterns: some pertinent distinctions. Journal of Theoretical Biology, 173 (3):231–240, April 7, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519385700598.

Repka:1995:ESS

[RG95] Joe Repka and Mart R. Gross. The evolutionarily stable strategy under individual condition and tactic frequency. Journal of Theoretical Biology, 176(1):27–31, September 7, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519385701725.

Ruxton:1997:MDD

[RGAP97] Graeme D. Ruxton, Jose L. Gonzalez-Andujar, and Joe N. Perry. Mortality during dispersal an the stability of a metapopulation. Journal of Theoretical Biology, 186(3):389–396, June 7, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519397904021.

Rosenberger:1991:MRL

[RGK91] R. F. Rosenberger, E. Gounaris, and E. Kolettas. Mechanisms responsible for the limited lifespan and immortal phenotypes in cultured mammalian cells. Journal of Theoretical Biology, 148(3):383–392, February 7, 1991. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305802437.

Ressayre:1998:MMA

[RGMG98] Adrienne Ressayre, Bernard Godelle, Agnès Mignot, and Pierre Henri Gouyon. A morphogenetic model accounting for pollen aperture pattern in flowering plants. Journal of Theoretical Biology, 193(2):321–334, July 21, 1998. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519398907044.

Rubin:1992:CTE

[RH92] Harvey N. Rubin and Mostafa N. Halim. Comparison of translation in eukaryotic mode with translation in prokaryote mode. Journal of Theoretical Biology, 159(3):377–379, December 7, 1992. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305807313.

Reed:1996:DAE

[RH96] Jeremy N. Reed and Laurence D. Hurst. Dynamic analysis of the evolution of a novel genetic system: the evolution of ciliate meiosis. Journal of Theoretical Biology, 178(4):355–368, February 21, 1996. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396900314.

Riek:1995:ECB

[RHS^{+}95] Peter R. Riek, Mark D. Handschumacher, Shen-Shu Sung, Ming Tan, and … Robert M. Graham. Evolutionary conservation of both the hydrophilic and hydrophobic nature of transmembrane residues. Journal of Theoretical Biology, 172(3):245–258, February 7, 1995. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519385700215.

Rice:1990:GME

[Ric90] Sean H. Rice. A geometric model for the evolution of development. Journal of Theoretical Biology, 143(3):319–342, April 5, 1990. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305800335.

Rice:1995:GTS

[Ric95] Sean H. Rice. A genetical theory of species selection. Journal of Theoretical Biology, 177(3):237–245, December 7, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251938570241X.

Ricci:1996:IPD

[Ric96] Juan C. Diaz Ricci. Influence of phosphoenolpyruvate on the dynamic behaviour of phosphofructokinase of Escherichia coli. Journal of Theoretical Biology, 178(2):145–150, January 21, 1996. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396900132.

Riggs:1994:CGI

[Rig94] Jack E. Riggs. Carcinogenesis, genetic instability and genomic entropy: Insight derived from malignant brain tumor age specific mortality rate dynamics. Journal of Theoretical Biology, 170(4):331–338, October 21, 1994. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384711957.

Rakocevic:1996:FST

[RJ96] Miloje Rakocević and Anja Jokić. Four stereochemical types of protein amino acids: Synchronic determination with chemical characteristics, atom and nucleon number. Journal of Theoretical Biology, 183(3):345–349, December 7, 1996. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251939690226X.

Ricard:1992:DAL

[RKM92] Jacques Ricard, Nicolas Kellershohn, and Guillermo Mulliert. Dynamics aspects of long distance functional interactions between membrane-bound enzymes. Journal of Theoretical Biology, 156(1):1–40, May 7, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251930580654X.

Richardson:1992:MM

[RL92] I. W. Richardson and A. H. Louie. Membranes and meters. Journal of Theoretical Biology, 154(1):9–26, January 7, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305801845.

Rajasekar:1994:BCS

[RL94] S. Rajasekar and M. Lakshmanan. Bifurcation, chaos and suppression of chaos in FitzHugh–Nagumo nerve conduction model equation. Journal of Theoretical Biology, 166(3):275–288, February 7, 1994. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519384710253.

Rosso:1993:UCB

[RLF93] L. Rosso, J. R. Lobry, and J. P. Flandrois. An unexpected correlation between cardinal temperatures of microbial growth highlighted by a new model. Journal of Theoretical Biology, 162(4):447–463, June 21, 1993. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519383710994.

Raghavan:1997:CMB

[RLRV97] R. Raghavan, Wayne Lawton, S. R. Ranjan, and R. R. Viswanathan. A continuum mechanics-based model for cortical growth. Journal of Theoretical Biology, 187(2):285–296, July 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519397904501.

Rose:1993:DFO

[RM93] Michael R. Rose and Chris Moore. A Darwinian function for the orbital cortex. Journal of Theoretical Biology, 161(1):119– 129, March 7, 1993. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519383710441.

Rohani:1995:IPC

[RM95] Pejman Rohani and Octavio Miramontes. Immigration and the persistence of chaos in population models. Journal of Theoretical Biology, 175(2):203–206, July 21, 1995. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385701336.

Ruxton:1997:IFD

[RM97] Graeme D. Ruxton and Anthony L. Moody. The ideal free distribution with kleptoparasitism. Journal of Theoretical Biology, 186(4):449–458, June 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519397904057.

Rohani:1996:MES

[RMH96] Pejman Rohani, Robert M. May, and Michael P. Hassell. Metapopulation and equilibrium stability: The effects of spatial structure. Journal of Theoretical Biology, 181(2):97–109, July 21, 1996. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396901186.

Read:1995:IBC

[RN95] Andrew F. Read and Sean Nee. Inference from binary comparative data. Journal of Theoretical Biology, 173(1):99–108, March 7, 1995. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519385700471.

Roberts:1996:MEE

[RN96] Bruce W. Roberts and M. E. J. Newman. A model for evolution and extinction. Journal of Theoretical Biology, 180(1): 39–54, May 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396900764.

Robinson:1990:PAC

[Rob90a] David Robinson. Phosphorus availability and cortical senescence in cereal roots. Journal of Theoretical Biology, 145 (2):257–265, July 24, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519305801304.

Robson:1990:EEG

[Rob90b] Arthur J. Robson. Efficiency in evolutionary games: Darwin, Nash and the secret handshake. Journal of Theoretical Biology, 144(3):379–396, June 7, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305800827.

Robertson:1991:FTD

[Rob91] Douglas S. Robertson. Feedback theory and Darwinian evolution. Journal of Theoretical Biology, 152(4):469–484, October 21, 1991. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305803935.

Rodriguez:1996:MEP

[Rod96] Diego J. Rodríguez. A model for the establishment of polyploidy in plants: Viable but infertile hybrids, iteroparity, and demographic stochasticity. Journal of Theoretical Biology, 180(3):189–196, June 7, 1996. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519396900958.

Rodriguez:1998:TDD

[Rod98] Diego J. Rodríguez. Time delays in density dependence are often not destabilizing. Journal of Theoretical Biology, 191(1):95–101, March 7, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519397905695.

Rognstad:1994:IEH

[Rog94] Robert Rognstad. Isotopic estimation of the hepatic glucose balance in vivo. Journal of Theoretical Biology, 168(2):161– 173, May 21, 1994. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519384710964.

Rognstad:1995:MLP

[Rog95] Robert Rognstad. Models of the liver pentose cycle. Journal of Theoretical Biology, 173(2):195–206, March 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385700550.

Roland:1994:ORT

[Rol94] P. E. Roland. Obstacles on the road towards a neuroscientific theory of mind. Journal of Theoretical Biology, 171(1): 19–28, July 1994. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519384712082.

Riccobene:1999:MAD

[ROL99] Todd A. Riccobene, Geneva M. Omann, and Jennifer J. Linderman. Modeling activation and desensitization of Gprotein coupled receptors provides insight into ligand efficacy. Journal of Theoretical Biology, 200(2):207–222, September 21, 1999. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519399909888.

Rodriguez:1990:GSM

[ROMM90] F. Rodríguez, J. L. Oliver, A. Marín, and J. R. Medina. The general stochastic model of nucleotide substitution. Journal of Theoretical Biology, 142(4):485–501, February 22, 1990. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305801043.

Rosen:1994:P

[Ros94a] Robert Rosen. On psychomimesis. Journal of Theoretical Biology, 171(1):87–92, July 1994. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519384712148.

Rossini:1994:QSU

[Ros94b] Gian Paolo Rossini. The quaternary structures of untransformed steroid hormone receptors: An open issue. Journal of Theoretical Biology, 166(3):339–353, February 7, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384710319.

Rose:1998:PND

[Ros98] Sam Rose. A proposal for a new direction to treat cancer. Journal of Theoretical Biology, 195(1):111–128, November 7, 1998. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519398907925.

Roterman:1995:MOS

[Rot95] I. Roterman. Modelling the optimal simulation path in the peptide chain folding-studies based on geometry of Alanine heptapeptide. Journal of Theoretical Biology, 177(3):283–288, December 7, 1995. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519385702457.

Roth:1998:PER

[Rot98] Bradley J. Roth. The pinwheel experiment revisited. Journal of Theoretical Biology, 190(4):389–393, February 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519397905658.

Roussel:1998:SRE

[Rou98] Marc R. Roussel. Slowly reverting enzyme inactivation: a mechanism for generating long-lived damped oscillations. Journal of Theoretical Biology, 195(2):233–244, November 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519398907883.

Ritchie:1996:SSD

[RP96] Raymond J. Ritchie and Tania Prvan. A simulation study on designing experiments to measure the K_{m} of Michaelis–Menten kinetics curves. Journal of Theoretical Biology, 178(3):239–254, February 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396900235.

Rispe:1998:CBC

[RP98] Claude Rispe and Jean-Sébastien Pierre. Coexistence between cyclical parthenogens, obligate parthenogens, and intermediates in a fluctuating environment. Journal of Theoretical Biology, 195(1):97–110, November 7, 1998. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519398907846.

Rakitzis:1999:KAP

[RP99] Emmanuel T. Rakitzis and Photini Papandreou. Kinetic analysis of parallel reactions, the initial reactant of which presents with two interconvertible isomeric forms (hydrolysis and hydroxylaminolysis of 6-phosphogluconolactone). Journal of Theoretical Biology, 200(4):427–434, October 21, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519399910068.

Ratnieks:1992:CSQ

[RR92] Francis L. W. Ratnieks and H. Kern Reeve. Conflict in single-queen hymenopteran societies: the structure of conflict and processes that reduce conflict in advanced eusocial species. Journal of Theoretical Biology, 158(1):33–65, September 7, 1992. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305806472.

Ruoff:1996:TCG

[RR96] Peter Ruoff and Ludger Rensing. The temperature-compensated Goodwin model simulates many circadian clock properties. Journal of Theoretical Biology, 179(4):275–285, April 21, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396900673.

Roberston:1998:DD

[RR98] David Roberston and Jennifer Robinson. Darwinian daisyworld. Journal of Theoretical Biology, 195(1):129–134, November 7, 1998. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519398907998.

Ryan:1990:QAS

[RRB90] George W. Ryan, J. L. Rouse, and L. A. Bursill. Quantitative analysis of sunflower seed packing. Journal of Theoretical Biology, 147(3):303–328, December 7, 1990. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305804904.

Ringel:1998:EDC

[RRG98] M. S. Ringel, M. Rees, and H. C. J. Godfray. The evolution of diapause in a coupled host-parasitoid system. Journal of Theoretical Biology, 194(2):195–204, September 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519398907548.

Robert:1997:CMI

[RS97] Mark E. Robert and James D. Sweeney. Computer model: Investigating role of filopodia-based steering in experimental neurite galvanotropism. Journal of Theoretical Biology, 188(3):277–288, October 7, 1997. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519397904719.

Rajagopal:1998:BIS

[RS98] K. Rajagopal and V. Sitaramam. Biology of the inner space: Voids in biopolymers. Journal of Theoretical Biology, 195(2):245–271, November 21, 1998. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251939890786X.

Russo:1992:DSY

[RSB92] Frank D. Russo, Isaac Scherson, and James R. Broach. Direct simulation of yeast 2-μ m circle plasmid amplification. Journal of Theoretical Biology, 155(3):369–385, April 7, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305806046.

Rohwer:1996:HRM

[RSW96] Johann M. Rohwer, Stefan Schuster, and Hans V. Westerhoff. How to recognize monofunctional units in a metabolic system. Journal of Theoretical Biology, 179(3):213–228, April 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519396900624.

Ruse:1997:OLP

[Rus97] Michael Ruse. The origin of life: Philosophical perspectives. Journal of Theoretical Biology, 187(4):473–482, August 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396903823.

Ruxton:1995:PMS

[Rux95] G. D. Ruxton. Population models with sexual reproduction show a reduced propensity to exhibit chaos. Journal of Theoretical Biology, 175(4):595–601, August 21, 1995. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385701683.

Ruxton:1996:SBI

[Rux96] Graeme D. Ruxton. Synchronisation between individuals and the dynamics of linked populations. Journal of Theoretical Biology, 183(1):47–54, November 7, 1996. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396902003.

Reinitz:1990:TEA

[RV90] John Reinitz and J. Rimas Vaisnys. Theoretical and experimental analysis of the phage lambda genetic switch implies missing levels of co-operativity. Journal of Theoretical Biology, 145(3):295–318, August 9, 1990. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305801110.

Rubelj:1999:SMC

[RV99] Ivica Rubelj and Zoran Vondracek. Stochastic mechanism of cellular aging-abrupt telomere shortening as a model for stochastic nature of cellular aging. Journal of Theoretical Biology, 197(4):425–438, April 21, 1999. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519398908864.

Baragar:1996:APD

[rvdBvdG96] F. A. Baragar (retired), A. van der Bilt, and H. W. van der Glas. An analytic probability density for particle size in human mastication. Journal of Theoretical Biology, 181(2):169–178, July 21, 1996. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251939690123X.

Ruoff:1999:GMS

[RVMR99] Peter Ruoff, Merete Vinsjevik, Saadat Mohsenzadeh, and Ludger Rensing. The Goodwin model: Simulating the effect of cycloheximide and heat shock on the sporulation rhythm of Neurospora crassa. Journal of Theoretical Biology, 196(4):483–494, February 21, 1999. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519398908463.

Reuter:1998:ETP

[RWB98] Max Reuter, Paul I. Ward, and Wolf U. Blanckenhorn. An ESS treatment of the pattern of female arrival at the mating site in the yellow dung fly scathophaga stercoraria (l.). Journal of Theoretical Biology, 195(3):363–370, December 7, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519398907950.

Rice:1998:MSR

[RWDM98] J. Jeremy Rice, Raimond L. Winslow, Julie Dekanski, and Elliot McVeigh. Model studies of the role of mechanosensitive currents in the generation of cardiac arrhythmias. Journal of Theoretical Biology, 190(4):295–312, February 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519397905385.

Recknagel:1993:SMF

[RWK93] Rolf-Dieter Recknagel, Klausdieter Weller, and Johann-Michael Köhler. A simple method to forecast good promoters of Escherichia coli by means of doublet frequencies. Journal of Theoretical Biology, 162(1):75–80, May 7, 1993. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519383710775.

Regoes:1998:VDE

[RWN98] Roland R. Regoes, Dominik Wodarz, and Martin A. Nowak. Virus dynamics: the effect of target cell limitation and immune responses on virus evolution. Journal of Theoretical Biology, 191(4):451–462, April 21, 1998. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519397906172.

Ryti:1990:TCD

[Ryt90] Randall T. Ryti. Theoretical considerations of densitydependent spacing of a mound-building ant (Formica altipetens). Journal of Theoretical Biology, 147(3):289–301, December 7, 1990. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305804898.

Rabinovitch:1999:BLP

[RZF^{+}99] Avinoam Rabinovitch, Arieh Zaritsky, Itzhak Fishov, Monica Einav, and Hilla Hadas. Bacterial lysis by phage — a theoretical model. Journal of Theoretical Biology, 201(3):209–213, December 7, 1999. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519399910299.

Sutherland:1993:PDI

[SA93] William J. Sutherland and Clive W. Anderson. Predicting the distribution of individuals and the consequences of habitat loss: The role of prey depletion. Journal of Theoretical Biology, 160(2):223–230, January 21, 1993. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519383710155.

Schank:1997:SOH

[SA97] Jeffrey C. Schank and Jeffrey R. Alberts. Self-organized huddles of rat pups modeled by simple rules of individual behavior. Journal of Theoretical Biology, 189(1):11–25, November 7, 1997. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519397904884.

Santillan:1997:TAC

[SAB97] M. Santillán and F. Angulo-Brown. A thermodynamic approach to the compromise between power and efficiency in muscle contraction. Journal of Theoretical Biology, 189(4):391–398, December 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519397905269.

Sachs:1994:VDB

[Sac94] T. Sachs. Variable development as a basis for robust pattern formation. Journal of Theoretical Biology, 170(4):423–425, October 21, 1994. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519384712033.

Sneyd:1993:MSP

[SAF^{+}93] J. Sneyd, A. Atri, M. W. J. Ferguson, M. A. Lewis, W. Seward, and J. D. Murray. A model for the spatial patterning of teeth primordia in the alligator: Initiation of the dental determinant. Journal of Theoretical Biology, 165(4):633–658, December 21, 1993. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519383712117.

Sun:1995:APD

[SAGB95] Jiong Sun, Farid Amellal, Leon Glass, and Jacques Billette. Alternans and period-doubling bifurcations in atrioventricular nodal conduc. Journal of Theoretical Biology, 173(1):79–91, March 7, 1995. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519385700458.

Szedlacsek:1996:TDC

[SAH96] Stefan E. Szedlacsek, Alexandru R. Aricescu, and Bent H. Havsteen. Time-dependent control of metabolic systems by external effectors. Journal of Theoretical Biology, 182(3): 341–350, October 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519396901733.

Saito:1995:AGP

[Sai95] Yutaka Saitô. “Altruistic gene” a priori involves two different deleterious gene effects: a reply to Keller. Journal of Theoretical Biology, 174(4):471–472, June 21, 1995. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305807337.

Sakai:1991:MAA

[Sak91] Satoki Sakai. A model analysis for the adaptive architecture of herbaceous plants. Journal of Theoretical Biology, 148(4):535–544, February 21, 1991. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305802358.

Sakai:1992:ALE

[Sak92] Satoki Sakai. Asynchronous leaf expansion and shedding in a seasonal environment: Result of a competitive game. Journal of Theoretical Biology, 154(1):77–90, January 7, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305801894.

Sakai:1994:OLP

[Sak94] Satoki Sakai. Optimal leaf phenology and photosynthetic capacity of herbs dependent on the genet density of herbs in their habitat. Journal of Theoretical Biology, 166(3):237–244, February 7, 1994. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519384710216.

Sakai:1995:ORA

[Sak95] Satoki Sakai. Optimal resource allocation to vegetative and sexual reproduction of a plant growing in a spatially varying environment. Journal of Theoretical Biology, 175(3):271–282, August 7, 1995. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519385701415.

Sakai:1996:OPE

[Sak96] Satoki Sakai. On ovule production in environments where pollinator or resource availability is unpredictable. Journal of Theoretical Biology, 183(3):317–327, December 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396902234.

Sakai:1997:AMF

[Sak97] S. Sakai. An analytical model for flower production in an environment where resource availability is unpredictable. Journal of Theoretical Biology, 189(4):449–452, December 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519397905336.

Salerno:1991:MIC

[Sal91] Costantino Salerno. Misleading interpretations of coupled enzyme kinetics. Journal of Theoretical Biology, 152(1):73–75, September 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305805144.

Salter:1996:DFC

[Sal96] Mark Salter. Determination of the flux control coefficient of nitric oxide synthase for nitric oxide synthesis in discrete brain regions in vivo. Journal of Theoretical Biology, 182(3):449–452, October 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251939690185X.

Stewart:1997:MPR

[SAM97] Paul D. Stewart, Carl Anderson, and David W. Macdonald. A mechanism for passive range exclusion: Evidence from the European badger (Meles meles). Journal of Theoretical Biology, 184(3):279–289, February 7, 1997. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396902489.

Santillan:1999:TOA

[San99] Moisés Santillán. A thermodynamic optimization analysis of a possible relation between the parameters that determine the energetics of muscle contraction in steady state. Journal of Theoretical Biology, 199(1):105–112, July 7, 1999. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519399909463.

Sandberg:1992:CMM

[SAS92] Sonja Sandberg, Tamara E. Awerbuch, and Andrew Spielman. A comprehensive multiple matrix model representing the life cycle of the tick that transmits agent of Lyme disease. Journal of Theoretical Biology, 157(2):203–220, July 21, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305806216.

Sasaki:1994:EAD

[Sas94] Akira Sasaki. Evolution of antigen drift/switching: Continuously evading pathogens. Journal of Theoretical Biology, 168(3):291–308, June 7, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519384711106.

Sasaki:1997:CDN

[Sas97] Akira Sasaki. Clumped distribution by neighbourhood competition. Journal of Theoretical Biology, 186(4):415–430, June 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519396903707.

Satulovsky:1996:LLV

[Sat96] Javier E. Satulovsky. Lattice Lotka–Volterra models and negative cross-diffusion. Journal of Theoretical Biology, 183 (4):381–389, December 21, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519396902295.

Saunders:1994:ENS

[Sau94] Peter T. Saunders. Evolution without natural selection: Further implications of the daisyworld parable. Journal of Theoretical Biology, 166(4):365–373, February 21, 1994. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384710332.

Savageau:1991:BST

[Sav91a] Michael A. Savageau. Biochemical systems theory: Operational differences among variant representations and their significance. Journal of Theoretical Biology, 151(4):509–530, August 21, 1991. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305803674.

Savageau:1991:MCI

[Sav91b] Michael A. Savageau. Metabolite channeling: Implications for regulation of metabolism and for quantitative description of reactions in vivo. Journal of Theoretical Biology, 152(1):85–92, September 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251930580517X.

Savageau:1992:DAM

[Sav92] Michael A. Savageau. Dominance according to metabolic control analysis: Major achievement or house of cards? Journal of Theoretical Biology, 154(1):131–136, January 7, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305801948.

Savageau:1995:MMM

[Sav95a] Michael A. Savageau. Michaelis–Menten mechanism reconsidered: implications of fractal kinetics. Journal of Theoretical Biology, 176(1):115–124, September 7, 1995. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385701816.

Savic:1995:MPF

[Sav95b] D. Savić. Model of pattern formation in animal coatings. Journal of Theoretical Biology, 172(4):299–303, February 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519385700264.

Saxton:1992:GEM

[Sax92] Michael J. Saxton. Gaps in the erythrocyte membrane skeleton: a stretched net model. Journal of Theoretical Biology, 155(4):517–536, April 21, 1992. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305806332.

Sayers:1994:CAI

[Say94] Jon R. Sayers. Computer aided identification of a potential 5’-3’ exonuclease gene encoded by Escherichia coli. Journal of Theoretical Biology, 170(4):415–421, October 21, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519384712021.

Smillie:1990:RSM

[SB90a] Frazer Smillie and William Bains. Repetition structure of mammalian nuclear DNA. Journal of Theoretical Biology, 142(4):463–471, February 22, 1990. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251930580102X.

Smith:1990:MNN

[SB90b] Daniel P. B. Smith and Theodore H. Bullock. Model nerve net can produce rectilinear, non-diffuse propagation as seen in the skin plexus of sea urchins. Journal of Theoretical Biology, 143(1):15–40, March 8, 1990. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305802863.

Seymour:1992:CTS

[SB92] R. M. Seymour and R. H. Bradbury. Is the crown-of-thorns starfish degrading the great barrier reef? Journal of Theoretical Biology, 159(1):111–133, November 7, 1992. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305807714.

Seymour:1994:SWS

[SB94] R. M. Seymour and R. H. Bradbury. Spatial waves or Synchronicity of outbreaks of crown-of-thorns starfish on the great barrier reef?: a Bayesian message from large scale data. Journal of Theoretical Biology, 166(4):453–460, February 21, 1994. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251938471040X.

Sole:1995:MCS

[SB95] Ricard V. Solé and Jordi Bascompte. Measuring chaos from spatial information. Journal of Theoretical Biology, 175(2): 139–147, July 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519385701269.

Singh:1997:MAU

[SB97] Prim B. Singh and David Brown. Modelling the activity of the ultrabithorax parasegment-specific regulatory domains around their anterior boundaries. Journal of Theoretical Biology, 186(4):397–413, June 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519396903689.

Sen:1992:CMR

[SBM92] Pali Sen, Denis Bell, and Donna Mohr. A calcium model with random absorption: a stochastic approach. Journal of Theoretical Biology, 154(4):485–493, February 21, 1992.

CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930580164X.

Solow:1999:ECS

[SBRG99] Daniel Solow, Apostolos Burnetas, Theresa Roeder, and Neil S. Greenspan. Evolutionary consequences of selected locus-specific variations in epistasis and fitness contribution in Kauffman’s NK model. Journal of Theoretical Biology, 196(2):181–196, January 21, 1999. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519398908323.

Sole:1992:SCS

[SBV92] Ricard V. Solé, Jordi Bascompte, and Joaquim Valls. Stability and complexity of spatially extended two-species competition. Journal of Theoretical Biology, 159(4):469–480, December 21, 1992. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305806915.

Subramaniam:1990:EID

[SC90] Bala Subramaniam and John S. Claudius. Effect of intravenous drug administration mode on drug distribution in a tumor slab: a finite Fourier transform analysis. Journal of Theoretical Biology, 143(1):55–76, March 8, 1990. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305802887.

Saunders:1992:EMA

[SC92a] K. B. Saunders and A. R. C. Cummin. Estimates of mean alveolar PCO 2 during steady-state exercise in man: a theoretical study. Journal of Theoretical Biology, 159(3):307–327, December 7, 1992. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305807271.

Seiden:1992:MSC

[SC92b] Philip E. Seiden and Franco Celada. A model for simulating cognate recognition and response in the immune system. Journal of Theoretical Biology, 158(3):329–357, October 7, 1992. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305807374.

Sibly:1993:AVL

[SC93] R. M. Sibly and R. N. Curnow. An allelocentric view of life-history evolution. Journal of Theoretical Biology, 160(4):533–546, February 21, 1993. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519383710349.

Schulte:1996:MDW

[SC96] Paul J. Schulte and David G. Costa. A mathematical description of water flow through plant tissues. Journal of Theoretical Biology, 180(1):61–70, May 7, 1996. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396900788.

Slot:1999:ERS

[SC99a] Liesbeth A. Bruins Slot and Francis C. Colpaert. Experimental realization of a signal transduction algorithm. Journal of Theoretical Biology, 200(1):39–48, September 7, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519399909736.

Stamenovic:1999:RPA

[SC99b] Dimitrije Stamenović and Mark F. Coughlin. The role of prestress and architecture of the cytoskeleton and deformability of cytoskeletal filaments in mechanics of adherent cells: a quantitative analysis. Journal of Theoretical Biology, 201(1):63–74, November 7, 1999. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519399910147.

Shumaker:1991:MPS

[SCG91] J. M. Shumaker, J. W. Clark, and W. R. Giles. A model of the phase-sensitivity of the pacemaking cell in the bullfrog heart. Journal of Theoretical Biology, 151(2):193–230, July 21, 1991. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305803613.

Schneider:1991:TMMa

[Sch91a] Thomas D. Schneider. Theory of molecular machines. I. Channel capacity of molecular machines. Journal of Theoretical Biology, 148(1):83–123, January 7, 1991. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305804667.

Schneider:1991:TMMb

[Sch91b] Thomas D. Schneider. Theory of molecular machines. II. energy dissipation from molecular machines. Journal of Theoretical Biology, 148(1):125–137, January 7, 1991. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305804679.

Schoutens:1994:PEP

[Sch94] Jacques E. Schoutens. Prediction of elastic properties of sperm flagella. Journal of Theoretical Biology, 171(2):163–177, November 21, 1994. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519384712215.

Schuster:1996:CAT

[Sch96] Stefan Schuster. Control analysis in terms of generalized variables characterizing metabolic systems. Journal of Theoretical Biology, 182(3):259–268, October 7, 1996. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396901630.

Schneider:1997:ICI

[Sch97a] T. D. Schneider. Information content of individual genetic sequences. Journal of Theoretical Biology, 189(4):427–441, December 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519397905403.

Schwartz:1997:SRP

[Sch97b] Alan W. Schwartz. Speculation on the RNA precursor problem. Journal of Theoretical Biology, 187(4):523–527, August 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519396903860.

Scheuring:1999:AEI

[Sch99a] István Scheuring. Allee effect increases the dynamical stability of populations. Journal of Theoretical Biology, 199(4):407–414, August 21, 1999. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519399909669.

Schneider:1999:MMI

[Sch99b] Thomas D. Schneider. Measuring molecular information. Journal of Theoretical Biology, 201(1):87–92, November 7, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519399910123.

seljkoDzakula:1991:BMPa

[sDA91] seljko Dzakula and Radoslav K. Andjus. Biophysical models of protein denaturation. I. An improvement of the model of two states. Journal of Theoretical Biology, 153(1):41–59, November 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305803522.

seljkoDzakula:1991:BMPb

[sDAB91] seljko Dzakula, Radoslav K. Andjus, and Milan Bozić. Biophysical models of protein denaturation. II. effects of denaturants and of pH. Journal of Theoretical Biology, 153(1):61–88, November 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305803534.

Skarka:1990:MMB

[SDB90] V. Skarka, J. L. Deneubourg, and M. R. Belić. Mathematical model of building behavior of Apis mellifera. Journal of Theoretical Biology, 147(1):1–16, November 7, 1990.

CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305802498.

Seaborg:1999:EFN

[Sea99] David M. Seaborg. Evolutionary feedback: a new mechanism for stasis and punctuated evolutionary change based on integration of the organism. Journal of Theoretical Biology, 198(1): 1–26, May 7, 1999. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519398908967.

Senger:1993:IDG

[SEB^{+}93] B. Senger, R. Ezzeddine, F. J. Bafaluy, P. Schaaf, F. J. G. Cuisinier, and J.-C. Voegel. Influence of diffusion and gravity on the adhesion of a two-component mixture of hard spheres on a flat surface. Journal of Theoretical Biology, 163(4):457–471, August 21, 1993. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519383711318.

Saunderson:1999:DMB

[SEB99] S. E. T. Saunderson, A. H. England, and C. J. Baker. A dynamic model of the behaviour of Sitka spruce in high winds. Journal of Theoretical Biology, 200(3):249–259, October 7, 1999. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519399909839.

Segel:1994:ELU

[Seg94] Irwin H. Segel. The effects of labeled and unlabeled impurities on the analysis of equilibrium binding and initial velocity data by means of scatchard plots. Journal of Theoretical Biology, 171(3):267–280, December 7, 1994. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384712306.

Sekimura:1995:DSM

[Sek95] Toshio Sekimura. The diversity in shoot morphology of herbaceous plants in relation to solar radiation captured by leaves. Journal of Theoretical Biology, 177(3):289–298, December 7, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519385702469.

Semenov:1996:NCM

[Sem96] Mikhail V. Semënov. New concept of microtubule dynamics and microtubule motor movement and new model of chromosome movement in mitosis. Journal of Theoretical Biology, 179(2):91–117, March 14, 1996. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396900521.

Sen:1996:SPM

[Sen96] Asok K. Sen. On the sign pattern of metabolic control coefficients. Journal of Theoretical Biology, 182(3):269–275, October 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396901642.

Sereno:1991:FAB

[Ser91] Martin I. Sereno. Four analogies between biological and cultural/linguistic evolution. Journal of Theoretical Biology, 151(4):467–507, August 21, 1991. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305803662.

Seyfarth:1991:RAL

[Sey91] Robert M. Seyfarth. Reciprocal altruism and the limits of correlational analysis. Journal of Theoretical Biology, 153(1):141–144, November 7, 1991. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305803571.

Siddons:1990:CNI

[SF90] R. C. Siddons and J. France. On correcting for the nonideal behaviour of the fluid-phase marker when determining digesta flow by continuous marker infusion. Journal of Theoretical Biology, 142(1):137–138, January 9, 1990. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305800189.

Sole:1999:RQD

[SFGG^{+}99] Ricard V. Solé, Ramon Ferrer, Isabel González-García, Josep Quer, and Esteban Domingo. Red Queen dynamics, competition and critical points in a model of RNA virus quasispecies. Journal of Theoretical Biology, 198(1):47–59, May 7, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519399909013.

Smith:1997:DSS

[SFHP97] Derek J. Smith, Stephanie Forrest, Ron R. Hightower, and Alan S. Perelson. Deriving shape space parameters from immunological data. Journal of Theoretical Biology, 189(2):141–150, November 21, 1997. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519397904951.

Saffre:1999:CDM

[SFKD99] F. Saffre, R. Furey, B. Krafft, and J. L. Deneubourg. Collective decision-making in social spiders: Dragline-mediated amplification process acts as a recruitment mechanism. Journal of Theoretical Biology, 198(4):507–517, June 21, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251939990927X.

Stamenovic:1996:MAC

[SFW^{+}96] Dimitrije Stamenović, Jeffrey J. Fredberg, Ning Wang, James P. Butler, and Donald E. Ingber. A microstructural approach to cytoskeletal mechanics based on tensegrity. Journal of Theoretical Biology, 181(2):125–136, July 21, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396901204.

Schurr:1990:AMM

[SG90] Tamar Schurr and N. B. Grover. Analysis of a model for minichromosome segregation in Escherichia coli. Journal of Theoretical Biology, 146(3):395–406, October 7, 1990. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305807490.

Shakhnovich:1991:IPM

[SG91] E. I. Shakhnovich and A. M. Gutin. Influence of point mutations on protein structure: Probability of a neutral mutation. Journal of Theoretical Biology, 149(4):537–546, April 21, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305800979.

Stadler:1993:AFL

[SG93] Peter F. Stadler and Walter Grüner. Anisotropy in fitness landscapes. Journal of Theoretical Biology, 165(3):373–388, December 7, 1993. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519383711951.

Schulze:1995:BMU

[SG95] H. Georg Schulze and Boris B. Gorzalka. Behavioral modeling: The use of chemical reaction kinetics to investigate lordosis behavior in female rats. Journal of Theoretical Biology, 174(4):355–369, June 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251938570104X.

Sole:1998:CDE

[SG98] Ricard V. Solé and Javier G. P. Gamarra. Chaos, dispersal and extinction in coupled ecosystems. Journal of Theoretical Biology, 193(3):539–541, August 7, 1998. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519398907160.

Sandblom:1995:FRI

[SGG95] John Sandblom, Paul George, and Juris Galvanovskis. The frequency response of ion channel currents in a combined diffusion and barrier type model. Journal of Theoretical Biology, 176(1):153–160, September 7, 1995. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385701841.

Sharan:1998:PAR

[SGP98] Maithili Sharan, Suman Gupta, and Aleksander S. Popel. Parametric analysis of the relationship between end-capillary and mean tissue PO 2 as predicted by a mathematical model. Journal of Theoretical Biology, 195(4):439–449, December 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519398908050.

Stirbet:1998:CFI

[SGSS98] Alexandrina Stirbet, Govindjee, Bruno J. Strasser, and Reto J. Strasser. Chlorophyll a fluorescence induction in higher plants: Modelling and numerical simulation. Journal of Theoretical Biology, 193(1):131–151, July 7, 1998. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519398906920.

Stone:1990:SBM

[SH90a] M. Stone and D. Harrison. A stratified binomial marker model for bone-marrow repopulation experiments. Journal of Theoretical Biology, 144(2):267–273, May 14, 1990. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930580325X.

Sugrue:1990:CBF

[SH90b] P. Sugrue and M. E. J. Holwill. Changes in beat frequency of reactivated flagella of Crithidia oncopelti inhibited with vanadate. Journal of Theoretical Biology, 143(3):421–423, April 5, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305800372.

Sekimura:1991:MLV

[SH91a] Toshio Sekimura and Hirokazu Hotani. The morphogenesis of liposomes viewed from the aspect of bending energy. Journal of Theoretical Biology, 149(3):325–337, April 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305803091.

Shiau:1991:DHB

[SH91b] L. J. Shiau and B. Hassard. Degenerate Hopf bifurcation and isolated periodic solutions of the Hodgkin–Huxley model with varying sodium ion concentration. Journal of Theoretical Biology, 148(2):157–173, January 21, 1991. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930580339X.

Sakai:1993:HPO

[SH93a] Satoki Sakai and Yasushi Harada. Habitat productivity and optimal growth schedule of annual plants in seasonal environments. Journal of Theoretical Biology, 163(3):347–363, August 7, 1993. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519383711240.

Sjerps:1993:IDO

[SH93b] Marjan Sjerps and Patsy Haccou. Information determines the optimal clutch sizes of competing insects: Stackelberg versus Nash equilibrium. Journal of Theoretical Biology, 163(4):473–483, August 21, 1993. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251938371132X.

Sozou:1994:PEM

[SH94] Peter D. Sozou and Alasdair I. Houston. Parental effort in a mating system involving two males and two females. Journal of Theoretical Biology, 171(3):251–266, December 7, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251938471229X.

Smith:1995:ASM

[SH95] Maynard J. Smith and D. G. C. Harper. Animal signals: Models and terminology. Journal of Theoretical Biology, 177(3):305–311, December 7, 1995. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519385702482.

Savill:1997:MMS

[SH97a] Nicholas J. Savill and Paulien Hogeweg. Modelling morphogenesis: From single cells to crawling slugs. Journal of Theoretical Biology, 184(3):229–235, February 7, 1997. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396902374.

Schmitt:1997:EED

[SH97b] Armin O. Schmitt and Hanspeter Herzel. Estimating the entropy of DNA sequences. Journal of Theoretical Biology, 188(3):369–377, October 7, 1997. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519397904938.

Schoniger:1990:STM

[SHB90] M. Schöniger, G. L. Hofacker, and B. Borstnik. Stochastic traits of molecular evolution — acceptance of point mutations in native actin genes. Journal of Theoretical Biology, 143(3):287–306, April 5, 1990. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305800311.

Shcherbak:1993:SAG

[Shc93a] V. I. Shcherbak. The symmetrical architecture of the genetic code systematization principle. Journal of Theoretical Biology, 162(3):395–398, June 7, 1993. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519383710957.

Shcherbak:1993:TCA

[Shc93b] V. I. Shcherbak. Twenty canonical amino acids of the genetic code: The arithmetical regularities. Part I. Journal of Theoretical Biology, 162(3):399–401, June 7, 1993. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519383710969.

Shcherbak:1994:SFT

[Shc94] V. I. Shcherbak. Sixty-four triplets and 20 canonical amino acids of the genetic code: The arithmetical regularities. part II. Journal of Theoretical Biology, 166(4):475–477, February 21, 1994. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519384710423.

Shen:1991:POS

[She91] Jianming Shen. On partitioning the opportunity for sexual selection. Journal of Theoretical Biology, 153(3):441–444, December 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305805818.

Sumida:1990:GAE

[SHMH90] B. H. Sumida, A. I. Houston, J. M. McNamara, and W. D. Hamilton. Genetic algorithms and evolution. Journal of Theoretical Biology, 147(1):59–84, November 7, 1990. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305802528.

Shulman:1991:CDH

[Shu91] Robert G. Shulman. Is channeling a disproveable hypothesis? Journal of Theoretical Biology, 152(1):133–134, September 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305805272.

Speakman:1994:HCD

[SHW94] J. R. Speakman, G. C. Hays, and P. I. Webb. Is hyperthermia a constraint on the diurnal activity of bats? Journal of Theoretical Biology, 171(3):325–339, December 7, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519384712355.

Sibly:1994:AAH

[Sib94] R. M. Sibly. An allelocentric analysis of Hamilton’s rule for overlapping generations. Journal of Theoretical Biology, 167(4):301–305, April 21, 1994. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519384710721.

Sibbald:1995:DPS

[Sib95] Peter R. Sibbald. Deducing protein structures using logic programming: exploiting minimum data of diverse types. Journal of Theoretical Biology, 173(4):361–375, April 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385700690.

Siller:1998:EHP

[Sil98] Steven Siller. The epistatic handicap principle does work. Journal of Theoretical Biology, 191(2):141–161, March 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519397905737.

Singh:1996:PDS

[Sin96] U. N. Singh. Polyribosome dynamics: Size-distribution as a function of attachment, translocation and release of ribosomes. Journal of Theoretical Biology, 179(2):147–159, March 14, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519396900557.

Stauffer:1994:SEK

[SJ94] Dietrich Stauffer and Naeem Jan. Size effects in Kauffman type evolution for rugged fitness landscapes. Journal of Theoretical Biology, 168(2):211–218, May 21, 1994. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384711003.

Schoner:1990:STQ

[SJK90] G. Schöner, W. Y. Jiang, and J. A. S. Kelso. A synergetic theory of quadrupedal gaits and gait transitions. Journal of Theoretical Biology, 142(3):359–391, February 9, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305805582.

Sjolander:1997:ERS

[Sjö97] Sverre Sjölander. On the evolution of reality — some biological prerequisites and evolutionary stages. Journal of Theoretical Biology, 187(4):595–600, August 21, 1997. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396903926.

Stein:1992:MIH

[SK92] G. I. Stein and B. N. Kudryavtsev. A method for investigating hepatocyte polyploidization kinetics during postnatal development in mammals. Journal of Theoretical Biology, 156(3):349– 363, June 7, 1992. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305806800.

Suematsu:1995:NPD

[SK95] Kazumi Suematsu and Minoru Kohno. A note on population dynamics. Journal of Theoretical Biology, 175(3):317–324, August 7, 1995. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519385701452.

Sachadyn:1996:PUM

[SK96a] Pawel Sachadyn and Józef Kur. A proposal for using modified site-specific recombination systems for making insertions into a chosen chromosomal site in vivo based on the analysis of I phage integration. Journal of Theoretical Biology, 180(1):55– 60, May 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396900776.

Sturis:1996:MSP

[SK96b] Jeppe Sturis and Carsten Knudsen. Modelling seasonal predator–prey interactions. Journal of Theoretical Biology, 178(1):99–103, January 7, 1996. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519396900090.

Sachadyn:1999:RNM

[SK99a] Pawel Sachadyn and Józef Kur. Reducing the number of microlocations in oligonucleotide microchip matrices by the application of degenerate oligonucleotides. Journal of Theoretical Biology, 197(3):393–401, April 7, 1999. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519398908803.

Suematsu:1999:AIG

[SK99b] Kazumi Suematsu and Minoru Kohno. Age invariant of Gompertz function and exponential decay of populations commensuration with CLOV experiments. Journal of Theoretical Biology, 201(4):231–238, December 21, 1999. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519399910263.

Satake:1998:NIR

[SKI98] Akiko Satake, Takuya Kubo, and Yoh Iwasa. Noise-induced regularity of spatial wave patterns in subalpine abies forests. Journal of Theoretical Biology, 195(4):465–479, December 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519398908074.

Saunders:1998:IRC

[SKW98] Peter T. Saunders, Johan H. Koeslag, and Jabus A. Wessels. Integral rein control in physiology. Journal of Theoretical Biology, 194(2):163–173, September 21, 1998. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519398907469.

Schnitzer:1991:EPB

[SL91a] Jan E. Schnitzer and Konstantine C. Lambrakis. Electrostatic potential and born energy of charged molecules interacting with phosholipid membranes: Calculation via 3-D numerical solution of the full Poisson equation. Journal of Theoretical Biology, 152(2):203–222, September 21, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305804539.

Singer:1991:APP

[SL91b] Debra F. Singer and Jennifer J. Linderman. Antigen processing and presentation: How can a foreign antigen be recognized in a sea of self proteins? Journal of Theoretical Biology, 151(3):385–404, August 7, 1991. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930580387X.

Stokes:1991:ARM

[SL91c] Cynthia L. Stokes and Douglas A. Lauffenburger. Analysis of the roles of microvessel endothelial cell random motility and chemotaxis in angiogenesis. Journal of Theoretical Biology, 152(3):377–403, October 7, 1991. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305802012.

Shea:1998:CRE

[SL98] Lonnie D. Shea and Jennifer J. Linderman. Compartmentalization of receptors and enzymes affects activation for a collision coupling mechanism. Journal of Theoretical Biology, 191(3):249–258, April 7, 1998. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519397905713.

Sept:1999:CKM

[SLBT99] D. Sept, H. J. Limbach, H. Bolterauer, and J. A. Tuszynski. A chemical kinetics model for microtubule oscillations. Journal of Theoretical Biology, 197(1):77–88, March 7, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251939890861X.

Solomovici:1997:DEC

[SLR97] Jean Solomovici, Thierry Lesnik, and Claude Reiss. Does Escherichia coli optimize the economics of the translation process? Journal of Theoretical Biology, 185(4):511–521, April 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519396903100.

Shankle:1998:EPD

[SLR^{+}98] W. Rodman Shankle, Benjamin H. Landing, Michael S. Rafii, Allen Schiano, James M. Chen, and Junko Hara. Evidence for a postnatal doubling of neuron number in the developing human cerebral cortex between 15 months and 6 years. Journal of Theoretical Biology, 191(2):115–140, March 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519397905701.

Saura:1993:OPP

[SLS93] Anssi Saura, Juhani Lokki, and Esko Suomalainen. Origin of polyploidy in parthenogenetic weevils. Journal of Theoretical Biology, 163(4):449–456, August 21, 1993. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519383711306.

Sansom:1990:AGS

[SM90a] M. S. P. Sansom and I. R. Mellor. Analysis of the gating of single ion channels using current-voltage surfaces. Journal of Theoretical Biology, 144(2):213–223, May 14, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305803212.

Sen:1990:KMC

[SM90b] Pali Sen and Donna Mohr. A kinetic model for calcium distribution. Journal of Theoretical Biology, 142(2):179–188, January 23, 1990. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305802206.

Schank:1992:COM

[SM92] Jeffrey C. Schank and Martha K. McClintock. A coupledoscillator model of ovarian-cycle synchrony among female rats. Journal of Theoretical Biology, 157(3):317–362, August 7, 1992. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305806149.

Sole:1995:RSO

[SM95] Ricard V. Solé and Susanna C. Manrubia. Are rainforests self-organized in a critical state? Journal of Theoretical Biology, 173(1):31–40, March 7, 1995. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385700409.

Schnell:1997:CFS

[SM97a] S. Schnell and C. Mendoza. Closed form solution for time-dependent enzyme kinetics. Journal of Theoretical Biology, 187(2):207–212, July 21, 1997. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519397904252.

Schnell:1997:ECT

[SM97b] S. Schnell and C. Mendoza. Enzymological considerations for the theoretical description of the quantitative competitive polymerase chain reaction (QC-PCR). Journal of Theoretical Biology, 184(4):433–440, February 21, 1997. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396902830.

Schnell:1997:TDP

[SM97c] S. Schnell and C. Mendoza. Theoretical description of the polymerase chain reaction. Journal of Theoretical Biology, 188(3):313–318, October 7, 1997. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519397904732.

Sitaramam:1997:EBO

[SM97d] V. Sitaramam and C. N. Madhavarao. The energetic basis of osmotolerance in plants: Physical principles. Journal of Theoretical Biology, 189(4):333–352, December 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519397905221.

Spier:1997:POD

[SM97e] Emmet Spier and David McFarland. Possibly optimal decisionmaking under self-sufficiency and autonomy. Journal of Theoretical Biology, 189(3):317–331, December 7, 1997. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519397905178.

Soares:1999:LIA

[SM99] Karol Maia Soares and Célia Cortez Maghelly. Linear instability analysis and mechanical interfacial tension. Journal of Theoretical Biology, 196(2):169–179, January 21, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519398908311.

Schroder:1995:KMN

[SMA95] Heinz C. Schröder, Werner E. G. Müller, and Paul S. Agutter. Kinetic models for nucleocytoplasmic transport of messenger RNA. Journal of Theoretical Biology, 174(2):169–177, May 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519385700896.

Smeller:1990:PDL

[Sme90] L. Smeller. The pressure dependence of the lipid phase transitions. effect of the pressure on the pre- and subtransition. Journal of Theoretical Biology, 142(4):453–462, February 22, 1990. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305801018.

Sole:1993:OCA

[SMG93] Ricard V. Solé, Octavio Miramontes, and Brian C. Goodwin. Oscillations and chaos in ant societies. Journal of Theoretical Biology, 161(3):343–357, April 7, 1993. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251938371060X.

Smith:1990:TSFa

[Smi90a] D. A. Smith. The theory of sliding filament models for muscle contraction. II. biochemically-based models of the contraction cycle. Journal of Theoretical Biology, 146(2):157–175, September 21, 1990. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251930580133X.

Smith:1990:TSFb

[Smi90b] D. A. Smith. The theory of sliding filament models for muscle contraction. III. dynamics of the five-state model. Journal of Theoretical Biology, 146(4):433–466, October 21, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305803728.

Smith:1990:MDI

[Smi90c] J. Maynard Smith. Models of a dual inheritance system. Journal of Theoretical Biology, 143(1):41–53, March 8, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305802875.

Smith:1991:SBT

[Smi91] W. John Smith. Singing is based on two markedly different kinds of signaling. Journal of Theoretical Biology, 152(2):241–253, September 21, 1991. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305804552.

Smith:1992:NNF

[Smi92] Charles E. Smith. A note on neuronal firing and input variability. Journal of Theoretical Biology, 154(3):271–275, February 7, 1992. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305801699.

Smith:1994:ADF

[Smi94] W. John Smith. Animal duets: Forcing a mate to be attentive. Journal of Theoretical Biology, 166(2):221–223, January 21, 1994. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519384710198.

Smith:1998:DME

[Smi98] Nick G. C. Smith. The dynamics of maternal-effect selfish genetic elements. Journal of Theoretical Biology, 191(2):173–180, March 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519397905798.

Smolen:1995:MGO

[Smo95] Paul Smolen. A model for glycolytic oscillations based on skeletal muscle phosphofructokinase kinetics. Journal of Theoretical Biology, 174(2):137–148, May 21, 1995. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385700872.

Soler:1998:MBE

[SMS98] Juan José Soler, Anders Pape Møller, and Manuel Soler. Mafia bahaviour and the evolution of faculatative virulence. Journal of Theoretical Biology, 191(3):267–277, April 7, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519397905993.

Stuckle:1992:POS

[SNG92] Erika E. Stückle, Peter J. Nielsen, and Ulrich Grob. Probability of occurrence of specific oligomers. Journal of Theoretical Biology, 159(3):299–306, December 7, 1992. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930580726X.

Stephani:1999:OSD

[SNH99] Amadeus Stephani, Juan Carlos Nuño, and Reinhart Heinrich. Optimal stoichiometric designs of ATP-producing systems as determined by an evolutionary algorithm. Journal of Theoretical Biology, 199(1):45–61, July 7, 1999. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519399909426.

Schreiner:1999:SSD

[SNK^{+}99] Wolfgang Schreiner, Friederike Neumann, Rudolf Karch, Martin Neumann, Susanne M. Roedler, and Adelheid End.

Shear stress distribution in arterial tree models, generated by constrained constructive optimization. Journal of Theoretical Biology, 198(1):27–45, May 7, 1999. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519399908986.

Schreiner:1996:SQB

[SNN^{+}96] Wolfgang Schreiner, Friederike Neumann, Martin Neumann, Adelheid End, and Michael R. Müller. Structural quantification and bifurcation symmetry in arterial tree models generated by constrained constructive optimization. Journal of Theoretical Biology, 180(2):161–174, May 21, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396900922.

Schreiner:1997:AVF

[SNN^{+}97] Wolfgang Schreiner, Friederike Neumann, Martin Neumann, Adelheid End, and Susanne M. Roedler. Anatomical variability and functional ability of vascular trees modeled by constrained constructive optimization. Journal of Theoretical Biology, 187(2):147–158, July 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519397904136.

Stephens:1995:EDI

[SNT95] D. W. Stephens, K. Nishimura, and K. B. Toyer. Error and discounting in the iterated Prisoner’s dilemma. Journal of Theoretical Biology, 176(4):457–469, October 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385702135.

Solc:1995:RIE

[Sol95] Milan Solc. Restricted intramolecular energy flow in the enzyme-substrate complex. Journal of Theoretical Biology, 175(1):57–61, July 7, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519385701208.

Sanyal:1991:IES

[SOR91] Nitish K. Sanyal, Rajendra Prasad Ojha, and Mihir Roychoudhury. Interaction energy studies on antibiotics nucleoside analogs 2-azaadenine. Journal of Theoretical Biology, 150(3):277–286, June 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519305804278.

Slominski:1990:TDH

[SP90] Andrzej Slominski and Ralf Paus. Are l-tyrosine and ldopa hormone-like bioregulators? Journal of Theoretical Biology, 143(1):123–138, March 8, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305802929.

Savinell:1992:NAIa

[SP92a] Joanne M. Savinell and Bernhard O. Palsson. Network analysis of intermediary metabolism using linear optimization. I. Development of mathematical formalism. Journal of Theoretical Biology, 154(4):421–454, February 21, 1992. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305801614.

Savinell:1992:NAIb

[SP92b] Joanne M. Savinell and Bernhard O. Palsson. Network analysis of intermediary metabolism using linear optimization: II. interpretation of hybridoma cell metabolism. Journal of Theoretical Biology, 154(4):455–473, February 21, 1992. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305801626.

Savinell:1992:OSMa

[SP92c] Joanne M. Savinell and Bernhard O. Palsson. Optimal selection of metabolic fluxes for in vivo measurement. I. Development of mathematical methods. Journal of Theoretical Biology, 155(2):201–214, March 21, 1992. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305805958.

Savinell:1992:OSMb

[SP92d] Joanne M. Savinell and Bernhard O. Palsson. Optimal selection of metabolic fluxes for in vivo measurement. II. application to Escherichia coli and hybridoma cell metabolism. Journal of Theoretical Biology, 155(2):215–242, March 21, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251930580596X.

Segel:1992:PCN

[SP92e] Lee A. Segel and Alan S. Perelson. Plasmid copy number control: a case study of the quasi-steady-state assumption. Journal of Theoretical Biology, 158(4):481–494, October 21, 1992. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305807118.

Sall:1994:MPA

[SP94] Torbjörn Säll and Peter Pettersson. A model of photosynthetic acclimation as a special case of reaction norms. Journal of Theoretical Biology, 166(1):1–8, January 7, 1994. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384710010.

Sanchez:1999:SDS

[SP99] Lucas Sánchez and André L. P. Perondini. Sex determination in sciarid flies: a model for the control of differential x-chromosome elimination. Journal of Theoretical Biology, 197(2):247–259, March 21, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519398908682.

Spaargaren:1992:TFB

[Spa92] D. H. Spaargaren. Transport function of branching structures and the “surface law” for basic metabolic rate. Journal of Theoretical Biology, 154(4):495–504, February 21, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305801651.

Speakman:1990:BFD

[Spe90] J. R. Speakman. On Blum’s four-dimensional geometric explanation for the 0·75 exponent in metabolic allometry. Journal of Theoretical Biology, 144(1):139–141, May 10, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305803054.

Spector:1994:DBC

[Spe94] David A. Spector. Definition in biology: The case of “bird song”. Journal of Theoretical Biology, 168(4):373–381, June 21, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519384711179.

Spivey:1991:ENC

[Spi91] H. Olin Spivey. Evidence of NADH channeling between dehydrogenases. Journal of Theoretical Biology, 152(1):103–107, September 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251930580520X.

Spirov:1993:CIS

[Spi93] A. V. Spirov. The change of initial symmetry in the patternform interaction model of sea urchin gastrulation. Journal of Theoretical Biology, 161(4):491–504, April 21, 1993. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519383710696.

Slominski:1993:MSR

[SPS93] Andrzej Slominski, Ralf Paus, and Dirk Schadendorf. Melanocytes as “sensory” and regulatory cells in the epidermis. Journal of Theoretical Biology, 164(1):103–120, September 7, 1993. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519383711422.

Spungin:1991:TSA

[Spu91] Ben Spungin. On the temporo-spatial array of dynein activity in the ciliary axoneme. Journal of Theoretical Biology, 151(3):313–321, August 7, 1991. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305803807.

Shapiro:1998:HFP

[SQ98] Bruce E. Shapiro and Hong Qian. Hysteresis in force probe measurements: a dynamical systems perspective. Journal of Theoretical Biology, 194(4):551–559, October 21, 1998. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519398907743.

Sherratt:1998:EGC

[SR98] Thomas N. Sherratt and Gilbert Roberts. The evolution of generosity and choosiness in cooperative exchanges. Journal of Theoretical Biology, 193(1):167–177, July 7, 1998. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519398907032.

Sherratt:1999:EQR

[SR99] Thomas N. Sherratt and Gilbert Roberts. The evolution of quantitatively responsive cooperative trade. Journal of Theoretical Biology, 200(4):419–426, October 21, 1999. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519399910056.

Srere:1991:CPC

[Sre91] Paul A. Srere. Channeling: the pathway that cannot be beaten. Journal of Theoretical Biology, 152(1):23, September 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305805016.

Savill:1997:SRS

[SRH97] Nicholas J. Savill, Pejman Rohani, and Paulien Hogeweg. Self-reinforcing spatial patterns enslave evolution in a hostparasitoid system. Journal of Theoretical Biology, 188(1): 11–20, September 7, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519397904483.

Srivastava:1991:PCE

[Sri91] D. K. Srivastava. Physiological constraints on evolution of enzymes for cellular metabolic pathways. Journal of Theoretical Biology, 152(1):93–100, September 7, 1991. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305805181.

Stekel:1995:CGS

[SRW95] D. Stekel, J. Rashbass, and E. D. Williams. A computer graphic simulation of squamous epithelium. Journal of Theoretical Biology, 175(3):283–293, August 7, 1995. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519395800115.

Stein:1990:TCT

[SS90a] Wilfred D. Stein and Aryeh D. Stein. Testing and characterizing the two-stage model of carcinogenesis for a wide range of human cancers. Journal of Theoretical Biology, 145(1):95– 122, July 9, 1990. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305805375.

Stillman:1990:OSP

[SS90b] Richard A. Stillman and William J. Sutherland. The optimal search path in a patchy environment. Journal of Theoretical Biology, 145(2):177–182, July 24, 1990. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305801237.

Svetina:1990:MBC

[Ss90c] Sasa Svetina and Bostjan seks. The mechanical behaviour of cell membranes as a possible physical origin of cell polarity. Journal of Theoretical Biology, 146(1):115–122, September 7, 1990. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305800475.

Smith:1993:OCS

[SS93a] J. Maynard Smith and E. Szathmáry. The origin of chromosomes I. Selection for linkage. Journal of Theoretical Biology, 164(4):437–446, October 21, 1993. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519383711653.

Smolle:1993:CST

[SS93b] Josef Smolle and Haro Stettner. Computer simulation of tumour cell invasion by a stochastic growth model. Journal of Theoretical Biology, 160(1):63–72, January 7, 1993. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519383710040.

Szathmary:1993:ECI

[SS93c] E. Szathmáry and J. Maynard Smith. The evolution of chromosomes II. molecular mechanisms. Journal of Theoretical Biology, 164(4):447–454, October 21, 1993. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519383711665.

Schole:1994:RTO

[SS94a] J. Schole and Chr. Schole. “Radical theory” of oxidative phosphorylation and photophosphorylation. Journal of Theoretical Biology, 169(2):197–207, July 21, 1994. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384711416.

Smolen:1994:PIR

[SS94b] Paul Smolen and Arthur Sherman. Phase independent resetting in relaxation and bursting oscillators. Journal of Theoretical Biology, 169(4):339–348, August 21, 1994. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384711568.

Stauffer:1994:HDS

[SS94c] Dietrich Stauffer and Muhammad Sahimi. High-dimensional simulation of simple immunological models. Journal of Theoretical Biology, 166(3):289–297, February 7, 1994. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384710265.

Suda:1994:MDS

[SS94d] Hitoshi Suda and Minoru Saito. Molecular dynamics simulations for actin monomers in solution. Journal of Theoretical Biology, 171(3):347–349, December 7, 1994. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384712379.

Sennerstam:1995:CCP

[SS95] Roland Sennerstam and Jan-Olof Strömberg. Cell cycle progression: computer simulation of uncoupled subcycles of DNA replication and cell growth. Journal of Theoretical Biology, 175(2):177–189, July 21, 1995. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385701300.

Schulz:1996:CAS

[SS96a] Arthur R. Schulz and Janos Südi. Control analysis of single enzyme sequences with abortive complexes and random substrate binding. Journal of Theoretical Biology, 182(3):397–403, October 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396901794.

Shiraishi:1996:TAC

[SS96b] Fumihide Shiraishi and Michael A. Savageau. The tricarboxylic acid cycle in Dictyostelium discoideum — two methods of analysis applied to the same model. Journal of Theoretical Biology, 178(2):219–222, January 21, 1996. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396900181.

Szathmary:1997:RRF

[SS97] Eörs Szathmáry and John Maynard Smith. From replicators to reproducers: the first major transitions leading to life. Journal of Theoretical Biology, 187(4):555–571, August 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519396903896.

Schole:1998:NMM

[SS98] J. Schole and C. Schole. A new model for the molecular mechanisms of photosynthetic water oxidation and photophosphorylation. Journal of Theoretical Biology, 192(2):247–260, May 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519398906592.

Shinbrot:1999:UVR

[SS99] Troy Shinbrot and Kathryn Scarbrough. Using variability to regulate long term biological rhythms. Journal of Theoretical Biology, 196(4):455–471, February 21, 1999. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519398908438.

Swinton:1994:TSA

[SSA94] Jonathan Swinton, A. Nicola Schweitzer, and Roy M. Anderson. Two signal activation as an explanation of high zone tolerance: a mathematical exploration of the nature of the second signal. Journal of Theoretical Biology, 169(1):23–30, July 7, 1994. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519384711271.

Sorribas:1993:MPC

[SSCC93] Albert Sorribas, Salvador Samitier, Enric I. Canela, and Marta Cascante. Metabolic pathway characterization from transient response data obtained in situ: Parameter estimation in Ssystem models. Journal of Theoretical Biology, 162(1):81–102, May 7, 1993. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519383710787.

Salvatori:1999:ETI

[SSCvdM99] V. Salvatori, A. K. Skidmore, F. Corsi, and F. van der Meer. Estimating temporal independence of radio-telemetry data on animal activity. Journal of Theoretical Biology, 198(4):567– 574, June 21, 1999. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519399909360.

Svrcek-Seiler:1998:MMB

[SSGR^{+}98] Wolfgang A. Svrcek-Seiler, Ilse C. Gebeshuber, Frank Rattay, Tamás S. Biro, and Harald Markum. Micromechanical models for the Brownian motion of hair cell stereocilia. Journal of Theoretical Biology, 193(4):623–630, August 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519398907299.

Sherratt:1993:CCE

[SSM93] Jonathan A. Sherratt, E. Helene Sage, and J. D. Murray. Chemical control of eukaryotic cell movement: a new model. Journal of Theoretical Biology, 162(1):23–40, May 7, 1993. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251938371074X.

Selvakumar:1993:MME

[SSS93] S. Selvakumar, M. Sharan, and M. P. Singh. A mathematical model for the elimination of carbon monoxide in humans. Journal of Theoretical Biology, 162(3):321–336, June 7, 1993. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519383710908.

Snyder:1995:GEI

[SSSF95] G. K. Snyder, Brandon Sheafor, David Scholnick, and Caroline Farrelly. Gas exchange in the insect tracheal system. Journal of Theoretical Biology, 172(3):199–207, February 7, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519385700161.

Shinoda:1991:ESD

[SST91] Takako Shinoda, Nobuyuki Shima, and Masaru Tsukada. Electronic structure of DNA by DV-Xα cluster calculations. I. d(GG) · d(CC), d(CG) 2, d(GC) 2 A and B conformations. Journal of Theoretical Biology, 151(4):433–466, August 21, 1991. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305803650.

Suda:1993:IFB

[ST93] Hitoshi Suda and Tad W. Taylor. Intermolecular forces between the motor protein and the filament. Journal of Theoretical Biology, 161(1):39–50, March 7, 1993. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519383710386.

Starzak:1990:KIG

[Sta90] Michael E. Starzak. The kinetics of interacting gates in potassium channels. a comparison of linear and non-linear models. Journal of Theoretical Biology, 144(1):1–14, May 10, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305802966.

Staudte:1992:BAM

[Sta92] Robert G. Staudte. A bifurcating autoregression model for cell lineages with variable generation means. Journal of Theoretical Biology, 156(2):183–195, May 21, 1992. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305806721.

Stadler:1996:HCS

[Sta96a] Bärbel M. R. Stadler. Heteroclinic cycles and segregation distortion. Journal of Theoretical Biology, 183(4):363–379, December 21, 1996. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396902283.

Stanfel:1996:NAC

[Sta96b] Larry E. Stanfel. A new approach to clustering the amino acid. Journal of Theoretical Biology, 183(2):195–205, November 21, 1996. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396902131.

Stanbro:1998:KMM

[Sta98] William D. Stanbro. A kinetic model of the myeloperoxidasehydrogen peroxidase-chloride ion system in phagolysosomes. Journal of Theoretical Biology, 193(1):59–68, July 7, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519398906865.

Stanbro:1999:KMS

[Sta99] William D. Stanbro. A kinetic model of the system: Tyrosyl radical-nitrogen oxide-superoxide ion. Journal of Theoretical Biology, 197(4):557–567, April 21, 1999. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519398908955.

Stein:1990:EDP

[Ste90a] W. D. Stein. Energetics and the design principles of the Na/ K-ATPase. Journal of Theoretical Biology, 147(2):145–159, November 21, 1990. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305800499.

Steinberg:1990:CSE

[Ste90b] Izchak Z. Steinberg. Computer simulations of electrical bistability in excitable cells due to non-inactivating sodium channels: Space- and time-dependent behavior. Journal of Theoretical Biology, 144(1):75–92, May 10, 1990. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305803005.

Steele:1991:CET

[Ste91a] John H. Steele. Can ecological theory cross the land-sea boundary? Journal of Theoretical Biology, 153(3):425–436, December 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251930580579X.

Stein:1991:RLS

[Ste91b] W. D. Stein. Rate-limiting steps in carcinogenesis and their bearing on the incidence/age data for cutaneous melanoma and hepatocellular carcinoma. Journal of Theoretical Biology, 153(4):519–529, December 21, 1991. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305801547.

Steinberg:1994:BME

[Ste94] Izchak Z. Steinberg. Brownian motion of the end-to-end distance in oligopeptide molecules: Numerical solution of the diffusion equations as coupled first order linear differential equations. Journal of Theoretical Biology, 166(2):173–187, January 21, 1994. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519384710162.

Stekel:1997:RIC

[Ste97] Dov J. Stekel. The role of inter-cellular adhesion in the recirculation of t lymphocytes. Journal of Theoretical Biology, 186(4):491–501, June 21, 1997. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519397904082.

Stitt:1991:CPS

[Sti91] M. Stitt. Channelling in photosynthesis: the sublime and the ridiculous. Journal of Theoretical Biology, 152(1):35–37, September 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305805041.

Suenobu:1993:IMO

[STI93] Katsuhiro Suenobu, Seiji Tani, and Akira Imamura. An ab initio molecular orbital study on effect of electric field strength on ionic selectivity of the sodium channel. Journal of Theoretical Biology, 161(3):395–403, April 7, 1993. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519383710635.

Stockhoff:1994:MDC

[Sto94] Brian A. Stockhoff. Maximization of daily canopy photosynthesis: Effects of herbivory on optimal nitrogen distribution. Journal of Theoretical Biology, 169(3):209–220, August 7, 1994. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519384711428.

Stocker:1995:RMD

[Stö95] Sabine Stöcker. Regarding mutations in daisyworld models. Journal of Theoretical Biology, 175(4):495–501, August 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519385701567.

Stormo:1998:ICF

[Sto98] Gary D. Stormo. Information content and free energy in DNA-protein interactions. Journal of Theoretical Biology, 195(1):135–137, November 7, 1998. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519398907858.

Strazewski:1991:PMT

[Str91] Peter Strazewski. Psoralen mutagenesis by transient misalignment? Journal of Theoretical Biology, 151(3):429–431, August 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251930580390X.

Streeter:1995:NMR

[Str95] John G. Streeter. A new model for the rapid effects of non-invasive treatments on nitrogenase and respiratory activity in legume nodules. Journal of Theoretical Biology, 174(4):441–442, IN3–IN4, 443–452, June 21, 1995. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385701105.

Suda:1990:MFA

[Sud90] Hitoshi Suda. Molecular friction in an actomyosin molecular machine. Journal of Theoretical Biology, 146(3):341–346, October 7, 1990. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305807441.

Sugimoto:1998:TAS

[Sug98] Takeshi Sugimoto. A theoretical analysis of sea-anchor soaring. Journal of Theoretical Biology, 192(3):393–402, June 7, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519398906695.

Sun:1995:RRA

[Sun95a] Shaojian Sun. Reduced representation approach to protein tertiary structure prediction: statistical potential and simulated annealing. Journal of Theoretical Biology, 172(1):13–32, January 7, 1995. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519385700021.

Sung:1995:CTS

[Sun95b] Shen-Shu Sung. Constant temperature simulations of helix folding. Journal of Theoretical Biology, 173(4):389–400, April 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519385700719.

Stewart:1990:DCI

[SV90] John Stewart and Francisco J. Varela. Dynamics of a class of immune networks II. oscillatory activity of cellular and humoral components. Journal of Theoretical Biology, 144(1):103– 115, May 10, 1990. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305803029.

Stewart:1991:MSS

[SV91] John Stewart and Francisco J. Varela. Morphogenesis in shapespace. elementary meta-dynamics in a model of the immune network. Journal of Theoretical Biology, 153(4):477–498, December 21, 1991. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305801523.

Sole:1992:SSC

[SV92] Ricard V. Solé and Joaquim Valls. On structural stability and chaos in biological systems. Journal of Theoretical Biology, 155(1):87–102, March 7, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305805508.

Serra:1997:MBD

[SV97] R. Serra and M. Villani. Modelling bacterial degradation of organic compounds with genetic networks. Journal of Theoretical Biology, 189(1):107–119, November 7, 1997. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519397904975.

Svensson:1994:RPI

[Sve94] Gunnar Svensson. Reflections on the problem of identifying mind and brain. Journal of Theoretical Biology, 171(1):93– 100, July 1994. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251938471215X.

Sulzer:1999:ACS

[SvH99] Bernhard Sulzer and J. Leo van Hemmen. Adaptive control: a strategy to treat autoimmunity. Journal of Theoretical Biology, 196(1):73–79, January 7, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519398908189.

Sanchez:1997:EDV

[SvHT97] L. Sánchez, J. van Helden, and D. Thieffry. Establishement of the dorso-ventral pattern during embryonic development of Drosophila melanogaster: a logical analysis. Journal of Theoretical Biology, 189(4):377–389, December 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519397905233.

Stickle:1994:FDC

[SVRF94] Douglas F. Stickle, Karen M. Vossen, Daniel A. Riley, and Michael G. Fried. Free DNA concentration in E. coli estimated by an analysis of competition for DNA binding proteins. Journal of Theoretical Biology, 168(1):1–12, May 7, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519384710824.

Schreibmayer:1991:DSV

[SW91] W. Schreibmayer and M. Waller. Domain specialization in voltage-dependent Na + and Ca^{2+} channels. Journal of Theoretical Biology, 151(1):141–143, July 7, 1991. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305801481.

Schulgasser:1992:SSS

[SW92] K. Schulgasser and A. Witztum. On the strength, stiffness and stability of tubular plant stems and leaves. Journal of Theoretical Biology, 155(4):497–510, IN1, 511–515, April 21, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305806320.

Spence:1995:PCW

[SW95a] Richard D. Spence and Hsin-I Wu. Plant cell wall elasticity III:: a polymer elastic interpretation of inverse pressurevolume curves. Journal of Theoretical Biology, 177(1):59–65, November 7, 1995. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251938570224X.

Stevens:1995:GCR

[SW95b] Ernest E. Stevens and R. Haven Wiley. Genetic consequences of restricted dispersal and incest avoidance in a cooperatively breeding wren. Journal of Theoretical Biology, 175(4):423–436, August 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251938570152X.

Strausbaugh:1996:HDS

[SW96] Linda D. Strausbaugh and Scott M. Williams. High density of an SAR-associated Motif differentiates heterochromatin from euchromatin. Journal of Theoretical Biology, 183(2):159–167, November 21, 1996. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251939690209X.

Sandler:1998:NCD

[SW98] U. Sandler and A. Wyler. Non-coding DNA can regulate gene transcription by its base Pair’s distribution. Journal of Theoretical Biology, 193(1):85–90, July 7, 1998. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251939890683X.

Smith:1991:MLE

[SWA91] Bruce R. Smith, J. Martin Wojtowicz, and Harold L. Atwood. Maximum likelihood estimation of non-uniform transmitter release probabilities at the crayfish neuromuscular junction. Journal of Theoretical Biology, 150(4):457–472, June 21, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305804400.

Skierczynski:1996:MHS

[SWKS96] B. A. Skierczynski, R. J. A. Wilson, W. B. Kristan, Jr., and R. Skalak. A model of the hydrostatic skeleton of the leech. Journal of Theoretical Biology, 181(4):329–342, August 21, 1996. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396901356.

Shabalina:1991:FNN

[SYK91] Svetlana A. Shabalina, Olga V. Yurieva, and Alexey S. Kondrashov. On the frequencies of nucleotides and nucleotide substitutions in conservative regulatory DNA sequences. Journal of Theoretical Biology, 149(1):43–54, March 7, 1991. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305800700.

Sykora:1992:MPM

[Sýk92] Peter Sýkora. Macroevolution of plasmids: a model for plasmid speciation. Journal of Theoretical Biology, 159(1):53–65, November 7, 1992. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305807672.

Symonds:1994:DMC

[Sym94] Neville Symonds. Directed mutation: a current perspective. Journal of Theoretical Biology, 169(4):317–322, August 21, 1994. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519384711532.

Szathmary:1992:NSD

[Sza92a] Eörs Szathmáry. Natural selection and dynamical coexistence of defective and complementing virus segments. Journal of Theoretical Biology, 157(3):383–406, August 7, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305806174.

Szathmary:1992:VSL

[Sza92b] Eörs Szathmáry. Viral sex, levels of selection, and the origin of life. Journal of Theoretical Biology, 159(1):99–109, November 7, 1992. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305807702.

Szathmary:1993:COD

[Sza93a] Eörs Szathmáry. Co-operation and defection: Playing the field in virus dynamics. Journal of Theoretical Biology, 165(3):341–356, December 7, 1993. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519383711938.

Szathmary:1993:NRD

[Sza93b] Eörs Szathmáry. A note on the reduction of the dynamics of multilocus diploid genetic systems with multiplicative fitness. Journal of Theoretical Biology, 164(3):351–358, October 7, 1993. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519383711598.

Szathmary:1994:TMS

[Sza94] Eörs Szathmäry. Toy models for simple forms of multicellularity, soma and germ. Journal of Theoretical Biology, 169(2):125– 132, July 21, 1994. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519384711349.

Szamado:1999:VHP

[Szá99] Szabolcs Számadó. The validity of the handicap principle in discrete action-response games. Journal of Theoretical Biology, 198(4):593–602, June 21, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519399909359.

Sunada:1998:RBD

[SZK98] Shigeru Sunada, Lijiang Zeng, and Keiji Kawachi. The relationship between dragonfly wing structure and torsional deformation. Journal of Theoretical Biology, 193(1):39–45, July 7, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519398906786.

Tainaka:1999:PPL

[TA99] Kei-Ichi Tainaka and Noriyuki Araki. Press perturbation in lattice ecosystems: Parity law and optimum strategy. Journal of Theoretical Biology, 197(1):1–13, March 7, 1999. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519398908293.

Tolkamp:1998:SSF

[TAA^{+}98] B. J. Tolkamp, D. J. Allcroft, E. J. Austin, B. L. Nielsen, and I. Kyriazakis. Satiety splits feeding behaviour into bouts. Journal of Theoretical Biology, 194(2):235–250, September 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519398907597.

Taber:1998:BGL

[Tab98] Larry A. Taber. Biomechanical growth laws for muscle tissue. Journal of Theoretical Biology, 193(2):201–213, July 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519397906184.

Takada:1995:ESI

[Tak95a] Takenori Takada. Evolution of semelparous and iteroparous perennial plants:: Comparison between the density-independent and density-dependent dynamics. Journal of Theoretical Biology, 173(1):51–60, March 7, 1995. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519385700422.

Takeda:1995:GEA

[Tak95b] Yasuhiko Takeda. A generalized epistasis analysis as a method of reconstruction of developmental process and its application to drosophila dorso-ventral axis formation. Journal of Theoretical Biology, 176(1):47–58, September 7, 1995. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385701750.

Takahasi:1998:ETB

[Tak98] Kiyosi Takahasi. Evolution of transmission bias in cultural inheritance. Journal of Theoretical Biology, 190(2):147–159, January 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519397905415.

Tanaka:1991:ESC

[Tan91] Yoshinari Tanaka. The evolution of social communication systems in a subdivided population. Journal of Theoretical Biology, 149(2):145–163, March 21, 1991. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305802747.

Tanaka:1994:DMC

[Tan94a] Masao Tanaka. Dynamics of a materially closed model ecosystem. Journal of Theoretical Biology, 168(2):131–140, May 21, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519384710939.

Tuomi:1994:DPD

[TAN94b] J. Tuomi, M. Augner, and P. Nilsson. A dilemma of plant defences: Is it really worth Killing the herbivore? Journal of Theoretical Biology, 170(4):427–430, October 21, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384712045.

Tanaka:1996:SSE

[Tan96] Yoshinari Tanaka. Sexual selection enhances population extinction in a changing environment. Journal of Theoretical Biology, 180(3):197–206, June 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251939690096X.

Tan:1999:IIV

[Tan99] Zheng Tan. Intramitotic and intraclonal variation in proliferative potential of human diploid cells: Explained by telomere shortening. Journal of Theoretical Biology, 198(2):259–268, May 21, 1999. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519399909141.

Tarutis:1994:MVA

[Tar94] William J. Tarutis. A mean-variance approach for describing organic matter decomposition. Journal of Theoretical Biology, 168(1):13–18, May 7, 1994. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519384710836.

Taylor:1991:OLH

[Tay91] Peter Taylor. Optimal life histories with age dependent tradeoff curves. Journal of Theoretical Biology, 148(1):33–48, January 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305804643.

Travis:1991:DSC

[TB91] Curtis C. Travis and Pamela D. Birkner. Druckery slope controlled by mitotic rate of enzymatically-altered foci. Journal of Theoretical Biology, 149(2):217–227, March 21, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305802784.

Theraulaz:1995:MCB

[TB95] Guy Theraulaz and Eric Bonabeau. Modelling the collective building of complex architectures in social insects with lattice swarms. Journal of Theoretical Biology, 177(4):381–400, December 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251938570255X.

z:1996:KSD

[tBázsáSH96] S. tefan Bal á z, K. Pir selov á, T. W. Schultz, and J. Hermens. Kinetics of subcellular distribution of multiply ionizable compounds: a mathematical description and its use in QSAR. Journal of Theoretical Biology, 178(1):7–16, January 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396900028.

Theraulaz:1995:SOH

[TBD95] Guy Theraulaz, Eric Bonabeau, and Jean-Louis Deneubourg. Self-organization of hierarchies in animal societies: the case of the primitively eusocial wasp polistes dominulus Christ. Journal of Theoretical Biology, 174(3):313–323, June 7, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385701014.

Takada:1997:OSM

[TC97] Takenori Takada and Hal Caswell. Optimal size at maturity in size-structured populations. Journal of Theoretical Biology, 187(1):81–93, July 7, 1997. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519397904203.

Tourtellot:1991:PMT

[TCB91] Michael K. Tourtellot, Robert D. Collins, and William J. Bell. The problem of movelength and turn definition in analysis of orientation data. Journal of Theoretical Biology, 150(3):287– 297, June 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251930580428X.

Taylor:1995:CUW

[TCE95] G. C. Taylor, J. A. Coles, and J. C. Eilbeck. Conditions under which Na + channels can boost conduction of small graded potentials. Journal of Theoretical Biology, 172(4):379–386, February 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519385700355.

Tchuraev:1991:NMA

[Tch91] Rustem N. Tchuraev. A new method for the analysis of the dynamics of the molecular genetic control systems. I. Description of the method of generalized threshold models. Journal of Theoretical Biology, 151(1):71–87, July 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305801444.

Tyson:1990:AKH

[TCLB90] John J. Tyson, Katherine C. Chen, Muriel Lederman, and Robert C. Bates. Analysis of the kinetic hairpin transfer model for parvoviral DNA replication. Journal of Theoretical Biology, 144(2):155–169, May 14, 1990. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305803169.

Thompson:1995:NNP

[TCZ95] Thomas B. Thompson, Kuo-Chen Chou, and Chong Zheng. Neural network prediction of the HIV-1 protease cleavage sites. Journal of Theoretical Biology, 177(4):369–379, December 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519385702548.

Thiebaux:1992:MAB

[TD92] M. L. Thiebaux and L. M. Dickie. Models of aquatic biomass size spectra and the common structure of their solutions. Journal of Theoretical Biology, 159(2):147–161, November 21, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251930580699X.

Takumi:1996:SAM

[TD96] Katsuhisa Takumi and Rob J. De Boer. Self assertion modeled as a network repertoire of multi-determinant antibodies. Journal of Theoretical Biology, 183(1):55–66, November 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396902015.

Taber:1996:TSS

[TE96] Larry A. Taber and Daniel W. Eggers. Theoretical study of stress-modulated growth in the aorta. Journal of Theoretical Biology, 180(4):343–357, June 21, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519396901071.

Theodoridis:1996:ESI

[TéS96] George C. Theodoridis, Antharvedi Ann é, and Lawrence Stark. On evolutive systems and the initial evolution of structure and function. Journal of Theoretical Biology, 178(1):61– 88, January 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396900077.

Tsonis:1991:PDC

[TET91] Anastasios A. Tsonis, James B. Elsner, and Panagiotis A. Tsonis. Periodicity in DNA coding sequences: Implications in gene evolution. Journal of Theoretical Biology, 151(3):323–331, August 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305803819.

Tsonis:1997:DL

[TET97] Anastasios A. Tsonis, James B. Elsner, and Panagiotis A. Tsonis. Is DNA a language? Journal of Theoretical Biology, 184(1):25–29, January 7, 1997. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519396902398.

Treinin:1993:UCD

[TF93] Millet Treinin and Dror G. Feitelson. Unequal cell division as a driving force during differentiation. Journal of Theoretical Biology, 160(1):85–95, January 7, 1993. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519383710064.

Thomas:1994:MCA

[TF94] Simon Thomas and David A. Fell. Metabolic control analysis: Sensitivity of control coefficients to experimentally determined variables. Journal of Theoretical Biology, 167(2):175–200, March 21, 1994. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519384710630.

Taylor:1996:HMK

[TF96a] Peter D. Taylor and Steven A. Frank. How to make a kin selection model. Journal of Theoretical Biology, 180(1):27–37, May 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396900752.

Thomas:1996:DMC

[TF96b] Simon Thomas and David A. Fell. Design of metabolic control for large flux changes. Journal of Theoretical Biology, 182(3):285–298, October 7, 1996. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519396901666.

Tanaka:1999:TCC

[TF99] Mark M. Tanaka and Marcus W. Feldman. Theoretical considerations of cross-immunity, recombination and the evolution of new parasitic strains. Journal of Theoretical Biology, 198(2):145–163, May 21, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519399909062.

Taylor:1993:SHT

[TGS93] Tad W. Taylor, Yoichi Goto, and Hiroyuki Suga. On the solutions of Huxley-type models in cardiac muscle fiber contractions. Journal of Theoretical Biology, 165(3):409–416, December 7, 1993. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519383711975.

Tseng:1990:CPE

[TH90] Shiojenn Tseng and Jyh-Ping Hsu. A comparison of the parameter estimating procedures for the Michaelis–Menten model. Journal of Theoretical Biology, 145(4):457–464, August 23, 1990. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305804813.

Takahashi:1998:OMP

[TH98a] Satoshi Takahashi and Michio Hori. Oscillation maintains polymorphisms — a model of lateral asymmetry in two competing scale-eating cichlids. Journal of Theoretical Biology, 195(1):1–12, November 7, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519398907445.

Takumi:1998:EIR

[TH98b] Katsuhisa Takumi and Pauline Hogeweg. Evolution of the immune repertoire with and without somatic DNA recombination. Journal of Theoretical Biology, 192(3):309–317, June 7, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519398906506.

Tofts:1992:AAL

[THF92] C. Tofts, M. Hatcher, and N. R. Franks. The autosynchronization of the ant leptothorax acervorum (fabricius): Theory, testability and experiment. Journal of Theoretical Biology, 157(1):71–82, July 7, 1992. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251930580757X.

Thierry:1990:FLB

[Thi90] Bernard Thierry. Feedback loop between kinship and dominance: the macaque model. Journal of Theoretical Biology, 145(4):511–522, August 23, 1990. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305804850.

Thierry:1993:ECC

[Thi93] B. Thierry. Emergent constraints condition the action of ultimate causes. Journal of Theoretical Biology, 160(3):403–405, February 7, 1993. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251938371026X.

Thomas:1991:RNS

[Tho91] René Thomas. Regulatory networks seen as asynchronous automata: a logical description. Journal of Theoretical Biology, 153(1):1–23, November 7, 1991. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305803509.

Thomas:1995:CPF

[Tho95] David J. Thomas. Chaos in protein folding: a simple onedimensional model. Journal of Theoretical Biology, 174(3):299– 311, June 7, 1995. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519385701002.

Thomas:1996:FBW

[Tho96a] Adrian L. R. Thomas. The flight of birds that have wings and a tail: Variable geometry expands the envelope of flight performance. Journal of Theoretical Biology, 183(3):237–245, December 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396902179.

Thomas:1996:WDB

[Tho96b] Adrian L. R. Thomas. Why do birds have tails? the tail as a drag reducing flap, and trim control. Journal of Theoretical Biology, 183(3):247–253, December 7, 1996. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396902180.

Tuszynski:1995:FBM

[THS^{+}95] J. A. Tuszyński, S. Hameroff, M. V. Satarić, B. Trpisová, and M. L. A. Nip. Ferroelectric behavior in microtubule dipole lattices: Implications for information processing, signaling and assembly/disassembly. Journal of Theoretical Biology, 174(4):371–380, June 21, 1995. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385701051.

Tanemura:1991:DDC

[THY91] Masaharu Tanemura, Hisao Honda, and Akihiro Yoshida. Distribution of differentiated cells in a cell sheet under the lateral inhibition rule of differentiation. Journal of Theoretical Biology, 153(3):287–300, December 7, 1991. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305805715.

Taylor:1993:DAA

[TJ93] William R. Taylor and David T. Jones. Deriving an amino acid distance matrix. Journal of Theoretical Biology, 164(1):65–83, September 7, 1993. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519383711409.

Tomiuk:1996:MLE

[TL96] Jürgen Tomiuk and Volker Loeschcke. A maximum-likelihood estimator of the genetic identity betweenn polyploid species. Journal of Theoretical Biology, 179(1):51–54, March 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396900478.

Tuckwell:1998:SME

[TL98] Henry C. Tuckwell and Emmanuelle Le Corfec. A stochastic model for early HIV-1 population dynamics. Journal of Theoretical Biology, 195(4):451–463, December 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519398908062.

Tomiuk:1994:OPP

[TLS94] J. Tomiuk, V. Loeschcke, and M. Schneider. On the origin of polyploid parthenogenetic races in the weevil polydrusus mollis (coleoptera: Curculionidae). Journal of Theoretical Biology, 167(1):89–92, March 7, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519384710538.

Tortolero:1992:MTF

[TM92a] M. Tortolero and J. R. Medina. Males and the twofold disadvantage of dioecy. Journal of Theoretical Biology, 159(4):431–437, December 21, 1992. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305806885.

Tranquillo:1992:CMF

[TM92b] Robert T. Tranquillo and J. D. Murray. Continuum model of fibroblast-driven wound contraction: Inflammation-mediation. Journal of Theoretical Biology, 158(2):135–172, September 21, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305807155.

Takesue:1998:CDR

[TMI98] Akio Takesue, Atsushi Mochizuki, and Yoh Iwasa. Celldifferentiation rules that generate regular Mosaic patterns: Modelling motivated by cone Mosaic formation in fish retina. Journal of Theoretical Biology, 194(4):575–586, October 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519398907779.

Tohya:1999:FCM

[TMI99] Shusaku Tohya, Atsushi Mochizuki, and Yoh Iwasa. Formation of cone Mosaic of zebrafish retina. Journal of Theoretical Biology, 200(2):231–244, September 21, 1999. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519399909906.

Tohya:1998:RSS

[TMII98] Shusaku Tohya, Atsushi Mochizuki, Shuhei Imayama, and Yoh Iwasa. On rugged shape of skin tumor (basal cell carcinoma). Journal of Theoretical Biology, 194(1):65–78, September 7, 1998. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519398907433.

Trommershauser:1999:SMC

[TMZ99] Julia Trommershäuser, Jörg Marienhagen, and Annette Zippelius. Stochastic model of central synapses: Slow diffusion of transmitter interacting with spatially distributed receptors and transporters. Journal of Theoretical Biology, 198(1):101– 120, May 7, 1999. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519399909050.

Takada:1996:OAS

[TN96] Takenori Takada and Hisao Nakajima. The optimal allocation for seed reproduction and vegetative reproduction in perennial plants: An application to the densitydependent transition matrix model. Journal of Theoretical Biology, 182(2):179–191, September 21, 1996. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396901538.

Topham:1990:GTT

[Top90] Christopher M. Topham. A generalized theoretical treatment of the kinetics of an enzyme-catalysed reaction in the presence of an unstable irreversible modifier. Journal of Theoretical Biology, 145(4):547–572, August 23, 1990. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305804886.

Toquenaga:1997:HSC

[Toq97] Yukihiko Toquenaga. Historicity of a simple competition model. Journal of Theoretical Biology, 187(2):175–181, July 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519397904288.

Torres:1990:IPT

[Tor90] José-Leonel Torres. Information, photosynthesis and temperature of the earth. Journal of Theoretical Biology, 144(1):131– 137, May 10, 1990. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305803042.

Torres:1996:MCA

[Tor96] Néstor V. Torres. Metabolic control analysis of inhibitory feedback interaction: Application to biotechnological processes. Journal of Theoretical Biology, 182(3):405–410, October 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396901800.

Tributsch:1993:COE

[TP93] H. Tributsch and L. Pohlmann. Co-operative electron transfer: Superconduction in biomolecular structures? Journal of Theoretical Biology, 165(2):225–250, November 21, 1993. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519383711872.

Talussot:1994:TND

[TP94] Corinne Talussot and Gabriel Ponsin. The transformation of nascent disc shaped high density lipoproteins into mature spherical high density lipoproteins: Conformational analysis of Apolipoprotein A-I in model intermediate particles. Journal of Theoretical Biology, 169(2):143–152, July 21, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519384711362.

Tributsch:1996:SME

[TP96] H. Tributsch and L. Pohlmann. Synergetic mechanisms in energy and signal transduction: Photo oscillating proton transport in Bacteriorhodopsin. Journal of Theoretical Biology, 178(1):17–28, January 7, 1996. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251939690003X.

Thuijsman:1995:AMF

[TPAS95] F. Thuijsman, B. Peleg, M. Amitai, and A. Shmida. Automata, matching and foraging behavior of bees. Journal of Theoretical Biology, 175(3):305–316, August 7, 1995. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385701440.

Turcotte:1998:NSB

[TPN98] D. L. Turcotte, J. D. Pelletier, and W. I. Newman. Networks with side branching in biology. Journal of Theoretical Biology, 193(4):577–592, August 21, 1998. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519398907238.

Treisman:1995:MSO

[Tre95] Michel Treisman. The multiregional and single origin hypotheses of the evolution of modern man:: a reconciliation. Journal of Theoretical Biology, 173(1):23–29, March 7, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519385700392.

Tributsch:1992:WCT

[Tri92]  Helmut Tributsch.  The water-cohesion-tension insuffi-

ciency syndrome of forest decline. Journal of Theoretical Biology, 156(2):235–267, May 21, 1992. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305806757.

Tuckwell:1996:TDS

[TRVL96] Henry C. Tuckwell, Jean-Pierre Rospars, Arthur Vermeulen, and Petr Lánsky. Time-dependent solutions for a cable model of an olfactory receptor neuron. Journal of Theoretical Biology, 181(1):25–31, July 7, 1996. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519396901113.

Trzeciakowski:1999:SAEa

[Trz99a] J. P. Trzeciakowski. Stimulus amplification, efficacy, and the operational model. part i — binary complex occupancy mechanisms. Journal of Theoretical Biology, 198(3):329–346, June 7, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519399909190.

Trzeciakowski:1999:SAEb

[Trz99b] J. P. Trzeciakowski. Stimulus amplification, efficacy, and the operational model. part II-ternary complex occupancy mechanisms. Journal of Theoretical Biology, 198(3):347–374, June 7, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519399909207.

Tawada:1991:PFE

[TS91] Katsuhisa Tawada and Ken Sekimoto. Protein friction exerted by motor enzymes through a weak-binding interaction. Journal of Theoretical Biology, 150(2):193–200, May 21, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305803315.

Tokuda:1994:SMC

[TS94] Hiroki Tokuda and Hiromi Seno. Some mathematical considerations on the parent-offspring conflict phenomenon. Journal of Theoretical Biology, 170(2):145–157, September 21, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251938471174X.

Tracy:1995:PES

[TS95] Nola D. Tracy and John W. Seaman, Jr. Properties of evolutionarily stable learning rules. Journal of Theoretical Biology, 177(2):193–198, November 21, 1995. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251938570238X.

Toha:1990:PRS

[TSC90] J. C. Tohá, M. A. Soto, and H. Chinga. Pattern recognition of the secondary structure of proteins (α-helix and β-structure).

Journal of Theoretical Biology, 146(2):145–156, September 21, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305801328.

Torres:1993:EFM

[TT93] José-Leonel Torres and L. E. H. Trainor. Energy function for a model of ant colonies. Journal of Theoretical Biology, 165(2):177–184, November 21, 1993. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519383711835.

Takahashi:1997:ESS

[TUHT97] Kengo Takahashi, Go Uchida, Zi-Song Hu, and Yoshimi Tsuchiya. Entrainment of the self-sustained oscillation in a Physarum polycephalum strand as a one-dimensionally coupled oscillator system. Journal of Theoretical Biology, 184(2):105–110, January 21, 1997. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251939690260X.

Tan:1997:DTH

[TW97] Yunho Tan and Zhi-Xin Wang. Derivation of the Tsao– Hanson equation using a statistical thermodynamic method. Journal of Theoretical Biology, 185(4):549–551, April 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519396903240.

Thorley:1999:TPS

[TW99] Joseph L. Thorley and Mark Wilkinson. Testing the phylogenetic stability of early tetrapods. Journal of Theoretical Biology, 200(3):343–344, October 7, 1999. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519399909992.

Tyson:1993:PCE

[TWB^{+}93] Rebecca Tyson, Wendell T. Welch, John Berreen, Craig Wells, and Carl E. Pearson. Pressure changes in the eye due to an injection of inert gases: a theoretical model.

Journal of Theoretical Biology, 164(1):15–36, September 7, 1993. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519383711380.

Travis:1990:IEP

[TWW90] Curtis C. Travis, Robin K. White, and Richard C. Ward. Interspecies extrapolation of pharmacokinetics. Journal of Theoretical Biology, 142(3):285–304, February 9, 1990. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305805545.

Tyvand:1993:EAT

[Tyv93] Peder A. Tyvand. An exact algebraic theory of genetic drift in finite diploid populations with random mating. Journal of Theoretical Biology, 163(3):315–331, August 7, 1993. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519383711227.

Tyvand:1994:IMC

[Tyv94] Peder A. Tyvand. Inhomogeneous Markov chain model of genetic drift in finite autotetraploid populations. Journal of Theoretical Biology, 171(4):397–407, December 21, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384712434.

Unrau:1995:CTR

[UJ95] P. Unrau and J. R. Johnson. Can tetraplex recombination models explain observations in induced mitotic gene conversion? Journal of Theoretical Biology, 177(1):73–86, November 7, 1995. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519385702263.

Ujhelyi:1996:TIS

[Ujh96] Maria Ujhelyi. Is there any intermediate stage between animal communication and language? Journal of Theoretical Biology, 180(1):71–76, May 7, 1996. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251939690079X.

Undeen:1990:PMG

[Und90] Albert H. Undeen. A proposed mechanism for the germination of microsporidian (protozoa: Microspora) spores. Journal of Theoretical Biology, 142(2):223–235, January 23, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305802231.

Uteshev:1996:PAS

[UP96] Vladimir V. Uteshev and Peter S. Pennefather. Phasic activation and state-dependent inhibition: an explicit solution for a three-state ion channel system. Journal of Theoretical Biology, 181(1):11–23, July 7, 1996. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396901101.

Ureta:1991:RIM

[Ure91] Tito Ureta. The role of isozymes in metabolite channelling. Journal of Theoretical Biology, 152(1):81–84, September 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305805168.

Uteshev:1993:BDE

[Ute93] Vladimir Uteshev. The binomial distribution and the evidence for independent action of ion channels. Journal of Theoretical Biology, 163(4):485–489, August 21, 1993. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519383711331.

Valone:1992:PEM

[Val92] Thomas J. Valone. Patch estimation via memory windows and the effect of travel time. Journal of Theoretical Biology, 157(2):243–251, July 21, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251930580623X.

Vandermeer:1990:IDI

[Van90] John Vandermeer. Indirect and difuse interactions: Complicated cycles in a population embedded in a large community. Journal of Theoretical Biology, 142(4):429–442, February 22, 1990. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305800992.

Vandermeer:1991:CGA

[Van91] John Vandermeer. Contributions to the global analysis of 3-D Lotka–Volterra equations: Dynamic boundedness and indirect interactions in the case of one predator and two prey. Journal of Theoretical Biology, 148(4):545–561, February 21, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251930580236X.

Vance:1992:OSGa

[Van92a] Richard R. Vance. Optimal somatic growth and reproduction in a limited, constant environment: Organisms with determinate growth. Journal of Theoretical Biology, 157(1):31– 50, July 7, 1992. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305807556.

Vance:1992:OSGb

[Van92b] Richard R. Vance. Optimal somatic growth and reproduction in a limited, constant environment: The general case. Journal of Theoretical Biology, 157(1):51–70, July 7, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305807568.

VanDooren:1999:EED

[Van99] Tom J. M. Van Dooren. The evolutionary ecology of dominance-recessivity. Journal of Theoretical Biology, 198 (4):519–532, June 21, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519399909293.

Varetto:1993:TAG

[Var93a] Louis Varetto. Typogenetics: An artificial genetic system. Journal of Theoretical Biology, 160(2):185–205, January 21, 1993. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519383710131.

Varmuza:1993:GIS

[Var93b] Susannah Varmuza. Gametic imprinting as a speciation mechanism in mammals. Journal of Theoretical Biology, 164(1):1–13, September 7, 1993. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519383711379.

Varetto:1998:SAL

[Var98] Louis Varetto. Studying artificial life with a molecular automaton. Journal of Theoretical Biology, 193(2):257–285, July 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519398906981.

Vasilakos:1993:END

[VB93] Konstantinon Vasilakos and Anne Beuter. Effects of noise on a delayed visual feedback system. Journal of Theoretical Biology, 165(3):389–407, December 7, 1993. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519383711963.

vanBaalen:1998:USV

[vBR98] Minus van Baalen and David A. Rand. The unit of selection in viscous populations and the evolution of altruism. Journal of Theoretical Biology, 193(4):631–648, August 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519398907305.

vanBoven:1998:ESD

[vBW98] Michiel van Boven and Franz J. Weissing. Evolution of segregation distortion: Potential for a high degree of polymorphism. Journal of Theoretical Biology, 192(2):131–142, May 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519397906287.

VanDerMeulen:1995:DMD

[VC95] Marjolein C. H. Van Der Meulen and Dennis R. Carter. Developmental mechanics determine long bone allometry. Journal of Theoretical Biology, 172(4):323–327, February 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251938570029X.

Volkert:1998:RWI

[VC98] L. Gwenn Volkert and Michael Conrad. The role of weak interactions in biological systems: the dual dynamics model. Journal of Theoretical Biology, 193(2):287–306, July 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519398907007.

vanderGlas:1992:SME

[vdGvdBB92] Hilbert W. van der Glas, Andries van der Bilt, and Frederik Bosman. A selection model to estimate the interaction between food particles and the post-canine teeth in human mastication. Journal of Theoretical Biology, 155(1):103–120, March 7, 1992. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251930580551X.

vanderHave:1996:ASE

[vdHdJ96] T. M. van der Have and G. de Jong. Adult size in ectotherms: Temperature effects on growth and differentiation. Journal of Theoretical Biology, 183(3):329–340, December 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519396902246.

vanderHoeven:1990:SER

[vdHH90] Nelly van der Hoeven and Lia Hemerik. Superparasitism as an ESS: to reject or not to reject, that is the question. Journal of Theoretical Biology, 146(4):467–482, October 21, 1990. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251930580373X.

vanderLaan:1995:SPM

[vdLLH95] Jan D. van der Laan, Ladislav Lhotka, and Pauline Hogeweg. Sequential predation: a multi-model study. Journal of Theoretical Biology, 174(2):149–167, May 21, 1995. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385700884.

vanderSteen:1990:CBS

[vdS90] Wim J. van der Steen. Concepts in biology: a survey of practical methodological principles. Journal of Theoretical Biology, 143(3):383–403, April 5, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305800359.

Verkhovod:1994:ADU

[Ver94] A. B. Verkhovod. Alphanumerical divisions of the universal genetic code: New divisions reveal new balances. Journal of Theoretical Biology, 170(3):327–330, October 7, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384711945.

Vesala:1998:CLB

[Ves98] T. Vesala. On the concept of leaf boundary layer resistance for forced convection. Journal of Theoretical Biology, 194(1):91–100, September 7, 1998. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519398907470.

Voit:1998:BMI

[VF98] Eberhard O. Voit and António E. N. Ferreira. Buffering in models of integrated biochemical systems. Journal of Theoretical Biology, 191(4):429–438, April 21, 1998. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519397906056.

VanSoest:1992:SST

[VFS92] P. J. Van Soest, J. France, and R. C. Siddons. On the steadystate turnover of compartments in the ruminant gastrointestinal tract. Journal of Theoretical Biology, 159(2):135–145, November 21, 1992. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305806988.

VanBatenburg:1995:APG

[VGP95] F. H. D. Van Batenburg, A. P. Gultyaev, and C. W. A. Pleij. An APL-programmed genetic algorithm for the prediction of RNA secondary structure. Journal of Theoretical Biology, 174(3):269–280, June 7, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519385700987.

Vork:1993:MLC

[VGV93] Michaël M. Vork, Jan F. C. Glatz, and Ger J. Van Der Vusse. On the mechanism of long chain fatty acid transport in cardiomyocytes as facilitated by cytoplasmic fatty acid-binding protein. Journal of Theoretical Biology, 160(2):207–222, January 21, 1993. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519383710143.

Varon:1990:KTP

[VH90] R. Varón and B. H. Havsteen. Kinetics of the transientphase and steady-state of the monocyclic enzyme cascades. Journal of Theoretical Biology, 144(3):397–413, June 7, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305800839.

Varon:1990:DTP

[VHG^{+}90] R. Varón, B. H. Havsteen, M. Garcia, E. Valero, and F. Garcia Cánovas. Derivation of the transient phase equations of enzyme mechanisms from those of other systems. Journal of Theoretical Biology, 143(2):251–268, March 22, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930580270X.

Varon:1992:KMA

[VHGV92] R. Varon, B. H. Havsteen, M. Garcia, and A. Vázquez. Kinetics of a model of autocatalysis, coupling of a reaction in which the enzyme acts on one of its substrates. Journal of Theoretical Biology, 154(2):261–270, January 21, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305804072.

Vits:1992:FMM

[VHSC92] Hugo Vits, Wei-Shou Hu, E. John Staba, and Todd J. Cooke. Formulation of a morphogenetic model for embryo development in dicotyledonous plants. Journal of Theoretical Biology, 157(2):221–241, July 21, 1992. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305806228.

Varon:1990:KTA

[VHV^{+}90] R. Varón, B. H. Havsteen, A. Vázquez, M. García, E. Valero, and F. García Cánovas. Kinetics of the trypsinogen activation by enterokinase and trypsin. Journal of Theoretical Biology, 145(1):123–131, July 9, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305805387.

Vickers:1991:SPT

[Vic91] G. T. Vickers. Spatial patterns and travelling waves in population genetics. Journal of Theoretical Biology, 150(3):329– 337, June 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305804321.

Vidybida:1991:SSI

[Vid91] A. K. Vidybida. Selectivity and sensitivity improvement in co-operative systems with a threshold in the presence of noise. Journal of Theoretical Biology, 152(2):159–164, September 21, 1991. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305804497.

Vinogradov:1998:BPP

[Vin98] Alexander E. Vinogradov. Buffering: a possible passivehomeostasis role for redundant DNA. Journal of Theoretical Biology, 193(2):197–199, July 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519397906299.

Visscher:1993:TAI

[Vis93] P. Kirk Visscher. A theoretical analysis of individual interests and intracolony conflict during swarming of honey bee colonies. Journal of Theoretical Biology, 165(2):191–212, November 21, 1993. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519383711859.

Vertosick:1991:ISN

[VK91] Frank T. Vertosick and Robert H. Kelly. The immune system as a neural network: a multi-epitope approach. Journal of Theoretical Biology, 150(2):225–237, May 21, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305803340.

VanDien:1998:DME

[VK98] Stephen J. Van Dien and J. D. Keasling. A dynamic model of the escherichia coli phosphate-starvation response. Journal of Theoretical Biology, 190(1):37–49, January 7, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519397905245.

Velikanov:1999:PCR

[VK99] Mikhail V. Velikanov and Raymond Kapral. Polymerase chain reaction: a Markov process approach. Journal of Theoretical Biology, 201(4):239–249, December 21, 1999. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519399910317.

Vico:1998:DSE

[VKH^{+}98] Pierre G. Vico, Soula Kyriacos, Olivier Heymans, Stéphane Louryan, and Louis Cartilier. Dynamic study of the extraembryonic vascular network of the chick embryo by fractal analysis. Journal of Theoretical Biology, 195(4):525–532, December 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519398908104.

vanLeeuwen:1991:OPO

[vL91] J. L. van Leeuwen. Optimum power output and structural design of sarcomeres. Journal of Theoretical Biology, 149(2):229–256, March 21, 1991. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305802796.

Volkov:1990:CTD

[Vol90] S. N. Volkov. Conformational transition dynamics and the mechanism of long-range effects in DNA. Journal of Theoretical Biology, 143(4):485–496, April 19, 1990. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305800256.

Vollmer:1997:USM

[Vol97] Robin T. Vollmer. Use of serum markers to measure acute myocardial infarct size: Lessons of a nonlinear dynamical model. Journal of Theoretical Biology, 187(2):195–205, July 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251939790429X.

vanOss:1996:SPF

[vOPH^{+}96] Catelijne van Oss, Alexandre V. Panfilov, Pauline Hogeweg, Florian Siegert, and Cornelis J. Weijer. Spatial pattern formation during aggregation of the slime mould Dictyostelium discoideum. Journal of Theoretical Biology, 181(3):203–213, August 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396901265.

vanOss:1997:EIN

[vOvO97] C. van Oss and A. van Ooyen. Effects of inhibition on neural network development through activity-dependent neurite outgrowth. Journal of Theoretical Biology, 185(2):263–280, March 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396903616.

vanOoyen:1994:ADO

[vOvP94] A. van Ooyen and J. van Pelt. Activity-dependent outgrowth of neurons and overshoot phenomena in developing neural networks. Journal of Theoretical Biology, 167(1):27–43, March 7, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519384710472.

vanOoyen:1996:CPB

[vOvP96] A. van Ooyen and J. van Pelt. Complex periodic behaviour in a neural network model with activity-dependent neurite outgrowth. Journal of Theoretical Biology, 179(3):229–242, April 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396900636.

vanOoyen:1999:PMI

[vOW99] Arjen van Ooyen and David J. Willshaw. Poly- and mononeuronal innervation in a model for the development of neuromuscular connections. Journal of Theoretical Biology, 196(4):495–511, February 21, 1999. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519398908499.

Varma:1993:MCEa

[VP93a] Amit Varma and Bernhard O. Palsson. Metabolic capabilities of Escherichia coli: I. Synthesis of biosynthetic precursors and cofactors. Journal of Theoretical Biology, 165(4):477–502, December 21, 1993. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519383712026.

Varma:1993:MCEb

[VP93b] Amit Varma and Bernhard O. Palsson. Metabolic capabilities of Escherichia coli II. optimal growth patterns. Journal of Theoretical Biology, 165(4):503–522, December 21, 1993. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519383712038.

vanPelt:1997:EPD

[vP97] Jaap van Pelt. Effect of pruning on dendritic tree topology. Journal of Theoretical Biology, 186(1):17–32, May 7, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519396903410.

Valdes-perez:1995:DMS

[VpM95] Raúl E. Valdés-pérez and Jonathan S. Minden. Drosophila melanogaster syncytial nuclear divisions are patterned: timelapse images, hypothesis and computational evidence. Journal of Theoretical Biology, 175(4):525–532, August 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385701609.

Vinet:1990:MSS

[VR90] Alain Vinet and Fernand A. Roberge. A model study of stability and oscillations in the myocardial cell membrane. Journal of Theoretical Biology, 147(3):377–412, December 7, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305804953.

Vinet:1994:AID

[VR94a] A. Vinet and F. A. Roberge. Analysis of an iterative difference equation model of the cardiac cell membrane. Journal of Theoretical Biology, 170(2):201–214, September 21, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519384711799.

Vinet:1994:ERI

[VR94b] A. Vinet and F. A. Roberge. Excitability and repolarization in an ionic model of the cardiac cell membrane. Journal of Theoretical Biology, 170(2):183–199, September 21, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519384711787.

Vrba:1998:MGM

[Vrb98] E. S. Vrba. Multiphasic growth models and the evolution of prolonged growth exemplified by human brain evolution.

Journal of Theoretical Biology, 190(3):227–239, February 7, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251939790549X.

vanRiel:1998:SMP

[vRGTV98] N. A. W. van Riel, M. L. F. Giuseppin, E. G. TerSchure, and C. T. Verrips. A structured, minimal parameter model of the central nitrogen metabolism in saccharomyces cerevisiae: the prediction of the behaviour of mutants. Journal of Theoretical Biology, 191(4):397–414, April 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519397906007.

Vega-Redondo:1993:GTM

[VRH93] Fernando Vega-Redondo and Oren Hasson. A game-theoretic model of predator–prey signaling. Journal of Theoretical Biology, 162(3):309–319, June 7, 1993. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519383710891.

Varela:1990:DCI

[VS90] Francisco J. Varela and John Stewart. Dynamics of a class of immune networks I. Global stability of idiotype interactions. Journal of Theoretical Biology, 144(1):93–101, May 10, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305803017.

Varghese:1997:CPE

[VS97] Anthony Varghese and George R. Sell. A conservation principle and its effect on the formulation of Na-Ca exchanger current in cardiac cells. Journal of Theoretical Biology, 189(1):33–40, November 7, 1997. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519397904872.

Vasiev:1997:HMD

[VSW97] Bakhtier Vasiev, Florian Siegert, and Cornelis J. Weller. A hydrodynamic model for Dictyostelium discoideum mound formation. Journal of Theoretical Biology, 184(4):441–450, February 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396902829.

Vincent:1996:UEM

[VV96] Tania L. S. Vincent and Thomas L. Vincent. Using the ESS maximum principle to explore root-shoot allocation, competition and coexistence. Journal of Theoretical Biology, 180(2):111–120, May 21, 1996. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519396900867.

VanOoyen:1995:IAD

[VVC95] A. Van Ooyen, J. Van Pelt, and M. A. Corner. Implications of activity dependent neurite outgrowth for neuronal morphology and network development. Journal of Theoretical Biology, 172(1):63–82, January 7, 1995. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385700057.

Vasieva:1994:MDD

[VVKZ94] O. O. Vasieva, B. N. Vasiev, V. A. Karpov, and A. N. Zaikin. A model of Dictyostelium discoideum aggregation. Journal of Theoretical Biology, 171(4):361–367, December 21, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519384712409.

vanVeen:1992:MOB

[vVvP92] M. van Veen and J. van Pelt. A model for outgrowth of branching neurites. Journal of Theoretical Biology, 159(1):1–23, November 7, 1992. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305807647.

Varghese:1994:DAP

[VW94a] A. Varghese and R. L. Winslow. Dynamics of abnormal pacemaking activity in cardiac Purkinje fibers. Journal of Theoretical Biology, 168(4):407–420, June 21, 1994. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384711210.

vonWright:1994:MM

[vW94b] Georg Henrik von Wright. On mind and matter. Journal of Theoretical Biology, 171(1):101–110, July 1994. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384712161.

vonWagenheim:1998:CCP

[vWP98] Karl-Hartmut von Wagenheim and Hans-Peter Peterson. Control of cell proliferation by progress in differentiation: Clues to mechanisms of aging, cancer causation and therapy. Journal of Theoretical Biology, 193(4):663–678, August 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519398907317.

Wright:1994:CMD

[WA94] Barbara E. Wright and Kathy R. Albe. Carbohydrate metabolism in Dictyostelium discoideum: I. Model construction. Journal of Theoretical Biology, 169(3):231–241, August 7, 1994. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519384711441.

Wachtershauser:1997:OLM

[Wäc97] Günter Wächtershäuser. The origin of life and its methodological challenge. Journal of Theoretical Biology, 187(4):483–494, August 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396903835.

Wagner:1993:PEP

[Wag93] Richard H. Wagner. The pursuit of extra-pair copulations by female birds: a new hypothesis of colony formation. Journal of Theoretical Biology, 163(3):333–346, August 7, 1993. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519383711239.

Winzor:1991:PLB

[WAJ91] Donald J. Winzor, Paul-Michael Agapow, and Craig M. Jackson. Preferential ligand binding to multi-state acceptor systems: Comparisons of the calcium-binding and dimerization characteristics of prothrombin and fragment 1. Journal of Theoretical Biology, 153(3):385–399, December 7, 1991. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305805776.

Walter:1991:WRP

[Wal91] G. H. Walter. What is resource partitioning? Journal of Theoretical Biology, 150(2):137–143, May 21, 1991. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305803273.

Walsh:1996:ARS

[Wal96] Peter D. Walsh. Area-restricted search and the scale dependence of path quality discrimination. Journal of Theoretical Biology, 183(4):351–361, December 21, 1996. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396902271.

Walker:1998:WVM

[Wal98] Michael M. Walker. On a wing and a vector: a model for magnetic navigation by homing pigeons. Journal of Theoretical Biology, 192(3):341–349, June 7, 1998. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519398906531.

Walker:1999:MPD

[Wal99a] Michael M. Walker. Magnetic position determination by homing pigeons. Journal of Theoretical Biology, 197(2):271–276, March 21, 1999. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251939890875X.

Wallraff:1999:MMH

[Wal99b] Hans G. Wallraff. The magnetic map of homing pigeons: an evergreen phantom. Journal of Theoretical Biology, 197(2):265–269, March 21, 1999. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519398908748.

Wang:1990:KSS

[Wan90a] Zhi-Xin Wang. Kinetics of suicide substrates. Journal of Theoretical Biology, 147(4):497–508, December 21, 1990. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305802619.

Wang:1990:SASa

[Wan90b] Zhi-Xin Wang. Some applications of statistical mechanics in enzymology 1. Elementary principle. Journal of Theoretical Biology, 143(4):445–453, April 19, 1990. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305800219.

Wang:1990:SASc

[Wan90c] Zhi-Xin Wang. Some applications of statistical mechanics in enzymology 3. Effects of nonspecific inhibitors and activators on enzyme reaction. Journal of Theoretical Biology, 143(4):465–472, April 19, 1990. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305800232.

Wang:1990:SASd

[Wan90d] Zhi-Xin Wang. Some applications of statistical mechanics in enzymology 4. Effects of pH on enzyme. Journal of Theoretical Biology, 143(4):473–484, April 19, 1990. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305800244.

Wang:1990:TTP

[Wan90e] Zhi-Xin Wang. Two theoretical problems concerning the irreversible modification kinetics of enzyme activity. Journal of Theoretical Biology, 142(4):551–563, February 22, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305801080.

Wang:1991:TCT

[Wan91] Zhi-Xin Wang. Theoretical considerations of the tsou plot. Journal of Theoretical Biology, 150(4):437–450, June 21, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305804382.

Wang:1999:PPP

[WAO^{+}99] Jin-Sheng Wang, Tetsuya Araki, Takahira Ogawa, Miho Sakai, Masayoshi Matsuoka, and Hideo Fukuda. Prediction of photosynthetic production rate of ethylene using a recombinant cyanobacterium. Journal of Theoretical Biology, 196(1):9–17, January 7, 1999. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251939890813X.

Watmough:1997:GMP

[Wat97] James Watmough. A general model of pheromone transmission within honey bee hives. Journal of Theoretical Biology, 189(2):159–170, November 21, 1997. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519397905063.

Wagner:1999:SFS

[WB99] H. Wagner and R. Blickhan. Stabilizing function of skeletal muscles: an analytical investigation. Journal of Theoretical Biology, 199(2):163–179, July 21, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519399909499.

Wohlfahrt:1999:URB

[WBC99] Georg Wohlfahrt, Michael Bahn, and Alexander Cernusca. The use of the ratio between the photosynthesis parameters P_{ml} and V_{cmax} for scaling up photosynthesis of C_{3} plants from leaves to canopies: a critical examination of different modelling approaches. Journal of Theoretical Biology, 200(2):163–181, September 21, 1999. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519399909852.

Wiehe:1995:EPR

[WBS95] Thomas Wiehe, Ellen Baake, and Peter Schuster. Error propagation in reproduction of diploid organisms: a case study on single peaked landscapes. Journal of Theoretical Biology, 177(1):1–15, November 7, 1995. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385702202.

Waelbroeck:1997:SHR

[WBS97] Magali Waelbroeck, Lamyae Boufrahi, and Stéphane Swillens. Seven helix receptors are enzymes catalysing G protein activation. What is the agonist K act? Journal of Theoretical Biology, 187(1):15–37, July 7, 1997. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519397904112.

Whittaker:1994:MTT

[WC94a] J. C. Whittaker and C. Cannings. Multiple trial two strategy conflicts. Journal of Theoretical Biology, 169(2):113–124, July 21, 1994. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519384711337.

Whittaker:1994:RAP

[WC94b] J. C. Whittaker and C. Cannings. A resource allocation problem. Journal of Theoretical Biology, 167(4):397–405, April 21, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519384710782.

Wagg:1995:SSR

[WC95a] Jonathan Wagg and Brian J. Chapman. Steady-state rate analysis: application to biological transport. Journal of Theoretical Biology, 174(1):45–59, May 7, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519385700781.

Watmough:1995:SOT

[WC95b] James Watmough and Scott Camazine. Self-organized thermoregulation of honeybee clusters. Journal of Theoretical Biology, 176(3):391–402, October 7, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251938570207X.

Ward:1991:DTP

[WCPR91] D. J. Ward, Y. Chen, E. Platt, and B. Robson. Development and testing of protocols for computer-aided design of peptide drugs, using oxytocin. Journal of Theoretical Biology, 148(2):193–227, January 21, 1991. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305803418.

Wagg:1995:FRA

[WCW95] Jonathan Wagg, Brian J. Chapman, and Susan A. Wood. A flux ratio analysis of the post-albers mechanism of the sodium pump. Journal of Theoretical Biology, 174(1):73–85, May 7, 1995. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251938570080X.

Wallet:1997:MCS

[WD97] France Wallet and Christophe Dussert. Multifactorial comparative study of spatial point pattern analysis methods. Journal of Theoretical Biology, 187(3):437–447, August 7, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519397904458.

Weisbuch:1990:LMI

[WDP90] Gérard Weisbuch, Rob J. De Boer, and Alan S. Perelson. Localized memories in idiotypic networks. Journal of Theoretical Biology, 146(4):483–499, October 21, 1990. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305803741.

Wein:1998:MAA

[WDP98] Lawrence M. Wein, Rebecca M. D’Amato, and Alan S. Perelson. Mathematical analysis of antiretroviral therapy aimed at HIV-1 eradication or maintenance of low viral loads. Journal of Theoretical Biology, 192(1):81–98, May 7, 1998. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519397906226.

Wynne-Edwards:1993:RGS

[WE93] V. C. Wynne-Edwards. A rationale for group selection. Journal of Theoretical Biology, 162(1):1–22, May 7, 1993. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519383710738.

Wolf:1998:RBG

[WE98] Denise M. Wolf and Frank H. Eeckman. On the relationship between genomic regulatory element organization and gene regulatory dynamics. Journal of Theoretical Biology, 195(2):167–186, November 21, 1998. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519398907901.

Weber:1999:BID

[Web99] Thomas P. Weber. Blissful ignorance? departure rules for migrants in a spatially heterogeneous environment. Journal of Theoretical Biology, 199(4):415–424, August 21, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519399909682.

Wedekind:1994:HOS

[Wed94] Claus Wedekind. Handicaps not obligatory in sexual selection for resistance genes. Journal of Theoretical Biology, 170(1):57–62, September 7, 1994. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519384711672.

Weinand:1990:SMA

[Wei90a] Richard G. Weinand. Somatic mutation, affinity maturation and the antibody repertoire: a computer model. Journal of Theoretical Biology, 143(3):343–382, April 5, 1990. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305800347.

Weisbuch:1990:SSA

[Wei90b] Gérard Weisbuch. A shape space approach to the dynamics of the immune system. Journal of Theoretical Biology, 143(4):507–522, April 19, 1990. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251930580027X.

Weiss:1997:NIT

[Wei97] Michael Weiss. A note on the interpretation of tracer dispersion in the liver. Journal of Theoretical Biology, 184(1):1–6, January 7, 1997. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396902659.

Watmough:1995:MFT

[WEK95] James Watmough and Leah Edelstein-Keshet. Modelling the formation of trail networks by foraging ants. Journal of Theoretical Biology, 176(3):357–371, October 7, 1995. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385702056.

Westerhoff:1991:ECA

[Wes91] Hans V. Westerhoff. Energetics and control aspects of channelling. Journal of Theoretical Biology, 152(1):123–130, September 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305805259.

West:1993:MAG

[Wes93] P. W. West. Model of above-ground assimilate partitioning and growth of individual trees in even-aged forest monoculture. Journal of Theoretical Biology, 161(3):369–394, April 7, 1993. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519383710623.

Wold:1992:FCM

[WG92] Matthew P. Wold and R. Igor Gamow. Fiber-composite model for helical growth in the phycomyces cell wall. Journal of Theoretical Biology, 159(1):39–51, November 7, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305807660.

West:1997:SRS

[WG97a] S. A. West and H. C. J. Godfray. Sex ratio strategies after perturbation of the stable age distribution. Journal of Theoretical Biology, 186(2):213–221, May 21, 1997. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396903690.

White:1997:RCH

[WG97b] K. A. J. White and B. T. Grenfell. Regulation of complex host dynamics by a macroparasite. Journal of Theoretical Biology, 186(1):81–91, May 7, 1997. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519396903446.

Whiteley:1998:MEM

[WGH98] J. P. Whiteley, D. J. Gavaghan, and C. E. W. Hahn. A mathematical evaluation of the multiple breath nitrogen washout (MBNW) technique and the multiple inert gas elimination technique (MIGET). Journal of Theoretical Biology, 194(4):517–539, October 21, 1998. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251939890772X.

Whiteley:1999:EWV

[WGH99] J. P. Whiteley, D. J. Gavaghan, and C. E. W. Hahn. The effect of the width of the ventilation-perfusion distribution on arterial blood oxygen content. Journal of Theoretical Biology, 201(4):271–279, December 21, 1999. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519399910329.

Witham:1992:CMG

[WH92] Francis H. Witham and Lawrence B. Hendry. Computer modeling of gibberellin-DNA binding. Journal of Theoretical Biology, 155(1):55–67, March 7, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251930580548X.

Wall:1994:ACH

[WH94a] Richard Wall and Julia Howard. Autosterilization for control of the housefly musca domestica. Journal of Theoretical Biology, 171(4):431–437, December 21, 1994. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384712471.

Weissing:1994:GCL

[WH94b] Franz J. Weissing and Jef Huisman. Growth and competition in a light gradient. Journal of Theoretical Biology, 168(3):323– 336, June 7, 1994. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519384711131.

Woodcock:1996:PEM

[WH96] Glenn Woodcock and Paul G. Higgs. Population evolution on a multiplicative single-peak fitness landscape. Journal of Theoretical Biology, 179(1):61–73, March 7, 1996. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396900491.

Weber:1997:GMT

[WH97] Thomas P. Weber and Alasdair I. Houston. A general model for time-minimising avian migration. Journal of Theoretical Biology, 185(4):447–458, April 21, 1997. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251939690329X.

Weber:1998:LHI

[WH98a] Thomas P. Weber and Alasdair I. Houston. A lifehistory invariant for migration. Journal of Theoretical Biology, 192(2):261–263, May 21, 1998. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519398906555.

Weiss:1998:CPS

[WH98b] Olaf Weiss and Hanspeter Herzel. Correlations in protein sequences and property codes. Journal of Theoretical Biology, 190(4):341–353, February 21, 1998. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519397905609.

Whitehead:1990:RRM

[Whi90] Hal Whitehead. Rules for roving males. Journal of Theoretical Biology, 145(3):355–368, August 9, 1990. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305801158.

White:1991:TAC

[Whi91] R. A. White. A theory for analysis of cell populations with non-cycling S phase cells. Journal of Theoretical Biology, 150(2):201–214, May 21, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305803327.

Whitehead:1994:DCB

[Whi94] Hal Whitehead. Delayed competitive breeding in roving males. Journal of Theoretical Biology, 166(2):127–133, January 21, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519384710113.

Whitaker:1997:RAV

[Whi97] Rupert E. D. Whitaker. Re-assessing the virological approach to HIV pathogenesis: Can it explain AIDS as an immunological disease? Journal of Theoretical Biology, 187(1):45– 56, July 7, 1997. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519397904161.

Whipple:1998:PBN

[Whi98] S. Whipple. Path-based network unfolding: a solution for the problem of mixed trophic and non-trophic processes in trophic dynamic analysis. Journal of Theoretical Biology, 190(3):263–276, February 7, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519397905518.

Watts:1993:MHL

[WHL93] Peter Watts, Stephen Hansen, and D. M. Lavigne. Models of heat loss by marine mammals: Thermoregulation below the zone of irrelevance. Journal of Theoretical Biology, 163(4):505–525, August 21, 1993. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519383711355.

Wick:1999:DCC

[Wic99] David Wick. The disappearing CD4 + T cells in HIV infection: a case of over-stimulation? Journal of Theoretical Biology, 197(4):507–516, April 21, 1999. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519398908918.

Wiegmann:1990:APE

[Wie90] Daniel D. Wiegmann. On assessing the potential for evolutionary change due to male-male competition and female choice in territorial species. Journal of Theoretical Biology, 144(2):203– 208, May 14, 1990. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305803194.

Wiens:1996:AMC

[Wie96] Darrell J. Wiens. An alternative model for cell sheet migration on fibronectin during heart formation. Journal of Theoretical Biology, 179(1):33–39, March 7, 1996. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396900454.

Williams:1992:MSC

[Wil92a] Brian Williams. The measurement of “sinuosity” in correlated random walks. Journal of Theoretical Biology, 155(4):437–442, April 21, 1992. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305806289.

Wills:1992:PPP

[Wil92b] Peter R. Wills. Potential pseudoknots in the PrP-encoding mRNA. Journal of Theoretical Biology, 159(4):523–527, December 21, 1992. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305806964.

Williams:1993:OCS

[Wil93a] J. M. Williams. Objective comparisons of song syllables: a dynamic programming approach. Journal of Theoretical Biology, 161(3):317–328, April 7, 1993. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519383710581.

Wills:1993:SOG

[Wil93b] Peter R. Wills. Self-organization of genetic coding. Journal of Theoretical Biology, 162(3):267–287, June 7, 1993. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519383710878.

Williams:1994:MTS

[Wil94] Brian Williams. Models of trap seeking by tsetse flies: Anemotaxis, klinokinesis and edge detection. Journal of Theoretical Biology, 168(1):105–115, May 7, 1994. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384710915.

Williamson:1995:ITA

[Wil95] Rufus M. Williamson. Information theory analysis of the relationship between primary sequence structure and ligand recognition among a class of facilitated transporters. Journal of Theoretical Biology, 174(2):179–188, May 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385700902.

Willson:1998:MIP

[Wil98] Stephen J. Willson. Measuring inconsistency in phylogenetic trees. Journal of Theoretical Biology, 190(1):15–36, January 7, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519397905270.

Wilson:1999:SDH

[Wil99] Hugh R. Wilson. Simplified dynamics of human and mammalian neocortical neurons. Journal of Theoretical Biology, 200(4):375–388, October 21, 1999. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519399910020.

Witting:1995:BMA

[Wit95] L. Witting. The body mass allometries as evolutionarily determined by the foraging of mobile organisms. Journal of Theoretical Biology, 177(2):129–137, November 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385702317.

Watve:1997:EDH

[WJ97] Milind G. Watve and Maithili M. Jog. Epidemic diseases and host clustering: An optimum cluster size ensures maximum survival. Journal of Theoretical Biology, 184(2):165–169, January 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396902672.

Weisser:1997:EPD

[WJH97] W. W. Weisser, V. A. A. Jansen, and M. P. Hassell. The effects of a pool of dispersers on host-parasitoid systems. Journal of Theoretical Biology, 189(4):413–425, December 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519397905294.

Wang:1990:SASb

[WK90] Zhi-Xin Wang and Hiroshi Kihara. Some applications of statistical mechanics in enzymology 2. Statistical mechanical explanation on allosteric enzyme models. Journal of Theoretical Biology, 143(4):455–464, April 19, 1990. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305800220.

Westerhoff:1996:WBK

[WK96] Hans V. Westerhoff and Douglas B. Kell. What biotechnologists knew all along … ? Journal of Theoretical Biology, 182(3):411–420, October 7, 1996. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519396901812.

Wang:1992:SRB

[WKSH92] K. Wang, R. I. Kitney, J. W. Seakins, and M. Hjelm. Study of the relationship between estimates of enzyme kinetic parameters. Journal of Theoretical Biology, 155(4):485–495, April 21, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305806319.

Warburton:1991:TDM

[WL91] Kevin Warburton and John Lazarus. Tendency-distance models of social cohesion in animal groups. Journal of Theoretical Biology, 150(4):473–488, June 21, 1991. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305804412.

Wilby:1994:DMS

[WLH94] Mark R. Wilby, Max J. Labs, and Arun V. Holden. Dynamical model of signal propagation in the heart. Journal of Theoretical Biology, 168(4):399–406, June 21, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384711209.

Wodarz:1999:DMC

[WLJN99] Dominik Wodarz, Alun L. Lloyd, Vincent A. A. Jansen, and Martin A. Nowak. Dynamics of macrophage and T cell infection by HIV. Journal of Theoretical Biology, 196(1):101–113, January 7, 1999. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519398908165.

White:1996:MWP

[WLM96] K. A. J. White, M. A. Lewis, and J. D. Murray. A model for Wolf-Pack territory formation and maintenance. Journal of Theoretical Biology, 178(1):29–43, January 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396900041.

Wright:1996:MMR

[WLR96] Barbara E. Wright, Angelika Longacre, and Jacqueline Reimers. Models of metabolism in rhizopus oryzae. Journal of Theoretical Biology, 182(3):453–457, October 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396901861.

Welch:1991:MCC

[WM91] G. R. Welch and P. R. Marmillot. Metabolic “channeling” and cellular physiology. Journal of Theoretical Biology, 152(1):29–33, September 7, 1991. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251930580503X.

Wikan:1995:PAS

[WM95] Arild Wikan and Einar Mjølhus. Periodicity of 4 in agestructured population models with density dependence. Journal of Theoretical Biology, 173(2):109–119, March 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385700483.

Wiegmann:1998:FSS

[WM98] Daniel D. Wiegmann and Kajal Makhopadhyay. The fixed sample search rule and use of an indicator character to evaluate mate quality. Journal of Theoretical Biology, 193(4):709–715, August 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519398907378.

Weiss:1996:CTCa

[WMLK96a] Jack M. Weiss, Paul H. Morgan, Michael W. Lutz, and Terry P. Kenakin. The cubic ternary complex receptoroccupancy model I. Model description. Journal of Theoretical Biology, 178(2):151–167, January 21, 1996. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396900144.

Weiss:1996:CTCb

[WMLK96b] Jack M. Weiss, Paul H. Morgan, Michael W. Lutz, and Terry P. Kenakin. The cubic ternary complex receptoroccupancy model II. understanding apparent affinity. Journal of Theoretical Biology, 178(2):169–182, January 21, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396900156.

Weiss:1996:CTCc

[WMLK96c] Jack M. Weiss, Paul H. Morgan, Michael W. Lutz, and Terry P. Kenakin. The cubic ternary complex receptoroccupancy model III. resurrecting efficacy. Journal of Theoretical Biology, 181(4):381–397, August 21, 1996. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396901393.

Wahl:1999:CPDa

[WN99a] Lindi M. Wahl and Martin A. Nowak. The continuous Prisoner’s dilemma: I. Linear reactive strategies. Journal of Theoretical Biology, 200(3):307–321, October 7, 1999. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519399909967.

Wahl:1999:CPDb

[WN99b] Lindi M. Wahl and Martin A. Nowak. The continuous Prisoner’s dilemma: II. linear reactive strategies with noise.

Journal of Theoretical Biology, 200(3):323–338, October 7, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519399909979.

Werdelin:1999:EST

[WN99c] Lars Werdelin and Åsa Nilsonne. The evolution of the scrotum and testicular descent in mammals: a phylogenetic view. Journal of Theoretical Biology, 196(1):61–72, January 7, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519398908219.

Wakano:1998:DMO

[WNY98] J. Y. Wakano, K. Nakata, and N. Yamamura. Dynamic model of optimal age polyethism in social insects under stable and fluctuating environments. Journal of Theoretical Biology, 193(1):153–165, July 7, 1998. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251939890697X.

Wendling:1999:SRC

[WOI99] Sylvie Wendling, Christian Oddou, and Daniel Isabey. Stiffening response of a cellular tensegrity model. Journal of Theoretical Biology, 196(3):309–325, February 7, 1999. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519398908414.

Wojciechowski:1992:EAM

[Woj92] A. P. Wojciechowski. Evolutionary aspects of mammalian secondary sexual characteristics. Journal of Theoretical Biology, 155(2):271–272, March 21, 1992. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305805983.

Wolfe:1991:MDR

[Wol91] Kenneth H. Wolfe. Mammalian DNA replication: Mutation biases and the mutation rate. Journal of Theoretical Biology, 149(4):441–451, April 21, 1991. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930580092X.

Wong:1990:DSG

[Won90] Albert J. Wong. Development of a spin-glass model of prebiotic evolution: Environmental effects on ensembles of genetic polymers. Journal of Theoretical Biology, 146(4):523–543, October 21, 1990. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305803777.

Wong:1994:MCE

[Won94] Pierre Wong. Mechanism of control of erythrocyte shape: a possible relationship to band 3. Journal of Theoretical Biology, 171(2):197–205, November 21, 1994. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384712239.

Wong:1999:BES

[Won99] Pierre Wong. A basis of echinocytosis and stomatocytosis in the disc-sphere transformations of the erythrocyte. Journal of Theoretical Biology, 196(3):343–361, February 7, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519398908451.

Wood:1993:HEL

[Woo93] Simon N. Wood. How to estimate life history stage durations from stage structured population data. Journal of Theoretical Biology, 163(1):61–76, July 1, 1993. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519383711070.

Wood:1997:FMR

[Woo97] P. N. Wood. A functional model for the ribosome. Journal of Theoretical Biology, 185(1):97–118, March 7, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396902933.

Wood:1999:AFD

[Woo99] N. B. Wood. Aspects of fluid dynamics applied to the larger arteries. Journal of Theoretical Biology, 199(2):137–161, July 21, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519399909530.

Woldenberg:1993:MGD

[WOQP93] M. J. Woldenberg, M. P. O’Neill, L. J. Quackenbush, and R. J. Pentney. Models for growth, decline and regrowth of the dendrites of rat Purkinje cells induced from magnitude and linklength analysis. Journal of Theoretical Biology, 162(4):403– 429, June 21, 1993. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519383710970.

Worden:1995:SLE

[Wor95] R. P. Worden. A speed limit for evolution. Journal of Theoretical Biology, 176(1):137–152, September 7, 1995. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251938570183X.

Whipple:1993:PNP

[WP93] Stuart J. Whipple and Bernard C. Patten. The problem of nontrophic processes in trophic ecology: Toward a network unfolding solution. Journal of Theoretical Biology, 163(3):393–411, August 7, 1993. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519383711276.

Weerappuli:1990:ECC

[WPP90] D. P. V. Weerappuli, R. N. Pittman, and A. S. Popel. Effect of convection in capillaries on oxygen removal from arterioles in striated muscle. Journal of Theoretical Biology, 147(2):275–288, November 21, 1990. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305800578.

Whitaker:1993:PCE

[WR93] L. Whitaker and A. M. Renton. On the plausibility of the clonal expansion theory of the immune system in terms of the combinatorial possibilities of amino-acids in antigen and selftolerance. Journal of Theoretical Biology, 164(4):531–536, October 21, 1993. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519383711719.

Weller:1994:PSP

[WR94] Klausdieter Weller and Rolf-Dieter Recknagel. Promoter strength prediction based on occurrence frequencies of consensus patterns. Journal of Theoretical Biology, 171(4):355–359, December 21, 1994. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519384712392.

Weinand:1999:MIE

[WRP99] Richard G. Weinand, Allen J. Rosenspire, and Howard R. Petty. Modeling the influence of ectodomain affinities on the spatial distribution of membrane receptors. Journal of Theoretical Biology, 197(2):217–225, March 21, 1999. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519398908694.

Williams:1990:MBS

[WS90] J. M. Williams and P. J. B. Slater. Modelling bird song dialects: the influence of repertoire size and numbers of neighbours. Journal of Theoretical Biology, 145(4):487–496, August 23, 1990. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305804837.

Weinberger:1993:WSF

[WS93] Edward D. Weinberger and Peter F. Stadler. Why some fitness landscapes are fractal. Journal of Theoretical Biology, 163(2):255–275, July 21, 1993. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519383711203.

Wade:1994:EPS

[WS94] Michael J. Wade and Lori Stevens. The effect of population subdivision on the rate of spread of parasite-mediated cytoplasmic incompatibility. Journal of Theoretical Biology, 167(1):81– 87, March 7, 1994. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519384710526.

Witztum:1995:MSE

[WS95] A. Witztum and K. Schulgasser. The mechanics of seed expulsion in acanthaceae. Journal of Theoretical Biology, 176(4):531–542, October 21, 1995. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519385702196.

Wilkins:1998:ISW

[WS98] Matthew Wilkins and James Sneyd. Intercellular spiral waves of calcium. Journal of Theoretical Biology, 191(3):299–308, April 7, 1998. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519397905853.

Webb:1999:MMT

[WSF99] S. D. Webb, J. A. Sherratt, and R. G. Fish. Mathematical modelling of tumor acidity: Regulation of intracellular pH. Journal of Theoretical Biology, 196(2):237–250, January 21, 1999. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519398908360.

Wang:1990:AMD

[WT90] Zhi-Xin Wang and Chen-Lu Tsou. An alternative method for determining inhibition rate constants by following the substrate reaction. Journal of Theoretical Biology, 142(4):531–549, February 22, 1990. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305801079.

Wiener:1994:MVE

[WT94] Pamela Wiener and Shripad Tuljapurkar. Migration in variable environments: Exploring life-history evolution using structured population models. Journal of Theoretical Biology, 166(1):75– 90, January 7, 1994. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251938471006X.

Wyss:1992:MCA

[WW92] M. Wyss and T. Wallimann. Metabolite channelling in aerobic energy metabolism. Journal of Theoretical Biology, 158(1):129–132, September 7, 1992. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305806502.

Weissbluth:1994:SID

[WW94] Linda Weissbluth and Marc Weissbluth. Sudden infant death syndrome: a genetically determined impaired maturation of the photoneuroendocrine system. a unifying hypothesis. Journal of Theoretical Biology, 167(1):13–25, March 7, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519384710460.

Wilkins:1997:CSP

[WW97] Marc R. Wilkins and Keith L. Williams. Cross-species protein identification using amino acid composition, peptide mass fingerprinting, isoelectric point and molecular mass: a theoretical evaluation. Journal of Theoretical Biology, 186(1):7– 15, May 7, 1997. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251939690346X.

Wallace:1998:ITS

[WW98] Rodrick Wallace and Robert G. Wallace. Information theory, scaling laws and the thermodynamics of evolution. Journal of Theoretical Biology, 192(4):545–559, June 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519398906804.

Wells:1990:IOA

[WWC90] Harrington Wells, Patrick H. Wells, and Peyton Cook. The importance of overwinter aggregation for reproductive success of monarch butterflies (Danaus plexippus L.). Journal of Theoretical Biology, 147(1):115–131, November 7, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305802553.

Wyszomirski:1992:DDS

[Wys92] Tomasz Wyszomirski. Detecting and displaying size bimodality: Kurtosis, skewness and bimodalizable distributions. Journal of Theoretical Biology, 158(1):109–128, September 7, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305806496.

Wang:1996:SIE

[WZ96] Ji-Hua Wang and Chun-Ting Zhang. Study on the isentropic equations of nucleotide sequences and their application. Journal of Theoretical Biology, 181(3):197–202, August 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396901253.

Weisbuch:1993:THR

[WZN93] Gérard Weisbuch, Rita Maria Zorzenon Dos Santos, and Avidan U. Neumann. Tolerance to hormones and receptors in an idiotypic network model. Journal of Theoretical Biology, 163(2):237–253, July 21, 1993. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519383711197.

Wein:1997:DMT

[WZN97] Lawrence M. Wein, Stefanos A. Zenios, and Martin A. Nowak. Dynamic multidrug therapies for HIV: a control theoretic approach. Journal of Theoretical Biology, 185(1):15–29, March 7, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519396902532.

Xu:1994:FGS

[XCC94] Jun Xu, Yang Chao, and Runsheng Chen. Fractal geometry study of DNA binding proteins. Journal of Theoretical Biology, 171(3):239–249, December 7, 1994. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384712288.

Xiao:1995:FDE

[XCS^{+}95] Yi Xiao, Runsheng Chen, Ruqun Shen, Jian Sun, and Jun Xu. Fractal dimension of exon and intron sequences. Journal of Theoretical Biology, 175(1):23–26, July 7, 1995. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385701178.

Yasuda:1997:ABF

[YA97] Kunio Yasuda and Akira Azuma. The autorotation boundary in the flight of samaras. Journal of Theoretical Biology, 185(3):313–320, April 7, 1997. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519396902994.

Yamamura:1993:DEC

[Yam93] Norio Yamamura. Different evolutionary conditions for worker and soldier castes: Genetic systems explaining caste distribution among eusocial insects. Journal of Theoretical Biology, 161(1):111–117, March 7, 1993. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251938371043X.

Yamauchi:1999:ECS

[Yam99] Atsushi Yamauchi. Evolution of cyclic sexual reproduction under host-parasite interactions. Journal of Theoretical Biology, 201(4):281–291, December 21, 1999. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519399910342.

Yariv:1996:CEC

[Yar96] Joseph Yariv. Circumstantial evidence for cytochrome b_{1} involvement in the functioning of lac-permease in respiring Escherichia coli. Journal of Theoretical Biology, 182(4):459–462, October 21, 1996. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396901873.

Yeakley:1991:OLA

[YC91] J. Alan Yeakley and William G. Cale. Organizational levels analysis: a key to understanding processes in natural systems. Journal of Theoretical Biology, 149(2):203–216, March 21, 1991. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305802772.

Yang:1999:AIS

[YC99] Hyun Mo Yang and Francisco Antonio Bezerra Coutinho. Acquired immunity of a schistosomiasis transmission modelanalysis of the stabilizing effects. Journal of Theoretical Biology, 196(4):473–482, February 21, 1999. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519398908372.

Yang:1997:AIS

[YCM97] Hyun Mo Yang, Francisco Antonio Bezerra Coutinho, and Eduardo Massad. Acquired immunity on a schistosomiasis transmission model — fitting the data. Journal of Theoretical Biology, 188(4):495–506, October 21, 1997. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519397904793.

Yee:1991:RAA

[Yee91]  Eugene Yee.  Reconstruction of the antibody affinity

distribution from experimental binding data by a minimum cross-entropy procedure. Journal of Theoretical Biology, 153(2):205–227, November 21, 1991. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305804230.

Yasin:1993:IPL

[YFG^{+}93] S. Yasin, M. Friedman, S. Goshen, A. Rabinovitch, and R. Thieberger. Intermittency and phase locking of the Bonhoeffer van der Pol model. Journal of Theoretical Biology, 160(2):179–184, January 21, 1993. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251938371012X.

Yuen:1994:CDS

[YG94] Ita S. Yuen and Richard H. Gomer. Cell density-sensing in Dictyostelium by means of the accumulation rate, diffusion coefficient and activity threshold of a protein secreted by starved cells. Journal of Theoretical Biology, 167(3):273–282, April 7, 1994. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519384710691.

Ying:1990:MRG

[YH90] Da-Jun Ying and Gwang-Tsi Ho. Mathematical representation with graphic reconstruction on a microcomputer for an arterial tree. Journal of Theoretical Biology, 146(2):201–207, September 21, 1990. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305801353.

Yamauchi:1995:CFF

[YI95] Atsushi Yamauchi and Yoh Iwasa. Coupling of fermentation and foraging strategies of herbivorous mammals. Journal of Theoretical Biology, 172(1):1–11, January 7, 1995. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251938570001X.

Yi:1998:FDS

[Yi98] Tao Yi. Frequency-dependent selection: a two-phenotype, two-allele, single-locus, two-species diploid model. Journal of Theoretical Biology, 191(3):289–294, April 7, 1998. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519397905828.

Yang:1998:MSC

[YLT98] S. Yang, X. Liu, and M. T. Tyree. A model of stomatal conductance in sugar maple (Acer saccharum marsh). Journal of Theoretical Biology, 191(2):197–211, March 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251939790583X.

Yockey:1995:CLT

[Yoc95] Hubert P. Yockey. Comments on “Let there be Life”; Thermodynamic Reflections on Biogenesis and Evolution“ by Avshalom C. Elitzur. Journal of Theoretical Biology, 176(3):349–355, October 7, 1995. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519385702044.

Yoshigaki:1999:SDG

[Yos99] Tomoyoshi Yoshigaki. Simulation of density gradients of astral microtubules at cell surface in cytokinesis of sea urchin eggs. Journal of Theoretical Biology, 196(2):211–224, January 21, 1999. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519398908359.

Yotsumoto:1993:DMP

[Yot93] Akira Yotsumoto. A diffusion model for phyllotaxis. Journal of Theoretical Biology, 162(2):131–151, May 21, 1993. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519383710805.

Yi:1997:ESS

[YQsZgZw97] Tao Yi, Yang Qi-sen, Jiang Zhi-gang, and Wang Zu-wang. Evolutionarily stable strategy, stable state, periodic cycle and chaos in a simple discrete time two-phenotype model. Journal of Theoretical Biology, 188(1):21–27, September 7, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519397904525.

Yi:1992:RDS

[YS92] Tao Yi and Zhao Songling. Resource-dependent selection. Journal of Theoretical Biology, 159(4):387–395, December 21, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251930580685X.

Youhei:1994:ISS

[YY94] Fujitani Youhei and Kenji Yamamoto. Influence of slight sequence changes on the free energy of a single stranded ribonucleic acid molecule. Journal of Theoretical Biology, 171(2):151–161, November 21, 1994. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519384712203.

You:1999:ASV

[YY99] Lingchong You and John Yin. Amplification and spread of viruses in a growing plaque. Journal of Theoretical Biology, 200(4):365–373, October 21, 1999. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519399910019.

Yan:1991:PNA

[YYY91] Johnson F. Yan, Alexander K. Yan, and Benjamin C. Yan. Prime numbers and the amino acid code: Analogy in coding properties. Journal of Theoretical Biology, 151(3):333–341, August 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305803820.

Yomo:1997:GES

[YYYU97] Tetsuya Yomo, Tsuyoshi Yamano, Keizo Yamamoto, and Itaru Urabe. General equation of steady-state enzyme kinetics using net rate constants and its applicaiton to the kinetic analysis of catalase reaction. Journal of Theoretical Biology, 188(3):301–312, October 7, 1997. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519397904744.

Yi:1994:NSC

[YZ94] Tao Yi and Wang Zuwang. The necessary and sufficient conditions of density-dependent evolutionary stable strategy (DDESS) in the two-phenotype model. Journal of Theoretical Biology, 167(3):257–262, April 7, 1994. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384710678.

Yi:1995:FDD

[YZ95] Tao Yi and Wang Zuwang. Frequency- and density-dependent selection in the diploid population with only two pure strategies. Journal of Theoretical Biology, 176(3):317–323, October 7, 1995. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519385702019.

Yi:1997:ETD

[YZ97] Tao Yi and Wang Zuwang. Effect of time delay and evolutionarily stable strategy. Journal of Theoretical Biology, 187(1):111– 116, July 7, 1997. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519397904276.

Zahalak:1996:NAM

[Zah96] George I. Zahalak. Non-axial muscle stress and stiffness. Journal of Theoretical Biology, 182(1):59–84, September 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519396901435.

Zamir:1999:FPA

[Zam99] M. Zamir. On fractal properties of arterial trees. Journal of Theoretical Biology, 197(4):517–526, April 21, 1999. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251939890892X.

Zieschang:1997:MRG

[ZBB97] H. E. Zieschang, P. Brain, and P. W. Barlow. Modelling of root growth and bending in two dimensions. Journal of Theoretical Biology, 184(3):237–246, February 7, 1997. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519396902593.

Zeng:1990:TCP

[ZCS^{+}90] W.-Z. Zeng, M. Courtemanche, L. Sehn, A. Shrier, and Leon Glass. Theoretical computation of phase locking in embryonic atrial heart cell aggregates. Journal of Theoretical Biology, 145(2):225–244, July 24, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305801286.

Zhang:1994:CLU

[ZGZ94] Shiping Zhang, Emanuel Goldman, and Geoffrey Zubay. Clustering of low usage codons and ribosome movement. Journal of Theoretical Biology, 170(4):339–354, October 21, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384711969.

Zhang:1997:ODM

[ZH97] H. Zhang and A. V. Holden. One-dimensional modelling of the vulnerability to re-entry to homogeneous atrial tissue. Journal of Theoretical Biology, 184(2):117–122, January 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396902805.

Zhang:1998:SRS

[ZH98] Da-Yong Zhang and Ilkka Hanski. Sexual reproduction and stable coexistence of identical competitors. Journal of Theoretical Biology, 193(3):465–473, August 7, 1998. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519398907123.

Zhang:1990:EBF

[Zha90] Chun-Ting Zhang. Equations between frequencies of amino acids in organisms. Journal of Theoretical Biology, 142 (2):281–284, January 23, 1990. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519305802292.

Zhang:1997:STD

[Zha97] Chun-Ting Zhang. A symmetrical theory of DNA sequences and its applications. Journal of Theoretical Biology, 187(3):297–306, August 7, 1997. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251939790401X.

Zhang:1998:ESRa

[Zha98] Da-Yong Zhang. Evolutionarily stable reproductive strategies in sexual organisms: IV. parent-offspring conflict and selection of seed size in perennial plants. Journal of Theoretical Biology, 192(2):143–153, May 21, 1998. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519397906263.

Zheng:1991:MMP

[Zhe91] Tan Zheng. A mathematical model of proliferation and aging of cells in culture. Journal of Theoretical Biology, 149(3):287– 315, April 7, 1991. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305803078.

Zhu:1991:TBT

[Zhu91] Cheng Zhu. A thermodynamic and biomechanical theory of cell adhesion part i: General formulism. Journal of Theoretical Biology, 150(1):27–50, May 7, 1991. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305804734.

Zhu:1996:RDR

[Zhu96] Bao Ting Zhu. Rational design of receptor partial agonists and possible mechanisms of receptor partial activation: a theory. Journal of Theoretical Biology, 181(3):273–291, August 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396901307.

Zhang:1994:HAN

[ZJ94] Da-Yong Zhang and Xin-Hua Jiang. Is heterozygous advantage necessary for polymorphism in one-locus two-allele systems? Journal of Theoretical Biology, 166(3):245–250, February 7, 1994. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519384710228.

Zhang:1995:LMC

[ZJ95] Da-Yong Zhang and Xin-Hua Jiang. Local mate competition promotes coexistence of similar competitors. Journal of Theoretical Biology, 177(2):167–170, November 21, 1995. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519385702354.

Zhang:1997:ESR

[ZJ97] Da-Yong Zhang and Xin-Hua Jiang. Evolutionarily stable reproductive strategies in sexual organisms: III. the effects of lottery density dependence and pollen limitation. Journal of Theoretical Biology, 185(2):223–231, March 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519396903045.

Zhang:1998:ESRb

[ZJ98] Da-Yong Zhang and Xin-Hua Jiang. Evolutionarily stable reproductive strategies in sexual organisms. Part V — joint effects of parent-offspring conflict and sibling conflict in perennial plants. Journal of Theoretical Biology, 192(3):275–281, June 7, 1998. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519397906019.

Zotin:1996:ABC

[ZL96] A. I. Zotin and I. Lamprecht. Aspects of bioenergetics and civilization. Journal of Theoretical Biology, 180(3):207–214, June 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396900971.

Zhang:1997:ECA

[ZL97] Da-Yong Zhang and Kui Lin. The effects of competitive asymmetry on the rate of competitive displacement: How robust is Hubbell’s community drift model? Journal of Theoretical Biology, 188(3):361–367, October 7, 1997. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519397904914.

Zhang:1998:NAD

[ZLYZ98] Chun-Ting Zhang, Zhe-Suai Lin, Ming Yan, and Ren Zhang. A novel approach to distinguish between intron-containing and intronless genes based on the format of Z curves. Journal of Theoretical Biology, 192(4):467–473, June 21, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519398906713.

Zhang:1995:AMS

[ZM95] M. Q. Zhang and T. G. Marr. Alignment of molecular sequences seen as random path analysis. Journal of Theoretical Biology, 174(2):119–129, May 21, 1995. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385700859.

Zonneveld:1996:SCC

[Zon96] C. Zonneveld. Sperm competition cannot eliminate protandry. Journal of Theoretical Biology, 178(1):105–111, January 7, 1996. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519396900107.

Zonneveld:1997:MEP

[Zon97]  C. Zonneveld.  Modeling effects of photoadaption on

the photosynthesis-irradiance curve. Journal of Theoretical Biology, 186(3):381–388, June 7, 1997. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519397904008.

Zonneveld:1998:PAP

[Zon98] C. Zonneveld. Photoinhibition as affected by photoacclimation in phytoplankton: a model approach. Journal of Theoretical Biology, 193(1):115–123, July 7, 1998. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519398906889.

Zhang:1991:SSE

[ZS91] Chun-Ting Zhang and Zuo-Xu Shang. The study of stacking energy for natural DNA sequences. Journal of Theoretical Biology, 149(2):257–263, March 21, 1991. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305802802.

Zhao:1992:KSR

[ZT92] Kang-Yuan Zhao and Chen-Lu Tsou. Kinetics of substrate reaction during irreversible modification of enzyme activity where the modifier is not in great excess of the enzyme. Journal of Theoretical Biology, 157(4):505–521, August 21, 1992. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305806666.

Zonneveld:1997:MCC

[ZvdBK97] C. Zonneveld, H. A. van den Berg, and S. A. L. M. Kooijman. Modeling carbon cell quota in light-limited phytoplankton. Journal of Theoretical Biology, 188(2):215–226, September 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519397904707.

Zou:1995:NND

[ZW95] Guangzhou Zou and Hsin-I Wu. Nearest-neighbor distribution of interacting biological entities. Journal of Theoretical Biology, 172(4):347–353, February 21, 1995. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251938570032X.

Zhao:1996:KAM

[ZW96] Kang-Yuan Zhao and Zhi-Xin Wang. Kinetic analysis of modification reactions at comparable enzyme and modifier concentrations. Journal of Theoretical Biology, 181(4):319–327, August 21, 1996. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519396901344.

Zhang:1998:REE

[ZWH98] H. Zhang, R. W. Winslow, and A. V. Holden. Re-entrant excitation initiated in models of inhomogeneous atrial tissue. Journal of Theoretical Biology, 191(3):279–287, April 7, 1998. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519397905981.

Zhang:1994:ADB

[ZZ94] Chun-Ting Zhang and Yong Zhan. Analysis on the distribution of bases in 1487 human protein coding sequences. Journal of Theoretical Biology, 167(2):161–165, March 21, 1994. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519384710605.

Zhang:1999:QDA

[ZZ99] Chun-Ting Zhang and Ren Zhang. A quadratic discriminant analysis of protein structure classification based on the helix/ strand content. Journal of Theoretical Biology, 201(3):189–199, December 7, 1999. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519399910238.